A Complete Bibliography of Publications in Journal of Computational Physics: 1990–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

04 September 2023 Version 1.16

Title word cross-reference

(2 + 1) [PZ98]. 0.5 cos 2nx [Kob91]. 0.5 cos Inx [Kob90]. 1 [BM91a, BM91b, FQ96, GBL^{+}92, KO98, LLD92, LLD93, TH96a, VAVB92, VAVB93]. 1 + 1 [YG94]. 11/2 [GBL^{+}92]. 2 [BP90, BBF^{+}99, BR97b, BDML94, COS^{+}98, CC98, Clé96, Dau92a, Dau92b, Dob99, Dvo91, DK92d, DK92a, DK92c, DK92b, Eça96, Efr97, FLD93, FWBS94, FG97, GNHP95, GM96, GHS93b, Hu95, IM95, Ji96, Kan92, KL92a, Kar97, KCJ95, KO98, LMP92a, LMP92b, LQK98, Mat94, Mau99, MW94, Nor96, PH94, Pet96, PV99, PV93, RF96, RS94b, Str96, Str97, TA91b, TAD92a, Wal94, Wes92a, Wes92b]. 2N [SH98b]. 3 [Bor91b, Bor92a, Bor92b, BIAV98, COS^{+}98, CDM98, CPB96, DH96, EL97a, FK97, FPB98, GHKH93, GS93, HKS98, ID96, KKH^{+}99, KH97, LL95, MKM99, MKM04, Mar90, OS97, Oos97, PK94, Sal94, Sha95a, SS99c, TZA96, TLW95, Val90, Xia99]. 3d^{N} [YR93]. ^{+} [Lio96]. _{2} [KE93]. Ai(x) [MPG92a, MPG92b]. α [ACT95]. β [CGKC92, CGKC93, RG90]. C exp(−λ|x|^{ν}) [BES90]. d/dx [Ano95u, Str94]. d = 2 + 1 [Kou96]. Δ [TV94, Gou96a, VDJ93]. δf [DK95, DMZ91, KT95]. E × B [BDML94].

F_{m}(x) [CP98b]. γ [JZQ^{+}95]. h^{4} [CGY92, CGY93]. Je_{0}(a, z) [DK90]. k [BH95, KL92b, LZ96]. K(2, 2) [dLMSS95]. LU [Shu91]. m [KM92]. N [JDC94, SDG94, OGBSK90]. ∇ · (σ∇u) ≈ f [ZY94]. O [PRMV90]. O(n) [WA95]. O(N^{−1}) [KTE93]. ω [LZ96]. ω_{ce}Δt [PB91]. P [ST97, CLO93]. P^{1} [ADS96]. ∂u∂t = (∂∂x)^{α}δGδu [Fur99]. QZ [vD97b]. r [Ano99-30, CHR99, MRS98]. S^{2} [PZ98]. S_{n} [MWS96]. U(1) [HN95]. ε [KL92b, TV94]. X^{T} R [DÜ92b, DÜ92a]. Z [PRMV90, MRS98].

-Adaptive [Ano99-30, CHR99]. -body [OGBSK90, JDC94, SDG94]. -d [Mar90, CDM98, CC98, FQ96, HKS98, ID96, LMP92a, OS97]. -dimensional [Val90]. -Functions [Gou96a]. -Space [BH95]. -Stable [ST97]. -Storage [SH98b]. -Version [CLO93].

/algebraic [KL92b].

148 [CNG17]. 149 [Ano99-30]. 152 [MKM04].

2-D [KPCJ97]. 244 [Ano99-30]. 2d [Lay90].

92 [LW96b].

ABCs [YP98]. abruptly [Lav90]. Absolute [CL98b]. Absorbing [BR92c, BR92d, Clé96, Col97, Hig92a, Hig92b, Hu96, JW90, PZC98, Pet98, Ren92a, Ren92b, RF96, Rom92a, Rom92b, SDI98, TAC98]. Absorption [Ber96, Ber94, MB92]. Abstract [Ano91a, Ano91b, Ano91c, Ano92a, Ano92b]. Abstracts [Ano90c, Ano90d, Ano90e, Ano90a, Ano90b, Ano90f, Ano91d, Ano91e, Ano91f, Ano92c, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano93h, Ano93i, Ano93j, Ano93k, Ano93l, Ano94a, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano94i, Ano94j, Ano94k, Ano94l, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano95i, Ano95j, Ano95k, Ano95l, Ano95m, Ano95n, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, Ano96a, Ano96i, Ano96j, Ano96k, Ano96l, Ano96m, Ano96n, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Ano97j, Ano97k, Ano97l, Ano97m, Ano97n, Ano97o, Ano97p, Ano97q, Ano97r, Ano98a, Ano98b, Ano98c, Ano98d]. Abstracts [Ano98e, Ano98f, Ano98g, Ano98h, Ano98i, Ano98j, Ano98k, Ano98l, Ano98m, Ano98n, Ano98o, Ano98p, Ano98q, Ano99a, Ano99b, Ano99c, Ano99d, Ano99e, Ano99r, Ano99f, Ano99g, Ano99h, Ano99i, Ano99j, Ano99k, Ano99l, Ano99m, Ano99n, Ano99o, Ano99p, Ano99q]. AC [LJG96, WCO94]. Accelerated [BGM91, PHL93]. Accelerating [BR94, TWV97]. Acceleration [AAL97, Azm99, BGB99, Eye96, SPC92, SPC93]. Acceptance [Nea94]. accepting [DS90]. access [CV92a, CV92b]. Accuracy [AS99b, BT97a, BIAV98, CNG99, CP95, CK93b, FMO97, KKH^{+}99, MW97, Maz97, Raw90, Raw91, Sil93, SVG93, Vre91a, Vre91b, XP99, YG94, Zha98, AB94, Arm94, BT92a, BT92b, CNG17, GL96, Gre90, Gup90, Gup91, RS91, Sil92, SVG92]. Accurate [AD97, BR97a, BR97b, BPT95, Bri95, DVD93, DM97b, Dor99, EG95, FM97, HT95, HEOC97, Hen94, IH95, Jac95, KJ98, KPP99, Kul95, MH94, MLS99, MB94, Nor98, Nor95, PM98, RK95, RS94a, RH98, Sco96, SNU98, SW96, SH97, TH96a, TP97, VM94, VB95, Yak96, YMUS99, vBvNW95, BM90b, BM91c, KS92, Tak92a, Tak92b]. Acknowledgment [Ano95o, Ano96o]. Acoustic [AK99, WJTP96, Hig92a, Hig92b]. Acoustics [AGH99, AK94a, AE98, HHM96, QG98, SH98b, TW93]. Action [KB99, RP90, RP91]. action-angle [RP90, RP91]. Active [WJC93]. Actuator [HH98]. Adaptation [SR97]. Adapted [YB98]. Adaption [Hag94]. Adaptive [ALTP98, ABC94, ABC^{+}98, Ano97-29, Ano99-30, APS97, AC93, BD96, Bau97, BCM95a, BGM95, BK97a, BFO99, Bra93, CZ98, CHR99, CBS93, Coo99, Dvi90, Dvi91, Edw96b, FGM97, FS97, GBCA99, GO98, GMP92a, GMP92b, Har94, HJ98, HT97, JP95, JFH^{+}98, Kho98, KJ98, KP98, LW95, LLD93, LT93, LH95, LT96, Mav90, MF92, MQS97, MA96, MG97b, NB98, PK94, PBC^{+}95, PPM92a, PPM92b, PRL^{+}99, PGD99, RPB99, SLMS98, SRF93, SB96b, Str97, SL98, ST99, SAB^{+}99, TYS95, VP96, VP97, ZP95, ZLP97, AM90, AS90, AS91, BKM90, But90a, But91, CW90, CGR99, CW92c, DT90a, DT90b, Dar92, Dur91, DEO92, FM92, Fra92, GL96, HGH90, HGH91, LLD92, Nad95, Wes94, YM91, YM92b, YM92a]. adaptive-grid [YM91, YM92b, YM92a]. Adaptively [DP93, PV99, DP92a]. adaptively-refined [DP92a]. Addition [lX96, lX97]. Additions [Lin95]. Additive [Zho96]. Adhesion [ALTP98]. ADI [AS94, AK98, CZ98, HLD92b, HLD92a]. ADI-Solver [AK98]. Adiabatic [Vu96, Weg97, WK96]. Adjacent [Azm99]. Adjacent-Cell [Azm99]. Adjoint [WEQ^{+}99, YZ98]. adjustable [Fri90]. Adjustment [WNY93]. Admittance [WEMH99]. admitting [Shu92a]. adoption [SV91]. Advance [Mat94]. Advanced [GNH96, Zan97, YP92b, YP92a]. Advances [Sch99]. Advancing [KP98, Mav95]. Advancing-Front [KP98]. Advection [Ano97-29, AE95, BYMZY95, Bec94, BT97a, EP96, Fey98b, FPQ93, HMOS95, Hol94, HKvV95, OS98, OP97, PA95a, PBD94a, SWT98, SB96b, Sun93, Thu96, VK98, Wan97, WEQ^{+}99, AP90, AP91b, Ben92d, Ben92c, SG90]. Advection-Diffusion [BYMZY95, Bec94, EP96, FPQ93, HMOS95, OP97, PBD94a, SWT98, Wan97]. Advection-Reaction [WEQ^{+}99]. Advective [AGH99]. Aeroacoustics [Eka99, MLBW97, ÖL96]. Aerodynamic [Ari97, BB97a, KTN^{+}94, STG^{+}96, TWV97]. aerodynamics [FS91, FS92]. Aeroelasticity [Shu95]. Aggregation [WF99, VG92b, VG92a]. Aided [MB99]. Air [Sun96]. Airflow [Hol94]. Airfoil [SS99c, vBKG97, HPS92a, HPS92b]. Airfoils [FH98b, MDB91a, MDB91b]. Airy [CJR91, CJR92a, CJR92b, MPG92a, MPG92b]. AlAs [LK94b]. ALE [Smi99, LKK99, PFS98]. Alfvén [BK99]. algebra [WS91, WS92a, WS92b].

Algebraic [CWF96, Gou96b, Mat97, Spe95, SDI98, Wha96, KL92b]. Algorithm [AL97, AB97, AWY99, AFH94, BS99b, BS99c, BD94b, Boa97, BN94, BO98, BFO99, CDM98, CC97, CL93, DS99, DK95, DFaIM98, DHSS94, DD95, Due93, Edw96a, EM95, FD99, FS97, Fuj95, GBCA99, GMM99, GKDT96, Gou96a, Gra95, GNH96, GR97, Gre94, Hol94, JH97, JFH^{+}98, Ji94, JMR94, JS93, Kar94, KF94, KM95, KT95, KNR99, Kul95, LeV98, Leh99, Lub93, MJP93, Mav95, MSV98, MDH98, NCF96, Nea94, NPV96, Ods93, ÖL96, Pal93, PDHS94, Pet94, QF99, Reb93, RF95, RH98, SD93a, SLMS98, SMJ98, Shy98, Shy99, SW99b, TS98, VP97, Vie94, Wan95, WL96, YL98, YG94, Zal97, ZC94, ZB96, ZDD99, Zin94a, Zin94b, BRR92, BG92, Bas92, BBK90, BBK91, Boy92b, BES90, BS91b, BS92a, BS92b, CR91, CW90, CGR99, CW92c]. algorithm [Dar93, DM91a, DM92a, DM92b, Dol92a, Dol92b, DS90, DLMN91, DLMN92b, DLMN92a, HB90a, HB91, JVR93, KH90, KH92, KHM92a, KHM92b, LDB96, LKLE90, LLEK91, LMP92a, LMP92b, LK90, LS92c, LS92d, Luc90, Luc91, LM92, MMS90a, Ods90a, Ods90b, Osb90, Osb91, PRS90, Par90, QS90, Rei92a, Rei92b, Row91, SG90, Sol92, SRSB93, Ver91, YSG91]. Algorithmic [SSB^{+}99]. Algorithms [AS95a, ARB96, BCM99a, BR94, Boy92c, BFGG94, BCDL97, BCT98, CK96, CBSW98, CMM93, DH96, DS97a, DL93, DM93, ETRW95, EP99, Goe95, HBF93, HKW97, Kar97, Kho98, LeV97, LML99, Liu96b, Löh95, MRC93, Maz97, Nat91, PDHS94, Pli95, Pri95, Rap93, Rav97, RFL93, Rie94b, Ruu98b, SKR^{+}93, SF96, SS95, SK96, SE90, SC97, SGW94, UL98, WS96, WD97, WH97, dM97a, Alm93, Bal95, BRL91, Dar92, FG90b, FG90a, GC90, HW92a, HW92b, Mak90a, MRD92, Nat90, RD91, Shu90, SLV90, TP92a, TP92b, Vre91a, Vre91b]. aliasing [Ben92a, Ben92b]. all-pairs [BF91]. Alloys [YZ98]. Almost [AKW93]. along [MB99]. Alternate [Taf95]. Alternating [DB98, MWW96, RY94, ZWS98]. Alternating-Directional [ZWS98]. Alternative [CO95, Pri94, Shu95]. Amalgamation [SP96]. Ambient [MSF94]. Amplitude [Nic99, ML92]. analogy [Glo92]. Analyses [Hwa94]. Analysis [AG97, Abg94, AN98, AR93, BK99, Can95, CS99, CB96, CSZ97, DH96, Efr97, FH98b, FQ96, Gho96, Gil97, GTA99, GBS99, HG94, Hes98, HB96, HI98, HHR99, ID96, JMDB99, KD97, KGH^{+}98, KM93b, KB96, LNR99, LW97, Liu96b, Liu96a, MB99, Mas96, Mat97, NGN97, PNC94, PVVS98, Per93, RCA93, RHB94, SS96b, SMW99, SDP97, SC96a, SW99c, SHA95b, TS98, VM96a, VSG95, WEMH99, YP98, YS98a, Yav97, AF90, AF91, Bal95, Cha90, CH90b, CH91, CK90b, CK91, Fab92, GM92, JC92, KBS90, LKLE90, LLEK91, LS92c, LS92d, MS90a, MS91a, ND90, Osb90, Osb91, PO90, PO91, RCA92, Raw90, Raw91, Sch92b, Sch92a, Sch92c, TA91b, TAD92a, TAD92b]. Analytic [GM95b, NCF96, Dom90]. Analytical [Bar94, dFD^{+}94]. Anatomical [JBA96]. and/or [CK96]. anelastic [Ful93]. Angle [AE98, NY98, DLMN91, DLMN92b, DLMN92a, RP90, RP91]. angle-dependent [DLMN91, DLMN92b, DLMN92a]. Angular [Ben95, UL98, Dvo91, DK92d, DK92c, Mik90, Mik91, RF90, RF91, WS91, WS92a, WS92b, Zan92a, Zan92b]. Anisotropic [ABBM96, Hag94, HKW97, MG96, Mav98, Oos97, TN98, Waj93, ZV97, vdVvdV98, AP91a]. Anisotropically [UL98]. Anisotropy [Kal97, TW94]. Annealing [CDR99, DS90]. Announcement [Ano90g, Ano92d, Bra99]. Announcements [Ano90h]. Annulus [CKSB97]. Anomalies [SR90]. anti [Ben92a, Ben92b]. anti-spatial-aliasing [Ben92a, Ben92b]. Antiplane [GT97]. Apparent [Jak93, DM92c]. Appear [Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93g, Ano93h, Ano93i, Ano93j, Ano93k, Ano93l, Ano94a, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano94i, Ano94k, Ano94l, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano95i, Ano95j, Ano95k, Ano95m, Ano95n, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, Ano96a, Ano96i, Ano96j, Ano96k, Ano96l, Ano96m, Ano96n, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Ano97j, Ano97k, Ano97l, Ano97m, Ano97n, Ano97o, Ano97p, Ano97q, Ano97r, Ano98a, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano98h, Ano98i, Ano98j, Ano98k, Ano98l, Ano98m, Ano98n, Ano98o, Ano98p, Ano98q]. Appear [Ano99a, Ano99b, Ano99c, Ano99d, Ano99e, Ano99r, Ano99f, Ano99g, Ano99h, Ano99i, Ano99j, Ano99k, Ano99l, Ano99m, Ano99n, Ano99o, Ano99p, Ano99q, Ano90c, Ano90d, Ano90e, Ano90a, Ano90b, Ano90f, Ano90p, Ano90q, Ano91a, Ano91d, Ano91b, Ano91e, Ano91f, Ano91c, Ano91q, Ano91o, Ano91p, Ano92a, Ano92c, Ano92b, Ano92k, Ano93f, Ano94j, Ano95l, Hor90]. appearing [DS90]. Application [ARTAAA97, AC92, AAP97, BFW98, BCRR98, BG92, Bas92, BS99d, BS99e, Boy94, Boy95a, Boy96b, Boy97a, BGP98, CGA94, CS99, CM93, CKE99, DBS94, DZ96, Fen99, GTD98, HCZ99, Hol94, HLOZ97, IH95, KvdVPG95, KE93, LS93a, MML98, Mas96, NPC93, NB98, Pal93, PFRB93, RG97, SB96a, Sak96, Shu95, Tol94, Tót97, Tsy95, WNY93, Wit96a, XS93, YS98a, YMB^{+}91, ZLP97, ZRR99, Dar90c, Dar91b, DM91a, DM92a, DM92b, GMP92a, GMP92b, KH90, KH92, LS92b, MDB91a, MDB91b, TSS92, YR98]. Applications [Ben95, BCT98, Don94, FH97, Kla96, Luk99, RvR93, SMD98, Sch99, vL97a, vL97b, Dar90b, Dar93, KL97, Nat90, Nat91, Rei92a, Rei92b]. Applied [ALW94, ABS96, BD93, BK97b, CAA93, Cra94, DFaIM98, ES94, EP96, FO93, Gla95, Gra95, KI99, Kno98, Man93b, NC99, PH95a, SK94, Sco96, Set94, WLC96, BB90a, BB91a, Dol92a, Dol92b, NSR92a, NSR92b, PC92b, PC92a, YP92b, YP92a]. Approach [Abg96, AS95a, AS95b, AS97, BKP96, Cha95, Coo99, FAMO99, FHKZ97, GDP96, HOS96, HRL99, KP97b, KNR99, LWT99, LLT94, Lya99, MBO94, Pap93, Rom97, Ruu98a, SFHD99, SSO94, SAB^{+}99, TS96, TSW95, Tou98, Wit96b, ZCMO96, ZMOW98, DHS91, Gra90, JB91, JBvK^{+}91, LKLE90, LLEK91, MIM90, Pap92, SW92, TSR92, VY91, VY92, VG92b, VG92a, YSG91]. approach-algorithm [LLEK91]. Approaches [CP95, PMR97, FS91, FS92].

Approved [Tay91]. Approximate [Bai97, BHJU99, DW94a, DD92b, Gla95, Gla92b, Gla92a, Gre94, Roe97a, Sai95, SCM98, SLL94, TK96, Tou92]. Approximate-Factorization [BNW96]. Approximately [KTN^{+}94]. Approximating [BS99d, CK90b, CK91, NM95, LO91a, LO91b]. Approximation [Ano94s, ACT95, ADS96, BCM95a, BGM95, Cor96, FH94, FPB98, FPQ93, Her93, HS93b, MP93, Ohw98, PR93b, Sai95, SNU98, Smo98, SVG93, Wer95, Car91, KS92, LM90c, NSR92a, NSR92b, SVG92, TEKC92]. Approximations [BCDL97, GTA99, HMOS95, SK94, SRVK96, Ano95u, BB90a, BB91a, BT92a, BT92b, MPG92a, MPG92b, Str94, WDH^{+}92a, WDH^{+}92b]. Arbitrary [AB97, AWY99, AD93, BIAV98, CG96, CNG99, Gra99, HAC97, Hum96a, Hwa94, KW93, LV98, Mar97, MG97b, PFS98, Smi99, Tab96, Tau94, UI99, UKSTST97, Ush96, CNG17, GL96, MS91b, MS92b, MS92a, Rob90b]. Arc [Poy92, WCO94]. Architectures [SDG94]. Arctan [Boy91b, Boy92a]. Arctan/Tan [Boy91b, Boy92a]. ARCTIC [MMW96]. Areas [YPH94]. Arguments [CP98b, CJR91, CJR92a, CJR92b]. arising [CGSS90, CGSS91, Nat92a]. Arnoldi [Nat92a, KvdVPG95]. Arnoldi-based [Nat92a]. Array [EA97, SK94]. Arrays [DS97a, BF91]. arrowhead [OS90]. Arteriolar [AMP^{+}98]. Article [Ano99-30, Van99, TT90]. Artifacts [SS99b, SMSS98]. Artificial [AD93, BL99, BH96, Boy98, CS98a, CSW98, Cot96, DR96, Efr97, TZA96, Tou97, Tou98, XMJ95, HLB94, MW93, Yan90]. Aspect [Kal97, SP98, VRD99]. Aspects [KKH^{+}99]. ASPEN [VBD99]. Assemblages [HM99, ZDG95]. assemblies [CPP93, LDB96]. Assessment [CP95, XP99, Sur94]. Assimilation [Lya99]. Associated [SPC93, MRD92, SPC92]. astronomical [Rei92a, Rei92b]. Asymptotic [AN98, CS93b, CM91, CJ95b, FEO95, JYH93, ZOI97, Bor91a, BLA92a, BLA92b, DK92a, DK92b]. Asymptotically [AD97]. asymptotics [Kle95]. Atmosphere [ALW94, Tol94]. Atmospheres [CMMH94]. Atmospheric [Ano97-29, FWBS94, Hol94, SPCD96, SB96b, SNL90]. Atom [Hof95, KKOF97]. Atom-Centered [Hof95]. Atomic [dFD^{+}94, LKB90, MAS92]. Atomistic [Had99]. Atomistic-Continuum [Had99]. Atoms [BK97b, HM99]. Attached [DW93]. attenuating [Nic93]. Augmented [Fra95]. AUSM [Lio96, Smi99]. Author [Ano90i, Ano90j, Ano90k, Ano90l, Ano90m, Ano90n, Ano91h, Ano91g, Ano91i, Ano91j, Ano91k, Ano91l, Ano92e, Ano92f, Ano92g, Ano92h, Ano92i, Ano92j, Ano93m, Ano93n, Ano93o, Ano93p, Ano93q, Ano93r, Ano94m, Ano94n, Ano94o, Ano94p, Ano94q, Ano95p, Ano95q, Ano95r, Ano95s, Ano95t, Ano96p, Ano96q, Ano96r, Ano96s, Ano96t, Ano96u, Ano96v, Ano96w, Ano97s, Ano97t, Ano97u, Ano97v, Ano97w, Ano97x, Ano97y, Ano97z, Ano97-27, Ano97-28, Ano98r, Ano98s, Ano98t, Ano98u, Ano98v, Ano98w, Ano98x, Ano98y, Ano98z, Ano98-27, Ano99s, Ano99t, Ano99u, Ano99v, Ano99w, Ano99x, Ano99y, Ano99z, Ano99-27, Ano99-28]. Automata [Kre98, RMO99, vWBS^{+}97, BRR92]. Automated [WS91, WS92a, WS92b]. Automatic [CMW95, GNH96, Nog93, STG^{+}96, Tiw98, LP90]. automation [BB90b]. Automaton [SBC99, BB91c, CMK90]. Auxiliary [LO93, OB95, PC92a]. auxiliary-heated [PC92a]. auxilliary [PC92b]. auxilliary-heated [PC92b]. Average [BLG97, OGBSK90, VM90]. Average-State [BLG97]. Averaged [AAL97, CW93, CPJ90, DHS91]. Averaging [Smo98]. axes [Lay90]. Axi [NB98, Sch95a]. Axi-symmetric [NB98, Sch95a]. Axially [DÜ92a, SLMS98, DÜ92b]. Axillary [KOS^{+}96]. Axisymmetric [AM96, ART95a, ART95b, KC93, MG97a, PBD94b, PS93b, RH99, WJC93, XS93, HCJ92a, HCJ92b, KBS90, KC92, PDC91, LS98]. Axisymmetrization [Kou97].

B [HB90b, KMS99, Man93a, SM98a]. B-Spline [KMS99, Man93a, SM98a]. B-splines [HB90b]. Backflow [RHT96]. Background [Tan94, GCMR90]. Backward [HIY95, Zhu95]. Backward-Facing [Zhu95]. Baffled [SK94]. Balance [AAL97, APS97, YM95]. Balancing [FLD93, LeV98, BBK90, BBK91]. Ball [Rap93]. Ballistically [LPR96]. Band [AK99, Dob99, FK97, För96]. Band-Gap [AK99]. Bands [GT97]. Baroclinic [Die95, HdS97, Hig99]. Barotropic [Die95, HdS97, Hig99, TG96, Gla91]. Barotropic-Baroclinic [HdS97, Hig99]. Based [Ano99-30, Bis95, BS96, CHR99, CG97b, CS97b, Cod99, Cor96, Dur93, Fuj95, GS97a, GC95, Hag94, HRR94, JLM^{+}96, LS96, LH95, MS98, NY98, OG97, Oos97, PMO^{+}99, Sai95, SZ97, SWS96, SFHD99, SF96, SC97, Spe95, SR97, Str99a, TZA96, WKMH97, WS93, Xu97, Xu99, YSD99, Dur91, DEO92, FS91, FS92, FVZ90, Kle95, LDB96, Luc90, Luc91, Mul92b, Mul92a, Nat92a, Ver91]. Bases [BK97a, KP97a]. Basic [GM95a, Rus90]. Basics [Dru99]. Basin [Boy96b, Boy97a, Ma93a, Ver97a, Ver97b]. Basis [BK94, CP98a, Gou96b, KOS^{+}96, MP93, QF99, UWSB90, UWSB91, VSG95]. Basis-spline [UWSB90, UWSB91]. bathymetry [LG94]. Battle [ZLP97]. Baxter [PRMV90]. BDF [vM95]. Be [AAK^{+}93]. Beach [GGP97]. Bead [JVS97]. Beam [KT95, LLR94, MMW96, YCTC97, Mik90, Mik91]. beam-beam [Mik90, Mik91]. Beams [Bör97, KST90, KST91]. Bean [Pri96]. Bearing [GDP96]. Beatwave [BGB99]. Beck [Liu96a]. becomes [DM92c]. bed [CGSS90, CGSS91]. Beds [vWBS^{+}97]. Behaved [AR97b]. Behavior [CL98a, Die95, KM97a, Vad95, ZMOW98, AF90, Rob90a]. behaviors [TZWH92b, TZWH92a]. behaviour [AF91]. Beltrami [Dar93]. BEM [vM95]. Bénard [CT98, Gel99]. Benchmark [CH90b, CH91, TO98]. Benchmarking [CK90a]. Bessel [Jab94, PDL93, Raw90, Raw91, SBKH92]. Beta [Bri95]. Beta-Function [Bri95]. Better [Lub93]. between [CDM98, CS97a, FV97, GBL^{+}92, LG90, LG91, MMS90a, ÖM95, RMO99]. Beware [Hil97]. Beyond [BH98]. BGK [FH98a, KKH^{+}99, KJ98, Xu97]. BGK-Based [Xu97]. Bi [ID96]. Bi-grid [ID96]. Bias [AT99]. Biased [Li97]. Biasing [UL98]. Biaxial [YP98]. biconjugate [JE90, KH90, KH92]. bifurcated [HG91, HG92]. Bifurcating [Bal94a]. Bifurcation [KD97, hYsY96, GGH90, She90, She91, VBB90]. Bifurcations [AAP97, PA95b]. Big [Ben96, Rie99b]. Biharmonic [Mar94a]. Billiard [Rap93]. Billiard-Ball [Rap93]. billiards [Lub91]. Binaries [SD93a]. Binary [ES98, FA97, Kre98, YZ98]. Binding [EA97, JMDB99]. Binning [SGW94]. Binomial [GÖAY95]. Biofilm [DFFG96, SCC93b]. Biological [Kla96]. Biomolecular [SS99b, Sch99]. Biophysics [Ano99r, SSB^{+}99, Sch99]. Biperiodic [Mau99]. Bird [BK97b]. Bisection [Ji96, Gre91, Gre92a, Gre92b]. Black [MDH98, dG96]. BLAS [DHSS94]. Blob [CLD^{+}96, DD95]. Blobs [MG96, HM95]. Bloch [GMPGV96]. Block [CN95, DF98, Fat99, WL96, ZS97, IM95]. block-matrix [IM95]. blowup [TSS92]. Bodies [ARTAAA97, Xia99, RIW90, Sal91, dG92a, dG92b]. Body [AKW93, BF97, Cra94, JDC94, JBA96, Lan97, MRC93, PM93, SDG94, Swe94, TWV97, dG96, HB90b, OGBSK90, PM92, YHW90, YHW91]. Boltzmann [Abe97, Anc94, BLT96, DGD90, FO93, HN95, HLD96, HD97, HCD98, HCZ99, HZC^{+}95, HSZ^{+}97, Kla99, LBC^{+}93, LM97, LW94, MS98, MLS99, MBI^{+}97, Ohw98, PHL93, PVVS98, RW96b, RSW98, SC96a, TV94, Tiw98, YH95]. Boltzmann-like [LBC^{+}93]. Born [Bar94]. bounce [DHS91]. bounce-averaged [DHS91]. Bound [Boy96a, Cra94, FK97, dAT95]. Bound-State [dAT95]. Boundaries [APV98, BFW98, DR96, FAH97, Joh93, JMR94, KKH^{+}99, Nak95, PP93, SF99, Wer95, YMUS99, ZHL96, CS91, HL90, HL91, Poi91, Wes94]. boundaries-adaptive [Wes94]. Boundary [APV98, AC95, AK95, AE98, AMP^{+}98, BH96, BPT95, BC99, Ber95, BJ95, BW98, Bot98, Boy98, CLO93, CP90, CGA93, CGA94, CMW95, CJR95, Clé96, Cor96, DR96, DW93, DNN93, DFFG96, DS94, EL96b, Eça96, EP99, Eve96, FH94, Fer95, FJ98, GGG94, GS96a, GC95, GGK95, GHS93a, Gri94, GK95, GK96, GK98, GZ95, GS97b, HG94, HIY95, Hew97, HR96, Hu95, Hu96, IH95, Ji94, JB93, JC98, Joh93, KKK95, KLP94, Leb99, LTE93, LEI96, LM90c, LE90, LE91, LL95, LLK98, LT96, LWS98, Mal96a, Mar93, MM95a, Mas96, MLS99, MA96, Nak95, Nic99, Nor95, NC99, OK94, PNC94, PPBC93, Pen95, PZC98, Pet98, PH95b, PL92b, Poi92, RM93, Ram97a, Rea97, RF96, RPB99, SSR96, SB96a, SY97, SVV98, SI94, SR98, SG98, SW99c, SDI98, SWD95, TS96]. Boundary [TW94, TAC98, TKT97, TG96, TH96b, Tou97, Tou98, Vie94, Wan97, Wan99, ZC94, ZW97, Zho98, ZCH96, Zin94a, ZOI97, ZRR99, Bey91, Bey92, Bor91b, Bor92a, Bor92b, Boy91b, Boy92a, BDR91, BR92c, BR92d, Cal90, Cal91, CW92a, CW92b, CWG90, CWG91, Che91, CS92b, CS92c, Dom90, Fer90a, Fer90b, For90, For92c, FMS92, Giv90, Giv91, Glo92, HLB94, HTK90, Hig92a, Hig92b, HS90, Jac90, JW90, Kan92, KL92a, Kar91, Kar92, KR95, LGC92, LG94, Mal90, Mit92, Or94, PPBC92, PTM92, Poy92, QR90, Ram92a, Ram92b, RR90, RR91, Ren92a, Ren92b, Rom92a, Rom92b, Ros90, RWC90, SKK90, SKK91, SH92, Sal91, SRSB93, TSR92, Tho90, UWSB90, UWSB91, VMK91, Wes92a, Wes92b, YHW90, YHW91, dG92a, dG92b, vdVMK92b, vdVMK92a]. Boundary-Fitted [BW98, SKK90, SKK91, SH92, Wes92a, Wes92b].

Boundary-Induced [Mar93]. Boundary-Layer [BJ95, Mal96a, MM95a, Mas96, Zho98]. boundary-value [Or94]. Bounded [Boy96b, Boy97a, KMM96, PFRB93, PPBC93, PR93b, TSW95, VD97a, VAVB93, AN90, PPBC92, VAVB92]. boundedness [Tas92]. Boussinesq [CM95, LMP92a, LMP92b]. Bowyer [Reb93]. Box [LS92c, LS92d, LL95, MDH98]. Box-counting [LS92c, LS92d]. Breakdown [RCA93, RCA92]. Breaking [AAP97, hYsY96, BLK90]. breakup [SE90]. Bridges [Hel96]. Brief [Nac96]. Broken [GS97b]. Broken-Ice [GS97b]. Brown [KD97]. Brownian [EA97, Yin96]. Bubble [Dar95, Dar96, DE99, HG94, TM93, SCA92a]. Bubble-Laden [DE99]. Bubble-Slug [TM93]. Bubbles [ZYKC98, ZMOW98, MD90]. Bubnov [AC92]. Buff [LPP99]. Bulk [HKW97, KB96]. Buoyancy [LSLG97]. Burger [Boy91b, Poy92, YP92b]. Burgers’ [MS91a, YP92a, AC93, Boy92a, CK90a, MS90a].

Calculate [AFH94]. calculated [SH92, YHW90, YHW91]. Calculating [BFO99, DL94b, oJ95, oJ97, KTN^{+}94, OBL93, OS98, RMSB96, SMT99, TS96, YB98, Zin94b, BE91, BE92a, BE92b, DLMN91, DLMN92b, DLMN92a, TSR92]. Calculation [BD94b, Bru93, ETRW95, För96, GQ97, Jac99, JV95, Lav90, PJ94, Pap93, Sch95a, Sho93, TWV97, TH96b, XS93, lX96, lX97, BS90, BS91a, BF91, Bru92, GLM^{+}90, GSY92, Got92, HHG91, HHG92, LDB96, MDB91a, MDB91b, ML92, Pap92, Rob90b]. Calculations [Abg96, AR97b, BGGT90, Boy96b, Boy97a, CL98a, CJ95b, Dob99, EB94, Fat99, Goe95, HL98, JB93, Kar94, LPP99, MP93, MWS96, NMW^{+}96, PS93a, PDS99, PR93a, PG97, RC95, RK99, RHT96, SK93, Ske98, WJC93, WP97, Wri98, WCSW99, Xu97, BLMR92, Blu92, CLJ^{+}90, CHL^{+}90, HPS92a, HPS92b, HSN90, HB90a, HB91, PH94, YTS92]. Calculus [DÜ92a, DÜ92b]. Cancellation [SMT99]. Canonical [LPP99, Sch95b, KF91]. Cao [Ano99-30]. Capable [SBC99]. Capacitance [BL93, BL92]. Capacitances [Rea97]. Capacity [LEI96, GSY92, Got92]. Capillary [CF94, HF98, HM90]. Capturing [AR97a, BKP96, CHMO96, DW95b, DM96, GMS^{+}99, HP93, KP97b, Lav93, Shy98, TH96a, Ton96, UI99, Yee97, YSD99, ZMOW98, HP92, kM90, kM91, YP92b, YP92a, YKM90]. Carcinogen [BH99]. Carcinogen-Damaged [BH99]. Cardinal [Boy92d]. cardiovascular [VY91, VY92]. Carlo [WONM96, AT99, BPS95, BRL91, Boy90, Boy91a, BGM91, CB96, COBA95, DR90, DB96, EM94, EHM97, GBCA99, Gre98, HBF93, KNR99, KE93, Kuh96, Lif92, MSV98, MC95, Nea94, Nic93, PC99a, Pri90, Pri91, Rie94b, STAS91, SH90, Sch95b, SFHD99, STS^{+}97, UL98, Val91a, Val91b, WVHS98, XP99]. Carrier [SB98]. Cartesian [ALTP98, CP95, DP92a, DP93, Fis94, FJ98, JC98, LZG99, PBC^{+}95, PR93b, UKSTST97, YMUS99]. Cartesian-Grid [FJ98]. Cartesian-Mesh [CP95]. Cartesian-Product [PR93b]. Case [AK99, DF98, Jun99, kM93a, PC99a, AP91a, Mao92b, RLM90, GACN93]. Cases [LS98, CS91]. Casing [LTE93]. Casting [LPR96, BBC^{+}90, BBC^{+}91].

CASTOR [BK99, KGH^{+}98]. Catastrophe [VSG95]. Cathodic [LTE93]. Cauchy [Abr91a, Abr91b, BGH^{+}97, BGH^{+}99, NM95]. Cauchy-Characteristic [BGH^{+}97, BGH^{+}99]. Causal [AS94]. Caused [BDML94, Ush96]. Caustic [Bas92, BG92]. cavitation [SCA92a, SCA92b]. Cavities [HK98]. Cavity [HZC^{+}95, JV95, MWM94, RWW94, GGH90, She90, She91]. Cebysev [GE92]. Cell [Azm99, BH98, Bot98, Bru93, CMM93, Edw96b, GH95, GH97, JLSM97, Kal97, OBL93, QF99, RHT96, TM94, Vu96, VBD99, Bra90, Bra91, Bru92, BSB92, FPRB90, FPRB91, KW92, MK90, MDHW92a, MDHW92b, MRS94, PC92b, PC92a, Wes94]. cell-centered [KW92, MDHW92a, MDHW92b]. Cells [ALTP98, BF97, Lan97, SSW98]. Cellular [Kre98, RMO99, vWBS^{+}97, BRR92, BB90b, BB91c, CMK90]. Center [SH99]. Centered [Edw96b, Hof95, Sjö95a, KW92, MDHW92a, MDHW92b]. Central [GO98, JT93, NT90, ST92b]. central-difference [ST92b]. Centre [KPCJ97]. Centred [GN96]. Ceramic [BMJ^{+}97]. Certain [Sen99]. CFD [GNH96, KH97, MG97b]. CG [VSM99]. Chain [CSZ97, JVS97, SVS90, SVS91]. Chakravarthy [Shu97]. Challenges [SSB^{+}99]. changing [Lav90]. Channel [BH96, CBL95, HG94, LL93a, MH93, PS93a, RHB94, Sie99, Dom90, HLB94, KHM92a, KHM92b]. Channels [VC99, TPL90]. Chaos [AAK^{+}93, HA93, PV93]. Chaotic [KCJ95, dM97a]. Chapman [GA98, KT99]. Characteristic [BGH^{+}97, BGH^{+}99, FMO98, HL98, LS92b, LS93a, Leh99, Mal96b, Nic99, RH98, SF96, YSD99, GWI92]. Characteristic-Based [SF96, YSD99]. Characteristics [MW97, MM95a, MQS95, WK99b, XPK90]. Characterization [GH97]. Charge [BH95, Fen99, Hum96b, BW90, BW91, GSY92, Got92, Hor90, Hor91, Mik90, Mik91]. Charged [Hum96b, KC93, LLR94, MMW96, VB95, CPP93, KC92]. Charged-Particle [Hum96b]. Chebychev [RF96]. Chebyshev [BT92a, BT92b, BCM95a, BCM95b, BBMB97, Boy92b, Boy92c, Boy95a, Boy98, BSPL96, BE91, BE92a, BE92b, CJ95a, DVD93, DD90, DD91, Fra95, GMCH92, KI99, KK96, Kop96, KTE93, LMP92a, LMP92b, LRW95, MBV93, Mal96a, MMB90, NPC93, NC93, PBD94a, RK93, SD93b, Sen99, SK93, TKKB92, Tes92, WNY93, Wer95, ZY94]. Check [FK90]. Chemical [Ano94s, BN94, EGS96, Gou96b, LW96b, RC95, SPCD96, WH97, Shu91, Shu92b, Shu92c, SCC93a]. Chemically [Edw96a, GSB98, CYD92]. Chemistry [EGS96, FMO97, NWK98, OBL93, CYD92, GC90, Shu90, SLV90]. Christoffel [Yak96]. Circuit [VD97a, VAVB93, VAVB92]. Circular [AR96, DNN93, HD97, PM98, Ver97a, Ver97b, hYsY96, IT90]. Circulating [ALTP98]. Circulation [Bry97, HdS97, Sem97, SH94, UWH99, Waj93, WH98, HWY90, HWY91]. Circulations [LBB94]. Class [Bar94, BKV98, Kob99, LPMS94, Nor98, VLM98, VSM99, For92c, FMS92, ISW92a, ISW92b, KL97, SG92a, SG92b]. Classical [Dru99, FHL97, Rei99, Sch95b, SW99b, Yak96, BG92, Bas92].

Clebsch [GÖAY95]. Cloot [Boy94]. Close [Zin94a]. Closed [KP97b, KP98, SVS90, SVS91, Sen99, SCC93b, VD90]. closed-end [VD90]. Closed-Form [Sen99]. Closet [SW94]. Closures [BHJU99]. Cloud [DS97b]. Clouds [BF97, KPCJ97, Lan97]. Clouds-in-Cells [BF97, Lan97]. Clouds-in-Clouds [BF97, Lan97]. Cluster [HK93b, BGM91, Lif92, VG92b, VG92a]. Clustering [Ser96]. Clusters [LW96b]. Coagulation [SGW94, Lif92]. Coalescing [SD93a]. Coarse [EL96c]. coarsening [DS99, Mul92b, Mul92a]. Coastal [SR99, TG96]. Coated [ART95a, ART95b]. Coating [BD94a, CS97a, CS99, RLT93]. cochlea [Bey91, Bey92]. Code [BBF^{+}99, Bar90, BO96, BDML94, CDM98, CW93, FLD93, GBS^{+}93, GRB^{+}95, KC93, LPC^{+}93, MMW96, OS97, PFS98, Swe94, Swi96, Tay94, TLW95, Vu96, Vu98, Web98, AN90, GB90, GBS^{+}90, KC92, LP90, MS91b, MS92b, MS92a, MK90, Nic93, Poi91, Val90]. Codes [JLM^{+}96, NMW^{+}96, PSL99, TO98, Tót97, VD97a, VAVB93, Duk91, DM92d, DM92e, GBL^{+}92, Par90, Pri90, Pri91, VAVB92]. coditions [Bas92]. Coefficient [Boy95b, Boy97b, EL96c, Mot98, SR99, WNY93]. Coefficients [AIV99, Boy94, CS93b, CSW95, GÖAY95, SZ97, SS96c, VSM99, lX96, lX97, lX98, ML92]. Coil [KWG94]. Collapse [Man93b, MIM90]. Collection [Eyr95, TRL99]. Colliding [MKM99, MKM04]. Collision [JLM^{+}96, NY98, Nun93, AAP^{+}95, LBC^{+}93]. Collision-Free [Nun93]. Collisional [GGZ93, MK97, Dem92a, Dem92b, RD92a, RD92b, Rus90]. Collisions [MLJ97, WONM96]. Collocated [LL96, Taf95]. Collocation [BM94, BBMB97, CG93, CD98, DVD93, DS94, Göt99, Hei98, HDL99, Hol96a, KOS^{+}96, Kou96, Lif96, MBV93, MG97a, MS93, RK93, VPS95, VP96, VP97, WNY93, DD90, DD91, EHW90, EHW91, KR95, LMP92a, LMP92b, SD93b, UWSB90, UWSB91]. Colocated [Dor99]. combinations [LJT92a, LJT92b]. Combined [BMZ95, CF98, CF99, JVS97, dFD^{+}94, BS90, yGmC92, GC92]. Combining [COBA95, WVHS98]. Combustion [PC98]. Comment [RNSV91, Wer98]. Comments [Abd95, Boy98, Lif96, Per95, Ske98]. Comminution [Eyr95]. Common [Maz97]. Communication [JJK92]. Commutative [VLM98]. Comp [MKM04]. Compact [AS96, CGA93, CGA94, CF98, CF99, DM97b, Dor99, EL96a, Eka99, GS97a, GKZ97a, HKW97, oJ97, Lel92, Mar94a, Rav97, RF95, Sha99, Tol94, TL98, Yav97, Yee97]. Compact-Difference [Sha99]. Compact-Difference-Based [GS97a]. Compact-ENO [AS96]. Compacton [dLMSS95]. Comparative [Eye96, SS99a, SF96, YP92b, YP92a]. Comparing [LDB98, UWH99]. Comparison [BLMR92, Blu92, CDM98, CS97a, CE97, CJ95b, DOOSM99, FV97, GBL^{+}92, GKZ97b, HHF95, HOS93, LBB94, LD93, Mak90a, Mei98, NMW^{+}96, NM95, PB98, PMR97, PC99b, ST95, TZA96, TO96, WEMH99, YP98, FS91, FS92, Hal90a, Hal90b, LRB^{+}90, LBC^{+}91, PC92b, PC92a]. Compatibility [HW96]. Compatible [CBSW98, VK98]. Compensating [Jai97]. Complement [CD95b]. Complementary [FMO98, Ram97a]. complete [HG91, HG92]. Complex [AD97, Adh92a, Adh92b, BMGM95, CBL95, CP98b, CSZ97, DECB98, DL94b, FM97, GP98a, GM95a, GS95, GS96a, JMDB99, JT99, KMS99, Ku95, MGM95, Sch99, VLM98, YMUS99, dAT95, CYD92, CM91, CJR91, CJR92a, CJR92b, Dar90b, KvdVPG95, SD93b, SBGM92]. Complexity [Goe95, WA95]. Component [SCT^{+}99, Lar91]. Components [oJ95, oJ97, SSW98]. Composite [ACLW97, CH90a, Fat99, GO98, GS96a, HW97, PM93, SW99a, WS93, PM92]. Composites [Hel95]. Composition [Hum96a]. compound [LR92]. Compressibility [TZA96]. Compressible [BD96, BR97a, BLG97, BW98, CHL^{+}90, CS97b, CCG95, CMM93, ES93, FB94, FAH97, FMM99, GK94, GHKH93, Han93a, Han93b, HL98, Kop94, KK96, Kop96, Kop98, Lee98b, LQK98, LKK99, LBL98, PBC^{+}95, PG97, RG97, Sam97, SA99, SBGK99, SMT99, SMJ98, Shy98, Shy99, Sjö95a, Sjö95b, Tid97, TO98, Tou97, Tou98, WNY93, WP97, Wel98, vBKG97, vdVvdV98, BB90a, BB91a, Dar90b, Dar90c, Dar91b, DeV91, ES92, HW92a, HW92b, Lar91, MOS92, PPM92a, PPM92b, PL92b, Poi92, Tid95]. compression [Yan90]. Compton [BKP91, Win95]. Comput [CNG17]. Computation [AR96, BCRR98, BAB90, Ben96, BO92, BGP98, CS91, CKE99, DGN97, DW93, DF96, DS97b, FG97, Gue94, GÖAY95, HHF95, HRL99, IH95, JSD95, KM93b, LNR99, LJG96, LY90b, LAE98, LW97, LLT94, MM95a, Mas96, Mei98, MBI^{+}97, MS99b, Nak95, NO93, NMB94, Nor98, Pai97, PP93, Pet96, PBD94b, PS93b, PGD99, Pry94, RS94a, RP90, RG90, RP91, RH98, Sch96, SCR97, TT90, TT91, UMS99, VM96b, WS96, WL93, WKHS97, Yak96, ZYKC98, CKQ^{+}93, DK90, Dvo91, DK92d, DK92a, DK92c, DK92b, HCJ92a, HCJ92b, HL92a, HL92b, KKD90, LH90, LP90, PA95b, SMM^{+}90, SV91, WR91]. Computational [AAK^{+}93, AN98, ARV99, Ano99r, Ara97, AK94b, BH99, CAA93, CM94, FH97, HHM96, nJWP96, Kel94, KM96, Lil97, MW97, Mar94a, MLBW97, MBP94, Nac96, ÖL96, PC98, QG98, Rie99a, Sam97, SSB^{+}99, Sch99, SH98b, TW93, TM94, VNC92a, VNC92b, WF99, Xia99, YDT93, ZHL96, AA90, AA91, Bey91, Bey92, Fau90, FS91, FS92, Gre90, NSR92a, NSR92b, YSG91, YMB^{+}91]. computational-mode [AA90, AA91]. Computations [BL99, Bar91a, BCM99a, Boy94, Boy96a, Boy96b, Boy97a, Boy97b, BCR99, CSW98, DZ96, Ena93, FCC97, Har94, HR96, Hwa96, Jak93, KC97, ST95, Str99a, TCS97, Tsy95, YH95, CW92c, DHS91, GMP92a, GMP92b, LL91, PPM92a, PPM92b]. Compute [CL93]. Computed [Shi93, AF90, AF91]. Computer [BGGT90, CR96, DM90, DM91b, JZQ^{+}95, KWG94, MB99, CYD92, MNR90, VY91, VY92]. Computers [BGB99, DB98, DLMN91, DLMN92b, DLMN92a]. Computing [AK99, BCM95b, BW98, Bör97, Boy95b, Cho93a, Edw96a, Fen99, For97, HAC97, Hob93, HKS98, KM97a, LLD93, Mar97, MOS92, OS90, PM92, PM93, SW96, SSO94, Vra95, WSK99, Zho96, vD97b, BP90, CPP93, Hob92, Hor90, Hor91, LLD92, Lar91, Mor92, VBB90]. Concentrated [Zin94b]. Concepts [Mat97]. Concurrent [FLD93, Swe94]. Condensed [MP96]. Condition [EL96b, HS93b, JB93, PPBC93, PZC98, SDI98, TW94, TAC98, WNY93, Wan99, AN90, Dar90c, Dar91b, HLB94, PPBC92, SC90]. condition-free [AN90]. Conditional [dNPT95]. Conditioning [NPC93]. Conditions [APV98, AE98, BH96, BPT95, BC99, BJ95, Boy98, CGA94, CJR95, Clé96, DR96, DNN93, FH94, Fer95, GN96, GC95, Gri94, GK95, GK96, GK98, GZ95, GS97b, Hu96, Ji94, Joh93, KLP94, Leb99, LT96, MB94, Nic99, NC99, PNC94, Pet98, PH95b, RF96, SY97, Set94, SI94, SWD95, TG96, Tou97, Tou98, Tsy95, Zin94a, BG92, BR92c, BR92d, CP90, DM92c, Fer90a, Fer90b, Giv90, Giv91, HTK90, Hig92a, Hig92b, JW90, LG94, PL92b, Poi92, RR90, RR91, Ren92a, Ren92b, Rom92a, Rom92b, Sal91, Tho90, dG92a, dG92b]. Conducting [ARTAAA97, Bis95, HR96, AN90, Poi91]. Conduction [CM99b, HIY95, LEI96, Liu96a, SCM98, ZWS98, dG96]. Conduction/Radiation [dG96]. conductivities [Yor90]. Conductivity [LEI96]. Conduits [SCC93b]. Cone [LB93]. Configuration [Gre98, OGBSK90]. Configurations [KK95, KCJ95, LH92, Nic93]. Confined [AM96]. confirmal [Dar90a]. Confluence [Sch99]. Conformal [BST95, DE93, Hum96b, KM96, Dar91a, LM92]. Conformational [SFHD99]. Conforming [ADS96, KT99]. Conjugate [ANL94, BL98, Che91, MBP94, Dav91, Dav92a, Dav92b, NP92]. Conjugated [PNC94]. Connected [KM95, vGVO98, BW90, BW91, GGM93]. Connecting [LLT94]. Connection [LK94b]. Conquer [WDE98]. Conservation [Anc94, BLK90, Bih96, BT97b, CBSW98, CG97a, Cha95, CWC99, CL98a, CS98b, CK93b, DW95a, DW96b, FB94, FAH97, Fur99, Han93b, Har97, Jin94, Jin95, JL96a, Lav93, Lax97, LWC93, LR94, Noe94, PLD97, PQ94, PP93, Sal94, YPH94, BWJ90, CGSS90, CGSS91, Col90, Dur91, DEO92, Har92a, Har92b, LO91a, LO91b, LY90a, MWJ91, NT90, RT92, SW92, Shu92a]. Conservative [Abg96, ABC^{+}98, BTW96, BCDL97, BC98, CNG99, CWG90, CWG91, CJ94, CDW99a, Epp94, FKM99, Gra99, HS93a, Hir97, HB97, JMT97, Jin94, KPP99, KK96, Kop96, KNR99, LW96a, MLVM98, Smi99, ST99, TLW95, TPRC96, Wan95, vL97c, CNG17]. Conserving [Oku95, Ruy93, SVV98, Flå92b, Flå92a]. Considerations [Bar94, MBI^{+}97, SC96b]. Considering [JM98]. Consistency [dNPT95]. Consistent [Coo99, Kar94, KTN^{+}94, LW94, MJPC99, Swe96, Waj93, DHS91]. consistently [HHG91, HHG92]. Constant [BLL99, EOS98, She95, Sor95, But90a, GCMR90]. Constants [BCM99b]. Constrained [CK93b, Jai97, LS94, MB92]. constraints [HM90]. Construct [MS99a]. Construction [AS99a, AD93, BWJ90, CBSW98, CG97a, Gou96b, Hes98, MWJ91]. Contact [Had99, Lav93]. Contact-Line [Had99]. Containers [KP97b]. Containing [CLO93, LO93, OB95, Ram92a, Ram92b]. Contaminant [GP98a]. Contamination [RHT96]. Context [Bot98]. Continuation [CMW95, NCF96, Nic98, vD97b, SK90]. continuity [SA90b, SA91]. Continuous [GEK^{+}97, PS93b]. Continuum [CW98, Had99, LCM96, RMO99, WF99, BKZ92]. Contour [BFW98, Dri97, LD93, VM94, VM97, ZHR97]. Contrasts [VSM99]. Contribution [IS96]. Contributions [WL93]. Control [BD97a, Bra93, Eça96, FT96, GB97, HR96, IR98, LB94, MG97b, OK94, Glo92, HL90, HL91, TA91a]. Controllability [BGP98]. Controlled [BD96]. Convection [BY94, BDS^{+}99, CZS98, CT98, GKZ97a, HOS93, HO96, Hwa96, nJ93, Kno98, KB99, LSLG97, OGWW98, Rig94, SW96, TC98, XMP97, Yav97, AM90, Fog92b, Fog92a, LMP92a, LMP92b, LT91, LT92, LE90, LE91]. Convection-Diffusion [CZS98, GKZ97a, HOS93, Hwa96, XMP97, Yav97, AM90, Fog92b, Fog92a, LE90, LE91]. Convection-Dominated [OGWW98]. Convective [CMMH94, CGY93, CL98b, HI98, Li97, Mal96b, MWW96, CGY92, JC92, Sal91, dG92a, dG92b]. convective/radiative [Sal91, dG92a, dG92b]. Convergence [Azm99, BR94, Boy94, Eye96, Löt94, SPC93, Ven95, Web98, Arm94, HHG91, HHG92, SPC92, VBB90]. convergence-improving [VBB90]. Convergent [HOS93]. Conversion [Bel97]. Convex [LO98]. Convolution [RMO99]. Convolution-Generated [RMO99]. Coordinate [CKR93, DM93, Hig99, HLL99, HC93, PFS98, Pri95, SH94, Zha94, BRL91, BLMR92, BM90b, BM91c, Cal90, Cal91, EHW90, EHW91, LT91, LT92, LH92, RKV90, RKV91, SKK90, SKK91, SH92]. coordinate-space [BLMR92]. coordinated [Blu92]. coordinated-space [Blu92]. Coordinates [BD97b, BW98, HD97, Jor99, MM95b, MS98, NGN97, Ruy93, SNU98, SW99a, STS98, VO96, YP98, ZSK94, CLJ^{+}90, HWY90, HWY91, MDA91, RP90, RP91, SV91]. Coriolis [Cod99]. Corner [AAL97, FC95]. Corner-Balance [AAL97]. Corners [DE93]. Corona [Fen99]. Coronal [LMS98]. Corrected [BB97b, BK97b, För96, GMM99, SS94, TO96, Zal97, ZWS98, DeV91, Gra90]. Correcting [AK94b]. Correction [Boy97a, JMT97, SS96a, TW94, Wer95, AB94, Kor90, Luc90, Luc91, Mul92b, Mul92a]. corrections [BS90]. Correlated [RS99b]. Correlation [BK97b, DD92a, För96]. Correspondence [OK94]. Corrigendum [Ano90o, CNG17, MKM04, VSG96]. Corrugated [SF99]. cosmic [MMR90]. Cost [FBM93]. Couette [Bar91a, CRRR97, Kup98, PC99b]. Coulomb [BFO99, GS94, GS97b, JLM^{+}96, Kou96, MLJ97, MBF94, NY98, SH99, WONM96]. Coulomb-Type [GS97b]. counterexample [Gre90]. counting [LS92c, LS92d]. Coupled [AK94a, Cra94, GKKR99, Gou96a, HKM96, Her93, HK93b, JSD95, LNR99, LZ96, SS96a, Sco96, STW97, SA90a, SL96, TS99, IM95, SCC93a, VNC92a, VNC92b, ZS92a, ZS92b]. Coupling [Ari97, AC93, BMZ95, BLT96, CLO93, CM99a, Clé96, LM97, Tiw98, YG95, ML92]. Courant [AN90]. Cover [CS97a]. Covering [BS99d]. CPA [BFGG94]. Crack [ARV99]. Crank [Alv91, Alv92]. CRAY [WKHS97, Par90]. Criteria [Lee98a, MW93]. Criterion [Sco98]. Critical [GS96a, Cha90, PRMV90]. Cross [EOS98, HC97, Hel95, hYsY96]. Cross-Section [EOS98]. Cross-Sectional [Hel95]. Crystal [AT92, BCRR98, HKW97, RT94, SS91, SS92, YR93]. Crystal-Crystal [HKW97]. Crystals [Dob99, FG97, LW96b, PCLC97]. Cube [Rea97]. Cubed [RIP96]. Cubic [Eyr94, RFH93]. Cumulative [Ano95t, Ano96w, Ano97-28, Ano98-27, Ano99-28, NY98, WH97]. Curl [BD97b, Mad95]. Current [FH97, FGM97, FCC97, Mar93, Mat94, MQS95, TSW95, vL97a, vL97b, LHI90, ND90, TA91b, TAD92a, TAD92b, XPK90]. Current-Voltage [MQS95, XPK90]. Currents [ARV99, Fen99]. Curvature [BW95, Cho93a, HKS98, NPV96, Ruu98b, Set94, vM95]. Curve [BW95, SR92b, SR92a]. Curved [Hew97, MLS99, hYsY96, SD93b]. curves [SR90]. Curvilinear [BD97b, HD97, Jor99, MS99a, MS98, YP98, ZSK94, Zha94, vBvNW95, BK91, BK92a, BK92b, HWY90, HWY91, LT91, LT92, SV91]. Cutting [TKT97]. CVD [TH96b]. Cyclic [Mat94]. Cyclo [CO95]. Cyclotrons [oJ97]. Cylinder [AR96, DQ95, HD97, HC93, JYH93, SB96a, IT90]. Cylinders [MB96, CG97c]. Cylindrical [BFW98, LMS98, LS98, MV99, PL93, Pri95, Ruy93, VO96, dG96]. Cytidine [LLLY99].

D [Ano99-30, Dvo91, DK92d, DK92a, DK92c, DK92b, MKM04, TA91b, TAD92a, Mar90, BP90, BBF^{+}99, BR97b, BM91a, BM91b, Bor91b, Bor92a, Bor92b, BIAV98, BDML94, COS^{+}98, CDM98, CPB96, CC98, Clé96, DH96, Dau92a, Dau92b, Dob99, EL97a, Eça96, Efr97, FLD93, FWBS94, FK97, FG97, FPB98, FQ96, GNHP95, GBL^{+}92, GM96, GHS93b, GHKH93, GS93, HKS98, Hu95, ID96, IM95, Ji96, KKH^{+}99, Kan92, KL92a, Kar97, KPCJ97, KCJ95, KH97, KO98, LLD92, LLD93, LMP92a, LMP92b, LL95, LQK98, MKM99, Mat94, Mau99, MW94, Nor96, OS97, Oos97, PK94, PH94, Pet96, PV99, PV93, RF96, RS94b, Sal94, Sha95a, SS99c, Str96, Str97, TZA96, TH96a, TLW95, VAVB92, VAVB93, Wal94, Wes92a, Wes92b, Xia99]. DADIPIC [GH95]. Damage [Maz94]. Damaged [BH99]. Damped [TWV97]. damping [Fri90]. Dancoff [MP93]. Dangers [BS93]. Darwin [OS97, DHS94, GH95, GH97, HLD92b, HLD92a, SAB95]. Darwin-EM [OS97]. Data [AK98, Ben95, BS96, Eve96, oJ95, LWS98, Lya99, MD95, Nor95, OG97, PBD94b, PGD99, WEMH99, DT90a, DT90b, GWI92]. Data-Dependent [OG97]. Data-Parallel [BS96]. Davidson [BO98, Mor90, Mor92, vD97b]. Davidson-Type [BO98]. Deaminase [LLLY99]. Decay [JZQ^{+}95]. Decaying [BC98, DM92c]. Decomposition [AC93, BGM95, BD97a, Ben95, DZ96, DK96, Dvo91, DK92c, HP99, LO98, MQS95, RG97, RF95, SR97, SRA98, Tiw98, Chi91, Chi92a, Chi92b, DD90, DD91, DK92d, GHS93b, SD93b, Tid95, Fey98b]. Deconvolution [Rob99, Rei92a, Rei92b]. Decoupling [SK94]. defect [AB94, Kor90, Mul92b, Mul92a, SVS90, SVS91]. defect-correction [AB94]. deferred [Luc90, Luc91]. deferred-correction [Luc90, Luc91]. Defined [KKK95]. Defining [Nic99]. definite [SG90]. Deflagration [FAX99]. Defocusing [Osb93]. Deformable [CS97a, CS99, CT98, SK98]. Deformation [ALTP98, Bot98, KWG94, LSLG97]. Deformations [CKR93]. degenerate [FP92]. Degree [CK96]. Delaunay [Her93, Mav95, Reb93]. Delta [Iva93]. Delta-Matrix [Iva93]. Deluge [Due93]. Delves [Boy97b]. Dendrites [Sch96]. Dendritic [JT96, WS96, Alm93, AT92, SS91, SS92]. Dense [WL96, CG97c]. Densities [SRVK96, EMRS91]. Density [ABC^{+}98, BCM99a, CDT98, EL97b, Gag98, PAB^{+}97, Rom97, Ruy93, SBGK99, Val91a, Val91b, WKMH97, BM92, BES90]. Density-Conserving [Ruy93]. Density-Matrix [BCM99a]. Density-scaling [Val91a, Val91b]. Dependence [Ben95, CK93a, Kal97, CK92]. Dependent [AN98, AIV99, BY94, CBL95, GG97, GP98b, GK96, HG94, HK94, JT93, LP97, NC97, NPV96, OG97, OP97, Rom97, SB95, SNU98, Sha99, SWD95, TCS97, ZSK94, CH90b, CH91, Dra90, DLMN91, DLMN92b, DLMN92a, LHI90, LRB^{+}90, LBC^{+}91, SM92, Sto92, TEKC92, Tho90, Wat92b, Wat92a, XMP97]. Deposition [AS95a, AS95b, AS97, EGS96, LPR96]. Depth [BY94, KS92]. Depth-Dependent [BY94]. Derivation [Abe97, Car91, MDA91]. Derivative [Cod99, CSW95, MJP93, MWW96, MS93]. Derivatives [BCM95b, GNH96, IH95, KTN^{+}94, STG^{+}96, YG94, BE91, BE92a, BE92b, RF90, RF91]. Derived [LPP99]. descent [HB90a, HB91]. described [MDA91]. describing [Nic93, TMR92a, TMR92b]. Description [RWTC95, VBD99]. Design [Ara97, Ari97, BB97a, ES98, IS96, JB95, Lee98a, Lee98b, LWT99, Lil97, SCR97, YZ98, FS91, FS92]. Designed [Mav95]. Designing [GO96, GO98]. Despite [Nor95]. Destabilizing [Bec94]. Detailed [EGS96]. detecting [MH92]. Detection [ARV99, Loe93, MML98]. Determinant [Wit96b]. Determination [CG97b, EOS98, MM98, SF99, LH92, LP90]. Determining [BH99, MG96, MSV98]. Deterministic [Ber95, Nor96, Rus93, Tad97]. Detonation [ABS96, FAX99, ZDD99, CKQ^{+}93]. Developing [AK97, LWS98, Bue91]. Development [KP97b, KCV99, SWT98, Swe94, YSG91]. Developments [vL97a, vL97b]. Device [KL93, LWT99]. Devices [FO93, Fen99, MQS95, PL98, XPK90]. DFT [WDE98]. Diagonal [Ano94t, ESM98]. Diagonalization [CG97b, YS98b]. Diagram [PJ94]. Diagrams [HK93b]. diatomic [Dav91, Dav92a, Dav92b, KEHK90, KKD90, KEHK91]. Dielectric [ARTAAA97, GGK95, GKG95, Rob90b]. Dielectrically [ART95a, ART95b]. dielectrics [AP91a]. Difference [Bai97, CGA93, CGA94, CO95, CJ94, CJS99, CGY93, CM95, CF98, CF99, CKR93, CDW99a, CL98b, DW98c, DDS97, EL96b, EL97a, EL97b, Epp94, FO93, Fat99, FMO98, FWBS94, FM97, Fur99, GS97a, GO96, GO98, GN96, God99, GO95, HT95, HOS93, HMOS95, Hir97, HWM96, HW96, Hui90, Hui91, HKvV95, Hwa96, oJ95, oJ97, JcW99, JT93, LBB94, Li97, Lin95, LL93a, Luk99, Mah98, kM93a, Mat97, MS98, MLVM98, NC97, NC99, PZC98, PBD94a, Rav97, Rig94, Roe97a, Sak96, SZ97, SPC93, Sco98, SA95, Sha99, SS95, Sjö95a, SM98b, TW93, TW94, TN98, Van99, VO96, VSG95, Weg97, XMP97, YL98, ZLOT98, ZC94, Zho98, ZL93, dLMSS95, vL97c, Ano95u, AP91a, BB90a, BB91a, Ben90, BY90, BY91, Car91, Cas90, CGY92, Dic90, For92c, FMS92]. difference [HT94, HGH90, HGH91, Lay90, Lel92, kM90, kM91, Mao92b, Mao92a, MH90b, MH91, MLB92, MMS90b, MMS91, NSR92a, NSR92b, RM90, RM91, Ren92a, Ren92b, Rob90a, SPC92, Str94, ST92b]. Difference-Based [MS98]. Differences [CO95, MQS97, Gla92b, Gla92a]. Differencing [MWS96, Sil93, Tol94, TL98, Fig92, HSN90, MDHW92a, MDHW92b, NT90, Shu91, Shu92b, Shu92c, Sil92]. Different [Gel99, Mei98, NMW^{+}96, LRB^{+}90, LBC^{+}91, Mak90a, Tid95]. Differential [BK97a, BPM98, CZS98, CGKC93, CD98, HKM96, HRR94, HS94, MQS97, OD99, Pet94, PMR97, QW93, RIP96, Tei99, VPS95, VP96, WD97, BB91b, BBF92a, BBF92b, Car91, CH90a, CGKC92, DM90, DM91b, FVZ90, HGH90, HGH91, HLD92b, HLD92a, ISW92a, ISW92b, KL97, LK90, MS90b, MS91c, Wat92b, Wat92a, YSG91]. Differentiation [BCM95b, GNH96, STG^{+}96, Sol92]. Diffractive [HDL99]. Diffusing [LR90, LR91]. Diffusing-vortex [LR90, LR91]. Diffusion [AD97, AIV99, BYMZY95, Bec94, Boy98, BBGL93, CZ98, CBS93, CGY93, CC97, CZS98, EP96, EM94, FD93, FPQ93, Gra95, GKZ97a, Hag94, Han93b, HOS93, HMOS95, HO96, HWM96, HGFH95, Hwa96, HSS97, KD97, MRS98, OBB98, OP97, PA95a, PBD94a, RKO99, RK99, Rig94, RT94, Ruu98a, Ruu98b, SvD96, SS96c, SWT98, SL96, Tei99, TN98, Wan97, Win95, WK99b, XS93, XMJ95, XMP97, Yav97, AM90, Bor91a, BLA92a, BLA92b, BB90b, BB91c, CW92a, CW92b, CGY92, CYD92, Fog92b, Fog92a, GM92, Lay90, LE90, LE91, MH92, MMS90b, MMS91, MB92, MDHW92a, MDHW92b, DW98b]. Diffusion-Convection [Rig94]. Diffusion-Generated [Ruu98a, Ruu98b]. diffusion-type [GM92]. Diffusive [FWBS94, MP97, nJ93]. Diffusivities [Rie94a]. Diffusivity [HB95, OP97, HCP93]. Dilute [BMGM95]. Dimension [CK96, PLD97, Dar92, MRS94, RT92, ZB92a, ZB92b]. Dimensional [AS95a, AS95b, AR93, ACLW97, Ara97, AD93, ADH^{+}93, BY94, BYMZY95, BGB99, Ber96, BO96, BS98b, Bro95, CS93a, CZ98, CS99, CA93, DW97, DR96, DW93, DIV98, DF98, DL94a, EO98, ES93, EGS96, Eyr94, FAH97, FKM99, Fis94, FJGP94, FCC97, Gel99, GG97, GB97, GM98, GLN^{+}99, HOS93, HKNT98, HK94, HB97, HLL99, IH95, IS96, JH97, KKK95, KT99, KPC93, KKOF97, LO96, LS96, LWT99, LJG96, LPC^{+}93, Lil97, LMB94, LL94, MD95, kM93a, MM96, MB97, MLBW97, NMB94, NGN97, Pai97, PB98, PBBS95, PDHS94, PZ98, Pri93, Pry94, RS94a, RCA93, RKJ97, Rig94, RHT96, Sch96, SLMS98, SVV98, SM98a, SBC99, Sim93, ST93, TSW95, Tan94, Tei99, TN98, TO98, VD97a, Ver97a, Ver97b, VO96, VS95, Vu98, Wel98, WSK99, Wha96]. Dimensional [WL93, WWW95, YYCH98, ZLOT98, ZB96, Zha98, ZZ98, ZYKC98, ARS91, AS90, AS91, AF90, AF91, BRR92, Bar91b, BKM90, Bil90, BM90a, Bor91a, BLA92a, BLA92b, Bue91, CS92a, CW92a, CW92b, CSW98, Cas90, CPJ90, CYD92, CKQ^{+}93, DGD90, ESM98, ES92, Fig92, Fog92b, Fog92a, FVZ90, GM95b, Gre91, Gre92a, Gre92b, yGmC92, GC92, GMCH92, HSN90, Hoo90, Hor90, Hor91, HL90, HL91, Kle95, KG90a, KG90b, KG91, KPC92, LS92a, Lav90, LG90, LG91, LHI90, LPvL93, LS92c, LS92d, LO98, Mao92b, MB92, PPM92a, PPM92b, RCA92, RH91, RCAM92a, RCAM92b, RG90, RIW90, Rus90, SC90, SH92, SBGK99, SBGM92, SWT98, ST92a, SA92a, SA92b, Val90, VMK91, WEQ^{+}99, XMP97, YR98, YTS92, vdVMK92b, vdVMK92a]. Dimensions [ABC94, AFH94, BLP98, Dri97, Nor98, Tan93, Wit96b, YG94, ZHR97, But90a, CGR99, DZ91, DP91, DP92b, DP92c, MF92, Rok90, SKK90, SKK91]. Dimer [BS99d]. Diminishing [TO96]. Diode [CCGJ95]. Dip [RLT93]. dipolar [CPP93]. Dirac [Alv91, Alv92, CK90b, CK91]. Direct [BK93, BCT98, CBL95, CB96, CR96, COBA95, DGN97, DB96, DE99, GBCA99, GH95, GH97, JMR94, KPC93, MM95a, Mau99, MB96, Pal93, RM90, RM91, RM93, SFHD99, She95, Sho93, SF99, BO92, GL96, KPC92, Lif92, Mar90, Osb90, Osb91, PL92b, Poi92, PA95b, Pri90, Pri91, PO90, PO91, RCAM92a, RCAM92b]. Direction [LT96, MWW96, RY94, HM95]. Direction-Adaptive [LT96]. Directional [Bra93, YZ98, ZWS98]. Directions [BM94, Cle97, MBV93, RL97, SMR91]. Dirichlet [GGM93, HPG98]. Dirichlet-to-Neumann [HPG98]. Discharge [Hag98, Kul95, LW94, YPH94]. Discharges [SG94]. Discontinuities [FAX99, Lav93, kM93a, kM90, kM91, Mao92b, Mao92a, RLM90, TP92a, TP92b]. Discontinuity [HML99, LK94a]. Discontinuous [BR97a, BR97b, CS98b, CSW95, HHR99, LW95, LKK99, MWS96, OBB98, SHF97, TSW95, WK99a, vdVvdV98, Bor91a, BLA92a, BLA92b]. Discrete [Abe97, AE98, Bec94, Edw96b, GACN93, Hag98, HW96, JR99, KLBD93, KI99, Kla99, KTN^{+}94, Mad95, NP94, PC99b, SY97, Sie99, SDI98, Tou98, TM93, WDE98, Arm94, HLB94, MGP91, MS90a, MS91a, Nad95, RNSV91, Sch92b, Sch92a, SA90b, SA91, Str91, Str92a, Str92b]. Discrete-Ordinates [GACN93, Sie99]. Discrete-Velocity [NP94, Nad95]. Discretisation [CSW95, Gil97]. Discretised [HO96, IABBG91]. Discretization [ABBM96, BKV98, COS^{+}98, CDF95, CS92b, CS92c, Cod99, CD95b, GN96, MRS98, RB95, RL95, SVV98, XMP97, vBvNW95, vGVO98, Bor91a, BLA92a, BLA92b, Gro91, Gro92, NSR92a, NSR92b, YSG91]. Discretizations [AHS96, GKZ97b, HS99c, MW97, MS99a, OGWW98, SP99, Tid95]. discretize [BK91, BK92a, BK92b]. Discretized [BH95, IGABB90]. Disks [Sch95a]. Dispersion [DS97b, FD99, HHM96, ST92a, ST93, SH98b, TW93, Zin94b, TMR92a, TMR92b]. Dispersion-Relation-Preserving [TW93]. dispersions [CG97c]. Dispersive [AQ98, Luk99, CWG90, CWG91, FM92, Fra92]. Dissipation [Fur99, HHM96, Lin95, SMD98, SH98b, SRT98, TYS95]. Dissipative [GA95, Leh99, WVHS98, YSD99]. dissociation [KKD90]. Distances [HA93]. Distortion [CS98a]. Distributed [SZ97]. Distribution [Eça96, GA98, LKLE90, LLEK91, Osc94, CK90b, CK91, HN95, MS91b, MS92b, MS92a]. Distributions [AH99, OP97, WKMH97]. Disturbance [MLBW97]. Disturbances [CRRR97]. Div [BD97b]. Div-Curl [BD97b]. Divergence [DL94a, Hil97, MBV93, Mad95, SP98, van93].

Divergence-Free [DL94a, MBV93, SP98]. Divergent [BDML94]. Divertor [Kno98, STS^{+}97]. Divide [WDE98]. Divide-and-Conquer [WDE98]. DNA [BH99, JMDB99, JVS97]. DNL [SDI98]. DNLS [Flå92b, Flå92a]. Domain [AK95, Ano94t, AC93, BGM95, BD97a, Bis95, Clé96, CJM97, DZ96, EOS98, Hag94, HKNT98, HDL99, Hwa94, KM95, MQS95, ÖLJ98, RG97, RF96, Rob99, RF95, SSR96, SA95, Sha95a, Sha96, SR97, SRA98, TS96, Tid95, Tiw98, VPS95, VP96, Wan99, YL98, Chi91, Chi92a, Chi92b, CH90b, CH91, DD90, DD91, GHS93b, Han93a, HPS92a, HPS92b, HJ98, Kan92, KL92a, LM90b, SD93b, TKKB92, Tes92]. Domains [BS98a, GPWZ96, JC98, Kel94, KW93, LKK99, MM97a, Mau99, PC96, PK95, TAC98, Tau94, TM94, vGVO98, CW92c, CJ95a, GGM93, LM90c]. Dominant [CDF95]. Dominated [AE95, KRGv93, OGWW98]. Don’t [Bar90]. Double [BES94, DÜ92a, Ji96, SMW99, Weg97, Wit96b, DÜ92b]. Double-Adiabatic [Weg97]. Double-Gyre [SMW99]. Double-Well [Wit96b]. doubling [NSR92a, NSR92b]. down [GGP97]. Drag [Mon97a]. Drift [GBS^{+}93, SCG94, Smo98, TN98]. Drift-Diffusion [TN98]. Drift-Kinetic [GBS^{+}93]. Drilling [GBBH96]. DRISM [LPP99]. Driven [MWM94, MRB94, SMW99, SG94, Vu96, Vu98, vM95, GGH90, MLB92, She90, She91]. Droplets [BMJ^{+}97]. Drops [PBBS95, ZMOW98, PBBS91]. Drying [FT96]. DSMC [Pri90, Pri91, Rie99a]. Dual [Osc94, MH90b, MH91]. Duality [DÜ92a, DÜ92b]. Duct [AK94a, FB94, BAB90]. Ducted [TAC98]. Due [AAK^{+}93, AR97a, Cor96, FCC97, Ske98, Wri98]. during [TH96b]. Dusty [Mon97a]. Dynamic [CL98a, Eyr95, FLD93, JVS97, KNR99, LB94, LPC^{+}93, MKM98, MKM99, MKM04, NPV96, PGD99, BBK90, BBK91, HLD92b, HLD92a, TKKB92, Tes92, Ver91]. Dynamical [LLT94, RG97, YSG91, YDT93, MMR90]. Dynamically [GDP96, Rie94a, VP96]. Dynamics [ABS96, BFW98, BGGT90, BC99, BN94, BLL99, CV97, DOOSM99, DM93, Dri97, DL94b, DMW99, Eka99, ESH93, EH93, GKDT96, GKL^{+}96, GBS99, HRL99, Jai97, KSB^{+}99, Kho98, Kla96, KT95, KKOF97, LPP99, MNP90, MW97, MS99a, MRC93, MSF94, Maz97, Maz94, Mon97a, PC99a, Pli95, RC95, Rap93, Rei99, RFL93, RMO99, SMD98, SS99b, SMSS98, SFHD99, SW99a, SHF97, SW99b, UKSTST97, VM94, Vie94, VM97, WH98, WVHS98, XH98, YCTC97, ZHR97, dM97a, vL97a, vL97b, Bil90, BBK90, BBK91, BM90a, Boy90, Boy91a, BGM91, JVR93, Kop90, Kop91, LRJ^{+}99, MH92, MM91a, MOS92, YSG91, YMB^{+}91]. Dynamo [KB99]. Dyson [För96].

Earth [JSD95]. Earthquakes [PM99]. easily [dFSS92]. Easy [HK93b]. Economization [Boy95a, LRW95]. EcoRI [JMDB99]. Eddy [ARV99, Coo99, DFN^{+}99, GMS^{+}99, GM95a, Gho96, Jor99, KM97b, LIH97, LHI90]. Eddy-Resolving [LIH97]. Edge [AR93, KPC93, KM95, VS95, Zan97, KPC92]. Edges [PH95a]. Edinburgh [Swe94]. Editor [KHW98, Wer98]. Editorial [Ano99-29, Bla95, Bra96a, Bra96b]. Effect [BJ95, DM92c, KM97b, LSLG97, RHB94, MS90b, MS91c, RCAM92a, RCAM92b, YHW90, YHW91]. Effective [APV98, Bel97, BH99, FH94, GG92b, Hel95, LK94a, WWW95]. Effects [BDC97, CM94, Duk95, GGZ93, HO96, JM98, JL96b, LK94a, Maz94, Rob99, YS98b, Dem92a, Dem92b, TPL90]. efficiency [HN95]. Efficient [AM96, AK98, AK99, BD94b, BSPL96, CC97, CZS98, DT90a, DT90b, DB97, DECB98, Dob99, EO98, ESH93, EH93, FMO98, GO96, GO98, Gou96a, Han93a, Hob93, JS96, JR96, KM93b, KKOF97, KS92, LPMS94, LS98, LLK98, LPR96, MV99, MRC93, MBP94, NC97, Nor98, ÖL96, Pry94, Reb93, RH98, Ruu98b, SH99, SPCD96, SMJ98, Shy98, SW96, Vra95, VSM99, Wan99, WLC96, lX98, YL98, ZCH96, Zin94b, BJ90, Dol92a, Dol92b, Hob92, JJK92, Kos92, LDB96, OGS90, OGS91, Par90]. Eigenequations [HKW97]. eigenfunction [KBS90]. Eigenfunctions [BK94, SW96, BS91a, Mit92]. Eigenmode [FH98b, Sch92c]. Eigenmodes [BK99]. Eigenpairs [DA94]. Eigensolutions [BSPL96]. Eigenstates [CG97b]. Eigenstructure [CDT98]. Eigensystems [WL96]. Eigenvalue [Bar94, Boy95a, Boy96b, Boy97a, DA94, HS94, Ji94, Ji96, KEHK90, MDB91a, MDB91b, NO93, OD99, BDR91, KEHK91, LP90, Mor90, PRS90, SK92]. Eigenvalues [BCRR98, BO98, DDD98, KM93b, NO93, SW96, WKHS97, vD97b, KKD90, MMB90, Mor92, MRD92, OS90]. Eigenvectors [DS96, MRD92, OS90]. Eighth [Jam99, ST97, ARS91]. Eighth-Order [ST97, ARS91]. Eikonal [Ben96]. Elastic [Ben95, CV97, JLL97, Kla96, PP93, ZS97, Hig92a, Hig92b, Sch92c, TP92a, TP92b]. elastic-plastic [TP92a, TP92b]. Elasticity [NMB94, GKM96, OB95, TKKB92, Tes92]. elastohydrodynamic [Ver91]. Electric [DDS97, JB91, JBvK^{+}91]. Electrical [HP99, RCA93, Yor90, YPH94, RCA92]. Electrically [HR96]. Electro [LZG99]. Electro-migration [LZG99]. Electrode [SG94]. Electrodynamics [Cra94]. Electrolyte [KB96]. Electromagnetic [AL97, ART95a, ART95b, ARTAAA97, AR93, ADS97, Ber96, CD95a, ETRW95, GRB^{+}95, KI99, PSL99, Sha96, SBC99, Tan93, lX96, lX97, YGH97, YP98, AN90, Ber94, BBC^{+}90, BBC^{+}91, GLM^{+}90, HTK90, KMB^{+}92b, KMB^{+}92a, Poi91]. Electromagnetics [ABLD94, nJWP96, Nac96]. Electron [FO93, KPCJ97, Kul95, SH99, Sco96, Vu96, AAP^{+}95, KDL^{+}92, LKB90, LL92a, LL92b, MM91a, PB91]. electron-molecule [AAP^{+}95]. electron-repulsion [LKB90]. Electronic [Fat99, Goe95, KKOF97, PDS99, WCSW99]. Electrostatic [FPB98, GGK95, GKG95, JLSM97, MBI^{+}97, WTH^{+}96, CG97c, Hor91, OsKB^{+}92]. Electrostatics [Hel96]. Element [ABLD94, Ano99-30, ADH^{+}93, ADS96, AK99, BD96, BYMZY95, BR97a, BR97b, BH98, Boy98, CLO93, CN95, CLNT98, CHR99, Cha95, CWC99, CD95b, DK98, Dur93, EP99, FA97, Fen99, FT96, Fis97, GMM99, GK94, Gir98, GP98b, GPWZ96, GGK95, GKG95, GHKH93, GDP96, HG94, HIY95, Han93b, HK95, HW97, Hum96a, IH95, nJ93, KKK95, KD97, KT99, Ku95, LW95, LBB94, LTE93, LEI96, LIH97, LK94b, LO93, Ma93a, Mat97, MA96, OBB98, PK95, PC99b, Poz99, Rea97, SSR96, SK93, SCR97, SL98, TSW95, TTI97, VS95, Wan97, WJTP96, YMAC99, YPH94, Zan97, ZW97, ZOI97, vGVO98, vdVvdV98, AAP^{+}95, Bor91b, Bor92a, Bor92b, CP90, DD90, DD91, FP91, GL95, yGmC92, GC92, HGH90, HGH91, HM90, KR95, KG92a, KG92b, LM90b, LE90, LE91, LG94, Mit92]. element [MB92, Nat90, Nat91, Nat92a, OB95, OGS90, OGS91, PTM92, SAB95, SNL90, VMK91, YMB^{+}91, ZB92a, ZB92b, vdVMK92b, vdVMK92a]. Element-FCT [GK94]. Element-Finite [Dur93]. Element/Asymptotic [ZOI97]. Elementary [AK95]. Elements [AR93, BD94b, BMZ95, Hei98, HDL99, RS99a, RMSB96, Sch96, SK96, YMAC99]. eletrostatic [Hor90]. Eliminating [Mar93]. Elimination [CS98a, Edw96b, MMB90, OsKB^{+}92, YS98b]. Elliptic [CLO93, CSW95, DB97, EL96c, HW97, JYH93, LF96, LO93, MB96, OD99, PK95, PH95b, SS95, Spe95, TKT97, Vad95, YMAC99, ZY94, KM92, Knu92, LM90c]. Elliptical [Kou97]. Ellipticity [Arm94]. Embedded [Hew97, IS96, JC98, KMM96, Mei98, BGM91]. Embedding [EP96, EP99, LW96b, SM98a]. Emission [Hum96b, KCJ95, Poi91]. Employing [DL94b, DHS91, Sal91, dG92a, dG92b]. Enclosures [Gel99, Gro91, Gro92, MRB94]. end [BG92, VD90]. ende [Bas92]. Endonuclease [JMDB99]. Energetic [BDC97, BK99]. Energy [CBSW98, FPB98, Fur99, JMDB99, Maz94, MBI^{+}97, MS99b, Oku95, PDS99, PR93b, Wit96b, YR93, CYD92, CPP93, CKQ^{+}93, Flå92b, Flå92a, Gre90, Sal91, dG92a, dG92b]. Energy-Bounded [PR93b]. Engquist [Shu97, LCM96]. Enhanced [Löt94, RWC90]. Enhancing [HS93a]. ENO [AS96, Bau97, CA93, HHF95, JS96, LO98, OG97, WM94, Yan90, YYCH98]. Ensemble [AWY99, Sch95b]. Enskog [GA98]. Ensure [BS99b, BS99c]. Ensuring [Glo92]. entering [BY92, Bra92]. Entire [Boy94]. entrainment [KG90b, KG91]. Entropic [BD99]. Entropy [BD96, BCDL97, BC98, Osc94, Set94, Sho93, BWJ90, MWJ91, Rei92a, Rei92b]. Equal [GA96, GG92a]. Equation [AD97, Abe97, AHS96, AHS97, AC95, AAP97, AK95, AC93, BK93, BB96, Ben96, BD97a, BB97a, Boy95a, Boy96a, Boy97b, BSPL96, BCR99, BCDL97, BC98, CAA93, Can96, COS^{+}98, CG93, CH99, CJS99, CGY93, CC97, CKSB97, CGKC93, CDT98, DVD93, DK96, Edw96b, EP96, EO98, Epp94, FO93, FEO95, För96, Fra95, GA96, GGZ93, GGG94, GBS^{+}93, GS94, GEK^{+}97, GKKR99, GS96b, HIY95, HMOS95, HO96, HI98, Hu95, Hwa96, Jam99, JFH^{+}98, JC98, Joh93, JMR94, KD97, KOS^{+}96, KB96, Kou96, LPMS94, LW95, LK94a, Lif96, Lya99, MKM98, NC97, Noe94, Nor96, Nor98, Ohw98, Osb93, OP97, Pap93, PHL93, PA95a, Pry94, QS98, RK93, RF96, RW96a, RW96b, RSW98, RFH93, SB95, SWT98, Shy99, Sim93, Sim99, SRBG99, SR99, STS98, Sun93, Swe96, Taf95, TW94]. Equation [TV94, Tei99, TN98, Tho97, Tol94, TRL99, UWH99, WH98, Wan99, Win97, XMP97, YWS96, Zha98, ZZ98, ZY94, dLMSS95, ARS91, Ben90, BO92, Bor91a, BLA92a, BLA92b, CK90a, CRS90, Cha90, CW90, CGY92, CGKC92, CHW90, Dar93, Dav91, Dav92a, Dav92b, DGD90, DR90, Flå92b, Flå92a, FG90b, FG90a, GG90, GG92a, GCMR90, Glo92, Gra90, GM92, GGM93, GKM96, Gro92, HWY90, HWY91, JC92, KL97, KS92, LRB^{+}90, LBC^{+}91, MSF^{+}95, MS90a, MS91a, Mik90, Mik91, NSR92a, NSR92b, OGS90, OGS91, Osb90, Osb91, PRS90, Pap92, PRMV90, PO90, PO91, RF90, RF91, RCAM92a, RCAM92b, RLM90, Sch92b, Sch92a, SA90b, SA91, SM92, Sto92, TEKC92, TMR92a, TMR92b, Tas92, TSS92, YP92b, YP92a]. Equations [AK94a, ANL94, AM96, ABC^{+}98, AAP97, Anc94, And94, Ara97, AS99b, AE98, AK97, ADH^{+}93, ADS96, AQ99, AIV99, BS98a, BR97a, BR97b, BK94, Bec94, BTW96, BK97a, BMZ95, BS99f, BT97a, Bis95, BL93, BLT96, BFGG94, BPM98, CS93a, CZ98, CJ94, Cha95, CBS93, CZS98, CB97, Cle97, Cod99, CP95, CDW99a, CDW99b, Col97, CD98, CD95b, Cra94, CMM93, CSW95, DW97, DW98c, DBS94, DS96, DS99, DP93, DF96, DH95, DÜ92a, DB97, DIV98, DF98, Dri97, EVP97, ETFS94, Efr97, EP99, Eve96, Eyr94, Eyr95, FKM99, Fer95, Fey98a, Fey98b, Fis97, FD99, GTD98, GA95, GNHP95, GO96, GM95a, GK94, GN96, Gla95, GKKR99, Göt99, GTA99, GK98, GHKH93, GMPGV96, GKZ97a, HKM96, Hag94, Han93b, Hei93, Hen94, Her93, Hes98, Hil97]. Equations [HKR99, HLR99, Hol96a, Hu96, HRR94, HB97, HLL99, Hwa94, HS99c, ID96, JcW99, JR99, KI99, KPC93, KM95, Kop98, LO96, LM97, LS93a, LCM96, Leh99, Lil97, LMB94, Liu96b, LZ96, LS98, Löt94, LDB98, Ma93a, Mad95, MV99, Mar94b, MWW96, MLBW97, MP97, Mot98, MQS97, MJPC99, Nak95, NGN97, NPC93, Nor95, NC99, PK94, Pet94, PZC98, PV99, PH95b, PBD94a, Pop95, QW93, RL95, RF95, RIP96, RvR93, Rus93, SP98, SZ97, SPC93, Sam97, SS96a, SY97, Sco96, SNU98, STW97, Sha95a, SF96, Sha99, SS96c, SI94, Sjö95a, Sjö95b, ST93, Spe95, SVG93, Str99b, SP99, SCW94, Taf95, TAC98, Tau94, TTI97, Tei99, Tiw98, Tou97, Tou98, Vad95, VPS95, VP96, VC99, Ven95, WNY93, WEQ^{+}99, WK99a, Weg97, WD97, Whi94, WS93, YH95, YYCH98]. Equations [YM95, YPH94, ZHR97, ZLOT98, ZSK94, Zha94, dFD^{+}94, Adh92a, Adh92b, AB94, Alv91, Alv92, Arm94, BB90a, BB91a, Bar91b, BB91b, BBF92a, BBF92b, BL92, BL90a, BL90b, BJ90, CS92a, Cas90, CPJ90, Che91, CH90a, CH90b, CH91, CJ95a, Cot90, Dau92a, Dau92b, DP92a, DD90, DD91, DÜ92b, Dic90, DM90, DM91b, DD92b, Fab92, Fig92, Fog92b, Fog92a, FM92, Fra92, Ful93, FVZ90, Gla92b, Gla92a, Gro91, yGmC92, GC92, GMCH92, Gup90, Gup91, HGH90, HGH91, HLD92b, HLD92a, HB90b, ISW92a, ISW92b, IM95, IGABB90, IABBG91, JW90, KM92, KL97, KIO91, KH90, KH92, KPC92, Kop90, Kop91, Kor90, LBC^{+}93, LM90a, LM91, LS92b, LM90b, LK90, LPvL93, LR90, LR91, LJT92a, LJT92b, Luc90, Luc91, ML92, MH90a, Mar90, MDA91, MH90b, MH91, MS90b, MS91c, Nad95]. equations [PM90, Raw90, Raw91, Rok90, RKV90, RKV91, Rus90, RS94b, SPC92, SD93b, SSB90, SSB91, Shu91, Shu92b, Shu92c, ST92a, SA90a, SMR91, Ste92b, Ste92a, SVG92, Tau92a, Tau92b, TfNW91, TNW92c, TNW92d, TKKB92, Tes92, Tid95, Wat92b, Wat92a, Whi90a, Whi90b, WDH^{+}92a, WDH^{+}92b, YSG91, ZC91, ZC92]. Equidistribution [SWS96]. Equilibria [VM96a, HB90a, HB91, RG90]. Equilibrium [KL93, MVZ97, NP94, NMW^{+}96, RC95, RKO99, TLES93, Web98, GC90, HG91, HG92, HSN90, Shu90, SLV90, Shu91, Shu92b, Shu92c, SCC93a, TT90, TT91, TA91b, TAD92a, TAD92b, VM90]. Erosion [SP96]. Erratum [Ano94r, Ano95u, Ano97-29, Ano99-30]. Error [AAK^{+}93, BCM95b, CB96, FBM93, GTA99, HI98, KHW98, MG97b, Wer98, CH90b, CH91, JC92, MNR90, NSR92a, NSR92b, PB91]. Errors [Edw96b, Gho96, KM97b, Mar93, RHB94, BE91, BE92a, BE92b, Duk91, DM92d, DM92e, LT91, LT92]. Essentially [Abg94, DW95b, ES94, EL96a, Fri98, HEOC97, HS99b, Jin94, LOC94, Shu97]. Estimated [MO95]. Estimating [Hel95, Rie99a]. Estimation [BCM99b, LWT99, Rie94a, LT91, LT92]. Estimator [OP97, WKMH97]. Etching [AS95a, AS95b, AS97, SKO94]. Euler [Dri97, And94, BR97b, BS99f, CS92a, CS93a, Cha95, CP95, DW97, DS96, DS99, DP92a, DP93, Dic90, Efr97, Fab92, Fer95, Fey98a, Fey98b, GO96, GK94, GM96, Han93b, Hes98, HLR99, Hu96, HLL99, ID96, JMT97, Kop90, Kop91, LS92b, LS93a, Lee98a, LW96a, LPvL93, Löt94, MDB91a, MDB91b, MP97, Mul92b, Mul92a, NC99, PK94, RvR93, RS94b, Sam97, Sjö95b, TCS97, TAC98, Tiw98, Ven95, ZHR97]. Euler/Navier [MDB91a, MDB91b]. Eulerian [Mar97, CHMO96, CCG95, DE99, FAMO99, GBL^{+}92, GBS^{+}93, GRB^{+}95, HOS96, HAC97, RH91, SE90, Smi99, Ush96, WEQ^{+}99]. Evaluate [CDM98, LKB90]. Evaluation [Ano94s, BN94, CP98a, EG95, Gag98, GM98, HSZ^{+}97, Jab94, MVZ97, PDHS94, QG98, SH99, lX98, BS91b, BS92a, BS92b, CG97c, CJR91, CJR92a, CJR92b, ESM98, May92b, May92a, SBKH92, ZB92a, ZB92b]. evaporation [CYD92]. event [LKLE90, LLEK91]. Evidence [VDJ93]. Evolution [BGH^{+}97, BGH^{+}99, BLP98, CM95, CRRR97, GWI92, Gre94, JLL97, SR99, TH96b, YDT93, ZLOT98, DZ91, TLES93]. Evolutionary [MQS97, MS90b, MS91c]. Evolving [KP98, LZG99, LL93a, NPV96, RM93, Rie94a, GG92b, Zan91]. Ewald [RN94]. Exact [AAP^{+}95, LSB96, LV98, MS90a, MS91a, Pri94, Sak96, SZ97, SLL94, Wan99, BM91a, BM91b]. Exchange [BK97b, AAP^{+}95, Mik90, Mik91]. Exchange-Correlation [BK97b]. Excimer [LW94]. Excitation [CAA93]. Excited [GSB^{+}93]. Expansion [Bar94, FEO95, MBF94, Sch95b]. Expansions [Boy92d, DL94a, Fra95, MD95]. expectation [BRL91]. Experience [GA95]. experimental [Nat90, Nat91]. Experiments [Mar94a, Pet94, Sjö95b, Sun93]. Explicit [AR96, Ano94s, BT97a, CD93, CK92, CK93a, CCG95, DW96a, DW96b, GM96, Han93b, Mur96, OS98, PC96, RFL93, Ske98, SCW94, TCS97, WNY93, Wri98, YM91, YM92b, YM92a, Yee97, ZWS98, dNPT95, Ben90, Ben92a, Ben92b, KL92b, MDA91]. Explicit/Implicit [OS98]. Exploration [dM97a]. Exponential [CGY93, Hwa96, UL98, CGY92, KS92, LKLE90, LLEK91]. exponential-linear-in-depth [KS92]. Exponentially [HA93, MW94, Sim99, Wan97, CRS90]. Exponents [SV97]. Expressions [Sen99]. Extended [Li97, Lio95, SNL90]. Extension [AS99a, DW94b, DHS94, MM96, SBGK99, Wir97, FK90, Kle95]. Exterior [AK95, DE93, DÜ92a, GP98b, SRA98, DÜ92b]. External [GHS93a, SW99b, Tsy95, BAB90, Fer90a, Fer90b, GHS93b]. Extracting [WJTP96]. Extraction [Tak92a, Tak92b, DK92a, DK92b]. Extrapolated [Cra94, GSY92, Got92]. Extrapolation [Rea97]. Extrapolative [SS99b]. Extreme [VSM99]. Extremum [Mas96].

Faces [SSW98]. Faceted [RT94]. Facing [Zhu95]. factor [Dra90]. Factored [KTN^{+}94]. Factorization [BNW96, Iva93, SNU98, CM91, DD92b, Shu91, Shu92b, Shu92c]. Factors [Bec94, Ruy93]. Faddeev [HB90b]. Falling [RKJ97]. Family [Mah98, WEQ^{+}99, GY91]. Far [Bas92, BJ95, Wan99, BG92, SC90]. Far-Field [BJ95, SC90]. Far-from-Caustic [Bas92, BG92]. Fast [AS95c, AS99a, ABC94, BCM99a, Boa97, Bör97, BFO99, BM90b, BM91c, Boy92c, BIAV98, BFGG94, BCDL97, CDM98, CL93, CD95b, DS97a, DD95, ESM98, EG95, FD99, GR97, Gue94, Iva93, JCA97, JMR94, JS93, KKH^{+}99, LBL98, LRW95, MCPR95, MGM95, PMO^{+}99, Pet96, Pli95, RS94b, She95, STS98, Str91, Str92a, Str92b, Str97, Str99a, VD90, VP97, Zha98, ZLP97, BP90, Boy92b, BL90a, BL90b, But90a, But91, CGR99, CPP93, Dar90c, Dar91b, Dar93, HLB94, HHG91, HHG92, JVR93, LM92, Mak99, Sol92, Val90, YR98, Boy92d, SBKH92, Sor95, Tak92a, Tak92b]. Fast-Fourier-Transform [Tak92a]. FBTCS [BD93]. FCT [GK94, GG92b, Ods90a, Ods90b, Ods93, Vel93]. FDTD [FHKZ97]. Features [Kel94, KE93]. Feedback [MSV98, WJC93]. Feigenbaum [VDJ93]. FEM [AD93, DT90a, DT90b, GQ97, PL98]. Fermi [Gag98, Gro91, Gro92, Hof95]. Ferraro [KT99]. Few [BCRR98, MRD92]. Feynman [Nog93, PJ94]. FFT [JJK92, LM92, PH95b]. Fibonacci [MCPR95]. Fibres [SG98]. Fictitious [CJM97, FPQ93, GPWZ96, HKNT98]. Field [BDS^{+}99, Ben96, BJ95, CSN93, DS97b, FV97, FC95, GSB^{+}93, GGK95, GKG95, GHS93a, Jac99, oJ95, oJ97, JSD95, KLB99, KB99, LO98, MM98, Nak95, Rob99, RT94, SSW98, TSW95, VB95, WS96, Wan99, YP98, YR93, BLMR92, Blu92, HLD92b, HLD92a, Hor90, Hor91, Poi91, RD91, RCAM92a, RCAM92b, SC90]. field-effect [RCAM92a, RCAM92b]. Fields [AR93, BS99b, BS99c, EM95, EHM97, FM97, GS97b, Knu95, MD95, Mei98, Shi93, SBC99, SW99b, SB98, TS96, Tan94, WJTP96, BAB90, CG97c, ESM98, GWI92, Hor91, RH91, VY91, VY92, VC90, VC91]. Fifth [Boy96a, Tol94, TL98]. Fifth-Order [Boy96a, Tol94, TL98]. Filamentation [KF94]. Filaments [HKS98, Qi93, Fau90]. Film [LNR99, MB99, Pri98]. Films [BCM99b, KK95, RKJ97, LY90b]. Filter [Boy98, CG97b, JCA97, LM95, LCM96, MO95, YS98b, Ben92a, Ben92b, SCMU92a, SCMU92b]. Filter-Diagonalization [CG97b]. Filtering [BL99, Boy98, Can95, LIH97, SD93a, LO91a, LO91b, MS90b, MS91c, YR98]. Filters [VLM98, YSD99]. Filtration [KF94, SI94]. Find [Rea97, MMS90a].

Finding [DA94, Hoo90]. Finite [AL97, ABLD94, Ano99-30, AP91a, ADH^{+}93, ADS96, AK99, BD96, BR97a, BR97b, BMZ95, BL94, CLO93, CN95, CLNT98, CHR99, CGA93, CGA94, CO95, CA93, CGY93, CGKC93, CKR93, CDW99a, CL98b, DW98c, DDS97, DK98, Dur93, EL96b, EL97a, EL97b, FO93, Fat99, FA97, Fen99, FT96, Fur99, Gag98, GS97a, GMM99, GO96, GO98, GN96, GP98b, GPWZ96, GHKH93, Han93b, HT95, HT94, HMOS95, Her93, HWM96, HW97, HW96, Hub99, Hum96a, HKvV95, oJ95, oJ97, nJ93, JcW99, KKK95, KD97, KT99, KPP99, Kob99, LW95, LBB94, Leb99, LWT99, LK94b, Li97, LL96, LWC93, LL93a, LO93, Luk99, Mah98, MW97, kM93a, Mat97, MS98, Mei98, MW94, MLVM98, MQS97, MJPC99, MG97b, NC97, Nat90, Nat91, NC99, OBB98, PK94, PS93a, PZC98, PBD94a, Ram95]. Finite [SSR96, Sai95, Sak96, SPC93, Sch96, SLMS98, SBGK99, SA95, SS95, SK96, SCR97, SM98b, SNL90, SL98, SC96b, Tad97, TW93, TW94, TSW95, Tan94, VPS95, VP96, VO96, VS95, Wan97, WJTP96, WK99b, XMJ95, XMP97, YL98, YMAC99, YPH94, Zan97, ZC94, Zho98, ZL93, ZRR99, dLMSS95, vGVO98, vdVvdV98, AAP^{+}95, Ano95u, BB90a, BB91a, Ben90, CP90, Cas90, CGY92, Chi91, Chi92a, Chi92b, CGKC92, DD90, DD91, For92c, FMS92, GL95, Gla92b, Gla92a, yGmC92, GC92, HGH90, HGH91, HHG91, HHG92, HM90, JC92, LM90b, Lel92, LG94, kM90, kM91, Mao92b, Mao92a, MLB92, MB92, Nat92a, NSR92a, NSR92b, OB95, OGS90, OGS91, PTM92, RM90, RM91, RF90, RF91, SPC92, SAB95, Str94, Tay91, TNW92a, TNW92b, YMB^{+}91, ZB92a, ZB92b]. finite-boundary [PTM92]. Finite-Difference [CGA93, CGA94, CKR93, DDS97, GN96, HWM96, HKvV95, Li97, NC99, SA95, SS95, SM98b, VO96, YL98, ZC94, Zho98, ZL93, AP91a, BB90a, BB91a, Ben90, HGH90, HGH91, NSR92a, NSR92b, RM90, RM91]. Finite-Element [ADH^{+}93, Fen99, GMM99, GP98b, Hum96a, TSW95, Nat90, SNL90, AAP^{+}95, HGH90, HGH91]. Finite-Grid [BL94, Ram95]. Finite-Temperature [Gag98]. Finite-Volume [AL97, CA93, GS97a, KPP99, LL96, LWC93, MJPC99, MG97b, PK94, SLMS98, SC96b, HHG91, HHG92]. Finite-Volume/Newton [LL96]. Finite-Volume/Particle [MJPC99]. Fire [MRB94]. Fire-Driven [MRB94]. First [Ben95, Ena93, Jab94, Mar94a, MR99, STG^{+}96, BY90, BY91, Hal90a, Hal90b]. First- [STG^{+}96, Hal90a]. First-Order [Ben95, Ena93, BY90, BY91]. Fisher [QS98]. Fit [CJ95b]. Fitted [BW98, MW94, Sim99, Wan97, CRS90, SKK90, SKK91, SH92, Wes92a, Wes92b]. Fitting [TPRC96]. five [Fig92]. five-point [Fig92]. Fix [AK97, FMM99]. Fixed [LZG99, UKSTST97, UMS99, Vad97, Dra90, PL92a]. fixed-stencil [PL92a]. FLAIR [AP91b, AP90]. Flame [XS93]. Flames [QTL98]. Flat [CPB96, HPS92a, HPS92b]. flaw [Bor91b, DM91a, DM92a, DM92b]. flawed [DZ91]. Flexible [SG98]. Flight [LK98]. FLIP [Bra90, Bra91, BSB92]. Flow [Abg96, And94, AR96, Ano97-29, Ara97, AR97b, AMP^{+}98, BST95, BD96, BH96, BS99e, BD94a, Ber95, BBMB97, Bro95, BDML94, CN95, CBL95, CJR95, CE97, CM94, Cho93a, Cho97, CCG95, CDT98, CRRR97, DGN97, DECB98, DK98, DS97b, Dur93, ES94, Fis94, FJGP94, FM97, GHS93a, HC97, HD97, HCZ99, HML99, HAC97, HH98, HT97, HZC^{+}95, HLOZ97, HW96, HL98, HC93, Jac95, JYH93, JV95, Kar94, KNW99, Ku95, KCV99, Kup98, LL99, LT93, Lee98b, LL96, Lil97, LL95, LF96, LLK98, MB99, MWM94, Mar97, Mav98, Mei98, MW94, Min96a, MB97, MB96, MH93, MLVM98, NWK98, NC93, Oos97, ÖLJ98, Pai97, Pap93, PBC^{+}95, PH95a, PM93, Pet96, PR93a, PG97, PMR97, Poz99, Pri93, Puc97, PV93, PR93b, RS95, RWW94, RS99a, RHB94, SB96a, ST95]. Flow [SBGK99, SK98, Set94, SK96, Shy99, SCG94, SR98, SB96b, SP96, SXG99, SSO94, TP97, Tsy95, VM96a, Ver97a, Ver97b, WK96, WKK98, Xu97, YH95, ZP95, Zhu95, vBKG97, BP90, Bar91a, BLK90, Bor92a, Bor92b, DM92c, DD92b, Fer90a, Fer90b, GKM96, HL92a, HL92b, IT90, IGABB90, IABBG91, KBS90, Kle95, KHM92a, KHM92b, KG92a, KG92b, LGC92, LG90, LG91, LH90, MNP90, Nat92b, Pap92, PPM92a, PPM92b, PM92, PDC91, RM90, RM91, Ram90a, Ram90b, RH91, RS90, RWC90, SBGM92, She90, She91, SV91, SA90b, SA91, Sul91, SB91, TMR92a, TMR92b, VY91, VY92, WR91]. Flow-Impedance [ÖLJ98]. Flowfields [JT99]. flowing [MD90]. Flows [APV98, ARB96, AR97a, AE95, BMGM95, BLG97, BPT95, BW98, BS96, BNW96, CDF95, CS97a, CS99, CPB96, CHMO96, CM99a, CM93, CK93a, Cle97, CS97b, CP99, CEG^{+}97, CL98b, DFaIM98, DE99, DL94a, EL96a, EL97a, EL97b, Edw96a, ES93, FB94, FAH97, FBM93, FMO97, FMM99, FAMO99, FH98b, FGM97, GSB98, GO96, GO98, GG97, GB97, GM95a, Gri94, GQ97, GLN^{+}99, Han93a, HK95, HS99a, HSZ^{+}97, IS96, IR98, Jac99, JR96, JSD95, Kar97, KP97b, Kno98, Kop94, KK96, Kop96, KMM96, KM97b, KMS99, LNS^{+}94, LNR99, LLD93, Leb99, LML99, LZS98, LL93b, LL94, LQK98, LKK99, LBL98, MD98, MBV93, MPR99, MG97a, MA95, Mau99, MRB94, MPPZ96, Mon94, MFZ97, MVZ97, MB94, MJPC99, NP94, PB98, PP93, PVQ97, Pop95, PC99b, Pri95, PM98]. Flows [PAB^{+}97, RG97, RL97, RH99, Sai95, SA99, SBGK99, SK98, SMT99, SMJ98, SM98b, SDP97, SCT^{+}99, SWD95, SAB^{+}99, SF99, TZA96, TH96a, Tid97, TM94, Ton96, TM93, Ush96, VM96b, VO96, Vie94, VRD99, Wel95, WP97, Wel98, WSK99, WL93, WLC96, WWW95, Xia99, XP99, YB98, hYsY96, YSMG97, YMUS99, ZB96, Zho96, vM95, vdVvdV98, BM92, BAB90, BY90, BY91, BY92, Bra92, Bue91, CPJ90, CS92b, CS92c, CK92, Dar90b, Dar90c, Dar91b, Dom90, ES92, Fis90, GG92b, GC90, GHS93b, GS93, HLB94, HPS92a, HPS92b, HLS94, HW92a, HW92b, Kal91, Kal92, LLD92, LS92a, Lar91, Lav90, MDB91a, MDB91b, MGP91, Mav90, MLB92, PTM92, PL92b, Poi92, RR90, RR91, RIW90, SC90, SK90, SCC93a, SA92a, SA92b, Tay91, TNW92a]. flows [TNW92b, UT91, UT92b, UT92a, YKM90, YTS92]. Fluctuation [Asl99, WK99b]. Fluid [Ara97, BH98, BS99e, BC99, BO96, CHMO96, CJR95, CDW99a, CDW99b, Cor96, DNN93, DE99, FAX99, FAMO99, GO96, GO98, GLN^{+}99, HR96, HW96, HC93, IR98, Iva93, Jac95, Kho98, KPC93, KM95, KM97a, LPMS94, LT93, Lil97, LW97, MW97, Mau99, Oos97, PAB^{+}97, RD91, RC95, RFL93, Sai95, SK94, Sco96, Shy99, SCT^{+}99, TK96, UI99, Ver97a, Ver97b, Vu96, Zal97, Bil90, Fig92, GL95, KPC92, KHM92a, KHM92b, KCV99, LL92a, LL92b, MNP90, RD92a, RD92b, Sch92c, SV91, UT91, UT92b, UT92a, VNC92a, VNC92b, YSG91, YMB^{+}91]. fluid-ion [LL92a, LL92b]. Fluid-Mixture [Shy99]. Fluid-Particle [Mau99]. Fluid-Solid [Iva93, GL95]. Fluidized [vWBS^{+}97, CGSS90, CGSS91]. Fluids [Kup98, Sch95b, SHF97, LRJ^{+}99]. fluorescence [LKLE90, LLEK91]. Flux [AAL97, AP90, AP91b, BB97b, CB97, DeV91, DM96, DFaIM98, GMM99, GVM99, GC90, LeV98, Lin95, LS93b, LDB98, MA95, PFS98, RK95, RvR93, SS94, SCM98, SS96a, Smi99, Thu96, TO96, XMJ95, Xu99, Zal97, ZWS98, BR92a, BR92b, CLJ^{+}90, Dic90, Gra90, Rob90a, SH90, Shu92a, Shu90, SLV90, VM90]. Flux-Corrected [BB97b, GMM99, SS94, Zal97, ZWS98, DeV91, Gra90]. Flux-Correction [SS96a]. Flux-Difference [Lin95, Dic90]. Flux-Limited [Thu96]. Flux-Split [DFaIM98, GC90]. flux-splitting [Shu90, SLV90]. Flux-Vector [CB97, LDB98, XMJ95, VM90]. Fluxes [AK97, BS99b, BS99c, CL98a, LM97, VK98, LVS90]. FMM [SB98]. Fock [Dav91, Dav92a, Dav92b, FG90b, FG90a, dFD^{+}94]. Fokker [BCDL97, BC98, Epp94, GGZ93, Mik90, Mik91, PC92b, PC92a, SVG92, SVG93]. Folding [AT99]. Following [Maz94, SH94, UWH99]. Force [GHS93a, GBS99, SS99b, Sch95a, TWV97]. Force-Splitting [SS99b]. Forced [GGG94, MD98]. Forces [CDM98, Cod99, Cor96, DLMN91, DLMN92b, DLMN92a, LDB96]. Ford [FK90]. Form [BTW96, CI97, HI98, Sen99, van93, Dau92a, Dau92b, SVS90, SVS91]. Formalism [SS96a]. Formation [RKJ97, RB95, Alm93, SCA92a, SCA92b]. Formula [DM96, Gra95, Jam99, ARS91, LT91, LT92, Vre91a, Vre91b]. Formulae [GC95]. Formulas [KU94, LV98, Hui90, Hui91]. Formulated [Hag94, HHG91, HHG92]. Formulation [BK97b, CHMO96, CZS98, Cle97, DH95, DE99, EL97a, EL97b, EP99, ES93, Gla95, KNW99, LM95, Leh99, LH95, LLK98, MKM98, Nak95, NPC93, Pri96, RS90, TK99, Waj93, Yee97, YMAC99, AC92, BSB92, Dic90, ES92, Hei92, LM90b, MH94, SAB95, Tou92]. Formulations [CSW98, CE97, Had99, Shu95, Taf95, WWW95, Tid95]. Fornberg [Aoy95]. Fortified [Fuj95]. FORTRAN [LP90]. Forward [Ano97-29, CS97a, CS99, CG97b, SB96b]. Forward-in-Time [Ano97-29, SB96b]. Four [AFH94]. Fourier [Boy92d, SBKH92, Sor95, Tak92a, Tak92b, Boy92b, Boy92c, Boy94, BS98b, Can95, CG93, CK90b, CK91, CJ95a, CD98, EHM97, GTD98, GMCH92, Liu96b, Osb90, Osb91, Pel91, PO90, PO91, RF90, RF91, SE98, TKKB92, Tes92, TC98]. Fourier-Wavelet [EHM97]. Fourier/Chebyshev [CJ95a]. Fourth [AD97, GKZ97b, GO95, Hen94, RF95, Yav97, dFSS92]. Fourth-Order [AD97, GKZ97b, GO95, Hen94, RF95, Yav97, dFSS92]. Fractal [EHM97, JMR94]. Fractional [Abd95, AS99b, JR96, Per93, Per95, Sha95a, Yin96, ZSK94, LM90a, LM91, MPG92a, MPG92b, RKV90, RKV91].

Fractional-Step [AS99b, JR96, Sha95a, LM90a]. fractions [Lar91]. Fractured [LLK98]. Fragmentation [LNS^{+}94]. Frame [Bar94, BTW96, HTK90]. Framework [BPS95]. Free [CS99, CJR95, CK93a, Clé96, CT98, DL94a, HS93a, HS99a, JMDB99, KP97b, KNR99, LSLG97, MBV93, MPR99, MA95, MBI^{+}97, Mon94, Nun93, Pet96, RLT93, RS99a, SP98, SSR96, SVV98, TH96b, TLW95, TM94, ZYKC98, AN90, BR92c, BR92d, Bue91, Cal90, Cal91, CS92b, CS92c, CK92, For92c, FMS92, Kar91, Kar92, LGC92, LBL98, PBBS91, PTM92, Ram90a, Ram90b, RR90, RR91, Rom92a, Rom92b, SRSB93]. Free-Boundary [TH96b, LGC92]. Free-Lagrangian [HS93a]. free-shear [Bue91]. Free-Surface [HS99a, RS99a, Ram90a]. Freeman [Boy97b]. frequencies [Tak92a, Tak92b]. Frequency [AN98, ARTAAA97, BDC97, DOOSM99, DDS97, FEO95, GRB^{+}95, GH95, PC96, SA95, Tan93, LL92a, LL92b]. Frequency-Dependent [AN98]. Frequency-Domain [SA95]. Friction [BLT96, PM99]. Frictional [GS97b]. frog [AA90]. Front [ALTP98, GT97, JT96, KP98, LS96, Mav95, MP93, QTL98, Boy92a, Poy92, RT92, UT91, UT92b, UT92a]. Front-Tracking [ALTP98, JT96, UT92b, UT92a]. Frontal [GTD98, RS94a]. Fronts [UMS99]. Froude [TSR92, TS96]. Full [BES94, CLJ^{+}90, ETRW95, XMP97]. Full-Wave [ETRW95, CLJ^{+}90]. Fully [Anc94, BY94, CLNT98, Kho98, MLVM98, MK97, RT94, SPCD96, STW97, VBD99, Wan95, MLB92, SA90a]. Fully-Lagrangian [Anc94]. Function [AFH94, AQ99, BPS95, Bri95, EP99, FD93, For97, MH94, MP93, OK94, Osc94, RvR93, Tad97, vGVO98, BG92, Bas92, MPG92a, MPG92b, MS91b, MS92b, MS92a, Raw91, SC90]. Function-Transport [FD93]. Functional [VC90, VC91]. Functionals [PDS99]. Functions [Boy92d, Boy94, CP98b, Gou96a, Hof95, Jab94, LZG99, MBF94, NCF96, PDHS94, PZ98, SRVK96, Sor95, BS91b, BS92a, BS92b, CR91, CJR91, CJR92a, CJR92b, Dom90, HS90, KF91]. fundamental [JB91, JBvK^{+}91, Kar91, Kar92]. Fusion [KL93]. Future [Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93g, Ano93h, Ano93i, Ano93j, Ano93k, Ano93l, Ano94a, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano94i, Ano94k, Ano94l, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano95i, Ano95j, Ano95k, Ano95m, Ano95n, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, Ano96a, Ano96i, Ano96j, Ano96k, Ano96l, Ano96m, Ano96n, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Ano97j, Ano97k, Ano97l, Ano97m, Ano97n, Ano97o, Ano97p, Ano97q, Ano97r, vL97a, vL97b, Ano90c, Ano90d, Ano90e, Ano90a, Ano90b, Ano90f, Ano91a, Ano91d, Ano91b, Ano91e, Ano91f, Ano91c, Ano92a, Ano92c, Ano92b]. future [Ano93f, Ano94j, Ano95l].

GaAs [LK94b]. Galerkin [AC92, AQ98, ABLD94, AQ99, CS98b, DS94, Eyr95, Fen99, Gil97, Gir97, Gir98, HHR99, JLB90, LKK99, Pri94, RL95, SO95, She95, WK99a, vdVvdV98].

Galerkin-Collocation [DS94]. Galerkin/Runge [Gil97]. Game [JH97]. Gamma [CP98b]. Gap [AK99, DL94a, RCA93, RCA92]. Gaps [CS97a]. Gas [APS97, BD96, BO96, DS97b, Eka99, FMO97, JMT97, Kul95, LL93b, MS99a, Mon97a, MS99b, MVZ97, NP94, Pap93, PC99a, PX93, SMD98, SW99a, SBC99, SXG99, TH96a, TS99, XP94, XMJ95, Xu99, YCTC97, YPH94, vL97a, vL97b, vWBS^{+}97, BM90a, Boy90, Boy91a, CMK90, CYD92, Kop90, Kop91, MOS92, Nic93, Pap92, VM90]. Gas-Dynamics [vL97a, vL97b, Kop90, Kop91]. Gas-Fluid [BO96]. Gas-Kinetic [PX93, SXG99, XMJ95, Xu99]. gas-phase [CYD92]. Gasdynamic [LMB94]. Gases [FWBS94, SLL94, LVS90, RNSV91, Shu90, SLV90]. Gauge [AFH94]. Gaunt [GÖAY95]. Gauss [BN94, LPMS94, Yak96]. GAUSSIAN [LW96b, CP98a, Chi91, Chi92a, Chi92b, FWWD94]. GEMMW [DHSS94]. General [AP91a, BRR92, GM98, KU94, RH98, SS95, TM94, Tót97, VLM98, WSK99, DS90, May92b, May92a, OGBSK90]. Generalisation [JB95]. Generalised [LH95]. Generalization [Boy92d]. Generalizations [Mor92, MVZ97]. Generalized [ART95a, ARTAAA97, AB97, BK94, BD97b, Can96, CJS99, CM95, HKR99, Jor99, NO93, OD99, SH94, VM90, Mor90, RKV90, RKV91, YR98, ART95b]. Generate [Gre94]. Generated [Ruu98a, Ruu98b, RMO99]. Generating [CKE99, Chi91, Chi92a, Chi92b, HN95]. Generation [And94, CPB96, CC98, CS93b, Eça96, GA98, HK93b, JS93, KM96, Knu95, LWS98, Man93a, Mur96, Nog93, OK94, Reb93, RS99b, SAS94, Set94, Spe95, TYS95, TA91a, AC92, AS90, AS91, BES90, Dar90a, Dar91a, Dar92, Dvi90, Dvi91, FK90, HL90, HL91, Kan92, KL92a, LP90, Mav90, SR90, SR92b, SR92a, Wes94]. Generator [CI97, MCPR95, ASTAS91, Knu92]. Genomics [SS99a]. GENSMAC [TM94]. Genuinely [CWC99, GNHP95]. Geodesic [Gir97]. Geometric [BB96, FH97, FHL97]. geometrical [BWJ90, MWJ91]. Geometrically [Smi99]. Geometries [AM96, CBL95, DÜ92a, DECB98, DL94a, Ku95, LS98, MV99, MGM95, SP98, UKSTST97, VLM98, DÜ92b, SD93b, SBGM92]. Geometry [Fis94, GM95a, LMS98, Pri95, RIP96, SS96b, SLMS98, UL98, VC99, BM91a, BM91b, CLJ^{+}90, CHL^{+}90, WDH^{+}92a, WDH^{+}92b]. Geophysical [BC99, SM98b]. Geostrophic [SMW99, Vie94, LBB94]. Ghost [FAX99, FAMO99]. Ghosting [YS98b]. Given [Gou96b]. Glaister [MIM90]. Global [Abd94, AK98, GS94, HW92a, HW92b, SSF92a, SSF92b, Swi96, WNY93, XMP97, BS91b, BS92a, BS92b, LJT92a, LJT92b]. Global-Scale [Swi96]. GMRES [KM92, Oos97]. GMRES-Based [Oos97]. GMT [ARTAAA97]. Godunov [Hir97, BS99b, BS99c, CCG95, DW95a, DW97, EMRS91, Hal90a, Hal90b, Kle95, KH97, LeV98, LH95, LH90, MP96, Min96a, MR98, SA99, Tab96, TS99, Van99, XH98, ZC91, ZC92, vL97a, vL97b]. Godunov-Projection [Min96a]. Godunov-type [EMRS91, MR98, XH98, Kle95, KH97]. Good [HP98, Mah98, Swa99]. Gordan [GÖAY95]. Gordon [AHS96, AHS97, CG93, DK96]. GPMT [ART95b]. Gradient [ANL94, BL98, BK97b, MBP94, BK91, BK92a, BK92b, Che91, JE90, KH90, KH92, NP92, ZB92a, ZB92b]. Gradient-Corrected [BK97b]. gradient-weighted [ZB92a, ZB92b]. Gradients [LeV98, vBvNW95]. Gradiometer [JB95]. Grand [LPP99, Sch95b]. Granular [PC99b, ZDG95]. Graph [Nog93, MMS90a]. Graphite [KE93]. graphs [FK90]. Gratings [ES98]. Gravitating [YDT93]. Gravitational [CDM98, JDC94, MML98, SD93a, Sch95a]. Gravity [BD93, Clé96, CS93c, Dol92a, Dol92b]. Gray [Win95]. Great [Due93, DA94]. Greater [KSB^{+}99]. Green [BPS95, Dom90, NCF96, Tad97]. Greengard [HM95, SB98]. Grenoble [Lya99]. Grid [And94, BK93, BH98, Boy92c, Bra93, BL94, BS98b, BSPL96, CS98a, CI97, CC98, COBA95, Eça96, Edw96b, EL96c, FH98a, FJ98, Hag94, Jak93, JC98, JLM^{+}96, JS93, JR96, KM96, KK96, Kop96, Kop98, LZG99, Man93a, MS99a, MA96, OBL93, OK94, PBC^{+}95, PR93b, Ram95, SS94, SAS94, SLMS98, Set94, SSW98, SBO93, SCR97, Spe95, SR97, SP99, ST99, Taf95, TYS95, Tan94, TRL99, WS93, YB98, YMUS99, ZSK94, ZC94, vdVvdV98, AM90, AC92, Boy92b, BBGL93, Cot90, Dar90a, Dar91a, Dar92, Dvi90, Dvi91, FVZ90, HTK90, HL90, HL91, ID96, Kal91, Kal92, Kan92, KL92a, Knu92, MH90a, Mit92, MF92, SV91, SR90, SR92b, SR92a, YM91, YM92b, YM92a, YMB^{+}91]. Grid-Based [JLM^{+}96]. Grid-Point [SP99]. Gridding [LT93]. Grids [ABBM96, AS94, AB97, ABLD94, And94, ARB96, AS99b, AD93, BD93, CG96, CKE99, CDW99b, EL97a, Fat99, Fri98, GO98, Gil97, Gir97, Gir98, HLD96, Hen94, Hub99, KPP99, KH97, KMM96, KMS99, LS93a, LL96, LO98, LV98, Löh95, LBL98, Mad95, Mei98, Min96b, Mur96, MG97b, PK94, PM93, PV99, SK98, SS96c, Tab96, TPRC96, UKSTST97, UMS99, Ven95, VM96b, VRD99, Wan95, WSK99, ZL93, vBvNW95, BK91, BK92a, BK92b, Lav90, LS92b, Löh90, LA90, PM92, SSB90, SSB91, SAB95, Ste92b, Ste92a, TA91a, Wat92b, Wat92a, Wes92a, Wes92b]. Griseus [DOOSM99]. Grooved [PS93a]. Gross [Eve96]. ground [FG90b, FG90a, Hoo90]. Group [FWWD94, ST93, For92a, For92b, ST92a, VG92b, VG92a]. Growing [KO98]. Growth [RT94, TH96b, AT92, SS91, SS92]. Guidelines [GH97]. Guides [HPG98]. Guiding [KPCJ97]. Gummel [Kul95]. Gust [SA95]. Gyration [KRGv93]. Gyration-Dominated [KRGv93]. Gyre [SMW99]. Gyrokinetic [BDML94, CW93, DL93].

Hairpin [Cho90a, Cho90b, Cho93b]. Half [LM97, YWS96]. Half-Space [YWS96]. Hall [Hum96a, KLBD93]. Hamilton [BS98a, LO96]. Hamiltonian [BS98b, CR91, Hol96a, LS94, MMS90b, MMS91, Row91]. Hankel [DK90, vNZ94]. Hard [MRC93]. Harmonic [Bor91b, Bor92a, Bor92b, HT95, Dvi90, Dvi91]. harmonics [Mik90, Mik91, YR98]. Harten [Shu97]. Hartmann [Leb99]. Hartree [Dav91, Dav92a, Dav92b, FG90b, FG90a, dFD^{+}94]. Heat [BL99, BR99, DW98a, HIY95, LL99, LG90, LG91, LEI96, LL96, Liu96a, Nor98, PNC94, SCM98, STW97, Sie99, ZWS98, dG96, CW92a, CW92b, LE90, LE91]. Heated [RKJ97, PC92b, PC92a]. Heating [CC97, Ram97b]. HECTOR [KC92, KC93]. Hele [BST95, HLOZ97, Whi90a, Whi90b, Whi94]. helium [FG90b, FG90a]. Helix [DMW99]. Helmholtz [AK95, BD97a, BL92, COS^{+}98, EVP97, EO98, Eve96, Hu95, Lif96, LJT92a, LJT92b, Lu99, Mar90, SRA98, TW94]. Helmholtz/Schrödinger [EVP97]. Helmhotz [BL93]. Helmhotz-Type [BL93]. Hermite [SH98a, TC98]. Hermitian [LPBS96, WL96]. Heterogeneous [HSS97, MML98, CYD92]. Heterojunctions [LK94b]. Heuristics [Due93]. Hexagon [SCR97]. Hexagon-Roll [SCR97]. Hierarchical [PL98, Ser96]. Hierarchy [SVG93, SVG92]. High [AS96, ARTAAA97, AR96, AR97b, BS99b, BS99c, BR97a, BR97b, CS93a, COS^{+}98, CGA93, CGA94, CO95, CA93, CWC99, DW95a, DDS97, DM97b, ES94, EVP97, Eka99, EB94, FEO95, FMO97, FM97, FCC97, GO96, GO98, GRB^{+}95, Gup90, Gup91, Har97, HEOC97, JcW99, KIO91, LO91a, LO91b, LO96, LNR99, Lax97, LS96, LeV98, Leb99, LM95, Li97, LPC^{+}93, LL93a, LO98, Mah98, MCPR95, Maz94, MP96, MS99b, Nor96, NC99, Pen95, PZC98, PAB^{+}97, Rig94, SP98, SO95, SW92, Sha99, SWT98, Shu97, Sjö95a, Tay94, TRL99, VRD99, WR91, YG94, YKM90, Yee97, YSD99, Zha98, Zho96, Zho98, AB94, BY92, Bra92, CS92a, CKQ^{+}93, DD92b, Mul92b, Mul92a, Tay91, TNW92a, TNW92b]. High-Accuracy [YG94]. High-Energy [Maz94]. High-Frequency [ARTAAA97, GRB^{+}95]. High-Order [BR97a, BR97b, CS93a, COS^{+}98, CGA93, CGA94, CO95, CA93, DW95a, DM97b, GO98, JcW99, LNR99, Li97, MP96, MS99b, NC99, Pen95, PZC98, PAB^{+}97, SO95, Sha99, YG94, YSD99, Zho98, KIO91]. High-Precision [EB94]. High-Resolution [AS96, AR97b, CWC99, DDS97, Eka99, LeV98, SWT98, Yee97, YKM90, Mul92b, Mul92a]. high-Reynolds [BY92, Bra92]. High-Speed [Zho96]. Higher [CK96, FJ98, Hwa96, MLVM98, Ohw98, OGWW98, Rav97, YR93, Nor91a, Nor91b, RS90, ZC91, ZC92]. Higher-Order [FJ98, Hwa96, ZC91, ZC92]. Highly [EL96c, HK98, MG96, Nor98, SSF92a]. Hilbert [GRB^{+}95]. Hill [Wit96b]. Histogram [KE93]. History [CK92, CK93a]. History-Dependence [CK93a, CK92]. Hole [GBBH96]. Homoclinic [AHS96]. Homogeneous [BT97a, PC99a, Sch90]. Homogenization [MDH98]. Homotopy [Eve96]. Hopf [AAP97, GGH90, She90, She91]. Hopping [Dru99]. horizontal [HWY90, HWY91]. horror [Mil91]. Hot [CKR93, Nun93]. Hourglass [CS98a]. Hourglass-Type [CS98a]. hp [LQK98, OBB98, RS99a, SK96]. hp/Spectral [RS99a]. Huang [Ano99-30]. Hugoniot [JM98]. Human [JBA96]. Humans [DDS97]. Hybrid [AS96, Bau97, BDC97, DW96a, DW96b, DDS97, FHKZ97, Had99, HP93, HK98, HLOZ97, KPP99, LT93, LJT92a, LJT92b, Mat94, MJPC99, Nea94, Nun93, OS97, Ram95, Ram97b, SH99, SK98, SFHD99, SCW94, Swi96, Vu98, ZOI97, Bue91, HP92, LS92a, PTM92, Ste92b, Ste92a]. Hydrodynamic [CCGJ95, GS95, KM93b, ST93, TO96, Zan92a, Zan92b, ST92a, SRSB93]. Hydrodynamics [BBF^{+}99, Bal94b, BH98, CBSW98, Chr97, CM99b, DW96a, DW98b, DW95b, FQ96, GS97b, LPC^{+}93, LSRS99, Man93b, MM96, Ods93, PL93, PX93, SKR^{+}93, SHA95b, WP97, Wha96, Zab97, Bal95, HCC91, MIM90]. hydrodynamics-suggestions [Bal95]. Hydrogen [KKOF97]. Hydrology [BL98]. Hyett [Cha90]. Hyperasymptotic [Boy95b, Boy97b]. Hyperbolic [AS94, BT97b, CGA94, CA93, CK93b, DW95a, DW96b, DS94, ER96, Ena93, FMO98, GTA99, GO95, Har97, Jin94, Jin95, JL96a, Lax97, LeV97, Noe94, PLD97, RL95, SO95, Sal94, TYS95, BWJ90, CGSS90, CGSS91, Col90, Dur91, DEO92, FP92, JW90, LO91a, LO91b, LY90a, MWJ91, NT90, PL92a, SW92, Shu92a, Tho90]. Hyperelliptic [Osb93]. Hypergeometric [For97]. Hypersonic [Edw96a, RS95, Zho98, LGC92, Mal90, MM91b, YKM90]. Hysteresis [CCGJ95].

IBVPs [GO95]. Ice [GS97b]. Icosahedral [SP99]. ID [Ano99-30]. Ideal [BD99, CG97a, DW94a, DW94b, JcW99, PV99, PS93b, PRL^{+}99, Rom97, Xu99, ZC91, ZC92, ZDG95]. Ideal-MHD [PS93b]. Identification [LEI96, DM91a, DM92a, DM92b]. II [AS95b, Cho93b, EH93, Fey98b, GO98, GKG95, GKL^{+}96, HL92a, KK95, KNW99, Kop96, MB97, Pri98, PO90, PO91, Rom97, Str91, Str92a, Str92b, SCA92b, Tho90, Tou98, WKK98]. III [Shu97, HEOC97, HL92b]. Image [TSW95]. imaging [Ben92a, Ben92b]. Immersed [AMP^{+}98, Bot98, DFFG96, PP93, RPB99, SG98, SW99c, YMUS99, Bey91, Bey92]. Impedance [HP99, ÖLJ98, WJTP96, Yor90]. Implants [Maz94]. implementable [dFSS92]. Implementation [Abg94, CW93, DW96a, DB96, ESH93, EH93, Hig99, JS96, KKH^{+}99, LM95, Nic98, PSL99, SSR96, SPCD96, Ser96, WONM96, Wan99, WDE98, WTH^{+}96, van93, HTK90, KM92, VNC92a, VNC92b]. Implementations [GM96]. Implemented [ZV97]. Implications [SW99c, YSG91]. Implicit [ARB96, BLG97, Bec94, BR92a, BR92b, BNW96, CM95, CCG95, DW96a, DW96b, Edw96a, Eka99, Ena93, Epp94, GSB98, GH95, GH97, JSE97, KMB^{+}92b, KMB^{+}92a, KLB99, KCV99, LW96a, LMB94, LBL98, MH90a, MWW96, Mei98, Mon97a, MK97, OS98, PC99a, RC95, RY94, SPC93, SPCD96, Sco96, SCW94, STS^{+}97, VM93, VM96b, WH98, WTH^{+}96, WM94, WMG^{+}91a, WMG^{+}91b, YYCH98, Yee97, Zho96, AN90, Ano91n, Car91, Cas90, CW93, DD92b, Fri90, Ful93, Kle95, KNW99, LL91, NWK98, RD91, RD92a, RD92b, SPC92, SCC93a, SH94, SA90a, VRD99]. Implicit-Explicit [CCG95, DW96a, DW96b, SCW94]. Implicit/Multigrid [ARB96]. implicitly [IGABB90, IABBG91, KvdVPG95]. Importance [BS99d, DS96]. Improved [DM96, Ena93, GMM99, HSN90, Kno98, LCM96, Löt94, MRD92, Nea94, Ods93, PP93, Rea97, Ton96, VC99, For90, TNW92a, TNW92b]. Improvement [HN95, LM90a, LM91]. Improvements [Loe93]. improving [VBB90]. Impulse [Cor96, CV97, EL97b]. Impulse-Based [Cor96]. Impulsive [JYH93]. impurities [Zan92a, Zan92b]. Impurity [STS^{+}97]. Inadequacy [BY90, BY91]. Incidence [WJTP96]. Inclined [MB99]. Include [Gou96a]. Including [Tan94, Wel95, CYD92]. inclusion [Dem92a, Dem92b]. Incomplete [CP98b, DK90]. Incompressibility [Wer95]. Incompressible [ABC^{+}98, ARB96, Ano97-29, Ara97, BH96, BK94, BTW96, BW98, BNW96, Bro95, CHMO96, Cho97, Cod99, CD95b, DF96, DK98, DIV98, ES94, EL96a, EL97a, EL97b, ETFS94, Fis97, FGM97, GTD98, GG97, GB97, GQ97, HCD98, HCZ99, HML99, Hen94, HW96, Hwa94, Jac95, JB93, Joh93, JR96, LL99, Lil97, LZS98, LF96, MBV93, MV99, Mar94b, MW94, Min96a, MB97, MLVM98, MFZ97, MB94, PR93a, PAB^{+}97, Puc97, RL97, SK98, SNU98, SR98, SB96b, SSO94, SAB^{+}99, Taf95, TZA96, Tau94, Ver97a, Ver97b, VO96, VRD99, WL93, WWW95, YYCH98, YMUS99, ZSK94, Zha94, AB94, BY92, Bra92, BJ90, CPJ90, CJ95a, DD92b, GS93, Gup90, Gup91, HLB94, KIO91, LM90a, LM91, LR90, LR91, MH90a, MGP91, Nat92b, PDC91, RKV90, RIW90, RKV91, RWC90, SC90, SBGM92, SA90a]. incompressible [SA90b, SA91, SA92a, SA92b, UT91, UT92b, UT92a]. Incorporating [KLBD93]. Incorporation [PM99]. Incremental [KTN^{+}94, STG^{+}96]. incurred [BE91, BE92a, BE92b]. Independent [SB98]. Index [Ano93m, Ano93n, Ano93p, Ano93q, Ano93r, Ano94m, Ano94n, Ano94o, Ano94q, Ano95p, Ano95q, Ano95s, Ano95t, Ano96p, Ano96q, Ano96r, Ano96s, Ano96t, Ano96u, Ano96v, Ano96w, Ano97s, Ano97t, Ano97u, Ano97v, Ano97w, Ano97x, Ano97y, Ano97z, Ano97-27, Ano97-28, Ano98r, Ano98s, Ano98t, Ano98u, Ano98v, Ano98w, Ano98x, Ano98y, Ano98z, Ano98-27, Ano99s, Ano99t, Ano99u, Ano99v, Ano99w, Ano99x, Ano99y, Ano99z, Ano99-27, Ano99-28, Ano90i, Ano90j, Ano90k, Ano90l, Ano90m, Ano90n, Ano91h, Ano91g, Ano91i, Ano91j, Ano91k, Ano91l, Ano92e, Ano92f, Ano92g, Ano92h, Ano92i, Ano92j, Ano93o, Ano94p, Ano95r]. Indirect [GGK95]. Induced [HK93a, KB96, Mar93, MG96, CWG90, CWG91, FA90, FA91, HK92, SE90, SCA92a, SCA92b]. Induction [DDS97, Sen99]. inductive [TA91b, TAD92a, TAD92b]. Inertia [BD93]. Inertia-Gravity [BD93]. Inexact [KM95, STW97]. INF [She95]. Infinite [Boy94, SK94, Wan99, CJ95a, HG94, Osb90, Osb91, PO90, PO91, SMR91]. infinite-interval [Osb90, Osb91, PO90, PO91]. Infinitely [Sch95a]. Inflow [LWS98]. Influence [Efr97, KHW98, MSF94, Wer95, Wer98, Dau92a, Dau92b, Tas92]. Influence-Matrix [Wer95]. Inherent [TYS95]. Inherit [Fur99]. Inhomogeneous [ABBM96, ADS96, GGK95, GKG95, JLL97, KM97a, Tan94, AP91a, Ben92a, Ben92b, DGD90]. Initial [DS94, ST97, DM92c, For90, SCA92a]. initial-boundary [For90]. Initial-Value [ST97]. Initio [AT99]. Injection [Zan99]. Inner [Wit96a, TfNW91, TNW92c, TNW92d]. Insect [LK98]. insight [BKP91]. Instabilities [CL98b, CB94, RKJ97, Vu96, Vu98, VBD99, vBKG97, DMZ91, GLM^{+}90, HCJ92a, HCJ92b, LL92a, LL92b, OsKB^{+}92]. Instability [Aoy95, ADS97, BL94, BDML94, Gel99, HCZ99, LJG96, LSB96, LL93a, Ram95, Ske98, VM96a, Wri98, AA90, AA91, MS90a, MS91a]. Integer [SCT^{+}99]. Integrable [AHS96]. Integral [ARV99, AK95, BMZ95, DNN93, Eyr94, GEK^{+}97, GKKR99, GKM96, Gue94, HBF93, Hof97, Hu95, KB96, LPMS94, LLK98, Mad95, Mot98, Nak95, NPC93, ZCH96, BL90a, BL90b, Che91, Dol92a, Dol92b, DK90, Dvo91, DK92d, DK92c, HB90b, Kan92, KL92a, LP90, Rok90]. Integrals [CP98a, GM98, NM95, QF99, SH99, Sen99, Dvo91, DK92d, DK92a, DK92c, DK92b, LBC^{+}93, LKB90, May92b, May92a]. integrals-decomposition [DK92d]. Integrated [PA95a, ZRR99]. Integrating [GS96b, LBB94]. Integration [AM96, Ara97, AKW93, CD93, DH96, FWWD94, GA95, Lil97, Mar94b, MC95, PZ98, Pop95, QF99, Sim93, ST97, Sim99, VM97, ARS91, CR91, HT94, HS90, MSF^{+}95, Mil91, dFSS92, tVB92a, tVB92b]. Integrations [LV98, SNL90]. Integrators [FHL97, HA93, LS94, Maz97, Oku95, For92a, For92b]. intensities [GCMR90]. Intensity [MSF94]. Interacting [KKOF97]. Interaction [AS96, DFN^{+}99, Fau90, GRB^{+}95, Gre98, HHF95, MBI^{+}97, SK98, SCR97, Zin94b, GL95]. Interactions [BFO99, Bru93, Cho93b, DQ95, DD95, DL94b, FCC97, ÖM95, Sch92c, BF91, Bru92, Cho90a, Cho90b, MMR90]. interactive [TSR92]. Interface [BMZ95, CNG99, CHMO96, Edw96b, FV97, FPQ93, GLN^{+}99, HKW97, HLOZ97, LL99, NC99, TH96b, UMS99, Wan95, Wel95, AP90, AP91b, CP90, CNG17, KW92, MOS92, SCA92b, ZS92a, ZS92b, SCA92b]. Interfaces [AS95c, CF94, FAMO99, HF98, NPV96, PAB^{+}97, RT94, Str99c, UI99, Kal91, Kal92]. interfacial [HLS94]. intermediate [AF90, AF91, IT90]. Internal [DM93, GG97, IS96, Leb99, Boy91b, Boy92a, KS92, MDA91, Poy92, QS90]. Interpolating [JBA96, LAE98]. Interpolation [Boy92c, Boy92b, GE92, Par90, RS91, SG92a, SG92b, SM92, Sto92, Fri98, GS96b, Hol96b, Löh95, MD95, OBL93, RHB94, SRF93, Swe96, Tad97]. Interpretation [Boy98, BWJ90, MWJ91]. Intersecting [Gra99]. Interval [Boy94, Boy95a, Chi91, Chi92a, Chi92b, Osb90, Osb91, PO90, PO91]. Interval-Searching [Boy95a]. intervening [Dra90]. intramolecular [DLMN91, DLMN92b, DLMN92a]. Introduction [Adl97, Bai97, Boa97, Dri97, Hir97, Lan97, Lax97, Lil97, Mar97, Puc97, Sem97, Tho97, Van99, Zab97, Zal97]. Invariant [Hob93, Hob92, MSF^{+}95]. Inverse [CC97, Cra94, FA97, För96, GM95b, KL93, KS94, Liu96a, Osb93, PDC91, VSH96, YZ98, Abr91a, Abr91b, Cal90, Cal91, Dar90c, Dar91b, DM91a, DM92a, DM92b, SSF92a, SSF92b]. Inverse-Variable [KL93]. Inversion [Ben95, ZZ98]. invert [KvdVPG95]. Investigation [Dar95, GM95b, JT99, LLLY99, LLR94, PM98, IT90]. Investigations [Kre98]. Inviscid [AC93, Kou97, Mei98, PBBS95, Shu90, SLV90, TH96a, YG95, vBKG97, vdVvdV98, LVS90, PBBS91, PTM92, YKM90]. Involving [Sen99, CHW90, Dra90]. Ion [OS97, RHT96, SKO94, Vu96, Vu98, KDL^{+}92, LL92a, LL92b]. Ion-Driven [Vu96, Vu98]. Ionic [LW96b]. ionisation [MM91a]. Ionized [JSD95]. Ions [BK99, YR93]. Irregular [Boy92c, JC98, PBC^{+}95, VC99, Boy92b]. irrotational [PDC91]. Ising [PH94]. islands [RG90]. Isobaric [FMM99]. Isolated [BST95]. Isometric [CC98]. Isomorphic [Mot98]. Isopycnic [Hig99]. Isotropic [BC98, Hel96, GKM96, HSS97]. Issue [Ano99r, Ano90a, Ano90b]. Issues [Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93g, Ano93h, Ano93i, Ano93j, Ano93k, Ano93l, Ano94a, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano94i, Ano94k, Ano94l, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano95i, Ano95j, Ano95k, Ano95m, Ano95n, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, Ano96i, Ano96j, Ano96k, Ano96l, Ano96m, Ano96n, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Ano97j, Ano97k, Ano97l, Ano97m, Ano97n, Ano97o, Ano97p, Ano97q, Ano97r, Dar95, EL96b, Ano90c, Ano90d, Ano90e, Ano90f, Ano91a, Ano91d, Ano91b, Ano91e, Ano91f, Ano91c, Ano92a, Ano92c, Ano92b, Ano93f, Ano94j, Ano95l]. ITER [Bot96]. Iteration [Ano94t, Fat99, HT97, RC95, MRS94]. Iterations [Fen99, TfNW91, TNW92c, TNW92d]. Iterative [Azm99, BBF^{+}99, BR94, Bör97, DW96a, DW96b, Dar92, EO98, Eye96, För96, KLB99, MH90b, MH91, SS96a, She95, STG^{+}96, WLC96, ZZ98, ZY94, GE92, HLB94, HHG91, HHG92, IM95, Nat92a, OK94, RLM90, VBB90]. Iteratively [nJ93].

J [CNG17, MKM04]. Jacobi [BS98a, LO96, Sen99, vD97b]. Jacobians [BLG97]. jcph.1998.6151 [Ano99-30]. Jet [Pet96, KBS90, SE90]. Jets [HC97]. Johnson [MJP93]. Jones [KKD90]. Joseph [Ske98]. Josephson [DS97a]. Journal [GDP96]. Junction [DS97a]. Junctions [MBO94].

Karhunen [BS91a]. KdV [HK92, HK93a]. Kepler [Boy91b, Boy92a, Poy92]. Kernel [LF96, SRVK96, BKP91]. Kernels [FQ96, SE98]. KFVS [Rav97]. Kind [Jab94]. Kinetic [CB97, CLD^{+}96, GBS^{+}93, Gou96b, HK94, Jun99, KLB99, KM97a, LB94, LDB98, MP97, MK97, PC99b, PX93, RC95, RW96a, Ros90, SXG99, SCW94, Tan93, TS98, TRL99, UL98, VBD99, XP94, XMJ95, Xu99, YCTC97, ZLOT98, Ano91n, GM92, LBC^{+}93, RNSV91, Rus90, TMR92a, TMR92b]. Kinetics [Ano94s, SPCD96, MM91a, VG92b, VG92a]. Kirchhoff [GC95]. Kirchhoff-Type [GC95]. Kirkwood [LPP99]. KKR [BFGG94]. Klein [CG93, DK96]. Kleiser [Wer98, KHW98, Wer95]. Knowledge [KS94]. Kohn [Adh92a, Adh92b, dAT95]. Korteweg [CK90a, Boy96a, HKR99, KL97, Osb90, Osb91, PO90, PO91, Sch92b, Sch92a]. Koster [KU94]. Krylov [ETFS94, Mei98, RKO99, Tid97, VSH96]. Kuramoto [AAP97]. Kutta [Gil97, AL97, AA90, AA91, CS98b, HHM96, Jin95, MR99, SH98b, SH97, Vad97, Wei92, Zho96].

laboratory [HTK90]. laboratory-frame [HTK90]. Laden [DE99]. Lagged [MCPR95]. Lagged-Fibonacci [MCPR95]. Lagrange [Gir97, Gir98, Pri94]. Lagrangian [Mar97, Anc94, BC99, BD99, Bil90, CS98a, CM99a, COBA95, DE99, Duk91, DM92d, DM92e, DL94b, Hal90a, Hal90b, HS93a, HAC97, HL92a, HL92b, LLD92, LLD93, LM95, LH95, LPC^{+}93, Lio95, LMB94, LH90, LL93b, LL94, MS99a, MG97a, MDHW92a, MDHW92b, Oli94, RH91, RHB94, SW99a, SBO93, Smi99, SRBG99, Str99b, Ush96, WEQ^{+}99]. Lagrangian-mesh [MDHW92a, MDHW92b]. Lake [UWH99, WH98]. Laminar [APV98, XS93, hYsY96]. Lanczos [Boy95a, BSPL96, CG97b, WL96, WCSW99, YS98b]. Lanczos-Based [CG97b]. Landau [BCDL97, BC98]. Langevin [MLJ97, SS99b]. Laplace [BB96, DM91a, DM92a, DM92b, GGM93, JMR94, OGS90, OGS91, Spe95, YS98a]. Laplacian [Hui90, Hui91, Taf95]. Large [BCRR98, CL93, Coo99, DFN^{+}99, DA94, GMS^{+}99, GM95a, Gho96, GKDT96, Hal97, HB96, Jor99, KM97b, KWT^{+}97, LML99, Lu99, SP96, TS96, Waj93, WS96, WDE98, Xia99, Mor92, MRD92, PB91]. Large-Domain [TS96]. Large-Eddy [DFN^{+}99, GMS^{+}99, Gho96, Jor99]. Large-Scale [GKDT96, Hal97, HB96, KWT^{+}97, Waj93]. Laser [GBBH96, GMPGV96, KKOF97, MSF94]. Lasers [LW94]. LASY [Tót97]. Latitude [STS98]. Latitude-Longitude [STS98]. Lattice [Abe97, Anc94, AFH94, FH98a, HLD96, HD97, HCD98, HCZ99, HZC^{+}95, HSZ^{+}97, KKH^{+}99, Kla99, KLM95, MS98, MLS99, Pap93, PVVS98, PM99, SBC99, SC96a, SCT^{+}99, vWBS^{+}97, CMK90, GY91, Pap92, UWSB90, UWSB91]. Lattice-BGK [FH98a, KKH^{+}99]. Lattice-Boltzmann [Anc94]. Lattice-Boltzmann-type [Kla99]. lattice-gas [Pap92]. Laugh [Bar90]. law [VC90, VC91]. Laws [Bih96, BT97b, CG97a, CWC99, CL98a, CS98b, CK93b, DW95a, DW96b, FB94, FAH97, Har97, Jin94, Jin95, JL96a, Lav93, Lax97, LWC93, LR94, Noe94, PLD97, PQ94, Sal94, BWJ90, CGSS90, CGSS91, Col90, Dur91, DEO92, Har92a, Har92b, LO91a, LO91b, LY90a, MWJ91, NT90, RT92, SW92, Shu92a]. Lax [Oli94, SCT^{+}99]. Layer [Ber96, Ber95, BJ95, CSN93, CRRR97, Die95, Hu96, LWS98, Mal96a, MM95a, Mas96, PSL99, PZC98, Pet98, QG98, RM93, Str91, TAC98, Zho98, Ber94, Boy91b, Boy92a, Dom90, GMP92a, GMP92b, KG90b, KG91, Mal90, MM91b, Poy92, Str92b, Str92a]. Layered [HKW97, SV97, VSM99]. Layers [AGH99, BCM95a, BGM95, Col97, DW93, HMOS95, Hes98, Iva93, LL95, RK95, SR98, TS96, Wan97, ZOI97, LGC92, Ros90, TSR92]. leap [AA90]. Leapfrog [Aoy95, CAA93, Mat94, Maz97, AA91]. Least [nJ93, OG97, PA95a, VSH96]. Least-Squares [nJ93, OG97, PA95a]. LED [KJ98]. Leer [VK98, PQ94, Swa99]. Legendre [Boy98, Kab96, NC93, She95]. Lehmer [Loe93]. Lemarie [ZLP97]. length [DD92a]. Lennard [KKD90]. Lennard-Jones [KKD90]. Letter [KHW98, Wer98]. Level [AS95c, AS95a, AS95b, AS97, AS99a, ABS96, BS98a, CHMO96, CMOS97, Cho93a, HOS96, HRL99, LZG99, MBO94, PMO^{+}99, PL98, Str99a, Str99b, SSO94, SAB^{+}99, ZCMO96, ZMOW98, ZS92a, ZS92b, DHSS94]. Level-Set [HRL99]. Levels [Wit96b, YR93]. Lewis [NSR92a, NSR92b]. lie [For92a, For92b]. Light [LB93, MP93, lX98, YL98, Kos92]. Light-Cone [LB93]. Light-Front [MP93]. Light-Scattering [lX98]. Lightning [Hag98]. like [AD93, Cra94, LPMS94, LBC^{+}93, Leh99, Lel92]. Likelihood [OP97]. Limit [DW98b, FV97, HCD98, Kla99, UMS99, Bor91a, BLA92a, BLA92b].

Limitations [Wha96]. Limited [Hum96b, MA98, Thu96]. Limiter [AR97b, JP95]. Limiters [Bis95, Hub99, Ven95]. Limiting [Swa99, BR92a, BR92b]. Line [Had99, OGWW98, VRD99, AP90, AP91b, HL92a]. Line-Method [VRD99]. line-segment [AP90, AP91b]. Linear [BBF^{+}99, Bel97, BT97a, CMW95, DK98, ER96, FD99, GS97a, GS94, Gra95, JVS97, LLLY99, Luk99, MWS96, MBP94, PVVS98, PBD94b, SO95, TS98, VM96a, BR99, BB91b, BBF92a, BBF92b, Bor91a, BLA92a, BLA92b, BDR91, BS91a, BB90b, BB91c, CHL^{+}90, FK90, GM92, KS92, MW93, SCMU92a, SCMU92b, VDJ93]. Linear-Discontinuous [MWS96]. Linear-Scaling [LLLY99]. Linearized [Alv91, Alv92, DL93, Hes98, Hu96, TAC98, TK96, Tou97, Tou98]. Linearly [KKOF97]. Lines [ZLP97, FVZ90, Hor90, Hor91]. Link [RMO99]. Linked [Bru93, Bru92]. Linked-Lists [Bru93, Bru92]. Liouville [Can96, JLB90, Ji94, Kob90, Kob91, Oku95, SFB91]. Lippman [GM95b]. Lippmann [Adh92a, Adh92b]. Lipschitz [DK90]. Liquid [AK94b, BCRR98, MWM94, UMS99, MD90, SCA92a, SCA92b]. Liquid-Liquid [MWM94]. list [FK90]. Lists [Bru93, Bru92]. Lithography [AS95a, AS95b, AS97]. LMTO [BS90]. Load [APS97, FLD93, BBK90, BBK91]. load-balancing [BBK91]. Loaded [GDP96, HK98, SK94]. Loading [CSN93]. Local [AAK^{+}93, BS99e, DS96, DS99, ID96, Lee98a, Lee98b, MRS98, PMO^{+}99, Pet98, SSW98, Ush96, Wel95, vdVvdV98, HW92a, HW92b, LJT92a, LJT92b, MF92]. Localization [SV97, Wes92a, Wes92b]. Localized [CM95, WEQ^{+}99, Hor90, Hor91, SVS90, SVS91]. Locally [CDW99b, Min96b, Sak96, SZ97, YMB^{+}91]. Locate [AB97]. Locating [Gre91, Gre92a, Gre92b, Vra95]. Loève [BS91a]. Log [MJP93, vNZ94]. Long [Ara97, AF90, AF91, CM95, JVS97, LNR99, Lil97, NC95, RN94, RS99b, Smo98, GG90, HT94]. Long-Range [NC95]. Long-Ranged [RN94]. Long-Term [Ara97, Lil97]. Long-Time [CM95, AF90, AF91, HT94]. Long-Wave [LNR99, GG90]. Longitude [STS98]. Longshore [RB95]. Lossy [ZLP97]. Low [BDC97, CP99, CJ95b, DOOSM99, DDS97, GH95, Goe95, HHM96, Jun99, JSD95, MFZ97, MP97, SMT99, SH98b, SDP97, SXG99, TEKC92, Tan93, TO98, VRD99, YSD99, EMRS91, IT90, Kle95, LL92a, LL92b]. Low-Dispersion [HHM96]. Low-Dissipation [HHM96]. Low-Dissipative [YSD99]. Low-Frequency [BDC97, DDS97, Tan93, LL92a, LL92b]. Low-Mach-Number [VRD99]. Low-order [TEKC92]. Low-Speed [SXG99]. Lower [BNW96]. Lower-Upper [BNW96]. Lower-Upper/Approximate-Factorization [BNW96]. lowest [MRD92]. LU [Shu92b, Shu92c, YTS92]. Lubrication [RK93, Ver91]. Lyapunov [CZS98, SV97].

MAC [BK91, BK92a, BK92b]. Mach [Jun99, Kle95, SBGK99, SMT99, SDP97, TO98, VRD99]. macro [TZWH92b, TZWH92a]. macromolecule [JB91, JBvK^{+}91]. macroscopic [LDB96]. Magnetic [BS99b, BS99c, CSN93, oJ95, oJ97, JB95, PFS98, SW99b, TSW95, VB95, ZV97, LH92, RP90, RG90, RP91]. Magnetised [Sco96]. Magnetization [Pri98]. magnetized [MS91b, MS92b, MS92a, OsKB^{+}92]. Magnetohydrodynamic [BS99a, BS99b, BS99c, FGM97, KB99, Leb99, SWD95, TO96, HSN90, HB90a, HB91]. Magnetohydrodynamical [DW98c]. Magnetohydrodynamics [AK94b, Asl99, BKP96, BD99, DW94a, DW94b, JcW99, JSE97, LMS98, MR98, PRL^{+}99, SLMS98, SL98, Xu99, Bra90, Bra91, DeV91, TLES93, YM91, YM92b, YM92a, ZC91, ZC92]. magnetohydrostatics [Cal90, Cal91]. magnetorheological [LRJ^{+}99]. Magnetosphere [KKK95]. Magnets [Bot96]. Magnus [YHW90, YHW91]. Maintaining [BS99a]. Manifolds [AHS96, Hob93, KO98, Dvi90, Dvi91, Hob92]. Mantle [BY94]. Many [BF97, EA97, EM94, EM95, Lan97, MRC93, Zan99]. Many-Body [BF97, Lan97, MRC93]. Many-Particle [EA97]. Map [BST95, TSW95, GGM93]. Mapped [GS95, PBD94a]. Mapping [Boy94, DE93, DE99, KM96, KW93, LL95, LO93, SSR96, DP91, DP92b, DP92c, Kos92, LM92, OB95]. Mappings [BT92a, BT92b, Vra95, Boy91b, Boy92a, Dar93, Poy92]. Maps [GS96a, HPG98, Hob93, KO98, Dvi90, Dvi91, Hob92]. Marangoni [CT98]. Marching [LZ96, TCS97, YB98, Wei92]. Marker [TM94]. Masking [SS99b]. Mass [AMP^{+}98, BR99, BSB92, Jai97, LK94a, STW97, TWV97, Wel95, CYD92, KDL^{+}92, Lar91]. Mass-Damped [TWV97]. Masses [CS98a]. Massive [MP93]. Massively [LPR96, SDG94, Vu98, SBGM92]. Master [SVG93, SVG92]. Matched [AGH99, Ber96, Col97, Hes98, Hu96, JYH93, MO95, PSL99, Pet98, QG98, TAC98, Ber94]. Matched-Phase [MO95]. Matching [ART95a, ART95b, PBD94b]. Material [Bis95, LPC^{+}93]. Material-Based [Bis95]. Materials [AK99, HW97, HSS97, DM91a, DM92a, DM92b, TZWH92b, TZWH92a]. Mathematical [AG97, SG94, Lav90]. Matrices [CG97a, LK94b, LR92, Mor92, MRD92, OS90]. Matrix [ASTAS91, BCRR98, BCM95b, BCM99a, BD94b, BL93, DHSS94, GS95, HL98, Iva93, Jai97, KHW98, LBL98, RH98, RMSB96, Wer95, Wer98, AAP^{+}95, BL92, BL90a, BL90b, BSB92, Dau92a, Dau92b, IM95, KH90, KH92, Nat92a, Wei92]. Matrix-free [LBL98]. Matter [Bör97, CK93a, CK92]. Maximum [Osc94, OP97, Rei92a, Rei92b]. Maxwell [ADH^{+}93, ADS96, Bis95, DF98, GK98, GMPGV96, Her93, Hil97, Hol96a, HS99c, LM90b, Liu96b, Mad95, PZC98, Sha95a, SF96, Sha99]. May [AAK^{+}93]. Mean [Fri98, GBS99, Ruu98b, BLMR92, Blu92]. mean-field [BLMR92, Blu92]. Means [Abe97, CS98a, JBA96, Reb93, RH99, Dau92a, Dau92b, GHS93b, HB90b]. measure [Gre90]. Measures [Kal97]. Mechanical [Bar94, KKOF97, LLLY99, SMM^{+}90]. Mechanics [HMD93, Ske98, Wri98, Val91a, Val91b, ZDG95]. Mechanisms [Gou96b].

Media [ABBM96, ADS96, BBGL93, CE97, Dur93, GP98a, GGK95, GKG95, HKW97, HT97, HW97, JLL97, LLK98, Pap93, SV97, UL98, ZS97, Ben92a, Ben92b, Bor91b, Bor92a, Bor92b, Dra90, Hig92a, Hig92b, Nic93, Pap92, RS90]. Median [oJ95]. Medium [FK97, Iva93, KI99]. Melt [BDS^{+}99]. Melting [KE93]. Melts [KWT^{+}97]. Membrane [CV97]. Membranes [DMW99, SK94]. memoriam [Ano91m]. memory [Nat90, Nat91, Par90]. memory-saving [Par90]. Merging [LNS^{+}94]. Meridians [Boy96b, Boy97a]. Mesh [Ano99-30, AE95, BS99b, BS99c, BCR99, CHR99, CH99, CP95, Coo99, DP93, Duk95, FGM97, GBCA99, HLD96, HT97, HRR94, Hum96b, JFH^{+}98, Knu95, LP97, LWC93, NPV96, PGD99, QS98, Reb93, SM98a, Tho97, Vel93, Win97, AS90, AS91, Ben92d, Ben92c, DP92a, DHS91, FM92, Fra92, MK90, Mav90, MDHW92a, MDHW92b, SW92, Wes94]. Meshes [Abg94, HMOS95, Her93, Hew97, HS99b, KJ98, LQK98, Mav98, MRS98, OS98, OG97, PG97, UI99, VM93, CH90a]. Metal [AK94b, CKR93]. Metals [GBBH96]. metastable [SCA92a, SCA92b]. meteorology [KM92]. Method [AG97, Abd95, Abe97, AS95c, AN98, ALTP98, AM96, ABC94, ABC^{+}98, AQ98, ABLD94, And94, AR96, ALW94, AAL97, Ano99-30, AS99b, AMP^{+}98, ADH^{+}93, ADS96, AQ99, AK99, BH95, BST95, BPS95, BY94, BR99, Bar94, BR97a, Bel97, BD97a, BD94a, Ber95, BMZ95, BW98, BL93, BLL99, BBMB97, Bör97, BB97a, Bot98, Boy95a, Boy95b, BL94, BS98b, BSPL96, BL98, BW95, Bru93, Bry97, BPM98, CLO93, CZ98, CLNT98, Can96, CG93, CHR99, CBL95, Cha95, CWF96, CWC99, CMOS97, CZS98, CM99a, Cho97, Chr97, CC98, CK93a, CKR93, CS97b, CS98b, CDW99a, CP99, CJM97, CLD^{+}96, CEG^{+}97, CDT98, CV97, CD95b, Cra94, CR99, DW94b, DBS94, DVD93, DDD98, DDS97, DZ96, DB98, DK96, DB96, DFFG96, DK98, Dob99, DS97b, DIV98, DF98, DA94, DL94a, DS94]. Method [EOS98, EP96, EM94, EHM97, Eyr94, FV97, FMO98, FAX99, FAMO99, FA97, Fen99, FKM99, Fey98a, Fis97, FPQ93, Fuj95, GACN93, Gar98, GTD98, GA96, GSB98, GK94, GH95, GH97, Gir98, GGK95, GKG95, GEK^{+}97, GKKR99, GHKH93, HIY95, Han93a, HS93a, HLD96, HD97, HCD98, HML99, Hen94, HJ98, Hew97, HK98, Hir97, HAC97, Hob93, HLR99, Hol96a, HGFH95, HZC^{+}95, HLOZ97, HW97, Hu95, HHR99, HS94, HK93b, IH95, IR98, oJ95, oJ97, Ji96, nJ93, JC98, JR99, JT96, JSD95, Kab96, KKH^{+}99, KD97, KOS^{+}96, KI99, KHW98, KPP99, KKOF97, KK96, Kop96, Kop98, KTE93, KM97a, KMS99, Ku95, KE93, KCV99, KP98, LPMS94, LPBS96, LW95, LLD93, LS93a, LEI96, LLR94, LL96, LWC93, Lio95, LMB94, LZ96, Loe93, LL94, LKK99, LS98, LDB98, LO93, Luk99, LBL98]. Method [MH94, Mal96a, Mal96b, Man93a, MV99, MKM99, Mar93, Mar97, MG96, MG97a, MM96, MA95, MM95b, MM97a, Mat94, MS98, MLS99, MPPZ96, MW94, MP96, Min96b, MS99b, MLBW97, MBP94, MA96, MJPC99, NP94, NC95, NC93, OS98, OBB98, Oli94, OP97, PC99a, PBC^{+}95, Pen95, PMO^{+}99, Per93, Per95, PQ94, PH95a, PCLC97, PVQ97, PL98, Pop95, Poz99, Pri93, Pri94, PAB^{+}97, Puc97, QTL98, QS98, RK93, RN94, Rea97, RF96, RL97, RKO99, Rie99a, RH99, RW96a, RW96b, RSW98, RPB99, RIP96, RY94, Rus93, SB96a, Sak96, SZ97, Sal94, SWS96, SA99, Sem97, SNU98, STW97, Sha95a, She95, SK93, ST97, SCR97, SRBG99, SG98, SW99c, SL98, Sun96, SRA98, STS^{+}97, SF99, Tab96, TKT97, TV94, TS99, Tau94, TTI97, TM94, TPRC96, TK99]. Method [TC98, UI99, VPS95, VP96, Vel93, VM96b, VRD99, VSG95, Vra95, VSH96, VSM99, WNY93, Wan97, WK99a, WDE98, Wer98, Win95, WEMH99, WS93, WCSW99, XS93, Xu99, YB98, YZ98, YMUS99, YMAC99, YPH94, Zab97, ZSK94, ZW97, ZZ98, ZY94, ZP95, ZWS98, Zhu95, ZCH96, ZOI97, ZV97, ZRR99, vD97b, vL97a, vL97b, vdVvdV98, AAP^{+}95, AN90, AS90, AS91, AB94, BP90, BB90a, BB91a, BKM90, BM92, BB91b, BBF92a, BBF92b, Bey91, Bey92, Bil90, BL92, Bor91b, Bor92a, Bor92b, Bra90, Bra91, BKZ92, BS90, BR92c, BR92d, Bru92, Bue91, BSB92, But90a, But91, CK90a, Cal90, Cal91, Car91, CRS90, Chi91, Chi92a, Chi92b, CPP93, CK92, Cot90, Dar90b, Dar90a, Dar91a, Dav91, Dav92a, Dav92b, DR90, Dom90, Dra90, DZ91, Flå92b, Flå92a, Fog92b]. method [Fog92a, For90, FM92, Fra92, Ful93, FVZ90, GCMR90, GL95, GSY92, Got92, Gre91, Gre92a, Gre92b, GHS93b, GS93, yGmC92, GC92, GMCH92, HM95, HLB94, Har92a, Har92b, Hei92, Hob92, Hoo90, HM90, HL92a, HL92b, IM95, Jac90, JLB90, JC92, KM92, Kar91, Kar92, KR95, KG90a, KEHK90, KKD90, KF91, KEHK91, KvdVPG95, KG92a, KG92b, LLD92, LS92a, LS92b, LL91, LKB90, Lif92, LH90, LP90, Mak99, MH92, MKM04, Mar90, MM91b, MMB90, Mel91, Mit92, Mor90, Mor92, Nat92b, NP92, ND90, OB95, OGS90, OGS91, PH94, PTM92, Pel91, PA95b, Ram92a, Ram92b, RF90, RF91, Raw90, Raw91, Rei92a, Rei92b, RT92, RNSV91, RKV90, RIW90, RKV91, RLM90, Rus90, SH90, SD93b, SSB90, SSB91, Shu92a, SCC93a, SA92a, SA92b]. method [Sul91, SB91, TMR92a, TMR92b, Tay91, TNW92a, TNW92b, TKKB92, Tes92, UWSB90, UWSB91, UT91, UT92b, UT92a, VY91, VY92, VBB90, Yan90, YR98, YHW90, YHW91, YMB^{+}91, ZC91, ZC92, ZB92a, ZB92b, dFSS92]. Methodological [Sch99]. Methodology [CGA94, Jor99]. Methods [AS99a, ANL94, AK95, Anc94, ABS96, Azm99, BM94, BBF^{+}99, BYMZY95, BLG97, BCM95a, BS93, BS96, Boy98, BBGL93, BGP98, CLO93, CS93a, CN95, Can95, CG96, CHMO96, CS93b, CK93b, Cot96, DE93, DH95, Don94, EL97a, ETFS94, Eye96, FMO97, FD93, FJ98, GS95, GS96a, Gir97, Gue94, GS96b, GLN^{+}99, GO95, HMD93, Han93b, HT95, HKNT98, Hei93, HK95, Her93, HKR99, Hol96a, HRR94, HW96, JDC94, Jin95, Kob99, KW93, KLP94, LO96, LL99, LBB94, LS96, LeV98, LD93, LP97, LL93a, LZS98, LQK98, Lu99, LRW95, Mad95, kM93a, Mat97, MR99, Mei98, Min96a, MH93, NM95, Nic98, Nor96, NC99, OBL93, Ohw98, ÖM95, PB98, PK95, PH95b, PG97, RCA93, SO95, Sam97, SPCD96, SS99b, Ser96, SvD96, SM98a, SRF93, STG^{+}96, Sil93, Sim99]. Methods [Sjö95a, SDP97, SC96a, Str95, Str96, Str97, Str99b, Str99c, SW96, Tad97, TCS97, TZA96, Tem96, TL98, Ton96, Vad97, Wal94, WF99, WEQ^{+}99, Whi94, WL93, Wir97, WLC96, XMJ95, XP99, YP98, Yee97, YSD99, Yin96, Zho96, ZHL96, vD97b, Bar91b, BM91a, BM91b, CS92a, CP90, Cas90, Che91, Col90, CJ95a, EMRS91, EHW90, EHW91, FVZ90, GY91, GKM96, GMP92a, GMP92b, ISW92a, ISW92b, JE90, KIO91, LO91a, LO91b, LM90a, LM91, LY90a, LM90c, LG94, Mal90, kM90, kM91, Mao92b, Mao92a, Mil91, MF92, MNR90, MRS94, Nor91a, Nor91b, PC92b, PC92a, RCA92, RS91, Rob90b, RS94b, SSF92a, SSF92b, SK90, Sil92, SMR91, SM92, Sto92, TfNW91, TNW92c, TNW92d, WMG^{+}91a, WMG^{+}91b, ZS92a, ZS92b]. Metric [Sho93]. MHD [AK97, Asl99, BDC97, Bra90, Bra91, CG97a, CHL^{+}90, HCJ92a, HCJ92b, KLBD93, KGH^{+}98, KvdVPG95, LL91, NMW^{+}96, PV99, PBD94b, PS93b, SS96b, TT90, TT91, Tan94, TA91b, TAD92a, TAD92b, Tay94, WK99a, van93]. micro [TZWH92b, TZWH92a]. microinstability [MS91b, MS92b, MS92a]. Microstructural [JLL97]. Microstructures [PGD99]. Microwave [CC97, Luk99, PL98]. migration [LZG99]. Mimetic [HS99c]. Minima [Abd94]. Minimal [Cho93a, PH94]. Minimization [AT99]. Minimizers [NMB94]. Minimizing [LR94]. minimum [Sal91, dG92a, dG92b]. MINRES [YS98b]. mirror [BLK90, PC92b, PC92a]. MIS [XPK90]. Mises [HN95]. Mistaken [LR98]. Mixed [CP98a, CSW95, DK98, Dru99, Dur93, EP99, Gar98, MWW96, YMAC99, LM90b, MLB92, Shu92a]. Mixing [CDR99, GE92, GMP92a, GMP92b, KG92a, KG92b, Mik90, Mik91]. Mixture [Shy99]. Mixtures [JMT97]. MKdV [GGG94]. möbius [HS90]. modal [AH99]. Mode [AK94a, Bel97, CAA93, CGKC93, DOOSM99, HK98, ID96, KGH^{+}98, MPPZ96, AA90, AA91, CGKC92]. Model [AS95a, AS95b, AS97, ARV99, Asl99, Bot96, CMMH94, CB94, COBA95, DHS94, Die95, FV97, FK97, GBBH96, GT97, GZ95, Hag98, HCD98, HM99, JVS97, JLM^{+}96, KLBD93, Kla99, LBB94, LTE93, LIH97, LZ96, Ma93a, MP93, MR98, PM99, Pop95, Pri96, RLT93, RB95, RW96a, SS99c, SMW99, Shu95, SR98, SH94, SG94, SP96, Tem96, TK96, TM93, Ver97a, Ver97b, VS95, VBD99, WH98, WJTP96, WP97, Xia99, YH95, YS98a, dM97a, AP90, AP91b, Bey91, Bey92, BB90b, BB91c, CWG90, CWG91, CYD92, CGSS90, CGSS91, FS91, FS92, HWY90, HWY91, JC92, KDL^{+}92, KST90, KST91, LG94, LL92a, LL92b, MGP91, MLB92, Nad95, PH94, SAB95, VNC92a, VNC92b, Lya99]. Modeling [AMP^{+}98, ABS96, BCRR98, BDS^{+}99, Bot98, DFFG96, Dur93, FH97, GHS93a, GNH96, Hal97, HDL99, HB96, HdS97, Hig99, Hum96b, Jac99, KKK95, KB99, MA98, MH94, MFZ97, PL98, RT94, SS99a, SBC99, Sun96, YL98, Zan97, ZLP97, Ben92a, Ben92b, BBC^{+}90, BKZ92, PC92a, TZWH92a, Zan92a, Zan92b]. Modelled [GGZ93]. Modelling [AHH94, BL93, CM99b, DDS97, LNS^{+}94, LW94, PC99b, TP97, ZS97, BBC^{+}91, BL92, LS92a, PC92b, TZWH92b]. Models [CS97a, Cot96, DB98, FH98a, HMD93, HB95, HSZ^{+}97, Hum96a, Kar97, LPMS94, LW97, Mur96, PVVS98, PMR97, ST95, Sil93, TG96, UWH99, Waj93, WF99, GY91, GG92b, HCP93, KL92b, Sil92, SNL90]. Moderately [DECB98]. Modes [AHH94, BM94, Gel99, PR93a, PBD94b, SVS90, SVS91]. Modification [SB98, Yan90]. Modified [Bar90, DÜ92a, Hag94, HBF93, KTE93, LCM96, PSL99, AN90, Cha90, DÜ92b, Gra90]. modulated [Bar91a]. Modulations [FCC97]. Mohr [GS97b]. Moisture [Tol94]. Molecular [Ano99r, BGGT90, BLL99, CP98a, CSZ97, DOOSM99, DM93, DL94b, DMW99, ESH93, EH93, GKDT96, GBS99, Jai97, KSB^{+}99, LPP99, MRC93, Maz97, Maz94, Pli95, Rei99, SMSS98, SSB^{+}99, Sch99, SHF97, SW99b, WDE98, BBK90, BBK91, CYD92, JVR93, MH92]. molecule [AAP^{+}95]. Molecules [WKHS97, Dav91, Dav92a, Dav92b]. Moment [BHJU99, ZV97]. Momentum [Ben92d, Ben92c, KLM95, dFD^{+}94, WS91, WS92a, WS92b, Zan92a, Zan92b]. Monolayer [KE93]. Monotone [Gro91, Leb99, TN98, Gro92, SG92a, SG92b]. Monotonic [COBA95, SBO93, GG92b, SM92, Sto92]. Monotonicity [SH97]. Monotonicity-Preserving [SH97]. Monte [AT99, BPS95, BRL91, Boy90, Boy91a, BGM91, CB96, COBA95, DR90, DB96, EM94, EHM97, GBCA99, Gre98, HBF93, KNR99, KE93, Kuh96, Lif92, MSV98, MC95, Nea94, Nic93, PC99a, Pri90, Pri91, Rie94b, STAS91, SH90, Sch95b, SFHD99, STS^{+}97, UL98, Val91a, Val91b, WONM96, WVHS98, XP99]. Motion [ALTP98, Ara97, BW95, Cor96, DOOSM99, DNN93, EB94, HG94, HOS96, HH98, Lil97, MBO94, NB98, QTL98, Ruu98a, Ruu98b, RMO99, SG98, VB95, Yin96, ZCMO96, CWG90, CWG91, MD90, MDA91, MOS92, TKKB92, Tes92]. Motions [CS98a]. movement [HGH90, HGH91]. mover [Fri90]. Moving [AS94, Ano99-30, BW95, BCR99, CHR99, FAH97, FHKZ97, GDP96, Had99, HLOZ97, HRR94, JL96b, KC97, LP97, LKK99, QS98, SSR96, SS94, Str99c, ZC94, ZW97, ZRR99, vM95, CW92a, CW92b, CS92b, CS92c, FVZ90, MB92, RR90, RR91, Rob90a, VD90, Wat92b, Wat92a, Wes94, ZB92a, ZB92b]. moving-finite-element [ZB92a, ZB92b]. moving-grid [FVZ90]. MP [Par90]. MPDATA [SM98b]. mufti [DHS91]. mufti-mesh [DHS91]. Mullins [ZCH96]. Multi [AH99, BBGL93, Bru93, CSW98, CH90b, CH91, FPRB90, FPRB91, GM95b, Han93a, HJ98, KCV99, LO98, PFRB93, RY94, SBGK99, SWT98, TKKB92, Tes92, WEQ^{+}99, BBK91, Bru92, DD90, DD91, JJK92, Lar91, SD93b, UT91, UT92b, UT92a, EH93]. multi-component [Lar91]. Multi-dimensional [CSW98, GM95b, LO98, SBGK99, SWT98, WEQ^{+}99]. Multi-domain [CH90b, CH91, Han93a, HJ98, TKKB92, DD90, DD91, SD93b]. Multi-fluid [KCV99, UT91, UT92b, UT92a]. Multi-grid [BBGL93]. Multi-modal [AH99]. Multi-particle [EH93]. Multi-processor [Bru93, BBK91, Bru92, JJK92]. Multi-scale [FPRB90, FPRB91, PFRB93]. Multi-zone [RY94]. Multiblock [ST95]. multicolor [CMK90]. Multicomponent [Abg96, BPT95, EG95, Kar94, Shy98, Shy99, Ton96, Xu97, CYD92, RS90]. Multicomputers [Sha96]. Multiconstrained [TLES93]. Multidimensional [BT97a, BT97b, CK96, CWC99, CS98b, Col90, DW94b, DW98c, Fey98a, Fey98b, GNHP95, HB97, Hub99, JM98, KP97b, LeV97, Noe94, PQ94, RvR93, SMD98, Swa99, Thu96, Tót97, VP97, Wit96a, Yee97, DeV91, Duk91, DM92d, DM92e, Mel91, SG90]. Multidisciplinary [Shu95].

Multidomain [DBS94, Kop90, Kop91, Kop94, KK96, Kop96, Kop98, LW96a, Mal96a, PVQ97, SP98]. Multifluid [SA99]. Multifrequency [Win95]. Multifrequency-Gray [Win95]. Multigrid [ARB96, Bar91b, BS96, BL98, CW90, CWF96, CSW95, DS99, DF96, Dic90, DIV98, Edw96a, GSB98, GKZ97a, GKZ97b, Hei93, JDC94, JV95, Kor90, LL99, LS93a, LL93a, LL95, LZ96, LZS98, LF96, Mav98, MDH98, Oos97, OGWW98, Pai97, PB98, PG97, RS95, RKO99, SMJ98, SR97, SDP97, VRD99, YM95, YSMG97, Zha94, Zha98, AB94, BY92, Bra92, Cal90, Cal91, Dav91, Dav92a, Dav92b, Hei92, KW92, LS92b, Luc90, Luc91, Mul92b, Mul92a, SK90]. multigrid-continuation [SK90]. Multilayer [LBB94]. Multilevel [Ano97-29, BL90a, BL90b, HT97, KLB99, SB96b, Tem96, VPS95, VP96]. Multimaterial [CS97b, FMM99, FAMO99, RKO99]. Multimaterials [DW98a]. multimodal [SSF92a, SSF92b]. Multiparticle [lX98, Zin94b]. Multiphase [HCZ99, HSZ^{+}97, LNS^{+}94, Ruu98a, SA99, UKSTST97, ZCMO96]. Multiple [AWY99, BCM95a, BGM95, BS93, Hof97, MBO94, MP96, Rei99, SCT^{+}99, ZC94, ZW97, ZLP97, Sch92b, Sch92a]. multiplication [BL90a, BL90b]. multiplicative [DM90, DM91b]. Multiply [DHSS94, KM95, vGVO98, BW90, BW91, GGM93]. Multiply-Connected [KM95]. Multipole [ARTAAA97, BFO99, Boy92d, CDM98, LRW95, CGR99, CPP93, LDB96, Mak99, YR98]. multipole-based [LDB96]. multipoles [Mak99]. multiproccessor [Nat90]. multiprocessor [BBK90, Nat91]. Multiresolution [Bih96, BS99f, Har94, Sjö95b]. Multiscale [HW97, LAE98, WA95, Wir97]. Multispecies [Mat94]. Multisphere [lX96, lX97]. Multisphere-Scattering [lX96, lX97]. Multistage [SC97]. Multivalued [Ben96]. Multiwavelets [AIV99]. MUSCL [Hub99]. MUSCL-Type [Hub99].

N [KE93, Swe94]. N-Body [Swe94]. NAMD2 [KSB^{+}99]. Nature [CL98b, DDD98, RR90, RR91]. Navier [Gil97, ANL94, AM96, ABC^{+}98, AB94, Arm94, AS99b, BB90a, BB91a, BR97a, BK94, BTW96, BLT96, BJ90, Cha95, CPJ90, CB97, CH90b, CH91, Cle97, Cod99, CJ95a, Cot90, CD95b, CMM93, DBS94, Dau92a, Dau92b, DF96, DD90, DD91, DH95, Dom90, DIV98, ETFS94, EP99, Fis97, GTD98, GB97, GHKH93, yGmC92, GC92, Gup90, Gup91, Hen94, HI98, Hwa94, ID96, IM95, Jac99, JB93, Joh93, JR99, KIO91, Kla99, Kop98, Kor90, KL92b, LM90a, LM91, LM97, Lee98b, LZ96, LS98, LDB98, LR90, LR91, Luc90, Luc91, MDB91a, MDB91b, MV99, MH90a, Mar94b, MH90b, MH91, MP97, NPC93, Nor96, Nor98, Nor95, NC99, Pri95, PM98, RF95, RKV90, RKV91, RvR93, Rus93, SP98, SPC92, SPC93, SD93b, SNU98, STW97, SS99c, Shu91, Shu92b, Shu92c, Sjö95a]. Navier [SA90a, SMR91, Taf95, Tau94, TfNW91, TNW92c, TNW92d, Tid95, Tou97, Tou98, WNY93, WS93, XP94, YYCH98, ZSK94, Zha94, ZWS98]. Near [Bas92, GS96a, JSD95, NP94, Rob99, ZYKC98, BG92, EMRS91, HS90, KKD90, Nak95, PRMV90, Rob90a, VBB90]. Near-Boundary [GS96a].

Near-Caustic [Bas92, BG92]. Near-Equilibrium [NP94]. Near-Field [Rob99]. Nearly [Ano97-29, Boy96a, SB96b]. Negative [CP98b]. neighborhood [Hor90, Hor91]. Neighborhoods [Swa99]. nesting [CH90b, CH91]. Network [ESH93, EH93, HGFH95, MML98]. Networks [Bot98, BW95, Kar97, ZC94]. Neumann [Bal95, GPWZ96, GGM93, HPG98, MBF94, Sco98, WNY93]. Neumann-Type [MBF94]. Neural [LK90]. Neutral [Azm99, BO96, PB98, Nic93, Val90, VNC92a, VNC92b]. Newton [Ano94t, BB90a, BB91a, Fen99, GS96b, KPC92, KPC93, KM95, LPMS94, LL96, PCLC97, RKO99, SSR96, STW97, Tid97, XS93]. Newton-like [LPMS94]. Newtonian [CK92, CK93a, MNP90]. Nicholson [Alv91, Alv92]. NLTE [KH90, KH92]. No [BGH^{+}99, GHS93a, KHW98, MB94, Wer95, Wer98, HLB94, LM92]. No-Slip [GHS93a, MB94, Wer95, HLB94]. Node [GN96, HGH90, HGH91]. Node-Centred [GN96]. node-movement [HGH90, HGH91]. Nodes [Kar97, MMS90a]. Noise [MO95]. Noises [RS99b]. Non [ABBM96, Abg94, BK93, BM94, BR99, Bar94, CK93a, DW95b, DA94, ES94, FAMO99, Fer95, Fis94, Fri98, Giv90, Giv91, HEOC97, HS99b, HSS97, nJ93, KT99, LPBS96, LLD93, LOC94, MBV93, Mad95, NT90, OK94, RCA93, Shu97, SK93, Sun96, VDJ93, ZSK94, vBvNW95, BB90b, BB91c, CK92, GM92, GC90, LLD92, LMP92a, LMP92b, MNP90, MH90a, PL92a, RCA92, Shu90, SLV90, Shu91, Shu92b, Shu92c, SCC93a, SCMU92a, SCMU92b, TA91b, TAD92a, TAD92b]. non-Boussinesq [LMP92a, LMP92b]. Non-Cartesian [Fis94]. Non-conforming [KT99]. Non-diffusive [nJ93]. non-equilibrium [GC90, Shu90, SLV90, Shu91, Shu92b, Shu92c, SCC93a]. Non-Hermitian [LPBS96]. non-inductive [TA91b, TAD92a, TAD92b]. Non-isotropic [HSS97]. Non-iterative [OK94]. Non-linear [BR99, VDJ93, BB90b, BB91c, GM92, SCMU92a, SCMU92b]. Non-Newtonian [CK93a, CK92, MNP90]. Non-Orthogonal [ABBM96, Mad95]. Non-Oscillatory [Fri98, Abg94, DW95b, ES94, FAMO99, HEOC97, HS99b, nJ93, LOC94, NT90, Shu97, SK93, PL92a]. Non-periodic [BM94, MBV93]. Non-reacting [LLD93, LLD92]. Non-Reflecting [Fer95, Giv90, Giv91]. Non-separable [Bar94]. Non-staggered [ZSK94, MH90a]. Non-symmetric [DA94]. Non-uniform [BK93, RCA93, vBvNW95, RCA92]. Nonanalog [UL98]. nonconforming [KR95]. Nonconservative [HP92, HP93]. Nonconvex [dG96]. Nonequilibrium [LGC92, Zho96, Boy90, Boy91a]. Nonevolutionary [BKP96]. Nonideal [GP98a, JSE97]. Noninterpolating [Oli94]. Nonisothermal [CF94, HF98]. Nonlinear [AQ98, AAL97, Aoy95, BL99, BYMZY95, BK97a, BKV98, BR94, Boy95a, Boy96a, Boy97b, BCR99, BCT98, CLNT98, Can95, CG93, CJS99, CBS93, CC97, DW98a, DZ96, DK96, DM97b, Die95, DIV98, FKM99, FD99, GP98b, HO96, HS99a, LM95, LEI96, MLBW97, ÖM95, Osb90, Osb91, Osb93, PNC94, PBBS91, PBBS95, PO90, PO91, RFL93, SH99, SMSS98, Str95, Tay94, Vad95, Vra95, WONM96, YH95, YG94, ZV97, dG96, Alv91, Alv92, BO92, CK90a, CWG90, CW90, CWG91, CHL^{+}90, CHW90, FM92, Fra92, GCMR90, GL95, LS92a, MS90a, MS91a, MLB92, PM90, RLM90, SW92, TSS92, YSG91]. Nonlocal [BH96, Boy95b, Boy96a, Boy97b, LPBS96, SDI98, Tsy95]. Nonnormality [HO96]. Nonoscillatory [LO96, SG90, Sur94]. Nonparametric [WEMH99]. Nonperiodic [Cle97, GTD98]. Nonpolynomial [Wit96a]. Nonprogressive [WJTP96]. Nonreacting [BS99e]. Nonreflecting [GC95, GK95, GK96, GK98, JB93, LT96]. Nonseparable [DB97]. Nonsmooth [ER96, PK95, WSK99]. Nonstaggered [EL97a, SSB90, SSB91]. Nonstationary [LW94]. Nonsymmetric [Wit96b, Mor92, Nat92a]. Nonuniform [CF99, Tho97, Win97, HLD96]. Nonzero [TS96, TSR92]. Normal [DOOSM99, KGH^{+}98, SSW98, WJTP96]. Normal-Mode [KGH^{+}98]. Normalisation [Bri95]. Nosé [BLL99]. Note [AC95, Ano91o, Ano91p, Ano91n, Ano92k, Ano94t, Ano94s, Rap93, vNZ94, LR92]. Notes [Ano90p, Ano90q, Ano91q, Hor90, Rie99b]. Notion [LR98]. Novel [HCD98, Nun93, TYS95, Wan97, DK92a, DK92b]. Nuclear [KP97a, BLMR92, Blu92, LBC^{+}93]. Nucleon [dAT95]. Number [BS99c, DA94, LB94, Leb99, MCPR95, MFZ97, RSW98, SBGK99, SMT99, SDP97, TS96, TO98, VDJ93, VRD99, lX97, vL97b, Ano99-30, CV92a, CV92b, Kle95, TSR92]. Numbers [Jun99, ASTAS91, BES90, IT90]. Numerical [AHS96, AHS97, AK94a, AC95, AHH94, AS94, AWY99, Ara97, AK97, ADS96, ADS97, BST95, BY94, BS98a, BR97a, Bel97, BD94a, BBC^{+}90, BBC^{+}91, BK97a, BD94b, Bot96, Boy96a, BDR91, BW95, BFGG94, Bry97, BCDL97, BC98, BCT98, BDML94, CN95, Can95, Can96, COS^{+}98, CW98, CH99, CBL95, CGA93, CF94, CM94, Cho97, Chr97, CGSS90, CGSS91, CGKC92, CGKC93, CKQ^{+}93, Cod99, CDW99a, CDW99b, CR96, CEG^{+}97, CK93b, CJR91, CJR92a, CJR92b, CL98b, CS93c, DW98a, DW98b, DE93, DNN93, Die95, DS97b, Don94, DQ95, DM93, DE99, DK90, Dvo91, DK92d, DK92a, DK92c, DK92b, ES94, ES98, EGS96, FEO95, FMO97, FT96, FPB98, FC95, GRB^{+}95, Gho96, GM95b, GKL^{+}96, GZ95, GMPGV96, HF98, HTK90, HKNT98, HML99, HA93, HK98, Hol96a, HS90, Hum96b, HC93]. Numerical [HSS97, Jab94, JT99, JL96b, JL96a, Kal91, Kal92, Kar97, KKK95, Kla99, KG90a, KCJ95, KNW99, KW93, KRGv93, KMM96, KM97b, KB99, LPMS94, LSLG97, LGC92, LBB94, LS94, LLR94, LK94a, LJG96, LZG99, Lil97, LLT94, Liu96b, LK98, LK99, LDB98, LR94, MD98, Mal90, Man93a, Mar94a, MPR99, MBI^{+}97, MW94, MMS90b, MMS91, MB96, MDH98, NWK98, NCF96, NMB94, Nun93, OS98, ÖM95, OD99, ÖLJ98, Pap92, Pap93, PB91, Pen95, Pet94, Pet96, PZ98, PVQ97, PMR97, PX93, Pri96, PO90, PO91, Puc97, Qi93, QW93, QS98, QR90, RM93, RCA92, RCA93, Ram90a, Ram90b, Ram97b, RL97, RH99, RW96a, RFH93, SB96a, Sak96, SZ97, SB95, SKR^{+}93, SA95, Sem97, SBKH92, SMW99, SKO94, SI94, Sim93, ST97, Sim99, Sjö95b, Ske98, ST93]. Numerical [SRBG99, SVG93, SK92, SFB91, Sun93, SWD95, SC96b, SRT98, SCC93b, Tad97, TG96, TA91b, TAD92a, TAD92b, Tau92a, Tau92b, TH96b, Tho97, TO96, TP92a, TP92b, TZWH92b, TZWH92a, TRL99, Ush96, Vad95, VDJ93, VSG95, VB95, Wal94, Wan99, Whi90a, Whi90b, Whi94, Win97, WK96, WKK98, Wri98, WCO94, XPK90, XP94, XP99, Zab97, ZB96, Zho98, ZS97, dG96, dM97a, dNPT95, dFD^{+}94, tVB92a, tVB92b, van93, ARS91, BLK90, BB91b, BBF92a, BBF92b, BW90, BW91, Bue91, CRS90, CG97c, CHW90, Dar90a, Dar90c, Dar91a, Dar91b, Dav91, Dav92a, Dav92b, DHS91, DZ91, Flå92b, Flå92a, FP92, FVZ90, GB90, GBS^{+}90, GBL^{+}92, GS93, HPS92a, HPS92b, HG91, HG92, Hoo90, HL90, HL91, IM95, IT90, JE90, LS92a, LY90a, LR90, LR91, MSF^{+}95, MD90, MM91b, Nat92b]. numerical [QS90, Ram92a, Ram92b, RS90, Ros90, Row91, SKK90, SKK91, Shu92a, SCMU92a, SCMU92b, ST92a, SVG92, Sul91, SB91, SRSB93, TMR92a, TMR92b, TSS92, Ver91, WDH^{+}92a, WDH^{+}92b, YSG91]. Numerically [FA90, FA91, HK92, HK93a, ZV97, AF90, AF91]. NUT [Val90]. Nyström [COS^{+}98].

O [AAP97]. Object [HM99]. Objects [ART95a, ART95b]. Observations [SWS96]. Obstacles [Pai97, PB98]. Obtaining [BSPL96, OP97, PDHS94, WH97, HG91, HG92]. Occurring [vD97b]. Ocean [BL93, Boy96b, Boy97a, Bry97, FCC97, Hal97, HB96, HdS97, Hig99, LBB94, LIH97, MA98, Mur96, Sem97, SMW99, SH94, TG96, Waj93, BL92, HWY90, HWY91, LG94]. Oceanic [DB98]. Octopus [Kos92]. ODE [Ano94s, Gou96a]. ODEs [BDR91]. Off [oJ95]. Off-Median [oJ95]. Old [FJGP94]. One [BYMZY95, CS93a, FK97, HKM96, Kle95, LMB94, Lu99, MM96, PLD97, Pri93, Pry94, RS94a, Rig94, RFH93, Sim93, Tei99, ARS91, AF90, AF91, CS92a, CW92a, CW92b, CYD92, Dar92, DGD90, FVZ90, HM95, HS90, Hoo90, Lav90, LG90, LG91, MB92, MRS94, RT92, Rus90, SMR91, YR98, ZB92a, ZB92b, DF98]. One- [CS93a, LMB94, CS92a]. One-Band [FK97]. One-Dimensional [BYMZY95, MM96, Pri93, Pry94, RS94a, Rig94, Sim93, Tei99, Kle95, ARS91, AF90, AF91, CW92a, CW92b, CYD92, DGD90, FVZ90, Hoo90, Lav90, LG90, LG91, MB92, Rus90, YR98, DF98]. One-Way [Lu99]. Onset [LSB96]. onto [Boy92b, Boy92c]. OOMPAA [HM99]. OOMPAA-Object-Oriented [HM99]. Open [BC99, CL98b, Fer90a, Fer90b, Gri94, GZ95, Joh93, KP98, PV93, TAC98, LG94]. Operating [WS96]. Operation [GH97]. Operator [Gag98, Gre94, HB96, HKR99, KNW99, PA95a, SS95, IGABB90, IABBG91]. Operator-Split [KNW99]. operator-splitting [IGABB90, IABBG91]. Operators [EL96c, LPBS96, MR99, MS93, MRS98, Ram97a, Rav97, SPC93, SC97, YS98a, AP91a, BRL91, CM91, ESM98, LJT92a, LJT92b, OGBSK90, Ren92a, Ren92b, SPC92]. Optical [BCM99b, HDL99, HKW97, KST90, KST91]. Optimal [AT99, BD97a, BB97a, ES98, FD93, GB97, GN96, HR96, IS96, Maz97, MR99, Rie94a, Wat92b, Wat92a, Bal95]. Optimal-Bias [AT99]. Optimisation [LG94]. Optimization [BCM99b, DM93, Due93, LWT99, MSV98, Pri90, Pri91, Shu95, BS91b, BS92a, BS92b, DS90, FS91, FS92, Mit92, ND90, SSF92a, SSF92b]. optimization-based [FS91, FS92]. Optimized [CK93b, DB96, GS97a, GKDT96, HJ98, Hof95, NC95, RN94, Ods90a, Ods90b]. option [SG90]. Orbit [CW93, JSD95, RMSB96, Smo98]. Orbit-Averaged [CW93]. Orbits [LLT94, Vra95, AF90, AF91, PB91, VBB90]. Order [AD97, AR96, BS99b, BS99c, BR97a, BR97b, Ben95, Boy94, Boy96a, Boy97b, BIAV98, CS93a, COS^{+}98, CGA93, CGA94, CO95, CA93, CF99, DW95a, DW96b, DW97, DM97b, ES94, EVP97, Ena93, FJ98, FM97, GO96, GO98, GN96, GKZ97b, GO95, HEOC97, Hen94, Hir97, Hof97, HL98, Hwa96, IR98, Jam99, JcW99, KPP99, LO96, LNR99, Li97, LL93a, LO98, Mah98, MR99, MP96, MS99b, MLVM98, MB94, Nor96, NC99, Ohw98, OGWW98, Pen95, PZC98, PAB^{+}97, Rav97, RWTC95, Rig94, RF95, SO95, Sco96, SNU98, Sha99, STG^{+}96, Shu97, ST97, Sjö95a, Tab96, TH96a, TS99, Tau94, TP97, Tol94, TL98, YG94, Yav97, YSD99, Zho98, Zhu95, ARS91, BM92, BY90, BY91, CS92a, CRS90, DD92b, For92a, For92b, Fri90, GL96, Hal90a, Hal90b, JW90, KIO91, LO91a]. order [LO91b, Nor91a, Nor91b, RS90, TEKC92, ZC91, ZC92, dFSS92]. ordinary [BB91b, BBF92a, BBF92b, KL97]. Ordinate [Abe97, RNSV91]. Ordinates [GACN93, Sie99]. Oriented [HM99]. Origin [AHH94, DDD98, nJWP96, KRGv93]. Ornstein [LPMS94]. Orr [AC95]. Orthogonal [ABBM96, CS93b, DP91, DP92b, DP92c, Eça96, Hew97, JB95, Kan92, KL92a, KR95, Mur96, OK94, AC92, Chi91, Chi92a, Chi92b, HWY90, HWY91, Mad95, TA91a]. Orthopositronium [JZQ^{+}95]. Orthorhombic [YR93]. Oscillating [HGFH95, BB91b, BBF92a, BBF92b, Fau90]. Oscillations [Abg96, AR97a, PBBS95, PS93a, VDJ93, PBBS91, SCMU92a, SCMU92b]. Oscillatory [EL96c, Fri98, Abg94, DW95b, ES94, FAMO99, HEOC97, HS99b, nJ93, LOC94, NT90, PL92a, Shu97, SK93]. Osher [Shu97]. other [CS92b, CS92c]. Outer [PBD94b]. Outflow [Nor95]. Outlet [JB93]. Overheating [FMM99]. Overlapped [Wan95]. Overlapping [Fis97, Hen94, CH90a, MF92]. Overlays [Gra99]. overposed [Dar90c, Dar91b]. oxygen [MB92].

P [Ske98, Sim99]. P-stable [Sim99]. Package [DÜ92a, LF96, DÜ92b]. Packing [Zin94a]. Packings [ACLW97]. Padé [Cra94, Kob99]. pages [BS99c, BGH^{+}99, lX97, vL97b]. Pair [Bru93, Bru92]. pairs [BF91]. Panel [DZ96, BR92c, BR92d]. Paper [Boy97a]. Papers [Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93g, Ano93h, Ano93i, Ano93j, Ano93k, Ano93l, Ano94a, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano94i, Ano94k, Ano94l, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano95i, Ano95j, Ano95k, Ano95m, Ano95n, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, Ano96a, Ano96i, Ano96j, Ano96k, Ano96l, Ano96m, Ano96n, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Ano97j, Ano97k, Ano97l, Ano97m, Ano97n, Ano97o, Ano97p, Ano97q, Ano97r, Ano98a, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano98h, Ano98i, Ano98j, Ano98k, Ano98l, Ano98m, Ano98n, Ano98o, Ano98p, Ano98q]. Papers [Ano99a, Ano99b, Ano99c, Ano99d, Ano99e, Ano99r, Ano99f, Ano99g, Ano99h, Ano99i, Ano99j, Ano99k, Ano99l, Ano99m, Ano99n, Ano99o, Ano99p, Ano99q, Ano90c, Ano90d, Ano90e, Ano90a, Ano90b, Ano90f, Ano91a, Ano91d, Ano91b, Ano91e, Ano91f, Ano91c, Ano92a, Ano92c, Ano92b, Ano93f, Ano94j, Ano95l]. Parabolic [AE98, DW94b, HMOS95, MM96, MWW96, Str95, YMAC99, ISW92a, ISW92b, MF92]. Paradigm [CWC99]. Parallel [AL97, APS97, BGB99, BS96, BO98, BL98, CN95, CDR99, CKE99, CR96, DB98, DF96, DB96, ESH93, EH93, FP91, GTD98, GM96, GM98, KSB^{+}99, LJG96, LF96, LPR96, MML98, MCPR95, MLBW97, Nic98, ÖL96, Pli95, SD93a, ST95, SDG94, Swe94, Vu98, Pel91, SBGM92]. Parallelization [KNR99]. Parallelizing [ESH93]. Parameter [Bai97, LWT99, Roe97a, SAS94, Ram92a, Ram92b]. Parameterized [CW92a, CW92b]. Parameters [DGN97, GNH96, MSV98, YWS96]. Parametric [CAA93, Hag94, JBA96, Vu96, Vu98, VBD99, WEMH99, DMZ91]. Parasitic [AHH94]. Paraxial [Col97]. Part [Lil97, HL92a, HL92b, Ara97, ESH93, EH93, Fey98a, Fey98b, FP92, HLD96]. Partial [BK97a, CD98, HRR94, KS94, MQS97, OD99, QW93, RIP96, Tei99, VPS95, VP96, WD97, CH90a, FVZ90, HGH90, HGH91, HLD92b, HLD92a, ISW92a, ISW92b, MS90b, MS91c, Wat92b, Wat92a]. Partially [DL93, KNR99, AN90]. participating [Dra90]. Particle [AB97, ACT95, AE95, ADS97, Azm99, Bal94a, BMGM95, BH98, Boa97, Bör97, CSN93, CM99b, CW93, Cot96, DH96, DGD90, DL93, DL94b, EA97, Eyr95, FBM93, Fog92b, Fog92a, FD93, FQ96, GH95, GH97, GR97, GS97b, HMD93, Her93, Hum96b, JLSM97, KP97a, LLR94, LSRS99, LDB98, LPR96, Man93b, MRC93, Mau99, MH93, MMW96, MJPC99, OBL93, PL93, Pop95, RH99, RW96a, RW96b, RSW98, RHT96, Ruy93, Smo98, SHA95b, TS98, TM93, UL98, VAVB93, VB95, Vu96, VBD99, WTH^{+}96, WP97, Wes94, WVHS98, WL93, AN90, Ano91n, Bal95, Bra90, Bra91, BSB92, Cot90, DMZ91, EH93, FPRB90, Fri90, FPRB91, HCC91, KMB^{+}92b, KMB^{+}92a, KDL^{+}92, LG90, LG91, Löh90, LA90, LRJ^{+}99, LL92a, LL92b, MK90, OsKB^{+}92, Poi91, PC92b, PC92a, Rus90, VAVB92]. particle [VG92b, VG92a, Zan92a, Zan92b]. particle-electron [LL92a, LL92b]. Particle-Finite [Her93]. Particle-Grid [OBL93, Cot90]. Particle-in-Cell [BH98, GH95, GH97, JLSM97, OBL93, RHT96, Vu96, VBD99, Wes94, Bra90, Bra91, BSB92, FPRB90, FPRB91, MK90, PC92b, PC92a]. Particle-Mesh [AE95]. Particle-method [Fog92b, Fog92a]. Particles [BDC97, CLD^{+}96, DS97b, GSB^{+}93, KC93, LB94, NY98, RSW98, SC96b, ZP95, CYD92, CPP93, KHM92a, KHM92b, Kos92, KC92, LDB96, Poi91, RH91, TMR92a, TMR92b]. Particular [BR99]. Particulate [CM99a, Poz99]. Partition [AFH94]. parts [Ano95u, Str94]. Passive [PL98, WJC93]. Past [HC93, MB96, NC93, ZP95, IT90]. patch [DZ91]. Patches [BPM98, But90a].

Path [DH96, HBF93]. paths [BG92, Bas92, MMS90a]. Pattern [Aoy95, Dar95, RMO99, Alm93]. PDE [BB97a, PMO^{+}99, YMAC99]. PDE-Based [PMO^{+}99]. PDEs [Ano99-30, BKV98, CHR99, FS97, Hol96a, LP97, PZ98, VP97, YG94]. PDF [Bri95, MJPC99, WP97, Wel98, XP99]. PDF/Monte [XP99]. PDF/SPH [Wel98]. peak [HS90]. Pearls [LPR96]. Pebble [JH97]. Penalty [TK99]. pencils [Nat92a]. Penetrating [Bör97]. Penetrative [TC98]. Peptides [AT99]. Percolation [JH97]. Perfect [FMO97]. Perfectly [AGH99, ARTAAA97, Ber96, Col97, Hes98, Hu96, PSL99, Pet98, QG98, TAC98, Ber94]. Performance [BBF^{+}99, BL92, BL93, BS96, Bro95, Fig92, MO95, MB97, NP92, Tay94, TL98, WH98]. Periodic [APV98, BGP98, CRRR97, FH94, FK97, Ji94, MH93, Osb93, ST97, Vra95, Wir97, ZLP97, Zin94a, BM94, Bor91b, Bor92a, Bor92b, Boy91b, Boy92a, CPP93, HM95, LDB96, MBV93, Poy92, SMR91, VBB90]. Periodicity [CDF95]. Permanent [BS99d]. Perspective [Nac96]. Perturbation [Boy97b, CDT98, EVP97, Hof97, LB93, Rie94b, ZV97]. Perturbational [CGY93, CGY92]. Perturbative [Boy95b, SMM^{+}90]. Perturbed [CMW95]. petroleum [BR92a, BR92b]. Phase [BDS^{+}99, BBMB97, CE97, CM99a, CEG^{+}97, CDT98, CJ95b, DA94, Dur93, FV97, HK93a, HRL99, Jac99, KS94, LL96, MO95, PMR97, Sai95, SSO94, SAB^{+}99, TP97, TM93, UMS99, WS96, Wel95, ZHL96, CS91, CYD92, HK92, Kos92, LP90, RS90]. Phase-Field [BDS^{+}99, FV97, Jac99]. Phase-Fluid [Sai95]. Phase-Shift [CJ95b]. Phase-Transition [HRL99]. Phases [MP96]. Phenomena [GS97a, GH95, RS94a, Tan93, Ano91n, KMB^{+}92b, KMB^{+}92a]. Photoconductor [dM97a]. Photographs [Hel95]. Photonic [AK99, Dob99, FG97]. Phys [CNG17, MKM04]. Physical [LLR94, PC96, BKP91, MW93, STAS91]. Physics [BF97, KP97a, Lan97, PM99, YMB^{+}91]. PIC [FLD93, JLM^{+}96, MLJ97, OS97, Par90, PSL99, SAB95]. Piecewise [DW94b, Hol96b, MM96, Sor95, Che91, Yor90]. Piecewise-Quintic [Hol96b]. Pipe [Kar97, PM98, Nic93]. Pipes [hYsY96]. Planar [Asl99, Hob93, Hum96a, JB95, LL93a, UL98, CS91, Hob92]. Planar-Geometry [UL98]. Planck [BCDL97, BC98, Epp94, GGZ93, Mik90, Mik91, PC92b, PC92a, SVG92, SVG93]. Plane [CP98a, CS97a, Dar95, EM95, oJ95, Oos97, Sie99, YP98, BM91a, BM91b, GKM96]. Plane-Strain [CS97a]. Plane-Wave [CP98a, YP98]. Plasma [BGB99, BF97, BO96, BL94, CSN93, CD93, CGKC93, CDW99a, CDW99b, CLD^{+}96, FLD93, GSB^{+}93, GRB^{+}95, GH95, HK94, JLSM97, JSD95, KL93, KPC93, KM95, Kno98, KLB99, LB94, Mat94, MSF94, Nun93, PFRB93, PPBC93, Ram97b, RHT96, Sco96, SCG94, STS^{+}97, Swi96, VD97a, VM96a, VAVB93, VS95, WCO94, Zan97, Ano91n, CGKC92, FPRB90, FPRB91, KMB^{+}92b, KMB^{+}92a, KDL^{+}92, KPC92, KST90, KST91, LG90, LG91, LL92a, LL92b, MS91b, MS92b, MS92a, PPBC92, RD91, RLM90, TT90, TT91, TA91b, TAD92a, TAD92b, VAVB92, VNC92a, VNC92b].

Plasma-Neutral [BO96, VNC92a, VNC92b]. Plasmas [AHH94, DHS94, GGZ93, KGH^{+}98, KRGv93, KC93, LMS98, LW94, MK97, PS93b, Tan93, TS98, WJC93, AN90, HCJ92a, HCJ92b, KC92, OsKB^{+}92, PC92b, PC92a, RD92a, RD92b, Rob90b, Zan92a, Zan92b]. plastic [TP92a, TP92b]. plate [HPS92a, HPS92b]. Platelet [WF99]. PLIM [RS94a]. Plume [CR96]. plus [GS94]. PML [AG97, YP98]. Poincaré [BLL99]. Point [AC95, ART95a, ART95b, CMW95, Chr97, CF98, CF99, Eça96, FWWD94, Pen95, SP99, Zab97, Fig92, Jac90, MPG92a, MPG92b, PRMV90, QR90]. Point-Matching [ART95a, ART95b]. Pointed [Dar96]. Points [GS96a, Mas96, Vad97, Zan99, BG92, Bas92, VBB90]. Poiseuille [FJGP94]. Poisson [Tho97, AKW93, BK93, BR99, Bar91b, BIAV98, CMK90, CZS98, CKSB97, DVD93, DR90, Fab92, GL96, GKZ97b, Her93, JC98, LJT92a, LJT92b, MGM95, MBI^{+}97, RCAM92a, RCAM92b, RLM90, She95, SFB91, Wan99, Win97, Zha98]. Poisson-type [Bar91b]. Pol [CAA93]. Polar [MM95b, SNU98, EHW90, EHW91]. Polarized [KKOF97]. Pole [HS93b]. polyatomic [tVB92a, tVB92b]. Polygons [MA93b, GSY92, Got92]. Polyhedra [Gra99, MA93b, GSY92, Got92]. Polyhedral [LV98]. Polymer [KWT^{+}97, SVS90, SVS91]. Polymeric [SHF97]. Polymerization [GTD98]. Polymers [För96, MDA91]. Polynomial [Boy95a, Eve96, RH98, SRVK96, Dic90, TEKC92, Tas92, Yor90]. Polynomials [CS93b, JB95, LV98, Sen99, BE91, BE92a, BE92b, Chi91, Chi92a, Chi92b]. Porous [CE97, Dur93, GP98a, HT97, HW97, LLK98, Pap93, SW96, Bor91b, Bor92a, Bor92b, Pap92, RS90]. Poroviscoelastic [CH99]. Portable [DHSS94]. Posed [BST95, AGH99, HM95]. posedness [Glo92]. Positive [HKvV95, SG90]. Positivity [BS99a, FWBS94, GVM99, Lar91]. possessing [MSF^{+}95]. Post [MQS97, SCA92b]. post-interface [SCA92b]. Post-processing [MQS97]. Postshock [AR97a]. Potential [DR96, EL97a, GS94, GM98, IH95, Jai97, KS94, Tan94, VD97a, VAVB93, ZZ98, dAT95, JB91, JBvK^{+}91, KKD90, MGP91, May92b, May92a, Str91, Str92a, Str92b, VAVB92]. Potentials [BD94b, BK97b, EH93, FV97, FH94, Gou96a, GBS99, NC95, RN94, SHF97, Wit96a, Wit96b, Kob90, Kob91, KF91, LDB96, Str91, Str92a, Str92b, Tas92]. Power [Cra94, HKM96, VC90, VC91]. power-law [VC90, VC91]. pp [Ano99-30]. Practical [GA95]. pre [SCA92a]. pre-bubble [SCA92a]. Precision [EB94]. Preconditioned [ANL94, HB90a, HB91, JE90, LZS98, PG97, VSH96, VSM99, KH90, KH92]. Preconditioner [EVP97, CP90]. Preconditioners [Azm99, DS96]. Preconditioning [BL98, CN95, CM93, DS99, Lee98a, Lee98b, PBD94a, Tid97, DD90, DD91, Mor90]. Prediction [Can95, Ush96]. Preface [Roe97b, Shu97]. Premixed [QTL98]. Preprocessor [Tót97]. Prescribed [MM98]. Presence [BK99, KS92]. Preservation [CW98]. preserve [Lar91]. Preserving [Boy98, FWBS94, Mad95, MS99a, SH97, TW93]. Pressure [Abg96, BS99a, BS96, Dor99, Edw96b, EOS98, LM95, MSF94, PR93a, SI94, SC97, Taf95, TZA96, WS93, BK91, BK92a, BK92b, RWC90]. Pressure-Based [BS96, SC97, TZA96, WS93]. Pressures [CS98a]. Prevent [Abg96]. Primitive [Kar94, LL96, UWH99, WH98, XS93, HWY90, HWY91, SA90b, SA91, SA92a, SA92b]. Principal [NM95, Lay90]. Principle [dAT95, Adh92a, Adh92b, BWJ90, MWJ91, Sal91, Ver91, dG92a, dG92b]. Principles [Osc94, SWS96]. Priori [MKM98]. Probabilities [WH97]. Probes [Hum96a]. Probing [HM99]. Problem [BS99d, BD97b, Boy95a, CS97b, Cra94, Dar95, FMM99, Gar98, GPWZ96, GM95b, GDP96, Had99, KS94, KT99, LWT99, LL96, Mal96a, Mar94a, OD99, SA95, SVV98, SMT99, Shu95, SDI98, SRA98, TWV97, TO98, Wan97, YZ98, ZCH96, Abr91a, Abr91b, BW90, BW91, FP91, FS91, FS92, FP92, Glo92, Hei92, KEHK90, KEHK91, LGC92, Mel91, MB92, PDC91, SKK90, SKK91, SH92]. Problems [AS96, AK98, BR99, BYMZY95, Bar94, BCM95a, Bel97, BD97a, BS99e, BMZ95, BBMB97, CLO93, CW98, CMW95, CC97, CMOS97, Cho97, CJM97, DR96, DA94, DS94, ES98, EL96c, FA97, FD93, FPQ93, GS95, GP98b, HOS93, HKNT98, HML99, HR96, HLOZ97, HW97, HHR99, HS93b, HS94, HSS97, IH95, Ji94, Ji96, nJ93, JT93, KP97a, KL94, LL99, LT93, LEI96, Liu96a, LO93, MW94, MB94, NO93, OBB98, Oos97, OGWW98, PFRB93, PK95, Pen95, Pri98, Puc97, RS95, Ram97a, RFL93, Rig94, SSR96, SS95, Shy98, ST97, SW96, TKT97, TS99, TH96b, TO96, VSH96, VSM99, Wir97, ZC94, ZW97, ZOI97, ZRR99, vGVO98, AM90, BDR91, CS92b, CS92c, Dar90c, Dar91b, DM91a, DM92a, DM92b, Dra90, Fer90a, Fer90b, For90, For92c, FMS92, HM95, Jac90]. problems [JLB90, Kar91, Kar92, KW92, Lay90, LHI90, LM90c, LE90, LE91, Mor90, Nat92a, Ods90a, Ods90b, OB95, Or94, QR90, Ram92a, Ram92b, SSF92a, SSF92b, SMM^{+}90, SK92, SRSB93, TLES93, UWSB90, UWSB91, WMG^{+}91a, WMG^{+}91b, dFSS92]. Procedure [AT99, AK94b, BTW96, HOS96, LT93, LIH97, Liu96a, MBV93, MG97b, Nea94, SvD96, Sim93, TYS95, AM90, BS91a, For92c, FMS92, MDB91a, MDB91b, SA90a, Wei92]. procedures [CH90b, CH91, HHG91, HHG92, KL92b]. Process [LL95, RLT93, SKO94, dG96, SCA92a, SCA92b]. Processes [DFFG96, LNR99, SL96, SCT^{+}99, SCC93b, BBC^{+}90, BBC^{+}91]. Processing [Shi93, MQS97, Pri90, Pri91, TZWH92b, TZWH92a]. processor [BBK91, Bru92, Bru93, JJK92, SBGM92]. Processors [ÖL96]. Produced [Gou96b]. Product [Gra95, Man93a, PR93b, Wit96a, Vre91a, Vre91b]. product-formula [Vre91a, Vre91b]. Program [LW96b]. Programming [ST95, Lav90]. Progress [HLD96, RS95]. Projection [ABC^{+}98, BCM99a, CP99, CR99, Eyr94, FMO98, GQ97, KPP99, LS98, Min96b, Min96a, PAB^{+}97, Tau94, XH98, ZS92a, ZS92b, Zhu95, BM92]. Projections [Pri94]. Projective [WL96]. Propagating [AS95c, ZHL96]. Propagation [AHH94, BD93, FHKZ97, HT95, HHR99, Iva93, KI99, LS96, LeV97, LeV98, Löt94, MMW96, Qi93, Ram97a, Ske98, WJTP96, Wri98, ZS97, AAP^{+}95, LRB^{+}90, LBC^{+}91, TPL90, WMG^{+}91a, WMG^{+}91b]. propagator [BG92, Bas92, CK90a]. propagators [TEKC92]. Proper [LR98].

Properties [BY94, CW98, För96, Hel95, JZQ^{+}95, MS93, NMW^{+}96, Rob90b, ZDG95]. Property [EG95, Fur99]. Proposed [TRL99]. Protease [DOOSM99]. Protection [LTE93]. Protein [JMDB99, SS99a]. Proteins [DOOSM99]. PS2M [AN90]. Pseudo [Ano94s, BCM95a, BGM95, LD93, Mar93, PM90, RG97, SSR96, Sun96, BKM90, BT92a, BT92b, Dom90, ND90]. Pseudo-Current [Mar93, ND90]. Pseudo-Non-Time-Splitting [Sun96]. Pseudo-Solid [SSR96]. Pseudo-Spectral [BCM95a, BGM95, RG97, LD93, PM90, BKM90, BT92a, BT92b, Dom90]. Pseudo-Steady-State [Ano94s]. Pseudocompressibility [SC97]. pseudoparticle [Mak99]. pseudoperiodic [RS91]. Pseudorandom [MCPR95, ASTAS91]. Pseudospectral [Boy92d, Boy94, Boy96b, Boy97a, Boy97b, CBL95, DECB98, Don94, DF98, GS96a, Han93a, HS93b, HS94, Kab96, KTE93, Ku95, Mal96a, MR99, MQS97, NPC93, PBD94a, Pri95, RF96, TC98, YP98, For90, Jac90, MS90b, MS91c, Pel91]. Pseudospectral/Delves [Boy97b]. Pulp [SG98]. pulsar [LS92c, LS92d]. Pulse [FHKZ97, KKOF97, KL94]. Pulses [Luk99]. Pumped [LW94]. Pure [FA97, PA95a]. Purpose [BGGT90, DS90]. pursued [YHW90, YHW91]. Pushing [Smo98].

Quadrature [CK96, CP98a, CZS98, CJ95b, LV98, NM95, NB98, Pri94, Str96, Yak96]. Quadrilateral [ABBM96, Gir98, MRS98, ST99]. Quality [MCPR95, Sun96]. Quantization [KLM95, LB93]. Quantum [Abr91a, Abr91b, Bar94, CCGJ95, Cra94, Dru99, KKOF97, LW96b, LLLY99, LK94a, LK94b, MH94, Rei99, Wit96a, SMM^{+}90, SFB91]. Quantum-Classical [Rei99]. Quantum-Mechanical [Bar94]. Quantum/Classical [Dru99]. Quasi [Abg96, Ano94t, CC98, KO98, LBB94, LeV98, MC95, OG97, PCLC97, Pri93, SS99c, SMW99, TO98, Wir97, Dar90a, Dar91a, MPG92a, MPG92b, ZDG95]. Quasi- [KO98, SS99c]. quasi-confirmal [Dar90a]. quasi-conformal [Dar91a]. Quasi-ENO [OG97]. quasi-fractional [MPG92a, MPG92b]. Quasi-Geostrophic [SMW99, LBB94]. Quasi-Isometric [CC98]. Quasi-Monte [MC95]. Quasi-Newton [Ano94t, PCLC97]. Quasi-Periodic [Wir97]. Quasi-Riemannian [Pri93]. quasi-static [ZDG95]. Quasi-Steady [LeV98]. Quasi-Two-Dimensional [TO98]. quasiconformal [Dar93]. Quasilinear [Tho97, Win97]. Quasineutral [DHS94, JLSM97, Ram95]. Quasiparticles [Rom97]. Quench [Bot96]. Question [Sco98]. quick [HHG91, HHG92]. Quintic [Hol96b, CHW90]. Quotient [Fat99].

R [AAP^{+}95]. R-matrix [AAP^{+}95]. RACAH [RMSB96]. Radau [BN94]. Radial [GEK^{+}97, CRS90, HSN90]. Radially [Nor98]. Radiation [BBF^{+}99, Boy95b, Boy97b, DW98b, Gra95, MSF94, RKO99, RK99, SCA92a, SCA92b, TW94, TG96, Win95, dG96, Dra90, YM91, YM92b, YM92a].

Radiation-Hydrodynamics [BBF^{+}99]. Radiation-induced [SCA92a, SCA92b]. radiation-magnetohydrodynamics [YM92b]. Radiative [AN98, Fra95, JFH^{+}98, KI99, MWS96, BKP91, Sal91, SM92, Sto92, dG92a, dG92b]. Raman [BGB99, GB90, GBS^{+}90, GBL^{+}92, GSB^{+}93]. Random [BBGL93, CV92a, CV92b, EM95, EHM97, HB95, KI99, Zan99, Zin94a, BES90, CG97c, HN95, HCP93, VC90, VC91]. Randomly [MD98, SV97]. Range [Lu99, NC95, Pli95, RS99b]. Ranged [RN94]. Rankine [JM98]. Raphson [SSR96]. Rapid [GM98, KL94, LH92, Rok90, SH99, HB90a, HB91, May92b, May92a]. Rapidly [KB99, BB91b, BBF92a, BBF92b]. Rarefied [APS97, FBM93, PC99a, YH95]. Rate [Boy94, Rom97, KH90, KH92]. Ratio [Kal97, SP98, VRD99]. Rational [LLT94]. Ray [ACT95, Ben96, FHKZ97, Rie99b]. Ray-FDTD [FHKZ97]. Rayleigh [Fat99, Gel99, HCZ99, KG92a, KG92b, LJG96]. Rayleigh-Bénard [Gel99]. Re [MH94]. Re-formulation [MH94]. Reacting [BS99e, Edw96a, GSB98, KNW99, LLD93, NWK98, Ton96, XP99, CYD92, IGABB90, IABBG91, LLD92, YTS92]. Reaction [CZ98, HO96, HGFH95, KD97, SL96, WEQ^{+}99, WH97, BB90b, BB91c, CYD92, MMS90b, MMS91]. Reaction-Convection-Diffusion [HO96]. Reaction-Diffusion [CZ98, HGFH95, KD97, SL96, MMS91]. Reactions [RC95, WH97]. Reactive [BPT95, BS99f, CR96, MJPC99, PDHS94, SMJ98, Zho96, BAB90]. Reactor [EGS96]. Real [CP98b, LL93b, MML98, MS99b, MVZ97, SLL94, ZC94, LVS90, Shu90, SLV90, VM90]. Real-Time [MML98]. Realistic [BD94b, Pri94]. Realization [SCT^{+}99]. Receptor [EA97]. Recirculating [LT93, BY90, BY91]. Recombining [Kno98]. Reconnection [AS94]. Reconstruct [TSW95]. Reconstructing [GBS99, RK98]. Reconstruction [CL98a, Mot98, OG97, Osc94, SSW98, YWS96, AP90, AP91b]. Record [Due93]. Record-to-Record [Due93]. recoupling [WS91, WS92a, WS92b]. Recovery [SR99]. Rectangles [Mar94a]. Rectangular [BS99e, Gel99, ZS97, Mar90, YMB^{+}91]. Recurrence [CS93b]. recursive [JVR93]. Rederived [MJP93]. Redistancing [Str99a]. Reduce [MM97b, DÜ92a, DÜ92b]. Reduced [Gou96b, IR98, VBD99]. Reduced-Description [VBD99]. Reduced-Order [IR98]. Reduction [FBM93, RSW98, SE98]. Reference [TRL99]. Refined [CDW99b, DP93, Min96b, PV99, DP92a, YMB^{+}91]. Refinement [BS99e, Coo99, FH98a, FGM97, GBCA99, HT97, JFH^{+}98, Kho98, PGD99, vdVvdV98, FM92, Fra92]. Reflected [LM95]. Reflecting [Fer95, Giv90, Giv91]. Reflection [Ben95, HKW97, TPL90]. Reflectionless [PZC98]. Reflections [DM96]. Reformulated [Hei93]. Refraction [HKW97]. Refrigerator [WK96, WKK98]. Region [Gra99, MA98, PBD94b, BW90, BW91]. Regions [DE93, GM98, JMR94, PBC^{+}95, Hor90, Hor91, Mar90, May92b, May92a]. Regridding [Löh96]. Regular [CC98, KCJ95, ZL93]. regularised [GG90]. Regularization [HOS96, PMR97, WKMH97, LKLE90, LLEK91].

Regularized [BLP98, She90, She91]. Regularizing [SBO93]. Rejection [KNR99]. Rejection-Free [KNR99]. Related [BD97a, EL96b]. Relation [Boy97b, TW93]. Relations [Gou96b]. Relativistic [Bal94b, Cra94, DW95b, DFaIM98, Man93b, MM96, YCTC97, BKP91, HTK90, SKR^{+}93]. Relaxation [GGP97, HMD93, Jin95, JL96a, KKH^{+}99, Kla99, LPBS96, LL95, Maz94, MS99b, PCLC97, ZB96, BG92, Bas92, GC90, Tay91, TfNW91, TNW92a, TNW92b, TNW92c, TNW92d]. Relevant [CGKC93, CGKC92]. Reliable [Pry94]. Remapping [Gra99]. Remeshing [SWS96, DT90a, DT90b, HW92a, HW92b, PPM92a, PPM92b]. Reminiscences [God99, Van99]. Removal [Cho93b, Jak93, Cho90a, Cho90b]. Removes [Rob99]. Removing [HLS94, HKS98, Kel94]. Representation [JBA96, MLJ97, VC90, VC91]. Representations [SH98a, PH94]. Representer [Lya99]. Representing [HK93b]. Reproducible [MCPR95]. Reproducing [HH98]. Republication [Shu97]. repulsion [LKB90]. Requirements [Rie99a]. reservoir [BR92a, BR92b, JE90]. Reservoirs [DGN97]. Residual [Ena93, SC97, TKT97, Dav91, Dav92a, Dav92b]. Residuals [LR94]. Resistance [SDI98]. Resistive [KGH^{+}98, PBD94b, SS96b, CHL^{+}90, HCJ92a, HCJ92b]. Resolution [AS96, AR97b, CWC99, CS97b, CEG^{+}97, Dau92a, Dau92b, DDS97, ES94, Eka99, FCC97, GO96, GO98, Har97, JCA97, Lax97, LS96, LeV98, Mah98, PPBC93, SWT98, SRBG99, TRL99, Yee97, DM92c, Har92a, Har92b, HL92a, Lel92, Mul92b, Mul92a, PPBC92, SW92, WR91, YKM90, DQ95]. Resolved [Kal97, Bro95, LKLE90, LLEK91, MB97]. Resolving [LIH97]. Resonance [SS99b, SMSS98]. Resonances [Pry94]. Resonant [CCGJ95, HK98]. Response [HH98, LPC^{+}93, SA95]. Responses [KL94]. Restart [WCSW99]. Restriction [KTE93]. Resulting [GH97]. Results [FJGP94, LLR94, ZDD99, Nat90, Nat91, Rus90]. Reversible [EA97]. Review [HGH90, HGH91, Giv90, TT90, TZWH92b, TZWH92a]. Revised [AN90]. Revisited [Dar96, Maz97, HN95]. Revolution [ARTAAA97]. Reweighted [nJ93]. Reynolds [BY92, Bra92, CPJ90, IT90, KL92b, MFZ97, RK93]. Reynolds-averaged [CPJ90]. rezoned [Hal90a, Hal90b]. Rezoning [Nor91a, Nor96, Nor91b]. RF [SG94]. Rheology [GS97b]. Ridges [RB95]. Riemann [Bai97, Bal94b, BHJU99, CS97b, DW94a, DW95a, FP92, Gla91, Gla95, JM98, LPvL93, MIM90, Mon97b, Roe97a, Sai95, SLL94, TWV97, Tou92, TK96]. Riemannian [Dvi90, Dvi91, Pri93]. Rigid [AKW93, BMJ^{+}97, Xia99, Sal91, dG92a, dG92b]. Rigidity [JH97]. Ring [OS97]. Rings [LSB96, MKM99, MKM04, OsKB^{+}92]. rings-and-spokes [OsKB^{+}92]. rise [MD90]. Rising [Dar95]. Rivulet [RKJ97]. Robbins [PH95b]. Robert [Ano99-30]. Robust [DS99, DB97, ETRW95, KJ98, Löh95, MP97, SDP97, Wal94, WD97, YM95, Knu92]. Robustness [Mav95]. Rocks [PM99]. Rods [Kla96]. Roe [CG97a, Gla95, MVZ97, PV99, Sai95, Tou92, VM90]. Roe-Type [Sai95].

Rokhlin [HM95, SB98]. Roll [CS97a, CS99, SCR97]. rollup [KG90b, KG91, KG92a, KG92b]. Root [Loe93]. roots [Gre91, Gre92a, Gre92b]. Rossby [Duk95]. Rotated [TN98, LPvL93]. Rotating [BTW96, KB99, PBBS95, SS99c, YSMG97, IT90, MLB92, Zan92a, Zan92b]. Rotational [HI98]. rotationally [FP92]. Rotations [HP98]. Rough [APV98, BES94, PR93b, SS96c]. Roundoff [BCM95b]. Rovibrational [WKHS97]. row [Gla91]. Rules [Str96, Yak96]. Runge [AL97, AA90, AA91, CS98b, Gil97, HHM96, Jin95, MR99, SH98b, SH97, Vad97, Wei92, Zho96]. Runs [Bar90]. Russell [Ano99-30]. Rys [CP98a].

S [Van99]. sampled [RS91]. Sampling [AH99, BS99d]. Sand [RB95]. Satisfying [DÜ92a, DÜ92b]. saving [Par90]. Scalability [BS96, KSB^{+}99]. Scalable [PDHS94]. Scalar [DB96, FEO95, LO96, Li97, GWI92, ZDG95, AK99]. Scale [GKDT96, Hal97, HB96, Kal97, KWT^{+}97, SP96, Swi96, Waj93, vNZ94, FPRB90, FPRB91, PFRB93]. Scaled [SH98a]. Scales [EM94, EM95, Rei99, Sch92b, Sch92a]. Scaling [LLLY99, Val91a, Val91b]. Scatter [BES94]. Scattered [AK98, MD95]. Scatterers [AWY99, Rob99]. Scattering [AL97, ART95a, ART95b, ARTAAA97, AWY99, BPS95, BGB99, BMZ95, BGP98, CL93, CJ95b, Eyr94, GSB^{+}93, GK96, HKNT98, Hof97, KI99, KS94, KL94, NY98, Osb93, PDHS94, Rom97, Sha96, UL98, VSH96, Win95, lX96, lX97, lX98, YGH97, YL98, dAT95, Abr91a, Abr91b, BKP91, BO92, GB90, GBS^{+}90, GBL^{+}92, HB90b, Mel91, Osb90, Osb91, PO90, PO91, Raw90, Raw91, Rok90]. Scharfetter [Kul95]. Scheme [AS96, AAL97, Ano97-29, ADS97, Bau97, BD93, BC98, CAA93, CD93, CA93, CJ94, CGY93, CD95a, CM95, CF98, CF99, CKE99, CCG95, DW96a, DW96b, DW97, DW98c, Epp94, FB94, FAH97, FM97, GO96, Göt99, HCZ99, Hir97, HWM96, Hol96b, HKvV95, Jac95, JcW99, Jin94, JSE97, Kla99, KNW99, Kno98, Leb99, LH95, LJG96, LS93b, MKM98, MWW96, Mei98, MWS96, MRS98, MP97, MVZ97, MA96, NWK98, NGN97, OK94, PK94, PV99, PRL^{+}99, RC95, Sak96, SZ97, SA95, Sco96, SW99a, SWT98, Sjö95b, Smi99, SB96b, SXG99, SCT^{+}99, Tan94, TH96a, TN98, TP97, Vad95, VO96, WTH^{+}96, Weg97, WM94, Xu97, YCTC97, Yav97, dLMSS95, dNPT95, vL97c, vM95, Alv91, Alv92, AA90, AA91, Boy90, Boy91a, BJ90, BW90, BW91, CWG90, CWG91, CGY92, Dem92a, Dem92b, Fis90]. scheme [HHG91, HHG92, Kle95, LH90, LR90, LR91, MDHW92a, MDHW92b, MRS94, Nat92a, Or94, PL92a, YTS92]. Schemes [AD97, Abg94, Ano94s, AR97a, AE95, Bai97, BS98a, BKV98, Bih96, BS99f, BT97a, BT97b, BNW96, CGA93, CGA94, CJS99, CEG^{+}97, DW95a, DM97b, DW95b, DL94b, ES94, EL96a, EL96b, EL97b, Eka99, Fat99, FMO98, FWBS94, Fri98, Fur99, GS97a, GMS^{+}99, GO98, God99, GVM99, Hal97, HHF95, HP93, Har94, Har97, HEOC97, HOS93, HHM96, HS99b, Hub99, Hwa96, JS96, JL96a, JT93, Jun99, KH97, Lax97, LP97, Li97, LOC94, LO98, Mah98, MD95, Mar94a, Mar94b, MS99b, MLVM98, MBP94, MR98, OG97, PC96, PLD97, PZC98, PC99b, RS95, Rig94, Roe97a, SMD98, Sco98, Sha99, Shu97, ST93, SH98b, SW99c, SH97, SRT98, TW93, Tei99, Thu96, TO96, Van99, VC99, VSG95, WK99b, XMP97, XH98, YYCH98, ZLOT98, Zho98, ZL93, BR92a, BR92b, BY90]. schemes [BY91, Car91, Dar92, Fig92, HP92, HT94, JLB90, KL97, Lay90, LRB^{+}90, LBC^{+}91, Lel92, MMS90b, MMS91, RM90, RM91, Rob90a, SG92a, SG92b, ST92a, SNL90, Sur94, ST92b, WR91, Wes92a, Wes92b, Yan90, YP92b, YKM90]. schrödinger [GCMR90, ARS91, BO92, Boy97b, BSPL96, BCR99, BCT98, CRS90, CW90, CJS99, CHW90, Cra94, EVP97, GS94, GEK^{+}97, GKKR99, Gou96a, Jam99, KOS^{+}96, Kou96, LRB^{+}90, LBC^{+}91, NC97, Osb93, PRS90, PM90, Pry94, RF90, RF91, RFH93, SB95, SY97, Sim93, Sim99, TEKC92, Tas92, TSS92]. Schrödinger-like [Cra94]. Schrödinger-Type [SY97]. Schumann [Wer98, KHW98, Wer95]. Schur [CD95b, Loe93]. Schwarz [DB98, Fis97]. Schwinger [Adh92a, Adh92b, GM95b, MP93]. Scour [Ush96]. Screened [BFO99]. Sea [GZ95]. Search [AB97, Löh95, MSV98]. Search-Locate [AB97]. Searching [Boy95a]. seas [LS92a]. Second [BHJU99, DW96b, DW97, GN96, GKZ97b, Hir97, Hof97, HL98, KPP99, MB94, RWTC95, Sco96, SNU98, STG^{+}96, Tab96, TH96a, TS99, Tau94, TP97, Zhu95, BM92, Fri90, Hal90a, Hal90b, JW90]. Second- [GKZ97b]. Second-Moment [BHJU99]. Second-Order [DW96b, DW97, GN96, Hir97, Hof97, HL98, MB94, RWTC95, Sco96, SNU98, STG^{+}96, Tab96, TH96a, TS99, Tau94, TP97, Zhu95, BM92, Fri90, Hal90a, JW90]. Second-Order-Accurate [KPP99]. Section [EOS98, hYsY96]. Sectional [Hel95]. Seeking [Abd94]. segment [AP90, AP91b]. Segmented [SR97]. Seismic [GM95b]. Sekerka [ZCH96]. Selected [FH97, NO93]. Selection [Dar95, RK99]. Selective [LB94]. Self [AK94b, AC93, BCR99, DHS91, LW94, PS93a, Sam97, YDT93, MM91b, YHW90, YHW91]. Self-Adaptive [AC93]. Self-Consistent [LW94, DHS91]. Self-Correcting [AK94b]. Self-Gravitating [YDT93]. self-pursued [YHW90, YHW91]. Self-Similar [BCR99, Sam97, MM91b]. Self-Sustained [PS93a]. Semi [Ano91n, Cas90, CW93, DS99, Ful93, HG94, Ji94, JR96, Kle95, KNW99, Kop96, NWK98, Oli94, SH94, SRBG99, Str99b, VRD99, WH98, Zho96, HWY90, HWY91, LL91, Mul92b, Mul92a]. Semi-coarsening [DS99, Mul92b, Mul92a]. Semi-Implicit [WH98, Zho96, Ano91n, Cas90, CW93, Ful93, Kle95, KNW99, NWK98, SH94, VRD99, LL91]. Semi-infinite [HG94]. Semi-Lagrangian [Oli94, SRBG99, Str99b]. Semi-periodic [Ji94]. semi-spectral [HWY90, HWY91]. Semi-staggered [JR96]. Semi-Structured [Kop96]. semiclassical [BG92, Bas92]. Semiconductor [FO93, Kul95, LWT99, MQS95, DGD90]. Semiconductors [BH95]. Semispectral [WH98]. Sensitivity [BB97a, GNH96, KTN^{+}94, STG^{+}96, DM92c]. Separable [EVP97, Bar94, CR91]. Separation [PH95a]. Sequel [Hir97, Lio96]. sequence [CV92a, CV92b]. Sequences [Eye96]. Series [GBS99, HKM96, Sho93, DD92a, LKLE90, LLEK91, RF90, RF91, VC90, VC91]. Set [AS95c, AS95a, AS95b, AS97, AS99a, ABS96, BS98a, CHMO96, CMOS97, Cho93a, GGP97, HOS96, HRL99, JCHW95, LZG99, MBO94, PMO^{+}99, Str99a, Str99b, SSO94, SAB^{+}99, ZCMO96, ZMOW98, WDH^{+}92a, WDH^{+}92b, ZS92a, ZS92b]. Set-down [GGP97]. Several [BO98, FKM99, Fra95, Wit96b]. Shadow [Ano94t]. Shadowing [GE92]. Shake [Pal93]. Shallow [BD93, CN95, GNHP95, Göt99, HB97, JCHW95, Ma93a, NGN97, Pri93, STS98, SP99, TTI97, VC99, YM95, Cas90, LS92a, Nad95, Ste92b, Ste92a, TPL90, WDH^{+}92a, WDH^{+}92b]. Shallow-Water [BD93, STS98, YM95]. Shape [HP99, Hum96a, IS96, Ruy93, MD90]. Shapes [AD97]. shared [Nat90, Nat91]. shared-memory [Nat90, Nat91]. Sharp [FV97, GO96, GO98, PH95a, UI99, UMS99, CS91, HS90]. Sharp-Interface [FV97]. SHASTA [Zal97]. Shaw [BST95, HLOZ97, Whi90a, Whi90b, Whi94]. shcemes [YP92a]. Shear [CSN93, DE99, GT97, Gri94, RK95, Bue91, GG92b, KG90b, KG91, MNP90, MM91b]. Sheared [SCG94]. Sheath [PFRB93, PPBC93, SG94, PPBC92]. Sheet [Ber95, BLP98, EB94, NB98, PH95a]. Sheets [Bis95, FGM97]. Shelf [GZ95]. Shell [KB99]. Shepard [SRF93]. Shift [CJ95b, HK93a, HK92, KvdVPG95]. shift-and-invert [KvdVPG95]. Ship [SDI98, TS96, TSR92]. Shock [AS96, AR97a, ABS96, BL99, BKP96, CSW98, DW95b, Don94, DQ95, DM96, DFN^{+}99, GMS^{+}99, GO96, GO98, HHF95, HP93, Har94, LH95, Ods93, RWTC95, Shy98, SK93, Ton96, TPRC96, Yee97, YSD99, ZDD99, BM91a, BM91b, Boy91b, Boy92a, HP92, LGC92, kM90, kM91, Ods90a, Ods90b, Poy92, Rob90a, YP92a, YKM90]. Shock-Adaptive [LH95]. Shock-Capturing [BKP96, GMS^{+}99, Shy98, Ton96, Yee97, YSD99, kM90, kM91, YP92a, YKM90]. Shock-Cylinder [DQ95]. Shock-Turbulence [AS96, HHF95]. Shock/Turbulence [DFN^{+}99]. shocked [Lav90]. Shocks [IS96, JL96b, KC97, LM95, RK95, HL92b, VD90]. shocl [YP92b]. shocl-capturing [YP92b]. Shooting [Ji94, KEHK90, KEHK91, Or94]. Short [Pli95, RS99b]. Short-Range [Pli95]. Sides [OK94]. sigma [HWY90, HWY91]. signals [RS91]. Similar [BCR99, Lub91, MM91b, Sam97]. Simple [CMOS97, DW95a, DW98c, Tem96, PA95b, Ver91]. Simplified [GGZ93, KEHK90, KEHK91]. Simulate [PM99, Lub91]. Simulated [CDR99, RHB94, DS90]. Simulating [AQ98, Luk99, MA95, Mon94, SG98, SC96b, Vu96, Vu98, VBD99, ZDD99]. Simulation [ARV99, ALW94, Bal94a, BMGM95, BGB99, BD94a, BMJ^{+}97, BF97, BN94, Bot96, CN95, CBL95, CF94, CCGJ95, CJR95, CB96, Chr97, CR96, Coo99, CLD^{+}96, CS93c, COBA95, DB96, DL93, DQ95, DE99, DFN^{+}99, DMW99, EA97, FBM93, FT96, Fis94, GBCA99, GMS^{+}99, GSB98, GG97, GRB^{+}95, GM95a, GH95, GKDT96, Gri94, GKL^{+}96, GHS93b, GS97b, HMD93, HC97, HF98, Han93a, HCZ99, HK95, Hew97, HK94, HS99a, HZC^{+}95, HC93, IR98, JZQ^{+}95, Jor99, JLSM97, KPCJ97, KF94, KLB99, KWG94, KG92a, KG92b, KLM95, Kuh96, Kul95, Kup98, Lan97, Leb99, LL95, LPR96, MD98, Mal96b, MM91a, MPR99, MSF94, MPPZ96, MQS95, MB96, Nun93, ÖLJ98, PFRB93, PPBC93, PVQ97, Pri95, RM93, RKJ97, RH99, RS99a, SB96a, Sal91, SKO94, SR98, SW99b]. Simulation [SL96, SXG99, STS^{+}97, Swi96, SCC93b, Tan93, Tan94, Tem96, VD97a, VAVB93, WTH^{+}96, Wel95, WK96, WKK98, WCO94, Yin96, Zab97, Zho98, ZDG95, dG92a, dG92b, dG96, vWBS^{+}97, Ano91n, BLK90, BR92a, BR92b, BBK90, BBK91, Boy90, Boy91a, FPRB90, FPRB91, GB90, GBS^{+}90, GBL^{+}92, GG92b, JVR93, JE90, KMB^{+}92b, KMB^{+}92a, KG90b, KG91, KHM92a, KHM92b, Lif92, MD90, Nat92b, PPBC92, Poi91, Pri90, Pri91, Ram90a, Ram90b, RD91, RD92a, RD92b, RWC90, STAS91, SFB91, VAVB92, VY91, VY92, XPK90]. Simulations [AS95a, AS95b, ACLW97, APS97, ADS97, BS99a, BS99b, BS99c, BD96, BDS^{+}99, BDC97, Boa97, BL94, Bro95, CD93, CD95a, Clé96, CL98b, DW98a, DW98b, DECB98, Dru99, DL94b, FBM93, Gho96, GR97, GBS99, HBF93, JVS97, Kho98, KMM96, KM97b, KMS99, KWT^{+}97, KCV99, LB94, LIH97, LWS98, LRJ^{+}99, MLJ97, Mat94, Mau99, Maz94, MB97, NY98, Pet94, Ram95, Ram97b, Rie99a, Ruy93, SS99b, SMSS98, Sch95b, SH98a, Sha96, SMW99, SK96, SDG94, TLW95, Wel98, Wha96, YGH97, YP98, BR92c, BR92d, But90b, CH90b, CH91, DGD90, DD92a, DMZ91, DM90, DM91b, KDL^{+}92, LG90, LG91, MNR90, OsKB^{+}92, PL92b, Poi92, RM90, RM91, Ros90, Sch90, Wes94]. Simultaneous [VD97a, VAVB92, VAVB93, YWS96, YG95, HLD92b, HLD92a]. Sinc [Boy92c, Kou96, Boy92b, JLB90]. sinc-Galerkin [JLB90]. Sine [AHS96, AHS97]. Single [AAK^{+}93, ACT95, BES94, HK98]. Singular [Boy96a, CGKC93, FGM97, HS93b, Nak95, NMB94, Str96, CGKC92, JLB90, KF91]. Singularities [BST95, CLO93, FB94, LO93, Pri95, BM90b, BM91c, EHW90, EHW91, OB95, OGS90, OGS91]. Singularity [FC95, Jak93, ZP95]. Singularly [CMW95]. Sinh [Boy94]. Sinh-Mapping [Boy94]. sintering [VMK91, vdVMK92b, vdVMK92a]. Sites [YR93]. Sivashinsky [AAP97]. Sixth [CF99, For92a, For92b, CRS90]. Sixth-Order [CF99, For92a, For92b, CRS90]. Size [CLD^{+}96, RK99, SGW94, WKMH97]. Skeletons [SW94]. Skyrme [CB94]. Slabs [SCM98]. Slanted [Poi91]. Slater [KU94, QF99]. slender [RIW90]. Slip [GHS93a, MB94, Wer95, HLB94, HL92a]. Slip-line [HL92a]. Slope [Hub99, Yan90]. Sloping [GGP97]. slow [TLES93]. Slowly [KC97, Rob90a, JL96b]. Slug [TM93]. Small [AF90, AF91, BCRR98, Eve96, HA93, Kel94, KLM95, LCM96, TMR92a, TMR92b]. smooth [Che91, HCC91]. Smoothed [CM99b, FQ96, GLN^{+}99, GS97b, LSRS99, Man93b, PL93, SHA95b, WP97, Bal95]. Smoother [Oos97, AA90, AA91, Luc90, Luc91]. Smoothers [OGWW98]. Smoothing [AK98, Ena93, FD93, SC97, ST99]. Smoothness [HS93a]. Smoothness-Enhancing [HS93a]. Snares [Boy96b, Boy97a]. sniffer [BS92a, BS92b, BS91b]. Solar [LMS98]. Solenoidal [BS99b, BS99c]. SOLGASMIX [Web98]. Solid [Iva93, NC93, PM99, SSR96, TS99, UMS99, WDE98, GL95, Sch92c].

Solid-Liquid [UMS99]. Solid-Water-Gas [TS99]. Solidification [BDS^{+}99, FA97, JT96, Kre98, PGD99, YZ98, Alm93, AT92, SKK90, SKK91, SH92, SS91, SS92]. solids [CKQ^{+}93, TP92a, TP92b]. Solitary [Boy95b, Boy97b, GG90, GG92a, CWG90, CWG91, QS90, TPL90]. Soliton [CSZ97, HK93a, HK92]. Solitons [Boy96a]. Solution [AHS96, AHS97, AK94a, AC95, AK95, And94, AIV99, BK93, BB96, BR99, BR97a, BR97b, BK94, BTW96, BD97b, BK97a, BKV98, BFGG94, Can96, COS^{+}98, CH99, CMW95, Cha95, CWC99, CZS98, CL93, CKSB97, Cod99, CDW99a, CDW99b, CSW95, DVD93, DNN93, DH95, EO98, ES98, Eyr95, FO93, FEO95, Fis97, För96, FS97, Fuj95, GP98a, GACN93, Gar98, GO96, GO98, GP98b, GM95b, GHKH93, HKM96, HIY95, HKNT98, HLD92b, HLD92a, HSS97, ID96, IGABB90, IABBG91, nJ93, JMR94, KT99, Kop94, KRGv93, Ku95, LPMS94, LL93b, LR94, MV99, Mik90, Mik91, MW94, MK97, MQS97, OD99, PHL93, PZC98, PBD94a, PRL^{+}99, Pri93, Pri96, Pri98, QW93, QS98, RK93, RFH93, RIP96, SB95, SA95, STW97, SRF93, SI94, Sie99, SR97, Ste92b, Ste92a, TH96b]. Solution [Tho97, Tid97, TRL99, VD97a, VAVB93, VB95, VSH96, VSM99, Wer95, Win97, WLC96, YB98, YM95, YSMG97, Zha98, AM90, Adh92a, Adh92b, Ben90, BWJ90, BDR91, BL90a, BL90b, BY92, Bra92, BW90, BW91, CW92a, CW92b, CRS90, Che91, CH90a, CGSS90, CGSS91, Dar90c, Dar91b, DM91a, DM92a, DM92b, Dav91, Dav92a, Dav92b, Dur91, DEO92, FP91, Fog92b, Fog92a, GY91, Gro91, Gro92, GS93, HG91, HG92, HB90b, Kar91, Kar92, Kop90, Kop91, KS92, LGC92, LE90, LE91, MH90a, MH90b, MH91, MWJ91, MMS90b, MMS91, MB92, MS90b, MS91c, NSR92a, NSR92b, PRS90, PM90, PRMV90, QR90, Ram92a, Ram92b, Raw90, Raw91, RCAM92a, RCAM92b, RS90, Rok90, RKV90, RKV91, SKK90, SKK91, SVS90, SVS91, SSB90, SSB91, SA90a, SA92a]. solution [SA92b, SM92, Sto92, SK92, Tau92a, Tau92b, TKKB92, Tes92, UWSB90, UWSB91, VMK91, VAVB92, Whi90a, Whi90b, vdVMK92b, vdVMK92a]. Solution-Adaptive [PRL^{+}99, SRF93]. Solution-Based [SR97]. Solutions [BKP96, BCM95a, Ben96, BGP98, BCR99, CPJ90, CM94, CM95, CK93b, ER96, Efr97, ETRW95, Fen99, GPWZ96, GS94, GMPGV96, HGFH95, JYH93, JCHW95, nJWP96, JT99, KB96, LCM96, LK94a, LK94b, LO93, NC97, NC93, Nor98, Nor95, SCM98, Sam97, SSO94, TW94, Ven95, XP94, BB91b, BBF92a, BBF92b, Boy91b, Boy92a, DD90, DD91, Gla92b, Gla92a, Gup90, Gup91, Hal90a, Hal90b, HGH90, HGH91, Kar91, Kar92, OB95, Poy92, SMM^{+}90, SA90b, SA91, TMR92a, TMR92b, YSG91]. Solvability [Dar90c, Dar91b]. Solve [LS93a, Oos97, dFD^{+}94, Dar93, LS92b, RLM90]. solvent [JB91, JBvK^{+}91]. Solver [Ano97-29, AK98, Bal94b, BD99, BIAV98, BHJU99, Cle97, DW94a, DW95a, DP93, DB97, DE99, GTD98, Gla95, GKZ97a, JM98, KJ98, KPC93, KLB99, Lif96, LW94, LF96, MGM95, SP98, Sai95, She95, SM98b, SB96b, SCW94, TK96, WA95, Zha94, CMK90, DP92a, Gla91, GL96, KPC92, LPvL93, MDB91a, MDB91b, MIM90, Mul92b, Mul92a, Tou92]. Solvers [Bai97, BBF^{+}99, GM96, GKZ97b, JMT97, LW96a, Mav98, Mon97b, Roe97a, SLL94, SBGK99, STS98, VM93, Dom90]. Solving [Anc94, ADH^{+}93, BL93, BBMB97, Boy95a, BPM98, CG93, CGA94, Cha95, CJS99, CWC99, CBS93, CC97, CMOS97, Cho97, DK96, Fey98a, Fra95, Gou96a, HML99, HS94, HLL99, Jam99, KOS^{+}96, KM95, Kou96, LO96, LHI90, LMB94, LZ96, Mar94a, Mot98, MB94, MBP94, PA95a, Puc97, Sha95a, SS96c, SG94, Taf95, TZA96, VPS95, VP96, WD97, YG94, YMAC99, ZY94, ZWS98, BB91b, BBF92a, BBF92b, BL92, BJ90, CW90, Dra90, yGmC92, GC92, GMCH92, IM95, LK90, LM90c, LR90, LR91, LJT92a, LJT92b, Mar90, Mel91, RIW90]. Some [Dar95, FG97, GA95, HLD96, Mat97, NM95, Pet94, SWS96, SRT98, TO96, Whi94, Wit96a, BY90, BY91, CH90b, CH91, HGH90, HGH91]. Sommerfeld [AC95, Dvo91, DK92d, DK92a, DK92c, DK92b]. Sonic [AK97]. Sound [KCJ95]. Source [CC97, JM98, LeV98, PLD97, SB98, VC99, CW92a, CW92b, CGSS90, CGSS91, LY90a]. sources [KS92]. Space [BH95, BYMZY95, BD93, Cha95, CWC99, DÜ92a, HB95, Hum96b, KRGv93, PLD97, PC99a, PA95a, RFH93, SAS94, YB98, YWS96, ZRR99, dFD^{+}94, BLMR92, Blu92, BW90, BW91, DÜ92b, HM95, HG91, HG92, Kos92, MF92, OsKB^{+}92, Wei92, ZB92a, ZB92b]. Space-Charge-Limited [Hum96b]. Space-Marching [YB98]. Spaces [Mot98]. spacing [TA91a]. Sparse [DR96, KH97, MPPZ96]. Sparse-Mode [MPPZ96]. Spatial [CNG99, EM94, EM95, GS96b, KM93b, MWS96, SWS96, YL98, YG94, Ben92a, Ben92b, CNG17, DM92c]. Spatially [LL93a, LWS98, RM93, SZ97, Bue91, GG92b, HCP93]. Spatially-Developing [LWS98, Bue91]. Special [Ano99r, BGGT90, HMOS95]. species [LG90, LG91]. Specific [NMW^{+}96]. Specified [OK94]. Spectra [AK99, FG97, MO95, Tak92a, Tak92b]. Spectral [ALW94, Ano94t, AQ99, BM94, BY94, BH96, BYMZY95, BCM95a, BGM95, BCM99a, Boy98, BCT98, CLO93, CS93a, CG96, CDF95, CMMH94, Cle97, CD95b, DBS94, DDD98, DB97, DL94a, EHW90, EHW91, EP96, EP99, FKM99, Fis97, FD99, FPQ93, FCC97, Gar98, GA96, GK94, Gir98, GDP96, Hei93, Hei98, HK95, HDL99, Hu95, ISW92a, ISW92b, JCHW95, KI99, KW93, Kop94, Kop98, LIH97, Lif96, LLT94, LQK98, LS98, Ma93a, Mah98, MM95b, MM97a, MPPZ96, NC93, Nic98, PK95, PR93a, PVQ97, Poz99, RK93, RG97, RS99a, SP98, SS96b, SSB90, SSB91, SK93, SRVK96, SMR91, Str95, SF99, Tay94, TTI97, TRL99, Ver97a, Ver97b, Wir97, YGH97, YS98a, ZY94, ZOI97, BKM90, BT92a, BT92b, BM91a, BM91b, BM90b, BM91c, CS92a]. spectral [CP90, CW92c, Dom90, FP91, Ful93, GCMR90, GMP92a, GMP92b, yGmC92, GC92, GMCH92, HWY90, HWY91, Hei92, KHM92a, KHM92b, Kop90, Kop91, LS92a, LD93, Lel92, MMB90, MLB92, PM90, RS91, Raw90, Raw91, Sol92, SK92, Sur94, TMR92a, TMR92b, WMG^{+}91a, WMG^{+}91b]. Spectral-Boundary [EP99]. Spectral-Element [Poz99]. spectral-finite [yGmC92, GC92]. spectral-like [Lel92]. Spectral-Projection [LS98]. Spectral/hp [LQK98]. Spectral/Wavelet [Gar98]. Spectrally [VM94]. Spectrum [GEK^{+}97, PS93b, SPC93, SPC92, Tas92]. Speed [CP99, SXG99, Zho96, Tay91, TNW92a, TNW92b]. Speeds [HAC97, Mar97, SCC93a]. Spencer [NSR92a, NSR92b]. SPH [CM97, CR99, LPC^{+}93, Mon94, Mon97a, Mon97b, MFZ97, MM97b, Wel98]. Sphere [NC93, PNC94, RIP96, SP96, SP99, TTI97, WKMH97, Bar91b, BM90a]. Spheres [CKE99, Zin94a, Zin94b]. Spherical [DL94a, Gir97, Göt99, Jab94, JCA97, KB99, Man93b, NGN97, PDL93, RIP96, Ruy93, SNU98, SBKH92, WDH^{+}92a, WDH^{+}92b, YR98]. Spherical-Gap [DL94a]. spherically [CS91]. Spin [FHL97, RMSB96]. Spin-Orbit [RMSB96]. Spine [MA95]. Spine-Flux [MA95]. Spinor [PJ94]. Splashing [BMJ^{+}97]. Spline [CZ98, Eyr94, Eyr95, FH94, FG90b, FG90a, Göt99, KOS^{+}96, KMS99, Mal96b, Man93a, QF99, SM98a, UWSB90, UWSB91]. Spline-Characteristic [Mal96b]. Spline-Galerkin [Eyr95]. Spline-Projection [Eyr94]. splines [HB90b]. Split [DFaIM98, GCMR90, Hal97, Jac95, KNW99, YP98, GC90, Wei92]. Split-Field [YP98]. split-matrix [Wei92]. Split-step [GCMR90]. Splitting [Asl99, CB97, GVM99, HA93, HB96, HdS97, Hig99, HKR99, KF94, Lin95, LVS90, LS93b, LDB98, RK95, SS99b, Smi99, Sun96, WK99b, XMJ95, Xu99, Dem92a, Dem92b, Dic90, DD92b, IGABB90, IABBG91, KIO91, Rob90a, Shu92a, Shu90, SLV90, VM90]. Splittings [CD98]. spokes [OsKB^{+}92]. Sponge [PZC98]. spreading [CS91]. Spring [CS97a]. Spruce [FT96]. Spurious [BM94, DDD98, nJWP96, LSRS99, PR93a, Vad95, Vad97, MMB90, YSG91]. Square [Rea97, RH98]. Squares [nJ93, OG97, PA95a, VSH96]. St. [Gla95]. Stability [AC95, BD93, BT97a, BBMB97, BK99, CGA93, CS99, DK98, Fab92, Gil97, GS95, GM92, HB96, KM93b, KL92b, LNR99, LLR94, LML99, Liu96a, MM95a, Mas96, Min96a, NMW^{+}96, PC96, PVVS98, PBD94b, SMD98, Sco98, Sil92, Sil93, SC96a, SHA95b, Tay94, VSG95, WH98, WJC93, Bal95, Mal90, Nat92a, Ren92a, Ren92b]. Stable [AD97, AS94, BMZ95, CGA94, CNG99, CMMH94, GO96, GQ97, Hal97, HS93a, HLR99, LW96a, RFL93, ST97, Tei99, CNG17, HHG91, HHG92, Sim99, GO98]. Stably [PB98, YSMG97]. Stack [WK96, WKK98]. stages [SCA92a, SCA92b]. Staggered [AS99b, BS99b, BS99c, CH99, KK96, Kop96, Kop98, LO98, PZC98, Vel93, WSK99, vBvNW95, Ben92d, Ben92c, JR96, MH90a, SW92, ZSK94]. Staggered-Grid [KK96, Kop96, Kop98]. Staging [HBF93]. stagnant [MD90]. standard [BS91b, BS92a, BS92b, WDH^{+}92a, WDH^{+}92b]. State [Ano94s, BLG97, BBGL93, CL98a, Cra94, Ena93, Fen99, JT93, Lav93, Löt94, Shy99, Swe96, Ven95, WS96, YR93, dAT95, BB90a, BB91a, FG90b, FG90a, Kar91, Kar92, Sal91, YSG91, dG92a, dG92b]. States [Boy96a, CDR99, FK97, Rom97, SRVK96, Hoo90]. Static [SWS96, Shu95, ZDG95]. Stationary [DÜ92a, KKK95, vBKG97, DÜ92b, GL95]. Statistical [CB96, MNR90, DD92a, Val91a, Val91b]. Steady [Ano94s, BBGL93, CL98a, CK93a, DNN93, Ena93, ES93, Fen99, Fer95, HB97, Hwa96, JT93, JV95, Lav93, LeV98, LL94, Löt94, RWW94, SDP97, TZA96, Ven95, YB98, Zha94, AB94, BB90a, BB91a, BJ90, CK92, Dic90, ES92, HLB94, HL92a, HL92b, IT90, JC92, Kar91, Kar92, Kor90, LH90, SKK90, SKK91, Sal91, SA90a, Tau92a, Tau92b, WR91, YSG91, dG92a, dG92b]. Steady-State [Fen99, Lav93, BB90a, BB91a, Kar91, Kar92, Sal91, YSG91, dG92a, dG92b]. Steered [GBS99]. Stefan [CMOS97, SVV98]. Stellar [MIM90]. Stellarator [NMW^{+}96, LH92]. stellarators [DHS91]. stencil [PL92a]. stencils [Dur91, DEO92]. Step [Abd95, AS99b, BS93, CM94, HL98, Jac95, JR96, KTE93, Lu99, Per93, Per95, RK99, Sha95a, Sim93, Smo98, ZSK94, Zhu95, GCMR90, KKD90, LM90a, LM91, RKV90, RKV91]. Stepping [CBS93, Hal97, SW99c, SH97]. Steps [Maz97, Ske98, Wri98]. Stiff [Ano94s, BPM98, Jin95, JL96a, SCW94, dM97a, LY90a, KNW99, NWK98]. Stiffness [HKS98, SW99c, HLS94]. Stimulated [GSB^{+}93, GB90, GBS^{+}90, GBL^{+}92]. Stochastic [ACLW97, DGN97, KLM95, Pet94, Pop95, RW96a, RW96b, RSW98, SL96, WD97, BB90b, BB91c, DM90, DM91b]. stochasticity [FA90, FA91]. Stokes [ANL94, AM96, ABC^{+}98, AB94, Arm94, AS99b, BB90a, BB91a, BR97a, BK94, BTW96, Bor91b, Bor92a, Bor92b, BLT96, BJ90, Cha95, CPJ90, CB97, CH90b, CH91, Cle97, Cod99, CJ95a, Cot90, CD95b, CMM93, DBS94, Dau92a, Dau92b, DF96, DD90, DD91, DH95, Dom90, DIV98, ETFS94, EP99, FP91, Fis97, Gar98, GTD98, GB97, Gil97, Glo92, GKM96, GHKH93, yGmC92, GC92, Gup90, Gup91, Hei92, Hei93, Hen94, HI98, Hwa94, ID96, IM95, Jac99, JB93, Joh93, JR99, KIO91, Kla99, Kop98, Kor90, KL92b, LM90a, LM91, LM97, Lee98b, LZ96, LS98, LDB98, LR90, LR91, Luc90, Luc91, MDB91a, MDB91b, MV99, MH90a, Mar94b, MH90b, MH91, MP97, NPC93, Nor96, Nor98, Nor95, NC99, Poz99, Pri95, PM98, RF95, RKV90, RKV91, RvR93, Rus93]. Stokes [SP98, SPC92, SPC93, SD93b, SSB90, SSB91, SNU98, STW97, She95, SS99c, SI94, Shu91, Shu92b, Shu92c, Sjö95a, SA90a, SMR91, Taf95, Tau92a, Tau92b, Tau94, TfNW91, TNW92c, TNW92d, Tid95, Tou97, Tou98, WNY93, WS93, XP94, YYCH98, ZSK94, Zha94, ZP95, ZWS98, vM95]. Stokes/Euler [MP97]. Storage [SH98b]. Störmer [Maz97]. story [Mil91]. straight [LH92]. Strain [CS97a, LCM96, LPC^{+}93, MAS92]. Strassen [DHSS94]. Strategies [BH99, IH95, Mav98]. Strategy [CMW95, CBS93, ESH93, GO96, GO98, Han93a, JR96, KTN^{+}94, MG97b, dFD^{+}94, KvdVPG95]. Stratified [ALW94, BD94a, CMMH94, LW97, Pai97, PB98, SR98, Vie94, YSMG97, YWS96, Hig92a, Hig92b, WKK98]. Stream [AQ99, EP99, vGVO98, SC90]. Stream-Function [vGVO98]. Streamfunction [HC93, Hei92]. Streamline [CK93a, CKR93, Han93b, CK92]. Streamline-Coordinate [CKR93]. Streamlined [DHS94, HLD92b, HLD92a]. Streptomyces [DOOSM99]. Stress [GLN^{+}99, Waj93, KL92b]. Stresses [Kal97]. Stretched [PG97]. Stretching [RF96]. strings [MMR90]. Strip [CL93]. Strong [BL99, BTW96, KKOF97, KB99, SW99b, Tan94, ZWS98, TP92a, TP92b, HL92b]. Strongly [CSW95, HSS97, LZ96, Sco96]. Structural [Ari97, SMW99].

Structure [Dob99, Fat99, Goe95, KB96, PDS99, RWTC95, SS99a, SK98, SBO93, WCSW99, RCAM92a, RCAM92b]. Structured [Kop96]. Structures [BH99, För96, JBA96, PGD99, DT90a, DT90b, DGD90]. Studies [KL93, OS97]. Study [AAP^{+}95, Bry97, CGKC93, CDT98, DÜ92a, Die95, Don94, ES94, EGS96, Eye96, GBS^{+}93, JZQ^{+}95, KCJ95, KC93, LSLG97, LJG96, LZG99, LK98, LK99, LDB98, Qi93, RW96a, Sem97, SF96, hYsY96, YDT93, ZB96, CGKC92, CHW90, DÜ92b, Fau90, FP92, FVZ90, KG90a, KC92, LY90a, NSR92a, NSR92b, QS90, TSS92, YP92b, YP92a, YSG91]. Sturm [JLB90, Ji94, Kob90, Kob91]. sub [VD90]. sub-subsonic [VD90]. subcell [Har92a, Har92b]. Subgrid [Kal97]. Submerged [PM93]. submergedj [PM92]. Subsonic [Gri94, VD90]. Subspace [Mei98, VSH96]. Substitution [CG97b]. Substrate [BMJ^{+}97]. Subsurface [BL98]. Subzonal [CS98a]. Sufficiency [Sco98]. suggestions [Bal95]. Suitable [PPBC93, PPBC92]. Summation [Ano95u, Str94]. Sums [JBA96, Str91, Str92a, Str92b]. SUP [She95]. Supercomputer [Swe94]. Superconducting [KK95]. Superconductivity [Pri96, Pri98, LY90b]. Superconductor [Rom97, ZV97]. Superconductors [GKL^{+}96]. Supercoolings [WS96]. superfluid [Sch90]. superheated [SCA92a, SCA92b]. superior [DS90]. superposition [Cot90]. Supersonic [LL93b, LL94, YB98, HL92a, HL92b, LH90, YTS92]. Support [MRS98, SS95]. Support-Operator [SS95]. Support-Operators [MRS98]. Suppress [BL94]. suppression [AA90, AA91, SCMU92a, SCMU92b]. Surface [CS99, CPB96, CJR95, Clé96, CT98, Dru99, GLN^{+}99, HML99, HS99a, KP97b, KM96, KB96, LSLG97, Löh96, Mad95, MPR99, MA95, Mon94, PDHS94, Pet96, RLT93, RS99a, SAS94, TWV97, TLW95, TM94, Vie94, ZYKC98, BKZ92, BR92c, BR92d, CYD92, CS92b, CS92c, Dol92a, Dol92b, GSY92, Got92, HLS94, Hui90, Hui91, PTM92, Ram90a, Ram90b, Rom92a, Rom92b, SE90, SR92b, SR92a]. Surface-Induced [KB96]. surface-integral [Dol92a, Dol92b]. Surfaces [BES94, Cho93a, CK93a, KP98, CK92, Dra90, HM90, RR90, RR91]. SURFER [LNS^{+}94]. Surgery [LD93]. Suspended [Xia99]. Suspension [SC96b, Sul91, SB91]. Sustained [PS93a]. SW [CZ98]. Sweeping [ZB96]. Swimming [LK99]. Swirl [LSB96, MG97a]. Switch [MM97b]. Symbolic [Fra95, LNR99, LP90, PJ94, VSG95, SMM^{+}90]. Symbolic-Numerical [VSG95]. Symmetric [AS94, DÜ92a, DA94, KOS^{+}96, LV98, Nor98, SLMS98, YS98a, CS91, DÜ92b, NB98, OS90, Sch95a]. symmetrical [OGBSK90]. Symmetrically [SG94]. Symmetries [FWWD94]. Symmetry [AAP97, CW98, MS99a, SE98, hYsY96, YR93, BLK90]. Symmetry-Breaking [hYsY96]. Symmetry-Preserving [MS99a]. Symplectic [CD93, HA93, LS94, Oku95, VM97, CR91, MMS90b, MMS91]. Synchronous [SS96a]. Synopsis [Rie94b]. System [BD94b, CM95, DW98a, HD97, HLL99, HC93, MR98, SH94, SCW94, Tan94, CGSS90, CGSS91, FP92, LT91, LT92, LH92, SH92, SV91, SFB91]. Systematic [CG97a, FK90, Rob90b]. Systems [AS94, Anc94, Ano94s, AKW93, BFW98, BMGM95, BL98, Bru93, BCT98, CG97a, CGA94, CA93, CSZ97, CS98b, CDT98, DW95a, DW96b, ER96, Ena93, Eve96, FMO98, FHL97, GKDT96, GTA99, HGFH95, JDC94, KOS^{+}96, LeV97, LS94, LLT94, MBP94, Noe94, PQ94, SO95, Str95, Tan94, TS99, WDE98, Wit96a, YDT93, Zha94, ZWS98, AF90, AF91, BB91b, BBF92a, BBF92b, BBK90, BBK91, Bru92, CK90a, Har92a, Har92b, Hoo90, Kob90, Kob91, LO91a, LO91b, Lub91, MMS90b, MMS91, MF92, PL92a, RKV90, RKV91, Row91, SKK90, SKK91, SW92, STAS91, Shu92a, Tho90, tVB92a, tVB92b]. Systolic [Bru92, Bru93, BF91].

T3D [WKHS97]. Tables [KU94, Swe96]. Tamm [MP93]. Tan [Boy91b, Boy92a]. Tangent [Poy92]. tangential [RLM90]. tangle [But90b, Sch90]. Target [SB98]. Tau [DDD98, MMB90, Wer95]. Taylor [AQ98, ABLD94, Bar91a, Dar95, Dar96, HCZ99, KG92a, KG92b, Kup98, LJG96, MD95, MD90, Mot98, SO95, Sch95b, SK90]. Taylor-Expansion [Sch95b]. Taylor-vortex [SK90]. Tcherenkov [ADS97]. Tearing [CGKC93, CGKC92]. Technical [Bar94]. Technique [AH99, ART95a, ART95b, ARTAAA97, CB96, Dur93, HP99, JR96, JLSM97, KL93, LW96b, Nun93, PJ94, Rea97, RG97, SSR96, SBO93, Wit96a, YL98, ZWS98, AAP^{+}95, Cha90, Dau92a, Dau92b, DK92a, DK92b, HG91, HG92, HL90, HL91, Kan92, KL92a, Val91a, Val91b]. Techniques [AS94, APS97, BH96, CN95, CO95, ID96, Tid97, WVHS98, BLMR92, Blu92, DeV91, HGH90, HGH91, SK92, ZS92a, ZS92b]. Telescoping [LRW95]. Temperature [BLL99, CJ95b, Gag98]. Temporal [AM96, CRRR97, RS99b]. temporally [Bar91a]. Tension [Pet96, BKZ92, HLS94, SE90]. Tensor [Man93a, Waj93, dAT95, MAS92, OGBSK90]. Term [Ara97, KLBD93, Lil97, CGSS90, CGSS91]. Termination [Pet98]. Terms [CSW95, Efr97, HI98, JM98, Jin95, JL96a, LeV98, MWW96, PLD97, VC99, CHW90, LT91, LT92, LY90a]. Terrain [PR93b]. tessellation [Zan91]. Test [JCHW95, BS91b, BS92a, BS92b, WDH^{+}92a, WDH^{+}92b]. Testing [BFGG94, MDA91]. Tests [LSRS99, MKM98, PO90, PO91]. Tetrahedral [Gil97, PK94, SK96]. Their [Boy97b]. Theories [KB96, lX98]. Theory [AFH94, Boy97b, CMK90, EVP97, Eyr94, GM98, Hof95, Hof97, LB93, MBI^{+}97, NY98, PC99b, PX93, UL98, VSG95, XP94, lX96, lX97, Xu99, LGC92, May92b, May92a, RNSV91, Rok90, Rus90, Str91, Str92a, Str92b]. Theory-Based [Xu99]. There [KHW98, Wer98]. Thermal [HMD93, HCD98, KRGv93, LEI96, Zin94b, CW92a, CW92b]. Thermally [FMO97, MLB92]. Thermo [PVVS98]. Thermo-Lattice [PVVS98]. Thermoacoustic [WK96, WKK98]. Thermocapillary [LSLG97]. Thermodynamically [Swe96]. Thermodynamics [LPP99, WCO94]. Thermomechanical [BMJ^{+}97]. Thick [WCSW99]. Thick-Restart [WCSW99]. Thickness [BCM99b]. Thin [BCM99b, Bis95, Hel96, LNR99, Pri98, Sch95a, LY90b]. Third [Boy97b, MS93, TL98]. Third- [TL98]. Third-Derivative [MS93].

Third-Order [Boy97b]. Thomas [Gro91, Gro92, Hof95]. Threaded [Kho98]. Three [AS95b, AR93, ABC94, AS99b, ADH^{+}93, BY94, Ber96, BO96, BM90a, BLP98, BS98b, CS99, CE97, CF98, CF99, DW97, DR96, DW93, DIV98, DL94a, EO98, ES93, EGS96, Fis94, Gel99, GG97, GB97, GM98, GLN^{+}99, KT99, KG90b, KG91, KKOF97, LJG96, LPC^{+}93, LL94, MD95, MLBW97, OK94, PDHS94, RKJ97, RHT96, SH99, Sch96, SLMS98, SBC99, ST93, Tan93, Tan94, VO96, Vu98, WSK99, WL93, WWW95, XMP97, YYCH98, YTS92, ZLOT98, ZB96, Zha98, ZYKC98, Bar91b, Bil90, Bue91, CPJ90, CGR99, ESM98, ES92, FVZ90, Gre91, Gre92a, Gre92b, HSN90, HB90b, KG90a, LS92a, PPM92a, PPM92b, RCAM92a, RCAM92b, RG90, RIW90, ST92a, SA92a, SA92b]. three-body [HB90b]. Three-Center [SH99]. Three-Dimensional [AS95b, AR93, ADH^{+}93, BY94, Ber96, BO96, BS98b, CS99, DW97, DR96, DW93, DIV98, DL94a, EO98, ES93, EGS96, Fis94, Gel99, GG97, GB97, GM98, GLN^{+}99, KKOF97, LJG96, LPC^{+}93, LL94, MD95, MLBW97, PDHS94, RKJ97, RHT96, SLMS98, SBC99, ST93, Tan94, VO96, Vu98, WSK99, WL93, WWW95, YYCH98, ZLOT98, ZB96, ZYKC98, BM90a, KG90b, KG91, XMP97, YTS92, Bar91b, Bil90, Bue91, CPJ90, ESM98, ES92, Gre91, Gre92a, Gre92b, HSN90, KG90a, LS92a, PPM92a, PPM92b, RCAM92a, RCAM92b, RG90, RIW90, ST92a, SA92a, SA92b]. Three-Phase [CE97]. Three-Point [CF98, CF99]. Threshold [DS90]. Thruster [RHT96]. Tidal [Lya99, YS98a]. Tide [KD97]. Tides [Boy96b, Boy97a, LG94]. Tikhonov [WKMH97]. Time [AN98, AS94, Ano97-29, AIV99, Ben96, BKV98, BS93, Bis95, BGP98, CBL95, CGA94, CDF95, CBS93, CM94, CM95, Clé96, Cod99, DZ96, EOS98, GA95, GG97, GP98b, Gre94, GK96, GBS99, Hal97, HG94, HT95, HPS92b, HDL99, HdS97, Hig99, HK94, HL98, JT93, KTE93, LP97, LZ96, MML98, Mar94b, Maz97, NC97, OP97, ÖLJ98, RD92a, RD92b, Rei99, RK99, Rob99, Rom97, SB95, SNU98, Sha95a, Sha96, Sha99, Sho93, Ske98, Smo98, SB96b, SW99c, SWD95, Sun96, SH97, TCS97, Tho90, VM97, Wri98, XMP97, ZLOT98, ZSK94, AF90, AF91, BYMZY95, Cha95, CWC99, CH90b, CH91, CM91, DÜ92a, DÜ92b, DD92a, Dra90, HT94, ISW92a, ISW92b, LKLE90, LLEK91, LM90b, LHI90, LRB^{+}90, LBC^{+}91, PA95a, RD91, SM92, Sto92, TEKC92, VC91, Wat92b, Wat92a]. time [YSG91, Zan91, ZRR99, dFSS92]. Time- [AN98]. Time-Dependent [AIV99, CBL95, GP98b, GK96, HG94, HK94, LP97, NC97, OP97, Rom97, SB95, SNU98, Sha99, SWD95, TCS97, ZSK94, Tho90, XMP97, CH90b, Dra90, LHI90, LRB^{+}90, TEKC92, Wat92b, Wat92a]. Time-Domain [Bis95, Clé96, DZ96, EOS98, HDL99, ÖLJ98, Rob99, Sha95a, Sha96]. Time-Harmonic [HT95]. Time-implicit [RD92a, RD92b, RD91]. Time-Marching [LZ96]. Time-Periodic [BGP98]. time-resolved [LKLE90, LLEK91]. Time-Stable [CGA94]. Time-Step [HL98, Smo98]. Time-Stepping [SW99c]. Time-Symmetric [AS94]. times [LKLE90, LLEK91, VC90]. timestep [AF90, AF91]. Timestepping [RH91]. Title [Ano95t, Ano96w, Ano97-28, Ano98-27, Ano99-28]. Today [Sch99]. tokamak [HG92, KPC92, KC92, TT90, TT91, Zan92a, Zan92b, HG91, KL93, KPC93, KM95, KC93, TS98, WJC93, Zan97]. Tokamaks [BK99, TA91b, GLM^{+}90, TAD92a, TAD92b]. Tomography [HP99, Yor90]. Tool [LNR99]. Topography [SH94, UWH99]. Topography-Following [SH94, UWH99]. Topology [Mat97, MG97b]. Toroidal [CGKC93, ESH93, EH93, KPCJ97, PBD94b, SS96b, VM96a, CLJ^{+}90, CHL^{+}90, CGKC92, HCJ92a, HCJ92b]. Total [CBSW98, Cod99, TO96]. totalistic [BRR92]. Tracer [OP97, Löh90, LA90]. Tracing [Ben96, Rie99b]. Track [BST95]. Tracking [ALTP98, BW95, DL94b, FB94, GT97, GLN^{+}99, JT96, LS96, PAB^{+}97, QTL98, RK98, Wel95, KHM92a, KHM92b, RT92, UT91, UT92b, UT92a]. Traditional [MW97]. Trajectories [Bal94a, GSB^{+}93]. Transfer [AN98, BL99, BR99, DW98a, Fra95, KI99, MWS96, PNC94, Sie99, Wel95, dG96, BKP91, Sal91, SM92, Sto92, dG92a, dG92b]. Transform [BCT98, JCHW95, Osb93, UL98, BO92, Osb90, Osb91, PO90, PO91, Boy92d, Tak92a, Tak92b]. Transformation [SPC93, SPC92]. Transformations [SH99, Spe95, GE92, HS90]. Transforms [Gue94, PDL93, vNZ94, SBKH92, SBKH92, Sor95]. Transient [NC93, PNC94, ESM98]. transistor [RCAM92a, RCAM92b]. Transition [BJ95, HRL99, LL95, PV93, RM93, ZDD99, Zho98]. Transitional [PM98]. transitions [Bar91a]. Translation [lX98, ESM98]. Transmembrane [DMW99]. Transmission [ZLP97]. Transonic [WLC96]. Transparent [AE98, SY97]. Transpiration [YG95]. Transport [ACT95, AMP^{+}98, Azm99, BH95, BB97b, EG95, FO93, Fey98a, FWBS94, FD93, GP98a, GMM99, JFH^{+}98, KL93, KRGv93, Kuh96, Kul95, KST90, KST91, LW95, LNR99, Li97, LSRS99, MH94, MKM98, PFS98, SZ97, SS94, SS96a, STW97, SCT^{+}99, STS^{+}97, Tol94, TO96, UL98, YWS96, Zal97, ZWS98, Bor91a, BLA92a, BLA92b, CYD92, DeV91, DHS91, Dra90, Fig92, Gra90, KG92a, KG92b, KS92, Lay90, Nic93, RS90, SG90, Val90, Zan92a, Zan92b, ZDG95]. transport-element [KG92a, KG92b]. transputer [BF91]. Trapped [GSB^{+}93]. Traps [Boy96b, Boy97a]. Travel [Ben96, Due93]. Traveling [Nic98]. traversals [Her90]. Treat [ADS97]. Treating [NC95, OGS90, OGS91]. Treatment [CNG99, Dor99, FC95, FJ98, LM95, kM93a, MLS99, PR93a, RF90, RF91, VC99, BM91a, BM91b, BM90b, BM91c, CNG17, Kal91, Kal92, LBC^{+}93, kM90, kM91, Mao92b, Mao92a, SE90]. Treatments [CGA93, LW96a]. Tree [Bar90, CDM98, Kho98, Str99a, Str99c, Her90, Mak90a]. Tree-Based [Str99a]. Tree-Code [CDM98]. Treecode [SW94, Mak90b]. Trefftz [BR99]. Trends [FH97]. Trial [BK94]. Triangle [Dur91, DEO92, Dur93, Tho97, Win97]. Triangular [DS97a, Hei98, HS99b, Hum96b, JV95, LWC93, MK90, PV99, RWW94, Wan97, YMAC99, ZL93]. Triangular-mesh [MK90]. Triangulated [BS98a, Hui90, Hui91, RS94b]. Triangulation [KP98, Mav95, Reb93, Mav90]. Triangulations [Löh96]. Trigonometric [JBA96]. Trinucleon [BD94b]. TRIPIC [MK90]. Truly [Oli94]. Truncating [YL98]. Truncation [HI98, JC92, LT91, LT92]. Tube [EOS98, ÖLJ98]. Tubes [Bal94a, MD90, VD90]. tunnel [XPK90]. Tunneling [CCGJ95]. Turbulence [AS96, DFN^{+}99, Gho96, HHF95, LZ96, Mal96b, RM93, Sco96, SCG94, Tem96, Wal94, KL92b, Sch90]. Turbulent [ARB96, BLG97, CM94, CR96, DW93, DS97b, DE99, EM94, Fis94, GM95a, Gri94, HK95, HS99a, KMM96, KM97b, KMS99, LWS98, MD98, MPPZ96, MJPC99, PVQ97, Pop95, Pri95, PM98, RHB94, SR98, SC96b, Ush96, WP97, Wel98, XP99, DM92c, KHM92a, KHM92b, RM90, RM91, TMR92a, TMR92b]. Turkel [NGN97]. TVD [AD93, AR97b, HHF95, JP95, JT93, LJG96, Tan94]. TVD-like [AD93]. TVD/AC [LJG96]. Twinning [HRL99]. Twisted [Kla96]. Two [AC95, AS95a, ACLW97, Ara97, AD93, BM94, BGB99, BBMB97, Boy98, Bro95, CS93a, CZ98, CMW95, CA93, CM99a, Clé96, Cle97, CEG^{+}97, CDT98, CRRR97, Cra94, DW97, Die95, Dri97, DE99, DA94, Dur93, Eyr94, FAH97, FKM99, FJGP94, FCC97, Gel99, GB97, HBF93, HOS93, HKNT98, Her93, HK94, HSZ^{+}97, HB97, HLL99, IH95, IS96, JH97, Jac99, KKK95, KPC93, LO96, LBB94, LS96, LL96, Lil97, LMB94, MBV93, kM93a, MPG92a, MPG92b, MB97, NMB94, NGN97, Nor98, Pai97, PB98, PBBS95, Pen95, PMR97, PL98, RCA93, RL97, RS90, SH99, Sai95, SH92, SVV98, SBGM92, SM98a, SW96, SSO94, SAB^{+}99, TSW95, TN98, TP97, TO98, TK96, TM93, UWH99, VD97a, Ver97a, Ver97b, VS95, VG92b, VG92a, Wel95, Wel98, Wha96, YWS96, ZHR97]. Two [ZZ98, dAT95, AS90, AS91, BRR92, Bar91b, BKM90, Bor91a, BLA92a, BLA92b, But90a, CS92a, Cas90, CKQ^{+}93, DZ91, DP91, DP92b, DP92c, Fig92, Fog92b, Fog92a, yGmC92, GC92, GMCH92, Hor90, Hor91, HL90, HL91, Jac90, KPC92, LHI90, LPvL93, Mak90a, Mao92b, MMS90a, MF92, QR90, RCA92, RH91, Rok90, SC90, SKK90, SKK91, SMR91, VMK91, vdVMK92b, vdVMK92a]. Two- [DW97, Gel99, GB97, Bar91b]. Two-Body [Cra94]. Two-Dimensional [AS95a, ACLW97, Ara97, AD93, BGB99, Bro95, CS93a, CZ98, CA93, Eyr94, FAH97, FKM99, FJGP94, FCC97, HOS93, HKNT98, HK94, HB97, HLL99, IH95, IS96, JH97, KKK95, KPC93, LS96, Lil97, LMB94, kM93a, MB97, NMB94, NGN97, Pai97, PB98, PBBS95, RCA93, SVV98, SM98a, TSW95, TN98, VD97a, Ver97a, Ver97b, VS95, Wel98, Wha96, ZZ98, SH92, SBGM92, AS90, AS91, BRR92, BKM90, Bor91a, BLA92a, BLA92b, CS92a, Cas90, CKQ^{+}93, Fig92, Fog92b, Fog92a, yGmC92, GC92, GMCH92, Hor90, Hor91, HL90, HL91, KPC92, LHI90, LPvL93, Mao92b, RCA92, SC90, VMK91, vdVMK92b, vdVMK92a]. Two-Fluid [DE99, TK96]. Two-group [VG92b, VG92a]. Two-Layer [CRRR97, Die95]. Two-Level [PL98]. Two-Nucleon [dAT95]. Two-Phase [BBMB97, CM99a, CEG^{+}97, CDT98, DA94, Jac99, LL96, PMR97, SSO94, SAB^{+}99, TP97, TM93, Wel95, RS90]. Two-Point [AC95, CMW95, Pen95, MPG92a, MPG92b, Jac90, QR90]. Two-Way [CM99a]. Type [BT97b, BL93, BO98, BCT98, CS98a, GC95, Gou96a, GKL^{+}96, GS97b, Hub99, KK95, MBF94, MR98, Pri98, Rom97, Sai95, SY97, Shy99, XH98, Bar91b, BL92, GM92, Kla99, Kle95, KH97, Mar90, Shu92a, Wei92, EMRS91]. Type-II [GKL^{+}96, KK95, Pri98].

Uhlenbeck [FK90]. Ultimate [Hir97, vL97c]. Ultra [CJ95b, SKR^{+}93]. Ultra-Low [CJ95b]. Ultra-relativistic [SKR^{+}93]. Ultrarelativistic [CM97]. Unbounded [HPG98, MM97a, RL97, CW92c, LM90c]. Uncertain [DGN97]. Unconditionally [GQ97, HLR99, RFL93]. Unconstrained [BCM99b, PDS99]. Unconventional [KL97]. Under-resolved [Bro95, MB97]. Underwater [AE98]. undetermined [Ram92a, Ram92b]. Undulatory [LK99]. Unfactored [ZB96]. Unfolding [WKMH97]. Unforced [PV93]. Uniaxial [YP98]. Unified [AS95a, AS95b, AS97, BW98, Fuj95, HLL99]. Uniform [CS92a, CS93a, JCA97, SB96a, SCM98, BK93, MS91b, MS92b, MS92a, RCA92, RCA93, vBvNW95]. Uniformly [HEOC97, HOS93, Shu97]. Unipolar [Fen99]. Unit [Rea97]. Unknown [LL99, Nor95]. Unlimited [OG97]. Unnested [LS93a, LS92b]. unpwind [Tay91]. Unsplit [DW97, PLD97, Sal94]. Unstable [KO98]. Unsteady [AR97a, AR97b, CJM97, DF96, EL96a, EOS98, FH98b, HC97, HF98, JB93, JR96, LEI96, LZS98, MW97, Mar94b, MW94, PBC^{+}95, PH95a, Rig94, SCM98, SK98, Taf95, VM96b, WL93, Xia99, CKQ^{+}93, Dol92a, Dol92b, yGmC92, GC92, GMCH92, HPS92a, HPS92b, MGP91, Ram90a, Ram90b, RKV90, RKV91, SH92, She90, She91, WK96]. Unstructured [Abg94, ABLD94, And94, ARB96, CKE99, Fri98, Gil97, Gir98, HK95, Hub99, KJ98, KPP99, LS93a, Löh95, LQK98, LBL98, Mad95, Mav98, Mei98, OS98, OG97, Reb93, SLMS98, Tan94, TPRC96, VM93, Ven95, VM96b, LS92b, Löh90, LA90, SAB95]. update [RWC90]. updated [KvdVPG95]. upon [Ano99-30, CHR99]. Upper/Approximate [BNW96]. Upwind [CL98a, CEG^{+}97, FO93, FMO98, Li97, LWC93, MRS94, OGWW98, PRL^{+}99, Rav97, Sal94, SMD98, Shu91, Shu92b, Shu92c, SM92, Sto92, TfNW91, TNW92c, TNW92d, Tol94, TL98, VC99, Weg97, ZB96, BY90, BY91, Col90, NSR92a, NSR92b, ST92b, TNW92a, TNW92b]. Upwind-Biased [Li97]. Upwind-Relaxation [ZB96]. Upwinding [Fey98a, Fey98b, GNHP95, HL98, HB97, LP97, LR98, Noe94]. Use [BL99, GMS^{+}99, Gou96b, KS94, LPvL93, Mit92, PSL99, PC98, Swi96, Waj93, BS91a, Gre90, HCP93, KKD90, SAB95]. Useful [KP97a]. Using [ART95a, ART95b, AR93, AMP^{+}98, AIV99, BS99b, BS99c, BH96, BCM95b, BKP91, Ben96, BCM99b, Bru93, CN95, COS^{+}98, CS93b, CM99b, CDW99b, Coo99, CEG^{+}97, CV97, DNN93, DÜ92a, DFFG96, DK98, DE99, GBCA99, GPWZ96, GÖAY95, Han93a, HOS96, Her93, Jac99, oJ95, JR96, KD97, KB96, Knu95, KW93, KLM95, LCM96, LEI96, LW96b, LK94a, LL96, LRW95, Mad95, MD95, MKM99, MS99a, Mat97, MQS95, MS99b, MFZ97, MLBW97, MQS97, MA96, PM93, PGD99, QS98, Rav97, RS99a, RT94, SH99, SK98, SH98a, SCR97, SHF97, SH94, SG98, TLW95, TRL99, WP97, XMP97, YH95, YSD99, ZC94, ZB96, ZMOW98, ZWS98, Bey91, Bey92, BE91, BE92a, BE92b, BF91, Bru92, BS91b, BS92a, BS92b, DD90, DD91, DÜ92b, ESM98, HWY90, HWY91, HLB94, HHG91]. using [HHG92, HLD92b, HLD92a, HS90, KG92a, KG92b, KL92b, MKM04, Mav90, MW93, PM92, RM90, RM91, SH92, SCMU92a, SCMU92b, Ste92b, Ste92a, SK92, TfNW91, TNW92c, TNW92d, YTS92]. Utilizing [CBSW98].

V [Hir97, CS98b]. vacancy [MH92]. Vacuum [TH96a, Hor90, Hor91]. Vaguelette [FS97]. Validation [JT99, VY91, VY92]. validity [Sch90]. Value [AC95, CLO93, CMW95, DS94, NM95, Pen95, ST97, TKT97, BDR91, For90, Jac90, Or94, QR90, Ram92a, Ram92b, UWSB90, UWSB91, YHW90, YHW91]. Valued [PZ98]. Values [Fri98, BRL91]. Vapor [EGS96]. Variable [ABC^{+}98, AIV99, CLD^{+}96, HKM96, KL93, OS98, PAB^{+}97, RFH93, Sak96, SBGK99, Sim93, SR99, XS93, BM92, Ben90, Dar90b, HCP93, KKD90, MH90b, MH91, SV91, SA90b, SA91, SA92a, SA92b]. variable-density [BM92]. Variable-Size [CLD^{+}96]. Variable-Step [Sim93, KKD90]. Variables [Dor99, Han93b, LL96, HN95, RP90, RP91]. Variant [DHSS94, PQ94]. Variation [TO96, Adh92a, Ver91]. Variational [Alm93, BD94a, CI97, GLM^{+}90, Lya99, Pri96, SR92b, SR92a, WDE98, ZCMO96, ZMOW98, dAT95, Adh92b, GL95, TLES93]. Variations [AS99b]. Various [BD93, CE97]. Varying [HB95, Ske98, Wri98]. Vector [CB97, EL97a, Eye96, GVM99, Knu95, LDB98, RK95, SSW98, Shi93, WONM96, XMJ95, lX98, DLMN91, DLMN92b, DLMN92a, MGP91, Par90, Pri90, Pri91, VM90]. vector-efficient [Par90]. Vectorising [HCC91]. Vectorization [Boy90, Boy91a, Her90, Mak90b]. Vectorized [Löh95, Löh90, LA90]. Vectors [Bai97, Roe97a, YR93]. Velocities [AS99a]. Velocity [CJR95, Cle97, DH95, EL97b, EM95, ES93, GN96, GHKH93, Hwa94, JR99, Kla99, MD95, MG96, MM98, NP94, OP97, Pal93, SH98a, ST93, SW99b, TK99, WWW95, Ben90, Dau92a, Dau92b, ES92, GS93, MD90, Nad95, OsKB^{+}92, SV91, ST92a]. Velocity-Impulse [EL97b]. Velocity-Scaled [SH98a]. velocity-space [OsKB^{+}92]. Velocity-Vorticity [Cle97, DH95, GHKH93, Dau92a, Dau92b]. Venant [Gla95]. Verlet [Gre94, Pal93, SW99b]. Version [CLO93, RPB99, Sch92b, Sch92a]. Vertex [CMM93, KW92, MRS94]. Vertex-centered [KW92]. vertical [HWY90, HWY91, MD90]. Very [LM95, SW96]. via [AAP^{+}95, ARV99, ABLD94, Azm99, BH98, BS99d, CG97b, Cho93a, EVP97, GACN93, Gla92b, Gla92a, GNH96, JB95, Kan92, KL92a, Kou96, OP97, SK90, SBKH92, STG^{+}96, YG95]. vibrational [GC90]. Victor [NP92]. View [Dra90]. View-factor [Dra90]. Virtual [SB96a, ZC94]. Viscoelastic [Bot98, HC93, Kup98]. Viscoplastic [Sil93, Sil92]. Viscosities [AD93, JL96b]. Viscosity [ALW94, BL99, Ben96, Bih96, Boy98, CSW98, Cot96, Efr97, Kab96, MM97b, ZWS98, GG92b, MW93]. Viscous [AC93, BH96, BLG97, BS96, CM93, Cho97, DNN93, EL96a, EL97a, Edw96a, GG97, GNH96, GQ97, Han93a, HW96, Jac95, JV95, Kop94, KLP94, LQK98, LKK99, MB99, MG97a, Mav98, Mei98, MB94, Puc97, RG97, RWW94, RH99, SDP97, TZA96, TPL90, Tsy95, VRD99, WK99a, WWW95, Xia99, YMUS99, YG95, BP90, Fis90, GHS93b, HLB94, Har92a, Har92b, Kal91, Kal92, LGC92, MGP91, MM91b, Mav90, Nat92b, PL92b, Poi92, Ram90a, Ram90b, RIW90, Sch92c, SBGM92, SA92a, SA92b, Tay91, TNW92a, TNW92b, UT91, UT92b, UT92a, VMK91, YKM90, vdVMK92b, vdVMK92a]. Viscous-Inviscid [YG95]. Viscous/Inviscid [AC93]. Visualization [Shi93]. Vlasov [GRB^{+}95, BGB99, GB90, GBS^{+}90, GBL^{+}92, GBS^{+}93, Her93, Hol96a, KDL^{+}92, KF94, MSF^{+}95, Nun93, SH98a, SRBG99]. Voiding [LZG99]. Voltage [MQS95, XPK90]. Volume [AL97, Ano93m, Ano93n, Ano93p, Ano93q, Ano93r, Ano94m, Ano94n, Ano94o, Ano94q, Ano95p, Ano95q, Ano95s, Ano96p, Ano96q, Ano96r, Ano96s, Ano96t, Ano96u, Ano96v, Ano97s, Ano97t, Ano97u, Ano97v, Ano97w, Ano97x, Ano97y, Ano97z, Ano97-27, Ano98s, Ano98t, Ano98u, Ano98v, Ano98w, Ano98x, Ano98y, Ano98z, Ano99s, Ano99t, Ano99u, Ano99v, Ano99w, Ano99x, Ano99y, Ano99z, Ano99-27, BS99c, BGH^{+}99, CA93, Dur93, FT96, GS97a, GLN^{+}99, Her93, Hub99, KPP99, Kob99, Leb99, LWC93, MW97, Mat97, Mei98, MW94, MG97b, PK94, PS93a, PP93, RK98, Sai95, SLMS98, SBGK99, SC96b, Tan94, WK99b, XMJ95, lX97, ZRR99, vL97b, Ano90i, Ano90j, Ano90k, Ano90l, Ano90m, Ano90n, Ano91h, Ano91g, Ano91i, Ano91j]. volume [Ano91k, Ano91l, Ano92e, Ano92f, Ano92g, Ano92h, Ano92i, Ano92j, Ano93o, Ano94p, Ano95r, HHG91, HHG92, HM90, JC92, May92b, May92a, Tay91, TNW92a, TNW92b, Ano99-30, GKG95]. Volume-of-Fluid [GLN^{+}99]. Volume/Newton [LL96]. Volume/Particle [MJPC99]. Voronoi [MA93b, Zan91]. Voronoĭ [Her93]. Vortex [ABC94, Ber95, BLP98, But90b, Cho93b, Cot96, DD95, EB94, GG97, GKL^{+}96, HOS96, KK95, KCJ95, KLP94, Kou97, LSB96, MKM99, MH93, NB98, Nor96, PH95a, Qi93, Rus93, SvD96, Str96, Str97, SP96, WL93, BP90, BLK90, But90a, But91, Cho90a, Cho90b, DZ91, Fis90, HM95, KG90a, KG90b, KG91, LR90, LR91, MKM04, Nor91a, Nor91b, RS94b, SK90, Sch90]. vortex-tangle [Sch90]. Vortical [RL97, HPS92a, HPS92b]. Vortices [Chr97, Zab97, Bar91a]. Vorticity [AQ99, CPB96, CM99a, Cle97, DH95, Duk91, DM92d, DM92e, EL96b, EL97a, EP99, ES92, ES93, FC95, GN96, GHKH93, MKM98, MG96, MG97a, MM98, NPC93, TK99, WWW95, But90a, Dau92a, Dau92b, GS93, GMCH92]. Vorticity-Stream [EP99]. Vorticity-Vector [EL97a]. Vorticity-Velocity [ES93, GN96, TK99, WWW95, ES92, GS93]. Vries [Boy96a, CK90a, HKR99, KL97, Osb90, Osb91, PO90, PO91, Sch92b, Sch92a].

Waals [Shy99, WKHS97]. WAF [BT97b]. WAF-Type [BT97b]. Wakes [TS96, TSR92]. Walk [HB95, Zan99, HCP93]. Wall [KMM96, CBL95, MB99]. Wall-Bounded [KMM96]. walls [AN90]. Warming [Cha90]. Water [AQ98, BD93, CN95, CLNT98, DZ96, GNHP95, Göt99, HB97, JCHW95, Luk99, Ma93a, NGN97, Nic98, ÖM95, Pri93, STS98, SP99, TH96a, TS99, TTI97, VC99, YM95, Cas90, Nad95, Ste92b, Ste92a, WDH^{+}92a, WDH^{+}92b]. Watson [Reb93]. Wave [AHH94, Asl99, BES94, CSW98, CH99, CP98a, CL93, Don94, ETRW95, FEO95, FKM99, FD99, FCC97, GS97a, GSB^{+}93, GP98b, HT95, HPG98, HHR99, KI99, LNR99, LS96, LeV97, LeV98, Leh99, Löt94, Lya99, Nic99, Osc94, Qi93, Ram97a, RF96, SCG94, SK93, Ske98, SDI98, TS96, TS99, WJTP96, Wri98, YGH97, YP98, ZZ98, ZS97, Ben90, BR92c, BR92d, CLJ^{+}90, ESM98, FM92, Fra92, GG90, GG92a, Glo92, Ods90a, Ods90b, WMG^{+}91a, WMG^{+}91b, dFSS92]. Wave-Current [FCC97]. Wave-like [Leh99]. Wave-Propagation [LeV98]. Wave-Scattering [CL93]. Wavecodes [Sil93, Sil92]. Waveforms [BGH^{+}97, BGH^{+}99]. Waveguides [Lu99]. Wavelet [BK97a, CZ98, EM94, EHM97, FS97, Gar98, GPWZ96, HJ98, RL95, VPS95, VP96, VP97, WA95]. Wavelet-Galerkin [RL95]. Wavelet-Vaguelette [FS97]. Wavelets [EM95, LAE98, PC98, QW93, ZLP97]. Wavenumber [Li97]. Wavenumber-Extended [Li97]. Waves [AQ98, ALW94, BD93, BDC97, Ber96, Boy95b, Boy97b, CLNT98, Clé96, CS93c, DZ96, Duk95, EM95, GRB^{+}95, LML99, Luk99, MML98, Nic98, ÖM95, SD93a, Ber94, BM91a, BM91b, CWG90, CWG91, CKQ^{+}93, Dol92a, Dol92b, GG90, GG92a, Hig92a, Hig92b, QS90, Rob90a, Rom92a, Rom92b, TPL90]. Wavy [MB99, SK90]. Way [CM99a, Lu99]. Weak [Gla95, KB99, ZWS98, Tou92]. Weakly [Boy95b, Boy96a, Boy97b, DA94, JSD95, Nak95]. Weideman [Boy94]. weight [Chi91, Chi92a, Chi92b]. Weighted [Fri98, HS99b, JS96, LOC94, NY98, RW96a, RW96b, RSW98, YYCH98, ZB92a, ZB92b]. Weights [Yak96]. Weiming [Ano99-30]. Weizhang [Ano99-30]. Well [AGH99, AR97b, BST95, LTE93, Wit96b, Glo92]. Well-Behaved [AR97b]. Well-Posed [BST95, AGH99]. well-posedness [Glo92]. Wells [LK94a, LK94b]. Wendroff [Oli94]. WENO [JcW99, MS99b]. Weyl [Noe94]. Which [Lub93, FVZ90]. White [BK97b]. Whole [LL95]. Whose [Gou96a]. Wide [AE98]. Width [GA96, GG92a]. Wigner [MH94]. Wind [GGP97, SMW99]. Wind-Driven [SMW99]. Window [FHKZ97]. Winograd [DHSS94]. Winslow [CI97]. within [Bar94, BM91a, BM91b, CP90, MBI^{+}97]. without [LO98, Mak99, MG97a, PPBC92, PPBC93, Smo98]. Works [Zal97]. Workstations [MML98]. World [Bry97, Sem97]. Wormlike [JVS97]. Wormlike-Chain [JVS97]. Wright [Ske98].

X [Par90]. X-MP [Par90].

Yukawa [BFO99].

Z [AFH94]. Zag [CSZ97]. Zakharov [CJ94, Gla92b, Gla92a]. ZD [TAD92b]. Zernike [LPMS94]. Zero [SBGK99, KDL^{+}92, RG90]. Zig [CSZ97]. Zig-Zag [CSZ97]. Zonal [Fuj95, KMM96, KMS99]. zone [RY94]. Zones [TH96a]. Zwas [NGN97].

References

Aoyagi:1990:RKS

[AA90] Akira Aoyagi and Kanji Abe. Runge–Kutta smoother for suppression of computational-mode instability of leap frog scheme. Journal of Computational Physics, 89(1):254–255, July 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090139R.

Aoyagi:1991:RKS

[AA91] Akira Aoyagi and Kanji Abe. Runge–Kutta smoother for suppression of computational-mode instability of leapfrog scheme. Journal of Computational Physics, 93(2):287–296, April 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190183L.

Adams:1993:CCM

[AAK^{+}93] E. Adams, W. F. Ames, W. Kühn, W. Rufeger, and H. Spreuer. Computational chaos may be due to a single local error. Journal of Computational Physics, 104(1):241–250, January 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710235.

Anistratov:1997:ANC

[AAL97] Dmitriy Y. Anistratov, Marvin L. Adams, and Edward W. Larsen. Acceleration of the nonlinear corner-balance scheme by the averaged flux method. Journal of Computational Physics, 135(1):66–75, July 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999197957351.

Abdolsalami:1995:SEM

[AAP^{+}95] Farzan Abdolsalami, Mehran Abdolsalami, Lennard Perez, Pedro Gomez, and Mark Silva. Study of electron-molecule collision via finite-element method and R-matrix propagation technique: Exact exchange. Journal of Computational Physics, 121(2):314–323, October 15, 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999195901531.

Amdjadi:1997:SBH

[AAP97] F. Amdjadi, P. J. Aston, and P. Plechác. Symmetry breaking Hopf bifurcations in equations with O(2) symmetry with application to the Kuramoto–Sivashinsky equation. Journal of Computational Physics, 131(1):181–192, February 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196955990.

Altas:1994:HAD

[AB94] Irfan Altas and Kevin Burrage. A high accuracy defect-correction multigrid method for the steady incompressible Navier–Stokes equations. Journal of Computational Physics, 114(2):227–233, October 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184711624.

Allievi:1997:GPS

[AB97] Alejandro Allievi and Rodolfo Bermejo. A generalized particle search-locate algorithm for arbitrary grids. Journal of Computational Physics, 132(2):157–166, April 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196956041.

Aavatsmark:1996:DNO

[ABBM96] I. Aavatsmark, T. Barkve, Ø. Bøe, and T. Mannseth. Discretization on non-orthogonal, quadrilateral grids for inhomogeneous, anisotropic media. Journal of Computational Physics, 127(1):2– 14, August 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196901540.

Almgren:1994:FAV

[ABC94] Ann S. Almgren, Thomas Buttke, and Phillip Colella. A fast adaptive vortex method in three dimensions. Journal of Computational Physics, 113(2):177–200, August 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711296.

Almgren:1998:CAP

[ABC^{+}98] Ann S. Almgren, John B. Bell, Phillip Colella, Louis H. Howell, and Michael L. Welcome. A conservative adaptive projection method for the variable density incompressible Navier–Stokes equations. Journal of Computational Physics, 142(1):1–46, May 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198958909.

Abdullah:1994:SGM

[Abd94] Wan Ahmad Tajuddin Wan Abdullah. Seeking global minima. Journal of Computational Physics, 110(2):320–326, February 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999184710291.

Abdallah:1995:CFS

[Abd95] Shaaban Abdallah. Comments on the fractional step method. Journal of Computational Physics, 117(1):179–180, March 1, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185710558.

Abe:1997:DLB

[Abe97] Takashi Abe. Derivation of the lattice Boltzmann method by means of the discrete ordinate method for the Boltzmann equation. Journal of Computational Physics, 131(1):241–246, February 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196955953.

Abgrall:1994:ENO

[Abg94] R. Abgrall. On essentially non-oscillatory schemes on unstructured meshes: Analysis and implementation. Journal of Computational Physics, 114(1):45–58, September 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918471148X.

Abgrall:1996:HPP

[Abg96] Rémi Abgrall. How to prevent pressure oscillations in multicomponent flow calculations: a quasi conservative approach. Journal of Computational Physics, 125(1):150–160, April 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900856.

Ambrosiano:1994:ETG

[ABLD94] John J. Ambrosiano, Scott T. Brandon, Rainald Löhner, and C. Richard DeVore. Electromagnetics via the Taylor–Galerkin finite element method on unstructured grids. Journal of Computational Physics, 110(2):310–319, February 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918471028X.

Abramov:1991:QISa

[Abr91a] D. I. Abramov. Quantum inverse scattering problem as a Cauchy problem. Journal of Computational Physics, 94(1):250, May 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190145B.

Abramov:1991:QISb

[Abr91b] D. I. Abramov. Quantum inverse scattering problem as a Cauchy problem. Journal of Computational Physics, 97(2):516–534, December 1991. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999191900119.

Aslam:1996:LSM

[ABS96] Tariq D. Aslam, John B. Bdzil, and D. Scott Stewart. Level set methods applied to modeling detonation shock dynamics. Journal of Computational Physics, 126(2):390–409, July 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919690145X.

Allievi:1992:ABG

[AC92] Alejandro Allievi and Sander M. Calisal. Application of Bubnov– Galerkin formulation to orthogonal grid generation. Journal of Computational Physics, 98(1):163–173, January 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290181W.

Arina:1993:SAD

[AC93] Renzo Arina and Claudio Canuto. A self-adaptive domain decomposition for the viscous/inviscid coupling. I. Burgers equation. Journal of Computational Physics, 105(2):290–300, April 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183710752.

Ache:1995:NTP

[AC95] Gerardo A. Ache and Debora Cores. Note on the two-point boundary value numerical solution of the Orr–Sommerfeld stability equation. Journal of Computational Physics, 116(1):180– 183, January 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999185710169.

Albrecht:1997:SST

[ACLW97] A. Albrecht, S. K. Cheung, K. S. Leung, and C. K. Wong. Stochastic simulations of two-dimensional composite packings. Journal of Computational Physics, 136(2):559–579, September 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197957818.

Apruzese:1995:SRA

[ACT95] J. P. Apruzese, R. W. Clark, and J. W. Thornhill. A single ray approximation for α particle transport. Journal of Computational Physics, 119(1):156–163, June 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711242.

Arminjon:1993:CTL

[AD93] Paul Arminjon and Alain Dervieux. Construction of TVD-like artificial viscosities on two-dimensional arbitrary FEM grids. Journal of Computational Physics, 106(1):176–198, May 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711010.

Abarbanel:1997:ASF

[AD97] Saul Abarbanel and Adi Ditkowski. Asymptotically stable fourthorder accurate schemes for the diffusion equation on complex shapes. Journal of Computational Physics, 133(2):279–288, May 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197956539.

Adhikari:1992:CKVa

[Adh92a] Sadhan K. Adhikari. Complex Kohn variation principle for the solution of Lippmann–Schwinger equations. Journal of Computational Physics, 101(2):452, August 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290023R.

Adhikari:1992:CKVb

[Adh92b] Sadhan K. Adhikari. Complex Kohn variational principle for the solution of Lippmann–Schwinger equations. Journal of Computational Physics, 103(2):415–421, December 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290412R.

Assous:1993:FEM

[ADH^{+}93] F. Assous, P. Degond, E. Heintze, P. A. Raviart, and J. Segre. On a finite-element method for solving the three-dimensional Maxwell equations. Journal of Computational Physics, 109(2):222–237, December 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183712143.

Adler:1997:I

[Adl97] Berni J. Adler. Introduction. Journal of Computational Physics, 135(2):100, August 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999197957582.

Assous:1996:NAM

[ADS96] F. Assous, P. Degond, and J. Segré. Numerical approximation of the Maxwell equations in inhomogeneous media by a P^{1} conforming finite element method. Journal of Computational Physics, 128(2):363–380, October 15, 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919690217X.

Assous:1997:NST

[ADS97] F. Assous, P. Degond, and J. Segré. A new scheme to treat the numerical Tcherenkov instability for electromagnetic particle simulations. Journal of Computational Physics, 138(1):171–192, November 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999197958186.

Arter:1995:PMS

[AE95] Wayne Arter and James W. Eastwood. Particle-mesh schemes for advection dominated flows. Journal of Computational Physics, 117(2):194–204, March 15, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999185710583.

Arnold:1998:DTB

[AE98]  Anton Arnold and Matthias Ehrhardt.  Discrete trans-

parent boundary conditions for wide angle parabolic equations in underwater acoustics. Journal of Computational Physics, 145(2):611–638, September 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960430.

Auerbach:1990:LTB

[AF90] Steven P. Auerbach and Alex Friedman. Long-time behavior of numerically computed orbits: Small and intermediate timestep analysis of one-dimensional systems. Journal of Computational Physics, 89(1):253–254, July 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090136O.

Auerbach:1991:LTB

[AF91] Steven P. Auerbach and Alex Friedman. Long-time behaviour of numerically computed orbits: Small and intermediate timestep analysis of one-dimensional systems. Journal of Computational Physics, 93(1):189–223, March 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190079Z.

Arisue:1994:ACP

[AFH94] Hiroaki Arisue, Toshiaki Fujiwara, and Yumi S. Hirata. Algorithm to calculate the partition function for Z(2) lattice gauge theory in four dimensions. Journal of Computational Physics, 111(1):156–164, March 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710515.

Abarbanel:1997:MAP

[AG97] Saul Abarbanel and David Gottlieb. A mathematical analysis of the PML method. Journal of Computational Physics, 134(2):357– 363, July 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199919795717X.

Abarbanel:1999:WPP

[AGH99] S. Abarbanel, D. Gottlieb, and J. S. Hesthaven. Well-posed perfectly matched layers for advective acoustics. Journal of Computational Physics, 154(2):266–283, September 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963131.

Abraham:1999:NTS

[AH99] K. J. Abraham and L. M. Haines. A new technique for sampling multi-modal distributions. Journal of Computational Physics, 155(2):380–386, November 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919996343X.

Alava:1994:OPM

[AHH94] Mikko J. Alava, Jukka A. Heikkinen, and Torbjörn Hellsten. On the origin of parasitic modes in numerical modelling of wave propagation in plasmas. Journal of Computational Physics, 114(1):85–99, September 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918471151X.

Ablowitz:1996:NSS

[AHS96] M. J. Ablowitz, B. M. Herbst, and Constance Schober. On the numerical solution of the Sine–Gordon equation: I. Integrable discretizations and homoclinic manifolds. Journal of Computational Physics, 126(2):299–314, July 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901394.

Ablowitz:1997:NSS

[AHS97] M. J. Ablowitz, B. M. Herbst, and C. M. Schober. On the numerical solution of the Sine–Gordon equation. Journal of Computational Physics, 131(2):354–367, March 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196956065.

Averbuch:1999:STD

[AIV99] A. Averbuch, M. Israeli, and L. Vozovoi. Solution of timedependent diffusion equations with variable coefficients using multiwavelets. Journal of Computational Physics, 150(2):394– 424, April 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999199961855.

Abrahamsson:1994:NSC

[AK94a] Leif Abrahamsson and Heinz-Otto Kreiss. Numerical solution of the coupled mode equations in duct acoustics. Journal of Computational Physics, 111(1):1–14, March 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710382.

Araseki:1994:SCP

[AK94b] Hideo Araseki and Shoji Kotake. A self-correcting procedure for computational liquid metal magnetohydrodynamics. Journal of Computational Physics, 110(2):301–309, February 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710278.

Amini:1995:SHE

[AK95] S. Amini and S. M. Kirkup. Solution of Helmholtz equation in the exterior domain by elementary boundary integral methods. Journal of Computational Physics, 118(2):208–221, May 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710935.

Aslan:1997:DNF

[AK97] Necdet Aslan and Terry Kammash. Developing numerical fluxes with new sonic fix for MHD equations. Journal of Computational Physics, 133(1):43–55, May 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956448.

Arge:1998:EAS

[AK98] Erlend Arge and Angela Kunoth. An efficient ADI-solver for scattered data problems with global smoothing. Journal of Computational Physics, 139(2):343–358, January 20, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958745.

Axmann:1999:EFE

[AK99] Waldemar Axmann and Peter Kuchment. An efficient finite element method for computing spectra of photonic and acoustic band-gap materials: I. Scalar case. Journal of Computational Physics, 150(2):468–481, April 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199961880.

Austin:1993:API

[AKW93] Mark A. Austin, P. S. Krishnaprasad, and Li-Sheng Wang. Almost Poisson integration of rigid body systems. Journal of Computational Physics, 107(1):105–117, July 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711289.

Ahuja:1997:PFV

[AL97] Vineet Ahuja and Lyle N. Long. A parallel finite-volume Runge– Kutta algorithm for electromagnetic scattering. Journal of Computational Physics, 137(2):299–320, November 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958022.

Almgren:1993:VAP

[Alm93] Robert Almgren. Variational algorithms and pattern formation in dendritic solidification. Journal of Computational Physics, 106(2):337–354, June 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999183711125.

Agresar:1998:ACF

[ALTP98] G. Agresar, J. J. Linderman, G. Tryggvason, and K. G. Powell. An adaptive, Cartesian, front-tracking method for the motion, deformation and adhesion of circulating cells. Journal of Computational Physics, 143(2):346–380, July 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959678.

Alvarez:1991:LCN

[Alv91] A. Alvarez. Linearized Crank–Nicholson scheme for nonlinear Dirac equations. Journal of Computational Physics, 97(2):585, December 1991. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919190034I.

Alvarez:1992:LCN

[Alv92] A. Alvarez. Linearized Crank–Nicholson scheme for nonlinear Dirac equations. Journal of Computational Physics, 99(2):348– 350, April 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290214J.

Andreassen:1994:SVM

[ALW94] Øyvind Andreassen, Ivar Lie, and Carl Erik Wasberg. The spectral viscosity method applied to simulation of waves in a stratified atmosphere. Journal of Computational Physics, 110(2):257– 273, February 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184710230.

Acharya:1990:AGS

[AM90] S. Acharya and F. H. Moukalled. An adaptive grid solution procedure for convection-diffusion problems. Journal of Computational Physics, 91(1):32–54, November 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090003J.

Akselvoll:1996:EMT

[AM96] Knut Akselvoll and Parviz Moin. An efficient method for temporal integration of the Navier–Stokes equations in confined axisymmetric geometries. Journal of Computational Physics, 125(2): 454–463, May 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196901072.

Arthurs:1998:MAF

[AMP^{+}98] Kayne M. Arthurs, Leon C. Moore, Charles S. Peskin, E. Bruce Pitman, and H. E. Layton. Modeling arteriolar flow and mass transport using the immersed boundary method. Journal of Computational Physics, 147(2):402–440, December 10, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960971.

Abe:1990:CCF

[AN90] Hirotada Abe and Sadanojyo Nakajima. A courant conditionfree modified partially implicit method. Revised electromagnetic particle code PS2M for bounded plasmas with conducting walls. Journal of Computational Physics, 86(2):259–269, February 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999190901016.

Adams:1998:AAC

[AN98] Marvin L. Adams and Paul F. Nowak. Asymptotic analysis of a computational method for time- and frequency-dependent radiative transfer. Journal of Computational Physics, 146(1):366–403, October 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198960636.

Ancona:1994:FLL

[Anc94] M. G. Ancona. Fully-Lagrangian and lattice-Boltzmann methods for solving systems of conservation equations. Journal of Computational Physics, 115(1):107–120, November 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711818.

Anderson:1994:GGF

[And94] W. Kyle Anderson. A grid generation and flow solution method for the Euler equations on unstructured grids. Journal of Computational Physics, 110(1):23–38, January 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710035.

Ajmani:1994:PCG

[ANL94] Kumud Ajmani, Wing-Fai Ng, and Meng-Sing Liou. Preconditioned conjugate gradient methods for the Navier–Stokes equations. Journal of Computational Physics, 110(1):68–81, January 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999184710060.

Anonymous:1990:APAd

[Ano90a] Anonymous. Abstracts of papers to appear in future issue. Journal of Computational Physics, 88(1):251–254, May 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090252V.

Anonymous:1990:APAe

[Ano90b] Anonymous. Abstracts of papers to appear in future issue. Journal of Computational Physics, 88(2):499–503, June 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190901946.

Anonymous:1990:APAa

[Ano90c] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 86(1):256–258, January 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090100F.

Anonymous:1990:APAb

[Ano90d] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 86(2):496–499, February 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919090112E.

Anonymous:1990:APAc

[Ano90e] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 87(1):245–247, March 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090238V.

Anonymous:1990:APAf

[Ano90f] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 90(2):506–509, October 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190901784.

Anonymous:1990:Aa

[Ano90g] Anonymous. Announcement. Journal of Computational Physics, 88(1):250, May 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919090251U.

Anonymous:1990:Ab

[Ano90h] Anonymous. Announcements. Journal of Computational Physics, 91(1):247–248, November 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919090015S.

Anonymous:1990:AIVa

[Ano90i] Anonymous. Author index for volume 86. Journal of Computational Physics, 86(2):500, February 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090113F.

Anonymous:1990:AIVb

[Ano90j] Anonymous. Author index for volume 87. Journal of Computational Physics, 87(2):497, April 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190902767.

Anonymous:1990:AIVc

[Ano90k] Anonymous. Author index for volume 88. Journal of Computational Physics, 88(2):504, June 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190901957.

Anonymous:1990:AIVd

[Ano90l] Anonymous. Author index for volume 89. Journal of Computational Physics, 89(2):492, August 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090166X.

Anonymous:1990:AIVe

[Ano90m] Anonymous. Author index for volume 90. Journal of Computational Physics, 90(2):510, October 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190901795.

Anonymous:1990:AIVf

[Ano90n] Anonymous. Author index for volume 91. Journal of Computational Physics, 91(2):498, December 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190900626.

Anonymous:1990:C

[Ano90o] Anonymous. Corrigendum. Journal of Computational Physics, 91(1):246, November 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919090014R.

Anonymous:1990:NAa

[Ano90p] Anonymous. Notes to appear. Journal of Computational Physics, 87(2):496, April 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999190902756.

Anonymous:1990:NAb

[Ano90q] Anonymous. Notes to appear. Journal of Computational Physics, 91(2):497, December 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999190900615.

Anonymous:1991:APAa

[Ano91a] Anonymous. Abstract of papers to appear in future issues. Journal of Computational Physics, 92(1):257–259, January 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190300A.

Anonymous:1991:APAc

[Ano91b] Anonymous. Abstract of papers to appear in future issues. Journal of Computational Physics, 94(2):494–497, June 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190233B.

Anonymous:1991:APAf

[Ano91c] Anonymous. Abstract of papers to appear in future issues. Journal of Computational Physics, 96(1):237–239, September 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190275P.

Anonymous:1991:APAb

[Ano91d] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 92(2):508–510, February 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190220F.

Anonymous:1991:APAd

[Ano91e] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 95(1):246–249, July 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190262J.

Anonymous:1991:APAe

[Ano91f] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 95(2):511–513, August 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190289W.

Anonymous:1991:AIVb

[Ano91g] Anonymous. Author index for volume 3. Journal of Computational Physics, 93(2):489, April 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999191902094.

Anonymous:1991:AIVa

[Ano91h] Anonymous. Author index for volume 92. Journal of Computational Physics, 92(2):511, February 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999191902216.

Anonymous:1991:AIVc

[Ano91i] Anonymous. Author index for volume 94. Journal of Computational Physics, 94(2):498, June 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190234C.

Anonymous:1991:AIVd

[Ano91j] Anonymous. Author index for volume 95. Journal of Computational Physics, 95(2):514, August 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999191902902.

Anonymous:1991:AIVe

[Ano91k] Anonymous. Author index for volume 96. Journal of Computational Physics, 96(2):494, October 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190249K.

Anonymous:1991:AIVf

[Ano91l] Anonymous. Author index for volume 97. Journal of Computational Physics, 97(2):586, December 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190035J.

Anonymous:1991:M

[Ano91m] Anonymous. In memoriam. Journal of Computational Physics, 95(2):251–253, August 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919190276Q.

Anonymous:1991:NSI

[Ano91n] Anonymous. Note: Semi-implicit particle simulation of kinetic plasma phenomena. Journal of Computational Physics, 97(1):224–234, November 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919190045M.

Anonymous:1991:NAb

[Ano91o] Anonymous. Note to appear. Journal of Computational Physics, 93(2):488, April 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999191902083.

Anonymous:1991:NAc

[Ano91p] Anonymous. Note to appear. Journal of Computational Physics, 94(1):254, May 1991. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919190159I.

Anonymous:1991:NAa

[Ano91q] Anonymous. Notes to appear. Journal of Computational Physics, 93(1):253, March 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999191900908.

Anonymous:1992:APAa

[Ano92a] Anonymous. Abstract of papers to appear in future issues. Journal of Computational Physics, 100(1):205–207, May 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290322P.

Anonymous:1992:APAc

[Ano92b] Anonymous. Abstract of papers to appear in future issues. Journal of Computational Physics, 103(2):478–479, December 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290418X.

Anonymous:1992:APAb

[Ano92c] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 102(1):225–226, September 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105800178.

Anonymous:1992:A

[Ano92d] Anonymous. Announcement. Journal of Computational Physics, 102(2):423, October 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919290384B.

Anonymous:1992:AIVa

[Ano92e] Anonymous. Author index for volume 100. Journal of Computational Physics, 100(2):436, June 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192902695.

Anonymous:1992:AIVb

[Ano92f] Anonymous. Author index for volume 101. Journal of Computational Physics, 101(2):455, August 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290037Y.

Anonymous:1992:AIVc

[Ano92g] Anonymous. Author index for volume 102. Journal of Computational Physics, 102(2):426, October 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290395F.

Anonymous:1992:AIVd

[Ano92h] Anonymous. Author index for volume 103. Journal of Computational Physics, 103(2):480, December 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290419Y.

Anonymous:1992:AIVe

[Ano92i] Anonymous. Author index for volume 98. Journal of Computational Physics, 98(2):350, February 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290167W.

Anonymous:1992:AIVf

[Ano92j] Anonymous. Author index for volume 99. Journal of Computational Physics, 99(2):353, April 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290228Q.

Anonymous:1992:NA

[Ano92k] Anonymous. Note to appear. Journal of Computational Physics, 103(1):187, November 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919290339Z.

Anonymous:1993:APAa

[Ano93a] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 104(1):285–286, January 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183710302.

Anonymous:1993:APAb

[Ano93b] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 104(2):473, February 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710466.

Anonymous:1993:APAc

[Ano93c] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 105(1):186, March 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710661.

Anonymous:1993:APAd

[Ano93d] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 105(2):371, April 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710843.

Anonymous:1993:APAe

[Ano93e] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 106(1):199, May 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711022.

Anonymous:1993:APAf

[Ano93f] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 106(2):400, June 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711186.

Anonymous:1993:APAg

[Ano93g] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 107(1):199, July 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711368.

Anonymous:1993:APAh

[Ano93h] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 107(2):406, August 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711551.

Anonymous:1993:APAi

[Ano93i] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 108(1):197, September 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711745.

Anonymous:1993:APAj

[Ano93j] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 108(2):377, October 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711927.

Anonymous:1993:APAk

[Ano93k] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 109(1):153, November 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183712076.

Anonymous:1993:APAl

[Ano93l] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 109(2):341, December 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183712222.

Anonymous:1993:AIVa

[Ano93m] Anonymous. Author index for volume 104. Journal of Computational Physics, 104(2):474, February 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710478.

Anonymous:1993:AIVb

[Ano93n] Anonymous. Author index for volume 105. Journal of Computational Physics, 105(2):372, April 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710855.

Anonymous:1993:AIVc

[Ano93o] Anonymous. Author index for volume 106. Journal of Computational Physics, 106(2):401, June 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711198.

Anonymous:1993:AIVd

[Ano93p] Anonymous. Author index for volume 107. Journal of Computational Physics, 107(2):407, August 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711563.

Anonymous:1993:AIVe

[Ano93q] Anonymous. Author index for volume 108. Journal of Computational Physics, 108(2):378, October 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711939.

Anonymous:1993:AIVf

[Ano93r] Anonymous. Author index for volume 109. Journal of Computational Physics, 109(2):342, December 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183712234.

Anonymous:1994:APAa

[Ano94a] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 110(1):200, January 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710175.

Anonymous:1994:APAb

[Ano94b] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 110(2):420, February 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710370.

Anonymous:1994:APAc

[Ano94c] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 111(1):204, March 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710552.

Anonymous:1994:APAd

[Ano94d] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 111(2):414–415, April 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710758.

Anonymous:1994:APAe

[Ano94e] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 112(1):210, May 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710928.

Anonymous:1994:APAf

[Ano94f] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 112(2):404, June 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711107.

Anonymous:1994:APAg

[Ano94g] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 113(1):154, July 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711260.

Anonymous:1994:APAh

[Ano94h] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 113(2):353, August 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711429.

Anonymous:1994:APAi

[Ano94i] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 114(1):160, September 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711569.

Anonymous:1994:APAj

[Ano94j] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 114(2):350, October 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711727.

Anonymous:1994:APAk

[Ano94k] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 115(1):248, November 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184712204.

Anonymous:1994:APAl

[Ano94l] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 115(2):553, December 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184712174.

Anonymous:1994:AIVa

[Ano94m] Anonymous. Author index for volume 111. Journal of Computational Physics, 111(2):416, April 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918471076X.

Anonymous:1994:AIVb

[Ano94n] Anonymous. Author index for volume 112. Journal of Computational Physics, 112(2):405, June 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711119.

Anonymous:1994:AIVc

[Ano94o] Anonymous. Author index for volume 113. Journal of Computational Physics, 113(2):354, August 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711430.

Anonymous:1994:AIVd

[Ano94p] Anonymous. Author index for volume 114. Journal of Computational Physics, 114(2):351, October 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184712198.

Anonymous:1994:AIVe

[Ano94q] Anonymous. Author index for volume 115. Journal of Computational Physics, 115(2):554, December 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184712186.

Anonymous:1994:E

[Ano94r] Anonymous. Erratum. Journal of Computational Physics, 110(2): 419, February 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184710369.

Anonymous:1994:NEE

[Ano94s] Anonymous. Note: An evaluation of explicit pseudo-steady-state approximation schemes for stiff ODE systems from chemical kinetics. Journal of Computational Physics, 113(2):347–352, August 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999184711417.

Anonymous:1994:NDS

[Ano94t] Anonymous. Note: “diagonal shadow” — a quasi-Newton iteration in spectral domain. Journal of Computational Physics, 111(2):410–413, April 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999184710746.

Anonymous:1995:APAa

[Ano95a] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 116(1):193, January 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710200.

Anonymous:1995:APAb

[Ano95b] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 116(2):388, February 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710388.

Anonymous:1995:APAc

[Ano95c] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 117(1):181, March 1, 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918571056X.

Anonymous:1995:APAd

[Ano95d] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 117(2):368, March 15, 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918571073X.

Anonymous:1995:APAe

[Ano95e] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 118(1):189, April 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918571090X.

Anonymous:1995:APAf

[Ano95f] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 118(2):401, May 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711114.

Anonymous:1995:APAg

[Ano95g] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 119(1):209–210, June 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711291.

Anonymous:1995:APAh

[Ano95h] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 119(2):390, July 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711448.

Anonymous:1995:APAi

[Ano95i] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 120(1):156, August 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711552.

Anonymous:1995:APAj

[Ano95j] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 120(2):383, September 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711746.

Anonymous:1995:APAk

[Ano95k] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 121(1):192, October 1, 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711904.

Anonymous:1995:APAl

[Ano95l] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 121(2):383, October 15, 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999195902554.

Anonymous:1995:APAm

[Ano95m] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 122(1):184, November 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185712065.

Anonymous:1995:APAn

[Ano95n] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 122(2):387, December 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185712259.

Anonymous:1995:A

[Ano95o] Anonymous. Acknowledgment. Journal of Computational Physics, 122(2):384–386, December 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0021999185712247.

Anonymous:1995:AIVa

[Ano95p] Anonymous. Author index for volume 117. Journal of Computational Physics, 117(2):369, March 15, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710741.

Anonymous:1995:AIVb

[Ano95q] Anonymous. Author index for volume 120. Journal of Computational Physics, 120(2):384, September 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711758.

Anonymous:1995:AIVc

[Ano95r] Anonymous. Author index for volume 121. Journal of Computational Physics, 121(2):384, October 15, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999195902724.

Anonymous:1995:AIVd

[Ano95s] Anonymous. Author index for volume 122. Journal of Computational Physics, 122(2):388, December 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185712260.

Anonymous:1995:CAT

[Ano95t] Anonymous. Cumulative author–title index 1990–1995. Journal of Computational Physics, 122(2):389–410, December 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185712272.

Anonymous:1995:E

[Ano95u] Anonymous. Erratum: “Summation by parts for finite difference approximations for d/dx”. Journal of Computational Physics, 116(1):192, January 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999185710194.

See [Str94].

Anonymous:1996:APAh

[Ano96a] Anonymous. Abstracts of papers to appear in future. Journal of Computational Physics, 126(2):479, July 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901515.

Anonymous:1996:APAa

[Ano96b] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 123(1):233, January 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900200.

Anonymous:1996:APAb

[Ano96c] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 123(2):495, February 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919690039X.

Anonymous:1996:APAc

[Ano96d] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 124(1):242, March 1, 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919690056X.

Anonymous:1996:APAd

[Ano96e] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 124(2):465, March 15, 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919690073X.

Anonymous:1996:APAe

[Ano96f] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 125(1):277–278, April 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900935.

Anonymous:1996:APAf

[Ano96g] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 125(2):526–527, May 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901138.

Anonymous:1996:APAg

[Ano96h] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 126(1):240–241, June 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901333.

Anonymous:1996:APAi

[Ano96i] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 127(1):226, August 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901710.

Anonymous:1996:APAj

[Ano96j] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 127(2):482–483, September 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196901898.

Anonymous:1996:APAk

[Ano96k] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 128(1):259, October 1, 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196902089.

Anonymous:1996:APAl

[Ano96l] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 128(2):498–499, October 15, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196902272.

Anonymous:1996:APAm

[Ano96m] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 129(1):243–244, November 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196902478.

Anonymous:1996:APAn

[Ano96n] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 129(2):435, December 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196902594.

Anonymous:1996:A

[Ano96o] Anonymous. Acknowledgment. Journal of Computational Physics, 129(2):431–434, December 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0021999196902612.

Anonymous:1996:AIVa

[Ano96p] Anonymous. Author index for volume 123. Journal of Computational Physics, 123(2):496, February 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900406.

Anonymous:1996:AIVb

[Ano96q] Anonymous. Author index for volume 124. Journal of Computational Physics, 124(2):466, March 15, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900741.

Anonymous:1996:AIVc

[Ano96r] Anonymous. Author index for volume 125. Journal of Computational Physics, 125(2):528, May 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919690114X.

Anonymous:1996:AIVd

[Ano96s] Anonymous. Author index for volume 126. Journal of Computational Physics, 126(2):480, July 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901527.

Anonymous:1996:AIVe

[Ano96t] Anonymous. Author index for volume 127. Journal of Computational Physics, 127(2):484, September 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901904.

Anonymous:1996:AIVf

[Ano96u] Anonymous. Author index for volume 128. Journal of Computational Physics, 128(2):500, October 15, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196902284.

Anonymous:1996:AIVg

[Ano96v] Anonymous. Author index for volume 129. Journal of Computational Physics, 129(2):436, December 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196902600.

Anonymous:1996:CAT

[Ano96w] Anonymous. Cumulative author–title index 1991–1996. Journal of Computational Physics, 129(2):437–460, December 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196902624.

Anonymous:1997:APAa

[Ano97a] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 130(1):160, January 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196956247.

Anonymous:1997:APAb

[Ano97b] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 130(2):283, January 15, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919695634X.

Anonymous:1997:APAc

[Ano97c] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 131(1):251, February 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956424.

Anonymous:1997:APAd

[Ano97d] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 131(2):378, March 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956540.

Anonymous:1997:APAe

[Ano97e] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 132(1):156, March 15, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956643.

Anonymous:1997:APAf

[Ano97f] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 132(2):414–415, April 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956850.

Anonymous:1997:APAg

[Ano97g] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 133(1):192, May 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919799999X.

Anonymous:1997:APAh

[Ano97h] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 133(2):365–366, May 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197957156.

Anonymous:1997:APAi

[Ano97i] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 134(1):201, June 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957247.

Anonymous:1997:APAj

[Ano97j] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 134(2):389, July 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957612.

Anonymous:1997:APAk

[Ano97k] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 135(1):97, July 15, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957703.

Anonymous:1997:APAl

[Ano97l] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 135(2):293, August 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957892.

Anonymous:1997:APAm

[Ano97m] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 136(1):229, September 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197958010.

Anonymous:1997:APAn

[Ano97n] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 136(2):660–661, September 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197958113.

Anonymous:1997:APAo

[Ano97o] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 137(1):245–246, October 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197958320.

Anonymous:1997:APAp

[Ano97p] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 137(2):406–407, November 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197958514.

Anonymous:1997:APAq

[Ano97q] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 138(1):248–249, November 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197958605.

Anonymous:1997:APAr

[Ano97r] Anonymous. Abstracts of papers to appear in future issues. Journal of Computational Physics, 138(2):991–992, December 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197958940.

Anonymous:1997:AIVa

[Ano97s] Anonymous. Author index for volume 130. Journal of Computational Physics, 130(2):284, January 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956357.

Anonymous:1997:AIVb

[Ano97t] Anonymous. Author index for volume 131. Journal of Computational Physics, 131(2):379, March 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956552.

Anonymous:1997:AIVc

[Ano97u] Anonymous. Author index for volume 132. Journal of Computational Physics, 132(2):416, April 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956849.

Anonymous:1997:AIVd

[Ano97v] Anonymous. Author index for volume 133. Journal of Computational Physics, 133(2):367, May 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957144.

Anonymous:1997:AIVe

[Ano97w] Anonymous. Author index for volume 134. Journal of Computational Physics, 134(2):390, July 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957600.

Anonymous:1997:AIVf

[Ano97x] Anonymous. Author index for volume 135. Journal of Computational Physics, 135(2):294, August 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957909.

Anonymous:1997:AIVg

[Ano97y] Anonymous. Author index for volume 136. Journal of Computational Physics, 136(2):662, September 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958125.

Anonymous:1997:AIVh

[Ano97z] Anonymous. Author index for volume 137. Journal of Computational Physics, 137(2):408, November 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958526.

Anonymous:1997:AIVi

[Ano97-27] Anonymous. Author index for volume 138. Journal of Computational Physics, 138(2):993, December 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958952.

Anonymous:1997:CAT

[Ano97-28] Anonymous. Cumulative author–title index 1992–1997. Journal of Computational Physics, 138(2):994–1033, December 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958964.

Anonymous:1997:EFT

[Ano97-29] Anonymous. Erratum: “A Forward-in-Time Advection Scheme and Adaptive Multilevel Flow Solver for Nearly Incompressible Atmospheric Flow”. Journal of Computational Physics, 132(2): 412–413, April 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999197956667. See [SB96b].

Anonymous:1998:APAa

[Ano98a] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 139(2):444–445, January 20, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197959118.

Anonymous:1998:APAb

[Ano98b] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 140(1):202–203, February 10, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959186.

Anonymous:1998:APAc

[Ano98c] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 140(2):481–482, March 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959411.

Anonymous:1998:APAd

[Ano98d] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 141(1):90–91, March 20, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959484.

Anonymous:1998:APAe

[Ano98e] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 141(2):225–226, April 10, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959630.

Anonymous:1998:APAf

[Ano98f] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 142(1):269–270, May 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959800.

Anonymous:1998:APAg

[Ano98g] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 142(2):615–616, May 20, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960041.

Anonymous:1998:APAh

[Ano98h] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 143(1):289–290, June 10, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919896020X.

Anonymous:1998:APAi

[Ano98i] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 143(2):682–683, July 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919896034X.

Anonymous:1998:APAj

[Ano98j] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 144(1):235–236, July 20, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960454.

Anonymous:1998:APAk

[Ano98k] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 144(2):749–750, August 10, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960545.

Anonymous:1998:APAl

[Ano98l] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 145(1):469–470, September 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960685.

Anonymous:1998:APAm

[Ano98m] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 145(2):760–761, September 20, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960831.

Anonymous:1998:APAn

[Ano98n] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 146(1):488–489, October 10, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960983.

Anonymous:1998:APAo

[Ano98o] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 146(2):775–776, November 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961095.

Anonymous:1998:APAp

[Ano98p] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 147(1):236–237, November 20, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961241.

Anonymous:1998:APAq

[Ano98q] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 147(2):610–611, December 10, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961435.

Anonymous:1998:AI

[Ano98r] Anonymous. Author index. Journal of Computational Physics, 139(2):446, January 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999197999988.

Anonymous:1998:AIVa

[Ano98s] Anonymous. Author index for volume 140. Journal of Computational Physics, 140(2):483, March 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198958193.

Anonymous:1998:AIVb

[Ano98t] Anonymous. Author index for volume 141. Journal of Computational Physics, 141(2):227, April 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959642.

Anonymous:1998:AIVc

[Ano98u] Anonymous. Author index for volume 142. Journal of Computational Physics, 142(2):617, May 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960053.

Anonymous:1998:AIVd

[Ano98v] Anonymous. Author index for volume 143. Journal of Computational Physics, 143(2):684, July 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960351.

Anonymous:1998:AIVe

[Ano98w] Anonymous. Author index for volume 144. Journal of Computational Physics, 144(2):751, August 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960557.

Anonymous:1998:AIVf

[Ano98x] Anonymous. Author index for volume 145. Journal of Computational Physics, 145(2):762, September 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960843.

Anonymous:1998:AIVg

[Ano98y] Anonymous. Author index for volume 146. Journal of Computational Physics, 146(2):777–778, November 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961101.

Anonymous:1998:AIVh

[Ano98z] Anonymous. Author index for volume 147. Journal of Computational Physics, 147(2):612, December 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961447.

Anonymous:1998:CAT

[Ano98-27] Anonymous. Cumulative author–title index 1993–1998. Journal of Computational Physics, 147(2):613–646, December 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198961459.

Anonymous:1999:APAa

[Ano99a] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 148(1):302–303, January 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961587.

Anonymous:1999:APAb

[Ano99b] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 148(2):675–676, January 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961745.

Anonymous:1999:APAc

[Ano99c] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 149(1):198–199, February 10, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199961892.

Anonymous:1999:APAd

[Ano99d] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 149(2):423–424, March 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962083.

Anonymous:1999:APAe

[Ano99e] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 150(1):299–300, March 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962307.

Anonymous:1999:APAg

[Ano99f] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 151(1):422–423, May 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962678.

Anonymous:1999:APAh

[Ano99g] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 151(2):1008–1009, May 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962770.

Anonymous:1999:APAi

[Ano99h] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 152(1):420–421, June 10, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962885.

Anonymous:1999:APAj

[Ano99i] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 152(2):796–797, July 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919996309X.

Anonymous:1999:APAk

[Ano99j] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 153(1):223–224, July 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963179.

Anonymous:1999:APAl

[Ano99k] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 153(2):672–673, August 10, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963258.

Anonymous:1999:APAm

[Ano99l] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 154(1):242–243, September 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963428.

Anonymous:1999:APAn

[Ano99m] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 154(2):528–529, September 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963520.

Anonymous:1999:APAo

[Ano99n] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 155(1):221–222, October 10, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963672.

Anonymous:1999:APAp

[Ano99o] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 155(2):499–500, November 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963805.

Anonymous:1999:APAq

[Ano99p] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 156(1):207–208, November 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963908.

Anonymous:1999:APAr

[Ano99q] Anonymous. Abstracts of papers to appear. Journal of Computational Physics, 156(2):399–400, December 10, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963994.

Anonymous:1999:APAf

[Ano99r] Anonymous. Abstracts of papers to appear in special issue on computational molecular biophysics. Journal of Computational Physics, 150(2):595–596, April 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962538.

Anonymous:1999:AIVa

[Ano99s] Anonymous. Author index for volume 148. Journal of Computational Physics, 148(2):677, January 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961757.

Anonymous:1999:AIVb

[Ano99t] Anonymous. Author index for volume 149. Journal of Computational Physics, 149(2):425, March 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962095.

Anonymous:1999:AIVc

[Ano99u] Anonymous. Author index for volume 150. Journal of Computational Physics, 150(2):597, April 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919996254X.

Anonymous:1999:AIVd

[Ano99v] Anonymous. Author index for volume 151. Journal of Computational Physics, 151(2):1010, May 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962782.

Anonymous:1999:AIVe

[Ano99w] Anonymous. Author index for volume 152. Journal of Computational Physics, 152(2):798, July 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963106.

Anonymous:1999:AIVf

[Ano99x] Anonymous. Author index for volume 153. Journal of Computational Physics, 153(2):674, August 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919996326X.

Anonymous:1999:AIVg

[Ano99y] Anonymous. Author index for volume 154. Journal of Computational Physics, 154(2):530, September 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963532.

Anonymous:1999:AIVh

[Ano99z] Anonymous. Author index for volume 155. Journal of Computational Physics, 155(2):501, November 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963817.

Anonymous:1999:AIVi

[Ano99-27] Anonymous. Author index for volume 156. Journal of Computational Physics, 156(2):401, December 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199964021.

Anonymous:1999:CAT

[Ano99-28] Anonymous. Cumulative author–title index 1994–1999. Journal of Computational Physics, 156(2):402–436, December 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199964033.

Anonymous:1999:E

[Ano99-29] Anonymous. Editorial. Journal of Computational Physics, 148(1):1, January 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999198961472.

Anonymous:1999:EVN

[Ano99-30] Anonymous. Erratum: Volume 149, number 2 (1999), in Article ID jcph.1998.6151, “An r-Adaptive Finite Element Method Based upon Moving Mesh PDEs,” by Weiming Cao, Weizhang Huang, and Robert D. Russell, pp. 221–244. Journal of Computational Physics, 156(1):206, November 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S002199919996391X. See erratum [CHR99].

Aoyagi:1995:NLI

[Aoy95] Akira Aoyagi. Nonlinear leapfrog instability for Fornberg’s pattern. Journal of Computational Physics, 120(2):316–322, September 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185711679.

Ashgriz:1990:FFL

[AP90] N. Ashgriz and J. Y. Poo. Flair: Flux line-segment model for advection and interface reconstruction. Journal of Computational Physics, 89(1):251–252, July 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090129O.

Asencor:1991:FDO

[AP91a] F. J. Asencor and M. Panizo. Finite-difference operators in anisotropic inhomogeneous dielectrics: General case. Journal of Computational Physics, 95(2):387–399, August 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190282P.

Ashgriz:1991:FFL

[AP91b] N. Ashgriz and J. Y. Poo. FLAIR: Flux line-segment model for advection and interface reconstruction. Journal of Computational Physics, 93(2):449–468, April 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190194P.

Antonov:1997:ALB

[APS97] S. Antonov, F.-J. Pfreundt, and J. Struckmeier. Adaptive load balance techniques in parallel rarefied gas simulations. Journal of Computational Physics, 138(2):400–418, December 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958253.

Achdou:1998:EBC

[APV98] Yves Achdou, O. Pironneau, and F. Valentin. Effective boundary conditions for laminar flows over periodic rough boundaries. Journal of Computational Physics, 147(1):187–218, November 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198960880.

Ambrosi:1998:TGM

[AQ98] D. Ambrosi and L. Quartapelle. A Taylor–Galerkin method for simulating nonlinear dispersive water waves. Journal of Computational Physics, 146(2):546–569, November 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960272.

Auteri:1999:GSM

[AQ99] F. Auteri and L. Quartapelle. Galerkin spectral method for the vorticity and stream function equations. Journal of Computational Physics, 149(2):306–332, March 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961551.

Albanese:1993:ATD

[AR93] R. Albanese and G. Rubinacci. Analysis of three-dimensional electromagnetic fields using edge elements. Journal of Computational Physics, 108(2):236–245, October 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711794.

Anderson:1996:HOE

[AR96] Christopher R. Anderson and Marc B. Reider. A high order explicit method for the computation of flow about a circular cylinder. Journal of Computational Physics, 125(1):207–224, April 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196900893.

Arora:1997:POD

[AR97a] Mohit Arora and Philip L. Roe. On postshock oscillations due to shock capturing schemes in unsteady flows. Journal of Computational Physics, 130(1):25–40, January 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196955345.

Arora:1997:WBT

[AR97b] Mohit Arora and Philip L. Roe. A well-behaved TVD limiter for high-resolution calculations of unsteady flow. Journal of Computational Physics, 132(1):3–11, March 15, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919695514X.

Arakawa:1997:CDL

[Ara97] Akio Arakawa. Computational design for long-term numerical integration of the equations of fluid motion: Twodimensional incompressible flow. Part I. Journal of Computational Physics, 135(2):103–114, August 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956977.

Anderson:1996:IMA

[ARB96] W. Kyle Anderson, Russ D. Rausch, and Daryl L. Bonhaus. Implicit/multigrid algorithms for incompressible turbulent flows on unstructured grids. Journal of Computational Physics, 128(2):391–408, October 15, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196902193.

Arian:1997:CAS

[Ari97] Eyal Arian. On the coupling of aerodynamic and structural design. Journal of Computational Physics, 135(1):83–96, July 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197957417.

Armfield:1994:EAC

[Arm94] S. W. Armfield. Ellipticity, accuracy, and convergence of the discrete Navier–Stokes equations. Journal of Computational Physics, 114(2):176–184, October 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711582.

Allison:1991:EOF

[ARS91] A. C. Allison, A. D. Raptis, and T. E. Simos. An eighth-order formula for the numerical integration of the one-dimensional Schrödinger equation. Journal of Computational Physics, 97(1):240–248, November 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190047O.

Al-Rizzo:1995:ESDa

[ART95a] H. M. Al-Rizzo and J. M. Tranquilla. Electromagnetic scattering from dielectrically coated axisymmetric objects using the generalized point-matching technique. Journal of Computational Physics, 119(2):342–355, July 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711394.

Al-Rizzo:1995:ESDb

[ART95b] H. M. Al-Rizzo and J. M. Tranquilla. Electromagnetic scattering from dielectrically coated axisymmetric objects using the Generalized Point-Matching Technique (GPMT). Journal of Computational Physics, 119(2):356–373, July 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711400.

Al-Rizzo:1997:AGM

[ARTAAA97] H. M. Al-Rizzo, J. M. Tranquilla, S. M. Al-Amri, and H. T. Alhafid. Application of the generalized multipole technique (GMT) to high-frequency electromagnetic scattering from perfectly conducting and dielectric bodies of revolution. Journal of Computational Physics, 136(1):1–18, September 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196956181.

Albanese:1999:ICM

[ARV99] Raffaele Albanese, Guglielmo Rubinacci, and Fabio Villone. An integral computational model for crack simulation and detection via eddy currents. Journal of Computational Physics, 152(2):736– 755, July 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999199962617.

Altas:1990:TDA

[AS90] Irfan Altas and John W. Stephenson. A two-dimensional adaptive mesh generation method. Journal of Computational Physics, 90(1):268, September 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919090212J.

Altas:1991:TDA

[AS91] Irfan Altas and John W. Stephenson. A two-dimensional adaptive mesh generation method. Journal of Computational Physics, 94(1):201–224, May 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999191901439.

Alcubierre:1994:TSA

[AS94] M. Alcubierre and Bernard F. Schutz. Time-symmetric ADI and causal reconnection: Stable numerical techniques for hyperbolic systems on moving grids. Journal of Computational Physics, 112 (1):44–77, May 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184710813.

Adalsteinsson:1995:LSAa

[AS95a] D. Adalsteinsson and J. A. Sethian. A level set approach to a unified model for etching, deposition, and lithography I: Algorithms and two-dimensional simulations. Journal of Computational Physics, 120(1):128–144, August 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711539.

Adalsteinsson:1995:LSAb

[AS95b] D. Adalsteinsson and J. A. Sethian. A level set approach to a unified model for etching, deposition, and lithography II: Three-dimensional simulations. Journal of Computational Physics, 122(2):348–366, December 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185712211.

Adalsteinsson:1995:FLS

[AS95c] David Adalsteinsson and James A. Sethian. A fast level set method for propagating interfaces. Journal of Computational Physics, 118(2):269–277, May 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710984.

Adams:1996:HRH

[AS96] N. A. Adams and K. Shariff. A high-resolution hybrid compact-ENO scheme for shock-turbulence interaction problems. Journal of Computational Physics, 127(1):27–51, August 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901564.

Adalsteinsson:1997:LSA

[AS97] D. Adalsteinsson and J. A. Sethian. A level set approach to a unified model for etching, deposition, and lithography. Journal of Computational Physics, 138(1):193–223, November 15, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958174.

Adalsteinsson:1999:FCE

[AS99a] D. Adalsteinsson and J. A. Sethian. The fast construction of extension velocities in level set methods. Journal of Computational Physics, 148(1):2–22, January 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960909.

Armfield:1999:FSM

[AS99b] S. Armfield and R. Street. The fractional-step method for the Navier–Stokes equations on staggered grids: The accuracy of three variations. Journal of Computational Physics, 153(2):660–665, August 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999199962757.

Aslan:1999:MFS

[Asl99] Necdet Aslan. MHD–A: a fluctuation splitting wave model for planar magnetohydrodynamics. Journal of Computational Physics, 153(2):437–466, August 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919996285X.

Akopov:1991:MGP

[ASTAS91] N. Z. Akopov, G. K. Savvidy, and N. G. Ter-Arutyunyan-Savvidy. Matrix generator of pseudorandom numbers. Journal of Computational Physics, 97(2):573–579, December 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190016E.

Asethian:1992:CGD

[AT92] James Asethian and John Train. Crystal growth and dendritic solidification. Journal of Computational Physics, 98(1):178–179, January 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999192901874.

Abagyan:1999:IFP

[AT99] Ruben A. Abagyan and Maxim Totrov. Ab initio folding of peptides by the optimal-bias Monte Carlo minimization procedure. Journal of Computational Physics, 151(1):402–421, May 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199962332.

Alvarez:1999:NAM

[AWY99] Alberto Alvarez, Cheng-Ching Wang, and Zhen Ye. A numerical algorithm of the multiple scattering from an ensemble of arbitrary scatterers. Journal of Computational Physics, 154(1):231–236, September 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999199963076.

Azmy:1999:ICA

[Azm99] Y. Y. Azmy. Iterative convergence acceleration of neutral particle transport methods via adjacent-cell preconditioners. Journal of Computational Physics, 152(1):359–384, June 10, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962514.

Ben-Artzi:1990:CRD

[BAB90] Matania Ben-Artzi and Amnon Birman. Computation of reactive duct flows in external fields. Journal of Computational Physics, 86(1):225–255, January 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090099M.

Baines:1997:IAR

[Bai97] M. J. Baines. Introduction to “Approximate Riemann Solvers, Parameter Vectors, and Difference Schemes”. Journal of Computational Physics, 135(2):249, August 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957284.

Balashazy:1994:SPT

[Bal94a] Imre Balásházy. Simulation of particle trajectories in bifurcating tubes. Journal of Computational Physics, 110(1):11–22, January 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999184710023.

Balsara:1994:RSR

[Bal94b] Dinshaw S. Balsara. Riemann solver for relativistic hydrodynamics. Journal of Computational Physics, 114(2):284–297, October 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999184711673.

Balsara:1995:NSA

[Bal95] Dinshaw S. Balsara. von Neumann stability analysis of smoothed particle hydrodynamics-suggestions for optimal algorithms. Journal of Computational Physics, 121(2):357–372, October 15, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002199919590221X.

Barnes:1990:MTC

[Bar90] Joshua E. Barnes. A modified tree code: Don’t laugh; it runs. Journal of Computational Physics, 87(1):161–170, March 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919090232P.

Barenghi:1991:CTT

[Bar91a] Carlo F. Barenghi. Computations of transitions and Taylor vortices in temporally modulated Taylor–Couette flow. Journal of Computational Physics, 95(1):175–194, July 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190258M.

Barros:1991:MMT

[Bar91b] Saulo R. M. Barros. Multigrid methods for two- and threedimensional Poisson-type equations on the sphere. Journal of Computational Physics, 92(2):313–348, February 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999191902135.

Barcza:1994:CNS

[Bar94] S. Barcza. On a class of non-separable quantum-mechanical eigenvalue problems: Analytical and technical considerations within the frame of a Born expansion method. Journal of Computational Physics, 110(2):242–256, February 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710229.

Basile:1992:RACb

[Bas92] A. G. Basile. A relaxation algorithm for classical paths as a function of ende points: Application to the semiclassical propagator for far-from-caustic and near-caustic coditions. Journal of Computational Physics, 99(2):352, April 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290224M.

Bauer:1997:HAE

[Bau97] Robert Bruce Bauer. A hybrid adaptive ENO scheme. Journal of Computational Physics, 136(1):180–196, September 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957661.

Bailey:1990:NMA

[BB90a] Harry E. Bailey and Richard M. Beam. Newton’s method applied to finite-difference approximations for the steady-state compressible Navier–Stokes equations. Journal of Computational Physics, 89(1):251, July 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919090126L.

Brieger:1990:SCA

[BB90b] Leesa M. Brieger and Ernesto Bonomi. A stochastic cellular automation model of non-linear diffusion and diffusion with reaction. Journal of Computational Physics, 90(1):269, September 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919090215M.

Bailey:1991:NMA

[BB91a] Harry E. Bailey and Richard M. Beam. Newton’s method applied to finite-difference approximations for the steadystate compressible Navier–Stokes equations. Journal of Computational Physics, 93(1):108–127, March 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190075V.

Bernstein:1991:NMS

[BB91b] Ira B. Bernstein and Leigh Brookshaw. A numerical method for solving systems of linear ordinary differential equations with rapidly oscillating solutions. Journal of Computational Physics, 94(1):251, May 1991. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919190148E.

Brieger:1991:SCA

[BB91c] Leesa M. Brieger and Ernesto Bonomi. A stochastic cellular automaton model of non-linear diffusion and diffusion with reaction. Journal of Computational Physics, 94(2):467–486, June 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999191902319.

Bakhoum:1996:GSL

[BB96] Ezzat G. Bakhoum and John A. Board, Jr. The geometric solution of Laplace’s equation. Journal of Computational Physics, 123(2):274–295, February 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900248.

Borggaard:1997:PSE

[BB97a] Jeff Borggaard and John Burns. A PDE sensitivity equation method for optimal aerodynamic design. Journal of Computational Physics, 136(2):366–384, September 15, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957430.

Boris:1997:FCT

[BB97b] Jay P. Boris and David L. Book. Flux-corrected transport. Journal of Computational Physics, 135(2):172–186, August 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957004.

Besson:1990:NME

[BBC^{+}90] O. Besson, J. Bourgeois, P.-A Chevalier, J. Rappaz, and R. Touzani. Numerical modeling of electromagnetic casting processes. Journal of Computational Physics, 89(1):253, July 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919090133L.

Besson:1991:NME

[BBC^{+}91] O. Besson, J. Bourgeois, P.-A Chevalier, J. Rappaz, and R. Touzani. Numerical modelling of electromagnetic casting processes. Journal of Computational Physics, 92(2):482–507, February 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190219B.

Bernstein:1992:NMSa

[BBF92a] Ira B. Bernstein, Leigh Brookshaw, and Peter A. Fox. A numerical method for solving systems of linear ordinary differential equations with rapidly oscillating solutions. Journal of Computational Physics, 98(1):179, January 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290191Z.

Bernstein:1992:NMSb

[BBF92b] Ira B. Bernstein, Leigh Brookshaw, and Peter A. Fox. A numerical method for solving systems of linear ordinary differential equations with rapidly oscillating solutions. Journal of Computational Physics, 98(2):269–284, February 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290142L.

Baldwin:1999:ILS

[BBF^{+}99] Chuck Baldwin, Peter N. Brown, Robert Falgout, Frank Graziani, and Jim Jones. Iterative linear solvers in a 2D radiationhydrodynamics code: Methods and performance. Journal of Computational Physics, 154(1):1–40, September 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962903.

Braess:1993:MGM

[BBGL93] Dietrich Braess, Michael Biebighäuser, Peter Grassberger, and Ricardo Leuverink. Multi-grid methods for steady state diffusion in random media. Journal of Computational Physics, 107(1): 118–123, July 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999183711290.

Boillat:1990:DLB

[BBK90] J. E. Boillat, F. Bruge, and P. G. Kropf. A dynamic load balancing algorithm for molecular dynamics simulation on multiprocessor systems. Journal of Computational Physics, 91(1):253, November 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919090029Z.

Boillat:1991:DLB

[BBK91] J. E. Boillat, F. Brugé, and P. G. Kropf. A dynamic loadbalancing algorithm for molecular dynamics simulation on multiprocessor systems. Journal of Computational Physics, 96(1):1– 14, September 1991. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919190263K.

Boomkamp:1997:CCM

[BBMB97] P. A. M. Boomkamp, B. J. Boersma, R. H. M. Miesen, and G. V. Beijnon. A Chebyshev collocation method for solving two-phase flow stability problems. Journal of Computational Physics, 132(2):191–200, April 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196955710.

Buet:1998:CED

[BC98] C. Buet and S. Cordier. Conservative and entropy decaying numerical scheme for the isotropic Fokker–Planck–Landau equation. Journal of Computational Physics, 145(1):228–245, September 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198960156.

Bennett:1999:OBC

[BC99] A. F. Bennett and B. S. Chua. Open boundary conditions for Lagrangian geophysical fluid dynamics. Journal of Computational Physics, 153(2):418–436, August 10, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962848.

Buet:1997:FAN

[BCDL97] C. Buet, S. Cordier, P. Degond, and M. Lemou. Fast algorithms for numerical, conservative, and entropy approximations of the Fokker–Planck–Landau equation. Journal of Computational Physics, 133(2):310–322, May 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956692.

Bayliss:1995:AAS

[BCM95a] Alvin Bayliss, Andreas Class, and Bernard J. Matkowsky. Adaptive approximation of solutions to problems with multiple layers by Chebyshev pseudo-spectral methods. Journal of Computational Physics, 116(1):160–172, January 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710145.

Bayliss:1995:REC

[BCM95b] Alvin Bayliss, Andreas Class, and Bernard J. Matkowsky. Roundoff error in computing derivatives using the Chebyshev differentiation matrix. Journal of Computational Physics, 116(2):380– 383, February 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999185710364.

Beylkin:1999:FSP

[BCM99a] Gregory Beylkin, Nicholas Coult, and Martin J. Mohlenkamp. Fast spectral projection algorithms for density-matrix computations. Journal of Computational Physics, 152(1):32–54, June 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199962150.

Birgin:1999:EOC

[BCM99b] Ernesto G. Birgin, Ivan Chambouleyron, and José Mario Martínez. Estimation of the optical constants and the thickness of thin films using unconstrained optimization. Journal of Computational Physics, 151(2):862–880, May 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962241.

Budd:1999:NSS

[BCR99] Chris J. Budd, Shaohua Chen, and Robert D. Russell. New self-similar solutions of the nonlinear Schrödinger equation with moving mesh computations. Journal of Computational Physics, 152(2):756–789, July 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962629.

Baglama:1998:CFS

[BCRR98] J. Baglama, D. Calvetti, L. Reichel, and A. Ruttan. Computation of a few small eigenvalues of a large matrix with application to liquid crystal modeling. Journal of Computational Physics, 146(1):203–226, October 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960648.

Burtsev:1998:NAD

[BCT98] S. Burtsev, R. Camassa, and I. Timofeyev. Numerical algorithms for the direct spectral transform with applications to nonlinear Schrödinger type systems. Journal of Computational Physics, 147(1):166–186, November 20, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960879.

Beckers:1993:SFS

[BD93] Jean-Marie Beckers and Eric Deleersnijder. Stability of a FBTCS scheme applied to the propagation of shallow-water inertia-gravity waves on various space grids. Journal of Computational Physics, 108(1):95–104, September 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711666.

Berghezan:1994:NSS

[BD94a] D. Berghezan and F. Dupret. Numerical simulation of stratified coating flow by a variational method. Journal of Computational Physics, 111(1):165–182, March 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710527.

Bhattacharyya:1994:ENA

[BD94b] Satadal Bhattacharyya and Tapan Kumar Das. An efficient numerical algorithm for calculation of matrix elements of trinucleon system with realistic potentials. Journal of Computational Physics, 114(2):308–311, October 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711697.

Banas:1996:ECA

[BD96] K. Banaś and L. Demkowicz. Entropy controlled adaptive finite element simulations for compressible gas flow. Journal of Computational Physics, 126(1):181–201, June 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901291.

Benamou:1997:DDM

[BD97a] Jean-David Benamou and Bruno Desprès. A domain decomposition method for the Helmholtz equation and related optimal control problems. Journal of Computational Physics, 136(1):68–82, September 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999197957429.

Bertagnolio:1997:SDC

[BD97b] F. Bertagnolio and O. Daube. Solution of the div-curl problem in generalized curvilinear coordinates. Journal of Computational Physics, 138(1):121–152, November 15, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958009.

Bezard:1999:ESI

[BD99] Fabienne Bezard and Bruno Després. An entropic solver for ideal Lagrangian magnetohydrodynamics. Journal of Computational Physics, 154(1):65–89, September 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963003.

Belova:1997:HSE

[BDC97] E. V. Belova, R. E. Denton, and A. A. Chan. Hybrid simulations of the effects of energetic particles on low-frequency MHD waves. Journal of Computational Physics, 136(2):324–336, September 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197957387.

Byers:1994:NIG

[BDML94] J. A. Byers, A. M. Dimits, Y. Matsuda, and A. B. Langdon. Numerical instability in a 2D gyrokinetic code caused by divergent E × B flow. Journal of Computational Physics, 115(2):352–365, December 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184712010.

Bramley:1991:NSE

[BDR91] S. Bramley, L. Dieci, and R. D. Russell. Numerical solution of eigenvalue problems for linear boundary value ODEs. Journal of Computational Physics, 94(2):382–402, June 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190226B.

Beckermann:1999:MMC

[BDS^{+}99] C. Beckermann, H.-J Diepers, I. Steinbach, A. Karma, and X. Tong. Modeling melt convection in phase-field simulations of solidification. Journal of Computational Physics, 154(2):468–496, September 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999199963234.

Breuer:1991:EIC

[BE91] Kenneth S. Breuer and Richard M. Everson. On the errors incurred calculating derivatives using Chebyshev polynomials. Journal of Computational Physics, 94(1):250–251, May 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190147D.

Breuer:1992:EICa

[BE92a] Kenneth S. Breuer and Richard M. Everson. On the errors incurred calculating derivatives using Chebyshev polynomials. Journal of Computational Physics, 98(2):347, February 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290149S.

Breuer:1992:EICb

[BE92b] Kenneth S. Breuer and Richard M. Everson. On the errors incurred calculating derivatives using Chebyshev polynomials. Journal of Computational Physics, 99(1):56–67, March 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192902743.

Beckers:1994:DIF

[Bec94] Jean-Marie Beckers. On destabilizing implicit factors in discrete advection-diffusion equations. Journal of Computational Physics, 111(2):260–265, April 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710618.

Bellan:1997:ENM

[Bel97]  P. M. Bellan.  An effective numerical method for lin-

ear mode conversion problems. Journal of Computational Physics, 136(2):654–659, September 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957922.

Benson:1990:EFD

[Ben90] Alvin K. Benson. An explicit finite-difference solution to the wave equation with variable velocity. Journal of Computational Physics, 87(1):47–60, March 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919090224O.

Benson:1992:ASAa

[Ben92a] Alvin K. Benson. An anti-spatial-aliasing filter for explicit modeling and imaging in inhomogeneous media. Journal of Computational Physics, 98(2):348, February 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290157T.

Benson:1992:ASAb

[Ben92b] Alvin K. Benson. An anti-spatial-aliasing filter for explicit modeling and imaging in inhomogeneous media. Journal of Computational Physics, 99(2):183–189, April 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192901999.

Benson:1992:MASb

[Ben92c] David J. Benson. Momentum advection on a staggered mesh. Journal of Computational Physics, 99(1):181, March 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290299E.

Benson:1992:MASa

[Ben92d] David J. Benson. Momentum advection on a staggered mesh. Journal of Computational Physics, 100(1):143–162, May 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290316Q.

Benson:1995:DIE

[Ben95] Alvin K. Benson. Decomposition and inversion of elastic reflection data: First-order angular dependence and applications. Journal of Computational Physics, 121(1):102–114, October 1, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185711825.

Benamou:1996:BRT

[Ben96] Jean-David Benamou. Big ray tracing: Multivalued travel time field computation using viscosity solutions of the eikonal equation. Journal of Computational Physics, 128(2):463–474, October 15, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196902247.

Berenger:1994:PML

[Ber94] Jean-Pierre Berenger. A perfectly matched layer for the absorption of electromagnetic waves. Journal of Computational Physics, 114(2):185–200, October 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711594.

Bernard:1995:DVS

[Ber95] Peter S. Bernard. A deterministic vortex sheet method for boundary layer flow. Journal of Computational Physics, 117(1):132– 145, March 1, 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999185710510.

Berenger:1996:TDP

[Ber96] Jean-Pierre Berenger. Three-dimensional perfectly matched layer for the absorption of electromagnetic waves. Journal of Computational Physics, 127(2):363–379, September 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901813.

Bralic:1990:AGR

[BES90] N. Bralić, R. Espinosa, and C. Saavedra. An algorithm for the generation of random numbers with density C exp(−λ|x|^{ν}). Journal of Computational Physics, 88(2):484–489, June 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190901913.

Bahar:1994:FWS

[BES94] E. Bahar and M. El-Shenawee. Full wave single and double scatter from rough surfaces. Journal of Computational Physics, 115(2):390–398, December 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184712046.

Beyer:1991:CMC

[Bey91] Richard P. Beyer, Jr. A computational model of the cochlea using the immersed boundary method. Journal of Computational Physics, 94(1):251–252, May 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190150J.

Beyer:1992:CMC

[Bey92] Richard P. Beyer, Jr. A computational model of the cochlea using the immersed boundary method. Journal of Computational Physics, 98(1):145–162, January 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192901807.

Bruge:1991:SCA

[BF91] Filippo Brugé and Sandro L. Fornili. On the systolic calculation of all-pairs interactions using transputer arrays. Journal of Computational Physics, 96(1):224–228, September 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190273N.

Birdsall:1997:CCC

[BF97] Charles K. Birdsall and Dieter Fuss. Clouds-in-clouds, cloudsin-cells physics for many-body plasma simulation. Journal of Computational Physics, 135(2):141–148, August 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957235.

Bruno:1994:FNS

[BFGG94] E. Bruno, G. M. Florio, B. Ginatempo, and E. S. Giuliano. Fast numerical solution of KKR–CPA equations: Testing new algorithms. Journal of Computational Physics, 111(2):248–255, April 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002199918471059X.

Boschitsch:1999:FAM

[BFO99] Alexander H. Boschitsch, Marcia O. Fenley, and Wilma K. Olson. A fast adaptive multipole algorithm for calculating screened Coulomb (Yukawa) interactions. Journal of Computational Physics, 151(1):212–241, May 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961769.

Backhaus:1998:ACD

[BFW98] E. Yu Backhaus, J. Fajans, and J. S. Wurtele. Application of contour dynamics to systems with cylindrical boundaries. Journal of Computational Physics, 145(1):462–468, September 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960247.

Basile:1992:RACa

[BG92] A. G. Basile and C. G. Gray. A relaxation algorithm for classical paths as a function of end points: Application to the semiclassical propagator for far-from-caustic and near-caustic conditions. Journal of Computational Physics, 101(1):80–93, July 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290044Y.

Begue:1999:TDV

[BGB99] M. L. Bégué, A. Ghizzo, and P. Bertrand. Two-dimensional Vlasov simulation of Raman scattering and plasma beatwave acceleration on parallel computers. Journal of Computational Physics, 151(2):458–478, May 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199961934.

Bakker:1990:SPC

[BGGT90] A. F. Bakker, G. H. Gilmer, M. H. Grabow, and K. Thompson. A special purpose computer for molecular dynamics calculations. Journal of Computational Physics, 90(2):313–335, October 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999190901692.

Bishop:1997:CCE

[BGH^{+}97] Nigel T. Bishop, Roberto Gómez, Paulo R. Holvorcem, Richard A. Matzner, Philippos Papadopoulos, and Jeffrey Winicour. Cauchy-characteristic evolution and waveforms. Journal of Computational Physics, 136(1):140–167, September 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957545.

Bishop:1999:CCE

[BGH^{+}99] Nigel T. Bishop, Roberto Gómez, Paulo R. Holvorcem, Richard A. Matzner, Philippos Papadopoulos, and Jeffrey Winicour. Cauchy-characteristic evolution and waveforms: Volume 136, no. 1 (1997), pages 140–167. Journal of Computational Physics, 148(1):299–301, January 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961393.

Brower:1991:AMC

[BGM91] R. C. Brower, N. A. Gross, and K. J. M. Moriarty. Accelerated Monte Carlo by embedded cluster dynamics. Journal of Computational Physics, 95(1):167–174, July 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190257L.

Bayliss:1995:APS

[BGM95] Alvin Bayliss, Marc Garbey, and Bernard J. Matkowsky. Adaptive pseudo-spectral domain decomposition and the approximation of multiple layers. Journal of Computational Physics, 119(1): 132–141, June 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999185711217.

Bristeau:1998:CMC

[BGP98] M. O. Bristeau, R. Glowinski, and J. Périaux. Controllability methods for the computation of time-periodic solutions: Application to scattering. Journal of Computational Physics, 147(2):265–292, December 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960442.

Bailey:1995:DSM

[BH95] D. W. Bailey and J. M. Higman. A discretized k-space method for charge transport in semiconductors. Journal of Computational Physics, 120(1):117–127, August 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711527.

Bao:1996:NAB

[BH96] Weizhu Bao and Houde Han. Nonlocal artificial boundary conditions for the incompressible viscous flow in a channel using spectral techniques. Journal of Computational Physics, 126(1): 52–63, June 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196901199.

Bateson:1998:GPH

[BH98] William B. Bateson and Dennis W. Hewett. Grid and particle hydrodynamics:: Beyond hydrodynamics via fluid element particlein-cell. Journal of Computational Physics, 144(2):358–378, August 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999197958241.

Broyde:1999:ECS

[BH99] S. Broyde and B. E. Hingerty. Effective computational strategies for determining structures of carcinogen-damaged DNA. Journal of Computational Physics, 151(1):313–332, May 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961721.

Brun:1999:ARS

[BHJU99] G. Brun, J.-M Hérard, D. Jeandel, and M. Uhlmann. An approximate Riemann solver for second-moment closures. Journal of Computational Physics, 151(2):990–996, May 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199961909.

Braverman:1998:FPS

[BIAV98] E. Braverman, M. Israeli, A. Averbuch, and L. Vozovoi. A fast 3D Poisson solver of arbitrary order accuracy. Journal of Computational Physics, 144(1):109–136, July 20, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960016.

Bihari:1996:MSC

[Bih96] Barna L. Bihari. Multiresolution schemes for conservation laws with viscosity. Journal of Computational Physics, 123(1):207– 225, January 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196900170.

Bilbao:1990:TDL

[Bil90] L. Bilbao. A three-dimensional Lagrangian method for fluid dynamics. Journal of Computational Physics, 91(2):361–380, December 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919090042Y.

Bishop:1995:MBL

[Bis95] David G. Bishop. Material-based limiters for the time-domain Maxwell equations with thin conducting sheets. Journal of Computational Physics, 121(1):115–128, October 1, 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711837.

Bruneau:1990:ESS

[BJ90] Charles-Henri Bruneau and Claude Jouron. An efficient scheme for solving steady incompressible Navier–Stokes equations. Journal of Computational Physics, 89(2):389–413, August 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090149U.

Bertolotti:1995:EFF

[BJ95] Fabio P. Bertolotti and Ronald D. Joslin. Effect of far-field boundary conditions on boundary-layer transition. Journal of Computational Physics, 118(2):392–395, May 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711096.

Bernard:1991:HDP

[BK91] Robert S. Bernard and Hartmut Kapitza. How to discretize the pressure gradient for curvilinear MAC grids. Journal of Computational Physics, 97(2):582, December 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190023E.

Bernard:1992:HDPa

[BK92a] Robert S. Bernard and Hartmut Kapitza. How to discretize the pressure gradient for curvilinear MAC grids. Journal of Computational Physics, 99(1):179–180, March 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290290F.

Bernard:1992:HDPb

[BK92b] Robert S. Bernard and Hartmut Kapitza. How to discretize the pressure gradient for curvilinear MAC grids. Journal of Computational Physics, 99(2):288–298, April 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290208G.

Babu:1993:DSP

[BK93] V. Babu and Seppo A. Korpela. On the direct solution of Poisson’s equation on a non-uniform grid. Journal of Computational Physics, 104(1):93–98, January 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710119.

Batcho:1994:GSE

[BK94] Paul F. Batcho and George Em Karniadakis. Generalized Stokes eigenfunctions: a new trial basis for the solution of incompressible Navier–Stokes equations. Journal of Computational Physics, 115(1):121–146, November 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918471182X.

Beylkin:1997:ANS

[BK97a] Gregory Beylkin and James M. Keiser. On the adaptive numerical solution of nonlinear partial differential equations in wavelet bases. Journal of Computational Physics, 132(2):233–259, April 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002199919695562X.

Bylander:1997:WBF

[BK97b] D. M. Bylander and Leonard Kleinman. White and Bird’s formulation of gradient-corrected exchange-correlation potentials applied to atoms. Journal of Computational Physics, 136 (2):599–602, September 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999197957843.

Borba:1999:CKS

[BK99] Duarte Borba and Wolfgang Kerner. CASTOR–K: Stability analysis of Alfvén eigenmodes in the presence of energetic ions in tokamaks. Journal of Computational Physics, 153(1):101–138, July 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199962642.

Bayliss:1990:TDA

[BKM90] A. Bayliss, R. Kuske, and B. J. Matkowsky. A twodimensional adaptive pseudo-spectral method. Journal of Computational Physics, 91(1):174–196, November 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090010X.

Beason:1991:UPI

[BKP91] J. Douglas Beason, David S. Kershaw, and Manoj K. Prasad. Using physical insight: The relativistic Compton scattering kernel for radiative transfer. Journal of Computational Physics, 95(2):497–504, August 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919190287U.

Barmin:1996:SCA

[BKP96] A. A. Barmin, A. G. Kulikovskiy, and N. V. Pogorelov. Shockcapturing approach and nonevolutionary solutions in magnetohydrodynamics. Journal of Computational Physics, 126(1):77–90, June 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196901217.

Beylkin:1998:NCT

[BKV98] Gregory Beylkin, James M. Keiser, and Lev Vozovoi. A new class of time discretization schemes for the solution of nonlinear PDEs. Journal of Computational Physics, 147(2):362–387, December 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198960934.

Brackbill:1992:CMM

[BKZ92] J. U. Brackbill, D. B. Kothe, and C. Zemach. A continuum method for modeling surface tension. Journal of Computational Physics, 100(2):335–354, June 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290240Y.

Brandt:1990:MMMa

[BL90a] A. Brandt and A. A. Lubrecht. Multilevel matrix multiplication and fast solution of integral equations. Journal of Computational Physics, 87(2):494, April 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190902642.

Brandt:1990:MMMb

[BL90b] A. Brandt and A. A. Lubrecht. Multilevel matrix multiplication and fast solution of integral equations. Journal of Computational Physics, 90(2):348–370, October 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090171V.

Blayo:1992:PCM

[BL92] E. Blayo and C. Le Provost. Performance of the capacitance matrix method for solving Helmholtz type equations in ocean modelling. Journal of Computational Physics, 102(2):424, October 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290388F.

Blayo:1993:PCM

[BL93] E. Blayo and C. LeProvost. Performance of the capacitance matrix method for solving Helmhotz-type equations in ocean modelling. Journal of Computational Physics, 104(2):347–360, February 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183710363.

Brackbill:1994:MSF

[BL94] J. W. Brackbill and Giovanni Lapenta. A method to suppress the finite-grid instability in plasma simulations. Journal of Computational Physics, 114(1):77–84, September 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711508.

Brieger:1998:PMP

[BL98] Leesa Brieger and Giuditta Lecca. Parallel multigrid preconditioning of the conjugate gradient method for systems of subsurface hydrology. Journal of Computational Physics, 142(1):148– 162, May 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198959162.

Bae:1999:UNF

[BL99] Song-Hyo Bae and Richard T. Lahey, Jr. On the use of nonlinear filtering, artificial viscosity, and artificial heat transfer for strong shock computations. Journal of Computational Physics, 153(2):575–595, August 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962964.

Borgers:1992:ADLa

[BLA92a] Christoph Borgers, Edward W. Larsen, and Marvin L. Adams. The asymptotic diffusion limit of a linear discontinuous discretization of a two-dimensional linear transport equation. Journal of Computational Physics, 98(1):179, January 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192901922.

Borgers:1992:ADLb

[BLA92b] Christoph Börgers, Edward W. Larsen, and Marvin L. Adams. The asymptotic diffusion limit of a linear discontinuous discretization of a two-dimensional linear transport equation. Journal of Computational Physics, 98(2):285–300, February 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290143M.

Blackbill:1995:E

[Bla95] Jerry Blackbill. Editorial. Journal of Computational Physics, 116(1):1–2, January 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999185710017.

Batten:1997:ASJ

[BLG97] P. Batten, M. A. Leschziner, and U. C. Goldberg. Averagestate Jacobians and implicit methods for compressible viscous and turbulent flows. Journal of Computational Physics, 137(1): 38–78, October 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999197957934.

Barten:1990:CBM

[BLK90] W. Barten, M. Lücke, and M. Kamps. Conservation and breaking of mirror symmetry in a numerical simulation of vortex flow. Journal of Computational Physics, 91(2):486–489, December 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919090050B.

Bond:1999:NPM

[BLL99] Stephen D. Bond, Benedict J. Leimkuhler, and Brian B. Laird. The Nosé–Poincaré method for constant temperature molecular dynamics. Journal of Computational Physics, 151(1):114– 134, May 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199919896171X.

Blum:1992:CCSa

[BLMR92] V. Blum, G. Lauritsch, J. A. Maruhn, and P.-G Reinhard. Comparison of coordinate-space techniques in nuclear mean-field calculations. Journal of Computational Physics, 100(2):364–376, June 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290242Q.

Brady:1998:RVS

[BLP98] M. Brady, A. Leonard, and D. I. Pullin. Regularized vortex sheet evolution in three dimensions. Journal of Computational Physics, 146(2):520–545, November 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959988.

Bourgat:1996:CBN

[BLT96] Jean-François Bourgat, Patrick Le Tallec, and Moulay Tidriri. Coupling Boltzmann and Navier–Stokes equations by friction. Journal of Computational Physics, 127(2):227–245, September 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196901722.

Blum:1992:CCSb

[Blu92] V. Blum. Comparison of coordinated-space techniques in nuclear mean-field calculations. Journal of Computational Physics, 99 (2):351, April 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290218N.

Bonazzola:1990:TDG

[BM90a] Silvano Bonazzola and Jean-Alain Marck. Three-dimensional gas dynamics in a sphere. Journal of Computational Physics, 87(1):201–230, March 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919090234R.

Bouaoudia:1990:FAS

[BM90b] S. Bouaoudia and P. S. Marcus. Fast and accurate spectral treatment of coordinate singularities. Journal of Computational Physics, 91(1):254, November 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090034X.

Bonazzola:1991:ETSa

[BM91a] Silvano Bonazzola and Jean-Alain Marck. A 1D exact treatment of shock waves within spectral methods in plane geometry. Journal of Computational Physics, 93(2):486, April 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190199U.

Bonazzola:1991:ETSb

[BM91b] Silvano Bonazzola and Jean-Alain Marck. A 1D exact treatment of shock waves within spectral methods in plane geometry. Journal of Computational Physics, 97(2):535–552, December 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190012A.

Bouaoudia:1991:FAS

[BM91c] S. Bouaoudia and P. S. Marcus. Fast and accurate spectral treatment of coordinate singularities. Journal of Computational Physics, 96(1):217–223, September 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190272M.

Bell:1992:SOP

[BM92] John B. Bell and Daniel L. Marcus. A second-order projection method for variable-density flows. Journal of Computational Physics, 101(2):334–348, August 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290011M.

Balachandar:1994:SMS

[BM94] S. Balachandar and Ravi K. Madabhushi. Spurious modes in spectral collocation methods with two non-periodic directions. Journal of Computational Physics, 113(1):151–153, July 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999184711259.

Baras:1995:PSC

[BMGM95] F. Baras, M. Malek Mansour, A. L. Garcia, and M. Mareschal. Particle simulation of complex flows in dilute systems. Journal of Computational Physics, 119(1):94–104, June 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711187.

Bertagnolli:1997:TSS

[BMJ^{+}97] Mauro Bertagnolli, Maurizio Marchese, Gianni Jacucci, Ioannis St. Doltsinis, and Swen Noelting. Thermomechanical simulation of the splashing of ceramic droplets on a rigid substrate. Journal of Computational Physics, 133(2):205–221, May 15, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196954686.

Bielak:1995:SCM

[BMZ95] Jacobo Bielak, Richard C. Maccamy, and Xiaogang Zeng. Stable coupling method for interface scattering problems by combined integral equations and finite elements. Journal of Computational Physics, 119(2):374–384, July 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711412.

Bolton:1994:EGR

[BN94] Kim Bolton and Sture Nordholm. An evaluation of the Gauss– Radau algorithm for the simulation of chemical dynamics. Journal of Computational Physics, 113(2):320–335, August 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711399.

Briley:1996:ILU

[BNW96] W. Roger Briley, Shyam S. Neerarambam, and David L. Whitfield. Implicit lower-upper/approximate-factorization schemes for incompressible flows. Journal of Computational Physics, 128(1):32–42, October 1, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999196901941.

Boffetta:1992:CDS

[BO92] G. Boffetta and A. R. Osborne. Computation of the direct scattering transform for the nonlinear Schrödinger equation. Journal of Computational Physics, 102(2):252–264, October 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290370E.

Birk:1996:TDP

[BO96] G. T. Birk and A. Otto. A three-dimensional plasma-neutral gas-fluid code. Journal of Computational Physics, 125(2):513– 525, May 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196901126.

Borges:1998:PDT

[BO98] Leonardo Borges and Suely Oliveira. A parallel Davidsontype algorithm for several eigenvalues. Journal of Computational Physics, 144(2):727–748, August 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919896003X.

Board:1997:IFA

[Boa97] John A. Board, Jr. Introduction to “A Fast Algorithm for Particle Simulations”. Journal of Computational Physics, 135(2): 279, August 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999197957624.

Borgers:1991:ADL

[Bor91a] C. Borgers. The asymptotic diffusion limit of a linear discontinuous discretization of a two-dimensional linear transport equation. Journal of Computational Physics, 94(1):254, May 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190158H.

Borne:1991:HSF

[Bor91b] Lionel Borne. Harmonic Stokes flaw through periodic porous media. A 3D boundary element method. Journal of Computational Physics, 97(2):582–583, December 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190026H.

Borne:1992:HSFa

[Bor92a] Lionel Borne. Harmonic Stokes flow through periodic porous media: a 3D boundary element method. Journal of Computational Physics, 99(1):179, March 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999192902868.

Borne:1992:HSFb

[Bor92b] Lionel Borne. Harmonic Stokes flow through periodic porous media: a 3D boundary element method. Journal of Computational Physics, 99(2):214–232, April 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290204C.

Borgers:1997:FIM

[Bör97] Christoph Börgers. A fast iterative method for computing particle beams penetrating matter. Journal of Computational Physics, 133(2):323–339, May 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956710.

Bottura:1996:NMS

[Bot96] L. Bottura. A numerical model for the simulation of quench in the ITER magnets. Journal of Computational Physics, 125(1): 26–41, April 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196900777.

Bottino:1998:MVN

[Bot98] Dean C. Bottino. Modeling viscoelastic networks and cell deformation in the context of the immersed boundary method.

Journal of Computational Physics, 147(1):86–113, November 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198960740.

Boyd:1990:VMC

[Boy90] Iain D. Boyd. Vectorization of a Monte Carlo simulation scheme for nonequilibrium gas dynamics. Journal of Computational Physics, 91(2):495, December 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190900534.

Boyd:1991:VMC

[Boy91a] Iain D. Boyd. Vectorization of a Monte Carlo simulation scheme for nonequilibrium gas dynamics. Journal of Computational Physics, 96(2):411–427, October 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190243E.

Boyd:1991:ATK

[Boy91b] John P. Boyd. The arctan/tan and Kepler–Burger mappings for periodic solutions with a shock, internal boundary layer. Journal of Computational Physics, 93(1):250, March 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190082V.

Boyd:1992:ATK

[Boy92a] John P. Boyd. The arctan/tan and Kepler–Burgers mappings for periodic solutions with a shock, front, or internal boundary layer. Journal of Computational Physics, 98(2):181–193, February 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290136M.

Boyd:1992:FACa

[Boy92b] John P. Boyd. A fast algorithm for Chebyshev, Fourier, and sinc interpolation onto an irregular grid. Journal of Computational Physics, 101(1):228, July 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192900668.

Boyd:1992:FACb

[Boy92c] John P. Boyd. A fast algorithms for Chebyshev, Fourier and sinc interpolation onto an irregular grid. Journal of Computational Physics, 103(2):243–257, December 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290399J.

Boyd:1992:MEP

[Boy92d] John P. Boyd. Multipole expansions and pseudospectral cardinal functions: A new generalization of the Fast Fourier Transform. Journal of Computational Physics, 103(1):184–186, November 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290333T.

Boyd:1994:RCF

[Boy94] John P. Boyd. The rate of convergence of Fourier coefficients for entire functions of infinite order with application to the Weideman–Cloot sinh-mapping for pseudospectral computations on an infinite interval. Journal of Computational Physics, 110(2):360–372, February 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710321.

Boyd:1995:CPI

[Boy95a] John P. Boyd. A Chebyshev polynomial interval-searching method (“Lanczos economization”) for solving a nonlinear equation with application to the nonlinear eigenvalue problem. Journal of Computational Physics, 118(1):1–8, April 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710753.

Boyd:1995:HPM

[Boy95b] John P. Boyd. A hyperasymptotic perturbative method for computing the radiation coefficient for weakly nonlocal solitary waves. Journal of Computational Physics, 120(1):15–32, August 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185711461.

Boyd:1996:NCN

[Boy96a] John P. Boyd. Numerical computations of a nearly singular nonlinear equation: Weakly nonlocal bound states of solitons for the fifth-order Korteweg–de Vries equation. Journal of Computational Physics, 124(1):55–70, March 1, 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900443.

Boyd:1996:TSE

[Boy96b] John P. Boyd. Traps and snares in eigenvalue calculations with application to pseudospectral computations of ocean tides in a basin bounded by meridians. Journal of Computational Physics, 126(1):11–20, June 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0021999196901163. See correction [Boy97a].

Boyd:1997:CPT

[Boy97a] John P. Boyd. A correction to the paper “Traps and Snares in Eigenvalue Calculations with Application to Pseudospectral Computations of Ocean Tides in a Basin Bounded by Meridians”. Journal of Computational Physics, 136(1):227–228, September 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197957533. See [Boy96b].

Boyd:1997:PDF

[Boy97b] John P. Boyd. Pseudospectral/Delves–Freeman computations of the radiation coefficient for weakly nonlocal solitary waves of the third-order nonlinear Schrödinger equation and their relation to hyperasymptotic perturbation theory. Journal of Computational Physics, 138(2):665–694, December 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919795840X.

Boyd:1998:TCF

[Boy98] John P. Boyd. Two comments on filtering (artificial viscosity) for Chebyshev and Legendre spectral and spectral element methods: Preserving boundary conditions and interpretation of the filter as a diffusion. Journal of Computational Physics, 143(1):283– 288, June 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198959617.

Baden:1990:FVM

[BP90] Scott B. Baden and Elbridge Gerry Puckett. A fast vortex method for computing 2D viscous flow. Journal of Computational Physics, 91(2):278–297, December 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190900383.

Brydon:1998:SSD

[BPM98] David Brydon, John Pearson, and Michael Marder. Solving stiff differential equations with the method of patches. Journal of Computational Physics, 144(2):280–298, August 10, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960089.

Bakker:1995:SFG

[BPS95] B. L. G. Bakker, M. I. Polikarpov, and L. V. Shevchenko. Scattering in the framework of the Green’s function Monte–Carlo method. Journal of Computational Physics, 116(2):277–280, February 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999185710261.

Baum:1995:ABC

[BPT95] M. Baum, T. Poinsot, and D. Thévenin. Accurate boundary conditions for multicomponent reactive flows. Journal of Computational Physics, 116(2):247–261, February 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710248.

Blunt:1992:IFLa

[BR92a] Martin Blunt and Barry Rubin. Implicit flux limiting schemes for petroleum reservoir simulation. Journal of Computational Physics, 101(1):227, July 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192900587.

Blunt:1992:IFLb

[BR92b] Martin Blunt and Barry Rubin. Implicit flux limiting schemes for petroleum reservoir simulation. Journal of Computational Physics, 102(1):194–210, September 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999105800154.

Broeze:1992:ABCa

[BR92c] J. Broeze and J. E. Romate. Absorbing boundary conditions for free surface wave simulations with a panel method. Journal of Computational Physics, 98(2):348, February 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290154Q.

Broeze:1992:ABCb

[BR92d] J. Broeze and J. E. Romate. Absorbing boundary conditions for free surface wave simulations with a panel method. Journal of Computational Physics, 99(1):146–158, March 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192902824.

Black:1994:ACN

[BR94] D. W. Black and A. P. Rothmayer. On accelerating the convergence of nonlinear iterative algorithms. Journal of Computational Physics, 111(2):324–335, April 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710667.

Bassi:1997:HOAa

[BR97a] F. Bassi and S. Rebay. A high-order accurate discontinuous finite element method for the numerical solution of the compressible Navier–Stokes equations. Journal of Computational Physics, 131(2):267–279, March 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196955722.

Bassi:1997:HOAb

[BR97b] F. Bassi and S. Rebay. High-order accurate discontinuous finite element solution of the 2D Euler equations. Journal of Computational Physics, 138(2):251–285, December 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197954541.

Balakrishnan:1999:PST

[BR99] Karthik Balakrishnan and Palghat A. Ramachandran. A particular solution Trefftz method for non-linear Poisson problems in heat and mass transfer. Journal of Computational Physics, 150(1):239–267, March 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961782.

Brackbill:1990:FMP

[Bra90] J. U. Brackbill. FLIP MHD: a particle-in-cell method for magnetohydrodynamics. Journal of Computational Physics, 91(2):497, December 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999190900589.

Brackbill:1991:FMP

[Bra91] J. U. Brackbill. FLIP MHD: a particle-in-cell method for magnetohydrodynamics. Journal of Computational Physics, 96(1):163– 192, September 1991. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919190270U.

Brandt:1992:MSHb

[Bra92] A. Brandt. On multigrid solution of high-Reynolds incompressible entering flows. Journal of Computational Physics, 99(2): 352, April 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290226O.

Brackbill:1993:AGD

[Bra93] J. U. Brackbill. An adaptive grid with directional control. Journal of Computational Physics, 108(1):38–50, September 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711617.

Brackbill:1996:Ea

[Bra96a] J. U. Brackbill. Editorial. Journal of Computational Physics, 126(1):v, June 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196901345.

Brackbill:1996:Eb

[Bra96b] J. U. Brackbill. Editorial. Journal of Computational Physics, 127 (1):1, August 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196901539.

Brackbill:1999:A

[Bra99] J. U. Brackbill. ANNOUNCEMENT. Journal of Computational Physics, 150(2):301, April 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962526.

Brizuela:1995:ANB

[Bri95] Edward Brizuela. Accurate normalisation of the beta-function PDF. Journal of Computational Physics, 119(2):385–387, July 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185711424.

Barnett:1991:MCA

[BRL91] R. N. Barnett, P. J. Reynolds, and W. A. Lester, Jr. Monte Carlo algorithms for expectation values of coordinate operators. Journal of Computational Physics, 96(2):258–276, October 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190236E.

Brown:1995:PUR

[Bro95] David L. Brown. Performance of under-resolved two-dimensional incompressible flow simulations. Journal of Computational Physics, 122(1):165–183, November 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0021999185712053.

Bagnoli:1992:GAT

[BRR92] Franco Bagnoli, Raúl Rechtman, and Stefano Ruffo. General algorithm for two-dimensional totalistic cellular automata. Journal of Computational Physics, 101(1):176–184, July 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290051Y.

Bruge:1992:SCP

[Bru92] F. Brugè. Systolic calculation of pair interactions using the cell linked-lists method on multi-processor systems. Journal of Computational Physics, 101(2):454, August 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290036X.

Bruge:1993:SCP

[Bru93] F. Brugé. Systolic calculation of pair interactions using the cell linked-lists method on multi-processor systems. Journal of Computational Physics, 104(1):263–266, January 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710259.

Bryan:1997:NMS

[Bry97] Kirk Bryan. A numerical method for the study of the circulation of the world ocean. Journal of Computational Physics, 135(2):154–169, August 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999197956990.

Bratkovsky:1990:CCC

[BS90] A. M. Bratkovsky and S. Yu. Savrasov. On the calculation of combined corrections in the LMTO method. Journal of Computational Physics, 88(1):243–249, May 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190902505.

Breuer:1991:UKL

[BS91a] Kenneth S. Breuer and Lawrence Sirovich. The use of the Karhunen–Loève procedure for the calculation of linear eigenfunctions. Journal of Computational Physics, 96(2):277–296, October 1991. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919190237F.

Butler:1991:ESG

[BS91b] Roger A. R. Butler and Edward E. Slaminka. An evaluation of the SNIFFER global optimization algorithm using standard test functions. Journal of Computational Physics, 97(2):580, December 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190019H.

Butler:1992:ESGa

[BS92a] Roger A. R. Butler and Edward E. Slaminka. An evaluation of the sniffer global optimization algorithm using standard test functions. Journal of Computational Physics, 98(1):179, January 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999192901885.

Butler:1992:ESGb

[BS92b] Roger A. R. Butler and Edward E. Slaminka. An evaluation of the sniffer global optimization algorithm using standard test functions. Journal of Computational Physics, 99(1):28–32, March 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290271Y.

Biesiadecki:1993:DMT

[BS93] Jeffrey J. Biesiadecki and Robert D. Skeel. Dangers of multiple time step methods. Journal of Computational Physics, 109(2):318–328, December 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999183712209.

Blosch:1996:SPD

[BS96] Edwin L. Blosch and Wei Shyy. Scalability and performance of data-parallel pressure-based multigrid methods for viscous flows. Journal of Computational Physics, 125(2):338–353, May 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196900984.

Barth:1998:NSH

[BS98a] Timothy J. Barth and James A. Sethian. Numerical schemes for the Hamilton–Jacobi and level set equations on triangulated domains. Journal of Computational Physics, 145(1):1–40, September 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198960077.

Brau:1998:TDF

[BS98b] F. Brau and C. Semay. The three-dimensional Fourier grid Hamiltonian method. Journal of Computational Physics, 139(1):127– 136, January 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999197958666.

Balsara:1999:MPP

[BS99a] Dinshaw S. Balsara and Daniel Spicer. Maintaining pressure positivity in magnetohydrodynamic simulations. Journal of Computational Physics, 148(1):133–148, January 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961083.

Balsara:1999:SMAa

[BS99b] Dinshaw S. Balsara and Daniel S. Spicer. A staggered mesh algorithm using high order Godunov fluxes to ensure solenoidal magnetic fields in magnetohydrodynamic simulations. Journal of Computational Physics, 149(2):270–292, March 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961538.

Balsara:1999:SMAb

[BS99c] Dinshaw S. Balsara and Daniel S. Spicer. A staggered mesh algorithm using high order Godunov fluxes to ensure solenoidal magnetic fields in magnetohydrodynamic simulations: Volume 149, number 1 (1999), pages 270–292. Journal of Computational Physics, 153(2):671, August 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963180.

Beichl:1999:API

[BS99d] Isabel Beichl and Francis Sullivan. Approximating the permanent via importance sampling with application to the dimer covering problem. Journal of Computational Physics, 149(1):128–147, February 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198961496.

Bennett:1999:LRR

[BS99e] Beth Anne V. Bennett and Mitchell D. Smooke. Local rectangular refinement with application to nonreacting and reacting fluid flow problems. Journal of Computational Physics, 151(2):684– 727, May 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999199962149.

Bihari:1999:MSR

[BS99f] Barna L. Bihari and Donald Schwendeman. Multiresolution schemes for the reactive Euler equations. Journal of Computational Physics, 154(1):197–230, September 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919996312X.

Burgess:1992:MMF

[BSB92] D. Burgess, D. Sulsky, and J. U. Brackbill. Mass matrix formulation of the FLIP particle-in-cell method. Journal of Computational Physics, 103(1):1–15, November 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290323Q.

Braun:1996:ECL

[BSPL96] M. Braun, S. A. Sofianos, D. G. Papageorgiou, and I. E. Lagaris. An efficient Chebyshev–Lanczos method for obtaining eigensolutions of the Schrödinger equation on a grid. Journal of Computational Physics, 126(2):315–327, July 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901400.

Baker:1995:WPN

[BST95] Gregory Baker, Michael Siegel, and Saleh Tanveer. A well-posed numerical method to track isolated conformal map singularities in Hele–Shaw flow. Journal of Computational Physics, 120(2):348– 364, September 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999185711709.

Bayliss:1992:MACa

[BT92a] Alvin Bayliss and Eli Turkel. Mappings and accuracy for Chebyshev pseudo-spectral approximations. Journal of Computational Physics, 101(1):227, July 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192900598.

Bayliss:1992:MACb

[BT92b] Alvin Bayliss and Eli Turkel. Mappings and accuracy for Chebyshev pseudo-spectral approximations. Journal of Computational Physics, 101(2):349–359, August 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290012N.

Billet:1997:ASE

[BT97a] S. J. Billet and E. F. Toro. On the accuracy and stability of explicit schemes for multidimensional linear homogeneous advection equations. Journal of Computational Physics, 131(1):247– 250, February 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196956107.

Billett:1997:WTS

[BT97b] S. J. Billett and E. F. Toro. On WAF-type schemes for multidimensional hyperbolic conservation laws. Journal of Computational Physics, 130(1):1–24, January 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196954704.

Beddhu:1996:SCF

[BTW96] Murali Beddhu, Lafayette K. Taylor, and David L. Whitfield. Strong conservative form of the incompressible Navier– Stokes equations in a rotating frame with a solution procedure. Journal of Computational Physics, 128(2):427–437, October 15, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196902211.

Buell:1991:HNM

[Bue91] Jeffrey C. Buell. A hybrid numerical method for threedimensional spatially-developing free-shear flows. Journal of Computational Physics, 95(2):313–338, August 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190279T.

Buttke:1990:FAV

[But90a] Thomas F. Buttke. A fast adaptive vortex method for patches of constant vorticity in two dimensions. Journal of Computational Physics, 89(1):161–186, July 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090121G.

Buttke:1990:VTS

[But90b] Thomas F. Buttke. Vortex tangle simulations. Journal of Computational Physics, 87(1):241–244, March 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090237U.

Buttke:1991:FAV

[But91] Thomas F. Buttke. The fast adaptive vortex method. Journal of Computational Physics, 93(2):485, April 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190198T.

Budd:1990:NSS

[BW90] C. J. Budd and A. A. Wheeler. A numerical scheme for the solution of the space charge problem on a multiply connected region. Journal of Computational Physics, 91(1):251, November 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919090023T.

Budd:1991:NSS

[BW91] C. J. Budd and A. A. Wheeler. A numerical scheme for the solution of the space charge problem on a multiply connected region. Journal of Computational Physics, 97(1):1–29, November 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190036K.

Bronsard:1995:NMT

[BW95] Lia Bronsard and Brian T. R. Wetton. A numerical method for tracking curve networks moving with curvature motion. Journal of Computational Physics, 120(1):66–87, August 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711497.

Bijl:1998:UMC

[BW98] Hester Bijl and Pieter Wesseling. A unified method for computing incompressible and compressible flows in boundary-fitted coordinates. Journal of Computational Physics, 141(2):153–173, April 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198959149.

Boing:1990:CES

[BWJ90] H. Boing, K. Werner, and H. Jackisch. Construction of the entropy solution of hyperbolic conservation laws by a geometrical interpretation of the conservation principle. Journal of Computational Physics, 91(1):250, November 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090019W.

Brandt:1990:IFO

[BY90] A. Brandt and Irad Yavneh. Inadequacy of first-order upwind difference schemes for some recirculating flows. Journal of Computational Physics, 89(1):251, July 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090127M.

Brandt:1991:IFO

[BY91] A. Brandt and I. Yavneh. Inadequacy of first-order upwind difference schemes for some recirculating flows. Journal of Computational Physics, 93(1):128–143, March 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190076W.

Brandt:1992:MSHa

[BY92] A. Brandt and I. Yavneh. On multigrid solution of high-Reynolds incompressible entering flows. Journal of Computational Physics, 101(1):151–164, July 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192900495.

Balachandar:1994:TDF

[BY94] S. Balachandar and D. A. Yuen. Three-dimensional fully spectral numerical method for mantle convection with depth-dependent properties. Journal of Computational Physics, 113(1):62–74, July 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999184711181.

Bar-Yoseph:1995:STS

[BYMZY95] Pinhas Bar-Yoseph, Eduard Moses, Uzi Zrahia, and Alexander L. Yarin. Space–time spectral element methods for onedimensional nonlinear advection-diffusion problems. Journal of Computational Physics, 119(1):62–74, June 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711163.

Casper:1993:FVH

[CA93] Jay Casper and H. L. Atkins. A finite-volume high-order ENO scheme for two-dimensional hyperbolic systems. Journal of Computational Physics, 106(1):62–76, May 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710910.

Cai:1993:PEC

[CAA93] DongSheng Cai, Akira Aoyagi, and Kanji Abe. Parametric excitation of computational mode of the leapfrog scheme applied to the van der Pol equation. Journal of Computational Physics, 107(1): 146–151, July 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999183711319.

Cally:1990:ICM

[Cal90] P. S. Cally. An inverse coordinate multigrid method for free boundary magnetohydrostatics. Journal of Computational Physics, 89(2):490, August 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919090164V.

Cally:1991:ICM

[Cal91] P. S. Cally. An inverse coordinate multigrid method for free boundary magnetohydrostatics. Journal of Computational Physics, 93(2):411–425, April 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919190192N.

Candela:1995:NMN

[Can95] D. Candela. Numerical methods for nonlinear Fourier analysis, prediction, and filtering. Journal of Computational Physics, 117(2):205–214, March 15, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999185710595.

Candy:1996:NMS

[Can96] J. Candy. A numerical method for solution of the generalized Liouville equation. Journal of Computational Physics, 129(1):160– 169, November 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196902405.

Caramana:1991:DID

[Car91] Edward J. Caramana. Derivation of implicit difference schemes by the method of differential approximation. Journal of Computational Physics, 96(2):484–493, October 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190248J.

Casulli:1990:SIF

[Cas90] Vincenzo Casulli. Semi-implicit finite difference methods for the two-dimensional shallow water equations. Journal of Computational Physics, 86(1):56–74, January 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090091E.

Crutchfield:1994:ISM

[CB94] William Y. Crutchfield and John B. Bell. Instabilities of the Skyrme model. Journal of Computational Physics, 110(2):234– 241, February 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184710217.

Chen:1996:SEA

[CB96] Gang Chen and Iain D. Boyd. Statistical error analysis for the direct simulation Monte Carlo technique. Journal of Computational Physics, 126(2):434–448, July 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901485.

Chou:1997:KFV

[CB97] S. Y. Chou and D. Baganoff. Kinetic flux-vector splitting for the Navier–Stokes equations. Journal of Computational Physics, 130(2):217–230, January 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196955795.

Carlson:1995:DNS

[CBL95] H. A. Carlson, G. Berkooz, and J. L. Lumley. Direct numerical simulation of flow in a channel with complex, timedependent Wall geometries: a pseudospectral method. Journal of Computational Physics, 121(1):155–175, October 1, 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711862.

Chen:1993:ATS

[CBS93] K. Chen, M. J. Baines, and P. K. Sweby. On an adaptive time stepping strategy for solving nonlinear diffusion equations. Journal of Computational Physics, 105(2):324–332, April 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710788.

Caramana:1998:CCH

[CBSW98] E. J. Caramana, D. E. Burton, M. J. Shashkov, and P. P. Whalen. The construction of compatible hydrodynamics algorithms utilizing conservation of total energy. Journal of Computational Physics, 146(1):227–262, October 10, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960296.

Chen:1997:EAS

[CC97] X. Chen and Y. M. Chen. Efficient algorithm for solving inverse source problems of a nonlinear diffusion equation in microwave heating. Journal of Computational Physics, 132(2):374– 383, April 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196956454.

Chumakov:1998:MDQ

[CC98] G. A. Chumakov and S. G. Chumakov. A method for the 2-D quasi-isometric regular grid generation. Journal of Computational Physics, 143(1):1–28, June 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919895968X.

Collins:1995:IEE

[CCG95] J. P. Collins, P. Colella, and H. M. Glaz. An implicit-explicit Eulerian Godunov scheme for compressible flow. Journal of Computational Physics, 116(2):195–211, February 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710212.

Chen:1995:QHS

[CCGJ95] Zhangxin Chen, Bernardo Cockburn, Carl L. Gardner, and Joseph W. Jerome. Quantum hydrodynamic simulation of hysteresis in the resonant tunneling diode. Journal of Computational Physics, 117(2):274–280, March 15, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710650.

Cary:1993:ESI

[CD93] J. R. Cary and I. Doxas. An explicit symplectic integration scheme for plasma simulations. Journal of Computational Physics, 107(1):98–104, July 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999183711277.

Chen:1995:NSE

[CD95a] K. R. Chen and J. M. Dawson. New scheme for electromagnetic simulations. Journal of Computational Physics, 118(1): 86–91, April 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999185710819.

Couzy:1995:FSC

[CD95b] W. Couzy and M. O. Deville. A fast Schur complement method for the spectral element discretization of the incompressible Navier–Stokes equations. Journal of Computational Physics, 116(1):135–142, January 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918571011X.

Costa:1998:FCS

[CD98] Bruno Costa and Lucia Dettori. Fourier collocation splittings for partial differential equations. Journal of Computational Physics, 142(2):562–580, May 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919895954X.

Carte:1995:STD

[CDF95] Gilles Carte, Jan Dusek, and Philippe Fraunié. A spectral time discretization for flows with dominant periodicity. Journal of Computational Physics, 120(2):171–183, September 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711576.

Capuzzo-Dolcetta:1998:CBF

[CDM98] R. Capuzzo-Dolcetta and P. Miocchi. A comparison between the fast multipole algorithm and the tree-code to evaluate gravitational forces in 3-D. Journal of Computational Physics, 143(1): 29–48, June 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198959496.

Chu:1999:PSA

[CDR99] King-Wai Chu, Yuefan Deng, and John Reinitz. Parallel simulated annealing by mixing of states. Journal of Computational Physics, 148(2):646–662, January 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961344.

Cortes:1998:DPM

[CDT98] J. Cortes, A. Debussche, and I. Toumi. A density perturbation method to study the eigenstructure of two-phase flow equation systems. Journal of Computational Physics, 147(2):463–484, December 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199919896096X.

Colella:1999:CFD

[CDW99a] Phillip Colella, Milo R. Dorr, and Daniel D. Wake. A conservative finite difference method for the numerical solution of plasma fluid equations. Journal of Computational Physics, 149(1):168–193, February 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198961368.

Colella:1999:NSP

[CDW99b] Phillip Colella, Milo R. Dorr, and Daniel D. Wake. Numerical solution of plasma fluid equations using locally refined grids. Journal of Computational Physics, 152(2):550–583, July 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962459.

Chen:1997:CVF

[CE97] Zhangxin Chen and Richard E. Ewing. Comparison of various formulations of three-phase flow in porous media. Journal of Computational Physics, 132(2):362–373, April 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196956417.

Coquel:1997:NMU

[CEG^{+}97] F. Coquel, K. El Amine, E. Godlewski, B. Perthame, and P. Rascle. A numerical method using upwind schemes for the resolution of two-phase flows. Journal of Computational Physics, 136(2):272–288, September 15, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957302.

Chen:1994:NSN

[CF94] C. Chen and J. M. Floryan. Numerical simulation of nonisothermal capillary interfaces. Journal of Computational Physics, 111(1):183–193, March 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999184710539.

Chu:1998:TPC

[CF98] Peter C. Chu and Chenwu Fan. A three-point combined compact difference scheme. Journal of Computational Physics, 140(2):370– 399, March 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198958995.

Chu:1999:TPS

[CF99] Peter C. Chu and Chenwu Fan. A three-point sixth-order nonuniform combined compact difference scheme. Journal of Computational Physics, 148(2):663–674, January 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961411.

Cao:1993:FCM

[CG93] Wei-Ming Cao and Ben-Yu Guo. Fourier collocation method for solving nonlinear Klein–Gordon equation. Journal of Computational Physics, 108(2):296–305, October 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711836.

Carpenter:1996:SMA

[CG96] Mark H. Carpenter and David Gottlieb. Spectral methods on arbitrary grids. Journal of Computational Physics, 129(1):74– 86, November 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199919690234X.

Cargo:1997:RMI

[CG97a] Patricia Cargo and Gérard Gallice. Roe matrices for ideal MHD and systematic construction of Roe matrices for systems of conservation laws. Journal of Computational Physics, 136 (2):446–466, September 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999197957739.

Chen:1997:DEL

[CG97b] Rongqing Chen and Hua Guo. Determination of eigenstates via Lanczos-based forward substitution and filter-diagonalization. Journal of Computational Physics, 136(2):494–502, September 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197957776.

Cheng:1997:NEE

[CG97c] Hongwei Cheng and Leslie Greengard. On the numerical evaluation of electrostatic fields in dense random dispersions of cylinders. Journal of Computational Physics, 136(2):629–639, September 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999197957879.

Carpenter:1993:SNB

[CGA93] Mark H. Carpenter, David Gottlieb, and Saul Abarbanel. The stability of numerical boundary treatments for compact high-order finite-difference schemes. Journal of Computational Physics, 108(2):272–295, October 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711824.

Carpenter:1994:TSB

[CGA94] Mark H. Carpenter, David Gottlieb, and Saul Abarbanel. Timestable boundary conditions for finite-difference schemes solving hyperbolic systems: Methodology and application to high-order compact schemes. Journal of Computational Physics, 111(2):220– 236, April 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184710576.

Chu:1992:NSS

[CGKC92] M. S. Chu, J. M. Greene, M. Klasky, and M. S. Chance. Numerical study of a singular differential equation relevant for the finite β tearing mode in a toroidal plasma. Journal of Computational Physics, 101(2):453, August 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290029X.

Chu:1993:NSS

[CGKC93] M. S. Chu, J. M. Greene, M. Klasky, and M. S. Chance. Numerical study of a singular differential equation relevant for the finite β tearing mode in a toroidal plasma. Journal of Computational Physics, 104(1):23–29, January 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918371003X.

Cheng:1999:FAM

[CGR99] H. Cheng, L. Greengard, and V. Rokhlin. A fast adaptive multipole algorithm in three dimensions. Journal of Computational Physics, 155(2):468–498, November 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963556.

Christie:1990:NSH

[CGSS90] I. Christie, G. H. Ganser, and J. M. Sanz-Serna. Numerical solution of a hyperbolic system of conservation laws with source term arising in a fluidized bed model. Journal of Computational Physics, 89(1):255, July 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090140V.

Christie:1991:NSH

[CGSS91] I. Christie, G. H. Ganser, and J. M. Sanz-Serna. Numerical solution of a hyperbolic system of conservation laws with source term arising in a fluidized bed model. Journal of Computational Physics, 93(2):297–311, April 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190184M.

Chen:1992:PEF

[CGY92] G. Q. Chen, Z. Gao, and Z. F. Yang. A perturbational h^{4} exponential finite difference scheme for the convective diffusion equation. Journal of Computational Physics, 101(2):454, August 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290035W.

Chen:1993:PEF

[CGY93] G. Q. Chen, Z. Gao, and Z. F. Yang. A perturbational h^{4} exponential finite difference scheme for the convective diffusion equation. Journal of Computational Physics, 104(1):129–139, January 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183710156.

Chesshire:1990:COM

[CH90a] G. Chesshire and W. D. Henshaw. Composite overlapping meshes for the solution of partial differential equations. Journal of Computational Physics, 90(1):1–64, September 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190901968.

Clark:1990:MDS

[CH90b] Terry L. Clark and William D. Hall. Multi-domain simulations of the time-dependent Navier–Stokes equations: Benchmark error analysis of some nesting procedures. Journal of Computational Physics, 89(1):252–253, July 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090132K.

Clark:1991:MDS

[CH91] Terry L. Clark and William D. Hall. Multi-domain simulations of the time dependent Navier–Stokes equations: Benchmark error analysis of some nesting procedures. Journal of Computational Physics, 92(2):456–481, February 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190218A.

Carcione:1999:NSP

[CH99] José M. Carcione and Hans B. Helle. Numerical solution of the poroviscoelastic wave equation on a staggered mesh. Journal of Computational Physics, 154(2):520–527, September 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199963210.

Chang:1990:CAM

[Cha90] Sin-Chung Chang. A critical analysis of the modified equation technique of Warming and Hyett. Journal of Computational Physics, 86(1):107–126, January 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090093G.

Chang:1995:MST

[Cha95] Sin-Chung Chang. The method of space–time conservation element and solution element — a new approach for solving the Navier–Stokes and Euler equations. Journal of Computational Physics, 119(2):295–324, July 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711370.

Chen:1991:CGM

[Che91] K. Chen. Conjugate gradient methods for the solution of boundary integral equations on a piecewise smooth boundary. Journal of Computational Physics, 97(1):127–143, November 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190042J.

Chin:1991:DDM

[Chi91] Raymond C. Y. Chin. A domain decomposition method for generating orthogonal polynomials for a Gaussian weight on a finite interval. Journal of Computational Physics, 97(2):582, December 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190024F.

Chin:1992:DDMa

[Chi92a] Raymond C. Y. Chin. A domain decomposition method for generating orthogonal polynomials for a Gaussian weight on a finite interval. Journal of Computational Physics, 99(1):180, March 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999192902938.

Chin:1992:DDMb

[Chi92b] Raymond C. Y. Chin. A domain decomposition method for generating orthogonal polynomials for a Gaussian weight on a finite interval. Journal of Computational Physics, 99(2):321–336, April 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290211G.

Charlton:1990:CLN

[CHL^{+}90] L. A. Charlton, J. A. Holmes, V. E. Lynch, B. A. Carreras, and T. C. Hender. Compressible linear and nonlinear resistive MHD calculations in toroidal geometry. Journal of Computational Physics, 86(2):270–293, February 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190901027.

Chang:1996:LSF

[CHMO96] Y. C. Chang, T. Y. Hou, B. Merriman, and S. Osher. A level set formulation of Eulerian interface capturing methods for incompressible fluid flows. Journal of Computational Physics, 124(2):449–464, March 15, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900728.

Chorin:1990:HRVa

[Cho90a] Alexander J. Chorin. Hairpin removal in vortex interactions. Journal of Computational Physics, 87(2):496, April 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190902723.

Chorin:1990:HRVb

[Cho90b] Alexandre Joel Chorin. Hairpin removal in vortex interactions. Journal of Computational Physics, 91(1):1–21, November 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919090001H.

Chopp:1993:CMS

[Cho93a] David L. Chopp. Computing minimal surfaces via level set curvature flow. Journal of Computational Physics, 106(1):77–91, May 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183710922.

Chorin:1993:HRV

[Cho93b] Alexandre Joel Chorin. Hairpin removal in vortex interactions II. Journal of Computational Physics, 107(1):1–9, July 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711204.

Chorin:1997:NMS

[Cho97] Alexandre Joel Chorin. A numerical method for solving incompressible viscous flow problems. Journal of Computational Physics, 135(2):118–125, August 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957168.

Christiansen:1997:NSH

[Chr97] J. P. Christiansen. Numerical simulation of hydrodynamics by the method of point vortices. Journal of Computational Physics, 135(2):189–197, August 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957016.

Cao:1999:AFE

[CHR99] Weiming Cao, Weizhang Huang, and Robert D. Russell. An radaptive finite element method based upon moving mesh PDEs. Journal of Computational Physics, 149(2):221–244, March 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198961514. See erratum [Ano99-30].

Cloot:1990:NSN

[CHW90] A. Cloot, B. M. Herbst, and J. A. C. Weideman. A numerical study of the nonlinear Schrödinger equation involving quintic terms. Journal of Computational Physics, 86(1):127–146, January 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919090094H.

Charakhchyan:1997:VFW

[CI97] A. A. Charakhch’yan and S. A. Ivanenko. A variational form of the Winslow grid generator. Journal of Computational Physics, 136(2):385–398, September 15, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957508.

Chang:1994:CDS

[CJ94] Qianshun Chang and Hong Jiang. A conservative difference scheme for the Zakharov equations. Journal of Computational Physics, 113(2):309–319, August 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711387.

Corral:1995:FCM

[CJ95a] Roque Corral and Javier Jiménez. Fourier/Chebyshev methods for the incompressible Navier–Stokes equations in infinite domains. Journal of Computational Physics, 121(2):261–270, October 15, 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999195900853.

Cote:1995:CPS

[CJ95b] R. Côté and M. J. Jamieson. Comparison of phase-shift calculations by asymptotic fit and quadrature in ultra-low temperature scattering. Journal of Computational Physics, 118(2):388–391, May 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185711084.

Collino:1997:FDM

[CJM97] Francis Collino, Patrick Joly, and Florence Millot. Fictitious domain method for unsteady problems. Journal of Computational Physics, 138(2):907–938, December 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958496.

Corless:1991:NEA

[CJR91] R. M. Corless, D. J. Jeffrey, and H. Rasmussen. Numerical evaluation of Airy functions with complex arguments. Journal of Computational Physics, 93(1):252–253, March 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999191900894.

Corless:1992:NEAa

[CJR92a] R. M. Corless, D. J. Jeffrey, and H. Rasmussen. Numerical evaluation of Airy functions with complex arguments. Journal of Computational Physics, 98(2):347, February 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290150W.

Corless:1992:NEAb

[CJR92b] R. M. Corless, D. J. Jeffrey, and H. Rasmussen. Numerical evaluation of Airy functions with complex arguments. Journal of Computational Physics, 99(1):106–114, March 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192902798.

Chen:1995:VBC

[CJR95] Shea Chen, David B. Johnson, and Peter E. Raad. Velocity boundary conditions for the simulation of free surface fluid flow. Journal of Computational Physics, 116(2):262–276, February 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002199918571025X.

Chang:1999:DSS

[CJS99] Qianshun Chang, Erhui Jia, and W. Sun. Difference schemes for solving the generalized nonlinear Schrödinger equation. Journal of Computational Physics, 148(2):397–415, January 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198961204.

Cacuci:1990:BPM

[CK90a] D. G. Cacuci and O. A. Karakashian. Benchmarking the propagator method for nonlinear systems: a Burgers–Korteweg–de Vries equation. Journal of Computational Physics, 89(1):63–79, July 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919090117J.

Cohen:1990:ADD

[CK90b] Stuart B. Cohen and Ivan N. Krischner. Approximating the Dirac distribution for Fourier analysis. Journal of Computational Physics, 89(1):255, July 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090141M.

Cohen:1991:ADD

[CK91] Stuart B. Cohen and Ivan N. Kirschner. Approximating the Dirac distribution for Fourier analysis. Journal of Computational Physics, 93(2):312–324, April 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190185N.

Chung:1992:ESM

[CK92] S. G. Chung and K. Kuwahara. Explicit streamline method for steady flows of non-Newtonian matter: History-dependence and free surfaces. Journal of Computational Physics, 102(2): 425, October 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290393D.

Chung:1993:ESM

[CK93a] S. G. Chung and K. Kuwahara. Explicit streamline method for steady flows of non-Newtonian matter: History-dependence and free surfaces. Journal of Computational Physics, 104(2):444– 450, February 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999183710429.

Coray:1993:AOM

[CK93b] C. Coray and J. Koebbe. Accuracy optimized methods for constrained numerical solutions of hyperbolic conservation laws. Journal of Computational Physics, 109(1):115–132, November 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183712040.

Capstick:1996:MQA

[CK96] Simon Capstick and B. D. Keister. Multidimensional quadrature algorithms at higher degree and/or dimension. Journal of Computational Physics, 123(2):267–273, February 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900236.

Chukkapalli:1999:SGU

[CKE99] Giri Chukkapalli, Steve R. Karpik, and C. Ross Ethier. A scheme for generating unstructured grids on spheres with application to parallel computation. Journal of Computational Physics, 149(1):114–127, February 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961460.

Clarke:1993:NCT

[CKQ^{+}93] J. F. Clarke, S. Karni, J. J. Quirk, P. L. Roe, L. G. Simmonds, and E. F. Toro. Numerical computation of two-dimensional unsteady detonation waves in high energy solids. Journal of Computational Physics, 106(2):215–233, June 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711046.

Chung:1993:SCF

[CKR93] S. G. Chung, K. Kuwahara, and O. Richmond. Streamlinecoordinate finite-difference method for hot metal deformations. Journal of Computational Physics, 108(1):1–7, September 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183711575.

Christopher:1997:SPE

[CKSB97] Ivar Christopher, George Knorr, Magdi Shoucri, and Pierre Bertrand. Solution of the Poisson equation in an annulus. Journal of Computational Physics, 131(2):323–326, March 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196955989.

Chew:1993:FAC

[CL93] W. C. Chew and C. C. Lu. A fast algorithm to compute the wave-scattering solution of a large strip. Journal of Computational Physics, 107(2):378–387, August 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711526.

Choi:1998:RUF

[CL98a] Hwajeong Choi and Jian-Guo Liu. The reconstruction of upwind fluxes for conservation laws: Its behavior in dynamic and steady state calculations. Journal of Computational Physics, 144(2):237–256, August 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959708.

Cossu:1998:CAN

[CL98b] C. Cossu and T. Loiseleux. On the convective and absolute nature of instabilities in finite difference numerical simulations of open flows. Journal of Computational Physics, 144(1):98–108, July 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198959903.

Coppa:1996:BMK

[CLD^{+}96] G. G. M. Coppa, G. Lapenta, G. Dellapiana, F. Donato, and V. Riccardo. Blob method for kinetic plasma simulation with variable-size particles. Journal of Computational Physics, 127(2):268–284, September 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901746.

Clement:1996:CTA

[Clé96] A. Clément. Coupling of two absorbing boundary conditions for 2D time-domain simulations of free surface gravity waves. Journal of Computational Physics, 126(1):139–151, June 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901266.

Clercx:1997:SSN

[Cle97] H. J. H. Clercx. A spectral solver for the Navier–Stokes equations in the velocity-vorticity formulation for flows with two nonperiodic directions. Journal of Computational Physics, 137(1):186– 211, October 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999197957995.

Carreras:1990:FWC

[CLJ^{+}90] B. A. Carreras, V. E. Lynch, E. F. Jaeger, D. B. Batchelor, and J. S. Tolliver. Full-wave calculations in flux coordinates for toroidal geometry. Journal of Computational Physics, 88(1):183– 204, May 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919090247X.

Cai:1998:FEM

[CLNT98] Xing Cai, Hans Petter Langtangen, Bjørn Fredrik Nielsen, and Aslak Tveito. A finite element method for fully nonlinear water waves. Journal of Computational Physics, 143(2):544–568, July 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198999971.

Cai:1993:CSM

[CLO93] W. Cai, H. C. Lee, and H.-S. Oh. Coupling of spectral methods and the p-version of the finite element method for elliptic boundary value problems containing singularities. Journal of Computational Physics, 108(2):314–326, October 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918371185X.

Cooke:1991:AFO

[CM91] Charlie H. Cooke and Andrew G. McMorran. Asymptotic factorization of operators in complex time. Journal of Computational Physics, 95(2):303–312, August 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190278S.

Choi:1993:APV

[CM93] Y.-H. Choi and C. L. Merkle. The application of preconditioning in viscous flows. Journal of Computational Physics, 105(2):207– 223, April 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999183710697.

Choi:1994:ECT

[CM94] Haecheon Choi and Parviz Moin. Effects of the computational time step on numerical solutions of turbulent flow. Journal of Computational Physics, 113(1):1–4, July 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711120.

Christov:1995:IDS

[CM95] C. I. Christov and G. A. Maugin. An implicit difference scheme for the long-time evolution of localized solutions of a generalized Boussinesq system. Journal of Computational Physics, 116(1): 39–51, January 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999185710042.

Chow:1997:US

[CM97] E. Chow and J. J. Monaghan. Ultrarelativistic SPH. Journal of Computational Physics, 134(2):296–305, July 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957089.

Chen:1999:LVM

[CM99a] H. Chen and J. S. Marshall. A Lagrangian vorticity method for two-phase particulate flows with two-way phase coupling. Journal of Computational Physics, 148(1):169–198, January 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961162.

Cleary:1999:CMU

[CM99b] Paul W. Cleary and Joseph J. Monaghan. Conduction modelling using smoothed particle hydrodynamics. Journal of Computational Physics, 148(1):227–264, January 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961186.

Chen:1990:TML

[CMK90] H. Chen, W. H. Matthaeus, and L. W. Klein. Theory of multicolor lattice gas: a cellular automaton Poisson solver. Journal of Computational Physics, 88(2):433–466, June 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190901887.

Crumpton:1993:CVA

[CMM93] P. I. Crumpton, J. A. Mackenzie, and K. W. Morton. Cell vertex algorithms for the compressible Navier–Stokes equations. Journal of Computational Physics, 109(1):1–15, November 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711940.

Chan:1994:SSM

[CMMH94] Kwing L. Chan, H. G. Mayr, J. G. Mengel, and I. Harris. A “stratified” spectral model for stable and convective atmospheres. Journal of Computational Physics, 113(2):165–176, August 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999184711284.

Chen:1997:SLS

[CMOS97] S. Chen, B. Merriman, S. Osher, and P. Smereka. A simple level set method for solving Stefan problems. Journal of Computational Physics, 135(1):8–29, July 15, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957211.

Cash:1995:ACS

[CMW95] J. R. Cash, G. Moore, and R. W. Wright. An automatic continuation strategy for the solution of singularly perturbed linear two-point boundary value problems. Journal of Computational Physics, 122(2):266–279, December 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185712120.

Cai:1995:PBP

[CN95] Y. Cai and I. M. Navon. Parallel block preconditioning techniques for the numerical simulation of the shallow water flow using finite element methods. Journal of Computational Physics, 122(1):39– 50, November 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999185711953.

Carpenter:1999:SCI

[CNG99] Mark H. Carpenter, Jan Nordström, and David Gottlieb. A stable and conservative interface treatment of arbitrary spatial accuracy. Journal of Computational Physics, 148(2):341–365, January 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198961149. See corrigendum [CNG17].

Carpenter:2017:CSC

[CNG17] Mark H. Carpenter, Jan Nordström, and David Gottlieb. Corrigendum to “A stable and conservative interface treatment of arbitrary spatial accuracy” [J. Comput. Phys. 148 (1999) 341– 365]. Journal of Computational Physics, 351(??):534, December 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117307179. See [CNG99].

Carpenter:1995:HOC

[CO95] Mark H. Carpenter and John Otto. High-order “cyclo–difference” techniques: An alternative to finite differences. Journal of Computational Physics, 118(2):242–260, May 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710960.

Cybyk:1995:CML

[COBA95] Bohdan Z. Cybyk, Elaine S. Oran, Jay P. Boris, and John D. Anderson, Jr. Combining the monotonic Lagrangian grid with a direct simulation Monte Carlo model. Journal of Computational Physics, 122(2):323–334, December 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185712181.

Codina:1999:NSI

[Cod99] Ramon Codina. Numerical solution of the incompressible Navier–Stokes equations with Coriolis forces based on the discretization of the total time derivative. Journal of Computational Physics, 148(2):467–496, January 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961265.

Colella:1990:MUM

[Col90] Phillip Colella. Multidimensional upwind methods for hyperbolic conservation laws. Journal of Computational Physics, 87(1):171– 200, March 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919090233Q.

Collino:1997:PMA

[Col97] Francis Collino. Perfectly matched absorbing layers for the paraxial equations. Journal of Computational Physics, 131(1):164– 180, February 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196955941.

Cook:1999:CAL

[Coo99] Andrew W. Cook. A consistent approach to large eddy simulation using adaptive mesh refinement. Journal of Computational Physics, 154(1):117–133, September 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963040.

Cortez:1996:IBA

[Cor96] Ricardo Cortez. An impulse-based approximation of fluid motion due to boundary forces. Journal of Computational Physics, 123(2):341–353, February 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900285.

Canino:1998:NSH

[COS^{+}98] Lawrence F. Canino, John J. Ottusch, Mark A. Stalzer, John L. Visher, and Stephen M. Wandzura. Numerical solution of the Helmholtz equation in 2D and 3D using a high-order Nyström discretization. Journal of Computational Physics, 146(2):627–663, November 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198960776.

Cottet:1990:PGS

[Cot90] Georges-Henri Cottet. A particle-grid superposition method for the Navier–Stokes equations. Journal of Computational Physics, 89(2):301–318, August 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919090146R.

Cottet:1996:AVM

[Cot96] G.-H. Cottet. Artificial viscosity models for vortex and particle methods. Journal of Computational Physics, 127(2):299–308, September 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199919690176X.

Canuto:1990:BIC

[CP90] Claudio Canuto and Paola Pietra. Boundary and interface conditions within a finite element preconditioner for spectral methods. Journal of Computational Physics, 91(2):310–343, December 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999190900408.

Coirier:1995:AAC

[CP95] William J. Coirier and Kenneth G. Powell. An accuracy assessment of Cartesian-mesh approaches for the Euler equations. Journal of Computational Physics, 117(1):121–131, March 1, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185710509.

Carsky:1998:EMI

[CP98a] Petr Cársky and Martin Polásek. Evaluation of molecular integrals in a mixed Gaussian and plane-wave basis by Rys quadrature. Journal of Computational Physics, 143(1):266–277, June 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198959769.

Carsky:1998:IGF

[CP98b] Petr Cársky and Martin Polásek. Incomplete gamma F_{m}(x) functions for real negative and complex arguments. Journal of Computational Physics, 143(1):259–265, June 10, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959757.

Colella:1999:PML

[CP99] Phillip Colella and Karen Pao. A projection method for low speed flows. Journal of Computational Physics, 149(2):245–269, March 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198961526.

Casciola:1996:VGF

[CPB96] C. M. Casciola, R. Piva, and P. Bassanini. Vorticity generation on a flat surface in 3D flows. Journal of Computational Physics, 129(2):345–356, December 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196902545.

Chen:1990:SRA

[CPJ90] H. C. Chen, V. C. Patel, and S. Ju. Solutions of Reynoldsaveraged Navier–Stokes equations for three-dimensional incompressible flows. Journal of Computational Physics, 88(2):305– 336, June 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919090182Z.

Christiansen:1993:FMM

[CPP93] Dorthe Christiansen, John W. Perram, and Henrik G. Petersen. On the fast multipole method for computing the energy of periodic assemblies of charged and dipolar particles. Journal of Computational Physics, 107(2):403–405, August 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918371154X.

Candy:1991:SIA

[CR91] J. Candy and W. Rozmus. A symplectic integration algorithm for separable Hamiltonian functions. Journal of Computational Physics, 92(1):230–256, January 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190299Z.

Cook:1996:DNS

[CR96] Andrew W. Cook and James J. Riley. Direct numerical simulation of a turbulent reactive plume on a parallel computer. Journal of Computational Physics, 129(2):263–283, December 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196902491.

Cummins:1999:SPM

[CR99] Sharen J. Cummins and Murray Rudman. An SPH projection method. Journal of Computational Physics, 152(2):584–607, July 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199962460.

Crater:1994:PEI

[Cra94] H. W. Crater. A Padé extrapolated inverse power method for coupled Schrödinger-like equations applied to the two-body relativistic bound state problem in quantum electrodynamics. Journal of Computational Physics, 115(2):470–484, December 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184712113.

Coward:1997:TEP

[CRRR97] Adrian V. Coward, Yuriko Y. Renardy, Michael Renardy, and John R. Richards. Temporal evolution of periodic disturbances in two-layer Couette flow. Journal of Computational Physics, 132(2):346–361, April 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196956405.

Cash:1990:SOE

[CRS90] J. R. Cash, A. D. Raptis, and T. E. Simos. A sixthorder exponentially fitted method for the numerical solution of the radial Schrödinger equation. Journal of Computational Physics, 91(2):413–423, December 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190900453.

Caginalp:1991:CSP

[CS91] G. Caginalp and E. A. Socolovsky. Computation of sharp phase boundaries by spreading: The planar and spherically symmetric cases. Journal of Computational Physics, 95(1):85–100, July 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190254I.

Cai:1992:UHO

[CS92a] Wei Cai and Chi-Wang Shu. Uniform high order spectral methods for one- and two-dimensional Euler equations. Journal of Computational Physics, 102(2):425, October 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290392C.

Christodoulou:1992:DFSa

[CS92b] K. N. Christodoulou and L. E. Scriven. Discretization of free surface flows and other moving boundary problems. Journal of Computational Physics, 98(2):347, February 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290148R.

Christodoulou:1992:DFSb

[CS92c] K. N. Christodoulou and L. E. Scriven. Discretization of free surface flows and other moving boundary problems. Journal of Computational Physics, 99(1):39–55, March 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192902732.

Cai:1993:UHO

[CS93a] Wei Cai and Chi-Wang Shu. Uniform high-order spectral methods for one- and two-dimensional Euler equations. Journal of Computational Physics, 104(2):427–443, February 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710417.

Clarke:1993:GOP

[CS93b] Andrew S. Clarke and Bernie Shizgal. On the generation of orthogonal polynomials using asymptotic methods for recurrence coefficients. Journal of Computational Physics, 104(1):140–149, January 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999183710168.

Craig:1993:NSG

[CS93c] W. Craig and C. Sulem. Numerical simulation of gravity waves. Journal of Computational Physics, 108(1):73–83, September 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183711642.

Carvalho:1997:FFD

[CS97a] M. S. Carvalho and L. E. Scriven. Flows in forward deformable roll coating gaps: Comparison between spring and plane-strain models of roll cover. Journal of Computational Physics, 138(2):449–479, December 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958265.

Cocchi:1997:RPB

[CS97b] Jean-Pierre Cocchi and Richard Saurel. A Riemann problem based method for the resolution of compressible multimaterial flows. Journal of Computational Physics, 137(2):265–298, November 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999197957685.

Caramana:1998:EAG

[CS98a] E. J. Caramana and M. J. Shashkov. Elimination of artificial grid distortion and hourglass-type motions by means of Lagrangian subzonal masses and pressures. Journal of Computational Physics, 142(2):521–561, May 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959526.

Cockburn:1998:RKD

[CS98b] Bernardo Cockburn and Chi-Wang Shu. The Runge–Kutta discontinuous Galerkin method for conservation laws V: Multidimensional systems. Journal of Computational Physics, 141(2):199–224, April 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999198958922.

Carvalho:1999:TDS

[CS99] M. S. Carvalho and L. E. Scriven. Three-dimensional stability analysis of free surface flows: Application to forward deformable roll coating. Journal of Computational Physics, 151(2):534–562, May 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999199961958.

Cai:1993:PLP

[CSN93] D. Cai, L. R. O. Storey, and T. Neubert. Particle loading for a plasma shear layer in a magnetic field. Journal of Computational Physics, 107(1):84–97, July 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711265.

Crumpton:1995:DMS

[CSW95] P. I. Crumpton, G. J. Shaw, and A. F. Ware. Discretisation and multigrid solution of elliptic equations with mixed derivative terms and strongly discontinuous coefficients. Journal of Computational Physics, 116(2):343–358, February 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710327.

Caramana:1998:FAV

[CSW98] E. J. Caramana, M. J. Shashkov, and P. P. Whalen. Formulations of artificial viscosity for multi-dimensional shock wave computations. Journal of Computational Physics, 144(1):70–97, July 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198959897.

Christiansen:1997:SAC

[CSZ97] P. L. Christiansen, A. V. Savin, and A. V. Zolotaryuk. Soliton analysis in complex molecular systems: a zig-zag chain. Journal of Computational Physics, 134(1):108–121, June 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919795676X.

Cliffe:1998:MBC

[CT98] K. A. Cliffe and S. J. Tavener. Marangoni–Bénard convection with a deformable free surface. Journal of Computational Physics, 145(1):193–227, September 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959952.

Canfield:1992:RARa

[CV92a] E. H. Canfield, Jr. and J. A. Viecelli. Random access to a random number sequence. Journal of Computational Physics, 98(2): 349, February 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290164T.

Canfield:1992:RARb

[CV92b] E. H. Canfield, Jr. and J. A. Viecelli. Random access to a random number sequence. Journal of Computational Physics, 99(1):176– 178, March 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999192902857.

Cortez:1997:DEM

[CV97] R. Cortez and D. A. Varela. The dynamics of an elastic membrane using the impulse method. Journal of Computational Physics, 138(1):224–247, November 15, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958423.

Chang:1990:MAA

[CW90] Qianshun Chang and Guobin Wang. Multigrid and adaptive algorithm for solving the nonlinear Schrödinger equation. Journal of Computational Physics, 88(2):362–380, June 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190901843.

Caramana:1992:PSOa

[CW92a] Edward J. Caramana and Robert B. Webster. Parameterized solution of one-dimensional thermal diffusion with a heat source and a moving boundary. Journal of Computational Physics, 98(1): 180, January 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999192901988.

Caramana:1992:PSOb

[CW92b] Edward J. Caramana and Robert B. Webster. Parameterized solution of one-dimensional thermal diffusion with a heat source and a moving boundary. Journal of Computational Physics, 98(2):342–346, February 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290147Q.

Cloot:1992:AAS

[CW92c] A. Cloot and J. A. C. Weideman. An adaptive algorithm for spectral computations on unbounded domains. Journal of Computational Physics, 102(2):398–406, October 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192903818.

Cohen:1993:ISI

[CW93] Bruce I. Cohen and Timothy J. Williams. Implementation of a semi-implicit orbit-averaged gyrokinetic particle code. Journal of Computational Physics, 107(2):282–290, August 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711435.

Caramana:1998:NPS

[CW98] E. J. Caramana and P. P. Whalen. Numerical preservation of symmetry properties of continuum problems. Journal of Computational Physics, 141(2):174–198, April 10, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959125.

Chang:1999:STC

[CWC99] Sin-Chung Chang, Xiao-Yen Wang, and Chuen-Yen Chow. The space–time conservation element and solution element method: a new high-resolution and genuinely multidimensional paradigm for solving conservation laws. Journal of Computational Physics, 156(1):89–136, November 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963544.

Chang:1996:AMM

[CWF96] Qianshun Chang, Yau Shu Wong, and Hanqing Fu. On the algebraic multigrid method. Journal of Computational Physics, 125(2):279–292, May 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999196900947.

Chang:1990:CSM

[CWG90] Qianshun Chang, Goubin Wang, and Boling Guo. Conservative scheme for a model of nonlinear dispersive waves and its solitary waves induced by boundary motion. Journal of Computational Physics, 89(2):489, August 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090158W.

Chang:1991:CSM

[CWG91] Qianshun Chang, Guobin Wang, and Boling Guo. Conservative scheme for a model of nonlinear dispersive waves and its solitary waves induced by boundary motion. Journal of Computational Physics, 93(2):360–375, April 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190189R.

Cho:1992:CMO

[CYD92] S. Y. Cho, R. A. Yetter, and F. L. Dryer. A computer model for one-dimensional mass and energy transport in and around chemically reacting particles, including complex gasphase chemistry, multicomponent molecular diffusion, surface evaporation, and heterogeneous reaction. Journal of Computational Physics, 102(1):160–179, September 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999105800130.

Cai:1998:ASW

[CZ98] Wei Cai and Wu Zhang. An adaptive spline wavelet ADI (SW– ADI) method for two-dimensional reaction-diffusion equations. Journal of Computational Physics, 139(1):92–126, January 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197958654.

Chen:1998:LFE

[CZS98] Wen Chen, Tingxiu Zhong, and C. Shu. A Lyapunov formulation for efficient solution of the Poisson and convection-diffusion equations by the differential quadrature method. Journal of Computational Physics, 141(1):78–84, March 20, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198958612.

Dul:1994:TPM

[DA94] Franciszek A. Dul and Krzysztof Arczewski. The two-phase method for finding a great number of eigenpairs of the symmetric or weakly non-symmetric large eigenvalue problems. Journal of Computational Physics, 111(1):89–109, March 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710473.

Daripa:1990:NMQ

[Dar90a] Prabir Daripa. On a numerical method for quasi-confirmal grid generation. Journal of Computational Physics, 91(1):254, November 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919090035Y.

Daripa:1990:ACV

[Dar90b] Prabir Daripa. On applications of a complex variable method in compressible flows. Journal of Computational Physics, 88(2): 337–361, June 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999190901832.

Daripa:1990:SCA

[Dar90c] Prabir Daripa. Solvability condition and its application to fast numerical solution of overposed inverse problems in compressible flows. Journal of Computational Physics, 91(1):253, November 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999190900305.

Daripa:1991:NMQ

[Dar91a] Prabir Daripa. On a numerical method for quasi-conformal grid generation. Journal of Computational Physics, 96(1):229–236, September 1991. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919190274O.

Daripa:1991:SCA

[Dar91b] Prabir Daripa. Solvability condition and its application to fast numerical solution of overposed inverse problems in compressible flows. Journal of Computational Physics, 95(2):436–449, August 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190284R.

Daripa:1992:ISA

[Dar92] Prabir Daripa. Iterative schemes and algorithms for adaptive grid generation in one dimension. Journal of Computational Physics, 100(2):284–293, June 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290235Q.

Daripa:1993:FAS

[Dar93] Prabir Daripa. A fast algorithm to solve the Beltrami equation with applications to quasiconformal mappings. Journal of Computational Physics, 106(2):355–365, June 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711137.

Daripa:1995:ISP

[Dar95] Prabir Daripa. An investigation of some pattern selection issues in the rising plane Taylor bubble problem. Journal of Computational Physics, 121(1):129–141, October 1, 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711849.

Daripa:1996:PTB

[Dar96] Prabir Daripa. Pointed Taylor bubble revisited. Journal of Computational Physics, 123(1):226–230, January 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900182.

deAraujo:1995:CKV

[dAT95] Carlos F. de Araujo, Jr., Sadhan K. Adhikari, and Lauro Tomio. Complex Kohn variational principle for two-nucleon bound-state and scattering with the tensor potential. Journal of Computational Physics, 118(2):200–207, May 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710923.

Daube:1992:RNSa

[Dau92a] O. Daube. Resolution of the 2D Navier–Stokes equations in velocity-vorticity form by means of an influence matrix technique. Journal of Computational Physics, 101(2):452, August 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290021P.

Daube:1992:RNSb

[Dau92b] O. Daube. Resolution of the 2D Navier–Stokes equations in velocity-vorticity form by means of an influence matrix technique. Journal of Computational Physics, 103(2):402–414, December 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290411Q.

Davstad:1991:MCR

[Dav91] Kjell Davstad. A multigrid conjugate residual method for the numerical solution of the Hartree–Fock equation for diatomic molecules. Journal of Computational Physics, 93(2):486, April 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190202V.

Davstad:1992:MCRa

[Dav92a] Kjell Davstad. A multigrid conjugate residual method for the numerical solution of the Hartree–Fock equation for diatomic molecules. Journal of Computational Physics, 98(1):180, January 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999192901977.

Davstad:1992:MCRb

[Dav92b] Kjell Davstad. A multigrid conjugate residual method for the numerical solution of the Hartree–Fock equation for diatomic molecules. Journal of Computational Physics, 99(1):33– 38, March 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290272Z.

Dietrich:1996:SPO

[DB96] Stefan Dietrich and Iain D. Boyd. Scalar and parallel optimized implementation of the direct simulation Monte Carlo method. Journal of Computational Physics, 126(2):328–342, July 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196901412.

Dimitropoulos:1997:ERS

[DB97] Costas D. Dimitropoulos and Antony N. Beris. An efficient and robust spectral solver for nonseparable elliptic equations. Journal of Computational Physics, 133(1):186–191, May 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956618.

Debreu:1998:SAM

[DB98] L. Debreu and E. Blayo. On the Schwarz alternating method for oceanic models on parallel computers. Journal of Computational Physics, 141(2):93–111, April 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198958983.

Danabasoglu:1994:ASM

[DBS94] G. Danabasoglu, S. Biringen, and C. L. Streett. Application of the spectral multidomain method to the Navier–Stokes equations. Journal of Computational Physics, 113(2):155–164, August 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999184711272.

Demaret:1990:CCS

[DD90] P. Demaret and M. O. Deville. Chebyshev collocation solutions of the Navier–Stokes equations using multi-domain decomposition and finite element preconditioning. Journal of Computational Physics, 91(1):254, November 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090033W.

Demaret:1991:CCS

[DD91] P. Demaret and M. O. Deville. Chebyshev collocation solutions of the Navier–Stokes equations using multi-domain decomposition and finite element preconditioning. Journal of Computational Physics, 95(2):359–386, August 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190281O.

Dietrich:1992:CLT

[DD92a] S. Dietrich and H. Dette. Correlation length of time series in statistical simulations. Journal of Computational Physics, 101(1): 224–226, July 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999192900554.

Dukowicz:1992:AFH

[DD92b] John K. Dukowicz and Arkady S. Dvinsky. Approximate factorization as a high order splitting for the implicit incompressible flow equations. Journal of Computational Physics, 102(2):336– 347, October 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290376A.

Draghicescu:1995:FAV

[DD95] Cristina I. Draghicescu and Mircea Draghicescu. A fast algorithm for vortex blob interactions. Journal of Computational Physics, 116(1):69–78, January 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710066.

Dawkins:1998:ONS

[DDD98] Paul T. Dawkins, Steven R. Dunbar, and Rod W. Douglass. The origin and nature of spurious eigenvalues in the spectral tau method. Journal of Computational Physics, 147(2):441–462, December 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198960958.

Dawson:1997:HFD

[DDS97] Trevor W. Dawson, Jan De Moerloose, and Maria A. Stuchly. Hybrid finite-difference method for high-resolution modelling of low-frequency electric induction in humans. Journal of Computational Physics, 136(2):640–653, September 15, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957910.

DeLillo:1993:NCM

[DE93] Thomas K. DeLillo and Alan R. Elcrat. Numerical conformal mapping methods for exterior regions with corners. Journal of Computational Physics, 108(2):199–208, October 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711757.

Druzhinin:1999:LEM

[DE99] O. A. Druzhinin and S. E. Elghobashi. A Lagrangian–Eulerian mapping solver for direct numerical simulation of bubble-laden turbulent shear flows using the two-fluid formulation. Journal of Computational Physics, 154(1):174–196, September 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963118.

Dimitropoulos:1998:EPF

[DECB98] Costas D. Dimitropoulos, Brian J. Edwards, Kyung-Sun Chae, and Antony N. Beris. Efficient pseudospectral flow simulations in moderately complex geometries. Journal of Computational Physics, 144(2):517–549, August 10, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960090.

Demeio:1992:ICEa

[Dem92a] L. Demeio. The inclusion of collisional effects in the splitting scheme. Journal of Computational Physics, 98(2):348, February 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290160Z.

Demeio:1992:ICEb

[Dem92b] L. Demeio. The inclusion of collisional effects in the splitting scheme. Journal of Computational Physics, 99(2):203–208, April 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290202A.

Durlofsky:1992:TBA

[DEO92] Louis J. Durlofsky, Bjorn Engquist, and Stanley Osher. Triangle based adaptive stencils for the solution of hyperbolic conservation laws. Journal of Computational Physics, 98(1):64–73, January 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290173V.

DeVore:1991:FCT

[DeV91] C. Richard DeVore. Flux-corrected transport techniques for multidimensional compressible magnetohydrodynamics. Journal of Computational Physics, 92(1):142–160, January 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190295V.

Degani:1996:PMC

[DF96] A. T. Degani and G. C. Fox. Parallel multigrid computation of the unsteady incompressible Navier–Stokes equations. Journal of Computational Physics, 128(1):223–236, October 1, 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196902053.

Driscoll:1998:BPM

[DF98] Tobin A. Driscoll and Bengt Fornberg. A block pseudospectral method for Maxwell’s equations: I. One-dimensional case. Journal of Computational Physics, 140(1):47–65, February 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198958831.

Donat:1998:FSA

[DFaIM98] R. Donat, J. A. Font, J. M. a Ibáñez, and A. Marquina. A flux-split algorithm applied to relativistic flows. Journal of Computational Physics, 146(1):58–81, October 10, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959551.

deWindt:1994:CAN

[dFD^{+}94] Laurent deWindt, Patrick Fischer, Mireille Defranceschi, Joseph Delhalle, and Joseph G. Fripiat. A combined analytical and numerical strategy to solve the atomic Hartree–Fock equations in momentum space. Journal of Computational Physics, 111(2): 266–274, April 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199918471062X.

Dillon:1996:MBP

[DFFG96] Robert Dillon, Lisa Fauci, Aaron Fogelson, and Donald Gaver III. Modeling biofilm processes using the immersed boundary method. Journal of Computational Physics, 129(1):57–73, November 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196902338.

Ducros:1999:LES

[DFN^{+}99] F. Ducros, V. Ferrand, F. Nicoud, C. Weber, D. Darracq, C. Gacherieu, and T. Poinsot. Large-eddy simulation of the shock/turbulence interaction. Journal of Computational Physics, 152(2):517–549, July 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962381.

deFrutos:1992:EIF

[dFSS92] J. de Frutos and J. M. Sanz-Serna. An easily implementable fourth-order method for the time integration of wave problems. Journal of Computational Physics, 103(1):160–168, November 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290331R.

daGama:1992:SSSa

[dG92a] Rogério Martins Saldanha da Gama. Simulation of the steadystate energy transfer in rigid bodies, with convective/radiative boundary conditions, employing a minimum principle. Journal of Computational Physics, 99(1):180, March 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192902927.

daGama:1992:SSSb

[dG92b] Rogério Martins Saldanha da Gama. Simulation of the steadystate energy transfer in rigid bodies, with convective/radiative boundary conditions, employing a minimum principle. Journal of Computational Physics, 99(2):310–320, April 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290210P.

daGama:1996:NSN

[dG96] Rogério Martins Saldanha da Gama. Numerical simulation of the (nonlinear) conduction/radiation heat transfer process in a nonconvex and black cylindrical body. Journal of Computational Physics, 128(2):341–350, October 15, 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196902156.

Degond:1990:PSS

[DGD90] P. Degond and F. Guyot-Delaurens. Particle simulations of the semiconductor Boltzmann equation for one-dimensional inhomogeneous structures. Journal of Computational Physics, 90(1):65– 97, September 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999190901979.

Dainton:1997:DCS

[DGN97] M. P. Dainton, M. H. Goldwater, and N. K. Nichols. Direct computation of stochastic flow in reservoirs with uncertain parameters. Journal of Computational Physics, 130(2):203–216, January 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196955783.

Dennis:1995:MSV

[DH95] S. C. R. Dennis and J. D. Hudson. Methods of solution of the velocity-vorticity formulation of the Navier–Stokes equations. Journal of Computational Physics, 122(2):300–306, December 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185712156.

Darmofal:1996:APP

[DH96] D. L. Darmofal and R. Haimes. An analysis of 3D particle path integration algorithms. Journal of Computational Physics, 123(1):182–195, January 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900157.

Dhaeseleer:1991:SCN

[DHS91] W. D. D’haeseleer, W. N. G. Hitchon, and J. L. Shohet. Self-consistent numerical “bounce-averaged” transport computations in stellarators employing a mufti-mesh approach. Journal of Computational Physics, 95(1):117–166, July 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190256K.

DiPeso:1994:ESD

[DHS94] Gregory DiPeso, Dennis W. Hewett, and Gregory F. Simonson. Extension of the streamlined Darwin model to quasineutral plasmas. Journal of Computational Physics, 111(2):237–247, April 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999184710588.

Douglas:1994:GPL

[DHSS94] Craig C. Douglas, Michael Heroux, Gordon Slishman, and Roger M. Smith. GEMMW: a portable Level 3 BLAS Winograd variant of Strassen’s matrix–matrix multiply algorithm. Journal of Computational Physics, 110(1):1–10, January 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710011.

Dick:1990:MFP

[Dic90] Erik Dick. Multigrid formulation of polynomial flux-difference splitting for steady Euler equations. Journal of Computational Physics, 91(1):161–173, November 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090009P.

Diebels:1995:NSB

[Die95] Stefan Diebels. Numerical study of barotropic and baroclinic behavior of a nonlinear two-layer model. Journal of Computational Physics, 117(1):114–120, March 1, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710492.

Drikakis:1998:NMM

[DIV98] D. Drikakis, O. P. Iliev, and D. P. Vassileva. A nonlinear multigrid method for the three-dimensional incompressible Navier–Stokes equations. Journal of Computational Physics, 146(1):301–321, October 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198960673.

Dvorak:1990:NCI

[DK90] Steven L. Dvorak and Edward F. Kuester. Numerical computation of the incomplete Lipschitz–Hankel integral Je_{0}(a, z). Journal of Computational Physics, 87(2):301–327, April 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090255Y.

Dvorak:1992:NCSb

[DK92a] Steven L. Dvorak and Edward F. Kuester. Numerical computation of 2D Sommerfeld integrals — a novel asymptotic extraction technique. Journal of Computational Physics, 98(1):178, January 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999192901852.

Dvorak:1992:NCSd

[DK92b] Steven L. Dvorak and Edward F. Kuester. Numerical computation of 2D Sommerfeld integrals — A novel asymptotic extraction technique. Journal of Computational Physics, 98(2):217– 230, February 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290139P.

Dvorak:1992:NCSc

[DK92c] Steven L. Dvorak and Edward F. Kuester. Numerical computation of 2D Sommerfeld integrals — decomposition of the angular integral. Journal of Computational Physics, 98(2):199–216, February 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290138O.

Dvorak:1992:NCSa

[DK92d] Steven L. Dvorak and Edward F. Kuester. Numerical computation of 2D Sommerfeld integrals-decomposition of the angular integral. Journal of Computational Physics, 98(1):178, January 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290184Z.

Denton:1995:A

[DK95] Richard E. Denton and M. Kotschenreuther. δf algorithm. Journal of Computational Physics, 119(2):283–294, July 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711369.

Deeba:1996:DMS

[DK96] E. Y. Deeba and S. A. Khuri. A decomposition method for solving the nonlinear Klein–Gordon equation. Journal of Computational Physics, 124(2):442–448, March 15, 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900716.

Ding:1998:LSI

[DK98] Yan Ding and Mutsuto Kawahara. Linear stability of incompressible flow using a mixed finite element method. Journal of Computational Physics, 139(2):243–273, January 20, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958769.

Dimits:1993:PLA

[DL93] A. M. Dimits and W. W. Lee. Partially linearized algorithms in gyrokinetic particle simulation. Journal of Computational Physics, 107(2):309–323, August 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711460.

Dumas:1994:DFS

[DL94a] Guy Dumas and Anthony Leonard. A divergence-free spectral expansions method for three-dimensional flows in sphericalgap geometries. Journal of Computational Physics, 111(2):205– 219, April 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184710564.

Dunn:1994:CCI

[DL94b] J. H. Dunn and S. G. Lambrakos. Calculating complex interactions in molecular dynamics simulations employing Lagrangian particle tracking schemes. Journal of Computational Physics, 111(1):15–23, March 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710394.

Dunn:1991:ACI

[DLMN91] J. H. Dunn, S. G. Lambrakos, P. G. Moore, and M. Nagumo. An algorithm for calculating intramolecular angle-dependent forces on vector computers. Journal of Computational Physics, 97(2): 584, December 1991. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919190030O.

Dunn:1992:ACIb

[DLMN92a] J. H. Dunn, S. G. Lambrakos, P. G. Moore, and M. Nagumo. An algorithm for calculating intramolecular angle-dependent forces on vector computers. Journal of Computational Physics, 99(1): 179, March 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290288A.

Dunn:1992:ACIa

[DLMN92b] J. H. Dunn, S. G. Lambrakos, P. G. Moore, and M. Nagumo. An algorithm for calculating intramolecular angle-dependent forces on vector computers. Journal of Computational Physics, 100(1): 17–24, May 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290306J.

deFrutos:1995:FDS

[dLMSS95] J. deFrutos, M. A. López-Marcos, and J. M. Sanz-Serna. A finite difference scheme for the K(2, 2) compacton equation. Journal of Computational Physics, 120(2):248–252, September 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711618.

Drummond:1990:CSM

[DM90] P. D. Drummond and I. K. Mortimer. Computer simulations of multiplicative stochastic differential equations. Journal of Computational Physics, 89(1):252, July 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090131J.

Das:1991:ASI

[DM91a] Shuvra Das and Ambar K. Mitra. An algorithm for the solution of inverse Laplace problems and its application in flaw identification in materials. Journal of Computational Physics, 94(1):254, May 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190157G.

Drummond:1991:CSM

[DM91b] P. D. Drummond and I. K. Mortimer. Computer simulations of multiplicative stochastic differential equations. Journal of Computational Physics, 93(1):144–170, March 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190077X.

Das:1992:ASIa

[DM92a] Shuvra Das and Ambar K. Mitra. An algorithm for the solution of inverse Laplace problems and its application in flaw identification in materials. Journal of Computational Physics, 98(1): 180, January 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999192901955.

Das:1992:ASIb

[DM92b] Shuvra Das and Ambar K. Mitra. An algorithm for the solution of inverse Laplace problems and its application in flaw identification in materials. Journal of Computational Physics, 99(1): 99–105, March 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999192902787.

Deissler:1992:ESR

[DM92c] Robert G. Deissler and Frank B. Molls. Effect of spatial resolution on apparent sensitivity to initial conditions of a decaying flow as it becomes turbulent. Journal of Computational Physics, 100(2): 430–432, June 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999192902503.

Dukowicz:1992:VEMa

[DM92d] John K. Dukowicz and Bertrand J. A. Meltz. Vorticity errors in multidimensional Lagrangian codes. Journal of Computational Physics, 98(2):347, February 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290151N.

Dukowicz:1992:VEMb

[DM92e] John K. Dukowicz and Bertrand J. A. Meltz. Vorticity errors in multidimensional Lagrangian codes. Journal of Computational Physics, 99(1):115–134, March 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290280C.

Dorofeyev:1993:ONA

[DM93] Vladimir E. Dorofeyev and Alexey K. Mazur. Optimization of numerical algorithms for internal coordinate molecular dynamics. Journal of Computational Physics, 107(2):359–366, August 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183711502.

Donat:1996:CSR

[DM96] Rosa Donat and Antonio Marquina. Capturing shock reflections: An improved flux formula. Journal of Computational Physics, 125(1):42–58, April 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999196900789.

deMarkus:1997:CAN

[dM97a] Alicia Serfaty de Markus. Chaotic algorithms: a numerical exploration of the dynamics of a stiff photoconductor model. Journal of Computational Physics, 132(2):409–411, April 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196956193.

Deng:1997:CHO

[DM97b] Xiaogang Deng and Hiroshi Maekawa. Compact high-order accurate nonlinear schemes. Journal of Computational Physics, 130(1):77–91, January 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999196955539.

Duong:1999:MDS

[DMW99] Tap Ha Duong, Ernest L. Mehler, and Harel Weinstein. Molecular dynamics simulation of membranes and a transmembrane helix. Journal of Computational Physics, 151(1):358–387, May 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199962228.

Dipeso:1991:PSP

[DMZ91] G. Dipeso, E. C. Morse, and R. W. Ziolkowski. δf and particle simulations of parametric instabilities. Journal of Computational Physics, 96(2):325–338, October 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190239H.

Dennis:1993:NSS

[DNN93] S. C. R. Dennis, M. Ng, and P. Nguyen. Numerical solution for the steady motion of a viscous fluid inside a circular boundary using integral conditions. Journal of Computational Physics, 108(1):142–152, September 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711691.

deNicola:1995:CCE

[dNPT95] C. de Nicola, G. Pinto, and R. Tognaccini. On the conditional consistency of an explicit numerical scheme. Journal of Computational Physics, 120(2):378–382, September 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711734.

Dobson:1999:EMB

[Dob99] David C. Dobson. An efficient method for band structure calculations in 2D photonic crystals. Journal of Computational Physics, 149(2):363–376, March 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961575.

Dold:1992:ESIa

[Dol92a] J. W. Dold. An efficient surface-integral algorithm applied to unsteady gravity waves. Journal of Computational Physics, 101 (1):227, July 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999192900576.

Dold:1992:ESIb

[Dol92b] J. W. Dold. An efficient surface-integral algorithm applied to unsteady gravity waves. Journal of Computational Physics, 103(1):90–115, November 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290327U.

Domaradzki:1990:AGF

[Dom90] J. Andrzej Domaradzki. An analytic Green’s functions method in pseudo-spectral Navier–Stokes solvers for boundary layer and channel flows. Journal of Computational Physics, 88(1):232–242, May 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919090249Z.

Don:1994:NSP

[Don94] Wai Sun Don. Numerical study of pseudospectral methods in shock wave applications. Journal of Computational Physics, 110(1):103–111, January 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999184710084.

Dauber-Osguthorpe:1999:LFM

[DOOSM99] Pnina Dauber-Osguthorpe, David J. Osguthorpe, Peter S. Stern, and John Moult. Low frequency motion in proteins: Comparison of normal mode and molecular dynamics of streptomyces griseus protease A. Journal of Computational Physics, 151(1):169–189, May 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999199962320.

Dormy:1999:ACT

[Dor99] Emmanuel Dormy. An accurate compact treatment of pressure for colocated variables. Journal of Computational Physics, 151(2):676–683, May 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999199962137.

Duraiswami:1991:OMT

[DP91] Ramani Duraiswami and Andrea Prosperetti. Orthogonal mapping in two dimensions. Journal of Computational Physics, 97(2): 580, December 1991. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919190018G.

DeZeeuw:1992:ARC

[DP92a] Darren De Zeeuw and Kenneth G. Powell. An adaptivelyrefined Cartesian mesh solver for the Euler equations. Journal of Computational Physics, 101(2):453–454, August 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290033U.

Duraiswami:1992:OMTa

[DP92b] Ramani Duraiswami and Andrea Prosperetti. Orthogonal mapping in two dimensions. Journal of Computational Physics, 98(1): 179, January 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999192901896.

Duraiswami:1992:OMTb

[DP92c]  Ramani Duraiswami and Andrea Prosperetti.  Orthogo-

nal mapping in two dimensions. Journal of Computational Physics, 98(2):254–268, February 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290141K.

DeZeeuw:1993:ARC

[DP93] Darren DeZeeuw and Kenneth G. Powell. An adaptively refined Cartesian mesh solver for the Euler equations. Journal of Computational Physics, 104(1):56–68, January 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710077.

Don:1995:NSS

[DQ95] Wai Sun Don and Carl B. Quillen. Numerical simulation of shock-cylinder interactions: I. Resolution. Journal of Computational Physics, 122(2):244–265, December 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185712119.

Delaurentis:1990:MCM

[DR90] J. M. Delaurentis and L. A. Romero. A Monte Carlo method for Poisson’s equation. Journal of Computational Physics, 90(1):123– 140, September 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919090199B.

Deakin:1996:SBC

[DR96] A. S. Deakin and H. Rasmussen. Sparse boundary conditions on artificial boundaries for three-dimensional potential problems. Journal of Computational Physics, 129(1):111–120, November 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196902375.

Drake:1990:VFM

[Dra90] Douglas J. Drake. View-factor method for solving timedependent radiation transport problems involving fixed surfaces with intervening, participating media. Journal of Computational Physics, 87(1):73–90, March 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090226Q.

Dritschel:1997:ICD

[Dri97] David G. Dritschel. Introduction to “Contour Dynamics for the Euler Equations in Two Dimensions”. Journal of Computational Physics, 135(2):217–219, August 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957466.

Drukker:1999:BSH

[Dru99] Karen Drukker. Basics of surface hopping in mixed quantum/classical simulations. Journal of Computational Physics, 153(2):225–272, August 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999199962873.

Dueck:1990:TAG

[DS90] Gunter Dueck and Tobias Scheuer. Threshold accepting: a general purpose optimization algorithm appearing superior to simulated annealing. Journal of Computational Physics, 90(1):161– 175, September 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919090201B.

Dutt:1994:GCM

[DS94] P. Dutt and A. K. Singh. The Galerkin-collocation method for hyperbolic initial boundary value problems. Journal of Computational Physics, 112(2):211–225, June 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918471093X.

Darmofal:1996:IEL

[DS96] D. L. Darmofal and P. J. Schmid. The importance of eigenvectors for local preconditioners of the Euler equations. Journal of Computational Physics, 127(2):346–362, September 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901801.

Datta:1997:FAT

[DS97a] Sujay Datta and Deshdeep Sahdev. Fast algorithms for triangular Josephson junction arrays. Journal of Computational Physics, 132(2):276–284, April 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919695620X.

Domelevo:1997:NMC

[DS97b] K. Domelevo and L. Sainsaulieu. A numerical method for the computation of the dispersion of a cloud of particles by a turbulent gas flow field. Journal of Computational Physics, 133(2):256– 278, May 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999197956606.

Darmofal:1999:RMA

[DS99] D. L. Darmofal and K. Siu. A robust multigrid algorithm for the Euler equations with local preconditioning and semicoarsening. Journal of Computational Physics, 151(2):728–756, May 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999199962162.

Dannelongue:1990:EDSa

[DT90a] H. H. Dannelongue and P. A. Tanguy. Efficient data structures for adaptive remeshing with the FEM. Journal of Computational Physics, 87(2):495, April 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190902697.

Dannelongue:1990:EDSb

[DT90b] H. H. Dannelongue and P. A. Tanguy. Efficient data structures for adaptive remeshing with the FEM. Journal of Computational Physics, 91(1):94–109, November 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090006M.

Dereli:1992:SSA

[DÜ92a] T. Dereli and G. Üçoluk. A study of stationary, axially symmetric space–time geometries satisfying modified double duality equations using the exterior calculus package X^{T R} for RE-DUCE. Journal of Computational Physics, 99(1):169–175, March 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999192902846.

Derili:1992:SSA

[DÜ92b] T. Derili and G. Üçoluk. A study of stationary, axially symmetric space–time geometries satisfying modified double duality equations using the exterior calculus package X^{T} R for Reduce. Journal of Computational Physics, 98(2):349, February 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290163S.

Dueck:1993:NOH

[Due93] Gunter Dueck. New optimization heuristics: The great deluge algorithm and the record-to-record travel. Journal of Computational Physics, 104(1):86–92, January 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710107.

Dukowicz:1991:VEM

[Duk91] John K. Dukowicz. Vorticity errors in multidimensional Lagrangian codes. Journal of Computational Physics, 93(1):251, March 1991. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919190085Y.

Dukowicz:1995:MER

[Duk95] John K. Dukowicz. Mesh effects for Rossby waves. Journal of Computational Physics, 119(1):188–194, June 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711266.

Durlofsky:1991:TBA

[Dur91] Louis J. Durlofsky. Triangle based adaptive stencils for the solution of hyperbolic conservation laws. Journal of Computational Physics, 94(1):252, May 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919190152B.

Durlofsky:1993:TBM

[Dur93] Louis J. Durlofsky. A triangle based mixed finite element-finite volume technique for modeling two phase flow through porous media. Journal of Computational Physics, 105(2):252–266, April 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183710727.

Dang-Vu:1993:ASP

[DVD93] H. Dang-Vu and C. Delcarte. An accurate solution of the Poisson equation by the Chebyshev collocation method. Journal of Computational Physics, 104(1):211–220, January 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710211.

Dvinsky:1990:AGG

[Dvi90] Arkady S. Dvinsky. Adaptive grid generation from harmonic maps on Riemannian manifolds. Journal of Computational Physics, 91(1):250, November 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190900202.

Dvinsky:1991:AGG

[Dvi91] Arkady S. Dvinsky. Adaptive grid generation from harmonic maps on Riemannian manifolds. Journal of Computational Physics, 95(2):450–476, August 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190285S.

Dvorak:1991:NCS

[Dvo91] Steven L. Dvorak. Numerical computation of 2D Sommerfeld integrals — decomposition of the angular integral. Journal of Computational Physics, 94(1):253–254, May 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190156F.

Degani:1993:CAT

[DW93] A. T. Degani and J. D. A. Walker. Computation of attached three-dimensional turbulent boundary layers. Journal of Computational Physics, 109(2):202–214, December 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918371212X.

Dai:1994:ARS

[DW94a] Wenlong Dai and Paul R. Woodward. An approximate Riemann solver for ideal magnetohydrodynamics. Journal of Computational Physics, 111(2):354–372, April 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710692.

Dai:1994:EPP

[DW94b] Wenlong Dai and Paul R. Woodward. Extension of the piecewise parabolic method to multidimensional ideal magnetohydrodynamics. Journal of Computational Physics, 115(2):485–514, December 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184712125.

Dai:1995:SRS

[DW95a] Wenlong Dai and Paul R. Woodward. A simple Riemann solver and high-order Godunov schemes for hyperbolic systems of conservation laws. Journal of Computational Physics, 121(1):51–65, October 1, 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999185711783.

Dolezal:1995:RHE

[DW95b] A. Dolezal and S. S. M. Wong. Relativistic hydrodynamics and essentially non-oscillatory shock capturing schemes. Journal of Computational Physics, 120(2):266–277, September 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711643.

Dai:1996:III

[DW96a] Wenlong Dai and Paul R. Woodward. Iterative implementation of an implicit-explicit hybrid scheme for hydrodynamics. Journal of Computational Physics, 124(1):217–229, March 1, 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900546.

Dai:1996:SOI

[DW96b] Wenlong Dai and Paul R. Woodward. A second-order iterative implicit-explicit hybrid scheme for hyperbolic systems of conservation laws. Journal of Computational Physics, 128(1):181–196, October 1, 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196902028.

Dai:1997:SOU

[DW97] Wenlong Dai and Paul R. Woodward. A second-order unsplit Godunov scheme for two- and three-dimensional Euler equations. Journal of Computational Physics, 134(2):261–281, July 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002199919795693X.

Dai:1998:NSN

[DW98a] Wenlong Dai and Paul R. Woodward. Numerical simulations for nonlinear heat transfer in a system of multimaterials. Journal of Computational Physics, 139(1):58–78, January 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958630.

Dai:1998:NSR

[DW98b] Wenlong Dai and Paul R. Woodward. Numerical simulations for radiation hydrodynamics. I. Diffusion limit. Journal of Computational Physics, 142(1):182–207, May 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919895940X.

Dai:1998:SFD

[DW98c] Wenlong Dai and Paul R. Woodward. A simple finite difference scheme for multidimensional magnetohydrodynamical equations. Journal of Computational Physics, 142(2):331–369, May 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198959447.

Dritschel:1991:NFN

[DZ91] David G. Dritschel and Norman J. Zabusky. A new, but flawed, numerical method for vortex patch evolution in two dimensions. Journal of Computational Physics, 93(2):481–484, April 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190197S.

DeHaas:1996:ADD

[DZ96] P. C. A. De Haas and P. J. Zandbergen. The application of domain decomposition to time-domain computations of nonlinear water waves with a panel method. Journal of Computational Physics, 129(2):332–344, December 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196902533.

Edelstein:1997:BSM

[EA97] Arieh L. Edelstein and Noam Agmon. Brownian simulation of many-particle binding to a reversible receptor array. Journal of Computational Physics, 132(2):260–275, April 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919695617X.

Ely:1994:HPC

[EB94] Jeffrey S. Ely and Gregory R. Baker. High-precision calculations of vortex sheet motion. Journal of Computational Physics, 111(2):275–281, April 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999184710631.

Eca:1996:OGG

[Eça96] Luis Eça. 2D orthogonal grid generation with boundary point distribution control. Journal of Computational Physics, 125(2): 440–453, May 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196901060.

Edwards:1996:IMA

[Edw96a] Jack R. Edwards. An implicit multigrid algorithm for computing hypersonic, chemically reacting viscous flows. Journal of Computational Physics, 123(1):84–95, January 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900078.

Edwards:1996:EAG

[Edw96b] Michael G. Edwards. Elimination of adaptive grid interface errors in the discrete cell centered pressure equation. Journal of Computational Physics, 126(2):356–372, July 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901436.

Efraimsson:1997:AIA

[Efr97] Gunilla Efraimsson. A 2D analysis of the influence of artificial viscosity terms on solutions of the Euler equations. Journal of Computational Physics, 138(1):103–120, November 15, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958289.

Ern:1995:FAM

[EG95] Alexandre Ern and Vincent Giovangigli. Fast and accurate multicomponent transport property evaluation. Journal of Computational Physics, 120(1):105–116, August 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711515.

Ern:1996:NST

[EGS96] Alexandre Ern, Vincent Giovangigli, and Mitchell D. Smooke. Numerical study of a three-dimensional chemical vapor deposition reactor with detailed chemistry. Journal of Computational Physics, 126(1):21–39, June 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999196901175.

Esselink:1993:EPIb

[EH93] K. Esselink and P. A. J. Hilbers. Efficient parallel implementation of molecular dynamics on a toroidal network. Part II. Multiparticle potentials. Journal of Computational Physics, 106(1): 108–114, May 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999183710958.

Elliott:1997:FWM

[EHM97] Frank W. Elliott, Jr., David J. Horntrop, and Andrew J. Majda. A Fourier-wavelet Monte Carlo method for fractal random fields. Journal of Computational Physics, 132(2):384–408, April 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196956478.

Eisen:1990:SCM

[EHW90] Henner Eisen, Wilhelm Heinrichs, and Kristian Witsch. Spectral collocation methods and polar coordinate singularities. Journal of Computational Physics, 91(1):251, November 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090022S.

Eisen:1991:SCM

[EHW91] Henner Eisen, Wilhelm Heinrichs, and Kristian Witsch. Spectral collocation methods and polar coordinate singularities. Journal of Computational Physics, 96(2):241–257, October 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190235D.

Ekaterinaris:1999:IHR

[Eka99] John A. Ekaterinaris. Implicit, high-resolution, compact schemes for gas dynamics and aeroacoustics. Journal of Computational Physics, 156(2):272–299, December 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919996360X.

E:1996:ECS

[EL96a] Weinan E. and Jian-Guo Liu. Essentially compact schemes for unsteady viscous incompressible flows. Journal of Computational Physics, 126(1):122–138, June 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901254.

E:1996:VBC

[EL96b] Weinan E. and Jian-Guo Liu. Vorticity boundary condition and related issues for finite difference schemes. Journal of Computational Physics, 124(2):368–382, March 15, 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900662.

Engquist:1996:NCG

[EL96c] Bjorn Engquist and Erding Luo. New coarse grid operators for highly oscillatory coefficient elliptic problems. Journal of Computational Physics, 129(2):296–306, December 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919690251X.

E:1997:FDM

[EL97a] Weinan E. and Jian-Guo Liu. Finite difference methods for 3D viscous incompressible flows in the vorticity-vector potential formulation on nonstaggered grids. Journal of Computational Physics, 138(1):57–82, November 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958150.

E:1997:FDS

[EL97b] Weinan E. and Jian-Guo Liu. Finite difference schemes for incompressible flows in the velocity-impulse density formulation. Journal of Computational Physics, 130(1):67–76, January 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196955370.

Elliott:1994:WMC

[EM94] Frank W. Elliott, Jr. and Andrew J. Majda. A wavelet Monte Carlo method for turbulent diffusion with many spatial scales.

Journal of Computational Physics, 113(1):82–111, July 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918471120X.

Elliott:1995:NAP

[EM95] Frank W. Elliott, Jr. and Andrew J. Majda. A new algorithm with plane waves and wavelets for random velocity fields with many spatial scales. Journal of Computational Physics, 117(1):146–162, March 1, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710522.

Einfeldt:1991:GTM

[EMRS91] B. Einfeldt, C. D. Munz, P. L. Roe, and B. Sjögreen. On Godunov-type methods near low densities. Journal of Computational Physics, 92(2):273–295, February 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999191902113.

Enander:1993:IIR

[Ena93] Rickard Enander. Improved implicit residual smoothing for steady state computations of first-order hyperbolic systems. Journal of Computational Physics, 107(2):291–296, August 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711447.

Elman:1998:EIS

[EO98] Howard C. Elman and Dianne P. O’Leary. Efficient iterative solution of the three-dimensional Helmholtz equation. Journal of Computational Physics, 142(1):163–181, May 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959393.

Ehrendorfer:1998:TDM

[EOS98] K. Ehrendorfer, F. Ottitsch, and H. Sockel. A time-domain method for the determination of unsteady pressure with a tube of constant cross-section. Journal of Computational Physics, 142(1): 67–79, May 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198959320.

Elghaoui:1996:SEM

[EP96] M. Elghaoui and R. Pasquetti. A spectral embedding method applied to the advection-diffusion equation. Journal of Computational Physics, 125(2):464–476, May 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901084.

Elghaoui:1999:MSB

[EP99] M. Elghaoui and R. Pasquetti. Mixed spectral-boundary element embedding algorithms for the Navier–Stokes equations in the vorticity-stream function formulation. Journal of Computational Physics, 153(1):82–100, July 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962654.

Epperlein:1994:ICD

[Epp94] E. M. Epperlein. Implicit and conservative difference scheme for the Fokker–Planck equation. Journal of Computational Physics, 112(2):291–297, June 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999184711016.

Eckhoff:1996:NSL

[ER96] Knut S. Eckhoff and Jens H. Rolfsnes. On nonsmooth solutions of linear hyperbolic systems. Journal of Computational Physics, 125 (1):1–15, April 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196900753.

Ern:1992:VVF

[ES92] A. Ern and M. D. Smooke. Vorticity-velocity formulation for three-dimensional steady compressible flows. Journal of Computational Physics, 103(1):187, November 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290337X.

Ern:1993:VVF

[ES93] A. Ern and M. D. Smooke. Vorticity-velocity formulation for three-dimensional steady compressible flows. Journal of Computational Physics, 105(1):58–71, March 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710533.

E:1994:NRS

[ES94] Weinan E. and Chi-Wang Shu. A numerical resolution study of high order essentially non-oscillatory schemes applied to incompressible flow. Journal of Computational Physics, 110(1):39– 46, January 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184710047.

Elschner:1998:NSO

[ES98] J. Elschner and G. Schmidt. Numerical solution of optimal design problems for binary gratings. Journal of Computational Physics, 146(2):603–626, November 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960715.

Esselink:1993:EPIa

[ESH93] K. Esselink, B. Smit, and P. A. J. Hilbers. Efficient parallel implementation of molecular dynamics on a toroidal network. Part I. Parallelizing strategy. Journal of Computational Physics, 106(1): 101–107, May 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999183710946.

Ergin:1998:FET

[ESM98] A. Arif Ergin, Balasubramaniam Shanker, and Eric Michielssen. Fast evaluation of three-dimensional transient wave fields using diagonal translation operators. Journal of Computational Physics, 146(1):157–180, October 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959083.

Edwards:1994:KMI

[ETFS94] W. S. Edwards, L. S. Tuckerman, R. A. Friesner, and D. C. Sorensen. Krylov methods for the incompressible Navier–Stokes equations. Journal of Computational Physics, 110(1):82–102, January 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184710072.

Efrat:1995:RAC

[ETRW95] I. Efrat, M. Tismenetsky, B. J. Rubin, and I. Webman. Robust algorithms for the calculation of full-wave electromagnetic solutions. Journal of Computational Physics, 117(1):163–170, March 1, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185710534.

Everett:1996:HPE

[Eve96] M. E. Everett. Homotopy, polynomial equations, gross boundary data, and small Helmholtz systems. Journal of Computational Physics, 124(2):431–441, March 15, 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900704.

Edlund:1997:HOP

[EVP97] Åke Edlund, Ilya Vorobeichik, and Uri Peskin. High order perturbation theory for Helmholtz/Schrödinger equations via a separable preconditioner. Journal of Computational Physics, 138(2):788–800, December 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958459.

Eyert:1996:CSM

[Eye96] V. Eyert. A comparative study on methods for convergence acceleration of iterative vector sequences. Journal of Computational Physics, 124(2):271–285, March 15, 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900595.

Eyre:1994:CSP

[Eyr94] D. Eyre. Cubic spline-projection method for two-dimensional integral equations of scattering theory. Journal of Computational Physics, 114(1):1–8, September 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711442.

Eyre:1995:SGS

[Eyr95] D. Eyre. Spline-Galerkin solution of dynamic equations for particle comminution and collection. Journal of Computational Physics, 120(2):305–315, September 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711667.

Freedman:1990:NIS

[FA90] Alex Freedman and Steven P. Auerbach. Numerically induced stochasticity. Journal of Computational Physics, 89(1):253, July 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919090135N.

Friedman:1991:NIS

[FA91] Alex Friedman and Steven P. Auerbach. Numerically induced stochasticity. Journal of Computational Physics, 93(1):171– 188, March 1991. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919190078Y.

Fedoseyev:1997:IFE

[FA97] Alexandre I. Fedoseyev and J. Iwan D. Alexander. An inverse finite element method for pure and binary solidification problems. Journal of Computational Physics, 130(2):243–255, January 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196955825.

Fabre:1992:SAE

[Fab92] Sylvie Fabre. Stability analysis of the Euler–Poisson equations. Journal of Computational Physics, 101(2):445–451, August 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290020Y.

Falcovitz:1997:TDC

[FAH97] J. Falcovitz, G. Alfandary, and G. Hanoch. A two-dimensional conservation laws scheme for compressible flows with moving boundaries. Journal of Computational Physics, 138(1):83–102, November 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999197958083.

Fedkiw:1999:NOE

[FAMO99] Ronald P. Fedkiw, Tariq Aslam, Barry Merriman, and Stanley Osher. A non-oscillatory Eulerian approach to interfaces in multimaterial flows (the ghost fluid method). Journal of Computational Physics, 152(2):457–492, July 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962368.

Fattebert:1999:FDS

[Fat99] Jean-Luc Fattebert. Finite difference schemes and block Rayleigh quotient iteration for electronic structure calculations on composite grids. Journal of Computational Physics, 149(1):75–94, February 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198961381.

Fauci:1990:IOF

[Fau90] Lisa J. Fauci. Interaction of oscillating filaments: a computational study. Journal of Computational Physics, 86(2):294–313, February 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999190901038.

Fedkiw:1999:GFM

[FAX99] Ronald P. Fedkiw, Tariq Aslam, and Shaojie Xu. The ghost fluid method for deflagration and detonation discontinuities. Journal of Computational Physics, 154(2):393–427, September 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199963209.

Falcovitz:1994:STC

[FB94] J. Falcovitz and A. Birman. A singularities tracking conservation laws scheme for compressible duct flows. Journal of Computational Physics, 115(2):431–439, December 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184712083.

Fallavollita:1993:RSC

[FBM93] M. A. Fallavollita, D. Baganoff, and J. D. McDonald. Reduction of simulation cost and error for particle simulations of rarefied flows. Journal of Computational Physics, 109(1):30–36, November 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183711964.

Floryan:1995:NTC

[FC95] J. M. Floryan and L. Czechowski. On the numerical treatment of corner singularity in the vorticity field. Journal of Computational Physics, 118(2):222–228, May 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710947.

Fusina:1997:HRC

[FCC97] R. A. Fusina, A. L. Cooper, and S. R. Chubb. High resolution computations of ocean wave spectral modulations due to two-dimensional wave-current interactions. Journal of Computational Physics, 132(2):215–225, April 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196955862.

Fogelson:1993:OSF

[FD93] Aaron L. Fogelson and Robert H. Dillon. Optimal smoothing in function-transport particle methods for diffusion problems. Journal of Computational Physics, 109(2):155–163, December 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183712088.

Fornberg:1999:FSA

[FD99] Bengt Fornberg and Tobin A. Driscoll. A fast spectral algorithm for nonlinear wave equations with linear dispersion. Journal of Computational Physics, 155(2):456–467, November 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963519.

Feng:1999:AGF

[Fen99] James Q. Feng. Application of Galerkin finite-element method with Newton iterations in computing steady-state solutions of unipolar charge currents in corona devices. Journal of Computational Physics, 151(2):969–989, May 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962290.

Fatemi:1995:NSH

[FEO95] E. Fatemi, B. Engquist, and S. Osher. Numerical solution of the high frequency asymptotic expansion for the scalar wave equation. Journal of Computational Physics, 120(1):145–155, August 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185711540.

Ferm:1990:OBCa

[Fer90a] Lars Ferm. Open boundary conditions for external flow problems. Journal of Computational Physics, 87(2):495, April 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090270B.

Ferm:1990:OBCb

[Fer90b] Lars Ferm. Open boundary conditions for external flow problems. Journal of Computational Physics, 91(1):55–70, November 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919090004K.

Ferm:1995:NRB

[Fer95] Lars Ferm. Non-reflecting boundary conditions for the steady Euler equations. Journal of Computational Physics, 122(2):307– 316, December 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999185712168.

Fey:1998:MUPa

[Fey98a] Michael Fey. Multidimensional upwinding. Part I. The method of transport for solving the Euler equations. Journal of Computational Physics, 143(1):159–180, June 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959587.

Fey:1998:MUPb

[Fey98b] Michael Fey. Multidimensional upwinding. Part II. Decomposition of the Euler equations into advection equations. Journal of Computational Physics, 143(1):181–199, June 10, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959599.

FroeseFischer:1990:SAHb

[FG90a] Charlotte Froese Fischer and W. Guo. Spline algorithms for the Hartree–Fock equation for the helium ground state. Journal of Computational Physics, 87(2):496, April 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190902745.

FroeseFischer:1990:SAH

[FG90b] Charlotte Froese Fischer and W. Guo. Spline algorithms for the Hartree–Fock equation for the helium ground state. Journal of Computational Physics, 90(2):486–496, October 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190901762.

Figotin:1997:CSS

[FG97] Alexander Figotin and Yuri A. Godin. The computation of spectra of some 2D photonic crystals. Journal of Computational Physics, 136(2):585–598, September 15, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957831.

Friedel:1997:AMR

[FGM97] Holger Friedel, Rainer Grauer, and Christiane Marliani. Adaptive mesh refinement for singular current sheets in incompressible magnetohydrodynamic flows. Journal of Computational Physics, 134(1):190–198, June 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956837.

Farouki:1994:SAE

[FH94] R. T. Farouki and S. Hamaguchi. Spline approximation of “effective” potentials under periodic boundary conditions. Journal of Computational Physics, 115(2):276–287, December 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711958.

Farin:1997:CTG

[FH97] Gerald Farin and Bernd Hamann. Current trends in geometric modeling and selected computational applications. Journal of Computational Physics, 138(1):1–15, November 15, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196956211.

Filippova:1998:GRL

[FH98a] Olga Filippova and Dieter Hänel. Grid refinement for lattice-BGK models. Journal of Computational Physics, 147(1):219–228, November 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198960892.

Florea:1998:EAU

[FH98b] Razvan Florea and Kenneth C. Hall. Eigenmode analysis of unsteady flows about airfoils. Journal of Computational Physics, 147(2):568–593, December 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961022.

Fidel:1997:HRF

[FHKZ97] B. Fidel, E. Heyman, R. Kastner, and R. W. Ziolkowski. Hybrid ray-FDTD moving window approach to pulse propagation. Journal of Computational Physics, 138(2):480–500, December 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197958277.

Frank:1997:GIC

[FHL97] Jason Frank, Weizhang Huang, and Benedict Leimkuhler. Geometric integrators for classical spin systems. Journal of Computational Physics, 133(1):160–172, May 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956722.

Figueiredo:1992:PFP

[Fig92] JoséRicardo Figueiredo. Performance of five-point differencing schemes for two-dimensional fluid transport equations. Journal of Computational Physics, 101(2):256–264, August 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290003H.

Fishelov:1990:NVS

[Fis90] Dalia Fishelov. A new vortex scheme for viscous flows. Journal of Computational Physics, 86(1):211–224, January 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090098L.

Fishelov:1994:STD

[Fis94] Dalia Fishelov. Simulation of three-dimensional turbulent flow in non-Cartesian geometry. Journal of Computational Physics, 115(2):249–266, December 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999184711934.

Fischer:1997:OSM

[Fis97] Paul F. Fischer. An overlapping Schwarz method for spectral element solution of the incompressible Navier–Stokes equations. Journal of Computational Physics, 133(1):84–101, May 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197956515.

Forrer:1998:HOB

[FJ98] Hans Forrer and Rolf Jeltsch. A higher-order boundary treatment for Cartesian-grid methods. Journal of Computational Physics, 140(2):259–277, March 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198958910.

Fortin:1994:ONR

[FJGP94] A. Fortin, M. Jardak, J. J. Gervais, and R. Pierre. Old and new results on the two-dimensional Poiseuille flow. Journal of Computational Physics, 115(2):455–469, December 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184712101.

Foidl:1990:SGL

[FK90] Ch Foidl and P. Kasperkovitz. Systematic generation of linear graphs — check and extension of the list of Uhlenbeck and Ford. Journal of Computational Physics, 89(1):246–250, July 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090125K.

Figotin:1997:BSO

[FK97] Alexander Figotin and Igor Khalfin. Bound states of a oneband model for 3D periodic medium. Journal of Computational Physics, 138(1):153–170, November 15, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958216.

Feng:1999:CSM

[FKM99] B.-F Feng, T. Kawahara, and T. Mitsui. A conservative spectral method for several two-dimensional nonlinear wave equations. Journal of Computational Physics, 153(2):467–487, August 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199962861.

Flaa:1992:NECb

[Flå92a] Tor Flå. A numerical energy conserving method for the DNLS equation. Journal of Computational Physics, 99(2):352, April 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290223L.

Flaa:1992:NECa

[Flå92b] Tor Flå. A numerical energy conserving method for the DNLS equation. Journal of Computational Physics, 101(1):71–79, July 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290043X.

Ferraro:1993:DLB

[FLD93] Robert D. Ferraro, Paulett C. Liewer, and Viktor K. Decyk. Dynamic load balancing for a 2D concurrent plasma PIC code. Journal of Computational Physics, 109(2):329–341, December 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183712210.

Fraga:1992:AMRa

[FM92] Eric S. Fraga and John Ll. Morris. An adaptive mesh refinement method for nonlinear dispersive wave equations. Journal of Computational Physics, 101(1):94–103, July 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290045Z.

Fu:1997:HOA

[FM97] Dexun Fu and Yanwen Ma. A high order accurate difference scheme for complex flow fields. Journal of Computational Physics, 134(1):1–15, June 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999196954923.

Fedkiw:1999:IFO

[FMM99] Ronald P. Fedkiw, Antonio Marquina, and Barry Merriman. An isobaric fix for the overheating problem in multimaterial compressible flows. Journal of Computational Physics, 148 (2):545–578, January 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999198961290.

Fedkiw:1997:HAN

[FMO97] Ronald P. Fedkiw, Barry Merriman, and Stanley Osher. High accuracy numerical methods for thermally perfect gas flows with chemistry. Journal of Computational Physics, 132(2):175– 190, April 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196956223.

Fedkiw:1998:ECP

[FMO98] Ronald P. Fedkiw, Barry Merriman, and Stanley Osher. Efficient characteristic projection in upwind difference schemes for hyperbolic systems: The complementary projection method. Journal of Computational Physics, 141(1):22–36, March 20, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959137.

Fornberg:1992:FDPb

[FMS92] Bengt Fornberg and Rita Meyer-Spasche. A finite difference procedure for a class of free boundary problems. Journal of Computational Physics, 102(1):72–77, September 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999105800063.

Fatemi:1993:UFD

[FO93] Emad Fatemi and Faroukh Odeh. Upwind finite difference solution of Boltzmann equation applied to electron transport in semiconductor devices. Journal of Computational Physics, 108(2):209–217, October 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711769.

Fogelson:1992:PMSb

[Fog92a] Aaron L. Fogelson. Particle-method solution of two-dimensional convection-diffusion equations. Journal of Computational Physics, 99(1):181, March 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919290300N.

Fogelson:1992:PMSa

[Fog92b] Aaron L. Fogelson. Particle-method solution of two-dimensional convection-diffusion equations. Journal of Computational Physics, 100(1):1–16, May 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919290305I.

Fornberg:1990:IPM

[For90] Bengt Fornberg. An improved pseudospectral method for initialboundary value problems. Journal of Computational Physics, 91(2):381–397, December 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919090043Z.

Forest:1992:SOLa

[For92a] Étienne Forest. Sixth-order lie group integrators. Journal of Computational Physics, 98(2):349, February 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290162R.

Forest:1992:SOLb

[For92b] Étienne Forest. Sixth-order lie group integrators. Journal of Computational Physics, 99(2):209–213, April 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290203B.

Fornberg:1992:FDPa

[For92c] Bengt Fornberg. A finite difference procedure for a class of free boundary problems. Journal of Computational Physics, 100(2): 434, June 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999192902592.

Forner:1996:PIS

[För96] Wolfgang Förner. The properties of the iterative solution of the inverse Dyson equation for the calculation of correlation corrected band structures of polymers. Journal of Computational Physics, 125(2):477–487, May 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901096.

Forrey:1997:CHF

[For97] Robert C. Forrey. Computing the hypergeometric function. Journal of Computational Physics, 137(1):79–100, October 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957946.

Fischer:1991:PSE

[FP91] Paul F. Fischer and Anthony T. Patera. Parallel spectral element solution of the Stokes problem. Journal of Computational Physics, 92(2):380–421, February 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999191902168.

Freistuhler:1992:NSR

[FP92] Heinrich Freistühler and E. Bruce Pitman. A numerical study of a rotationally degenerate hyperbolic system. Part I. The Riemann problem. Journal of Computational Physics, 100(2):306– 321, June 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290238T.

Finocchiaro:1998:NAE

[FPB98] Daniele Finocchiaro, Marco Pellegrini, and Paolo Bientinesi. On numerical approximation of electrostatic energy in 3D. Journal of Computational Physics, 146(2):707–725, November 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960806.

Frati:1993:SAA

[FPQ93] A. Frati, F. Pasquarelli, and A. Quarteroni. Spectral approximation to advection-diffusion problems by the fictitious interface method. Journal of Computational Physics, 107(2):201–212, August 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002199918371137X.

Friedman:1990:MSP

[FPRB90] A. Friedman, S. E. Parker, S. L. Ray, and C. K. Birdsall. Multi-scale particle-in-cell plasma simulation. Journal of Computational Physics, 91(1):252, November 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090027X.

Friedman:1991:MSP

[FPRB91] A. Friedman, S. E. Parker, S. L. Ray, and C. K. Birdsall. Multi-scale particle-in-cell plasma simulation. Journal of Computational Physics, 96(1):54–70, September 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190265M.

Fulk:1996:ADS

[FQ96] David A. Fulk and Dennis W. Quinn. An analysis of 1-D smoothed particle hydrodynamics kernels. Journal of Computational Physics, 126(1):165–180, June 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919690128X.

Fraga:1992:AMRb

[Fra92] Eric S. Fraga. An adaptive mesh refinement method for nonlinear dispersive wave equations. Journal of Computational Physics, 99 (2):352, April 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290225N.

Frankel:1995:SSA

[Fra95] Jay I. Frankel. Several symbolic augmented Chebyshev expansions for solving the equation of radiative transfer. Journal of Computational Physics, 117(2):350–363, March 15, 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710716.

Friedman:1990:SOI

[Fri90] Alex Friedman. A second-order implicit particle mover with adjustable damping. Journal of Computational Physics, 90(2):292– 312, October 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919090168Z.

Friedrich:1998:WEN

[Fri98] Oliver Friedrich. Weighted essentially non-oscillatory schemes for the Interpolation of mean values on unstructured grids. Journal of Computational Physics, 144(1):194–212, July 20, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959885.

Frank:1991:COB

[FS91] Paul D. Frank and Gregory R. Shubin. A comparison of optimization-based approaches for a model computational aerodynamics design problem. Journal of Computational Physics, 93 (2):487, April 1991. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919190203W.

Frank:1992:COB

[FS92] Paul D. Frank and Gregory R. Shubin. A comparison of optimization-based approaches for a model computational aerodynamics design problem. Journal of Computational Physics, 98(1):74–89, January 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919290174W.

Frohlich:1997:AWV

[FS97] Jochen Fröhlich and Kai Schneider. An adaptive waveletvaguelette algorithm for the solution of PDEs. Journal of Computational Physics, 130(2):174–190, January 15, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196955734.

Ferguson:1996:CVF

[FT96] W. J. Ferguson and I. W. Turner. A control volume finite element numerical simulation of the drying of spruce. Journal of Computational Physics, 125(1):59–70, April 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900790.

Fujii:1995:UZM

[Fuj95] Kozo Fujii. Unified zonal method based on the fortified solution algorithm. Journal of Computational Physics, 118(1):92– 108, April 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999185710820.

Fulton:1993:SIS

[Ful93] Scott R. Fulton. A semi-implicit spectral method for the anelastic equations. Journal of Computational Physics, 106(2):299–305, June 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183711095.

Furihata:1999:FDS

[Fur99] Daisuke Furihata. Finite difference schemes for ∂u∂t = (∂∂x)^{α}δGδu that inherit energy conservation or dissipation property. Journal of Computational Physics, 156(1):181–205, November 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999199963775.

Fabbri:1997:PFM

[FV97] M. Fabbri and V. R. Voller. The phase-field method in the sharp-interface limit: a comparison between model potentials. Journal of Computational Physics, 130(2):256–265, January 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196955850.

Furzeland:1990:NST

[FVZ90] R. M. Furzeland, J. G. Verwer, and P. A. Zegeling. A numerical study of three moving-grid methods for one-dimensional partial differential equations which are based on the method of lines. Journal of Computational Physics, 89(2):349–388, August 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919090148T.

Fiebig-Wittmaack:1994:PPD

[FWBS94] Melitta Fiebig-Wittmaack and Wolfgang Börsch-Supan. Positivity preserving in difference schemes for the 2D diffusive transport of atmospheric gases. Journal of Computational Physics, 115(2):524–529, December 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184712149.

Fernando:1994:PGS

[FWWD94] G. W. Fernando, M. Weinert, R. E. Watson, and J. W. Davenport. Point group symmetries and Gaussian integration. Journal of Computational Physics, 112(2):282–290, June 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711004.

Garcia-Archilla:1995:SPE

[GA95] Bosco García-Archilla. Some practical experience with the time integration of dissipative equations. Journal of Computational Physics, 122(1):25–29, November 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918571193X.

Garcia-Archilla:1996:SME

[GA96] Bosco García-Archilla. A spectral method for the equal width equation. Journal of Computational Physics, 125(2):395–402, May 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196901011.

Garcia:1998:GCE

[GA98] Alejandro L. Garcia and Berni J. Alder. Generation of the Chapman–Enskog distribution. Journal of Computational Physics, 140(1):66–70, February 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198958892.

Ganguly:1993:DOM

[GACN93] K. Ganguly, E. J. Allen, E. Coskun, and S. Nielsen. On the discrete-ordinates method via Case’s solution. Journal of Computational Physics, 107(1):66–83, July 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711253.

Gagel:1998:FTE

[Gag98] Florian Gagel. Finite-temperature evaluation of the Fermi density operator. Journal of Computational Physics, 139(2):399–405, January 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199919795871X.

Garba:1998:MSW

[Gar98] A. Garba. A mixed spectral/wavelet method for the solution of the Stokes problem. Journal of Computational Physics, 145 (1):297–315, September 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999198960314.

Ghizzo:1990:VCNa

[GB90] A. Ghizzo and P. Bertrand. A Vlasov code for the numerical simulation of stimulated Raman scattering. Journal of Computational Physics, 87(2):495, April 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190902686.

Ghattas:1997:OCT

[GB97] Omar Ghattas and Jai-Hyeong Bark. Optimal control of twoand three-dimensional incompressible Navier–Stokes flows. Journal of Computational Physics, 136(2):231–244, September 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197957442.

Ganesh:1996:MLH

[GBBH96] Ram K. Ganesh, Wallace W. Bowley, Robert R. Bellantone, and Yukap Hahn. A model for laser hole drilling in metals. Journal of Computational Physics, 125(1):161–176, April 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900868.

Garcia:1999:AMA

[GBCA99] Alejandro L. Garcia, John B. Bell, William Y. Crutchfield, and Berni J. Alder. Adaptive mesh and algorithm refinement using direct simulation Monte Carlo. Journal of Computational Physics, 154(1):134–155, September 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963052.

Ghizzo:1992:CBE

[GBL^{+}92] A. Ghizzo, P. Bertrand, J. Lebas, M. Shoucri, T. Johnston, E. Fijalkow, and M. R. Feix. Comparison between 1D and 11/2D Eulerian Vlasov codes for the numerical simulation of stimulated Raman scattering. Journal of Computational Physics, 102(2):417– 422, October 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290383A.

Ghizzo:1990:VCNb

[GBS^{+}90] A. Ghizzo, P. Bertrand, M. M. Shoucri, T. W. Johnston, E. Fualkow, and M. R. Feix. A Vlasov code for the numerical simulation of stimulated Raman scattering. Journal of Computational Physics, 90(2):431–457, October 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090174Y.

Ghizzo:1993:ECS

[GBS^{+}93] A. Ghizzo, P. Bertrand, M. Shoucri, E. Fijalkow, and M. R. Feix. An Eulerian code for the study of the drift-kinetic Vlasov equation. Journal of Computational Physics, 108(1):105–121, September 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183711678.

Gullingsrud:1999:RPM

[GBS99] Justin R. Gullingsrud, Rosemary Braun, and Klaus Schulten. Reconstructing potentials of mean force through time series analysis of steered molecular dynamics simulations. Journal of Computational Physics, 151(1):190–211, May 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962186.

Grossman:1990:FSA

[GC90] B. Grossman and P. Cinnella. Flux-split algorithms for flows with non-equilibrium chemistry and vibrational relaxation. Journal of Computational Physics, 88(1):131–168, May 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090245V.

Guo:1992:CSFb

[GC92] Ben-Yu Guo and Wei-Ming Cao. A combined spectral-finite element method for solving two-dimensional unsteady Navier– Stokes equations. Journal of Computational Physics, 101(2):375– 385, August 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290015Q.

Givoli:1995:NBC

[GC95] Dan Givoli and Dan Cohen. Nonreflecting boundary conditions based on Kirchhoff-type formulae. Journal of Computational Physics, 117(1):102–113, March 1, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710480.

Geisler:1990:SSS

[GCMR90] T. Geisler, P. L. Christiansen, J. Mørk, and P. S. Ramanujam. Split-step spectral method for nonlinear schrödinger equation with constant background intensities. Journal of Computational Physics, 86(2):492–495, February 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090111D.

Gwynllyw:1996:MSE

[GDP96] D. Rh. Gwynllyw, A. R. Davies, and T. N. Phillips. A moving spectral element approach to the dynamically loaded journal bearing problem. Journal of Computational Physics, 123(2):476– 494, February 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196900388.

Gavelek:1992:SII

[GE92] D. Gavelek and T. Erber. Shadowing and iterative interpolation for Čebyšev mixing transformations. Journal of Computational Physics, 101(1):25–50, July 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192900406.

Gonzales:1997:IEM

[GEK^{+}97] R. A. Gonzales, J. Eisert, I. Koltracht, M. Neumann, and G. Rawitscher. Integral equation method for the continuous spectrum radial Schrödinger equation. Journal of Computational Physics, 134(1):134–149, June 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956795.

Gelfgat:1999:DMR

[Gel99] Alexander Yu. Gelfgat. Different modes of Rayleigh-Bénard instability in two- and three-dimensional rectangular enclosures. Journal of Computational Physics, 156(2):300–324, December 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199963635.

Gardner:1990:SWR

[GG90] L. R. T. Gardner and G. A. Gardner. Solitary waves of the regularised long-wave equation. Journal of Computational Physics, 91(2):441–459, December 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190900475.

Gardner:1992:SWE

[GG92a] L. R. T. Gardner and G. A. Gardner. Solitary waves of the equal width wave equation. Journal of Computational Physics, 101(1): 218–223, July 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999192900543.

Grinstein:1992:EVS

[GG92b] F. F. Grinstein and R. H. Guirguis. Effective viscosity in the simulation of spatially evolving shear flows with monotonic FCT models. Journal of Computational Physics, 101(1):165–175, July 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999192900509.

Gharakhani:1997:TDV

[GG97] Adrin Gharakhani and Ahmed F. Ghoniem. Three-dimensional vortex simulation of time dependent incompressible internal viscous flows. Journal of Computational Physics, 134(1):75–95, June 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197956679.

Gardner:1994:BFM

[GGG94] L. R. T. Gardner, G. A. Gardner, and T. Geyikli. The boundary forced MKdV equation. Journal of Computational Physics, 113 (1):5–12, July 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184711132.

Goodrich:1990:HBD

[GGH90] John W. Goodrich, Karl Gustafson, and Kadosa Halasi. Hopf bifurcation in the driven cavity. Journal of Computational Physics, 90(1):219–261, September 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090204E.

Goel:1995:EFIa

[GGK95] Narenda S. Goel, Fengshi Gang, and Zhonglin Ko. Electrostatic field in inhomogeneous dielectric media I. Indirect boundary element method. Journal of Computational Physics, 118(1):172– 179, April 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999185710881.

Greenbaum:1993:LED

[GGM93] A. Greenbaum, L. Greengard, and G. B. McFadden. Laplace’s equation and the Dirichlet–Neumann map in multiply connected domains. Journal of Computational Physics, 105(2):267– 278, April 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999183710739.

Gelb:1997:WSR

[GGP97] Anne Gelb, David Gottlieb, and Nathan Paldor. Wind setdown relaxation on a sloping beach. Journal of Computational Physics, 138(2):644–664, December 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919795837X.

Gardner:1993:CEP

[GGZ93] L. R. T. Gardner, G. A. Gardner, and S. I. Zaki. Collisional effects in plasmas modelled by a simplified Fokker–Planck equation. Journal of Computational Physics, 107(1):40–50, July 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918371123X.

Gibbons:1995:DDI

[GH95] Matthew R. Gibbons and Dennis W. Hewett. The Darwin direct implicit particle-in-cell (DADIPIC) method for simulation of low frequency plasma phenomena. Journal of Computational Physics, 120(2):231–247, September 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711606.

Gibbons:1997:CDD

[GH97] Matthew R. Gibbons and Dennis W. Hewett. Characterization of the Darwin direct implicit particle-in-cell method and resulting guidelines for operation. Journal of Computational Physics, 130(1):54–66, January 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196954753.

Guevremont:1993:FES

[GHKH93] G. Guevremont, W. G. Habashi, P. L. Kotiuga, and M. M. Hafez. Finite element solution of the 3D compressible Navier– Stokes equations by a velocity-vorticity method. Journal of Computational Physics, 107(1):176–187, July 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711344.

Ghosal:1996:ANE

[Gho96] Sandip Ghosal. An analysis of numerical errors in large-eddy simulations of turbulence. Journal of Computational Physics, 125(1): 187–206, April 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196900881.

Goldstein:1993:MNS

[GHS93a] D. Goldstein, R. Handler, and L. Sirovich. Modeling a noslip flow boundary with an external force field. Journal of Computational Physics, 105(2):354–366, April 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710818.

Guermond:1993:SEV

[GHS93b] J.-L. Guermond, S. Huberson, and W.-Z. Shen. Simulation of 2D external viscous flows by means of a domain decomposition method. Journal of Computational Physics, 108(2):343– 352, October 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999183711873.

Giles:1997:SAG

[Gil97] M. B. Giles. Stability analysis of a Galerkin/Runge–Kutta Navier–Stokes discretisation on unstructured tetrahedral grids. Journal of Computational Physics, 132(2):201–214, April 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0021999196956168.

Giraldo:1997:LGM

[Gir97] Francis X. Giraldo. Lagrange–Galerkin methods on spherical geodesic grids. Journal of Computational Physics, 136 (1):197–213, September 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999197957715.

Giraldo:1998:LGS

[Gir98] Francis X. Giraldo. The Lagrange–Galerkin spectral element method on unstructured quadrilateral grids. Journal of Computational Physics, 147(1):114–146, November 20, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960788.

Givoli:1990:NRB

[Giv90] Dan Givoli. Non-reflecting boundary conditions: a review. Journal of Computational Physics, 90(1):270, September 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090219Q.

Givoli:1991:NRB

[Giv91] Dan Givoli. Non-reflecting boundary conditions. Journal of Computational Physics, 94(1):1–29, May 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999191901358.

Giannakouros:1994:SEF

[GK94] John Giannakouros and George Em Karniadakis. A spectral element-FCT method for the compressible Euler equations. Journal of Computational Physics, 115(1):65–85, November 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002199918471179X.

Grote:1995:NBC

[GK95] Marcus J. Grote and Joseph B. Keller. On nonreflecting boundary conditions. Journal of Computational Physics, 122(2):231–243, December 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999185712107.

Grote:1996:NBC

[GK96] Marcus J. Grote and Joseph B. Keller. Nonreflecting boundary conditions for time-dependent scattering. Journal of Computational Physics, 127(1):52–65, August 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901576.

Grote:1998:NBC

[GK98] Marcus J. Grote and Joseph B. Keller. Nonreflecting boundary conditions for Maxwell’s equations. Journal of Computational Physics, 139(2):327–342, January 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958812.

Glikman:1996:OAM

[GKDT96] E. Glikman, I. Kelson, N. V. Doan, and H. Tietze. An optimized algorithm for molecular dynamics simulation of large-scale systems. Journal of Computational Physics, 124(1):85–92, March 1, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196900467.

Goel:1995:EFIb

[GKG95] Narendra S. Goel, Zhonglin Ko, and Fengshi Gang. Electrostatic field in inhomogeneous dielectric media II. Volume element method. Journal of Computational Physics, 118(1):180– 188, April 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999185710893.

Gonzales:1999:IEM

[GKKR99] R. A. Gonzales, S.-Y. Kang, I. Koltracht, and G. Rawitscher. Integral equation method for coupled Schrödinger equations. Journal of Computational Physics, 153(1):160–202, July 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199962721.

Gropp:1996:NSV

[GKL^{+}96] William D. Gropp, Hans G. Kaper, Gary K. Leaf, David M. Levine, Mario Palumbo, and Valerii M. Vinokur. Numerical simulation of vortex dynamics in type-II superconductors. Journal of Computational Physics, 123(2):254–266, February 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900224.

Greengard:1996:IEM

[GKM96] Leslie Greengard, Mary Catherine Kropinski, and Anita Mayo. Integral equation methods for Stokes flow and isotropic elasticity in the plane. Journal of Computational Physics, 125(2):403–414, May 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196901023.

Gupta:1997:CMS

[GKZ97a] Murli M. Gupta, Jules Kouatchou, and Jun Zhang. A compact multigrid solver for convection-diffusion equations. Journal of Computational Physics, 132(1):123–129, March 15, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196956272.

Gupta:1997:CSF

[GKZ97b] Murli M. Gupta, Jules Kouatchou, and Jun Zhang. Comparison of second- and fourth-order discretizations for multigrid Poisson solvers. Journal of Computational Physics, 132(2):226–232, April 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196954662.

Ghattas:1995:VFE

[GL95] Omar Ghattas and Xiaogang Li. A variational finite element method for stationary nonlinear fluid-solid interaction. Journal of Computational Physics, 121(2):347–356, October 15, 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919590204X.

Greengard:1996:DAP

[GL96] Leslie Greengard and June-Yub Lee. A direct adaptive Poisson solver of arbitrary order accuracy. Journal of Computational Physics, 125(2):415–424, May 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901035.

Glaister:1991:RSB

[Gla91] P. Glaister. A Riemann solver for “barotropic” row. Journal of Computational Physics, 93(2):477–480, April 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190196R.

Glassey:1992:ASZb

[Gla92a] R. T. Glassey. Approximate solutions to the Zakharov equations via finite differences. Journal of Computational Physics, 99(2): 351, April 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290219O.

Glassey:1992:ASZa

[Gla92b] R. T. Glassey. Approximate solutions to the Zakharov equations via finite differences. Journal of Computational Physics, 100(2): 377–383, June 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290243R.

Glaister:1995:WFR

[Gla95] P. Glaister. A weak formulation of Roe’s approximate Riemann solver applied to the St. Venant equations. Journal of Computational Physics, 116(1):189–191, January 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710182.

Garbet:1990:VCE

[GLM^{+}90] X. Garbet, L. Laurent, F. Mourgues, J. P. Roubin, and A. Samain. Variational calculation of electromagnetic instabilities in tokamaks. Journal of Computational Physics, 87(2):249– 269, April 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919090253W.

Gueyffier:1999:VFI

[GLN^{+}99] Denis Gueyffier, Jie Li, Ali Nadim, Ruben Scardovelli, and Stéphane Zaleski. Volume-of-fluid interface tracking with smoothed surface stress methods for three-dimensional flows. Journal of Computational Physics, 152(2):423–456, July 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002199919896168X.

Glowinski:1992:EWP

[Glo92] R. Glowinski. Ensuring well-posedness by analogy: Stokes problem and boundary control for the wave equation. Journal of Computational Physics, 103(2):189–221, December 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290396G.

Granek:1992:SAN

[GM92] Henry Granek and Bruce H. J. McKellar. Stability analysis of a non-linear diffusion-type kinetic equation. Journal of Computational Physics, 100(2):253–269, June 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290233O.

Ghosal:1995:BEL

[GM95a] S. Ghosal and P. Moin. The basic equations for the large eddy simulation of turbulent flows in complex geometry. Journal of Computational Physics, 118(1):24–37, April 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710777.

Gray:1995:NIA

[GM95b] David B. Gray and George A. McMechan. Numerical investigation of an analytic solution of a multi-dimensional Lippman–Schwinger seismic inverse problem. Journal of Computational Physics, 119(1):195–205, June 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711278.

Giraud:1996:PIE

[GM96] L. Giraud and G. Manzini. Parallel implementations of 2D explicit Euler solvers. Journal of Computational Physics, 123(1):111–118, January 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999196900091.

Greenbaum:1998:RPE

[GM98] A. Greenbaum and A. Mayo. Rapid parallel evaluation of integrals in potential theory on general three-dimensional regions. Journal of Computational Physics, 145(2):731–742, September 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002199919896048X.

Guo:1992:FCS

[GMCH92] Ben-Yu Guo, He-Ping Ma, Wei-Ming Cao, and Hui Huang. The Fourier–Chebyshev spectral method for solving twodimensional unsteady vorticity equations. Journal of Computational Physics, 101(1):207–217, July 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192900532.

Georghiou:1999:IFE

[GMM99] G. E. Georghiou, R. Morrow, and A. C. Metaxas. An improved finite-element flux-corrected transport algorithm. Journal of Computational Physics, 148(2):605–620, January 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961320.

Guillard:1992:ASMa

[GMP92a] H. Guillard, J. M. Male, and R. Peyret. Adaptive spectral methods with application to mixing layer computations. Journal of Computational Physics, 100(2):435, June 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192902684.

Guillard:1992:ASMb

[GMP92b] H. Guillard, J. M. Malé, and R. Peyret. Adaptive spectral methods with application to mixing layer computations. Journal of Computational Physics, 102(1):114–127, September 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999105800105.

Guo:1996:NSM

[GMPGV96] Ben-Yu Guo, I. Martín, Victor M. Pérez-García, and Luis Vázquez. Numerical solutions of the Maxwell–Bloch laser equations. Journal of Computational Physics, 129(1):181–189, November 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196902429.

Garnier:1999:USC

[GMS^{+}99] Eric Garnier, Michele Mossi, Pierre Sagaut, Pierre Comte, and Michel Deville. On the use of shock-capturing schemes for large-eddy simulation. Journal of Computational Physics, 153(2):273–311, August 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199919996268X.

Giannattasio:1996:OVC

[GN96] P. Giannattasio and M. Napolitano. Optimal vorticity conditions for the node-centred finite-difference discretization of the second-order vorticity-velocity equations. Journal of Computational Physics, 127(1):208–217, August 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901692.

Green:1996:SDA

[GNH96] Lawrence L. Green, Perry A. Newman, and Kara J. Haigler. Sensitivity derivatives for advanced CFD algorithm and viscous modeling parameters via automatic differentiation. Journal of Computational Physics, 125(2):313–324, May 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900960.

Garcia-Navarro:1995:GMU

[GNHP95] P. Garcia-Navarro, M. E. Hubbard, and A. Priestley. Genuinely multidimensional upwinding for the 2D shallow water equations. Journal of Computational Physics, 121(1):79–93, October 1, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185711801.

Gustafsson:1995:FOD

[GO95] Bertil Gustafsson and Pelle Olsson. Fourth-order difference methods for hyperbolic IBVPs. Journal of Computational Physics, 117(2):300–317, March 15, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710686.

Gerritsen:1996:DES

[GO96] Margot Gerritsen and Pelle Olsson. Designing an efficient solution strategy for fluid flows: 1. A stable high order finite difference scheme and sharp shock resolution for the Euler equations. Journal of Computational Physics, 129(2):245–262, December 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002199919690248X.

Gerritsen:1998:DES

[GO98] Margot Gerritsen and Pelle Olsson. Designing an efficient solution strategy for fluid flows: II. Stable high-order central finite difference schemes on composite adaptive grids with sharp shock resolution. Journal of Computational Physics, 147(2):293–317, December 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198960569.

Guseinov:1995:CCG

[GÖAY95] I. I. Guseinov, A. Özmen, Ü. Atav, and H. Yüksel. Computation of Clebsch–Gordan and Gaunt coefficients using binomial coefficients. Journal of Computational Physics, 122(2):343–347, December 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199918571220X.

Godunov:1999:RAD

[God99] Sergei Konstantinovich Godunov. Reminiscences about difference schemes. Journal of Computational Physics, 153(1):6–25, July 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002199919996271X. See introduction [Van99].

Goedecker:1995:LCA

[Goe95] S. Goedecker. Low complexity algorithms for electronic structure calculations. Journal of Computational Physics, 118(2):261–268, May 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185710972.

Goto:1992:ESCb

[Got92] E. Goto. Extrapolated surface charge method for capacity calculation of polygons and polyhedra. Journal of Computational Physics, 99(1):182, March 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290304H.

Gottelmann:1999:SCS

[Göt99] Jochen Göttelmann. A spline collocation scheme for the spherical shallow water equations. Journal of Computational Physics, 148(1):291–298, January 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961113.

Gousheh:1996:EAS

[Gou96a] Siamak S. Gousheh. An efficient algorithm for solving coupled Schrödinger type ODE’s, whose potentials include δ-functions. Journal of Computational Physics, 123(1):162–168, January 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196900133.

Goussis:1996:CUR

[Gou96b] D. A. Goussis. On the construction and use of reduced chemical kinetic mechanisms produced on the basis of given algebraic relations. Journal of Computational Physics, 128(2):261–273, October 15, 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196902090.

Gambolati:1998:CSN

[GP98a] Giuseppe Gambolati and Giorgio Pini. Complex solution to nonideal contaminant transport through porous media. Journal of Computational Physics, 145(2):538–554, September 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198960387.

Givoli:1998:FES

[GP98b] Dan Givoli and Igor Patlashenko. Finite-element solution of nonlinear time-dependent exterior wave problems. Journal of Computational Physics, 143(1):241–258, June 10, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198999995.

Glowinski:1996:WFE

[GPWZ96] Roland Glowinski, T. W. Pan, Raymond O. Wells, Jr., and Xiaodong Zhou. Wavelet and finite element solutions for the Neumann problem using fictitious domains. Journal of Computational Physics, 126(1):40–51, June 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901187.

Guermond:1997:CIV

[GQ97] J.-L. Guermond and L. Quartapelle. Calculation of incompressible viscous flows by an unconditionally stable projection FEM. Journal of Computational Physics, 132(1):12–33, March 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196955874.

Greengard:1997:FAP

[GR97] L. Greengard and V. Rokhlin. A fast algorithm for particle simulations. Journal of Computational Physics, 135(2):280–292, August 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197957065.

Grandjouan:1990:MEA

[Gra90] N. Grandjouan. The modified equation approach to fluxcorrected transport. Journal of Computational Physics, 91 (2):424–440, December 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999190900464.

Graziani:1995:PFA

[Gra95] Frank R. Graziani. The product formula algorithm applied to linear and radiation diffusion. Journal of Computational Physics, 118(1):9–23, April 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999185710765.

Grandy:1999:CRR

[Gra99] Jeffrey Grandy. Conservative remapping and region overlays by intersecting arbitrary polyhedra. Journal of Computational Physics, 148(2):433–466, January 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961253.

Ghizzo:1995:EVH

[GRB^{+}95] A. Ghizzo, T. Réveillé, P. Bertrand, T. W. Johnston, J. Lebas, and M. Shoucri. An Eulerian Vlasov–Hilbert code for the numerical simulation of the interaction of high-frequency electromagnetic waves with plasma. Journal of Computational Physics, 118(2):356–365, May 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999185711059.

Greenspan:1990:CUE

[Gre90] Donald Greenspan. A counterexample of the use of energy as a measure of computational accuracy. Journal of Computational Physics, 91(2):490–494, December 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190900512.

Greene:1991:LTD

[Gre91] John M. Greene. Locating three-dimensional roots by a bisection method. Journal of Computational Physics, 94(1):253, May 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190155E.

Greene:1992:LTDa

[Gre92a] John M. Greene. Locating three-dimensional roots by a bisection method. Journal of Computational Physics, 98(1):178, January 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999192901863.

Greene:1992:LTDb

[Gre92b] John M. Greene. Locating three-dimensional roots by a bisection method. Journal of Computational Physics, 98(2):194–198, February 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290137N.

Greer:1994:ATE

[Gre94]  J. C. Greer.  An approximate time evolution operator

to generate the Verlet algorithm. Journal of Computational Physics, 115(1):245–247, November 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711909.

Greer:1998:MCC

[Gre98] J. C. Greer. Monte Carlo configuration interaction. Journal of Computational Physics, 146(1):181–202, October 10, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959538.

Grinstein:1994:OBC

[Gri94] Fernando F. Grinstein. Open boundary conditions in the simulation of subsonic turbulent shear flows. Journal of Computational Physics, 115(1):43–55, November 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711776.

Grossman:1991:EST

[Gro91] C. Grossman. Enclosures of the solution of the Thomas– Fermi equations by monotone discretization. Journal of Computational Physics, 93(1):252, March 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999191900883.

Grossmann:1992:EST

[Gro92] C. Grossmann. Enclosures of the solution of the Thomas– Fermi equation by monotone discretization. Journal of Computational Physics, 98(1):26–32, January 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192901704.

Guj:1993:VVM

[GS93] G. Guj and F. Stella. A vorticity-velocity method for the numerical solution of 3D incompressible flows. Journal of Computational Physics, 106(2):286–298, June 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711083.

Gomez:1994:GSS

[GS94] F. J. Gomez and J. Sesma. Global solutions of the Schrödinger equation with Coulomb plus linear potential. Journal of Computational Physics, 115(2):296–301, December 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711971.

Gill:1995:CMM

[GS95] Andrew W. Gill and G. E. Sneddon. Complex mapped matrix methods in hydrodynamic stability problems. Journal of Computational Physics, 122(1):13–24, November 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711928.

Gill:1996:PMC

[GS96a] Andrew W. Gill and G. E. Sneddon. Pseudospectral methods and composite complex maps for near-boundary critical points. Journal of Computational Physics, 129(1):1–7, November 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196902296.

Gueron:1996:SIM

[GS96b] Shay Gueron and David Shalloway. Spatial Interpolation methods for integrating Newton’s equation. Journal of Computational Physics, 129(1):87–100, November 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196902351.

Gaitonde:1997:OCD

[GS97a] Datta Gaitonde and J. S. Shang. Optimized compact-difference-based finite-volume schemes for linear wave phenomena. Journal of Computational Physics, 138(2):617–643, December 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958368.

Gutfraind:1997:SPH

[GS97b] Ricardo Gutfraind and Stuart B. Savage. Smoothed particle hydrodynamics for the simulation of broken-ice fields: Mohr– Coulomb-type rheology and frictional boundary conditions. Journal of Computational Physics, 134(2):203–215, July 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956813.

Ghizzo:1993:TTP

[GSB^{+}93] A. Ghizzo, M. Shoucri, P. Bertrand, T. Johnston, and J. Lebas. Trajectories of trapped particles in the field of a plasma wave excited by a stimulated Raman scattering. Journal of Computational Physics, 108(2):373–376, October 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711915.

Gerlinger:1998:IMM

[GSB98] P. Gerlinger, P. Stoll, and D. Brüggemann. An implicit multigrid method for the simulation of chemically reacting flows. Journal of Computational Physics, 146(1):322–345, October 10, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960612.

Goto:1992:ESCa

[GSY92] E. Goto, Y. Shi, and N. Yoshida. Extrapolated surface charge method for capacity calculation of polygons and polyhedra. Journal of Computational Physics, 100(1):105–115, May 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290313N.

Garaizar:1997:FTS

[GT97] F. Xabier Garaizar and John Trangenstein. Front tracking for shear bands in an antiplane shear model. Journal of Computational Physics, 131(1):54–69, February 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919695456X.

Gottlieb:1999:AEA

[GTA99] D. Gottlieb, E. Turkel, and S. Abarbanel. Analysis of the error for approximations to systems of hyperbolic equations. Journal of Computational Physics, 151(2):997–1007, May 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919996196X.

Garbey:1998:NPS

[GTD98] M. Garbey and D. Tromeur-Dervout. A new parallel solver for the nonperiodic incompressible Navier–Stokes equations with a Fourier method: Application to frontal polymerization. Journal of Computational Physics, 145(1):316–331, September 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002199919896017X.

Gueron:1994:MFC

[Gue94] Shay Gueron. Methods for fast computation of integral transforms. Journal of Computational Physics, 110(1):164–170, January 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999184710126.

Gupta:1990:HAS

[Gup90] Murli M. Gupta. High accuracy solutions of incompressible Navier–Stokes equations. Journal of Computational Physics, 89(2):488–489, August 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919090157V.

Gupta:1991:HAS

[Gup91] Murli M. Gupta. High accuracy solutions of incompressible Navier–Stokes equations. Journal of Computational Physics, 93(2):343–359, April 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919190188Q.

Gressier:1999:PFV

[GVM99] Jérémie Gressier, Philippe Villedieu, and Jean-Marc Moschetta. Positivity of flux vector splitting schemes. Journal of Computational Physics, 155(1):199–220, October 10, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963374.

Gomez:1992:ESF

[GWI92] R. Gómez, J. Winicour, and R. Isaacson. Evolution of scalar fields from characteristic data. Journal of Computational Physics, 98(1):11–25, January 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290169Y.

Gong-Yan:1991:FMS

[GY91] Lei Gong-Yan. A family of methods for the solution of lattice models. Journal of Computational Physics, 92(1):106–141, January 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190294U.

Guo:1995:OBC

[GZ95] Dong-Jian Guo and Qing-Cun Zeng. Open boundary conditions for a numerical shelf sea model. Journal of Computational Physics, 116(1):97–102, January 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918571008X.

Herbst:1993:NCS

[HA93] B. M. Herbst and Mark J. Ablowitz. Numerical chaos, symplectic integrators, and exponentially small splitting distances. Journal of Computational Physics, 105(1):122–132, March 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710582.

Hirt:1997:ALE

[HAC97] C. W. Hirt, A. A. Amsden, and J. L. Cook. An arbitrary Lagrangian–Eulerian computing method for all flow speeds. Journal of Computational Physics, 135(2):203–216, August 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957028.

Hadjiconstantinou:1999:HAC

[Had99] Nicolas G. Hadjiconstantinou. Hybrid atomistic-continuum formulations and the moving contact-line problem. Journal of Computational Physics, 154(2):245–265, September 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963027.

Hagmeijer:1994:GAB

[Hag94] R. Hagmeijer. Grid adaption based on modified anisotropic diffusion equations formulated in the parametric domain. Journal of Computational Physics, 115(1):169–183, November 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711855.

Hager:1998:DML

[Hag98] William W. Hager. A discrete model for the lightning discharge. Journal of Computational Physics, 144(1):137–150, July 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002199919895985X.

Hall:1990:CFSa

[Hal90a] M. S. Hall. A comparison of first- and second-order rezoned and Lagrangian Godunov solutions. Journal of Computational Physics, 87(2):496, April 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999190902712.

Hall:1990:CFSb

[Hal90b] M. S. Hall. A comparison of first and second order rezoned and Lagrangian Godunov solutions. Journal of Computational Physics, 90(2):458–485, October 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090175Z.

Hallberg:1997:SST

[Hal97] Robert Hallberg. Stable split time stepping schemes for largescale ocean modeling. Journal of Computational Physics, 135(1): 54–65, July 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199919795734X.

Hanley:1993:SES

[Han93a] Patrick Hanley. A strategy for the efficient simulation of viscous compressible flows using a multi-domain pseudospectral method. Journal of Computational Physics, 108(1):153–158, September 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183711708.

Hansbo:1993:ESD

[Han93b] Peter Hansbo. Explicit streamline diffusion finite element methods for the compressible Euler equations in conservation variables. Journal of Computational Physics, 109(2):274–288, December 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999183712179.

Harabetian:1992:SRMa

[Har92a] Eduard Harabetian. A subcell resolution method for viscous systems of conservation laws. Journal of Computational Physics, 101 (1):228, July 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999192900613.

Harabetian:1992:SRMb

[Har92b] Eduard Harabetian. A subcell resolution method for viscous systems of conservation laws. Journal of Computational Physics, 103(2):350–358, December 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290406O.

Harten:1994:AMS

[Har94] Ami Harten. Adaptive multiresolution schemes for shock computations. Journal of Computational Physics, 115(2):319–338, December 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184711995.

Harten:1997:HRS

[Har97] Ami Harten. High resolution schemes for hyperbolic conservation laws. Journal of Computational Physics, 135(2):260–278, August 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197957132.

Hirshman:1990:PDA

[HB90a] S. P. Hirshman and O. Betancourt. Preconditioned descent algorithm for rapid calculations of magnetohydrodynamic equilibria. Journal of Computational Physics, 91(1):249, November 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919090017U.

Huizing:1990:SFI

[HB90b] A. J. Huizing and B. L. G. Bakker. The solution of Faddeev integral equations for three-body scattering by means of B-splines. Journal of Computational Physics, 90(1):200–218, September 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919090203D.

Hirshman:1991:PDA

[HB91] S. P. Hirshman and O. Betancourt. Preconditioned descent algorithm for rapid calculations of magnetohydrodynamic equilibria. Journal of Computational Physics, 96(1):99–109, September 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190267O.

Heemink:1995:RWM

[HB95] Arnold W. Heemink and Pieter A. Blokland. On random walk models with space varying diffusivity. Journal of Computational Physics, 119(2):388–389, July 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711436.

Higdon:1996:SAO

[HB96] Robert L. Higdon and Andrew F. Bennett. Stability analysis of operator splitting for large-scale ocean modeling. Journal of Computational Physics, 123(2):311–329, February 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900261.

Hubbard:1997:CMU

[HB97] M. E. Hubbard and M. J. Baines. Conservative multidimensional upwinding for the steady two-dimensional shallow water equations. Journal of Computational Physics, 138(2):419–448, December 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199919795823X.

Hansson:1993:TMS

[HBF93] L. Cruzeiro Hansson, J. O. Baum, and J. L. Finney. Two modified staging algorithms for path integral Monte Carlo simulations. Journal of Computational Physics, 104(1):110–117, January 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183710132.

Husain:1993:NSV

[HC93] I. Husain and O. P. Chandna. Numerical simulation of viscoelastic fluid flow past a cylinder on a streamfunction coordinate system. Journal of Computational Physics, 108(2):226–235, October 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183711782.

Hahn:1997:USJ

[HC97] Seonghyeon Hahn and Haecheon Choi. Unsteady simulation of jets in a cross flow. Journal of Computational Physics, 134(2):342–356, July 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999197957120.

Haddad:1991:VSP

[HCC91]  B. Haddad, F. Clausset, and F. Combes.  Vectorising

the smooth particle hydrodynamics. Journal of Computational Physics, 97(1):103–126, November 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190041I.

He:1998:NTM

[HCD98] Xiaoyi He, Shiyi Chen, and Gary D. Doolen. A novel thermal model for the lattice Boltzmann method in incompressible limit. Journal of Computational Physics, 146(1):282–300, October 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198960570.

Harley:1992:CRMa

[HCJ92a] T. R. Harley, C. Z. Cheng, and S. C. Jardin. The computation of resistive MHD instabilities in axisymmetric toroidal plasmas. Journal of Computational Physics, 101(2):452, August 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290022Q.

Harley:1992:CRMb

[HCJ92b] T. R. Harley, C. Z. Cheng, and S. C. Jardin. The computation of resistive MHD instabilities in axisymmetric toroidal plasmas. Journal of Computational Physics, 103(1):43–62, November 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290325S.

Hunter:1993:URW

[HCP93] J. R. Hunter, P. D. Craig, and H. E. Phillips. On the use of random walk models with spatially variable diffusivity. Journal of Computational Physics, 106(2):366–376, June 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711149.

He:1999:LBS

[HCZ99] Xiaoyi He, Shiyi Chen, and Raoyang Zhang. A lattice Boltzmann scheme for incompressible multiphase flow and its application in simulation of Rayleigh–Taylor instability. Journal of Computational Physics, 152(2):642–663, July 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962575.

He:1997:LBM

[HD97] Xiaoyi He and Gary Doolen. Lattice Boltzmann method on curvilinear coordinates system: Flow around a circular cylinder. Journal of Computational Physics, 134(2):306–315, July 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957090.

Hesthaven:1999:SCT

[HDL99] J. S. Hesthaven, P. G. Dinesen, and J. P. Lynov. Spectral collocation time-domain modeling of diffractive optical elements. Journal of Computational Physics, 155(2):287–306, November 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199963337.

Higdon:1997:BBT

[HdS97] Robert L. Higdon and Roland A. de Szoeke. Barotropicbaroclinic time splitting for ocean circulation modeling. Journal of Computational Physics, 135(1):30–53, July 15, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957338.

Heinrichs:1992:SMM

[Hei92] Wilhelm Heinrichs. A spectral multigrid method for the Stokes problem in streamfunction formulation. Journal of Computational Physics, 102(2):310–318, October 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192903748.

Heinrichs:1993:SMM

[Hei93] Wilhelm Heinrichs. Spectral multigrid methods for the reformulated Stokes equations. Journal of Computational Physics, 107(2):213–224, August 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999183711381.

Heinrichs:1998:SCT

[Hei98] Wilhelm Heinrichs. Spectral collocation on triangular elements. Journal of Computational Physics, 145(2):743–757, September 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198960521.

Helsing:1995:EEP

[Hel95] Johan Helsing. Estimating effective properties of composites from cross-sectional photographs. Journal of Computational Physics, 117(2):281–288, March 15, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710662.

Helsing:1996:TBI

[Hel96] Johan Helsing. Thin bridges in isotropic electrostatics. Journal of Computational Physics, 127(1):142–151, August 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901643.

Henshaw:1994:FOA

[Hen94] William D. Henshaw. A fourth-order accurate method for the incompressible Navier–Stokes equations on overlapping grids. Journal of Computational Physics, 113(1):13–25, July 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711144.

Harten:1997:UHO

[HEOC97] Ami Harten, Bjorn Engquist, Stanley Osher, and Sukumar R. Chakravarthy. Uniformly high order accurate essentially non-oscillatory schemes, III. Journal of Computational Physics, 131(1):3–47, February 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196956326.

Hernquist:1990:VTT

[Her90] Lars Hernquist. Vectorization of tree traversals. Journal of Computational Physics, 87(1):137–147, March 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090230X.

Hermeline:1993:TCP

[Her93] F. Hermeline. Two coupled particle-finite volume methods using Delaunay–Voronoĭ meshes for the approximation of Vlasov–Poisson and Vlasov–Maxwell equations. Journal of Computational Physics, 106(1):1–18, May 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710867.

Hesthaven:1998:ACP

[Hes98] J. S. Hesthaven. On the analysis and construction of perfectly matched layers for the linearized Euler equations. Journal of Computational Physics, 142(1):129–147, May 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959381.

Hewett:1997:ECB

[Hew97] Dennis W. Hewett. The embedded curved boundary method for orthogonal simulation meshes. Journal of Computational Physics, 138(2):585–616, December 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958356.

Hamed:1998:NSU

[HF98] M. S. Hamed and J. M. Floryan. Numerical simulation of unsteady nonisothermal capillary interfaces. Journal of Computational Physics, 145(1):110–140, September 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960235.

Helton:1991:NTO

[HG91] F. J. Helton and J. M. Greene. A numerical technique for obtaining the complete bifurcated equilibrium solution space for a Tokamak. Journal of Computational Physics, 93(2):487, April 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190204X.

Helton:1992:NTO

[HG92] F. J. Helton and J. M. Greene. A numerical technique for obtaining the complete bifurcated equilibrium solution space for a tokamak. Journal of Computational Physics, 98(1):101–107, January 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290176Y.

Halpern:1994:BEA

[HG94] D. Halpern and D. P. Gaver III. Boundary element analysis of the time-dependent motion of a semi-infinite bubble in a channel. Journal of Computational Physics, 115(2):366–375, December 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999184712022.

Horno:1995:NMS

[HGFH95] J. Horno, C. F. González-Fernández, and A. Hayas. The network method for solutions of oscillating reaction-diffusion systems. Journal of Computational Physics, 118(2):310–319, May 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185711011.

Hawken:1990:RSA

[HGH90] D. F. Hawken, J. J. Gottlieb, and J. S. Hansen. Review of some adaptive node-movement techniques in finite-element and finite-difference solutions of partial differential equations. Journal of Computational Physics, 91(1):249, November 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090016T.

Hawken:1991:RSA

[HGH91] D. F. Hawken, J. J. Gottlieb, and J. S. Hansen. Review of some adaptive node-movement techniques in finite-element and finite-difference solutions of partial differential equations. Journal of Computational Physics, 95(2):254–302, August 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190277R.

Hofmann:1998:RFR

[HH98] Lorenz M. Hofmann and Thorwald Herbert. Reproducing the flow response to actuator motion. Journal of Computational Physics, 142(1):264–268, May 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959459.

Hannappel:1995:CET

[HHF95] Ralf Hannappel, Thomas Hauser, and Rainer Friedrich. A comparison of ENO and TVD schemes for the computation of shock-turbulence interaction. Journal of Computational Physics, 121(1):176–184, October 1, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711874.

Hayase:1991:CFQ

[HHG91] T. Hayase, J. A. C. Humphrey, and R. Greif. A consistently formulated quick scheme for fast and stable convergence using finite-volume iterative calculation procedures. Journal of Computational Physics, 94(1):252, May 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190151A.

Hayase:1992:CFQ

[HHG92] T. Hayase, J. A. C. Humphrey, and R. Greif. A consistently formulated QUICK scheme for fast and stable convergence using finite-volume iterative calculation procedures. Journal of Computational Physics, 98(1):108–118, January 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290177Z.

Hu:1996:LDL

[HHM96] F. Q. Hu, M. Y. Hussaini, and J. L. Manthey. Low-dissipation and low-dispersion Runge–Kutta schemes for computational acoustics. Journal of Computational Physics, 124(1):177–191, March 1, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196900522.

Hu:1999:ADG

[HHR99] Fang Q. Hu, M. Y. Hussaini, and Patrick Rasetarinera. An analysis of the discontinuous Galerkin method for wave propagation problems. Journal of Computational Physics, 151(2):921– 946, May 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999199962277.

Horiuti:1998:TEA

[HI98] Kiyosi Horiuti and Takao Itami. Truncation error analysis of the rotational form for the convective terms in the Navier–Stokes equation. Journal of Computational Physics, 145(2):671–692, September 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198960508.

Higdon:1992:ABCa

[Hig92a] Robert L. Higdon. Absorbing boundary conditions for acoustic and elastic waves in stratified media. Journal of Computational Physics, 100(2):433–434, June 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290255W.

Higdon:1992:ABCb

[Hig92b] Robert L. Higdon. Absorbing boundary conditions for acoustic and elastic waves in stratified media. Journal of Computational Physics, 101(2):386–418, August 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290016R.

Higdon:1999:IBB

[Hig99] Robert L. Higdon. Implementation of a barotropic-baroclinic time splitting for isopycnic coordinate ocean modeling. Journal of Computational Physics, 148(2):579–604, January 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961307.

Hillion:1997:BMD

[Hil97] Pierre Hillion. Beware of Maxwell’s divergence equations. Journal of Computational Physics, 132(1):154–155, March 15, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196956296.

Hirsch:1997:ITU

[Hir97] Ch. Hirsch. Introduction to “Towards the Ultimate Conservative Difference Scheme. V. A Second-Order Sequel to Godunov’s Method”. Journal of Computational Physics, 135(2):227– 228, August 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999197957570.

Han:1995:BEM

[HIY95] H. Han, D. B. Ingham, and Y. Yuan. The boundary element method for the solution of the backward heat conduction equation. Journal of Computational Physics, 116(2):292–299, February 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185710285.

Hesthaven:1998:WOA

[HJ98] J. S. Hesthaven and L. M. Jameson. A wavelet optimized adaptive multi-domain method. Journal of Computational Physics, 145(1):280–296, September 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960120.

Herman:1992:NIP

[HK92] R. L. Herman and C. J. Knickerbocker. Numerically induced phase shift in the KdV soliton. Journal of Computational Physics, 101(2):453, August 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919290032T.

Herman:1993:NIP

[HK93a] R. L. Herman and C. J. Knickerbocker. Numerically induced phase shift in the KdV soliton. Journal of Computational Physics, 104(1):50–55, January 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710065.

Hughes:1993:EGD

[HK93b] S. R. Hughes and Uzi Kaldor. On the easy generation of the diagrams representing the coupled cluster method. Journal of Computational Physics, 109(1):108–114, November 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183712039.

Hitchon:1994:KST

[HK94] W. N. G. Hitchon and E. R. Keiter. Kinetic simulation of a time-dependent two-dimensional plasma. Journal of Computational Physics, 112(2):226–233, June 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710941.

Henderson:1995:USE

[HK95] Ronald D. Henderson and George Em Karniadakis. Unstructured spectral element methods for simulation of turbulent flows. Journal of Computational Physics, 122(2):191–217, December 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185712089.

Hile:1998:HNM

[HK98] Cheryl V. Hile and Gregory A. Kriegsmann. A hybrid numerical method for loaded highly resonant single mode cavities. Journal of Computational Physics, 142(2):506–520, May 20, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959514.

Haftel:1996:PSS

[HKM96] M. Haftel, R. Krivec, and V. B. Mandelzweig. Power series solution of coupled differential equations in one variable. Journal of Computational Physics, 123(1):149–161, January 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900121.

Heikkola:1998:FDM

[HKNT98] Erkki Heikkola, Yuri A. Kuznetsov, Pekka Neittaanmäki, and Jari Toivanen. Fictitious domain methods for the numerical solution of two-dimensional scattering problems. Journal of Computational Physics, 145(1):89–109, September 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960144.

Holden:1999:OSM

[HKR99] Helge Holden, Kenneth Hvistendahl Karlsen, and Nils Henrik Risebro. Operator splitting methods for generalized Korteweg–de Vries equations. Journal of Computational Physics, 153(1):203– 222, July 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999199962733.

Hou:1998:RSC

[HKS98] Thomas Y. Hou, Isaac Klapper, and Helen Si. Removing the stiffness of curvature in computing 3-D filaments. Journal of Computational Physics, 143(2):628–664, July 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959770.

Hundsdorfer:1995:PFD

[HKvV95] W. Hundsdorfer, B. Koren, M. vanLoon, and J. G. Verwer. A positive finite-difference advection scheme. Journal of Computational Physics, 117(1):35–46, March 1, 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918571042X.

Hodgkinson:1997:ECA

[HKW97] I. J. Hodgkinson, S. Kassam, and Q. H. Wu. Eigenequations and compact algorithms for bulk and layered anisotropic optical media: Reflection and refraction at a crystal-crystal interface. Journal of Computational Physics, 133(1):75–83, May 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956485.

Hsu:1990:NTT

[HL90] K. Hsu and S. L. Lee. A numerical technique for two-dimensional grid generation with grid control at all of the boundaries. Journal of Computational Physics, 91(2):496, December 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190900567.

Hsu:1991:NTT

[HL91] K. Hsu and S. L. Lee. A numerical technique for two-dimensional grid generation with grid control at all of the boundaries. Journal of Computational Physics, 96(2):451–469, October 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190245G.

Hui:1992:NLMa

[HL92a] W. H. Hui and C. Y. Loh. A new Lagrangian method for steady supersonic flow computation part II. Slip-line resolution. Journal of Computational Physics, 103(2):450–464, December 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290415U.

Hui:1992:NLMb

[HL92b] W. H. Hui and C. Y. Loh. A new Lagrangian method for steady supersonic flow computation part III. Strong shocks. Journal of Computational Physics, 103(2):465–471, December 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290416V.

Huang:1998:SOU

[HL98] Ying Huang and Alain Lerat. Second-order upwinding through a characteristic time-step matrix for compressible flow calculations. Journal of Computational Physics, 142(2):445–472, May 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198959356.

Han:1994:DAB

[HLB94] Houde Han, Jinfu Lu, and Weizhu Bao. A discrete artificial boundary condition for steady incompressible viscous flows in a no-slip channel using a fast iterative method. Journal of Computational Physics, 114(2):201–208, October 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711600.

Hewett:1992:SSPb

[HLD92a] D. W. Hewett, D. J. Larson, and S. Doss. Solution of simultaneous partial differential equations using dynamic ADI: Solution of the streamlined Darwin field equations. Journal of Computational Physics, 99(2):351, April 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290216L.

Hewett:1992:SSPa

[HLD92b] D. W. Hewett, D. J. Larson, and S. Doss. Solution of simultaneous partial differential equations using dynamic ADI: Solution of the streamlined Darwin field equations. Journal of Computational Physics, 101(1):11–24, July 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192900392.

He:1996:SPL

[HLD96] Xiaoyi He, Li-Shi Luo, and Micah Dembo. Some progress in lattice Boltzmann method. Part I. Nonuniform mesh grids. Journal of Computational Physics, 129(2):357–363, December 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196902557.

Hui:1999:UCS

[HLL99] W. H. Hui, P. Y. Li, and Z. W. Li. A unified coordinate system for solving the two-dimensional Euler equations. Journal of Computational Physics, 153(2):596–637, August 10, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962952.

Hou:1997:HMM

[HLOZ97] Thomas Y. Hou, Zhilin Li, Stanley Osher, and Hongkai Zhao. A hybrid method for moving interface problems with application to the Hele–Shaw flow. Journal of Computational Physics, 134(2):236–252, July 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956898.

Holden:1999:USM

[HLR99] Helge Holden, Knut-Andreas Lie, and Nils Henrik Risebro. An unconditionally stable method for the Euler equations. Journal of Computational Physics, 150(1):76–96, March 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961642.

Hou:1994:RSI

[HLS94] Thomas Y. Hou, John S. Lowengrub, and Michael J. Shelley. Removing the stiffness from interfacial flows with surface tension. Journal of Computational Physics, 114(2):312–338, October 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999184711703.

Hornung:1990:FEM

[HM90] Ulrich Hornung and Hans D. Mittelmann. A finite element method for capillary surfaces with volume constraints. Journal of Computational Physics, 87(1):126–136, March 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090229T.

Hamilton:1995:RGM

[HM95] John T. Hamilton and George Majda. On the Rokhlin– Greengard method with vortex blobs for problems posed in all space or periodic in one direction. Journal of Computational Physics, 121(1):29–50, October 1, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711771.

Huber:1999:OOO

[HM99] Gary A. Huber and J. Andrew McCammon. OOMPAA-object-oriented model for probing assemblages of atoms. Journal of Computational Physics, 151(1):264–282, May 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199961995.

Haas:1993:MTR

[HMD93] Brian L. Haas, Jeffrey D. McDonald, and Leonardo Dagum. Models of thermal relaxation mechanics for particle simulation methods. Journal of Computational Physics, 107(2):348–358, August 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183711496.

Helenbrook:1999:NMS

[HML99] B. T. Helenbrook, L. Martinelli, and C. K. Law. A numerical method for solving incompressible flow problems with a surface of discontinuity. Journal of Computational Physics, 148(2):366–396, January 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198961150.

Hegarty:1995:SMF

[HMOS95] Alan F. Hegarty, John J. H. Miller, Eugene O’Riordan, and G. I. Shishkin. Special meshes for finite difference approximations to an advection-diffusion equation with parabolic layers. Journal of Computational Physics, 117(1):47–54, March 1, 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710431.

Hattori:1995:IEG

[HN95] Tetsuya Hattori and Hideo Nakajima. Improvement of efficiency in generating random U(1) variables with Boltzmann distribution (von Mises distribution revisited). Journal of Computational Physics, 121(2):238–245, October 15, 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999195900519.

Higham:1996:NED

[HO96] Desmond J. Higham and Brynjulf Owren. Nonnormality effects in a discretised nonlinear reaction-convection-diffusion equation. Journal of Computational Physics, 124(2):309–323, March 15, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196900625.

Hobson:1992:EMC

[Hob92] Dana Hobson. An efficient method for computing invariant manifolds of planar maps. Journal of Computational Physics, 101(2): 453, August 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290028W.

Hobson:1993:EMC

[Hob93] Dana Hobson. An efficient method for computing invariant manifolds of planar maps. Journal of Computational Physics, 104(1): 14–22, January 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999183710028.

Hoffman:1995:ACF

[Hof95] Gary G. Hoffman. Atom-centered functions in the optimized Thomas–Fermi theory. Journal of Computational Physics, 116(1):154–159, January 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999185710133.

Hoffman:1997:ISO

[Hof97] Gary G. Hoffman. An integral for second-order multiple scattering perturbation theory. Journal of Computational Physics, 130(1):129–135, January 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196955746.

Holmgren:1994:AAA

[Hol94] Pär Holmgren. An advection algorithm and an atmospheric airflow application. Journal of Computational Physics, 115(1):27– 42, November 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184711764.

Holloway:1996:NMH

[Hol96a] James Paul Holloway. On numerical methods for Hamiltonian PDEs and a collocation method for the Vlasov–Maxwell equations. Journal of Computational Physics, 129(1):121–133, November 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196902387.

Holnicki:1996:PQI

[Hol96b] Piotr Holnicki. A piecewise-quintic Interpolation scheme. Journal of Computational Physics, 127(2):316–329, September 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901783.

Hood:1990:NMF

[Hoo90] Kevin Hood. A numerical method for finding the ground states of one-dimensional systems. Journal of Computational Physics, 89(1):187–206, July 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919090122H.

Horowitz:1990:NAC

[Hor90] E. J. Horowitz. Notes to appear on computing eletrostatic field lines for two-dimensional vacuum field in the neighborhood of localized regions of charge. Journal of Computational Physics, 90(1):270, September 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090220U.

Horowitz:1991:CEF

[Hor91] Eric J. Horowitz. On computing electrostatic field lines for twodimensional vacuum fields in the neighborhood of localized regions of charge. Journal of Computational Physics, 97(2):553– 558, December 1991. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919190013B.

Hegarty:1993:CUC

[HOS93] Alan F. Hegarty, Eugene O’Riordan, and Martin Stynes. A comparison of uniformly convergent difference schemes for two-dimensional convection-diffusion problems. Journal of Computational Physics, 105(1):24–32, March 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710508.

Harabetian:1996:EAV

[HOS96] Eduard Harabetian, Stanley Osher, and Chi-Wang Shu. An Eulerian approach for vortex motion using a level set regularization procedure. Journal of Computational Physics, 127(1):15–26, August 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196901552.

Harabetian:1992:NHS

[HP92] Eduard Harabetian and Robert Pego. Nonconservative hybrid shock capturing schemes. Journal of Computational Physics, 103(1):187, November 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919290334U.

Harabetian:1993:NHS

[HP93] Eduard Harabetian and Robert Pego. Nonconservative hybrid shock capturing schemes. Journal of Computational Physics, 105(1):1–13, March 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199918371048X.

Henon:1998:GR

[HP98] M. Hénon and J-M. Petit. Good rotations. Journal of Computational Physics, 146(1):420–435, October 10, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960661.

Han:1999:SDT

[HP99] David K. Han and Andrea Prosperetti. A shape decomposition technique in electrical impedance tomography. Journal of Computational Physics, 155(1):75–95, October 10, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963301.

Harari:1998:DNM

[HPG98] Isaac Harari, Igor Patlashenko, and Dan Givoli. Dirichletto-Neumann maps for unbounded wave guides. Journal of Computational Physics, 143(1):200–223, June 10, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959605.

Hariharan:1992:DNC

[HPS92a] S. I. Hariharan, Yu Ping, and J. R. Scott. The domain numerical calculations of unsteady vortical flows about a flat plate airfoil. Journal of Computational Physics, 100(2):434–435, June 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290263X.

Hariharan:1992:TDN

[HPS92b] S. I. Hariharan, Yu Ping, and J. R. Scott. Time domain numerical calculations of unsteady vortical flows about a flat plate airfoil. Journal of Computational Physics, 101(2):419–430, August 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290017S.

Hou:1996:CBO

[HR96] L. S. Hou and S. S. Ravindran. Computations of boundary optimal control problems for an electrically conducting fluid. Journal of Computational Physics, 128(2):319–330, October 15, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196902132.

Hou:1999:LSA

[HRL99] Thomas Y. Hou, Phoebus Rosakis, and Philippe LeFloch. A levelset approach to the computation of twinning and phase-transition dynamics. Journal of Computational Physics, 150(2):302–331, April 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198961794.

Huang:1994:MMM

[HRR94] Weizhang Huang, Yuhe Ren, and Robert D. Russell. Moving mesh methods based on moving mesh partial differential equations. Journal of Computational Physics, 113(2):279–290, August 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999184711351.

Homeier:1990:NIF

[HS90] Herbert H. H. Homeier and E. Otto Steinborn. Numerical integration of functions with a sharp peak at or near one boundary using möbius transformations. Journal of Computational Physics, 87(1):61–72, March 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090225P.

Hazins:1993:CSS

[HS93a] V. M. Hazins and V. V. Svetsov. A conservative stable smoothness-enhancing free-Lagrangian method. Journal of Computational Physics, 105(2):187–198, April 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710673.

Huang:1993:PCS

[HS93b] Weizhang Huang and David M. Sloan. Pole condition for singular problems: The pseudospectral approximation. Journal of Computational Physics, 107(2):254–261, August 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711411.

Huang:1994:PMS

[HS94] Weizhang Huang and David M. Sloan. The pseudospectral method for solving differential eigenvalue problems. Journal of Computational Physics, 111(2):399–409, April 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710734.

Hodges:1999:STN

[HS99a] Ben R. Hodges and Robert L. Street. On simulation of turbulent nonlinear free-surface flows. Journal of Computational Physics, 151(2):425–457, May 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961666.

Hu:1999:WEN

[HS99b] Changqing Hu and Chi-Wang Shu. Weighted essentially nonoscillatory schemes on triangular meshes. Journal of Computational Physics, 150(1):97–127, March 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961654.

Hyman:1999:MDM

[HS99c] James M. Hyman and Mikhail Shashkov. Mimetic discretizations for Maxwell’s equations. Journal of Computational Physics, 151(2):881–909, May 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962253.

Hirshman:1990:IRD

[HSN90] S. P. Hirshman, U. Schwenn, and J. Nührenberg. Improved radial differencing for three-dimensional magnetohydrodynamic equilibrium calculations. Journal of Computational Physics, 87(2):396– 407, April 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999190902594.

Hyman:1997:NSD

[HSS97] James Hyman, Mikhail Shashkov, and Stanly Steinberg. The numerical solution of diffusion problems in strongly heterogeneous non-isotropic materials. Journal of Computational Physics, 132(1):130–148, March 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196956338.

Hou:1997:ETL

[HSZ^{+}97] Shuling Hou, Xiaowen Shan, Qisu Zou, Gary D. Doolen, and Wendy E. Soll. Evaluation of two lattice Boltzmann models for multiphase flows. Journal of Computational Physics, 138(2):695– 713, December 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999197958393.

Haras:1994:FDS

[HT94] Zigo Haras and Shlomo Ta’asan. Finite difference schemes for long-time integration. Journal of Computational Physics, 114(2):265–279, October 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199918471165X.

Harari:1995:AFD

[HT95] Isaac Harari and Eli Turkel. Accurate finite difference methods for time-harmonic wave propagation. Journal of Computational Physics, 119(2):252–270, July 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711345.

Hornung:1997:AMR

[HT97] Richard D. Hornung and John A. Trangenstein. Adaptive mesh refinement and multilevel iteration for flow in porous media. Journal of Computational Physics, 136(2):522–545, September 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002199919795779X.

Harfoush:1990:NIR

[HTK90] Fady Harfoush, Allen Taflove, and Gregory A. Kriegsmann. Numerical implementation of relativistic electromagnetic boundary conditions in a laboratory-frame grid. Journal of Computational Physics, 89(1):80–94, July 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919090118K.

Hu:1995:SBI

[Hu95] Fang Q. Hu. A spectral boundary integral equation method for the 2D Helmholtz equation. Journal of Computational Physics, 120(2):340–347, September 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999185711692.

Hu:1996:ABC

[Hu96] Fang Q. Hu. On absorbing boundary conditions for linearized Euler equations by a perfectly matched layer. Journal of Computational Physics, 129(1):201–219, November 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196902442.

Hubbard:1999:MSL

[Hub99] M. E. Hubbard. Multidimensional slope limiters for MUSCLtype finite volume schemes on unstructured grids. Journal of Computational Physics, 155(1):54–74, October 10, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963295.

Huiskamp:1990:DFS

[Hui90] Geertjan Huiskamp. Difference formulas for the surface Laplacian on a triangulated surface. Journal of Computational Physics, 91(1):252–253, November 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090028Y.

Huiskamp:1991:DFS

[Hui91] Geertjan Huiskamp. Difference formulas for the surface Laplacian on a triangulated surface. Journal of Computational Physics, 95(2):477–496, August 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190286T.

Humphries:1996:FEM

[Hum96a] Stanley Humphries, Jr. Finite-element models for planar Hall probes of arbitrary shape and composition. Journal of Computational Physics, 128(2):438–444, October 15, 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196902223.

Humphries:1996:NMS

[Hum96b] Stanley Humphries, Jr. Numerical modeling of space-charge-limited charged-particle emission on a conformal triangular mesh. Journal of Computational Physics, 125(2):488–497, May 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196901102.

Hwang:1992:GLRa

[HW92a] C. J. Hwang and S. J. Wu. Global and local remeshing algorithms for compressible flows. Journal of Computational Physics, 100 (2):434, June 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290261V.

Hwang:1992:GLRb

[HW92b] C. J. Hwang and S. J. Wu. Global and local remeshing algorithms for compressible flows. Journal of Computational Physics, 102(1):98–113, September 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999105800099.

Huang:1996:DCF

[HW96] Huaxiong Huang and Brian R. Wetton. Discrete compatibility in finite difference methods for viscous incompressible fluid flow. Journal of Computational Physics, 126(2):468–478, July 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196901503.

Hou:1997:MFE

[HW97] Thomas Y. Hou and Xiao-Hui Wu. A multiscale finite element method for elliptic problems in composite materials and porous media. Journal of Computational Physics, 134(1):169–189, June 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197956825.

Hwang:1994:ADV

[Hwa94] Yao-Hsin Hwang. Arbitrary domain velocity analyses for the incompressible Navier–Stokes equations. Journal of Computational Physics, 110(1):134–149, January 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710102.

Hwang:1996:HOE

[Hwa96] Yao-Hsin Hwang. Higher-order exponential difference schemes for the computations of the steady convection-diffusion equation. Journal of Computational Physics, 129(1):134–159, November 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196902399.

Hobson:1996:NFD

[HWM96] J. M. Hobson, N. Wood, and P. J. Mason. A new finite-difference diffusion scheme. Journal of Computational Physics, 125(1):16– 25, April 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196900765.

Haidvogel:1990:SSP

[HWY90] Dale B. Haidvogel, John L. Wilkin, and Roberta Young. A semispectral primitive equation ocean circulation model using vertical sigma and orthogonal curvilinear horizontal coordinates. Journal of Computational Physics, 90(1):268, September 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090210R.

Haidvogel:1991:SSP

[HWY91] Dale B. Haidvogel, John L. Wilkin, and Roberta Young. A semispectral primitive equation ocean circulation model using vertical sigma and orthogonal curvilinear horizontal coordinates. Journal of Computational Physics, 94(1):151–185, May 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999191901417.

Yang:1996:SBB

[hYsY96] Zhong hua Yang and Rui song Ye. Symmetry-breaking and bifurcation study on the laminar flows through curved pipes with a circular cross section. Journal of Computational Physics, 127(1): 73–87, August 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199919690159X.

Hou:1995:SCF

[HZC^{+}95] Shuling Hou, Qisu Zou, Shiyi Chen, Gary Doolen, and Allen C. Cogley. Simulation of cavity flow by the lattice Boltzmann method. Journal of Computational Physics, 118(2):329–347, May 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185711035.

Issa:1991:SID

[IABBG91] R. I. Issa, B. Ahmadi-Befrui, K. R. Beshay, and A. D. Gosman. Solution of the implicitly discretised reacting flow equations by operator-splitting. Journal of Computational Physics, 93(2):388– 410, April 1991. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919190191M.

Ibraheem:1996:BGL

[ID96] S. O. Ibraheem and A. O. Demuren. On bi-grid local mode analysis of solution techniques for 3-D Euler and Navier–Stokes equations. Journal of Computational Physics, 125(2):354–377, May 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196900996.

Issa:1990:SID

[IGABB90] R. I. Issa, A. D. Gosman, B. Ahmadi-Befrui, and K. R. Beshay. Solution of the implicitly discretized reacting flow equations by operator-splitting. Journal of Computational Physics, 90(1):267, September 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919090207H.

Igarashi:1995:SAC

[IH95] Hajime Igarashi and Toshihisa Honma. Strategies for the accurate computation of potential derivatives in boundary element method: Application to two-dimensional problems. Journal of Computational Physics, 119(2):244–251, July 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711333.

Iliev:1995:BMI

[IM95] Oleg P. Iliev and Mikhail M. Makarov. A block-matrix iterative numerical method for coupled solving 2D Navier–Stokes equations. Journal of Computational Physics, 121(2):324–330, October 15, 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999195901701.

Ito:1998:ROM

[IR98] K. Ito and S. S. Ravindran. A reduced-order method for simulation and control of fluid flows. Journal of Computational Physics, 143(2):403–425, July 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959435.

Iollo:1996:COS

[IS96] Angelo Iollo and Manuel D. Salas. Contribution to the optimal shape design of two-dimensional internal flows with embedded shocks. Journal of Computational Physics, 125(1):124–134, April 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196900832.

Ierley:1992:SMTa

[ISW92a] Glenn Ierley, Brian Spencer, and Rodney Worthing. Spectral methods in time for a class of parabolic partial differential equations. Journal of Computational Physics, 100(2):434, June 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290262W.

Ierley:1992:SMTb

[ISW92b] Glenn Ierley, Brian Spencer, and Rodney Worthing. Spectral methods in time for a class of parabolic partial differential equations. Journal of Computational Physics, 102(1):88–97, September 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999105800087.

Ingham:1990:NIS

[IT90] D. B. Ingham and T. Tang. A numerical investigation into the steady flow past a rotating circular cylinder at low and intermediate Reynolds numbers. Journal of Computational Physics, 87(1): 91–107, March 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919090227R.

Ivansson:1993:DMF

[Iva93] Sven Ivansson. Delta-matrix factorization for fast propagation through solid layers in a fluid-solid medium. Journal of Computational Physics, 108(2):357–367, October 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711897.

Jablonski:1994:NES

[Jab94] Aleksander Jablonski. Numerical evaluation of spherical Bessel functions of the first kind. Journal of Computational Physics, 111(2):256–259, April 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999184710606.

Jacobs:1990:PMT

[Jac90] S. J. Jacobs. A pseudospectral method for two-point boundary value problems. Journal of Computational Physics, 88(1):169– 182, May 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919090246W.

Jacobs:1995:ASS

[Jac95] S. J. Jacobs. An accurate split step scheme for viscous incompressible fluid flow. Journal of Computational Physics, 119(1): 26–33, June 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999185711138.

Jacqmin:1999:CTP

[Jac99] David Jacqmin. Calculation of two-phase Navier–Stokes flows using phase-field modeling. Journal of Computational Physics, 155(1):96–127, October 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963325.

Jain:1997:CMM

[Jai97]  Abhinandan Jain.  Compensating mass matrix potential

for constrained molecular dynamics. Journal of Computational Physics, 136(2):289–297, September 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957314.

Jakubovics:1993:RAS

[Jak93] J. P. Jakubovics. Removal of apparent singularity in grid computations. Journal of Computational Physics, 104(1):274–276, January 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183710284.

Jamieson:1999:EOF

[Jam99] M. J. Jamieson. On an eighth order formula for solving a Schrödinger equation. Journal of Computational Physics, 149 (1):194–197, February 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999198961216.

Juffer:1991:EPMa

[JB91] Andre H. Juffer and Herman J. C. Berendsen. The electric potential of a macromolecule in a solvent: a fundamental approach. Journal of Computational Physics, 93(1):251, March 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190084X.

Jin:1993:NOB

[JB93] G. Jin and M. Braza. A nonreflecting outlet boundary condition for incompressible unsteady Navier–Stokes calculations. Journal of Computational Physics, 107(2):239–253, August 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918371140X.

Jones:1995:GPM

[JB95] A. E. Jones and R. J. P. Bain. A generalisation of planar magnetic gradiometer design via orthogonal polynomials. Journal of Computational Physics, 118(2):191–199, May 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710911.

Jimenez:1996:PRA

[JBA96] J. C. Jimenez, R. Biscay, and E. Aubet. Parametric representation of anatomical structures of the human body by means of trigonometric interpolating sums. Journal of Computational Physics, 126(2):243–250, July 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901357.

Juffer:1991:EPMb

[JBvK^{+}91] AndréH Juffer, Eugen F. F. Botta, Bert A. M. van Keulen, Auke van der Ploeg, and Herman J. C. Berendsen. The electric potential of a macromolecule in a solvent: a fundamental approach. Journal of Computational Physics, 97(1):144–171, November 1991. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919190043K.

Jeng:1992:TEA

[JC92] Yih Nen Jeng and Jiann Lin Chen. Truncation error analysis of the finite volume method for a model steady convective equation. Journal of Computational Physics, 100(1):64–76, May 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290310U.

Johansen:1998:CGE

[JC98] Hans Johansen and Phillip Colella. A Cartesian grid embedded boundary method for Poisson’s equation on irregular domains. Journal of Computational Physics, 147(1):60–85, November 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198959654.

Jakob-Chien:1997:FSF

[JCA97] Rüdiger Jakob-Chien and Bradley K. Alpert. A fast spherical filter with uniform resolution. Journal of Computational Physics, 136(2):580–584, September 15, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919795782X.

Jakob-Chien:1995:STS

[JCHW95] Ruediger Jakob-Chien, James J. Hack, and David L. Williamson. Spectral transform solutions to the shallow water test set. Journal of Computational Physics, 119(1):164–187, June 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711254.

Jiang:1999:HOW

[JcW99] Guang-Shan Jiang and Cheng chin Wu. A high-order WENO finite difference scheme for the equations of ideal magnetohydrodynamics. Journal of Computational Physics, 150(2):561–594, April 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199962071.

Jessop:1994:MMB

[JDC94] Chris Jessop, Martin Duncan, and W. Y. Chau. Multigrid methods for N-body gravitational systems. Journal of Computational Physics, 115(2):339–351, December 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184712009.

Joly:1990:PBG

[JE90] P. Joly and R. Eymard. Preconditioned biconjugate gradient methods for numerical reservoir simulation. Journal of Computational Physics, 91(2):298–309, December 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190900394.

Jessee:1998:AMR

[JFH^{+}98] J. Patrick Jessee, Woodrow A. Fiveland, Louis H. Howell, Phillip Colella, and Richard B. Pember. An adaptive mesh refinement algorithm for the radiative transport equation. Journal of Computational Physics, 139(2):380–398, January 20, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958708.

Jacobs:1997:ATD

[JH97] Donald J. Jacobs and Bruce Hendrickson. An algorithm for twodimensional rigidity percolation: The pebble game. Journal of Computational Physics, 137(2):346–365, November 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958095.

Ji:1994:SAS

[Ji94] Xingzhi Ji. On a shooting algorithm for Sturm–Liouville eigenvalue problems with periodic and semi-periodic boundary conditions. Journal of Computational Physics, 111(1):74–80, March 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002199918471045X.

Ji:1996:BMD

[Ji96] Xingzhi Ji. On 2D bisection method for double eigenvalue problems. Journal of Computational Physics, 126(1):91–98, June 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196901229.

Jin:1994:ECS

[Jin94] Bao Xia Jin. On an essentially conservative scheme for hyperbolic conservation laws. Journal of Computational Physics, 112(2): 308–315, June 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199918471103X.

Jin:1995:RKM

[Jin95] Shi Jin. Runge–Kutta methods for hyperbolic conservation laws with stiff relaxation terms. Journal of Computational Physics, 122(1):51–67, November 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711965.

Johnsson:1992:CEM

[JJK92] S. Lennart Johnsson, Michel Jacquemin, and Robert L. Krawitz. Communication efficient multi-processor FFT. Journal of Computational Physics, 102(2):381–397, October 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290380H.

Jin:1996:NSH

[JL96a] Shi Jin and C. David Levermore. Numerical schemes for hyperbolic conservation laws with stiff relaxation terms. Journal of Computational Physics, 126(2):449–467, July 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901497.

Jin:1996:ENV

[JL96b] Shi Jin and Jian-Guo Liu. The effects of numerical viscosities: I. Slowly moving shocks. Journal of Computational Physics, 126(2): 373–389, July 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196901448.

Jarratt:1990:GSS

[JLB90] Mary Jarratt, John Lund, and Kenneth L. Bowers. Galerkin schemes and the sinc-Galerkin method for singular Sturm– Liouville problems. Journal of Computational Physics, 89(1): 41–62, July 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919090116I.

Jou:1997:MEI

[JLL97] H.-J. Jou, P. H. Leo, and J. S. Lowengrub. Microstructural evolution in inhomogeneous elastic media. Journal of Computational Physics, 131(1):109–148, February 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196955813.

Jones:1996:GBC

[JLM^{+}96] Michael E. Jones, Don S. Lemons, Rodney J. Mason, Vincent A. Thomas, and Dan Winske. A grid-based Coulomb collision model for PIC codes. Journal of Computational Physics, 123(1):169– 181, January 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196900145.

Joyce:1997:EPC

[JLSM97] Glenn Joyce, Martin Lampe, Steven P. Slinker, and Wallace M. Manheimer. Electrostatic particle-in-cell simulation technique for quasineutral plasma. Journal of Computational Physics, 138(2):540–562, December 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958332.

Jenny:1998:RHR

[JM98] P. Jenny and B. Müller. Rankine–Hugoniot–Riemann solver considering source terms and multidimensional effects. Journal of Computational Physics, 145(2):575–610, September 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198960375.

Jayaram:1999:FEA

[JMDB99] B. Jayaram, K. J. McConnell, Surjit B. Dixit, and D. L. Beveridge. Free energy analysis of protein–DNA binding: The EcoRI endonuclease–DNA complex. Journal of Computational Physics, 151(1):333–357, May 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961733.

Jones:1994:FDA

[JMR94] P. Jones, J. Ma, and V. Rokhlin. A fast direct algorithm for the solution of the Laplace equation on regions with fractal boundaries. Journal of Computational Physics, 113(1):35–51, July 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999184711168.

Jenny:1997:CCE

[JMT97] P. Jenny, B. Müller, and H. Thomann. Correction of conservative Euler solvers for gas mixtures. Journal of Computational Physics, 132(1):91–107, March 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196956259.

Johansson:1993:BCO

[Joh93] B. Christer V. Johansson. Boundary conditions for open boundaries for the incompressible Navier–Stokes equation. Journal of Computational Physics, 105(2):233–251, April 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710715.

Jordan:1999:LES

[Jor99] Stephen. A. Jordan. A large-eddy simulation methodology in generalized curvilinear coordinates. Journal of Computational Physics, 148(2):322–340, January 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961125.

Jeng:1995:ATL

[JP95] Yih Nen Jeng and Uon Jan Payne. An adaptive TVD limiter. Journal of Computational Physics, 118(2):229–241, May 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185710959.

Jordan:1996:EFS

[JR96] Stephen A. Jordan and Saad A. Ragab. An efficient fractionalstep technique for unsteady incompressible flows using a semistaggered grid strategy. Journal of Computational Physics, 127(1):218–225, August 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999196901709.

Junk:1999:NDV

[JR99] Michael Junk and S. V. Raghurama Rao. A new discrete velocity method for Navier–Stokes equations. Journal of Computational Physics, 155(1):178–198, October 10, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963362.

Jordan:1993:FAG

[JS93] Stephen A. Jordan and Malcolm L. Spaulding. A fast algorithm for grid generation. Journal of Computational Physics, 104(1):118–128, January 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999183710144.

Jiang:1996:EIW

[JS96] Guang-Shan Jiang and Chi-Wang Shu. Efficient implementation of weighted ENO schemes. Journal of Computational Physics, 126(1):202–228, June 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901308.

Justiz:1995:MNF

[JSD95] Charles R. Justiz, Ronald M. Sega, and Charles Dalton. A method for near field computation of coupled weakly ionized plasma flows in low Earth orbit. Journal of Computational Physics, 118(2):278–293, May 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710996.

Jones:1997:ISN

[JSE97] Ogden S. Jones, Uri Shumlak, and D. Scott Eberhardt. An implicit scheme for nonideal magnetohydrodynamics. Journal of Computational Physics, 130(2):231–242, January 15, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196955801.

Jorgenson:1993:CDT

[JT93] P. Jorgenson and E. Turkel. Central difference TVD schemes for time dependent and steady state problems. Journal of Computational Physics, 107(2):297–308, August 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711459.

Juric:1996:FTM

[JT96] Damir Juric and Grétar Tryggvason. A front-tracking method for dendritic solidification. Journal of Computational Physics, 123(1):127–148, January 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199919690011X.

Jiang:1999:IVN

[JT99] Z. Jiang and K. Takayama. An investigation into the validation of numerical solutions of complex flowfields. Journal of Computational Physics, 151(2):479–497, May 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199961867.

Junk:1999:KSC

[Jun99] Michael Junk. Kinetic schemes in the case of low Mach numbers. Journal of Computational Physics, 151(2):947–968, May 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199962289.

Jyotsna:1995:MCS

[JV95] R. Jyotsna and S. P. Vanka. Multigrid calculation of steady, viscous flow in a triangular cavity. Journal of Computational Physics, 122(1):107–117, November 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185712004.

Jain:1993:FRA

[JVR93] A. Jain, N. Vaidehi, and G. Rodriguez. A fast recursive algorithm for molecular dynamics simulation. Journal of Computational Physics, 106(2):258–268, June 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918371106X.

Jian:1997:CWC

[JVS97] Hongmei Jian, Alexander V. Vologodskii, and Tamar Schlick. A combined wormlike-chain and bead model for dynamic simulations of long linear DNA. Journal of Computational Physics, 136(1):168–179, September 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919795765X.

Jiang:1990:ABC

[JW90] Hong Jiang and Yau Shu Wong. Absorbing boundary conditions for second-order hyperbolic equations. Journal of Computational Physics, 88(1):205–231, May 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090248Y.

Jafroudi:1993:MAS

[JYH93] Hamid Jafroudi, Hsun-Tiao Yang, and Jerry Hermel. Matched asymptotic solutions of impulsive flow over an elliptic cylinder. Journal of Computational Physics, 109(2):289–305, December 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183712180.

Jinrong:1995:CSS

[JZQ^{+}95] Cheng Jinrong, Zhang Ziping, Zhou Qin, Yi Youmin, Gao Yongchun, and Ynag Detian. Computer simulation study of the properties of orthopositronium 3 γ decay. Journal of Computational Physics, 118(2):396–400, May 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711102.

Kaber:1996:LPV

[Kab96] S. M. Ould Kaber. A Legendre pseudospectral viscosity method. Journal of Computational Physics, 128(1):165–180, October 1, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196902016.

Kalinderis:1991:NTG

[Kal91] Yannis Kalinderis. Numerical treatment of grid interfaces for viscous flows. Journal of Computational Physics, 93(1):252, March 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190086Z.

Kallinderis:1992:NTG

[Kal92] Yannis Kallinderis. Numerical treatment of grid interfaces for viscous flows. Journal of Computational Physics, 98(1):129– 144, January 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999192901793.

Kaltenbach:1997:CAR

[Kal97] Hans-Jakob Kaltenbach. Cell aspect ratio dependence of anisotropy measures for resolved and subgrid scale stresses. Journal of Computational Physics, 136(2):399–410, September 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197957557.

Kang:1992:OGGa

[Kan92]  I. S. Kang.  Orthogonal grid generation in a 2D do-

main via the boundary integral technique. Journal of Computational Physics, 100(2):434, June 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192902606.

Karageorghis:1991:MFS

[Kar91] Andreas Karageorghis. The method of fundamental solutions for the solution of steady-state free boundary problems. Journal of Computational Physics, 93(2):486, April 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190201U.

Karageorghis:1992:MFS

[Kar92] Andreas Karageorghis. The method of fundamental solutions for the solution of steady-state free boundary problems. Journal of Computational Physics, 98(1):119–128, January 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192901782.

Karni:1994:MFC

[Kar94] Smadar Karni. Multicomponent flow calculations by a consistent primitive algorithm. Journal of Computational Physics, 112(1): 31–43, May 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184710801.

Karlin:1997:NAF

[Kar97] Vladimir Karlin. Numerical algorithms for flows in the nodes of 2D models of pipe networks. Journal of Computational Physics, 132(1):62–77, March 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196956090.

Kinoshita:1996:ABS

[KB96] M. Kinoshita and D. R. Bérard. Analysis of the bulk and surfaceinduced structure of electrolyte solutions using integral equation theories. Journal of Computational Physics, 124(1):230– 241, March 1, 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196900558.

Kuang:1999:NMM

[KB99] Weijia Kuang and Jeremy Bloxham. Numerical modeling of magnetohydrodynamic convection in a rapidly rotating spherical shell: Weak and strong field dynamo action. Journal of Computational Physics, 153(1):51–81, July 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962745.

Kirby:1990:EAA

[KBS90] M. Kirby, J. Boris, and L. Sirovich. An eigenfunction analysis of axisymmetric jet flow. Journal of Computational Physics, 90(1):98–122, September 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090198A.

Kovanen:1992:HCS

[KC92] M. A. Kovanen and W. G. F. Core. HECTOR: a code for the study of charged particles in axisymmetric tokamak plasmas. Journal of Computational Physics, 103(1):187, November 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290335V.

Kovanen:1993:HCS

[KC93] M. A. Kovanen and W. G. F. Core. HECTOR: a code for the study of charged particles in axisymmetric Tokamak plasmas. Journal of Computational Physics, 105(1):14–23, March 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183710491.

Karni:1997:CSM

[KC97] Smadar Karni and Sunćica Canić. Computations of slowly moving shocks. Journal of Computational Physics, 136(1):132–139, September 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199919795751X.

Knio:1995:NSS

[KCJ95] Omar M. Knio, Luc Collorec, and Daniel Juvé. Numerical study of sound emission by 2D regular and chaotic vortex configurations. Journal of Computational Physics, 116(2):226–246, February 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185710236.

Kunz:1999:DIM

[KCV99] R. F. Kunz, W. K. Cope, and S. Venkateswaran. Development of an implicit method for multi-fluid flow simulations. Journal of Computational Physics, 152(1):78–101, June 10, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962356.

Kawahara:1997:BAB

[KD97] Mutsuto Kawahara and Yan Ding. Bifurcation analysis of Brown tide by reaction-diffusion equation using finite element method. Journal of Computational Physics, 131(2):253–266, March 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196956144.

Kazeminezhad:1992:VPI

[KDL^{+}92] F. Kazeminezhad, J. M. Dawson, J. N. Leboeuf, R. Sydora, and D. Holland. A Vlasov particle ion zero mass electron model for plasma simulations. Journal of Computational Physics, 102(2):277–296, October 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192903726.

Kuchta:1993:FHM

[KE93] B. Kuchta and R. D. Etters. Features of the histogram Monte Carlo method: Application to N_{2} monolayer melting on graphite. Journal of Computational Physics, 108(2):353–356, October 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183711885.

Kobeissi:1990:SSM

[KEHK90] Hafez Kobeissi, Ali El-Hajj, and Munif Kobersi. A simplified shooting method for the diatomic eigenvalue problem. Journal of Computational Physics, 91(1):254, November 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090036Z.

Kobeissi:1991:SSM

[KEHK91] Hafez Kobeissi, Ali El-Hajj, and Munif Kobersi. A simplified shooting method for the diatomic eigenvalue problem. Journal of Computational Physics, 96(2):470–474, October 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190246H.

Keller:1994:RSF

[Kel94] Joseph B. Keller. Removing small features from computational domains. Journal of Computational Physics, 113(1):148–150, July 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999184711247.

Kobeissi:1991:CFM

[KF91]  H. Kobeissi and K. Fakhreddine.  The canonical func-

tions method and singular potentials. Journal of Computational Physics, 95(2):505–510, August 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190288V.

Klimas:1994:SAV

[KF94] A. J. Klimas and W. M. Farrell. A splitting algorithm for Vlasov simulation with filamentation filtration. Journal of Computational Physics, 110(1):150–163, January 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710114.

Knio:1990:NST

[KG90a] Omar M. Knio and Ahmed F. Ghoniem. Numerical study of a three-dimensional vortex method. Journal of Computational Physics, 86(1):75–106, January 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090092F.

Knio:1990:TDV

[KG90b] Omar M. Knio and Ahmed F. Ghoniem. Three-dimensional vortex simulation of rollup and entrainment in a shear layer. Journal of Computational Physics, 91(2):495–496, December 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190900545.

Knio:1991:TDV

[KG91] Omar M. Knio and Ahmed F. Ghoniem. Three-dimensional vortex simulation of rollup and entrainment in a shear layer. Journal of Computational Physics, 97(1):172–223, November 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190044L.

Krishnan:1992:SRMa

[KG92a] Anantha Krishnan and Ahmed F. Ghoniem. Simulation of rollup and mixing in Rayleigh–Taylor flow using the transport-element method. Journal of Computational Physics, 98(2):347, February 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290152O.

Krishnan:1992:SRMb

[KG92b] Anantha Krishnan and Ahmed F. Ghoniem. Simulation of rollup and mixing in Rayleigh–Taylor flow using the transport-element method. Journal of Computational Physics, 99(1):1–27, March 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999192902709.

Kerner:1998:CNM

[KGH^{+}98] W. Kerner, J. P. Goedbloed, G. T. A. Huysmans, S. Poedts, and E. Schwarz. CASTOR: Normal-mode analysis of resistive MHD plasmas. Journal of Computational Physics, 142(2):271– 303, May 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198959101.

Kaushik:1990:APB

[KH90] Sumanth Kaushik and Peter L. Hagelstein. The application of the preconditioned biconjugate gradient algorithm to NLTE rate matrix equations. Journal of Computational Physics, 91(2):496, December 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999190900578.

Kaushik:1992:APB

[KH92] Sumanth Kaushik and Peter L. Hagelstein. The application of the preconditioned biconjugate gradient algorithm to NLTE rate matrix equations. Journal of Computational Physics, 101(2):360– 367, August 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290013O.

Koren:1997:GTS

[KH97] B. Koren and P. W. Hemker. Godunov-type schemes, sparse grids and 3D CFD. Journal of Computational Physics, 137(??), ???? 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). To appear.

Kontomaris:1992:ATFa

[KHM92a] K. Kontomaris, T. J. Hanratty, and J. B. McLaughlin. An algorithm for tracking fluid particles in a spectral simulation of turbulent channel flow. Journal of Computational Physics, 101 (1):228, July 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999192900657.

Kontomaris:1992:ATFb

[KHM92b] K. Kontomaris, T. J. Hanratty, and J. B. McLaughlin. An algorithm for tracking fluid particles in a spectral simulation of turbulent channel flow. Journal of Computational Physics, 103(2):231– 242, December 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290398I.

Khokhlov:1998:FTT

[Kho98] A. M. Khokhlov. Fully threaded tree algorithms for adaptive refinement fluid dynamics simulations. Journal of Computational Physics, 143(2):519–543, July 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198999983.

Kleiser:1998:LET

[KHW98] Leonhard Kleiser, Carlos Härtel, and Torsten Wintergerste. Letter to the Editor: There is no error in the Kleiser– Schumann influence matrix method. Journal of Computational Physics, 141(1):85–87, March 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0021999198959241. See comment [Wer98].

Kim:1999:CSM

[KI99] Arnold D. Kim and Akira Ishimaru. A Chebyshev spectral method for radiative transfer equations applied to electromagnetic wave propagation and scattering in a discrete random medium. Journal of Computational Physics, 152(1):264–280, June 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999199962472.

Karniadakis:1991:HOS

[KIO91] George Em Karniadakis, Moshe Israeli, and Steven A. Orszag. High-order splitting methods for the incompressible Navier– Stokes equations. Journal of Computational Physics, 97(2):414– 443, December 1991. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999191900078.

Kim:1998:RAL

[KJ98] Chongam Kim and Antony Jameson. A robust and accurate LED–BGK solver on unstructured adaptive meshes. Journal of Computational Physics, 143(2):598–627, July 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959733.

Kaper:1995:VCT

[KK95] Hans G. Kaper and Man Kam Kwong. Vortex configurations in type-II superconducting films. Journal of Computational Physics, 119(1):120–131, June 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711205.

Kopriva:1996:CSGa

[KK96] David A. Kopriva and John H. Kolias. A conservative staggered-grid Chebyshev multidomain method for compressible flows. Journal of Computational Physics, 125(1):244–261, April 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196900911.

Kobeissi:1990:UVS

[KKD90] Hafez Kobeissi, Majida Kobeissi, and Mounzer Dagher. On the use of a variable-step method for the computation of diatomic eigenvalues near dissociation: The Lennard-Jones potential. Journal of Computational Physics, 86(2):487–491, February 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919090110M.

Kandhai:1999:IAL

[KKH^{+}99] D. Kandhai, A. Koponen, A. Hoekstra, M. Kataja, J. Timonen, and P. M. A. Sloot. Implementation aspects of 3D lattice-BGK: Boundaries, accuracy, and a new fast relaxation method. Journal of Computational Physics, 150(2):482–501, April 10, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199961910.

Kartalev:1995:FEN

[KKK95] M. D. Kartalev, M. S. Kaschiev, and D. K. Koitchev. Finite element numerical modeling of stationary two-dimensional magnetosphere with defined boundary. Journal of Computational Physics, 119(2):220–230, July 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918571131X.

Kono:1997:EQM

[KKOF97] Hirohiko Kono, Akihisa Kita, Yukiyoshi Ohtsuki, and Yuichi Fujimura. An efficient quantum mechanical method for the electronic dynamics of the three-dimensional hydrogen atom interacting with a linearly polarized strong laser pulse. Journal of Computational Physics, 130(1):148–159, January 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196955886.

Kang:1992:OGGb

[KL92a] I. S. Kang and L. G. Leal. Orthogonal grid generation in a 2D domain via the boundary integral technique. Journal of Computational Physics, 102(1):78–87, September 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999105800075.

Kunz:1992:SEN

[KL92b] Robert F. Kunz and Budugur Lakshminarayana. Stability of explicit Navier–Stokes procedures using k–ε and k–ε / algebraic Reynolds stress turbulence models. Journal of Computational Physics, 103(1):141–159, November 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192903302.

Khayrutdinov:1993:SPE

[KL93] R. R. Khayrutdinov and V. E. Lukash. Studies of plasma equilibrium and transport in a Tokamak fusion device with the inverse-variable technique. Journal of Computational Physics, 109(2):193–201, December 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999183712118.

Kriegsmann:1994:RPR

[KL94] Gregory A. Kriegsmann and Jonathan H. C. Luke. Rapid pulse responses for scattering problems. Journal of Computational Physics, 111(2):390–398, April 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710722.

Kahan:1997:USC

[KL97] William Kahan and Ren-Chang Li. Unconventional schemes for a class of ordinary differential equations—with applications to the Korteweg–de Vries equation. Journal of Computational Physics, 134(2):316–331, July 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957107.

Klapper:1996:BAD

[Kla96] I. Klapper. Biological applications of the dynamics of twisted elastic rods. Journal of Computational Physics, 125(2):325–337, May 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196900972.

Klar:1999:RSL

[Kla99] Axel Klar. Relaxation scheme for a lattice-Boltzmann-type discrete velocity model and numerical Navier–Stokes limit. Journal of Computational Physics, 148(2):416–432, January 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002199919896123X.

Knoll:1999:MIF

[KLB99] D. A. Knoll, G. Lapenta, and J. U. Brackbill. A multilevel iterative field solver for implicit, kinetic, plasma simulation. Journal of Computational Physics, 149(2):377–388, March 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961599.

Kazeminezhad:1993:DMM

[KLBD93] F. Kazeminezhad, J. N. Leboeuf, F. Brunel, and J. M. Dawson. A discrete model for MHD incorporating the Hall term. Journal of Computational Physics, 104(2):398–417, February 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710399.

Klein:1995:SIE

[Kle95] R. Klein. Semi-implicit extension of a Godunov-type scheme based on low Mach number asymptotics I: One-dimensional flow. Journal of Computational Physics, 121(2):213–237, October 15, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999195900349.

Kroger:1995:MLS

[KLM95] H. Kröger, S. Lantagne, and K. J. M. Moriarty. Momentum lattice simulation on a small lattice using stochastic quantization. Journal of Computational Physics, 122(2):335–342, December 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185712193.

Koumoutsakos:1994:BCV

[KLP94] P. Koumoutsakos, A. Leonard, and F. Pépin. Boundary conditions for viscous vortex methods. Journal of Computational Physics, 113(1):52–61, July 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918471117X.

Mao:1990:TDS

[kM90] De kang Mao. A treatment of discontinuities in shock-capturing finite difference methods. Journal of Computational Physics, 89 (1):252, July 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919090130S.

Mao:1991:TDS

[kM91] De kang Mao. A treatment of discontinuities in shockcapturing finite difference methods. Journal of Computational Physics, 92(2):422–455, February 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999191902179.

Kadioglu:1992:IGM

[KM92] Mikdat Kadioğlu and Stephen Mudrick. On the implementation of the GMRES(m ) method to elliptic equations in meteorology. Journal of Computational Physics, 102(2):348–359, October 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290377B.

Mao:1993:TDF

[kM93a] De kang Mao. A treatment of discontinuities for finite difference methods in the two-dimensional case. Journal of Computational Physics, 104(2):377–397, February 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710387.

Khorrami:1993:ECS

[KM93b] Mehdi R. Khorrami and Mujeeb R. Malik. Efficient computation of spatial eigenvalues for hydrodynamic stability analysis. Journal of Computational Physics, 104(1):267–272, January 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710260.

Knoll:1995:INA

[KM95] D. A. Knoll and P. R. McHugh. An inexact Newton algorithm for solving the Tokamak edge plasma fluid equations on a multiply-connected domain. Journal of Computational Physics, 116(2):281–291, February 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710273.

Khamayseh:1996:CCM

[KM96] Ahmed Khamayseh and C. Wayne Mastin. Computational conformal mapping for surface grid generation. Journal of Computational Physics, 123(2):394–401, February 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900327.

Kotelnikov:1997:KMC

[KM97a] Alexei D. Kotelnikov and David C. Montgomery. A kinetic method for computing inhomogeneous fluid behavior. Journal of Computational Physics, 134(2):364–388, July 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919795720X.

Kravchenko:1997:ENE

[KM97b] A. G. Kravchenko and P. Moin. On the effect of numerical errors in large eddy simulations of turbulent flows. Journal of Computational Physics, 131(2):310–322, March 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196955977.

Kamimura:1992:IPSb

[KMB^{+}92a] T. Kamimura, E. Montalvo, D. C. Barnes, J. N. Leboeuf, and T. Tajima. Implicit particle simulation of electromagnetic plasma phenomena. Journal of Computational Physics, 99(1):181, March 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290298D.

Kamimura:1992:IPSa

[KMB^{+}92b] T. Kamimura, E. Montalvo, D. C. Barnes, J. N. Leboeuf, and T. Tajima. Implicit particle simulation of electromagnetic plasma phenomena. Journal of Computational Physics, 100(1):77–90, May 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290311L.

Kravchenko:1996:ZEG

[KMM96] A. G. Kravchenko, P. Moin, and R. Moser. Zonal embedded grids for numerical simulations of wall-bounded turbulent flows. Journal of Computational Physics, 127(2):412–423, September 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196901849.

Kravchenko:1999:BSM

[KMS99] Arthur G. Kravchenko, Parviz Moin, and Karim Shariff. B-spline method and zonal grids for simulations of complex turbulent flows. Journal of Computational Physics, 151(2):757–789, May 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199962174.

Knoll:1998:ICS

[Kno98] D. A. Knoll. An improved convection scheme applied to recombining divertor plasma flows. Journal of Computational Physics, 142(2):473–488, May 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959472.

Korniss:1999:PDM

[KNR99] G. Korniss, M. A. Novotny, and P. A. Rikvold. Parallelization of a dynamic Monte Carlo algorithm: a partially rejection-free conservative approach. Journal of Computational Physics, 153(2):488–508, August 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962915.

Knupp:1992:REG

[Knu92] Patrick M. Knupp. A robust elliptic grid generator. Journal of Computational Physics, 100(2):409–418, June 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290247V.

Knupp:1995:MGU

[Knu95] P. Knupp. Mesh generation using vector fields. Journal of Computational Physics, 119(1):142–148, June 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711229.

Knio:1999:SIN

[KNW99] Omar M. Knio, Habib N. Najm, and Peter S. Wyckoff. A semi-implicit numerical scheme for reacting flow: II. Stiff, operator-split formulation. Journal of Computational Physics, 154(2):428–467, September 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963222.

Krauskopf:1998:GQU

[KO98] Bernd Krauskopf and Hinke Osinga. Growing 1D and quasi-2D unstable manifolds of maps. Journal of Computational Physics, 146(1):404–419, October 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960594.

Kobayashi:1990:SLS

[Kob90] Mei Kobayashi. Sturm–Liouville systems with potentials: 0.5 cos Inx. Journal of Computational Physics, 90(1):270, September 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919090221L.

Kobayashi:1991:SLS

[Kob91] Mei Kobayashi. Sturm–Liouville systems with potentials 0.5 cos 2nx. Journal of Computational Physics, 94(2):487–493, June 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190232A.

Kobayashi:1999:CPF

[Kob99] Marcelo H. Kobayashi. On a class of Padé finite volume methods. Journal of Computational Physics, 156(1):137–180, November 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199963763.

Kopriva:1990:MSS

[Kop90] David A. Kopriva. Multidomain spectral solution of the Euler gas-dynamics equations. Journal of Computational Physics, 91(2):497, December 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919090059A.

Kopriva:1991:MSS

[Kop91] David A. Kopriva. Multidomain spectral solution of the Euler gas-dynamics equations. Journal of Computational Physics, 96(2):428–450, October 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919190244F.

Kopriva:1994:MSS

[Kop94] David A. Kopriva. Multidomain spectral solution of compressible viscous flows. Journal of Computational Physics, 115(1):184–199, November 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184711867.

Kopriva:1996:CSGb

[Kop96] David A. Kopriva. A conservative staggered-grid Chebyshev multidomain method for compressible flows. II. A semi-structured method. Journal of Computational Physics, 128(2):475–488, October 15, 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196902259.

Kopriva:1998:SGM

[Kop98] David A. Kopriva. A staggered-grid multidomain spectral method for the compressible Navier–Stokes equations. Journal of Computational Physics, 143(1):125–158, June 10, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959563.

Koren:1990:MDC

[Kor90] Barry Koren. Multigrid and defect correction for the steady Navier–Stokes equations. Journal of Computational Physics, 87(1):25–46, March 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919090223N.

Kosower:1992:OEP

[Kos92] David A. Kosower. Octopus: An efficient phase space mapping for light particles. Journal of Computational Physics, 102(1):18– 38, September 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105800026.

Kegley:1996:BSC

[KOS^{+}96] D. R. Kegley, Jr., V. E. Oberacker, M. R. Strayer, A. S. Umar, and J. C. Wells. Basis spline collocation method for solving the Schrödinger equation in axillary symmetric systems. Journal of Computational Physics, 128(1):197–208, October 1, 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919690203X.

Koures:1996:SCS

[Kou96] Vasilios G. Koures. Solving the Coulomb Schrödinger equation in d = 2 + 1 via sinc collocation. Journal of Computational Physics, 128(1):1–5, October 1, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901916.

Koumoutsakos:1997:IAE

[Kou97] P. Koumoutsakos. Inviscid axisymmetrization of an elliptical vortex. Journal of Computational Physics, 138(2):821–857, December 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197957491.

Keister:1997:UBP

[KP97a] B. D. Keister and W. N. Polyzou. Useful bases for problems in nuclear and particle physics. Journal of Computational Physics, 134(2):231–235, July 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956886.

Kelecy:1997:DFS

[KP97b] F. J. Kelecy and R. H. Pletcher. The development of a free surface capturing approach for multidimensional free surface flows in closed containers. Journal of Computational Physics, 138(2):939– 980, December 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999197958472.

Kwak:1998:ATE

[KP98] S. Kwak and C. Pozrikidis. Adaptive triangulation of evolving, closed, or open surfaces by the advancing-front method. Journal of Computational Physics, 145(1):61–88, September 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960302.

Knoll:1992:DNS

[KPC92] D. A. Knoll, A. K. Prinja, and R. B. Campbell. A direct Newton solver for the two-dimensional tokamak edge plasma fluid equations. Journal of Computational Physics, 102(2):424–425, October 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290391B.

Knoll:1993:DNS

[KPC93] D. A. Knoll, A. K. Prinja, and R. B. Campbell. A direct Newton solver for the two-dimensional Tokamak edge plasma fluid equations. Journal of Computational Physics, 104(2):418–426, February 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183710405.

Khirwadkar:1997:DGC

[KPCJ97] S. S. Khirwadkar, P. S. Pathak, S. Chaturvedi, and P. I. John. 2-D guiding centre simulation of toroidal electron clouds. Journal of Computational Physics, 132(2):291–298, April 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196956363.

Kobayashi:1999:CFV

[KPP99] Marcelo H. Kobayashi, José M. C. Pereira, and José C. F. Pereira. A conservative finite-volume second-order-accurate projection method on hybrid unstructured grids. Journal of Computational Physics, 150(1):40–75, March 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961630.

Karur:1995:OCN

[KR95] Sriganesh R. Karur and P. A. Ramachandran. Orthogonal collocation in the nonconforming boundary element method. Journal of Computational Physics, 121(2):373–382, October 15, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999195902384.

Kremeyer:1998:CAI

[Kre98] Kevin Kremeyer. Cellular automata investigations of binary solidification. Journal of Computational Physics, 142(1):243– 262, May 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198959265.

Korosmezey:1993:TGD

[KRGv93] Á. Körösmezey, C. E. Rasmussen, T. I. Gombosi, and Bram vanLeer. Transport of gyration-dominated space plasmas of thermal origin 2: Numerical solution. Journal of Computational Physics, 109(1):16–29, November 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711952.

Kylling:1992:EYA

[KS92] Arve Kylling and Knut Stamnes. Efficient yet accurate solution of the linear transport equation in the presence of internal sources: The exponential-linear-in-depth approximation. Journal of Computational Physics, 102(2):265–276, October 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192903715.

Klibanov:1994:UPK

[KS94] Michael V. Klibanov and Paul E. Sacks. Use of partial knowledge of the potential in the phase problem of inverse scattering. Journal of Computational Physics, 112(2):273–281, June 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710990.

Kale:1999:NGS

[KSB^{+}99] Laxmikant Kalé, Robert Skeel, Milind Bhandarkar, Robert Brunner, Attila Gursoy, Neal Krawetz, James Phillips, Aritomo Shinozaki, Krishnan Varadarajan, and Klaus Schulten. NAMD2: Greater scalability for parallel molecular dynamics. Journal of Computational Physics, 151(1):283–312, May 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962010.

Kurki-Suonio:1990:TMO

[KST90] T. Kurki-Suonio and T. Tajima. Transport model of optical beams in a plasma. Journal of Computational Physics, 90(1): 268, September 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919090211I.

Kurki-Suonio:1991:TMO

[KST91] T. Kurki-Suonio and T. Tajima. Transport model of optical beams in a plasma. Journal of Computational Physics, 94(1): 186–200, May 1991. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999191901428.

Koga:1995:ABD

[KT95] J. K. Koga and T. Tajima. The δf algorithm for beam dynamics. Journal of Computational Physics, 116(2):314–329, February 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185710303.

Kloucek:1999:TDN

[KT99] Petr Kloucek and Frank R. Toffoletto. The three dimensional non-conforming finite element solution of the Chapman–Ferraro problem. Journal of Computational Physics, 150(2):549–560, April 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199919996206X.

Kosloff:1993:MCP

[KTE93] Dan Kosloff and Hillel Tal-Ezer. A modified Chebyshev pseudospectral method with an O(N^{−1}) time step restriction. Journal of Computational Physics, 104(2):457–469, February 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710442.

Korivi:1994:AFI

[KTN^{+}94] Vamshi Mohan Korivi, Arthur C. Taylor III, Perry A. Newman, Gene W. Hou, and Henry E. Jones. An approximately factored incremental strategy for calculating consistent discrete aerodynamic sensitivity derivatives. Journal of Computational Physics, 113(2):336–346, August 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711405.

Kollar:1994:GFS

[KU94] J. Kollár and B. Ujfalussy. General formulas for the Slater–Koster tables. Journal of Computational Physics, 110(1):187–189, January 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002199918471014X.

Ku:1995:SFC

[Ku95] Hwar-Ching Ku. Solution of flow in complex geometries by the pseudospectral element method. Journal of Computational Physics, 117(2):215–227, March 15, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710601.

Kuhl:1996:MCS

[Kuh96] Nelson M. Kuhl. Monte Carlo simulation of transport. Journal of Computational Physics, 129(1):170–180, November 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196902417.

Kulikovsky:1995:MAS

[Kul95] A. A. Kulikovsky. A more accurate Scharfetter–Gummel algorithm of electron transport for semiconductor and gas discharge simulation. Journal of Computational Physics, 119(1):149–155, June 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185711230.

Kupferman:1998:SVF

[Kup98] Raz Kupferman. Simulation of viscoelastic fluids: Couette– Taylor flow. Journal of Computational Physics, 147(1):22–59, November 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198959344.

Kooper:1995:AIU

[KvdVPG95] M. N. Kooper, H. A. van der Vorst, S. Poedts, and J. P. Goedbloed. Application of the implicitly updated Arnoldi method with a complex shift-and-invert strategy in MHD. Journal of Computational Physics, 118(2):320–328, May 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711023.

Khalil:1992:VCC

[KW92] M. Khalil and P. Wesseling. Vertex-centered and cellcentered multigrid for interface problems. Journal of Computational Physics, 98(1):1–10, January 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290168X.

Koomullil:1993:NMA

[KW93] G. P. Koomullil and Z. U. A. Warsi. Numerical mapping of arbitrary domains using spectral methods. Journal of Computational Physics, 104(1):251–262, January 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710247.

Kreitmeier:1994:CSC

[KWG94] S. Kreitmeier, M. Wittkop, and D. Göritz. Computer simulation of coil deformation. Journal of Computational Physics, 112(2): 267–272, June 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184710989.

Kreitmeier:1997:LSS

[KWT^{+}97] Stefan Kreitmeier, Markus Wittkop, Hans L. Trautenberg, Thomas Hölzl, and Dietmar Göritz. Large-scale simulations on polymer melts. Journal of Computational Physics, 133(1):181– 185, May 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999197956503.

Lohner:1990:VPTb

[LA90] Rainald Löhner and John Ambrosiano. A vectorized particle tracer for unstructured grids. Journal of Computational Physics, 91(1):22–31, November 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090002I.

Lippert:1998:MCI

[LAE98] Ross A. Lippert, T. A. Arias, and Alan Edelman. Multiscale computation with interpolating wavelets. Journal of Computational Physics, 140(2):278–310, March 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198958855.

Langdon:1997:ICC

[Lan97] A. Bruce Langdon. Introduction to “Clouds-in-Clouds, Cloudsin-Cells Physics for Many-Body Simulation”. Journal of Computational Physics, 135(2):139–140, August 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957636.

Larrouturou:1991:HPM

[Lar91] B. Larrouturou. How to preserve the mass fractions positivity when computing compressible multi-component flows. Journal of Computational Physics, 95(1):59–84, July 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190253H.

Lavery:1990:CSO

[Lav90] John E. Lavery. Calculation of shocked one-dimensional flows on abruptly changing grids by mathematical programming. Journal of Computational Physics, 86(1):1–17, January 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090089J.

Lavery:1993:CCD

[Lav93] John E. Lavery. Capturing contact discontinuities in steady-state conservation laws. Journal of Computational Physics, 108(1):59– 72, September 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999183711630.

Lax:1997:IHR

[Lax97] Peter Lax. Introduction to “High Resolution Schemes for Hyperbolic Conservation Laws”. Journal of Computational Physics, 135(2):259, August 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999197957259.

Layton:1990:PAD

[Lay90] William Layton. On the principal axes of diffusion in difference schemes for 2d transport problems. Journal of Computational Physics, 90(2):336–347, October 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190901706.

Langnau:1993:PTL

[LB93] Alex Langnau and Stanley J. Brodsky. Perturbation theory in light-cone quantization. Journal of Computational Physics, 109(1):84–92, November 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999183712015.

Lapenta:1994:DSC

[LB94] Giovanni Lapenta and Jeremiah U. Brackbill. Dynamic and selective control of the number of particles in kinetic plasma simulations. Journal of Computational Physics, 115(1):213–227, November 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184711880.

LeProvost:1994:CTN

[LBB94] Christian LeProvost, Christine Bernier, and Eric Blayo. A comparison of two numerical methods for integrating a quasigeostrophic multilayer model of ocean circulations: Finite element and finite difference methods. Journal of Computational Physics, 110(2):341–359, February 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918471031X.

Leforestier:1991:CDP

[LBC^{+}91] C. Leforestier, R. H. Bisseling, C. Cerjan, M. D. Feit, R. Friesner, A. Guldberg, A. Hammerich, G. Jolicard, W. Karrlein, H.-D Meyer, N. Lipkin, O. Roncero, and R. Kosloff. A comparison of different propagation schemes for the time dependent Schrödinger equation. Journal of Computational Physics, 94(1):59–80, May 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190137A.

Lang:1993:NTB

[LBC^{+}93] Andreas Lang, Hans Babovsky, Wolfgang Cassing, Ulrich Mosel, Hans-Georg Reusch, and Klaus Weber. A new treatment of Boltzmann-like collision integrals in nuclear kinetic equations. Journal of Computational Physics, 106(2):391–396, June 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183711162.

Luo:1998:FMF

[LBL98] Hong Luo, Joseph D. Baum, and Rainald Löhner. A fast, matrixfree implicit method for compressible flows on unstructured grids. Journal of Computational Physics, 146(2):664–690, November 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198960764.

Lee:1996:ISS

[LCM96] C. K. B. Lee, R. C. Crawford, and J. M. McDonough. Improved solutions to the small strain continuum equations using a modified Engquist filter. Journal of Computational Physics, 123(1):119–126, January 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900108.

Legras:1993:CCS

[LD93] Bernard Legras and David G. Dritschel. A comparison of the contour surgery and pseudo-spectral methods. Journal of Computational Physics, 104(2):287–302, February 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710314.

Lambert:1996:MBA

[LDB96] Christophe G. Lambert, Thomas A. Darden, and John A. Board, Jr. A multipole-based algorithm for efficient calculation of forces and potentials in macroscopic periodic assemblies of particles. Journal of Computational Physics, 126(2):274–285, July 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196901370.

Lou:1998:NSC

[LDB98] T. Lou, D. C. Dahlby, and D. Baganoff. A numerical study comparing kinetic flux-vector splitting for the Navier– Stokes equations with a particle method. Journal of Computational Physics, 145(2):489–510, September 20, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960405.

Li:1990:BES

[LE90] B. Q. Li and J. W. Evans. Boundary element solution of heat convection-diffusion problems. Journal of Computational Physics, 89(1):254, July 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919090137P.

Li:1991:BES

[LE91] B. Q. Li and J. W. Evans. Boundary element solution of heat convection-diffusion problems. Journal of Computational Physics, 93(2):255–272, April 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190181J.

Leboucher:1999:MSB

[Leb99] L. Leboucher. Monotone scheme and boundary conditions for finite volume simulation of magnetohydrodynamic internal flows at high Hartmann number. Journal of Computational Physics, 150(1):181–198, March 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961708.

Lee:1998:DCL

[Lee98a] Dohyung Lee. Design criteria for local Euler preconditioning. Journal of Computational Physics, 144(2):423–459, August 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198959939.

Lee:1998:DLN

[Lee98b]  Dohyung Lee.  The design of local Navier–Stokes pre-

conditioning for compressible flow. Journal of Computational Physics, 144(2):460–483, August 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959940.

Lehner:1999:DAW

[Leh99] Luis Lehner. A dissipative algorithm for wave-like equations in the characteristic formulation. Journal of Computational Physics, 149(1):59–74, February 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919896137X.

Lesnic:1996:ITC

[LEI96] D. Lesnic, L. Elliott, and D. B. Ingham. Identification of the thermal conductivity and heat capacity in unsteady nonlinear heat conduction problems using the boundary element method. Journal of Computational Physics, 126(2):410–420, July 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901461.

Lele:1992:CFD

[Lel92] Sanjiva K. Lele. Compact finite difference schemes with spectrallike resolution. Journal of Computational Physics, 103(1):16–42, November 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290324R.

LeVeque:1997:WPA

[LeV97] Randall J. LeVeque. Wave propagation algorithms for multidimensional hyperbolic systems. Journal of Computational Physics, 131(2):327–353, March 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919695603X.

LeVeque:1998:BST

[LeV98] Randall J. LeVeque. Balancing source terms and flux gradients in high-resolution Godunov methods: The quasisteady wave-propagation algorithm. Journal of Computational Physics, 146(1):346–365, October 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960582.

Lou:1996:PIF

[LF96] John Z. Lou and Robert Ferraro. A parallel incompressible flow solver package with a parallel multigrid elliptic kernel. Journal of Computational Physics, 125(1):225–243, April 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919690090X.

Lawson:1990:HFB

[LG90] William S. Lawson and Perry C. Gray. Heat flow between species in one-dimensional particle plasma simulations. Journal of Computational Physics, 91(1):252, November 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090025V.

Lawson:1991:HFB

[LG91] William S. Lawson and Perry C. Gray. Heat flow between species in one-dimensional particle plasma simulations. Journal of Computational Physics, 95(1):195–211, July 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190259N.

Lyard:1994:OMB

[LG94] F. Lyard and M. L. Genco. Optimisation methods for bathymetry and open boundary conditions in a finite element model of ocean tides. Journal of Computational Physics, 114(2):234–256, October 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999108800010.

Laboudigue:1992:NSF

[LGC92] Bruno Laboudigue, Vincent Giovangigli, and Sébastien Candel. Numerical solution of a free-boundary problem in hypersonic flow theory: Nonequilibrium viscous shock layers. Journal of Computational Physics, 102(2):297–309, October 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192903737.

Loh:1990:NLM

[LH90] C. Y. Loh and W. H. Hui. A new Lagrangian method for steady supersonic flow computation I. Godunov scheme. Journal of Computational Physics, 89(1):207–240, July 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090123I.

Lee:1992:RDS

[LH92] D. K. Lee and S. P. Hirshman. Rapid determination of a straight magnetic coordinate system for stellarator configurations. Journal of Computational Physics, 101(2):330–333, August 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290010V.

Lepage:1995:SAG

[LH95] C. Y. Lepage and W. H. Hui. A shock-adaptive Godunov scheme based on the generalised Lagrangian formulation. Journal of Computational Physics, 122(2):291–299, December 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185712144.

Lee:1990:STD

[LHI90] Min Eig Lee, S. I. Hariharan, and Nathan Ida. Solving timedependent two-dimensional eddy current problems. Journal of Computational Physics, 89(2):319–348, August 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090147S.

Li:1997:WEH

[Li97] Yuguo Li. Wavenumber-extended high-order upwind-biased finite-difference schemes for convective scalar transport. Journal of Computational Physics, 133(2):235–255, May 15, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956497.

Liffman:1992:DSM

[Lif92] Kurt Liffman. A direct simulation Monte–Carlo method for cluster coagulation. Journal of Computational Physics, 100(1):116– 127, May 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290314O.

Liffman:1996:CCS

[Lif96] Kurt Liffman. Comments on a collocation spectral solver for the Helmholtz equation. Journal of Computational Physics, 128(1):254–258, October 1, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999196902077.

Levin:1997:SFP

[LIH97] Julia G. Levin, Mohamed Iskandarani, and Dale B. Haidvogel. A spectral filtering procedure for eddy-resolving simulations with a spectral element ocean model. Journal of Computational Physics, 137(1):130–154, October 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957971.

Lilly:1997:ICD

[Lil97] Douglas K. Lilly. Introduction to “Computational Design for Long-Term Numerical Integration of the Equations of Fluid Motion: Two-Dimensional Incompressible Flow. Part I”. Journal of Computational Physics, 135(2):101–102, August 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957223.

Lin:1995:DAF

[Lin95] Hong-Chia Lin. Dissipation additions to flux-difference splitting. Journal of Computational Physics, 117(1):20–27, March 1, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185710406.

Liou:1995:ELM

[Lio95] Meng-Sing Liou. An extended Lagrangian method. Journal of Computational Physics, 118(2):294–309, May 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918571100X.

Liou:1996:SAA

[Lio96] Meng-Sing Liou. A sequel to AUSM: AUSM^{+}. Journal of Computational Physics, 129(2):364–382, December 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196902569.

Liu:1996:SAB

[Liu96a] Jun Liu. A stability analysis on Beck’s procedure for inverse heat conduction problems. Journal of Computational Physics, 123(1): 65–73, January 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196900054.

Liu:1996:FAN

[Liu96b] Yen Liu. Fourier analysis of numerical algorithms for the Maxwell equations. Journal of Computational Physics, 124(2):396–416, March 15, 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196900686.

Li:1996:NST

[LJG96] X. L. Li, B. X. Jin, and J. Glimm. Numerical study for the three-dimensional Rayleigh–Taylor instability through the TVD/AC scheme and parallel computation. Journal of Computational Physics, 126(2):343–355, July 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901424.

Lu:1992:HCGa

[LJT92a] I.-Tai Lu, H. K. Jung, and C. M. Tsai. Hybrid combinations of global and local operators for solving Helmholtz and Poisson equations. Journal of Computational Physics, 101(2):452, August 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290026U.

Lu:1992:HCGb

[LJT92b] I-Tai Lu, H. K. Jung, and C. M. Tsai. Hybrid combinations of global and local operators for solving Helmholtz and Poisson equations. Journal of Computational Physics, 103(2):390–401, December 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290410Z.

Lee:1990:NAS

[LK90] Hyuk Lee and In Seok Kang. Neural algorithm for solving differential equations. Journal of Computational Physics, 91(1):110– 131, November 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919090007N.

Li:1994:EMD

[LK94a] Tsung L. Li and Kelin J. Kuhn. Effects of mass discontinuity on the numerical solutions to quantum wells using the effective mass equation. Journal of Computational Physics, 110(2):292– 300, February 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184710266.

Li:1994:FES

[LK94b] Tsung L. Li and Kelin J. Kuhn. Finite element solutions to GaAs–AlAs quantum wells with connection matrices at heterojunctions. Journal of Computational Physics, 115(2):288–295, December 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199918471196X.

Liu:1998:NSI

[LK98] H. Liu and K. Kawachi. A numerical study of insect flight. Journal of Computational Physics, 146(1):124–156, October 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198960193.

Liu:1999:NSU

[LK99] H. Liu and K. Kawachi. A numerical study of undulatory swimming. Journal of Computational Physics, 155(2):223–247, November 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999199963416.

Ley-Koo:1990:NME

[LKB90] E. Ley-Koo and Carlos F. Bunge. New method to evaluate atomic electron-repulsion integrals. Journal of Computational Physics, 91(1):240–245, November 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090013Q.

Lomtev:1999:DGA

[LKK99] I. Lomtev, R. M. Kirby, and G. E. Karniadakis. A discontinuous Galerkin ALE method for compressible viscous flows in moving domains. Journal of Computational Physics, 155(1):128–159, October 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999199963313.

Landl:1990:DET

[LKLE90] Gerhard Landl, F. Kauffmann, Thomas Langthaler, and Heinz W. Engi. Distribution of event times in time-resolved fluorescence: the exponential series approach — algorithm, regularization, analysis. Journal of Computational Physics, 90(1): 269, September 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919090214L.

Lerbinger:1991:NSI

[LL91] K. Lerbinger and J. F. Luciani. A new semi-implicit method for MHD computations. Journal of Computational Physics, 97(2):444–459, December 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999191900089.

Lyster:1992:FIPa

[LL92a] P. M. Lyster and J.-N Leboeuf. A fluid-ion and particleelectron model for low-frequency plasma instabilities. Journal of Computational Physics, 101(1):227, July 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192900565.

Lyster:1992:FIPb

[LL92b] P. M. Lyster and J.-N. Leboeuf. A fluid-ion and particleelectron model for low-frequency plasma instabilities. Journal of Computational Physics, 102(1):180–193, September 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999105800142.

Liu:1993:HOF

[LL93a] C. Liu and Z. Liu. High order finite difference and multigrid methods for spatially evolving instability in a planar channel. Journal of Computational Physics, 106(1):92–100, May 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710934.

Loh:1993:LSS

[LL93b] Ching-Yuen Loh and Meng-Sing Liou. Lagrangian solution of supersonic real gas flows. Journal of Computational Physics, 104(1):150–161, January 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918371017X.

Loh:1994:NLM

[LL94] Ching-Yuen Loh and Meng-Sing Liou. A new Lagrangian method for three-dimensional steady supersonic flows. Journal of Computational Physics, 113(2):224–248, August 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711326.

Liu:1995:MMB

[LL95] Chaoqun Liu and Zhining Liu. Multigrid mapping and box relaxation for simulation of the whole process of flow transition in 3D boundary layers. Journal of Computational Physics, 119(2): 325–341, July 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999185711382.

Liang:1996:FVN

[LL96] M. C. Liang and C. W. Lan. A finite-volume/Newton method for a two-phase heat flow problem using primitive variables and collocated grids. Journal of Computational Physics, 127(2):330– 345, September 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196901795.

Lan:1999:MMI

[LL99] C. W. Lan and M. C. Liang. Multigrid methods for incompressible heat flow problems with an unknown interface. Journal of Computational Physics, 152(1):55–77, June 10, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962344.

Lappas:1992:ALM

[LLD92] Tasso Lappas, Anthony Leonard, and Paul E. Dimotakis. An adaptive Lagrangian method for computing 1D reacting and non-reacting flows. Journal of Computational Physics, 102(2): 424, October 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290389G.

Lappas:1993:ALM

[LLD93] Tasso Lappas, Anthony Leonard, and Paul E. Dimotakis. An adaptive Lagrangian method for computing 1D reacting and nonreacting flows. Journal of Computational Physics, 104(2):361– 376, February 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999183710375.

Landl:1991:DET

[LLEK91] Gerhard Landl, Thomas Langthaler, Heinz W. Englt, and Harald F. Kauffmann. Distribution of event times in time-resolved fluorescence: The exponential series approach-algorithm, regularization, analysis. Journal of Computational Physics, 95(1): 1–28, July 1991. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919190250O.

Lough:1998:EBI

[LLK98] M. F. Lough, S. H. Lee, and J. Kamath. An efficient boundary integral formulation for flow through fractured porous media. Journal of Computational Physics, 143(2):462–483, July 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198958582.

Lewis:1999:LSQ

[LLLY99] James P. Lewis, Shubin Liu, Tai-Sung Lee, and Weitao Yang. A linear-scaling quantum mechanical investigation of cytidine deaminase. Journal of Computational Physics, 151(1):242–263, May 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999199962198.

Leyman:1994:ICP

[LLR94] Vladimir G. Leyman, Svetlana P. Litvintseva, and Igor D. Rodionov. Investigation of charged particle beam stability: Numerical method and physical results. Journal of Computational Physics, 115(1):86–106, November 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711806.

Liu:1994:NCC

[LLT94] Yingjie Liu, Lixin Liu, and Tao Tang. The numerical computation of connecting orbits in dynamical systems: a rational spectral approach. Journal of Computational Physics, 111(2):373– 380, April 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184710709.

Le:1990:IFS

[LM90a] Hung Le and Parviz Moin. An improvement of fractionalstep methods for the incompressible Navier–Stokes equations. Journal of Computational Physics, 89(1):253, July 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090134M.

Lee:1990:MFE

[LM90b] Robert L. Lee and Niel K. Madsen. A mixed finite element formulation for Maxwell’s equations in the time domain. Journal of Computational Physics, 88(2):284–304, June 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090181Y.

Li:1990:BAM

[LM90c] Zi-Cai Li and Rudolf Mathon. Boundary approximation methods for solving elliptic problems on unbounded domains. Journal of Computational Physics, 89(2):414–431, August 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090150Y.

Le:1991:IFS

[LM91] Hung Le and Parviz Moin. An improvement of fractional step methods for the incompressible Navier–Stokes equations. Journal of Computational Physics, 92(2):369–379, February 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999191902157.

Luchini:1992:FCM

[LM92] P. Luchini and F. Manzo. A fast conformal mapping algorithm with no FFT. Journal of Computational Physics, 101(2):368– 374, August 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290014P.

Lee:1995:INF

[LM95] C. K. B. Lee and J. M. McDonough. Implementation of a nonlinear filter with a Lagrangian formulation for the treatment of very high pressure reflected shocks. Journal of Computational Physics, 117(2):289–299, March 15, 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710674.

LeTallec:1997:CBN

[LM97] Patrick Le Tallec and François Mallinger. Coupling Boltzmann and Navier–Stokes equations by half fluxes. Journal of Computational Physics, 136(1):51–67, September 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957296.

Liu:1994:ILM

[LMB94] F. Liu, A. C. McIntosh, and J. Brindley. An implicit Lagrangian method for solving one- and two-dimensional gasdynamic equations. Journal of Computational Physics, 110(1):112–133, January 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999184710096.

Lionello:1999:SAW

[LML99] Roberto Lionello, Zoran Mikić, and Jon A. Linker. Stability of algorithms for waves with large flows. Journal of Computational Physics, 152(1):346–358, June 10, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962502.

LeQuere:1992:CCAa

[LMP92a] P. Le Quéré, R. Masson, and P. Perrot. A Chebyshev collocation algorithm for 2-D non-Boussinesq convection. Journal of Computational Physics, 101(1):229, July 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192900716.

LeQuere:1992:CCAb

[LMP92b] P. Le Quéré, R. Masson, and P. Perrot. A Chebyshev collocation algorithm for 2D non-Boussinesq convection. Journal of Computational Physics, 103(2):320–335, December 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290404M.

Lionello:1998:MSC

[LMS98] Roberto Lionello, Zoran Mikić, and Dalton D. Schnack. Magnetohydrodynamics of solar coronal plasmas in cylindrical geometry. Journal of Computational Physics, 140(1):172–201, February 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198958417.

Lange:1999:SCT

[LNR99] U. Lange, K. Nandakumar, and H. Raszillier. Symbolic computation as a tool for high-order long-wave stability analysis of thin film flows with coupled transport processes. Journal of Computational Physics, 150(1):1–16, March 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961605.

Lafaurie:1994:MMF

[LNS^{+}94] Bruno Lafaurie, Carlo Nardone, Ruben Scardovelli, Stéphane Zaleski, and Gianluigi Zanetti. Modelling merging and fragmentation in multiphase flows with SURFER. Journal of Computational Physics, 113(1):134–147, July 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711235.

Lafon:1991:HOFa

[LO91a] F. Lafon and S. Osher. High order filtering methods for approximating hyperbolic systems for conservation laws. Journal of Computational Physics, 93(1):250–251, March 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190083W.

Lafon:1991:HOFb

[LO91b] F. Lafon and S. Osher. High order filtering methods for approximating hyperbolic systems of conservation laws. Journal of Computational Physics, 96(1):110–142, September 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190268P.

Lucas:1993:MAM

[LO93] T. R. Lucas and H. S. Oh. The method of auxiliary mapping for the finite element solutions of elliptic problems containing singularities. Journal of Computational Physics, 108(2):327–342, October 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183711861.

Lafon:1996:HOT

[LO96] F. Lafon and S. Osher. High order two dimensional nonoscillatory methods for solving Hamilton–Jacobi scalar equations. Journal of Computational Physics, 123(2):235–253, February 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900212.

Liu:1998:CEH

[LO98] Xu-Dong Liu and Stanley Osher. Convex ENO high order multidimensional schemes without field by field decomposition or staggered grids. Journal of Computational Physics, 142(2):304–330, May 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199919895937X.

Liu:1994:WEN

[LOC94] Xu-Dong Liu, Stanley Osher, and Tony Chan. Weighted essentially non-oscillatory schemes. Journal of Computational Physics, 115(1):200–212, November 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711879.

Loewenthal:1993:ILS

[Loe93] Dan Loewenthal. Improvements on the Lehmer–Schur root detection method. Journal of Computational Physics, 109(2):164–168, December 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199918371209X.

Lohner:1990:VPTa

[Löh90] Rainald Löhner. A vectorized particle tracer for unstructured grids. Journal of Computational Physics, 87(2):496, April 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999190902734.

Lohner:1995:RVS

[Löh95] Rainald Löhner. Robust, vectorized search algorithms for Interpolation on unstructured grids. Journal of Computational Physics, 118(2):380–387, May 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999185711072.

Lohner:1996:RST

[Löh96] Rainald Löhner. Regridding surface triangulations. Journal of Computational Physics, 126(1):1–10, June 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901151.

Lotstedt:1994:ICS

[Löt94] Per Lötstedt. Improved convergence to the steady state of the Euler equations by enhanced wave propagation. Journal of Computational Physics, 114(1):34–44, September 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711478.

Luppi:1990:SCA

[LP90] Jussi Luppi and Petri Pajunen. Symbolic computation and automatic FORTRAN code generation for eigenvalue determination by phase integral method. Journal of Computational Physics, 88 (1):15–30, May 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999190902402.

Li:1997:MMM

[LP97] Shengtai Li and Linda Petzold. Moving mesh methods with upwinding schemes for time-dependent PDEs. Journal of Computational Physics, 131(2):368–377, March 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196956119.

Lagaris:1996:RMN

[LPBS96] I. E. Lagaris, D. G. Papageorgiou, M. Braun, and S. A. Sofianos. A relaxation method for nonlocal and non-Hermitian operators. Journal of Computational Physics, 126(1):229–236, June 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002199919690131X.

Libersky:1993:HSL

[LPC^{+}93] Larry D. Libersky, Albert G. Petschek, Theodore C. Carney, Jim R. Hipp, and Firooz A. Allahdadi. High strain Lagrangian hydrodynamics: a three-dimensional SPH code for dynamic material response. Journal of Computational Physics, 109(1):67–75, November 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199918371199X.

Labik:1994:EGN

[LPMS94] Stanislav Labík, Roman Pospísil, Anatol Malijevský, and William Robert Smith. An efficient Gauss–Newton-like method for the numerical solution of the Ornstein–Zernike integral equation for a class of fluid models. Journal of Computational Physics, 115(1):12–21, November 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711740.

Lynch:1999:KBT

[LPP99] Gillian C. Lynch, John S. Perkyns, and B. Montgomery Pettitt. Kirkwood–Buff thermodynamics derived from grand canonical molecular dynamics and DRISM calculations. Journal of Computational Physics, 151(1):135–145, May 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961836.

Lubachevsky:1996:CPB

[LPR96] Boris D. Lubachevsky, Vladimir Privman, and Subhas C. Roy. Casting pearls ballistically: Efficient massively parallel simulation of particle deposition. Journal of Computational Physics, 126(1):152–164, June 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901278.

Levy:1993:URR

[LPvL93] David W. Levy, Kenneth G. Powell, and Bram van Leer. Use of a rotated Riemann solver for the two-dimensional Euler equations. Journal of Computational Physics, 106(2):201–214, June 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183711034.

Lomtev:1998:SHM

[LQK98] I. Lomtev, C. B. Quillen, and G. E. Karniadakis. Spectral/ hp methods for viscous compressible flows on unstructured 2D meshes. Journal of Computational Physics, 144(2):325–357, August 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999197958319.

Lu:1990:DVN

[LR90] Zhi Yun Lu and Timothy J. Ross. Diffusing-vortex numerical scheme for solving incompressible Navier–Stokes equations. Journal of Computational Physics, 91(1):250–251, November 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919090021R.

Lu:1991:DVN

[LR91] Zhi Yun Lu and Timothy J. Ross. Diffusing-vortex numerical scheme for solving incompressible Navier–Stokes equations. Journal of Computational Physics, 95(2):400–435, August 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190283Q.

Lindsay:1992:NCM

[LR92] K. A. Lindsay and C. E. Rooney. A note on compound matrices. Journal of Computational Physics, 103(2):472–477, December 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290417W.

Lowrie:1994:NSC

[LR94] R. B. Lowrie and P. L. Roe. On the numerical solution of conservation laws by minimizing residuals. Journal of Computational Physics, 113(2):304–308, August 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711375.

Linde:1998:MNP

[LR98] Timur Linde and Philip Roe. On a mistaken notion of “proper upwinding”. Journal of Computational Physics, 142(2):611–614, May 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198959502.

Leforestier:1990:CDP

[LRB^{+}90] C. Leforestier, O. Roncero, R. Bisseling, C. Cerjan, M. D. Feit, R. Friesner, A. Guldberg, A. Hammerich, R. Kosloff, G. Jolicard, W. Karrlein, H.-D Meyer, and N. Lipkin. A comparison of different propagation schemes for the time-dependent Schrödinger equation. Journal of Computational Physics, 89(2):490–491, August 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919090165W.

Ly:1999:SPD

[LRJ^{+}99] H. V. Ly, F. Reitich, M. R. Jolly, H. T. Banks, and K. Ito. Simulations of particle dynamics in magnetorheological fluids. Journal of Computational Physics, 155(1):160–177, October 10, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963350.

Lustig:1995:TFM

[LRW95] Steven R. Lustig, Sanjeev Rastogi, and Norman Wagner. Telescoping fast multipole methods using Chebyshev economization. Journal of Computational Physics, 122(2):317–322, December 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002199918571217X.

Lardner:1992:HSM

[LS92a] R. W. Lardner and Y. Song. A hybrid spectral method for the three-dimensional numerical modelling of nonlinear flows in shallow seas. Journal of Computational Physics, 100(2):322–334, June 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290239U.

Leclercq:1992:CMM

[LS92b] M. P. Leclercq and B. Stoufflet. Characteristic multigrid method application to solve the Euler equations with unstructured and unnested grids. Journal of Computational Physics, 102(2):424, October 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290387E.

Ling:1992:BCAa

[LS92c] F. H. Ling and G. Schmidt. Box-counting algorithm and dimensional analysis of a pulsar. Journal of Computational Physics, 98(2):348, February 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919290159V.

Ling:1992:BCAb

[LS92d] F. H. Ling and G. Schmidt. Box-counting algorithm and dimensional analysis of a pulsar. Journal of Computational Physics, 99(2):196–202, April 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999192902019.

Leclercq:1993:CMM

[LS93a] M. P. Leclercq and B. Stoufflet. Characteristic multigrid method application to solve the Euler equations with unstructured and unnested grids. Journal of Computational Physics, 104(2):329– 346, February 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999183710351.

Liou:1993:NFS

[LS93b] Meng-Sing Liou and Christopher J. Steffen, Jr. A new flux splitting scheme. Journal of Computational Physics, 107(1):23–39, July 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183711228.

Leimkuhler:1994:SNI

[LS94] Benedict J. Leimkuhler and Robert D. Skeel. Symplectic numerical integrators in constrained Hamiltonian systems. Journal of Computational Physics, 112(1):117–125, May 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710850.

LeVeque:1996:TDF

[LS96] Randall J. LeVeque and Keh-Ming Shyue. Two-dimensional front tracking based on high resolution wave propagation methods. Journal of Computational Physics, 123(2):354–368, February 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196900297.

Lopez:1998:ESP

[LS98] J. M. Lopez and Jie Shen. An efficient spectral-projection method for the Navier–Stokes equations in cylindrical geometries: I. Axisymmetric cases. Journal of Computational Physics, 139(2):308–326, January 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958721.

Lifschitz:1996:OIE

[LSB96] Alexander Lifschitz, W. Henry Suters, and J. Thomas Beale. The onset of instability in exact vortex rings with swirl. Journal of Computational Physics, 129(1):8–29, November 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196902302.

Labonia:1997:NSE

[LSLG97] G. Labonia, F. Stella, E. Leonardi, and G. Guj. A numerical study of the effect of free surface deformation on buoyancy and thermocapillary convection. Journal of Computational Physics, 132(1):34–50, March 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919695593X.

Lombardi:1999:TST

[LSRS99] James C. Lombardi, Jr., Alison Sills, Frederic A. Rasio, and Stuart L. Shapiro. Tests of spurious transport in smoothed particle hydrodynamics. Journal of Computational Physics, 152(2):687– 735, July 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999199962563.

Lee:1991:FET

[LT91] D. Lee and Y. M. Tsuei. A formula for estimation of truncation errors of convection terms in a curvilinear coordinate system. Journal of Computational Physics, 93(2):488, April 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190206Z.

Lee:1992:FET

[LT92] D. Lee and Y. M. Tsuei. A formula for estimation of truncation errors of convection terms in a curvilinear coordinate system. Journal of Computational Physics, 98(1):90–100, January 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290175X.

Lee:1993:HAG

[LT93] D. Lee and Y. M. Tsuei. A hybrid adaptive gridding procedure for recirculating fluid flow problems. Journal of Computational Physics, 108(1):122–141, September 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918371168X.

Luchini:1996:DAN

[LT96] Paolo Luchini and Renato Tognaccini. Direction-adaptive nonreflecting boundary conditions. Journal of Computational Physics, 128(1):121–133, October 1, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901990.

Lee:1993:BEM

[LTE93] S. H. Lee, D. W. Townley, and K. O. Eshun. A boundary element model of cathodic well casing protection. Journal of Computational Physics, 107(2):338–347, August 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711484.

Lu:1999:OWL

[Lu99] Ya Yan Lu. One-way large range step methods for Helmholtz waveguides. Journal of Computational Physics, 152(1):231–250, June 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999199962435.

Lubachevsky:1991:HSB

[Lub91] Boris D. Lubachevsky. How to simulate billiards and similar systems. Journal of Computational Physics, 94(2):255–283, June 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999191902227.

Lubachevsky:1993:WAB

[Lub93] B. D. Lubachevsky. Which algorithm is better? Journal of Computational Physics, 105(2):369–370, April 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710831.

Luchini:1990:DCM

[Luc90] Paolo Luchini. A deferred-correction multigrid algorithm based on a new smoother for the Navier–Stokes equations. Journal of Computational Physics, 89(1):251, July 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090128N.

Luchini:1991:DCM

[Luc91] Paolo Luchini. A deferred-correction multigrid algorithm based on a new smoother for the Navier–Stokes equations. Journal of Computational Physics, 92(2):349–368, February 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999191902146.

Luke:1999:FDM

[Luk99] Jonathan H. C. Luke. A finite difference method for dispersive linear waves with applications to simulating microwave pulses in water. Journal of Computational Physics, 148(1):199–226, January 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198961174.

Liu:1998:EIP

[LV98] Yen Liu and Marcel Vinokur. Exact integrations of polynomials and symmetric quadrature formulas over arbitrary polyhedral grids. Journal of Computational Physics, 140(1):122–147, February 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198958843.

Liou:1990:SIF

[LVS90] Meng-Sing Liou, Bram Van Leer, and Jian-Shun Shuen. Splitting of inviscid fluxes for real gases. Journal of Computational Physics, 87(1):1–24, March 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919090222M.

Loffhagen:1994:NNB

[LW94] D. Loffhagen and R. Winkler. A new nonstationary Boltzmann solver in self-consistent modelling of discharge pumped plasmas for excimer lasers. Journal of Computational Physics, 112(1): 91–101, May 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184710837.

Lang:1995:ADF

[LW95] Jens Lang and Artur Walter. An adaptive discontinuous finite element method for the transport equation. Journal of Computational Physics, 117(1):28–34, March 1, 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710418.

Lerat:1996:SCM

[LW96a] A. Lerat and Z. N. Wu. Stable conservative multidomain treatments for implicit Euler solvers. Journal of Computational Physics, 123(1):45–64, January 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900042.

Lewandowski:1996:TEC

[LW96b] A. C. Lewandowski and T. M. Wilson. A technique for embedding clusters in ionic crystals using the GAUSSIAN 92 quantum chemical program. Journal of Computational Physics, 129(1):233–242, November 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196902466.

Liska:1997:ACS

[LW97] Richard Liska and Burton Wendroff. Analysis and computation with stratified fluid models. Journal of Computational Physics, 137(1):212–244, October 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919795806X.

Lin:1993:UFV

[LWC93] San-Yih Lin, Tsuen-Muh Wu, and Yan-Shin Chin. Upwind finite-volume method with a triangular mesh for conservation laws. Journal of Computational Physics, 107(2):324–337, August 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183711472.

Lund:1998:GTI

[LWS98] Thomas S. Lund, Xiaohua Wu, and Kyle D. Squires. Generation of turbulent inflow data for spatially-developing boundary layer simulations. Journal of Computational Physics, 140(2):233– 258, March 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199919895882X.

Lee:1999:OAF

[LWT99] W. R. Lee, S. Wang, and K. L. Teo. An optimization approach to a finite dimensional parameter estimation problem in semiconductor device design. Journal of Computational Physics, 156(2):241–256, December 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963581.

Xu:1996:CAC

[lX96] Yu lin Xu. Calculation of the addition coefficients in electromagnetic multisphere-scattering theory. Journal of Computational Physics, 127(2):285–298, September 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901758.

Xu:1997:CAC

[lX97] Yu lin Xu. Calculation of the addition coefficients in electromagnetic multisphere-scattering theory: Volume 127, number 2 (1996), pages 285–298. Journal of Computational Physics, 134 (1):200, June 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999197956874.

Xu:1998:EEV

[lX98] Yu lin Xu. Efficient evaluation of vector translation coefficients in multiparticle light-scattering theories. Journal of Computational Physics, 139(1):137–165, January 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958678.

Leveque:1990:SNM

[LY90a] R. J. Leveque and H. C. Yee. A study of numerical methods for hyperbolic conservation laws with stiff source terms. Journal of Computational Physics, 86(1):187–210, January 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090097K.

Lin:1990:CST

[LY90b] San Yih Lin and Yisong Yang. Computation of superconductivity in thin films. Journal of Computational Physics, 89(2):257– 275, August 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919090144P.

Lyard:1999:DAW

[Lya99] F. H. Lyard. Data assimilation in a wave equation: a variational representer approach for the Grenoble Tidal Model. Journal of Computational Physics, 149(1):1–31, February 10, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959666.

Liu:1996:SCT

[LZ96] Feng Liu and Xiaoqing Zheng. A strongly coupled timemarching method for solving the Navier–Stokes and k–ω turbulence model equations with multigrid. Journal of Computational Physics, 128(2):289–300, October 15, 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196902119.

Li:1999:NSE

[LZG99] Zhilin Li, Hongkai Zhao, and Huajian Gao. A numerical study of electro-migration voiding by evolving level set functions on a fixed Cartesian grid. Journal of Computational Physics, 152(1):281– 304, June 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999199962496.

Liu:1998:PMM

[LZS98] C. Liu, X. Zheng, and C. H. Sung. Preconditioned multigrid methods for unsteady incompressible flows. Journal of Computational Physics, 139(1):35–57, January 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958599.

Ma:1993:SEB

[Ma93a] Hong Ma. A spectral element basin model for the shallow water equations. Journal of Computational Physics, 109(1):133–149, November 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999183712052.

Moore:1993:VPP

[MA93b] R. E. M. Moore and I. O. Angell. Voronoi polygons and polyhedra. Journal of Computational Physics, 105(2):301–305, April 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183710764.

Mashayek:1995:SFM

[MA95] F. Mashayek and N. Ashgriz. A spine-flux method for simulating free surface flows. Journal of Computational Physics, 122(2):367– 379, December 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999185712223.

Munipalli:1996:AGS

[MA96] Ramakanth Munipalli and Dale A. Anderson. An adaptive grid scheme using the boundary element method. Journal of Computational Physics, 127(2):452–463, September 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901862.

Mahadevan:1998:MLR

[MA98] Amala Mahadevan and David Archer. Modeling a limited region of the ocean. Journal of Computational Physics, 145(2):555–574, September 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198960417.

Madsen:1995:DPD

[Mad95] Niel K. Madsen. Divergence preserving discrete surface integral methods for Maxwell’s curl equations using non-orthogonal unstructured grids. Journal of Computational Physics, 119(1): 34–45, June 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199918571114X.

Mahesh:1998:FHO

[Mah98] Krishnan Mahesh. A family of high order finite difference schemes with good spectral resolution. Journal of Computational Physics, 145(1):332–358, September 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960223.

Making:1990:CTD

[Mak90a] Junichiro Making. Comparison of two different tree algorithms. Journal of Computational Physics, 88(2):393–408, June 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999190901865.

Makino:1990:VT

[Mak90b] Junichiro Makino. Vectorization of a treecode. Journal of Computational Physics, 87(1):148–160, March 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090231O.

Makino:1999:YAF

[Mak99] Junichiro Makino. Yet another fast multipole method without multipoles—pseudoparticle multipole method. Journal of Computational Physics, 151(2):910–920, May 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962265.

Malik:1990:NMH

[Mal90] M. R. Malik. Numerical methods for hypersonic boundary layer stability. Journal of Computational Physics, 86(2):376– 413, February 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919090106B.

Malara:1996:CPM

[Mal96a] Francesco Malara. A Chebyshev pseudospectral multidomain method for a boundary-layer problem. Journal of Computational Physics, 124(2):254–270, March 15, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900583.

Malevsky:1996:SCM

[Mal96b] Andrei V. Malevsky. Spline-characteristic method for simulation of convective turbulence. Journal of Computational Physics, 123(2):466–475, February 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900376.

Manke:1993:TPB

[Man93a] Joseph W. Manke. A tensor product B-spline method for numerical grid generation. Journal of Computational Physics, 108(1):15– 26, September 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999183711599.

Mann:1993:SPH

[Man93b] Patrick J. Mann. Smoothed particle hydrodynamics applied to relativistic spherical collapse. Journal of Computational Physics, 107(1):188–198, July 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711356.

Mao:1992:TDFb

[Mao92a] De-Kang Mao. A treatment of discontinuities for finite difference methods. Journal of Computational Physics, 103(2):359– 369, December 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290407P.

Mao:1992:TDFa

[Mao92b] De-Kang Mao. A treatment of discontinuities for finite difference methods in the two-dimensional case. Journal of Computational Physics, 102(2):424, October 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290390K.

Marinos:1990:DMS

[Mar90] Aristides Th. Marinos. On a direct method for solving Helmholtz’s type equations in 3-d rectangular regions. Journal of Computational Physics, 88(1):62–85, May 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090242S.

Marder:1993:EBI

[Mar93] Barry Marder. Eliminating boundary-induced errors in the pseudo-current method. Journal of Computational Physics, 104(1):273, January 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999183710272.

Marinos:1994:NEC

[Mar94a] Aristides Th. Marinos. Numerical experiments on compact computational schemes for solving the first biharmonic problem in rectangles. Journal of Computational Physics, 115(2):406–422, December 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199918471206X.

Marx:1994:TIS

[Mar94b] Yves P. Marx. Time integration schemes for the unsteady incompressible Navier–Stokes equations. Journal of Computational Physics, 112(1):182–209, May 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710916.

Margolin:1997:IAL

[Mar97] L. G. Margolin. Introduction to “An Arbitrary Lagrangian– Eulerian Computing Method for All Flow Speeds”. Journal of Computational Physics, 135(2):198–202, August 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957272.

Mott:1992:AST

[MAS92] Peter H. Mott, Ali S. Argon, and Ulrich W. Suter. The atomic strain tensor. Journal of Computational Physics, 101(1):140–150, July 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999192900484.

Masad:1996:ACE

[Mas96] J. A. Masad. Analysis and computation of extremum points with application to boundary-layer stability. Journal of Computational Physics, 127(2):464–472, September 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901874.

Matthews:1994:CAM

[Mat94] Alan P. Matthews. Current advance method and cyclic leapfrog for 2D multispecies hybrid plasma simulations. Journal of Computational Physics, 112(1):102–116, May 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710849.

Mattiussi:1997:AFV

[Mat97] Claudio Mattiussi. An analysis of finite volume, finite element, and finite difference methods using some concepts from algebraic topology. Journal of Computational Physics, 133(2):289– 309, May 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999197956564.

Maury:1999:DSF

[Mau99] B. Maury. Direct simulations of 2D fluid-particle flows in biperiodic domains. Journal of Computational Physics, 156(2):325–351, December 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999199963659.

Mavriplis:1990:AMG

[Mav90] Dimitri J. Mavriplis. Adaptive mesh generation for viscous flows using triangulation. Journal of Computational Physics, 90(2):271–291, October 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919090167Y.

Mavriplis:1995:AFD

[Mav95] Dimitri J. Mavriplis. An advancing front Delaunay triangulation algorithm designed for robustness. Journal of Computational Physics, 117(1):90–101, March 1, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710479.

Mavriplis:1998:MSV

[Mav98] D. J. Mavriplis. Multigrid strategies for viscous flow solvers on anisotropic unstructured meshes. Journal of Computational Physics, 145(1):141–165, September 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960363.

Mayo:1992:REVb

[May92a] Anita Mayo. The rapid evaluation of volume integrals of potential theory on general regions. Journal of Computational Physics, 99 (1):181, March 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290301E.

Mayo:1992:REVa

[May92b] Anita Mayo. The rapid evaluation of volume integrals of potential theory on general regions. Journal of Computational Physics, 100(2):236–245, June 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290231M.

Mazzone:1994:MDS

[Maz94] A. M. Mazzone. Molecular dynamics simulations of the effects of damage relaxation following high-energy implants. Journal of Computational Physics, 112(2):247–252, June 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710965.

Mazur:1997:CMD

[Maz97] Alexy K. Mazur. Common molecular dynamics algorithms revisited: Accuracy and optimal time steps of Störmer–leapfrog integrators. Journal of Computational Physics, 136(2):354–365, September 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999197957405.

Moody:1992:CMF

[MB92] R. O. Moody and M. J. Baines. A constrained moving finite element solution of the one-dimensional oxygen diffusion with absorption problem. Journal of Computational Physics, 103(2):442– 449, December 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290414T.

Mukhopadhyay:1994:SOA

[MB94] Supratik Mukhopadhyay and B. C. Basu. Second-order accurate no-slip conditions for solving problems of incompressible viscous flows. Journal of Computational Physics, 111(1):53–61, March 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999184710436.

Mittal:1996:DNS

[MB96] R. Mittal and S. Balachandar. Direct numerical simulation of flow past elliptic cylinders. Journal of Computational Physics, 124(2):351–367, March 15, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900650.

Minion:1997:PUR

[MB97] Michael L. Minion and David L. Brown. Performance of underresolved two-dimensional incompressible flow simulations, II. Journal of Computational Physics, 138(2):734–765, December 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197958435.

Malamataris:1999:CAA

[MB99] N. A. Malamataris and V. Bontozoglou. Computer aided analysis of viscous film flow along an inclined wavy Wall. Journal of Computational Physics, 154(2):372–392, September 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963192.

Marksteiner:1994:NTE

[MBF94] P. Marksteiner, E. Badralexe, and A. J. Freeman. Neumanntype expansion of Coulomb functions. Journal of Computational Physics, 111(1):49–52, March 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710424.

Micu:1997:NCC

[MBI^{+}97] Alexandru M. Micu, Babak Bagheri, Andrew V. Ilin, Ridgway Scott, and B. Montgomery Pettitt. Numerical considerations in the computation of the electrostatic free energy of interaction within the Poisson–Boltzmann theory. Journal of Computational Physics, 136(2):263–271, September 15, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957077.

Merriman:1994:MMJ

[MBO94] Barry Merriman, James K. Bence, and Stanley J. Osher. Motion of multiple junctions: a level set approach. Journal of Computational Physics, 112(2):334–363, June 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711053.

Mulyarchik:1994:ECS

[MBP94] Stepan G. Mulyarchik, Stanislav S. Bielawski, and Andrew V. Popov. Efficient computational schemes of the conjugate gradient method for solving linear systems. Journal of Computational Physics, 110(2):201–211, February 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710187.

Madabhushi:1993:DFC

[MBV93] Ravi K. Madabhushi, S. Balachandar, and S. P. Vanka. A divergence-free Chebyshev collocation procedure for incompressible flows with two non-periodic directions. Journal of Computational Physics, 105(2):199–206, April 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710685.

Morokoff:1995:QMC

[MC95] William J. Morokoff and Russel E. Caflisch. Quasi-Monte Carlo integration. Journal of Computational Physics, 122(2):218–230, December 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999185712090.

Mascagni:1995:FHQ

[MCPR95] Michael Mascagni, Steven A. Cuccaro, Daniel V. Pryor, and M. L. Robinson. A fast, high quality, and reproducible parallel lagged-Fibonacci pseudorandom number generator. Journal of Computational Physics, 119(2):211–219, July 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711308.

Mao:1990:MTB

[MD90] Zai-Sha Mao and A. E. Dukler. The motion of Taylor bubbles in vertical tubes. I. A numerical simulation for the shape and rise velocity of Taylor bubbles in stagnant and flowing liquid. Journal of Computational Physics, 91(1):132–160, November 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090008O.

Malik:1995:IST

[MD95] Nadeem A. Malik and Th. Dracos. Interpolation schemes for three-dimensional velocity fields from scattered data using Taylor expansions. Journal of Computational Physics, 119(2):231–243, July 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185711321.

Machiels:1998:NSR

[MD98] L. Machiels and M. O. Deville. Numerical simulation of randomly forced turbulent flows. Journal of Computational Physics, 145(1):246–279, September 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960065.

Mazur:1991:DTE

[MDA91] A. K. Mazur, V. E. Dorofeev, and R. A. Abagyan. Derivation and testing of explicit equations of motion for polymers described by internal coordinates. Journal of Computational Physics, 92(2):261–272, February 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190210C.

Mahajan:1991:ECPa

[MDB91a] Aparajit J. Mahajan, Earl H. Dowell, and Donald B. Bliss. Eigenvalue calculation procedure for an Euler/Navier–Stokes solver with application to flows over airfoils. Journal of Computational Physics, 93(2):488, April 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999191902072.

Mahajan:1991:ECPb

[MDB91b] Aparajit J. Mahajan, Earl H. Dowell, and Donald B. Bliss. Eigenvalue calculation procedure for an Euler/Navier–Stokes solver with application to flows over airfoils. Journal of Computational Physics, 97(2):398–413, December 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999191900067.

Moulton:1998:BBM

[MDH98] J. David Moulton, Joel E. Dendy, Jr., and James M. Hyman. The black box multigrid numerical homogenization algorithm. Journal of Computational Physics, 142(1):80–108, May 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959113.

Morel:1992:CCLa

[MDHW92a] J. E. Morel, J. E. Dendy, Jr., Michael L. Hall, and Stephen W. White. A cell-centered Lagrangian-mesh diffusion differencing scheme. Journal of Computational Physics, 101(1):229, July 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290069B.

Morel:1992:CCLb

[MDHW92b] J. E. Morel, J. E. Dendy, Jr., Michael L. Hall, and Stephen W. White. A cell-centered Lagrangian-mesh diffusion differencing scheme. Journal of Computational Physics, 103(2):286–299, December 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290402K.

Meister:1998:CDK

[Mei98] Andreas Meister. Comparison of different Krylov subspace methods embedded in an implicit finite volume scheme for the computation of viscous and inviscid flow fields on unstructured grids. Journal of Computational Physics, 140(2):311–345, March 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198958624.

Melezhik:1991:NMS

[Mel91] V. S. Melezhik. New method for solving multidimensional scattering problem. Journal of Computational Physics, 92(1):67– 81, January 1991. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919190292S.

Moore:1992:ALO

[MF92] Peter K. Moore and Joseph E. Flaherty. Adaptive local overlapping grid methods for parabolic systems in two space dimensions. Journal of Computational Physics, 98(1):54–63, January 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290172U.

Morris:1997:MLR

[MFZ97] Joseph P. Morris, Patrick J. Fox, and Yi Zhu. Modeling low Reynolds number incompressible flows using SPH. Journal of Computational Physics, 136(1):214–226, September 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957764.

Marshall:1996:MDV

[MG96] J. S. Marshall and J. R. Grant. A method for determining the velocity induced by highly anisotropic vorticity blobs. Journal of Computational Physics, 126(2):286–298, July 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901382.

Marshall:1997:LVC

[MG97a] J. S. Marshall and J. R. Grant. A Lagrangian vorticity collocation method for viscous, axisymmetric flows with and without swirl. Journal of Computational Physics, 138(2):302–330, December 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197958204.

Muzaferija:1997:FVC

[MG97b] Samir Muzaferija and David Gosman. Finite-volume CFD procedure and adaptive error control strategy for grids of arbitrary topology. Journal of Computational Physics, 138(2):766–787, December 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999197958538.

McKenney:1995:FPS

[MGM95] A. McKenney, L. Greengard, and A. Mayo. A fast Poisson solver for complex geometries. Journal of Computational Physics, 118(2):348–355, May 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999185711047.

Mansutti:1991:DVP

[MGP91] D. Mansutti, G. Graziani, and R. Piva. A discrete vector potential model for unsteady incompressible viscous flows. Journal of Computational Physics, 92(1):161–184, January 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190296W.

Mansour:1990:ISI

[MH90a] M. L. Mansour and A. Hamed. Implicit solution of the incompressible Navier–Stokes equations on a non-staggered grid. Journal of Computational Physics, 86(1):147–167, January 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090095I.

Mesina:1990:ISN

[MH90b] George Mesina and Charles Hall. Iterative solution of Navier– Stokes dual variable difference equations. Journal of Computational Physics, 91(1):252, November 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090026W.

Mesina:1991:ISN

[MH91] George Mesina and Charles Hall. Iterative solution of Navier– Stokes dual variable difference equations. Journal of Computational Physics, 96(1):71–98, September 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190266N.

Manninen:1992:MDV

[MH92] M. Manninen and H. Häkkinen. A method for detecting vacancy diffusion in molecular dynamics. Journal of Computational Physics, 100(1):197–199, May 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290320X.

Monaghan:1993:VPM

[MH93] J. J. Monaghan and R. J. Humble. Vortex particle methods for periodic channel flow. Journal of Computational Physics, 107(1):

152–159, July 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999183711320.

Mains:1994:ARF

[MH94] R. K. Mains and G. I. Haddad. An accurate re-formulation of the Wigner function method for quantum transport modeling. Journal of Computational Physics, 112(1):149–161, May 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710886.

Mikkelsen:1990:SFP

[Mik90] D. R. Mikkelsen. Solution of the Fokker–Planck equation with mixing of angular harmonics by beam-beam charge exchange. Journal of Computational Physics, 89(1):256, July 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090143O.

Mikkelsen:1991:SFP

[Mik91] D. R. Mikkelsen. Solution of the Fokker–Planck equation with mixing of angular harmonics by beam-beam charge exchange. Journal of Computational Physics, 93(2):336–342, April 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190187P.

Miller:1991:HSA

[Mil91] R. H. Miller. A horror story about integration methods. Journal of Computational Physics, 93(2):469–476, April 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190195Q.

Marti:1990:SHG

[MIM90] José M. Martí, José M. Ibáñez, and Juan A. Miralles. Stellar hydrodynamics with Glaister’s Riemann solver: An approach to the stellar collapse. Journal of Computational Physics, 90(1):262– 266, September 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919090205F.

Minion:1996:SGP

[Min96a] Michael L. Minion. On the stability of Godunov-projection methods for incompressible flow. Journal of Computational Physics, 123(2):435–449, February 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900352.

Minion:1996:PML

[Min96b] Michael L. Minion. A projection method for locally refined grids. Journal of Computational Physics, 127(1):158–178, August 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196901667.

Mitra:1992:UEO

[Mit92] Ambar K. Mitra. Use of eigenfunctions in the optimization of the grid for the boundary element method. Journal of Computational Physics, 100(2):246–252, June 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290232N.

Manolopoulos:1993:JLD

[MJP93] David E. Manolopoulos, Michael J. Jamieson, and Atul D. Pradhan. Johnson’s log derivative algorithm rederived. Journal of Computational Physics, 105(1):169–172, March 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710624.

Muradoglu:1999:CHF

[MJPC99] Metin Muradoglu, Patrick Jenny, Stephen B. Pope, and David A. Caughey. A consistent hybrid finite-volume/particle method for the PDF equations of turbulent reactive flows. Journal of Computational Physics, 154(2):342–371, September 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963167.

Matsumoto:1990:TTM

[MK90] Masami Matsumoto and Shigeo Kawata. TRIPIC: Triangularmesh particle-in-cell code. Journal of Computational Physics, 87(2):488–493, April 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919090262Y.

Mousseau:1997:FIK

[MK97] V. A. Mousseau and D. A. Knoll. Fully implicit kinetic solution of collisional plasmas. Journal of Computational Physics, 136(2):308–323, September 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957363.

Mansfield:1998:DSV

[MKM98] John R. Mansfield, Omar M. Knio, and Charles Meneveau. A dynamic LES scheme for the vorticity transport equation: Formulation and a priori tests. Journal of Computational Physics, 145(2):693–730, September 20, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919896051X.

Mansfield:1999:DCV

[MKM99] John R. Mansfield, Omar M. Knio, and Charles Meneveau. Dynamic LES of colliding vortex rings using a 3D vortex method. Journal of Computational Physics, 152(1):305–345, June 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199962587. See erratum [MKM04].

Mansfield:2004:CDC

[MKM04] John R. Mansfield, Omar M. Knio, and Charles Meneveau. Corrigendum to “Dynamic LES of colliding vortex rings using a 3D vortex method” [J. Comp. Phys. 152 (1999) 305–345]. Journal of Computational Physics, 197(2):779–780, July 1, 2004. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999103006569. See [MKM99].

Mahalov:1992:CCC

[ML92] Alex Mahalov and Sidney Leibovich. On the calculation of coupling coefficients in amplitude equations. Journal of Computational Physics, 101(2):441–444, August 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290019U.

Miller:1992:FNM

[MLB92] Timothy L. Miller, Huei-Iin Lu, and Karen A. Butler. A fully nonlinear, mixed spectral and finite difference model for thermally driven, rotating flows. Journal of Computational Physics, 101(2):265–275, August 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290004I.

Morris:1997:PTD

[MLBW97] Philip J. Morris, Lyle N. Long, Ashok Bangalore, and Qunzhen Wang. A parallel three-dimensional computational aeroacoustics method using nonlinear disturbance equations. Journal of Computational Physics, 133(1):56–74, May 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956461.

Manheimer:1997:LRC

[MLJ97] Wallace M. Manheimer, Martin Lampe, and Glenn Joyce. Langevin representation of Coulomb collisions in PIC simulations. Journal of Computational Physics, 138(2):563–584, December 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197958344.

Mei:1999:ACB

[MLS99] Renwei Mei, Li-Shi Luo, and Wei Shyy. An accurate curved boundary treatment in the lattice Boltzmann method. Journal of Computational Physics, 155(2):307–330, November 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963349.

Morinishi:1998:FCH

[MLVM98] Y. Morinishi, T. S. Lund, O. V. Vasilyev, and P. Moin. Fully conservative higher order finite difference schemes for incompressible flow. Journal of Computational Physics, 143(1):90–124, June 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198959629.

Marchand:1991:SID

[MM91a] R. Marchand and J. P. Matte. Simulation of ionisation dynamics and electron kinetics. Journal of Computational Physics, 97(2):352–365, December 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999191900045.

Matarrese:1991:NMS

[MM91b] M. D. Matarrese and A. F. Messiter. A numerical method for the self-similar hypersonic viscous shear layer. Journal of Computational Physics, 96(2):475–483, October 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190247I.

Masad:1995:DCB

[MM95a] J. A. Masad and M. R. Malik. Direct computation of boundary-layer stability characteristics. Journal of Computational Physics, 117(2):228–239, March 15, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710613.

Matsushima:1995:SMP

[MM95b] T. Matsushima and P. S. Marcus. A spectral method for polar coordinates. Journal of Computational Physics, 120(2):365–374, September 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999185711710.

Marti:1996:EPP

[MM96] José M. Martí and Ewald Müller. Extension of the piecewise parabolic method to one-dimensional relativistic hydrodynamics. Journal of Computational Physics, 123(1):1–14, January 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196900017.

Matsushima:1997:SMU

[MM97a] T. Matsushima and P. S. Marcus. A spectral method for unbounded domains. Journal of Computational Physics, 137 (2):321–345, November 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999197958046.

Morris:1997:SRS

[MM97b] J. P. Morris and J. J. Monaghan. A switch to reduce SPH viscosity. Journal of Computational Physics, 136(1):41–50, September 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197956904.

Mohring:1998:DVF

[MM98] Leonhard Möhring and Willi Möhring. On the determination of a velocity field with prescribed vorticity. Journal of Computational Physics, 147(1):229–235, November 20, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960855.

McFadden:1990:ESE

[MMB90] G. B. McFadden, B. T. Murray, and R. F. Boisvert. Elimination of spurious eigenvalues in the Chebyshev tau spectral method. Journal of Computational Physics, 91(1):228–239, November 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919090012P.

Marano:1998:RTP

[MML98] Stefano Marano, Mario Medugno, and Maurizio Longo. A real-time parallel application:: The detection of gravitational waves by a network of heterogeneous workstations. Journal of Computational Physics, 139(1):15–34, January 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958575.

Moriarty:1990:DIC

[MMR90] K. J. M. Moriarty, Eric Myers, and Claudio Rebbi. The dynamical interactions of cosmic strings. Journal of Computational Physics, 88(2):467–476, June 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190901898.

Migliore:1990:AFA

[MMS90a] M. Migliore, V. Martorana, and F. Sciortino. An algorithm to find all paths between two nodes in a graph. Journal of Computational Physics, 87(1):231–236, March 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090235S.

Mitchell:1990:NSH

[MMS90b] A. R. Mitchell, B. A. Murray, and B. D. Sleeman. Numerical solution of Hamiltonian systems in reaction diffusion by symplectic difference schemes. Journal of Computational Physics, 91(1): 253, November 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919090031U.

Mitchell:1991:NSH

[MMS91] A. R. Mitchell, B. A. Murray, and B. D. Sleeman. Numerical solution of Hamiltonian systems in reaction-diffusion by symplectic difference schemes. Journal of Computational Physics, 95(2):339– 358, August 1991. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919190280X.

Mostrom:1996:ACP

[MMW96] Michael A. Mostrom, Dushan Mitrovich, and Dale R. Welch. The ARCTIC charged particle beam propagation code. Journal of Computational Physics, 128(2):489–497, October 15, 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196902260.

Malkus:1990:DSF

[MNP90] David S. Malkus, John A. Nohel, and Bradley J. Plohr. Dynamics of shear flow of a non-Newtonian fluid. Journal of Computational Physics, 87(2):464–487, April 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090261X.

Morales:1990:SEM

[MNR90] Juan J. Morales, María J. Nuevo, and Luis F. Rull. Statistical error methods in computer simulations. Journal of Computational Physics, 89(2):432–438, August 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090151P.

McDonald:1995:PMP

[MO95] B. E. McDonald and Gregory J. Orbis. Performance of the matched-phase noise filter with estimated noise spectra. Journal of Computational Physics, 122(2):185–190, December 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185712077.

Monaghan:1994:SFS

[Mon94] J. J. Monaghan. Simulating free surface flows with SPH. Journal of Computational Physics, 110(2):399–406, February 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710345.

Monaghan:1997:ISD

[Mon97a] J. J. Monaghan. Implicit SPH drag and dusty gas dynamics. Journal of Computational Physics, 138(2):801–820, December 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197958460.

Monaghan:1997:SRS

[Mon97b] J. J. Monaghan. SPH and Riemann solvers. Journal of Computational Physics, 136(2):298–307, September 15, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957326.

Morgan:1990:DMP

[Mor90] Ronald B. Morgan. Davidson’s method and preconditioning for generalized eigenvalue problems. Journal of Computational Physics, 89(1):241–245, July 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919090124J.

Morgan:1992:GDM

[Mor92] Ronald B. Morgan. Generalizations of Davidson’s method for computing eigenvalues of large nonsymmetric matrices. Journal of Computational Physics, 101(2):287–291, August 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290006K.

Mulder:1992:CIM

[MOS92] W. Mulder, S. Osher, and James A. Sethian. Computing interface motion in compressible gas dynamics. Journal of Computational Physics, 100(2):209–228, June 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290229R.

Motin:1998:SIE

[Mot98] H. C. Motin. Solving integral equations by reconstruction in isomorphic Taylor coefficient spaces. Journal of Computational Physics, 143(2):291–311, July 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959873.

Mo:1993:BFC

[MP93] Yizhang Mo and Robert J. Perry. Basis function calculations for the massive Schwinger model in the light-front Tamm–Dancoff approximation. Journal of Computational Physics, 108(1):159– 174, September 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199918371171X.

Miller:1996:HOG

[MP96] Gregory Hale Miller and Elbridge Gerry Puckett. A high-order Godunov method for multiple condensed phases. Journal of Computational Physics, 128(1):134–164, October 1, 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196902004.

Moschetta:1997:RLD

[MP97] Jean-Marc Moschetta and D. I. Pullin. A robust low diffusive kinetic scheme for the Navier–Stokes/Euler equations. Journal of Computational Physics, 133(2):193–204, May 15, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956734.

Martin:1992:TPQa

[MPG92a] Pablo Martin, Ricardo Pérez, and Antonio L. Guerrero. Twopoint quasi-fractional approximations to the Airy function Ai(x). Journal of Computational Physics, 98(2):349, February 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290165U.

Martin:1992:TPQb

[MPG92b] Pablo Martín, Ricardo Pérez, and Antonio L. Guerrero. Twopoint quasi-fractional approximations to the Airy function Ai(x). Journal of Computational Physics, 99(2):337–340, April 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290212H.

Meneguzzi:1996:SMS

[MPPZ96] M. Meneguzzi, H. Politano, A. Pouquet, and M. Zolver. A sparsemode spectral method for the simulation of turbulent flows. Journal of Computational Physics, 123(1):32–44, January 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900030.

Maronnier:1999:NSF

[MPR99] V. Maronnier, M. Picasso, and J. Rappaz. Numerical simulation of free surface flows. Journal of Computational Physics, 155(2):439–455, November 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999199963465.

Micheletti:1995:CVC

[MQS95] S. Micheletti, A. Quarteroni, and R. Sacco. Current-voltage characteristics simulation of semiconductor devices using domain decomposition. Journal of Computational Physics, 119(1):46–61, June 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185711151.

Mulholland:1997:SEP

[MQS97] L. S. Mulholland, Y. Qiu, and D. M. Sloan. Solution of evolutionary partial differential equations using adaptive finite differences with pseudospectral post-processing. Journal of Computational Physics, 131(2):280–298, March 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196955928.

Myong:1998:GTS

[MR98] R. S. Myong and P. L. Roe. On Godunov-type schemes for magnetohydrodynamics: 1. A model system. Journal of Computational Physics, 147(2):545–567, December 10, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961010.

Mead:1999:ORK

[MR99] J. L. Mead and R. A. Renaut. Optimal Runge–Kutta methods for first order pseudospectral operators. Journal of Computational Physics, 152(1):404–419, June 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962605.

McGrattan:1994:FDF

[MRB94] K. B. McGrattan, R. G. Rehm, and H. R. Baum. Fire-driven flows in enclosures. Journal of Computational Physics, 110(2):285– 291, February 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184710254.

Marin:1993:EAM

[MRC93] Mauricio Marín, Dino Risso, and Patricio Cordero. Efficient algorithms for many-body hard particle molecular dynamics. Journal of Computational Physics, 109(2):306–317, December 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183712192.

Murray:1992:IAL

[MRD92] Christopher W. Murray, Stephen C. Racine, and Ernest R. Davidson. Improved algorithms for the lowest few eigenvalues and associated eigenvectors of large matrices. Journal of Computational Physics, 103(2):382–389, December 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290409R.

Morton:1994:UIM

[MRS94] K. W. Morton, M. A. Rudgyard, and G. J. Shaw. Upwind iteration methods for the cell vertex scheme in one dimension. Journal of Computational Physics, 114(2):209–226, October 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711612.

Morel:1998:LSO

[MRS98] J. E. Morel, Randy M. Roberts, and Mikhail J. Shashkov. A local support-operators diffusion discretization scheme for quadrilateral r–z meshes. Journal of Computational Physics, 144(1): 17–51, July 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198959812.

Maritz:1990:EAN

[MS90a] M. F. Maritz and S. W. Schoombie. Exact analysis of nonlinear instability in a discrete Burgers’ equation. Journal of Computational Physics, 91(2):496, December 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190900556.

Mulholland:1990:EFP

[MS90b] L. S. Mulholland and D. M. Sloan. The effect of filtering on the pseudospectral solution of evolutionary partial differential equations. Journal of Computational Physics, 91(2):495, December 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999190900523.

Maritz:1991:EAN

[MS91a] M. F. Maritz and S. W. Schoombie. Exact analysis of nonlinear instability in a discrete Burgers’ equation. Journal of Computational Physics, 97(1):73–90, November 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190039N.

Matsuda:1991:MCU

[MS91b] Y. Matsuda and Gary R. Smith. A microinstability code for a uniform magnetized plasma with an arbitrary distribution function. Journal of Computational Physics, 97(2):582, December 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190025G.

Mulholland:1991:EFP

[MS91c] L. S. Mulholland and D. M. Sloan. The effect of filtering on the pseudospectral solution of evolutionary partial differential equations. Journal of Computational Physics, 96(2):369–390, October 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190241C.

Matsuda:1992:MCUb

[MS92a] Y. Matsuda and Gary R. Smith. A microinstability code for a uniform magnetized plasma with an arbitrary distribution function. Journal of Computational Physics, 99(1):180, March 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290296B.

Matsuda:1992:MCUa

[MS92b] Y. Matsuda and Gary R. Smith. A microinstability code for a uniform magnetized plasma with an arbitrary distribution function. Journal of Computational Physics, 100(2):229–235, June 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290230V.

Merryfield:1993:PCT

[MS93] William J. Merryfield and Bernie Shizgal. Properties of collocation third-derivative operators. Journal of Computational Physics, 105(1):182–185, March 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918371065X.

Mei:1998:FDB

[MS98] Renwei Mei and Wei Shyy. On the finite difference-based lattice Boltzmann method in curvilinear coordinates. Journal of Computational Physics, 143(2):426–448, July 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959848.

Margolin:1999:UCG

[MS99a] Len Margolin and Mikhail Shashkov. Using a curvilinear grid to construct symmetry-preserving discretizations for Lagrangian gas dynamics. Journal of Computational Physics, 149(2):389– 417, March 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198961617.

Montarnal:1999:RGC

[MS99b] Philippe Montarnal and Chi-Wang Shu. Real gas computation using an energy relaxation method and high-order WENO schemes. Journal of Computational Physics, 148(1):59–80, January 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198961058.

Mazhukin:1994:SLP

[MSF94] V. Mazhukin, I. Smurov, and G. Flamant. Simulation of laser plasma dynamics: Influence of ambient pressure and intensity of laser radiation. Journal of Computational Physics, 112(1): 78–90, May 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184710825.

Manfredi:1995:NIV

[MSF^{+}95] G. Manfredi, M. Shoucri, M. R. Feix, P. Bertrand, E. Fijalkow, and A. Ghizzo. The numerical integration of the Vlasov equation possessing an invariant. Journal of Computational Physics, 121(2):298–313, October 15, 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999195901361.

Morey:1998:FAD

[MSV98] Chris Morey, John Scales, and Erik S. Van Vleck. A feedback algorithm for determining search parameters for Monte Carlo optimization. Journal of Computational Physics, 146(1):263–281, October 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199919896065X.

Mulder:1992:HREb

[Mul92a] Wim A. Mulder. A high-resolution Euler solver based on multigrid, semi-coarsening, and defect correction. Journal of Computational Physics, 99(1):181, March 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290303G.

Mulder:1992:HREa

[Mul92b] Wim A. Mulder. A high-resolution Euler solver based on multigrid, semi-coarsening, and defect correction. Journal of Computational Physics, 100(1):91–104, May 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290312M.

Murray:1996:EGO

[Mur96] Ross J. Murray. Explicit generation of orthogonal grids for ocean models. Journal of Computational Physics, 126(2):251–273, July 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196901369.

Manna:1999:EMS

[MV99] M. Manna and A. Vacca. An efficient method for the solution of the incompressible Navier–Stokes equations in cylindrical geometries. Journal of Computational Physics, 151(2):563–584, May 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199961971.

Mottura:1997:ERS

[MVZ97] Lorenzo Mottura, Luigi Vigevano, and Marco Zaccanti. An evaluation of Roe’s scheme generalizations for equilibrium real gas flows. Journal of Computational Physics, 138(2):354–399, December 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197958381.

Moss:1993:NPC

[MW93] William C. Moss and John W. White. New physical criteria for using linear artificial viscosity. Journal of Computational Physics, 106(2):397–399, June 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711174.

Miller:1994:EFF

[MW94] John J. H. Miller and Song Wang. An exponentially fitted finite volume method for the numerical solution of 2D unsteady incompressible flow problems. Journal of Computational Physics, 115(1):56–64, November 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711788.

Manson:1997:ACT

[MW97] J. Russell Manson and Steve G. Wallis. Accuracy characteristics of traditional finite volume discretizations for unsteady computational fluid dynamics. Journal of Computational Physics, 132(1):149–153, March 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196956235.

Ming:1991:CES

[MWJ91] H. Ming, K. Werner, and H. Jackisch. Construction of the entropy solution of hyperbolic conservation laws by a geometrical interpretation of the conservation principle. Journal of Computational Physics, 95(1):40–58, July 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190252G.

Mansell:1994:LLD

[MWM94] G. Mansell, J. Walter, and E. Marschall. Liquid-liquid driven cavity flow. Journal of Computational Physics, 110(2):274–284, February 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184710242.

Morel:1996:LDS

[MWS96] J. E. Morel, Todd A. Wareing, and Kenneth Smith. A lineardiscontinuous spatial differencing scheme for S_{n} radiative transfer calculations. Journal of Computational Physics, 128(2):445–462, October 15, 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196902235.

McKee:1996:ADI

[MWW96] S. McKee, D. P. Wall, and S. K. Wilson. An alternating direction implicit scheme for parabolic equations with mixed derivative and convective terms. Journal of Computational Physics, 126(1): 64–76, June 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196901205.

Nachman:1996:BPC

[Nac96] Arje Nachman. A brief perspective on computational electromagnetics. Journal of Computational Physics, 126(1):237–239, June 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196901321.

Nadiga:1995:ADV

[Nad95] B. T. Nadiga. An adaptive discrete-velocity model for the shallow water equations. Journal of Computational Physics, 121(2):271–280, October 15, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999195901027.

Nakagawa:1995:WSF

[Nak95] Norio Nakagawa. A weakly singular formulation of boundary integral equations and field computation near boundaries. Journal of Computational Physics, 122(1):143–149, November 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918571203X.

Natarajan:1990:FEA

[Nat90] Ramesh Natarajan. Finite-element applications on a sharedmemory multiproccessor — algorithms and experimental results. Journal of Computational Physics, 90(1):270, September 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919090218P.

Natarajan:1991:FEA

[Nat91] Ramesh Natarajan. Finite element applications on a sharedmemory multiprocessor: Algorithms and experimental results. Journal of Computational Physics, 94(2):352–381, June 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190225A.

Natarajan:1992:ABI

[Nat92a] Ramesh Natarajan. An arnoldi-based iterative scheme for nonsymmetric matrix pencils arising in finite element stability problems. Journal of Computational Physics, 100(1):128–142, May 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290315P.

Natarajan:1992:NMI

[Nat92b] Ramesh Natarajan. A numerical method for incompressible viscous flow simulation. Journal of Computational Physics, 100(2): 384–395, June 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290244S.

Nie:1998:AAQ

[NB98] Qing Nie and Greg Baker. Application of adaptive quadrature to axi-symmetric vortex sheet motion. Journal of Computational Physics, 143(1):49–69, June 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959721.

Nguyen:1993:CLS

[NC93] Hoa D. Nguyen and Jacob N. Chung. A Chebyshev–Legendre spectral method for the transient solutions of flow past a solid sphere. Journal of Computational Physics, 104(2):303–312, February 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999183710326.

Natoli:1995:OMT

[NC95] Vincent Natoli and David M. Ceperley. An optimized method for treating long-range potentials. Journal of Computational Physics, 117(1):171–178, March 1, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710546.

Nash:1997:EFD

[NC97] Patrick L. Nash and L. Y. Chen. Efficient finite difference solutions to the time-dependent Schrödinger equation. Journal of Computational Physics, 130(2):266–268, January 15, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196955898.

Nordstrom:1999:BIC

[NC99] Jan Nordström and Mark H. Carpenter. Boundary and interface conditions for high-order finite-difference methods applied to the Euler and Navier–Stokes equations. Journal of Computational Physics, 148(2):621–645, January 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961332.

Natoli:1996:NNA

[NCF96] Vincent D. Natoli, Morrel H. Cohen, and Bengt Fornberg. A new numerical algorithm for the analytic continuation of Green’s functions. Journal of Computational Physics, 126(1):99–108, June 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196901230.

Nielsen:1990:AOP

[ND90] Dale E. Nielsen, Jr. and Adam T. Drobot. An analysis and optimization of the pseudo-current method. Journal of Computational Physics, 89(1):31–40, July 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090115H.

Neal:1994:IAP

[Nea94] Radford M. Neal. An improved acceptance procedure for the hybrid Monte Carlo algorithm. Journal of Computational Physics, 111(1):194–203, March 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710540.

Neta:1997:ATZ

[NGN97] B. Neta, F. X. Giraldo, and I. M. Navon. Analysis of the Turkel–Zwas scheme for the two-dimensional shallow water equations in spherical coordinates. Journal of Computational Physics, 133(1):102–112, May 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956576.

Nicolai:1993:MCC

[Nic93] A. Nicolai. A Monte Carlo code describing the neutral gas transport in pipe configurations with attenuating media. Journal of Computational Physics, 106(2):377–390, June 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711150.

Nicholls:1998:TWW

[Nic98] David P. Nicholls. Traveling water waves: Spectral continuation methods with parallel implementation. Journal of Computational Physics, 143(1):224–240, June 10, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959575.

Nicoud:1999:DWA

[Nic99] F. Nicoud. Defining wave amplitude in characteristic boundary conditions. Journal of Computational Physics, 149(2):418–422, March 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198961319.

Jiang:1993:NON

[nJ93] Bo nan Jiang. Non-oscillatory and non-diffusive solution of convection problems by the iteratively reweighted least-squares finite element method. Journal of Computational Physics, 105(1):108– 121, March 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999183710570.

Jiang:1996:OSS

[nJWP96] Bo nan Jiang, Jie Wu, and L. A. Povinelli. The origin of spurious solutions in computational electromagnetics. Journal of Computational Physics, 125(1):104–123, April 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900820.

Natarajan:1995:CSQ

[NM95] A. Natarajan and N. Mohankumar. A comparison of some quadrature methods for approximating Cauchy principal value integrals. Journal of Computational Physics, 116(2):365–368, February 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999185710340.

Negron-Marrero:1994:NCS

[NMB94] Pablo V. Negrón-Marrero and Octavio Betancourt. The numerical computation of singular minimizers in two-dimensional elasticity. Journal of Computational Physics, 113(2):291–303, August 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999184711363.

Nakamura:1996:CCS

[NMW^{+}96] Y. Nakamura, T. Matsumoto, M. Wakatani, S. A. Galkin, V. V. Drozdov, A. A. Martynov, Yu. Yu. Poshekhonov, K. Ichiguchi, L. Garcia, B. A. Carreras, C. Nührenberg (née Schwab), W. A. Cooper, and J. L. Johnson. Comparison of the calculations of the stability properties of a specific stellarator equilibrium with different MHD stability codes. Journal of Computational Physics, 128(1):43–57, October 1, 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901953.

Nayar:1993:CSE

[NO93] Narinder Nayar and James M. Ortega. Computation of selected eigenvalues of generalized eigenvalue problems. Journal of Computational Physics, 108(1):8–14, September 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711587.

Noelle:1994:HSC

[Noe94] Sebastian Noelle. Hyperbolic systems of conservation laws, the Weyl equation, and multidimensional upwinding. Journal of Computational Physics, 115(1):22–26, November 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711752.

Nogueira:1993:AFG

[Nog93] P. Nogueira. Automatic Feynman graph generation. Journal of Computational Physics, 105(2):279–289, April 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710740.

Nordmark:1991:RHO

[Nor91a] Henrik O. Nordmark. Rezoning for higher order vortex methods. Journal of Computational Physics, 97(2):366–397, December 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999191900056.

Normark:1991:RHO

[Nor91b] Henrik O. Normark. Rezoning for higher order vortex methods. Journal of Computational Physics, 93(1):250, March 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190081U.

Nordstrom:1995:ASN

[Nor95] Jan Nordström. Accurate solutions of the Navier–Stokes equations despite unknown outflow boundary data. Journal of Computational Physics, 120(2):184–205, September 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711588.

Nordmark:1996:DHO

[Nor96] Henrik O. Nordmark. Deterministic high order vortex methods for the 2D Navier–Stokes equation with rezoning. Journal of Computational Physics, 129(1):41–56, November 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196902326.

Nordmark:1998:EHA

[Nor98] Henrik O. Nordmark. Efficient and highly accurate computation of a class of radially symmetric solutions of the Navier– Stokes equation and the heat equation in two dimensions. Journal of Computational Physics, 142(2):428–444, May 20, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959332.

Natarajan:1992:PCG

[NP92] Ramesh Natarajan and Pratap Pattnaik. Performance of the conjugate gradient method on Victor. Journal of Computational Physics, 100(2):396–401, June 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290245T.

Nadiga:1994:MNE

[NP94] B. T. Nadiga and D. I. Pullin. A method for nearequilibrium discrete-velocity gas flows. Journal of Computational Physics, 112(1):162–172, May 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710898.

Nguyen:1993:AVI

[NPC93] Hoa D. Nguyen, Seungho Paik, and Jacob N. Chung. Application of vorticity integral conditioning to Chebyshev pseudospectral formulation for the Navier–Stokes equations. Journal of Computational Physics, 106(1):115–124, May 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918371096X.

Nochetto:1996:DMA

[NPV96] R. H. Nochetto, M. Paolini, and C. Verdi. A dynamic mesh algorithm for curvature dependent evolving interfaces. Journal of Computational Physics, 123(2):296–310, February 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919690025X.

Nelson:1992:CSDa

[NSR92a] P. Nelson, D. L. Seth, and A. K. Ray. A computational study of the discretization error in the solution of the Spencer–Lewis equation by doubling applied to the upwind finite-difference approximation. Journal of Computational Physics, 101(2):452, August 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290025T.

Nelson:1992:CSDb

[NSR92b] P. Nelson, D. L. Seth, and A. K. Ray. A computational study of the discretization error in the solution of the Spencer–Lewis equation by doubling applied to the upwind finite-difference approximation. Journal of Computational Physics, 103(2):370–381, December 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290408Q.

Nessyahu:1990:NOC

[NT90] Haim Nessyahu and Eitan Tadmor. Non-oscillatory central differencing for hyperbolic conservation laws. Journal of Computational Physics, 87(2):408–463, April 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190902608.

Nunn:1993:NTN

[Nun93] D. Nunn. A novel technique for the numerical simulation of hot collision-free plasma: Vlasov hybrid simulation. Journal of Computational Physics, 108(1):180–196, September 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711733.

Najm:1998:SIN

[NWK98] Habib N. Najm, Peter S. Wyckoff, and Omar M. Knio. A semi-implicit numerical scheme for reacting flow: I. Stiff chemistry. Journal of Computational Physics, 143(2):381–402, July 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197958563.

Nanbu:1998:WPC

[NY98] K. Nanbu and S. Yonemura. Weighted particles in Coulomb collision simulations based on the theory of a cumulative scattering angle. Journal of Computational Physics, 145(2):639–654, September 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198960491.

Oh:1995:MAM

[OB95] Hae-Soo Oh and Ivo Babuška. The method of auxiliary mapping for the finite element solutions of elasticity problems containing singularities. Journal of Computational Physics, 121(2):193–212, October 15, 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999195900179.

Oden:1998:DHF

[OBB98] J. Tinsley Oden, Ivo Babuška, and Carlos Erik Baumann. A discontinuous hp finite element method for diffusion problems. Journal of Computational Physics, 146(2):491–519, November 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198960326.

ORourke:1993:PGI

[OBL93] Peter J. O’Rourke, J. U. Brackbill, and Bernard Larrouturou. On particle-grid Interpolation and calculating chemistry in particlein-cell methods. Journal of Computational Physics, 109(1):37– 52, November 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999183711976.

Otto:1999:NSG

[OD99] S. R. Otto and James P. Denier. Numerical solution of a generalized elliptic partial differential eigenvalue problem. Journal of Computational Physics, 156(2):352–359, December 10, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963647.

Odstrcil:1990:NOFa

[Ods90a] Dusan Odstrcil. A new optimized FCT algorithm for shock wave problems. Journal of Computational Physics, 87(2):495, April 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999190902675.

Odstrcil:1990:NOFb

[Ods90b] Dusan Odstrcil. A new optimized FCT algorithm for shock wave problems. Journal of Computational Physics, 91(1):71–93, November 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919090005L.

Odstrcil:1993:IFA

[Ods93] Dus Odstrcil. Improved FCT algorithm for shock hydrodynamics. Journal of Computational Physics, 108(2):218–225, October 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183711770.

Ollivier-Gooch:1997:QES

[OG97] Carl F. Ollivier-Gooch. Quasi-ENO schemes for unstructured meshes based on unlimited data-dependent least-squares reconstruction. Journal of Computational Physics, 133(1):6–17, May 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196955849.

Oreg:1990:CAG

[OGBSK90] J. Oreg, W. H. Goldstein, A. Bar-Shalom, and M. Klapisch. Configuration average of general n-body symmetrical tensor operators. Journal of Computational Physics, 91(2):460–477, December 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999190900486.

Olson:1990:EFE

[OGS90] Lorraine G. Olson, Georgios C. Georgiou, and William W. Schultz. An efficient finite element method for treating singularities in Laplace’s equation. Journal of Computational Physics, 91(2):497, December 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919090060E.

Olson:1991:EFE

[OGS91] Lorraine G. Olson, Georgios C. Georgiou, and William W. Schultz. An efficient finite element method for treating singularities in Laplace’s equation. Journal of Computational Physics, 96(2):391–410, October 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190242D.

Oosterlee:1998:MLS

[OGWW98] C. W. Oosterlee, F. J. Gaspar, T. Washio, and R. Wienands. Multigrid line smoothers for higher order upwind discretizations of convection-dominated problems. Journal of Computational Physics, 139(2):274–307, January 20, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919795854X.

Ohwada:1998:HOA

[Ohw98] Taku Ohwada. Higher order approximation methods for the Boltzmann equation. Journal of Computational Physics, 139(1):1–14, January 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999197958691.

Jeon:1995:FDM

[oJ95] Dong o Jeon. Finite difference method for calculating magnetic field components off-median plane using median plane field data. Journal of Computational Physics, 117(1):55–66, March 1, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185710443.

Jeon:1997:CFD

[oJ97] Dong o Jeon. Compact finite difference method for calculating magnetic field components of cyclotrons. Journal of Computational Physics, 132(2):167–174, April 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196956351.

Oh:1994:NIS

[OK94] H. J. Oh and I. S. Kang. A non-iterative scheme for orthogonal grid generation with control function and specified boundary correspondence on three sides. Journal of Computational Physics, 112(1):138–148, May 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710874.

Okunbor:1995:ECL

[Oku95] Daniel I. Okunbor. Energy conserving, Liouville, and symplectic integrators. Journal of Computational Physics, 120(2):375–378, September 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999185711722.

Ozyoruk:1996:NEA

[ÖL96] Yusuf Özyörük and Lyle N. Long. A new efficient algorithm for computational aeroacoustics on parallel processors. Journal of Computational Physics, 125(1):135–149, April 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900844.

Olim:1994:TNS

[Oli94] M. Olim. A truly noninterpolating semi-Lagrangian Lax– Wendroff method. Journal of Computational Physics, 112(2): 253–266, June 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184710977.

Ozyoruk:1998:TDN

[ÖLJ98] Yusuf Özyörük, Lyle N. Long, and Michael G. Jones. Timedomain numerical simulation of a flow-impedance tube. Journal of Computational Physics, 146(1):29–57, October 10, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959198.

Olmez:1995:NMN

[ÖM95] H. S. Ölmez and J. H. Milgram. Numerical methods for nonlinear interactions between water waves. Journal of Computational Physics, 118(1):62–72, April 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710790.

Oosterlee:1997:GBP

[Oos97] C. W. Oosterlee. A GMRES-based plane smoother in multigrid to solve 3D anisotropic fluid flow problems. Journal of Computational Physics, 130(1):41–53, January 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919695442X.

Ostrovskii:1997:NMO

[OP97] A. G. Ostrovskii and L. I. Piterbarg. A new method for obtaining velocity and diffusivity from time-dependent distributions of a tracer via the maximum likelihood estimator for the advection-diffusion equation. Journal of Computational Physics, 133(2):340–360, May 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956746.

Or:1994:SSB

[Or94] A. C. Or. A shooting scheme for boundary-value problems. Journal of Computational Physics, 114(2):280–283, October 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999184711661.

OLeary:1990:CEE

[OS90] D. P. O’Leary and G. W. Stewart. Computing the eigenvalues and eigenvectors of symmetric arrowhead matrices. Journal of Computational Physics, 90(2):497–505, October 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190901773.

Omelchenko:1997:DDE

[OS97] Yu. A. Omelchenko and R. N. Sudan. A 3-D darwin-EM hybrid PIC code for ion ring studies. Journal of Computational Physics, 133(1):146–159, May 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956709.

ORourke:1998:VEI

[OS98] P. J. O’Rourke and M. S. Sahota. A variable explicit/implicit numerical method for calculating advection on unstructured meshes. Journal of Computational Physics, 143(2):312–345, July 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198959034.

Osborne:1990:NFA

[Osb90] A. R. Osborne. Nonlinear Fourier analysis for the infinite-interval Korteweg–de Vries equation I: An algorithm for the direct scattering transform. Journal of Computational Physics, 89(2):489, August 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919090159X.

Osborne:1991:NFA

[Osb91] A. R. Osborne. Nonlinear Fourier analysis for the infinite-interval Korteweg–de Vries equation I: An algorithm for the direct scattering transform. Journal of Computational Physics, 94(2):284– 313, June 1991. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999191902238.

Osborne:1993:HIS

[Osb93] A. R. Osborne. The hyperelliptic inverse scattering transform for the periodic, defocusing nonlinear Schrödinger equation. Journal of Computational Physics, 109(1):93–107, November 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183712027.

Oscarsson:1994:DPM

[Osc94] Tord Oscarsson. Dual principles in maximum entropy reconstruction of the wave distribution function. Journal of Computational Physics, 110(2):221–233, February 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710205.

Otani:1992:EVS

[OsKB^{+}92] Niels F. Otani, Jin soo Kim, Charles K. Birdsall, Bruce T. Cohen, William Mccay Nevins, and Neil Maron. Elimination of velocity-space rings-and-spokes instabilities in magnetized electrostatic particle simulations of plasmas. Journal of Computational Physics, 102(2):227–235, October 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192903689.

Perrochet:1995:STI

[PA95a] Pierre Perrochet and Pascal Azérad. Space–time integrated leastsquares: Solving a pure advection equation with a pure diffusion operator. Journal of Computational Physics, 117(2):183–193, March 15, 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999185710571.

Poliashenko:1995:DMC

[PA95b] Maxim Poliashenko and Cyrus K. Aidun. A direct method for computation of simple bifurcations. Journal of Computational Physics, 121(2):246–260, October 15, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999195900683.

Puckett:1997:HOP

[PAB^{+}97] Elbridge Gerry Puckett, Ann S. Almgren, John B. Bell, Daniel L. Marcus, and William J. Rider. A high-order projection method for tracking fluid interfaces in variable density incompressible flows. Journal of Computational Physics, 130(2):269–282, January 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196955904.

Paisley:1997:MCS

[Pai97] M. F. Paisley. Multigrid computation of stratified flow over two-dimensional obstacles. Journal of Computational Physics, 136(2):411–424, September 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957648.

Palmer:1993:DAS

[Pal93] Bruce J. Palmer. Direct application of shake to the velocity Verlet algorithm. Journal of Computational Physics, 104(2):470– 472, February 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999183710454.

Papatzacos:1992:NCE

[Pap92] Paul Papatzacos. Numerical calculation of the equation of flow in porous media: The lattice-gas approach. Journal of Computational Physics, 102(2):424, October 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290385C.

Papatzacos:1993:NCE

[Pap93] Paul Papatzacos. Numerical calculation of the equation of flow in porous media: The lattice gas approach. Journal of Computational Physics, 104(2):313–320, February 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710338.

Paruolo:1990:VEM

[Par90] Giuseppe Paruolo. A vector-efficient and memory-saving interpolation algorithm for PIC codes on a Cray X-MP. Journal of Computational Physics, 89(2):462–482, August 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090153R.

Parker:1991:NEE

[PB91] S. E. Parker and C. K. Birdsall. Numerical error in electron orbits with large ω_{ce}Δt. Journal of Computational Physics, 97(1):91– 102, November 1991. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919190040R.

Paisley:1998:CMM

[PB98] M. F. Paisley and N. M. Bhatti. Comparison of multigrid methods for neutral and stably stratified flows over two-dimensional obstacles. Journal of Computational Physics, 142(2):581–610, May 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198959150.

Patzek:1991:NOI

[PBBS91] T. W. Patzek, R. E. Benner, Jr., O. A. Basaran, and L. E. Scriven. Nonlinear oscillations of inviscid free drops. Journal of Computational Physics, 97(2):489–515, December 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190010I.

Patzek:1995:NOT

[PBBS95] T. W. Patzek, O. A. Basaran, R. E. Benner, and L. E. Scriven. Nonlinear oscillations of two-dimensional, rotating inviscid drops. Journal of Computational Physics, 116(1):3–25, January 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185710029.

Pember:1995:ACG

[PBC^{+}95] Richard B. Pember, John B. Bell, Phillip Colella, William Y. Curtchfield, and Michael L. Welcome. An adaptive Cartesian grid method for unsteady compressible flow in irregular regions. Journal of Computational Physics, 120(2):278–304, September 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185711655.

Pinelli:1994:CPS

[PBD94a] A. Pinelli, C. Benocci, and M. Deville. Chebyshev pseudospectral solution of advection-diffusion equations with mapped finite difference preconditioning. Journal of Computational Physics, 112 (1):1–11, May 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184710771.

Pletzer:1994:LSR

[PBD94b] A. Pletzer, A. Bondeson, and R. L. Dewar. Linear stability of resistive MHD modes: Axisymmetric toroidal computation of the outer region matching data. Journal of Computational Physics, 115(2):530–549, December 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184712150.

Procassini:1992:CPCb

[PC92a] Richard J. Procassini and Bruce I. Cohen. A comparison of particle-in-cell and Fokker–Planck methods as applied to the modeling of auxiliary-heated mirror plasmas. Journal of Computational Physics, 102(1):39–48, September 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999105800038.

Procassini:1992:CPCa

[PC92b] Richard J. Procassini and Bruce I. Cohen. A comparison of particle-in-cell and Fokker–Planck methods as applied to the modelling of auxilliary-heated mirror plasmas. Journal of Computational Physics, 100(2):433, June 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290252T.

Palaniswamy:1996:SES

[PC96] Sampath Palaniswamy and Sukumar R. Chakravarthy. Stability of explicit schemes in the physical and frequency domains. Journal of Computational Physics, 124(1):162–176, March 1, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196900510.

Prosser:1998:UWC

[PC98] R. Prosser and R. S. Cant. On the use of wavelets in computational combustion. Journal of Computational Physics, 147 (2):337–361, December 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999198960922.

Pareschi:1999:IMC

[PC99a] Lorenzo Pareschi and Russel E. Caflisch. An implicit Monte Carlo method for rarefied gas dynamics: I. The space homogeneous case. Journal of Computational Physics, 154(1):90–116, September 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999199963015.

Popken:1999:CKT

[PC99b] Lars Popken and Paul W. Cleary. Comparison of kinetic theory and discrete element schemes for modelling granular Couette flows. Journal of Computational Physics, 155(1):1–25, October 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199962927.

Pfrommer:1997:RCQ

[PCLC97] Bernd G. Pfrommer, Michel Côté, Steven G. Louie, and Marvin L. Cohen. Relaxation of crystals with the quasi-Newton method. Journal of Computational Physics, 131(1):233–240, February 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196956120.

Prasad:1991:IPI

[PDC91] Surya Prasad, S.-K Dinavahi, and S.-K Chow. Inverse problem in incompressible, irrotational axisymmetric flow. Journal of Computational Physics, 94(2):419–436, June 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190228D.

Pendergast:1994:SAT

[PDHS94] Phil Pendergast, Zareh Darakjian, Edward F. Hayes, and Danny C. Sorensen. Scalable algorithms for three-dimensional reactive scattering: Evaluation of a new algorithm for obtaining surface functions. Journal of Computational Physics, 113(2):201– 214, August 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184711302.

Pettitt:1993:SBT

[PDL93] Brian A. Pettitt, Werner Danchura, and Donna Labun. Spherical Bessel transforms. Journal of Computational Physics, 105(1):178–181, March 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999183710648.

Pfrommer:1999:UEF

[PDS99] Bernd G. Pfrommer, James Demmel, and Horst Simon. Unconstrained energy functionals for electronic structure calculations.

Journal of Computational Physics, 150(1):287–298, March 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198961812.

Pelz:1991:PFP

[Pel91] Richard B. Pelz. The parallel Fourier pseudospectral method. Journal of Computational Physics, 92(2):296–312, February 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999191902124.

Peng:1995:HON

[Pen95] Dianyun Peng. High-order numerical method for two-point boundary value problems. Journal of Computational Physics, 120(2):253–259, September 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199918571162X.

Perot:1993:AFS

[Per93] J. Blair Perot. An analysis of the fractional step method. Journal of Computational Physics, 108(1):51–58, September 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711629.

Perot:1995:CFS

[Per95] J. Blair Perot. Comments on the fractional step method. Journal of Computational Physics, 121(1):190–191, October 1, 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711898.

Petersen:1994:SEN

[Pet94] W. P. Petersen. Some experiments on numerical simulations of stochastic differential equations and a new algorithm. Journal of Computational Physics, 113(1):75–81, July 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711193.

Petersson:1996:FNC

[Pet96] N. Anders Petersson. Fast numerical computation of 2D free surface jet flow with surface tension. Journal of Computational Physics, 128(2):409–426, October 15, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919690220X.

Petropoulos:1998:TPM

[Pet98] Peter G. Petropoulos. On the termination of the perfectly matched layer with local absorbing boundary conditions. Journal of Computational Physics, 143(2):665–673, July 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959794.

Parker:1993:BMS

[PFRB93] S. E. Parker, A. Friedman, S. L. Ray, and C. K. Birdsall. Bounded multi-scale plasma simulation: Application to sheath problems. Journal of Computational Physics, 107(2):388–402, August 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999183711538.

Peterkin:1998:TMF

[PFS98] Robert E. Peterkin, Jr., Michael H. Frese, and Carl R. Sovinec. Transport of magnetic flux in an arbitrary coordinate ALE code. Journal of Computational Physics, 140(1):148–171, February 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198958806.

Pierce:1997:PMM

[PG97] Niles A. Pierce and Michael B. Giles. Preconditioned multigrid methods for compressible flow calculations on stretched meshes. Journal of Computational Physics, 136(2):425–445, September 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197957727.

Provatas:1999:AMR

[PGD99] Nikolas Provatas, Nigel Goldenfeld, and Jonathan Dantzig. Adaptive mesh refinement computation of solidification microstructures using dynamic data structures. Journal of Computational Physics, 148(1):265–290, January 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961228.

Parlett:1994:MMR

[PH94] Beresford N. Parlett and Wee-Liang Heng. The method of minimal representations in 2D Ising model calculations. Journal of Computational Physics, 114(2):257–264, October 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711648.

Peters:1995:VSM

[PH95a] M. C. A. M. Peters and H. W. M. Hoeijemakers. A vortex sheet method applied to unsteady flow separation from sharp edges. Journal of Computational Physics, 120(1):88–104, August 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185711503.

Pickering:1995:RBC

[PH95b] W. M. Pickering and P. J. Harley. On Robbins boundary conditions, elliptic equations, and FFT methods. Journal of Computational Physics, 122(2):380–383, December 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185712235.

Parker:1993:ASB

[PHL93] G. J. Parker, W. N. G. Hitchon, and J. E. Lawler. Accelerated solution of the Boltzmann equation. Journal of Computational Physics, 106(1):147–154, May 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710995.

Pang:1994:STS

[PJ94] Alex C.-Y. Pang and Chueng-Ryong Ji. A spinor technique in symbolic Feynman diagram calculation. Journal of Computational Physics, 115(2):267–275, December 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711946.

P:1994:FVS

[PK94] Vijayan P. and Y. Kallinderis. A 3D finite-volume scheme for the Euler equations on adaptive tetrahedral grids. Journal of Computational Physics, 113(2):249–267, August 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711338.

Pathria:1995:SEM

[PK95] D. Pathria and G. E. Karniadakis. Spectral element methods for elliptic problems in nonsmooth domains. Journal of Computational Physics, 122(1):83–95, November 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711989.

Pan:1992:FSN

[PL92a] Dartzi Pan and S. C. Lee. A fixed-stencil non-oscillatory scheme for hyperbolic systems. Journal of Computational Physics, 100(1):200–204, May 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919290321O.

Poinsot:1992:BCDa

[PL92b] T. J. Poinsot and S. K. Lelef. Boundary conditions for direct simulations of compressible viscous flows. Journal of Computational Physics, 101(1):104–129, July 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192900462.

Petschek:1993:CSP

[PL93] A. G. Petschek and L. D. Libersky. Cylindrical smoothed particle hydrodynamics. Journal of Computational Physics, 109(1):76– 83, November 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999183712003.

Polstyanko:1998:TLH

[PL98] Sergey V. Polstyanko and Jin-Fa Lee. Two-level hierarchical FEM method for modeling passive microwave devices. Journal of Computational Physics, 140(2):400–420, March 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198958880.

Papalexandris:1997:USH

[PLD97] Miltiadis V. Papalexandris, Anthony Leonard, and Paul E. Dimotakis. Unsplit schemes for hyperbolic conservation laws with source terms in one space dimension. Journal of Computational Physics, 134(1):31–61, June 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956928.

Plimpton:1995:FPA

[Pli95] Steve Plimpton. Fast parallel algorithms for short-range molecular dynamics. Journal of Computational Physics, 117(1):1–19, March 1, 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199918571039X.

Pathria:1990:PSS

[PM90] D. Pathria and J. Ll. Morris. Pseudo-spectral solution of nonlinear Schrödinger equations. Journal of Computational Physics, 87(1):108–125, March 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919090228S.

Petersson:1992:CFA

[PM92] N. Anders Petersson and Johan F. Malmliden. Computing the flow around a submergedj body using composite grids. Journal of Computational Physics, 103(1):187, November 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290336W.

Petersson:1993:CFA

[PM93] N. Anders Petersson and Johan F. Malmliden. Computing the flow around a submerged body using composite grids. Journal of Computational Physics, 105(1):47–57, March 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710521.

Priymak:1998:ANS

[PM98] V. G. Priymak and T. Miyazaki. Accurate Navier–Stokes investigation of transitional and turbulent flows in a circular pipe. Journal of Computational Physics, 142(2):370–411, May 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198959319.

Place:1999:LSM

[PM99] David Place and Peter Mora. The lattice solid model to simulate the physics of rocks and earthquakes: Incorporation of friction. Journal of Computational Physics, 150(2):332–372, April 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199961843.

Peng:1999:PBF

[PMO^{+}99] Danping Peng, Barry Merriman, Stanley Osher, Hongkai Zhao, and Myungjoo Kang. A PDE-based fast local level set method. Journal of Computational Physics, 155(2):410–438, November 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199963453.

Pokharna:1997:RTP

[PMR97] Himanshu Pokharna, Michitsugu Mori, and Victor H. Ransom. Regularization of two-phase flow models: a comparison of numerical and differential approaches. Journal of Computational Physics, 134(2):282–295, July 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956953.

Paik:1994:TCH

[PNC94] Seungho Paik, Hoa D. Nguyen, and Jacob N. Chung. Transient conjugated heat transfer analysis from a sphere with nonlinear boundary conditions. Journal of Computational Physics, 115(2):423–430, December 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999184712071.

Provenzale:1990:NFA

[PO90] A. Provenzale and A. R. Osborne. Nonlinear Fourier analysis for the infinite-interval Korteweg–de Vries equation II: Numerical tests of the direct scattering transform. Journal of Computational Physics, 89(2):489, August 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190901603.

Provenzale:1991:NFA

[PO91] A. Provenzale and A. R. Osborne. Nonlinear Fourier analysis for the infinite-interval Korteweg–de Vries equation II: Numerical tests of the direct scattering transform. Journal of Computational Physics, 94(2):314–351, June 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999191902249.

Pointon:1991:SCB

[Poi91] T. D. Pointon. Slanted conducting boundaries and field emission of particles in an electromagnetic particle simulation code. Journal of Computational Physics, 96(1):143–162, September 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190269Q.

Poinsot:1992:BCDb

[Poi92] T. J. Poinsot. Boundary conditions for direct simulations of compressible viscous flows. Journal of Computational Physics, 99(2): 352, April 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290227P.

Pope:1995:PMT

[Pop95] S. B. Pope. Particle method for turbulent flows: Integration of stochastic model equations. Journal of Computational Physics, 117(2):332–349, March 15, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710704.

Poyd:1992:ATK

[Poy92] John Poyd. The arc tangent and Kepler–Burger mappings for periodic solutions with a shock, front, or internal boundary layer. Journal of Computational Physics, 98(1):178, January 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290183Y.

Pozrikidis:1999:SEM

[Poz99] C. Pozrikidis. A spectral-element method for particulate Stokes flow. Journal of Computational Physics, 156(2):360–381, December 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999199963751.

Peskin:1993:IVC

[PP93] Charles S. Peskin and Beth Feller Printz. Improved volume conservation in the computation of flows with immersed elastic boundaries. Journal of Computational Physics, 105(1):33– 46, March 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199918371051X.

Parker:1992:SBC

[PPBC92] S. E. Parker, R. J. Procassini, C. K. Birdsall, and B. I. Cohen. A suitable boundary condition for bounded plasma simulation without sheath resolution. Journal of Computational Physics, 101(2): 453, August 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290031S.

Parker:1993:SBC

[PPBC93] S. E. Parker, R. J. Procassini, C. K. Birdsall, and B. I. Cohen. A suitable boundary condition for bounded plasma simulation without sheath resolution. Journal of Computational Physics, 104(1): 41–49, January 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999183710053.

Peraire:1992:ARTa

[PPM92a] J. Peraire, J. Peiró, and K. Morgan. Adaptive remeshing for three-dimensional compressible flow computations. Journal of Computational Physics, 101(1):229, July 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290068A.

Peraire:1992:ARTb

[PPM92b] J. Peraire, J. Peiró, and K. Morgan. Adaptive remeshing for three-dimensional compressible flow computations. Journal of Computational Physics, 103(2):269–285, December 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290401J.

Perthame:1994:VVL

[PQ94] Benoît Perthame and Youchun Qiu. A variant of van Leer’s method for multidimensional systems of conservation laws. Journal of Computational Physics, 112(2):370–381, June 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711077.

Phillips:1993:TSP

[PR93a] Timothy N. Phillips and Gareth W. Roberts. The treatment of spurious pressure modes in spectral incompressible flow calculations. Journal of Computational Physics, 105(1):150–164, March 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183710600.

Purnell:1993:EBF

[PR93b] Don K. Purnell and Michael J. Revell. Energy-bounded flow approximation on a Cartesian-product grid over rough terrain. Journal of Computational Physics, 107(1):51–65, July 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711241.

Prisco:1990:ODS

[Pri90] G. Prisco. Optimization of direct simulation Monte Carlo (DSMC) codes for vector processing. Journal of Computational Physics, 90(1):269, September 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090213K.

Prisco:1991:ODS

[Pri91] G. Prisco. Optimization of direct simulation Monte Carlo (DSMC) codes for vector processing. Journal of Computational Physics, 94(2):454–466, June 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190230I.

Priestley:1993:QRM

[Pri93] A. Priestley. A quasi-Riemannian method for the solution of one-dimensional shallow water flow. Journal of Computational Physics, 106(1):139–146, May 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710983.

Priestley:1994:EPL

[Pri94] A. Priestley. Exact projections and the Lagrange–Galerkin method: a realistic alternative to quadrature. Journal of Computational Physics, 112(2):316–333, June 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711041.

Priymak:1995:PAN

[Pri95] V. G. Priymak. Pseudospectral algorithms for Navier–Stokes simulation of turbulent flows in cylindrical geometry with coordinate singularities. Journal of Computational Physics, 118(2):366–379, May 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185711060.

Prigozhin:1996:BMS

[Pri96] Leonid Prigozhin. The bean model in superconductivity: Variational formulation and numerical solution. Journal of Computational Physics, 129(1):190–200, November 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196902430.

Prigozhin:1998:STF

[Pri98] Leonid Prigozhin. Solution of thin film magnetization problems in type-II superconductivity. Journal of Computational Physics, 144(1):180–193, July 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959782.

Powell:1999:SAU

[PRL^{+}99] Kenneth G. Powell, Philip L. Roe, Timur J. Linde, Tamas I. Gombosi, and Darren L. De Zeeuw. A solution-adaptive upwind scheme for ideal magnetohydrodynamics. Journal of Computational Physics, 154(2):284–309, September 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919996299X.

Pinero:1990:BSE

[PRMV90] Petra Pinero, Luis F. Rull, Juan J. Morales, and JoséM Velarde. Baxter solution to the O–Z equation near the critical point. Journal of Computational Physics, 88(2):490–494, June 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190901924.

Papageorgiou:1990:ASE

[PRS90] C. D. Papageorgiou, A. D. Raptis, and T. E. Simos. An algorithm for the solution of the eigenvalue Schrödinger equation. Journal of Computational Physics, 88(2):477–483, June 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090190C.

Pryce:1994:ERC

[Pry94] John D. Pryce. Efficient, reliable computation of resonances of the one-dimensional Schrödinger equation. Journal of Computational Physics, 112(2):234–246, June 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710953.

Pereira:1993:FVC

[PS93a] J. C. F. Pereira and J. M. M. Sousa. Finite volume calculations of self-sustained oscillations in a grooved channel. Journal of Computational Physics, 106(1):19–29, May 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710879.

Poedts:1993:CIM

[PS93b] Stefaan Poedts and Elisabeth Schwartz. Computation of the ideal-MHD continuous spectrum in axisymmetric plasmas. Journal of Computational Physics, 105(1):165–168, March 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710612.

Pasik:1999:MPM

[PSL99] Michael F. Pasik, David B. Seidel, and Raymond W. Lemke. A modified perfectly matched layer implementation for use in electromagnetic PIC codes. Journal of Computational Physics, 148(1):125–132, January 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999198961071.

Pelekasis:1992:HFB

[PTM92] N. A. Pelekasis, J. A. Tsamopoulos, and G. D. Manolis. A hybrid finite-boundary element method for inviscid flows with free surface. Journal of Computational Physics, 101(2):231–251, August 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290001F.

Puckett:1997:INM

[Puc97] Gerry Puckett. Introduction to “A Numerical Method for Solving Incompressible Viscous Flow Problems”. Journal of Computational Physics, 135(2):115–117, August 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957697.

Pulliam:1993:TCO

[PV93] Thomas H. Pulliam and John A. Vastano. Transition to chaos in an open unforced 2D flow. Journal of Computational Physics, 105(1):133–149, March 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710594.

Peyrard:1999:RSI

[PV99] P. F. Peyrard and P. Villedieu. A Roe scheme for ideal MHD equations on 2D adaptively refined triangular grids. Journal of Computational Physics, 150(2):373–393, April 10, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199961983.

Pinelli:1997:SMM

[PVQ97] A. Pinelli, A. Vacca, and A. Quarteroni. A spectral multidomain method for the numerical simulation of turbulent flows. Journal of Computational Physics, 136(2):546–558, September 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197957806.

Pavlo:1998:LSA

[PVVS98] Pavol Pavlo, George Vahala, Linda Vahala, and Min Soe. Linear stability analysis of thermo-lattice Boltzmann models. Journal of Computational Physics, 139(1):79–91, January 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958642.

Prendergast:1993:NHG

[PX93] Kevin H. Prendergast and Kun Xu. Numerical hydrodynamics from gas-kinetic theory. Journal of Computational Physics, 109(1):53–66, November 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711988.

Piette:1998:NID

[PZ98] B. Piette and W. J. Zakrzewski. Numerical integration of (2 + 1) dimensional PDEs for S^{2} valued functions. Journal of Computational Physics, 145(1):359–381, September 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960132.

Petropoulos:1998:RSL

[PZC98] Peter G. Petropoulos, Li Zhao, and Andreas C. Cangellaris. A reflectionless sponge layer absorbing boundary condition for the solution of Maxwell’s equations with high-order staggered finite difference schemes. Journal of Computational Physics, 139(1):184– 208, January 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999197958551.

Qiu:1999:ICA

[QF99] Yanghui Qiu and Charlotte Froese Fischer. Integration by cell algorithm for Slater integrals in a spline basis. Journal of Computational Physics, 156(2):257–271, December 10, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963611.

Qi:1998:EPM

[QG98] Quan Qi and Thomas L. Geers. Evaluation of the perfectly matched layer for computational acoustics. Journal of Computational Physics, 139(1):166–183, January 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919795868X.

Qi:1993:NSW

[Qi93] Anmin Qi. Numerical study of wave propagation on vortex filaments. Journal of Computational Physics, 104(1):185–210, January 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002199918371020X.

Quartapelle:1990:NST

[QR90] L. Quartapelle and S. Rebay. Numerical solution of twopoint boundary value problems. Journal of Computational Physics, 86(2):314–354, February 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999190901049.

Quon:1990:NAS

[QS90] Charles Quon and Helmuth Sandstrom. A numerical algorithm to study internal solitary waves. Journal of Computational Physics, 86(1):168–186, January 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090096J.

Qiu:1998:NSF

[QS98] Y. Qiu and D. M. Sloan. Numerical solution of Fisher’s equation using a moving mesh method. Journal of Computational Physics, 146(2):726–746, November 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960818.

Qian:1998:FTM

[QTL98] J. Qian, G. Tryggvason, and C. K. Law. A front tracking method for the motion of premixed flames. Journal of Computational Physics, 144(1):52–69, July 20, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959915.

Qian:1993:WNS

[QW93] Sam Qian and John Weiss. Wavelets and the numerical solution of partial differential equations. Journal of Computational Physics, 106(1):155–175, May 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711009.

Ramaswamy:1990:NSUa

[Ram90a] Balasubramaniam Ramaswamy. Numerical simulation of unsteady viscous free-surface flow. Journal of Computational Physics, 87(2):495, April 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999190902664.

Ramaswamy:1990:NSUb

[Ram90b]  Balasubramaniam Ramaswamy.  Numerical simulation of

unsteady viscous free surface flow. Journal of Computational Physics, 90(2):396–430, October 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090173X.

Ramachandran:1992:NSMa

[Ram92a] P. A. Ramachandran. A numerical solution method for boundary value problems containing an undetermined parameter. Journal of Computational Physics, 100(2):433, June 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290254V.

Ramachandran:1992:NSMb

[Ram92b] P. A. Ramachandran. A numerical solution method for boundary value problems containing an undetermined parameter. Journal of Computational Physics, 102(1):63–71, September 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999105800051.

Rambo:1995:FGI

[Ram95] P. W. Rambo. Finite-grid instability in quasineutral hybrid simulations. Journal of Computational Physics, 118(1):152–158, April 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185710868.

Ramahi:1997:CBO

[Ram97a] Omar M. Ramahi. Complementary boundary operators for wave propagation problems. Journal of Computational Physics, 133(1):113–128, May 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999197956588.

Rambo:1997:NHH

[Ram97b] P. W. Rambo. Numerical heating in hybrid plasma simulations. Journal of Computational Physics, 133(1):173–180, May 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197956783.

Rapaport:1993:NAB

[Rap93] D. C. Rapaport. A note on algorithms for billiard-ball dynamics. Journal of Computational Physics, 105(2):367–368, April 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002199918371082X.

Ravichandran:1997:HOK

[Rav97] K. S. Ravichandran. Higher order KFVS algorithms using compact upwind difference operators. Journal of Computational Physics, 130(2):161–173, January 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196955618.

Rawitscher:1990:AAB

[Raw90] George H. Rawitscher. Accuracy analysis of a Bessel spectral on method for the solution of scattering equations. Journal of Computational Physics, 90(1):267, September 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090206G.

Rawitscher:1991:AAB

[Raw91] George H. Rawitscher. Accuracy analysis of a Bessel spectral function method for the solution of scattering equations. Journal of Computational Physics, 94(1):81–101, May 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190138B.

Restrepo:1995:DMF

[RB95] Juan Mario Restrepo and Jerry L. Bona. Discretization of a model for the formation of longshore sand ridges. Journal of Computational Physics, 122(1):129–142, November 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185712028.

Ramshaw:1995:ISI

[RC95] J. D. Ramshaw and C. H. Chang. Iteration scheme for implicit calculations of kinetic and equilibrium chemical reactions in fluid dynamics. Journal of Computational Physics, 116(2):359–364, February 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999185710339.

Ramakrishna:1992:NMT

[RCA92] K. Ramakrishna, I. M. Cohen, and P. S. Ayyaswamy. Numerical methods for two-dimensional analysis of electrical breakdown in a non-uniform gap. Journal of Computational Physics, 101(2): 454, August 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290034V.

Ramakrishna:1993:NMT

[RCA93] K. Ramakrishna, I. M. Cohen, and P. S. Ayyaswamy. Numerical methods for two-dimensional analysis of electrical breakdown in a non-uniform gap. Journal of Computational Physics, 104(1):173– 184, January 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999183710193.

Reid:1992:DSPa

[RCAM92a] Donald Reid, Andrew Chan, and Mustafa Al-Mudares. A direct solution of Poisson’s equation in a three-dimensional field-effect transistor structure. Journal of Computational Physics, 98(2): 349, February 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290161Q.

Reid:1992:DSPb

[RCAM92b] Donald Reid, Andrew Chan, and Mustafa Al-Mudares. A direct solution of Poisson’s equation in a three-dimensional field-effect transistor structure. Journal of Computational Physics, 99(1): 79–83, March 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999192902765.

Rambo:1991:FFA

[RD91] P. W. Rambo and J. Denavit. Fluid and field algorithms for time-implicit plasma simulation. Journal of Computational Physics, 92(1):185–212, January 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190297X.

Rambo:1992:TIFa

[RD92a] P. W. Rambo and J. Denavit. Time-implicit fluid simulation of collisional plasmas. Journal of Computational Physics, 98(1): 180, January 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999192901944.

Rambo:1992:TIFb

[RD92b] P. W. Rambo and J. Denavit. Time-implicit fluid simulation of collisional plasmas. Journal of Computational Physics, 98(2):317–331, February 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919290145O.

Read:1997:IET

[Rea97] F. H. Read. Improved extrapolation technique in the boundary element method to find the capacitances of the unit square and cube. Journal of Computational Physics, 133(1):1–5, May 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196955199.

Rebay:1993:EUM

[Reb93] S. Rebay. Efficient unstructured mesh generation by means of Delaunay triangulation and Bowyer–Watson algorithm. Journal of Computational Physics, 106(1):125–138, May 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710971.

Reiter:1992:ADMa

[Rei92a] Johann Reiter. An algorithm for deconvolution by the maximum entropy method with astronomical applications. Journal of Computational Physics, 101(1):228, July 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192900624.

Reiter:1992:ADMb

[Rei92b] Johann Reiter. An algorithm for deconvolution by the maximum entropy method with astronomical applications. Journal of Computational Physics, 103(1):169–183, November 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290332S.

Reich:1999:MTS

[Rei99]  Sebastian Reich.  Multiple time scales in classical and

quantum-classical molecular dynamics. Journal of Computational Physics, 151(1):49–73, May 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961423.

Renaut:1992:ABCa

[Ren92a] R. A. Renaut. Absorbing boundary conditions, difference operators, and stability. Journal of Computational Physics, 100(2):

434, June 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290257Y.

Renaut:1992:ABCb

[Ren92b] R. A. Renaut. Absorbing boundary conditions, difference operators, and stability. Journal of Computational Physics, 102(2):236–251, October 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919290369A.

Ratowsky:1990:TAD

[RF90] R. P. Ratowsky and J. A. Fleck, Jr. Treatment of angular derivatives in the Schrödinger equation by the finite Fourier series method. Journal of Computational Physics, 89(2):490, August 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919090163U.

Ratowsky:1991:TAD

[RF91] R. P. Ratowsky and J. A. Fleck, Jr. Treatment of angular derivatives in the Schrödinger equation by the finite Fourier series method. Journal of Computational Physics, 93(2):376–387, April 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190190V.

Rokicki:1995:DDC

[RF95] J. Rokicki and J. M. Floryan. Domain decomposition and the compact fourth-order algorithm for the Navier–Stokes equations. Journal of Computational Physics, 116(1):79–96, January 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185710078.

Renaut:1996:PCM

[RF96] Rosemary Renaut and Jochen Fröhlich. A pseudospectral Chebychev method for the 2D wave equation with domain stretching and absorbing boundary conditions. Journal of Computational Physics, 124(2):324–336, March 15, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900637.

Robinson:1993:NSC

[RFH93] M. P. Robinson, G. Fairweather, and B. M. Herbst. On the numerical solution of the cubic Schrödinger equation in one space variable. Journal of Computational Physics, 104(1):277–284, January 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183710296.

Richardson:1993:USE

[RFL93] John L. Richardson, Robert C. Ferrell, and Lyle N. Long. Unconditionally stable explicit algorithms for nonlinear fluid dynamics problems. Journal of Computational Physics, 104(1):69–74, January 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183710089.

Reiman:1990:CZT

[RG90] A. H. Reiman and H. S. Greenside. Computation of zero β three-dimensional equilibria with magnetic islands. Journal of Computational Physics, 87(2):349–365, April 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190902572.

Renaud:1997:DPS

[RG97] François Renaud and Serge Gauthier. A dynamical pseudospectral domain decomposition technique: Application to viscous compressible flows. Journal of Computational Physics, 131(1):89– 108, February 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199919695576X.

Ramsden:1991:TLP

[RH91] Dave Ramsden and Greg Holloway. Timestepping Lagrangian particles in two dimensional Eulerian flow fields. Journal of Computational Physics, 95(1):101–116, July 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190255J.

Rombouts:1998:AEA

[RH98] S. Rombouts and K. Heyde. An accurate and efficient algorithm for the computation of the characteristic polynomial of a general square matrix. Journal of Computational Physics, 140(2):453– 458, March 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198959095.

Rivoalen:1999:NSA

[RH99] E. Rivoalen and S. Huberson. Numerical simulation of axisymmetric viscous flows by means of a particle method. Journal of Computational Physics, 152(1):1–31, June 10, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962101.

Rovelstad:1994:EIE

[RHB94] Amy L. Rovelstad, Robert A. Handler, and Peter S. Bernard. The effect of Interpolation errors on the Lagrangian analysis of simulated turbulent channel flow. Journal of Computational Physics, 110(1):190–195, January 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710151.

Roy:1996:TDP

[RHT96] Robie I. Samanta Roy, Daniel E. Hastings, and Steven Taylor. Three-dimensional plasma particle-in-cell calculations of ion thruster backflow contamination. Journal of Computational Physics, 128(1):6–18, October 1, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901928.

Riedel:1994:OED

[Rie94a] Kurt S. Riedel. Optimal estimation of dynamically evolving diffusivities. Journal of Computational Physics, 115(1):1–11, November 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999184711739.

Rief:1994:SMC

[Rie94b] Herbert Rief. A synopsis of Monte Carlo perturbation algorithms. Journal of Computational Physics, 111(1):33–48, March 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999184710412.

Rieffel:1999:MEC

[Rie99a] Marc A. Rieffel. A method for estimating the computational requirements of DSMC simulations. Journal of Computational Physics, 149(1):95–113, February 10, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919896140X.

Rietdijk:1999:NBR

[Rie99b] R. H. Rietdijk. Notes on big ray tracing. Journal of Computational Physics, 148(1):149–168, January 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961137.

Rigal:1994:HOD

[Rig94] Alain Rigal. High order difference schemes for unsteady onedimensional diffusion-convection problems. Journal of Computational Physics, 114(1):59–76, September 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711491.

Ronchi:1996:CSN

[RIP96] C. Ronchi, R. Iacono, and P. S. Paolucci. The “cubed sphere”: a new method for the solution of partial differential equations in spherical geometry. Journal of Computational Physics, 124(1):93–114, March 1, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900479.

Rosenfeld:1990:MST

[RIW90] Moshe Rosenfeld, Moshe Israeli, and Micha Wolfshtfin. A method for solving three-dimensional viscous incompressible flows over slender bodies. Journal of Computational Physics, 88(2):255– 283, June 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999190901809.

Raad:1993:CSC

[RK93] Peter E. Raad and Andreas Karageorghis. A Chebyshev spectral collocation method for the solution of the Reynolds equation of lubrication. Journal of Computational Physics, 106(1):

42–51, May 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999183710892.

Radespiel:1995:AFV

[RK95] R. Radespiel and N. Kroll. Accurate flux vector splitting for shocks and shear layers. Journal of Computational Physics, 121(1):66–78, October 1, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711795.

Rider:1998:RVT

[RK98] William J. Rider and Douglas B. Kothe. Reconstructing volume tracking. Journal of Computational Physics, 141(2):112– 152, April 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199919895906X.

Rider:1999:TSS

[RK99] William J. Rider and Dana A. Knoll. Time step size selection for radiation diffusion calculations. Journal of Computational Physics, 152(2):790–795, July 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962666.

Ramaswamy:1997:TDS

[RKJ97] B. Ramaswamy, S. Krishnamoorthy, and S. W. Joo. Threedimensional simulation of instabilities and rivulet formation in heated falling films. Journal of Computational Physics, 131(1):70–88, February 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999196955126.

Rider:1999:MNK

[RKO99] William J. Rider, Dana A. Knoll, and Gordon L. Olson. A multigrid Newton–Krylov method for multimaterial equilibrium radiation diffusion. Journal of Computational Physics, 152(1):164– 191, June 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199919996240X.

Rosenfeld:1990:FSS

[RKV90] Moshe Rosenfeld, Dochan Kwak, and Marxel Vinokur. A fractional step solution method for the unsteady incompressible Navier–Stokes equations in generalized coordinate systems. Journal of Computational Physics, 90(1):267, September 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090208I.

Rosenfeld:1991:FSS

[RKV91] Moshe Rosenfeld, Dochan Kwak, and Marcel Vinokur. A fractional step solution method for the unsteady incompressible Navier–Stokes equations in generalized coordinate systems. Journal of Computational Physics, 94(1):102–137, May 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190139C.

Restrepo:1995:WGD

[RL95] Juan Mario Restrepo and Gary K. Leaf. Wavelet-Galerkin discretization of hyperbolic equations. Journal of Computational Physics, 122(1):118–128, November 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185712016.

Rennich:1997:NMI

[RL97] Steven C. Rennich and Sanjiva K. Lele. Numerical method for incompressible vortical flows with two unbounded directions. Journal of Computational Physics, 137(1):101–129, October 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002199919795796X.

Roth:1990:IMS

[RLM90] M. Roth, J. Lemaire, and A. Misson. An iterative method to solve the nonlinear Poisson’s equation in the case of plasma tangential discontinuities. Journal of Computational Physics, 86(2):466– 486, February 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919090109E.

Reglat:1993:NFS

[RLT93] O. Réglat, R. Labrie, and P. A. Tanguy. A new free surface model for the dip coating process. Journal of Computational Physics, 109(2):238–246, December 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183712155.

Rai:1990:DST

[RM90] Man Mohan Rai and Parvis Moin. Direct simulations of turbulent flow using finite-difference schemes. Journal of Computational Physics, 91(1):251, November 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090024U.

Rai:1991:DST

[RM91] Man Mohan Rai and Parviz Moin. Direct simulations of turbulent flow using finite-difference schemes. Journal of Computational Physics, 96(1):15–53, September 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190264L.

Rai:1993:DNS

[RM93] Man Mohan Rai and Parviz Moin. Direct numerical simulation of transition and turbulence in a spatially evolving boundary layer. Journal of Computational Physics, 109(2):169–192, December 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183712106.

Ruuth:1999:CGM

[RMO99] Steven J. Ruuth, Barry Merriman, and Stanley Osher. Convolutiongenerated motion as a link between cellular automata and continuum pattern dynamics. Journal of Computational Physics, 151(2):836–861, May 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199919996223X.

Ross:1996:CSO

[RMSB96] Hughan J. Ross, Luke F. McAven, Kiminari Shinagawa, and Philip H. Butler. Calculating spin-orbit matrix elements with RACAH. Journal of Computational Physics, 128(2):331–340, October 15, 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196902144.

Rajagopal:1994:OEM

[RN94]  G. Rajagopal and R. J. Needs.  An optimized Ewald

method for long-ranged potentials. Journal of Computational Physics, 115(2):399–405, December 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184712058.

Robson:1991:CDO

[RNSV91] R. E. Robson, K. F. Ness, G. E. Sneddon, and Larry A. Viehland. Comment on the discrete ordinate method in the kinetic theory of gases. Journal of Computational Physics, 92(1):213–229, January 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190298Y.

Roberts:1990:BFD

[Rob90a] Thomas W. Roberts. The behavior of flux difference splitting schemes near slowly moving shock waves. Journal of Computational Physics, 90(1):141–160, September 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090200K.

Robinson:1990:SMC

[Rob90b] P. A. Robinson. Systematic methods for calculation of the dielectric properties of arbitrary plasmas. Journal of Computational Physics, 88(2):381–392, June 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190901854.

Roberts:1999:TDD

[Rob99] Thomas M. Roberts. Time-domain deconvolution removes the effects of near-field scatterers. Journal of Computational Physics, 149(2):293–305, March 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919896154X.

Roe:1997:ARS

[Roe97a] P. L. Roe. Approximate Riemann solvers, parameter vectors, and difference schemes. Journal of Computational Physics, 135(2):250–258, August 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999197957053.

Roe:1997:P

[Roe97b] Phil Roe. Preface. Journal of Computational Physics, 135(2): 99, August 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999197957594.

Rokhlin:1990:RSI

[Rok90] Vladimir Rokhlin. Rapid solution of integral equations of scattering theory in two dimensions. Journal of Computational Physics, 86(2):414–439, February 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090107C.

Romate:1992:ABCa

[Rom92a] J. E. Romate. Absorbing boundary conditions for free surface waves. Journal of Computational Physics, 98(2):347–348, February 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290153P.

Romate:1992:ABCb

[Rom92b] J. E. Romate. Absorbing boundary conditions for free surface waves. Journal of Computational Physics, 99(1):135–145, March 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999192902813.

Rom:1997:QDS

[Rom97] A. Y. Rom. Quasiparticles density of states and the scattering rate in ideal type II superconductor by a time-dependent approach. Journal of Computational Physics, 135(1):76–82, July 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197957375.

Rostand:1990:KBL

[Ros90] Philippe Rostand. Kinetic boundary layers, numerical simulations. Journal of Computational Physics, 86(1):18–55, January 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919090090N.

Rowlands:1991:NAH

[Row91] G. Rowlands. A numerical algorithm for Hamiltonian systems. Journal of Computational Physics, 97(1):235–239, November 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190046N.

Reiman:1990:CMC

[RP90] H. Reiman and N. Pomphrey. Computation of magnetic coordinates and action-angle variables. Journal of Computational Physics, 90(1):270, September 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090217O.

Reiman:1991:CMC

[RP91] A. H. Reiman and N. Pomphrey. Computation of magnetic coordinates and action-angle variables. Journal of Computational Physics, 94(1):225–249, May 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190144A.

Roma:1999:AVI

[RPB99] Alexandre M. Roma, Charles S. Peskin, and Marsha J. Berger. An adaptive version of the immersed boundary method. Journal of Computational Physics, 153(2):509–534, August 10, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962939.

Renardy:1990:NBC

[RR90] Michael Renardy and Yuriko Renardy. On the nature of boundary conditions for flows with moving free surfaces. Journal of Computational Physics, 89(1):255, July 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090142N.

Renardy:1991:NBC

[RR91] Michael Renardy and Yuriko Renardy. On the nature of boundary conditions for flows with moving free surfaces. Journal of Computational Physics, 93(2):325–335, April 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190186O.

Roberts:1990:TPF

[RS90] Lawrence J. Roberts and Kenneth S. Sorbie. Two-phase flow in porous media with multicomponent transport: Formulation and higher order numerical solution. Journal of Computational Physics, 86(2):440–465, February 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090108D.

Ramona:1991:ASI

[RS91] D. R. Ramona and P. Sulmont. On the accuracy of the spectral interpolation methods for sampled pseudoperiodic signals. Journal of Computational Physics, 97(1):30–52, November 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190037L.

Rajamaki:1994:AOD

[RS94a] Markku Rajamäki and Mika Saarinen. Accurate one-dimensional computation of frontal phenomena by PLIM. Journal of Computational Physics, 111(1):62–73, March 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710448.

Russo:1994:FTV

[RS94b] Giovanni Russo and John A. Strain. Fast triangulated vortex methods for the 2D Euler equations. Journal of Computational Physics, 111(2):291–323, April 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710655.

Radespiel:1995:PMS

[RS95] R. Radespiel and R. C. Swanson. Progress with multigrid schemes for hypersonic flow problems. Journal of Computational Physics, 116(1):103–122, January 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710091.

Robertson:1999:FSF

[RS99a] Iain Robertson and Spencer Sherwin. Free-surface flow simulation using hp/spectral elements. Journal of Computational Physics, 155(1):26–53, October 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963283.

Romero:1999:GSL

[RS99b] Aldo H. Romero and Jose M. Sancho. Generation of short and long range temporal correlated noises. Journal of Computational Physics, 156(1):1–11, November 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963477.

Rjasanow:1998:RNP

[RSW98] Sergej Rjasanow, Thomas Schreiber, and Wolfgang Wagner. Reduction of the number of particles in the stochastic weighted particle method for the Boltzmann equation. Journal of Computational Physics, 145(1):382–405, September 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960181.

Risebro:1992:FTM

[RT92] N. H. Risebro and A. Tveito. A front tracking method for conservation laws in one dimension. Journal of Computational Physics, 101(1):130–139, July 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192900473.

Roosen:1994:MCG

[RT94] Andrew R. Roosen and Jean E. Taylor. Modeling crystal growth in a diffusion field using fully faceted interfaces. Journal of Computational Physics, 114(1):113–128, September 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711533.

Russo:1990:PMC

[Rus90] G. Russo. A particle method for collisional kinetic equations. I. Basic theory and one-dimensional results. Journal of Computational Physics, 87(2):270–300, April 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090254X.

Russo:1993:DVM

[Rus93] Giovanni Russo. A deterministic vortex method for the Navier– Stokes equations. Journal of Computational Physics, 108(1):84– 94, September 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999183711654.

Ruuth:1998:DGA

[Ruu98a] Steven J. Ruuth. A diffusion-generated approach to multiphase motion. Journal of Computational Physics, 145(1):166–192, September 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198960284.

Ruuth:1998:EAD

[Ruu98b] Steven J. Ruuth. Efficient algorithms for diffusion-generated motion by mean curvature. Journal of Computational Physics, 144(2):603–625, August 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999198960259.

Ruyten:1993:DCS

[Ruy93] Wilhelmus M. Ruyten. Density-conserving shape factors for particle simulations in cylindrical and spherical coordinates. Journal of Computational Physics, 105(2):224–232, April 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710703.

Rumsey:1993:MFF

[RvR93] Christopher L. Rumsey, Bram vanLeer, and Philip L. Roe. A multidimensional flux function with applications to the Euler and Navier–Stokes equations. Journal of Computational Physics, 105(2):306–323, April 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999183710776.

Rjasanow:1996:NSS

[RW96a] Sergej Rjasanow and Wolfgang Wagner. Numerical study of a stochastic weighted particle method for a model kinetic equation. Journal of Computational Physics, 128(2):351–362, October 15, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196902168.

Rjasanow:1996:SWP

[RW96b] Sergej Rjasanow and Wolfgang Wagner. A stochastic weighted particle method for the Boltzmann equation. Journal of Computational Physics, 124(2):243–253, March 15, 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900571.

Rothberg:1990:EBP

[RWC90] R. H. Rothberg, W. F. Walker, and A. J. Chapman. Enhanced boundary pressure update for incompressible flow simulation. Journal of Computational Physics, 88(2):495–498, June 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999190901935.

Reese:1995:SOD

[RWTC95] J. M. Reese, L. C. Woods, F. J. P. Thivet, and S. M. Candel. A second-order description of shock structure. Journal of Computational Physics, 117(2):240–250, March 15, 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710625.

Ribbens:1994:SVF

[RWW94] Calvin J. Ribbens, Layne T. Watson, and C.-Y. Wang. Steady viscous flow in a triangular cavity. Journal of Computational Physics, 112(1):173–181, May 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710904.

Rosenfeld:1994:ADM

[RY94] Moshe Rosenfeld and Yuval Yassour. The alternating direction multi-zone implicit method. Journal of Computational Physics, 110(2):212–220, February 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710199.

Sotiropoulos:1990:CFI

[SA90a] F. Sotiropoulos and S. Abdallah. Coupled fully implicit solution procedure for the steady incompressible Navier–Stokes equations. Journal of Computational Physics, 87(2):328–348, April 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919090256Z.

Sotiropoulos:1990:DCE

[SA90b] F. Sotiropoulos and S. Abdallah. The discrete continuity equation in primitive variable solutions of incompressible flow. Journal of Computational Physics, 91(1):249, November 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090018V.

Sotiropoulos:1991:DCE

[SA91] F. Sotiropoulos and S. Abdallah. The discrete continuity equation in primitive variable solutions of incompressible flow. Journal of Computational Physics, 95(1):212–227, July 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190260R.

Sotiropoulos:1992:PVMa

[SA92a] F. Sotiropoulos and S. Abdallah. A primitive variable method for the solution of three-dimensional incompressible viscous flows. Journal of Computational Physics, 101(1):229, July 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192900727.

Sotiropoulos:1992:PVMb

[SA92b] F. Sotiropoulos and S. Abdallah. A primitive variable method for the solution of three-dimensional incompressible viscous flows. Journal of Computational Physics, 103(2):336–349, December 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290405N.

Scott:1995:FDF

[SA95] James R. Scott and Hafiz M. Atassi. A finite-difference, frequency-domain numerical scheme for the solution of the gust response problem. Journal of Computational Physics, 119(1): 75–93, June 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999185711175.

Saurel:1999:MGM

[SA99] Richard Saurel and Rémi Abgrall. A multiphase Godunov method for compressible multifluid and multiphase flows. Journal of Computational Physics, 150(2):425–467, April 10, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199961879.

Sonnendrucker:1995:FEF

[SAB95] Eric Sonnendrücker, John J. Ambrosiano, and Scott T. Brandon. A finite element formulation of the Darwin PIC model for use on unstructured grids. Journal of Computational Physics, 121(2):281–297, October 15, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999195901191.

Sussman:1999:ALS

[SAB^{+}99] Mark Sussman, Ann S. Almgren, John B. Bell, Phillip Colella, Louis H. Howell, and Michael L. Welcome. An adaptive level set approach for incompressible two-phase flows. Journal of Computational Physics, 148(1):81–124, January 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919896106X.

Sainsaulieu:1995:FVA

[Sai95] Lionel Sainsaulieu. Finite volume approximation of two phasefluid flows based on an approximate Roe-type Riemann solver. Journal of Computational Physics, 121(1):1–28, October 1, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002199918571176X.

Sakai:1996:NFV

[Sak96] Katsuhiro Sakai. A new finite variable difference method with application to locally exact numerical scheme. Journal of Computational Physics, 124(2):301–308, March 15, 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900613.

SaldanhadaGama:1991:SSS

[Sal91] Rogerlo Martins Saldanha da Gama. Simulation of the steadystate energy transfer in rigid bodies, with convective/radiative boundary conditions, employing a minimum principle. Journal of Computational Physics, 97(2):583, December 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190027I.

Saltzman:1994:UUM

[Sal94] Jeff Saltzman. An unsplit 3D upwind method for hyperbolic conservation laws. Journal of Computational Physics, 115(1):153– 168, November 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184711843.

Samtaney:1997:CMS

[Sam97] Ravi Samtaney. Computational methods for self-similar solutions of the compressible Euler equations. Journal of Computational Physics, 132(2):327–345, April 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196956399.

Samareh-Abolhassani:1994:SGG

[SAS94] Jamshid Samareh-Abolhassani and John E. Stewart. Surface grid generation in a parameter space. Journal of Computational Physics, 113(1):112–121, July 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711211.

Sulsky:1991:NMSb

[SB91] Deborah Sulsky and J. U. Brackbill. A numerical method for suspension flow. Journal of Computational Physics, 96(2):339– 368, October 1991. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919190240L.

Schafer:1995:NST

[SB95] R. Schäfer and R. Blendowske. On the numerical solution of the time-dependent Schrödinger equation. Journal of Computational Physics, 119(1):206–208, June 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918571128X.

Saiki:1996:NSC

[SB96a] E. M. Saiki and S. Biringen. Numerical simulation of a cylinder in uniform flow: Application of a virtual boundary method. Journal of Computational Physics, 123(2):450–465, February 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900364.

Stevens:1996:FTA

[SB96b] David E. Stevens and Christopher S. Bretherton. A forwardin-time advection scheme and adaptive multilevel flow solver for nearly incompressible atmospheric flow. Journal of Computational Physics, 129(2):284–295, December 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0021999196902508. See erratum [Ano97-29].

Strickland:1998:MCG

[SB98] James H. Strickland and Roy S. Baty. Modification of the Carrier, Greengard, and Rokhlin FMM for independent source and target fields. Journal of Computational Physics, 142(1):123–128, May 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198959071.

Simons:1999:LGA

[SBC99] Neil R. S. Simons, Greg E. Bridges, and Michel Cuhaci. A lattice gas automaton capable of modeling three-dimensional electromagnetic fields. Journal of Computational Physics, 151(2):816– 835, May 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999199962216.

Schneider:1999:EFV

[SBGK99] T. Schneider, N. Botta, K. J. Geratz, and R. Klein. Extension of finite volume compressible flow solvers to multi-dimensional, variable density zero Mach number flows. Journal of Computational Physics, 155(2):248–286, November 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963271.

Sethian:1992:TDV

[SBGM92] J. A. Sethian, Jean-Philippe Brunet, Adam Greenberg, and Jill P. Mesirov. Two-dimensional, viscous, incompressible flow in complex geometries on a massively parallel processor. Journal of Computational Physics, 101(1):185–206, July 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290052Z.

Sharafeddin:1992:NES

[SBKH92] Omar A. Sharafeddin, H. Ferrel Bowen, Donald J. Kouri, and David K. Hoffman. Numerical evaluation of spherical Bessel transforms via Fast Fourier Transforms. Journal of Computational Physics, 100(2):294–296, June 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290236R.

Sinkovits:1993:TRS

[SBO93] R. S. Sinkovits, J. P. Boris, and E. S. Oran. A technique for regularizing the structure of a monotonic Lagrangian grid. Journal of Computational Physics, 108(2):368–372, October 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711903.

Sa:1990:FFS

[SC90] Jong-Youb Sa and Keun-Shik Chang. On far-field stream function condition for two-dimensional incompressible flows. Journal of Computational Physics, 91(2):398–412, December 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190900442.

Sterling:1996:SAL

[SC96a] James D. Sterling and Shiyi Chen. Stability analysis of lattice Boltzmann methods. Journal of Computational Physics, 123(1):196–206, January 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999196900169.

Sundaram:1996:NCS

[SC96b] Shivshankar Sundaram and Lance R. Collins. Numerical considerations in simulating a turbulent suspension of finite-volume particles. Journal of Computational Physics, 124(2):337–350, March 15, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196900649.

Sotiropoulos:1997:PBR

[SC97] F. Sotiropoulos and G. Constantinescu. Pressure-based residual smoothing operators for multistage pseudocompressibility algorithms. Journal of Computational Physics, 133(1):129–145, May 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002199919795662X.

Sun:1992:RICa

[SCA92a] Y. Y. Sun, B. T. Chu, and R. E. Apfel. Radiation-induced cavitation process in a metastable superheated liquid I. Initial and pre-bubble formation stages. Journal of Computational Physics, 103(1):116–125, November 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290328V.

Sun:1992:RICb

[SCA92b] Y. Y. Sun, B. T. Chu, and R. E. Apfel. Radiation-induced cavitation process in a metastable superheated liquid II. Interface formation and post-interface formation stages. Journal of Computational Physics, 103(1):126–140, November 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290329W.

Shuen:1993:CIM

[SCC93a] Jian-Shun Shuen, Kuo-Huey Chen, and Yunho Choi. A coupled implicit method for chemical non-equilibrium flows at all speeds. Journal of Computational Physics, 106(2):306–318, June 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183711101.

Szego:1993:NSB

[SCC93b] Sandor Szego, Pasquale Cinnella, and Al B. Cunningham. Numerical simulation of biofilm processes in closed conduits. Journal of Computational Physics, 108(2):246–263, October 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711800.

Sidikman:1994:DWT

[SCG94] K. L. Sidikman, B. A. Carreras, and L. Garcia. Drift wave turbulence in a plasma with sheared flow. Journal of Computational Physics, 114(1):100–112, September 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711521.

Schwarz:1990:VVT

[Sch90] K. W. Schwarz. On the validity of vortex-tangle simulations of homogeneous superfluid turbulence. Journal of Computational Physics, 87(1):237–240, March 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090236T.

Schoombie:1992:DMSb

[Sch92a] S. W. Schoombie. A discrete multiple scales analysis of a discrete version of the Korteweg–de Vries equation. Journal of Computational Physics, 99(2):352, April 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290222K.

Schoombie:1992:DMSa

[Sch92b] S. W. Schoombie. A discrete multiple scales analysis of a discrete version of the Korteweg–de Vries equation. Journal of Computational Physics, 101(1):55–70, July 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290042W.

Schulkes:1992:IES

[Sch92c] R. M. S. M. Schulkes. Interactions of an elastic solid with a viscous fluid: Eigenmode analysis. Journal of Computational Physics, 100(2):270–283, June 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290234P.

Schmitz:1995:CGF

[Sch95a] F. Schmitz. On the calculation of the gravitational force of axi-symmetric infinitely thin disks. Journal of Computational Physics, 121(1):185–189, October 1, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711886.

Schoen:1995:TEM

[Sch95b] Martin Schoen. Taylor-expansion Monte Carlo simulations of classical fluids in the canonical and grand canonical ensemble. Journal of Computational Physics, 118(1):159–171, April 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002199918571087X.

Schmidt:1996:CTD

[Sch96] Alfred Schmidt. Computation of three dimensional dendrites with finite elements. Journal of Computational Physics, 125(2):293– 312, May 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196900959.

Schlick:1999:CMB

[Sch99] Tamar Schlick. Computational molecular biophysics today: a confluence of methodological advances and complex biomolecular applications. Journal of Computational Physics, 151(1):1–8, May 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199962630.

Salazar:1998:ASU

[SCM98] Abraham Salazar, Antonio Campo, and Biagio Morrone. On approximate solutions for unsteady conduction in slabs with uniform heat flux. Journal of Computational Physics, 144 (2):402–422, August 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199919895999X.

Shyy:1992:SNOa

[SCMU92a] W. Shyy, M.-H. Chen, R. Mittal, and H. S. Udaykumar. On the suppression of numerical oscillations using a non-linear filter. Journal of Computational Physics, 100(2):435, June 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192902662.

Shyy:1992:SNOb

[SCMU92b] W. Shyy, M.-H. Chen, R. Mittal, and H. S. Udaykumar. On the suppression of numerical oscillations using a non-linear filter. Journal of Computational Physics, 102(1):49–62, September 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910580004X.

Scott:1996:SOA

[Sco96] Bruce D. Scott. A second-order accurate implicit scheme for strongly coupled fluid equations applied to fluid electron turbulence in a magnetised plasma. Journal of Computational Physics, 124(1):71–84, March 1, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900455.

Scobelev:1998:QAS

[Sco98] B. Yu. Scobelev. On the question about the sufficiency of the von Neumann criterion for stability of difference schemes. Journal of Computational Physics, 143(1):278–282, June 10, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959368.

Skeldon:1997:GDC

[SCR97] A. C. Skeldon, K. A. Cliffe, and D. S. Riley. Grid design for the computation of a hexagon-roll interaction using a finite element method. Journal of Computational Physics, 133(1):18–26, May 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196956016.

Succi:1999:ILR

[SCT^{+}99] Sauro Succi, Hudong Chen, Chris Teixeira, Gino Bella, A. De Maio, and Kim Molvig. An integer lattice realization of a Lax scheme for transport processes in multiple component fluid flows. Journal of Computational Physics, 152(2):493–516, July 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199962423.

Sun:1994:IEH

[SCW94] Pu Sun, David P. Chock, and Sandra L. Winkler. An implicitexplicit hybrid solver for a system of stiff kinetic equations. Journal of Computational Physics, 115(2):515–523, December 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999184712137.

Sathyaprakash:1993:PAF

[SD93a] B. S. Sathyaprakash and S. V. Dhurandhar. A parallel algorithm for filtering gravitational waves from coalescing binaries. Journal of Computational Physics, 109(2):215–221, December 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183712131.

Schneidesch:1993:CCM

[SD93b] C. R. Schneidesch and M. O. Deville. Chebyshev collocation method and multi-domain decomposition for Navier– Stokes equations in complex curved geometries. Journal of Computational Physics, 106(2):234–257, June 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711058.

Stiller:1994:BSM

[SDG94] L. Stiller, L. L. Daemen, and J. E. Gubernatis. N-body simulations on massively parallel architectures. Journal of Computational Physics, 115(2):550–552, December 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184712162.

Storti:1998:ADN

[SDI98] M. Storti, J. D’Elía, and S. Idelsohn. Algebraic discrete nonlocal (DNL) absorbing boundary condition for the ship wave resistance problem. Journal of Computational Physics, 146(2):570–602, November 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198960697.

Steelant:1997:ARM

[SDP97] J. Steelant, E. Dick, and S. Pattijn. Analysis of robust multigrid methods for steady viscous low Mach number flows. Journal of Computational Physics, 136(2):603–628, September 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197957867.

Shokoohi:1990:AET

[SE90] F. Shokoohi and H. G. Elrod. Algorithms for Eulerian treatment of jet breakup induced by surface tension. Journal of Computational Physics, 89(2):483–487, August 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090154S.

Samson:1998:SRF

[SE98] J. H. Samson and G. A. Evans. Symmetry reduction of Fourier kernels. Journal of Computational Physics, 142(1):109–122, May 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198959204.

Semtner:1997:INM

[Sem97] Albert J. Semtner. Introduction to “A Numerical Method for the Study of the Circulation of the World Ocean”. Journal of Computational Physics, 135(2):149–153, August 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957181.

Sen:1999:CFE

[Sen99] R. Sen. Closed-form expressions for certain induction integrals involving Jacobi and Chebyshev polynomials. Journal of Computational Physics, 156(2):393–398, December 10, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963623.

Serna:1996:IHC

[Ser96] Arturo Serna. Implementation of hierarchical clustering methods. Journal of Computational Physics, 129(1):30–40, November 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196902314.

Sethian:1994:CFE

[Set94] J. A. Sethian. Curvature flow and entropy conditions applied to grid generation. Journal of Computational Physics, 115(2):440– 454, December 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184712095.

Shang:1996:CSC

[SF96] J. S. Shang and Robert M. Fithen. A comparative study of characteristic-based algorithms for the Maxwell equations. Journal of Computational Physics, 125(2):378–394, May 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919690100X.

Szumbarski:1999:DSM

[SF99] J. Szumbarski and J. M. Floryan. A direct spectral method for determination of flows over corrugated boundaries. Journal of Computational Physics, 153(2):378–402, August 10, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962824.

Suh:1991:NSQ

[SFB91] Nam-Duk Suh, Marl R. Feix, and Pierre Bertrand. Numerical simulation of the quantum Liouville–Poisson system. Journal of Computational Physics, 94(2):403–418, June 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190227C.

Schutte:1999:DAC

[SFHD99] Ch Schütte, A. Fischer, W. Huisinga, and P. Deuflhard. A direct approach to conformational dynamics based on hybrid Monte Carlo. Journal of Computational Physics, 151(1):146–168, May 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199962319.

Smolarkiewicz:1990:MPD

[SG90] Piotr K. Smolarkiewicz and Wojciech W. Grabowski. The multidimensional positive definite advection transport algorithm: nonoscillatory option. Journal of Computational Physics, 86(2):355–375, February 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919090105A.

Smolarkiewicz:1992:CMIa

[SG92a] Piotr K. Smolarkiewicz and Georg A. Grell. A class of monotone interpolation schemes. Journal of Computational Physics, 100 (2):435, June 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290265Z.

Smolarkiewicz:1992:CMIb

[SG92b] Piotr K. Smolarkiewicz and Georg A. Grell. A class of monotone interpolation schemes. Journal of Computational Physics, 101(2):431–440, August 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919290018T.

Sternberg:1994:SMM

[SG94] Natalia Sternberg and Valery A. Godyak. Solving the mathematical model of the electrode sheath in symmetrically driven RF discharges. Journal of Computational Physics, 111(2):347– 353, April 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184710680.

Stockie:1998:SMF

[SG98] John M. Stockie and Sheldon I. Green. Simulating the motion of flexible pulp fibres using the immersed boundary method. Journal of Computational Physics, 147(1):147–165, November 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198960867.

Stratton:1994:CAS

[SGW94] David M. Stratton, Jason Gans, and Eric Williams. Coagulation algorithms with size binning. Journal of Computational Physics, 112(2):364–369, June 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711065.

Schaefer:1990:MCF

[SH90] G. Schaefer and P. Hui. The Monte Carlo flux method. Journal of Computational Physics, 89(1):1–30, July 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090114G.

Saitou:1992:TDU

[SH92] M. Saitou and A. Hirata. Two-dimensional unsteady solidification problem calculated by using the boundary-fitted coordinate system. Journal of Computational Physics, 100(1):188–196, May 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290319T.

Song:1994:SIO

[SH94] Yuhe Song and Dale Haidvogel. A semi-implicit ocean circulation model using a generalized topography-following coordinate system. Journal of Computational Physics, 115(1):228–244, November 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999184711892.

Suresh:1997:AMP

[SH97] A. Suresh and H. T. Huynh. Accurate monotonicity-preserving schemes with Runge–Kutta time stepping. Journal of Computational Physics, 136(1):83–99, September 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957454.

Schumer:1998:VSU

[SH98a] Joseph W. Schumer and James Paul Holloway. Vlasov simulations using velocity-scaled Hermite representations. Journal of Computational Physics, 144(2):626–661, August 10, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959253.

Stanescu:1998:SLD

[SH98b] D. Stanescu and W. G. Habashi. 2N-storage low dissipation and dispersion Runge–Kutta schemes for computational acoustics. Journal of Computational Physics, 143(2):674–681, July 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198959861.

Safouhi:1999:ERE

[SH99] H. Safouhi and P. E. Hoggan. Efficient and rapid evaluation of three-center two electron Coulomb and hybrid integrals using nonlinear transformations. Journal of Computational Physics, 155(2):331–347, November 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963398.

Shang:1995:FSM

[Sha95a] J. S. Shang. A fractional-step method for solving 3D, time-domain Maxwell equations. Journal of Computational Physics, 118(1):109–119, April 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999185710832.

Swegle:1995:SPH

[SHA95b] J. W. Swegle, D. L. Hicks, and S. W. Attaway. Smoothed particle hydrodynamics stability analysis. Journal of Computational Physics, 116(1):123–134, January 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710108.

Shang:1996:TDE

[Sha96] J. S. Shang. Time-domain electromagnetic scattering simulations on multicomputers. Journal of Computational Physics, 128(2):381–390, October 15, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999196902181.

Shang:1999:HOC

[Sha99] J. S. Shang. High-order compact-difference schemes for time-dependent Maxwell equations. Journal of Computational Physics, 153(2):312–333, August 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962794.

Shen:1990:HBU

[She90] Jie Shen. Hopf bifurcation of the unsteady regularized driven cavity flow. Journal of Computational Physics, 91(1):254, November 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919090032V.

Shen:1991:HBU

[She91] Jie Shen. Hopf bifurcation of the unsteady regularized driven cavity flow. Journal of Computational Physics, 95(1):228–245, July 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190261I.

Shen:1995:FDP

[She95] Jie Shen. On fast direct Poisson solver, INF–SUP constant and iterative Stokes solver by Legendre–Galerkin method. Journal of Computational Physics, 116(1):184–188, January 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710170.

Smith:1997:MDP

[SHF97] Steven W. Smith, Carol K. Hall, and Benny D. Freeman. Molecular dynamics for polymeric fluids using discontinuous potentials. Journal of Computational Physics, 134(1):16–30, June 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196955102.

Shirayama:1993:PCV

[Shi93] Susumu Shirayama. Processing of computed vector fields for visualization. Journal of Computational Physics, 106(1):30–41, May 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183710880.

Short:1993:DCM

[Sho93] Kevin M. Short. Direct calculation of metric entropy from time series. Journal of Computational Physics, 104(1):162–172, January 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183710181.

Shuen:1990:IFSa

[Shu90] Jian-Shun Shuen. Inviscid flux-splitting algorithms for real gases with non-equilibrium chemistry. Journal of Computational Physics, 87(2):494, April 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999190902653.

Shuen:1991:UDF

[Shu91] Jian-Shun Shuen. Upwind differencing and LU factorization for chemical non-equilibrium Navier–Stokes equations. Journal of Computational Physics, 97(2):584, December 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190031F.

Shu:1992:NMS

[Shu92a] Chi-Wang Shu. A numerical method for systems of conservation laws of mixed type admitting hyperbolic flux splitting. Journal of Computational Physics, 100(2):424–429, June 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290249X.

Shuen:1992:UDLa

[Shu92b] Jian-Shun Shuen. Upwind differencing and LU factorization for chemical non-equilibrium Navier–Stokes equations. Journal of Computational Physics, 99(1):179, March 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192902879.

Shuen:1992:UDLb

[Shu92c] Jian-Shun Shuen. Upwind differencing and LU factorization for chemical non-equilibrium Navier–Stokes equations. Journal of Computational Physics, 99(2):233–250, April 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290205D.

Shubin:1995:AAM

[Shu95] Gregory R. Shubin. Application of alternative multidisciplinary optimization formulations to a model problem for static aeroelasticity. Journal of Computational Physics, 118(1):73–85, April 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185710807.

Shu:1997:PRU

[Shu97] Chi-Wang Shu. Preface to the republication of “Uniformly High Order Essentially Non-oscillatory Schemes, III,” by Harten, Engquist, Osher, and Chakravarthy. Journal of Computational Physics, 131(1):1–2, February 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196956302.

Shyue:1998:ESC

[Shy98] Keh-Ming Shyue. An efficient shock-capturing algorithm for compressible multicomponent problems. Journal of Computational Physics, 142(1):208–242, May 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959307.

Shyue:1999:FMT

[Shy99] Keh-Ming Shyue. A fluid-mixture type algorithm for compressible multicomponent flow with van der Waals equation of state. Journal of Computational Physics, 156(1):43–88, November 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199963490.

Shopov:1994:NSS

[SI94] Peter J. Shopov and Yuriy I. Iordanov. Numerical solution of Stokes equations with pressure and filtration boundary conditions. Journal of Computational Physics, 112(1):12–23, May 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999184710783.

Siewert:1999:DOS

[Sie99] C. E. Siewert. A discrete-ordinates solution for heat transfer in a plane channel. Journal of Computational Physics, 152(1):251– 263, June 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999199962447.

Silling:1992:SAD

[Sil92] S. A. Silling. Stability and accuracy of differencing methods for viscoplastic models in wavecodes. Journal of Computational Physics, 101(2):453, August 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999192900303.

Silling:1993:SAD

[Sil93] S. A. Silling. Stability and accuracy of differencing methods for viscoplastic models in wavecodes. Journal of Computational Physics, 104(1):30–40, January 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710041.

Simos:1993:NVS

[Sim93] T. E. Simos. New variable-step procedure for the numerical integration of the one-dimensional Schrödinger equation. Journal of Computational Physics, 108(1):175–179, September 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711721.

Simos:1999:PSE

[Sim99] T. E. Simos. P-stable exponentially fitted methods for the numerical integration of the Schrödinger equation. Journal of Computational Physics, 148(2):305–321, January 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961034.

Sjogreen:1995:HOC

[Sjö95a] Björn Sjögreen. High order centered difference methods for the compressible Navier–Stokes equations. Journal of Computational Physics, 117(1):67–78, March 1, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710455.

Sjogreen:1995:NEM

[Sjö95b] Björn Sjögreen. Numerical experiments with the multiresolution scheme for the compressible Euler equations. Journal of Computational Physics, 117(2):251–261, March 15, 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710637.

Schroder:1990:WTV

[SK90] W. Schröder and H. B. Keller. Wavy Taylor-vortex flows via multigrid-continuation methods. Journal of Computational Physics, 91(1):197–227, November 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090011O.

Su:1992:NSE

[SK92] Y. Y. Su and B. Khomami. Numerical solution of eigenvalue problems using spectral techniques. Journal of Computational Physics, 100(2):297–305, June 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290237S.

Sidilkover:1993:NOS

[SK93] David Sidilkover and George Em Karniadakis. Non-oscillatory spectral element Chebyshev method for shock wave calculations. Journal of Computational Physics, 107(1):10–22, July 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711216.

Scandrett:1994:DAA

[SK94] C. L. Scandrett and G. A. Kriegsmann. Decoupling approximations applied to an infinite array of fluid loaded baffled membranes. Journal of Computational Physics, 111(2):282–290, April 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999184710643.

Sherwin:1996:THF

[SK96] S. J. Sherwin and G. E. Karniadakis. Tetrahedral hp finite elements: Algorithms and flow simulations. Journal of Computational Physics, 124(1):14–45, March 1, 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919690042X.

Schulz:1998:UFS

[SK98] Karl W. Schulz and Yannis Kallinderis. Unsteady flow structure interaction for incompressible flows using deformable hybrid grids. Journal of Computational Physics, 143(2):569–597, July 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198959691.

Skeel:1998:CNI

[Ske98] Robert D. Skeel. Comments on “Numerical Instability due to Varying Time Steps in Explicit Wave Propagation and Mechanics Calculations” by Joseph P. Wright. Journal of Computational Physics, 145(2):758–759, September 20, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960727. See [Wri98].

Saitou:1990:NSS

[SKK90] Masatoshi Saitou, Elizaburo Kanda, and Mitsuo Kawashima. A numerical solution of the steady solidification problem in two dimensions by boundary-fitted coordinate systems. Journal of Computational Physics, 90(1):268, September 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090209J.

Saitou:1991:NSS

[SKK91] Masatoshi Saitou, Eizaburo Kanda, and Mitsuo Kawashima. A numerical solution of the steady solidification problem in two dimensions by boundary-fitted coordinate systems. Journal of Computational Physics, 94(1):138–150, May 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190140G.

Sherwin:1994:NSI

[SKO94] S. J. Sherwin, G. E. Karniadakis, and S. A. Orszag. Numerical simulation of the ion etching process. Journal of Computational Physics, 110(2):373–398, February 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710333.

Schneider:1993:NAU

[SKR^{+}93] V. Schneider, U. Katscher, D. H. Rischke, B. Waldhauser, J. A. Maruhn, and C.-D. Munz. New algorithms for ultra-relativistic numerical hydrodynamics. Journal of Computational Physics, 105(1):92–107, March 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710569.

Stundzia:1996:SSC

[SL96] Audrius B. Stundzia and Charles J. Lumsden. Stochastic simulation of coupled reaction-diffusion processes. Journal of Computational Physics, 127(1):196–207, August 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901680.

Strauss:1998:AFE

[SL98] H. R. Strauss and D. W. Longcope. An adaptive finite element method for magnetohydrodynamics. Journal of Computational Physics, 147(2):318–336, December 10, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960910.

Saurel:1994:EAR

[SLL94] Richard Saurel, Michel Larini, and Jean Claude Loraud. Exact and approximate Riemann solvers for real gases. Journal of Computational Physics, 112(1):126–137, May 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710862.

Schnack:1998:FVA

[SLMS98] D. D. Schnack, I. Lottati, Z. Mikić, and P. Satyanarayana. A finite-volume algorithm for three-dimensional magnetohydrodynamics on an unstructured, adaptive grid in axially symmetric geometry. Journal of Computational Physics, 140(1):71–121, February 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198958739.

Shuen:1990:IFSb

[SLV90] Jian-Shun Shuen, Meng-Sing Liou, and Bram Van Leer. Inviscid flux-splitting algorithms for real gases with non-equilibrium chemistry. Journal of Computational Physics, 90(2):371–395, October 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919090172W.

Stone:1992:UMIa

[SM92] James M. Stone and Dimitri Mihalas. Upwind monotonic interpolation methods for the solution of the time dependent radiative transfer equation. Journal of Computational Physics, 100(2): 402–408, June 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290246U.

Shariff:1998:TDM

[SM98a] Karim Shariff and Robert D. Moser. Two-dimensional mesh embedding for B-spline methods. Journal of Computational Physics, 145(2):471–488, September 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960533.

Smolarkiewicz:1998:MFD

[SM98b] Piotr K. Smolarkiewicz and Len G. Margolin. MPDATA: a finite-difference solver for geophysical flows. Journal of Computational Physics, 140(2):459–480, March 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959010.

Sanders:1998:MDU

[SMD98] Richard Sanders, Eric Morano, and Marie-Claude Druguet. Multidimensional dissipation for upwind schemes: Stability and applications to gas dynamics. Journal of Computational Physics, 145(2):511–537, September 20, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960478.

Smith:1999:AAG

[Smi99] Richard W. Smith. AUSM(ALE): a geometrically conservative arbitrary Lagrangian–Eulerian flux splitting scheme. Journal of Computational Physics, 150(1):268–286, March 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961800.

Sheffer:1998:EMA

[SMJ98] Scott G. Sheffer, Luigi Martinelli, and Antony Jameson. An efficient multigrid algorithm for compressible reactive flows. Journal of Computational Physics, 144(2):484–516, August 10, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960004.

Scott:1990:PSQ

[SMM^{+}90] T. C. Scott, R. A. Moore, M. B. Monagan, G. J. Fee, and E. R. Vrscay. Perturbative solutions of quantum mechanical problems by symbolic computation. Journal of Computational Physics, 87(2):366–395, April 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999190902583.

Smolsky:1998:LTS

[Smo98] Michael V. Smolsky. Long time-step particle pushing in drift approximation without orbit averaging. Journal of Computational Physics, 145(1):41–60, September 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960028.

Spalart:1991:SMN

[SMR91] Philippe R. Spalart, Robert D. Moser, and Michael M. Rogers. Spectral methods for the Navier–Stokes equations with one infinite and two periodic directions. Journal of Computational Physics, 96(2):297–324, October 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190238G.

Schlick:1998:NRA

[SMSS98] Tamar Schlick, Margaret Mandziuk, Robert D. Skeel, and K. Srinivas. Nonlinear resonance artifacts in molecular dynamics simulations. Journal of Computational Physics, 140(1):1–29, February 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199919895879X.

Sesterhenn:1999:CPC

[SMT99] Jörn Sesterhenn, Bernhard Müller, and Hans Thomann. On the cancellation problem in calculating compressible low Mach number flows. Journal of Computational Physics, 151(2):597– 615, May 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999199962113.

Shen:1999:WDD

[SMW99] Jie Shen, T. Tachim Medjo, and S. Wang. On a winddriven, double-gyre, quasi-geostrophic ocean model: Numerical simulations and structural analysis. Journal of Computational Physics, 155(2):387–409, November 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963441.

Steppeler:1990:FES

[SNL90] J. Steppeler, I. M. Navon, and H-I Lu. Finite-element schemes for extended integrations of atmospheric models. Journal of Computational Physics, 89(1):95–124, July 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090119L.

Sha:1998:ASO

[SNU98] Weiming Sha, Koichi Nakabayashi, and Hiromasa Ueda. An accurate second-order approximation factorization method for timedependent incompressible Navier–Stokes equations in spherical polar coordinates. Journal of Computational Physics, 142(1):47– 66, May 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198959216.

Safjan:1995:HOT

[SO95] A. Safjan and J. T. Oden. High-order Taylor–Galerkin methods for linear hyperbolic systems. Journal of Computational Physics, 120(2):206–230, September 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918571159X.

Solomonoff:1992:FAS

[Sol92] Alex Solomonoff. A fast algorithm for spectral differentiation. Journal of Computational Physics, 98(1):174–177, January 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290182X.

Sorets:1995:FFT

[Sor95] Eugene Sorets. Fast Fourier Transforms of piecewise constant functions. Journal of Computational Physics, 116(2):369–379, February 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999185710352.

Stuhne:1996:VEA

[SP96] G. R. Stuhne and W. R. Peltier. Vortex erosion and amalgamation in a new model of large scale flow on the sphere. Journal of Computational Physics, 128(1):58–81, October 1, 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901965.

Sabbah:1998:DFM

[SP98] C. Sabbah and R. Pasquetti. A divergence-free multidomain spectral solver of the Navier–Stokes equations in geometries of high aspect ratio. Journal of Computational Physics, 139(2):359–379, January 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999197958757.

Stuhne:1999:NIG

[SP99] G. R. Stuhne and W. R. Peltier. New icosahedral grid-point discretizations of the shallow water equations on the sphere. Journal of Computational Physics, 148(1):23–58, January 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961198.

Saleem:1992:ACS

[SPC92] M. Saleem, T. Pulliam, and A. Y. Cheer. Acceleration of convergence and spectrum transformation of implicit finite difference operators associated with Navier–Stokes equations. Journal of Computational Physics, 101(2):453, August 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290027V.

Saleem:1993:ACS

[SPC93] M. Saleem, T. Pulliam, and A. Y. Cheer. Acceleration of convergence and spectrum transformation of implicit finite difference operators associated with Navier–Stokes equations. Journal of Computational Physics, 104(1):1–13, January 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710016.

Sandu:1996:EIF

[SPCD96] A. Sandu, F. A. Potra, G. R. Carmichael, and V. Damian. Efficient implementation of fully implicit methods for atmospheric chemical kinetics. Journal of Computational Physics, 129(1):101– 110, November 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196902363.

Spekreijse:1995:EGG

[Spe95] S. P. Spekreijse. Elliptic grid generation based on Laplace equations and algebraic transformations. Journal of Computational Physics, 118(1):38–61, April 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999185710789.

Steinberg:1990:AGG

[SR90] Stanly Steinberg and Patrick J. Roache. Anomalies in grid generation on curves. Journal of Computational Physics, 91(2):255– 277, December 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999190900372.

Steinberg:1992:VCSb

[SR92a] Stanley Steinberg and Patrick Roache. Variational curve and surface grid generation. Journal of Computational Physics, 99 (1):181, March 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290302F.

Steinberg:1992:VCSa

[SR92b] Stanly Steinberg and Patrick Roache. Variational curve and surface grid generation. Journal of Computational Physics, 100(1): 163–178, May 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290317R.

Srinivasan:1997:SBG

[SR97] Kumar Srinivasan and Stanley G. Rubin. Solution-based grid adaptation through segmented multigrid domain decomposition. Journal of Computational Physics, 136(2):467–493, September 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197957740.

Slinn:1998:MST

[SR98] Donald N. Slinn and J. J. Riley. A model for the simulation of turbulent boundary layers in an incompressible stratified flow. Journal of Computational Physics, 144(2):550–602, August 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197959830.

Spivack:1999:RVC

[SR99] M. Spivack and D. E. Reeve. Recovery of a variable coefficient in a coastal evolution equation. Journal of Computational Physics, 151(2):585–596, May 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962046.

Susan-Resiga:1998:DDM

[SRA98] Romeo F. Susan-Resiga and Hafiz M. Atassi. A domain decomposition method for the exterior Helmholtz problem. Journal of Computational Physics, 147(2):388–401, December 10, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960946.

Sonnendrucker:1999:SLM

[SRBG99] Eric Sonnendrücker, Jean Roche, Pierre Bertrand, and Alain Ghizzo. The semi-Lagrangian method for the numerical resolution of the Vlasov equation. Journal of Computational Physics, 149(2):201–220, March 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961484.

Shen:1993:SIS

[SRF93] Cherng-Yeu Shen, Helen L. Reed, and Thomas A. Foley. Shepard’s Interpolation for solution-adaptive methods. Journal of Computational Physics, 106(1):52–61, May 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710909.

Szymczak:1993:NAH

[SRSB93] William G. Szymczak, Joel C. W. Rogers, Jay M. Solomon, and Alan E. Bergert. A numerical algorithm for hydrodynamic free boundary problems. Journal of Computational Physics, 106(2): 319–336, June 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999183711113.

Swanson:1998:SND

[SRT98] R. C. Swanson, R. Radespiel, and E. Turkel. On some numerical dissipation schemes. Journal of Computational Physics, 147(2):518–544, December 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999198961009.

Silver:1996:KPA

[SRVK96] R. N. Silver, H. Roeder, A. F. Voter, and J. D. Kress. Kernel polynomial approximations for densities of states and spectral functions. Journal of Computational Physics, 124(1):115–130, March 1, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196900480.

Sethian:1991:CGD

[SS91] James A. Sethian and John Strain. Crystal growth and dendritic solidification. Journal of Computational Physics, 97(2): 581, December 1991. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919190020L.

Sethian:1992:CGD

[SS92] James A. Sethian and John Straint. Crystal growth and dendritic solidification. Journal of Computational Physics, 98(2):231–253, February 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290140T.

Salari:1994:FCT

[SS94] Kambiz Salari and Stanly Steinberg. Flux-corrected transport in a moving grid. Journal of Computational Physics, 111(1):24– 32, March 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184710400.

Shashkov:1995:SOF

[SS95] Mikhail Shashkov and Stanly Steinberg. Support-operator finitedifference algorithms for general elliptic problems. Journal of Computational Physics, 118(1):131–151, April 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710856.

Schar:1996:SIF

[SS96a] Christoph Schär and Piotr K. Smolarkiewicz. A synchronous and iterative flux-correction formalism for coupled transport equations. Journal of Computational Physics, 128(1):101–120, October 1, 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196901989.

Schellhase:1996:SAR

[SS96b]  A. R. Schellhase and R. G. Storer.  Spectral analysis

of resistive MHD in toroidal geometry. Journal of Computational Physics, 123(1):15–31, January 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900029.

Shashkov:1996:SDE

[SS96c] Mikhail Shashkov and Stanly Steinberg. Solving diffusion equations with rough coefficients in rough grids. Journal of Computational Physics, 129(2):383–405, December 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196902570.

Sanchez:1999:CPS

[SS99a] Roberto Sánchez and Andrej Sali. Comparative protein structure modeling in genomics. Journal of Computational Physics, 151(1):388–401, May 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999199962009.

Sandu:1999:MRA

[SS99b] Adrian Sandu and Tamar Schlick. Masking resonance artifacts in force-splitting methods for biomolecular simulations by extrapolative Langevin dynamics. Journal of Computational Physics, 151(1):74–113, May 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962022.

Shen:1999:QNS

[SS99c] Wen Zhong Shen and Jens Nørkær Sørensen. Quasi-3D Navier–Stokes model for a rotating airfoil. Journal of Computational Physics, 150(2):518–548, April 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962034.

Schumack:1990:SMS

[SSB90] Mark R. Schumack, William W. Schultz, and John P. Boyd. Spectral method solution of the Stokes equations on nonstaggered grids. Journal of Computational Physics, 89(2):489–490, August 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919090161S.

Schumack:1991:SMS

[SSB91] Mark R. Schumack, William W. Schultz, and John P. Boyd. Spectral method solution of the Stokes equations on nonstaggered grids. Journal of Computational Physics, 94(1):30–58, May 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999191901369.

Schlick:1999:ACC

[SSB^{+}99] Tamar Schlick, Robert D. Skeel, Axel T. Brunger, Laxmikant V. Kalé, John A. Board, Jr., Jan Hermans, and Klaus Schulten. Algorithmic challenges in computational molecular biophysics. Journal of Computational Physics, 151(1):9–48, May 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198961824.

Scales:1992:GOMa

[SSF92a] John A. Scales, Martin L. Smith, and Terri L. Fischer. Global optimization methods for highly multimodal inverse problems. Journal of Computational Physics, 101(1):228–229, July 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999192900679.

Scales:1992:GOMb

[SSF92b] John A. Scales, Martin L. Smith, and Terri L. Fischer. Global optimization methods for multimodal inverse problems. Journal of Computational Physics, 103(2):258–268, December 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290400S.

Sussman:1994:LSA

[SSO94] Mark Sussman, Peter Smereka, and Stanley Osher. A level set approach for computing solutions to incompressible two-phase flow. Journal of Computational Physics, 114(1):146–159, September 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999184711557.

Sackinger:1996:NRP

[SSR96] P. A. Sackinger, P. R. Schunk, and R. R. Rao. A Newton– Raphson pseudo-solid domain mapping technique for free and moving boundary problems: a finite element implementation. Journal of Computational Physics, 125(1):83–103, April 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196900819.

Shashkov:1998:LRV

[SSW98] Mikhail Shashkov, Blair Swartz, and Burton Wendroff. Local reconstruction of a vector field from its normal components on the faces of grid cells. Journal of Computational Physics, 139(2):406–409, January 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958770.

Song:1992:DGV

[ST92a] Yuhe Song and Tao Tang. Dispersion and group velocity in numerical schemes for three-dimensional hydrodynamic equations. Journal of Computational Physics, 103(1):187, November 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290338Y.

Swanson:1992:CDU

[ST92b] R. C. Swanson and Eli Turkel. On central-difference and upwind schemes. Journal of Computational Physics, 101(2):292– 306, August 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290007L.

Song:1993:DGV

[ST93] Yuhe Song and Tao Tang. Dispersion and group velocity in numerical schemes for three-dimensional hydrodynamic equations. Journal of Computational Physics, 105(1):72–82, March 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183710545.

Sawley:1995:CPP

[ST95] M. L. Sawley and J. K. Tegnér. A comparison of parallel programming models for multiblock flow computations. Journal of Computational Physics, 122(2):280–290, December 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185712132.

Simos:1997:SEO

[ST97] T. E. Simos and Ch. Tsitouras. A P -stable eighth-order method for the numerical integration of periodic initial-value problems. Journal of Computational Physics, 130(1):123–128, January 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196955679.

Sun:1999:CSA

[ST99] M. Sun and K. Takayama. Conservative smoothing on an adaptive quadrilateral grid. Journal of Computational Physics, 150(1):143–180, March 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999198961678.

Savvidy:1991:MCS

[STAS91] G. K. Savvidy and N. G. Ter-Arutyunyan-Savvidy. On the Monte Carlo simulation of physical systems. Journal of Computational Physics, 97(2):566–572, December 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190015D.

Steppeler:1992:SSWb

[Ste92a] J. Steppeler. Solution of the shallow water equations using hybrid grids. Journal of Computational Physics, 98(2):349, February 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290166V.

Steppeler:1992:SSWa

[Ste92b] J. Steppeler. Solution of the shallow water equations using hybrid grids. Journal of Computational Physics, 100(2):419–423, June 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290248W.

Sherman:1996:FSO

[STG^{+}96] Laura L. Sherman, Arthur C. Taylor III, Larry L. Green, Perry A. Newman, Gene W. Hou, and Vamshi Mohan Korivi. First- and second-order aerodynamic sensitivity derivatives via automatic differentiation with incremental iterative methods. Journal of Computational Physics, 129(2):307–331, December 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196902521.

Stone:1992:UMIb

[Sto92] James M. Stone. Upwind monotonic interpolation methods for the solution of the time dependent radiative transfer equation.

Journal of Computational Physics, 99(2):351, April 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290220S.

Strain:1991:FPT

[Str91] John Strain. Fast potential theory II: Layer potentials and discrete sums. Journal of Computational Physics, 97(2):583, December 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190028J.

Strain:1992:FPTa

[Str92a] John Strain. Fast potential theory. II. Layer potentials and discrete sums. Journal of Computational Physics, 99(1):179, March 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290289B.

Strain:1992:FPTb

[Str92b] John Strain. Fast potential theory. II. layer potentials and discrete sums. Journal of Computational Physics, 99(2):251– 270, April 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290206E.

Strand:1994:SPF

[Str94] Bo Strand. Summation by parts for finite difference approximations for d/dx. Journal of Computational Physics, 110(1):47– 67, January 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999184710059. See erratum

[Ano95u].

Strain:1995:SMN

[Str95] John Strain. Spectral methods for nonlinear parabolic systems. Journal of Computational Physics, 122(1):1–12, November 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185711916.

Strain:1996:VMS

[Str96] John Strain. 2D vortex methods and singular quadrature rules. Journal of Computational Physics, 124(1):131–145, March 1, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196900492.

Strain:1997:FAV

[Str97] John Strain. Fast adaptive 2D vortex methods. Journal of Computational Physics, 132(1):108–122, March 15, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196956260.

Strain:1999:FTB

[Str99a] John Strain. Fast tree-based redistancing for level set computations. Journal of Computational Physics, 152(2):664–686, July 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199962599.

Strain:1999:SLM

[Str99b] John Strain. Semi-Lagrangian methods for level set equations. Journal of Computational Physics, 151(2):498–533, May 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199961946.

Strain:1999:TMM

[Str99c] John Strain. Tree methods for moving interfaces. Journal of Computational Physics, 151(2):616–648, May 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962058.

Suzuki:1997:IMC

[STS^{+}97] Akiko Suzuki, Tomonori Takizuka, Katsuhiro Shimizu, Nobuhiko Hayashi, Akiyoshi Hatayama, and Masatada Ogasawara. An implicit Monte Carlo method for simulation of impurity transport in divertor plasma. Journal of Computational Physics, 131(1):193– 198, February 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196956004.

Spotz:1998:FSW

[STS98] William F. Spotz, Mark A. Taylor, and Paul N. Swarztrauber. Fast shallow-water equation solvers in latitude-longitude coordinates. Journal of Computational Physics, 145(1):432–444, September 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198960260.

Shadid:1997:INM

[STW97] John N. Shadid, Ray S. Tuminaro, and Homer F. Walker. An inexact Newton method for fully coupled solution of the Navier– Stokes equations with heat and mass transport. Journal of Computational Physics, 137(1):155–185, October 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957983.

Sulsky:1991:NMSa

[Sul91] Deborah Sulsky. A numerical method for suspension flow. Journal of Computational Physics, 94(1):250, May 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190146C.

Sun:1993:NEA

[Sun93] Wen-Yih Sun. Numerical experiments for advection equation. Journal of Computational Physics, 108(2):264–271, October 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183711812.

Sun:1996:PNT

[Sun96] Pu Sun. A pseudo-non-time-splitting method in air quality modeling. Journal of Computational Physics, 127(1):152–157, August 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196901655.

Suresh:1994:ASN

[Sur94] Ambady Suresh. An assessment of spectral nonoscillatory schemes. Journal of Computational Physics, 114(2):339–349, October 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184711715.

Shyy:1991:AVV

[SV91] Wei Shyy and Thi C. Vu. On the adoption of velocity variable and grid system for fluid flow computation in curvilinear coordinates. Journal of Computational Physics, 92(1):82–105, January 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190293T.

Scales:1997:LEL

[SV97] John A. Scales and Erik S. Van Vleck. Lyapunov exponents and localization in randomly layered media. Journal of Computational Physics, 133(1):27–42, May 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956436.

Shankar:1996:NDP

[SvD96] S. Shankar and L. van Dommelen. A new diffusion procedure for vortex methods. Journal of Computational Physics, 127(1):88– 109, August 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196901606.

Stone:1992:NAF

[SVG92] C. A. Stone, M. Vicanek, and N. M. Ghoniem. On the numerical accuracy of the Fokker–Planck approximation to the hierarchy of master equations. Journal of Computational Physics, 102(2): 425, October 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290394E.

Stone:1993:NAF

[SVG93] C. A. Stone, M. Vicanek, and N. M. Ghoniem. On the numerical accuracy of the Fokker–Planck approximation to the hierarchy of master equations. Journal of Computational Physics, 104(2):451– 456, February 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999183710430.

Saxena:1990:CFS

[SVS90] V. K. Saxena, L. L. Van Zandt, and W. K. Schroll. Closed form solution for localized modes on a polymer chain with a defect. Journal of Computational Physics, 89(1):254, July 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090138Q.

Saxena:1991:CFS

[SVS91] V. K. Saxena, L. L. Van Zandt, and W. K. Schroll. Closed form solution for localized modes on a polymer chain with a defect. Journal of Computational Physics, 93(2):273–286, April 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190182K.

Segal:1998:CDF

[SVV98] Guus Segal, Kees Vuik, and Fred Vermolen. A conserving discretization for the free boundary in a two-dimensional Stefan problem. Journal of Computational Physics, 141(1):1–21, March 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198959009.

Sanders:1992:HRS

[SW92] Richard Sanders and Alan Weiser. High resolution staggered mesh approach for nonlinear hyperbolic systems of conservation laws. Journal of Computational Physics, 101(2):314–329, August 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290009N.

Salmon:1994:STC

[SW94] John K. Salmon and Michael S. Warren. Skeletons from the treecode closet. Journal of Computational Physics, 111(1):136– 155, March 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL ftp://ftp.cacr.caltech.edu/
nbody/skeletons.ps.Z;   http://www.ccsf.caltech.edu/~
johns/ftp/nbody/skeletons.ps.Z;   http://www.sciencedirect.
com/science/article/pii/S0021999184710503.

Straughan:1996:TVA

[SW96] B. Straughan and D. W. Walker. Two very accurate and efficient methods for computing eigenvalues and eigenfunctions in porous convection problems. Journal of Computational Physics, 127(1):128–141, August 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901631.

Shashkov:1999:CSG

[SW99a] Mikhail Shashkov and Burton Wendroff. A composite scheme for gas dynamics in Lagrangian coordinates. Journal of Computational Physics, 150(2):502–517, April 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199961922.

Spreiter:1999:CMD

[SW99b] Q. Spreiter and M. Walter. Classical molecular dynamics simulation with the velocity Verlet algorithm at strong external magnetic fields. Journal of Computational Physics, 152(1):102–119, June 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199919996237X.

Stockie:1999:ASI

[SW99c] John M. Stockie and Brian R. Wetton. Analysis of stiffness in the immersed boundary method and implications for time-stepping schemes. Journal of Computational Physics, 154(1):41–64, September 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999199962976.

Swartz:1999:GNM

[Swa99] Blair Swartz. Good neighborhoods for multidimensional Van Leer limiting. Journal of Computational Physics, 154(1):237–241, September 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999199963088.

Sun:1995:TDN

[SWD95] M. T. Sun, S. T. Wu, and Murray Dryer. On the time-dependent numerical boundary conditions of magnetohydrodynamic flows.

Journal of Computational Physics, 116(2):330–342, February 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185710315.

Sweatman:1994:DPB

[Swe94] W. L. Sweatman. The development of a parallel N-body code for the Edinburgh Concurrent Supercomputer. Journal of Computational Physics, 111(1):110–119, March 1, 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710485.

Swesty:1996:TCI

[Swe96] F. Douglas Swesty. Thermodynamically consistent Interpolation for equation of state tables. Journal of Computational Physics, 127(1):118–127, August 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919690162X.

Swift:1996:UHC

[Swi96] Daniel W. Swift. Use of a hybrid code for global-scale plasma simulation. Journal of Computational Physics, 126(1):109–121, June 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196901242.

Saucez:1996:SOS

[SWS96] P. Saucez, A. Vande Wouwer, and W. E. Schiesser. Some observations on a static spatial remeshing method based on equidistribution principles. Journal of Computational Physics, 128(2):274–288, October 15, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999196902107.

Sheu:1998:DHR

[SWT98] Tony W. H. Sheu, S. K. Wang, and S. F. Tsai. Development of a high-resolution scheme for a multi-dimensional advectiondiffusion equation. Journal of Computational Physics, 144(1): 1–16, July 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198959927.

Su:1999:LSF

[SXG99] Mingde Su, Kun Xu, and M. S. Ghidaoui. Low-speed flow simulation by the gas-kinetic scheme. Journal of Computational Physics, 150(1):17–39, March 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961629.

Schmidt:1997:DTB

[SY97] Frank Schmidt and David Yevick. Discrete transparent boundary conditions for Schrödinger-type equations. Journal of Computational Physics, 134(1):96–107, June 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956758.

Sakai:1997:NST

[SZ97] Katsuhiro Sakai and Gen Sheng Zhang. A numerical scheme for transport equations with spatially distributed coefficients based on locally exact difference method. Journal of Computational Physics, 134(2):332–341, July 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957119.

Tamamidis:1991:GOG

[TA91a] P. Tamamidis and D. N. Assanis. Generation of orthogonal grids with control of spacing. Journal of Computational Physics, 94(2): 437–453, June 1991. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919190229E.

Tani:1991:NAM

[TA91b] K. Tani and M. Azumi. Numerical analysis of 2D MHD equilibrium with non-inductive plasma current in tokamaks. Journal of Computational Physics, 93(2):487, April 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190205Y.

Tabak:1996:SOG

[Tab96] Esteban G. Tabak. A second-order Godunov method on arbitrary grids. Journal of Computational Physics, 124(2):383–395, March 15, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196900674.

Tam:1998:PML

[TAC98] Christopher K. W. Tam, Laurent Auriault, and Francesco Cambuli. Perfectly matched layer as an absorbing boundary condition for the linearized Euler equations in open and ducted domains. Journal of Computational Physics, 144(1):213–234, July 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198959976.

Tani:1992:NAM

[TAD92a] K. Tani, M. Azumi, and R. S. Devoto. Numerical analysis of 2D MHD equilibrium with non-inductive plasma current in tokamaks. Journal of Computational Physics, 98(2):332–341, February 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290146P.

Tani:1992:NAZ

[TAD92b] K. Tani, M. Azumi, and R. S. Devoto. Numerical analysis of ZD MHD equilibrium with non-inductive plasma current in tokamaks. Journal of Computational Physics, 98(1):180, January 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999192901966.

Taddei:1997:FIG

[Tad97] S. Taddei. Finite Interpolation in Green function deterministic numerical methods. Journal of Computational Physics, 134(1): 62–74, June 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999197956527.

Tafti:1995:AFP

[Taf95] Danesh Tafti. Alternate formulations for the pressure equation Laplacian on a collocated grid for solving the unsteady incompressible Navier–Stokes equations. Journal of Computational Physics, 116(1):143–153, January 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710121.

Takatsuka:1992:EAFa

[Tak92a] Kazuo Takatsuka. Extraction of accurate frequencies from the Fast-Fourier-Transform spectra. Journal of Computational Physics, 101(1):228, July 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999192900635.

Takatsuka:1992:EAFb

[Tak92b] Kazuo Takatsuka. Extraction of accurate frequencies from the Fast Fourier Transform spectra. Journal of Computational Physics, 102(2):374–380, October 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290379D.

Tanaka:1993:SLF

[Tan93] Motohiko Tanaka. A simulation of low-frequency electromagnetic phenomena in kinetic plasmas of three dimensions. Journal of Computational Physics, 107(1):124–145, July 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711307.

Tanaka:1994:FVT

[Tan94] T. Tanaka. Finite volume TVD scheme on an unstructured grid system for three-dimensional MHD simulation of inhomogeneous systems including strong background potential fields. Journal of Computational Physics, 111(2):381–389, April 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710710.

Taseli:1992:IBP

[Tas92] H. Taseli. The influence of the boundedness of polynomial potentials on the spectrum of the Schrödinger equation. Journal of Computational Physics, 101(2):252–255, August 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290002G.

Tau:1992:NSSa

[Tau92a] Eric Yu Tau. Numerical solution of the steady Stokes equations. Journal of Computational Physics, 98(2):348, February 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290158U.

Tau:1992:NSSb

[Tau92b] Eric Yu Tau. Numerical solution of the steady Stokes equations. Journal of Computational Physics, 99(2):190–195, April 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290200I.

Tau:1994:SOP

[Tau94] Eric Yu Tau. A second-order projection method for the incompressible Navier–Stokes equations in arbitrary domains. Journal of Computational Physics, 115(1):147–152, November 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711831.

Taylor:1991:AUF

[Tay91] Arthur C. Taylor III. Approved unpwind finite volume relaxation method for high speed viscous flows. Journal of Computational Physics, 94(1):253, May 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190154D.

Taylor:1994:HPS

[Tay94] Mark Taylor. A high performance spectral code for nonlinear MHD stability. Journal of Computational Physics, 110(2):407– 418, February 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184710357.

Tse:1998:FHP

[TC98] K. L. Tse and J. R. Chasnov. A Fourier–Hermite pseudospectral method for penetrative convection. Journal of Computational Physics, 142(2):489–505, May 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959460.

Tai:1997:ETM

[TCS97] C. H. Tai, D. C. Chiang, and Y. P. Su. Explicit time marching methods for the time-dependent Euler computations. Journal of Computational Physics, 130(2):191–202, January 15, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196955771.

Teixeira:1999:SSP

[Tei99] João Teixeira. Stable schemes for partial differential equations: The one-dimensional diffusion equation. Journal of Computational Physics, 153(2):403–417, August 10, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962836.

Tal-Ezer:1992:LOP

[TEKC92] Hillel Tal-Ezer, Ronnie Kosloff, and Charles Cerjan. Low-order polynomial approximation of propagators for the time-dependent Schrödinger equation. Journal of Computational Physics, 100(1): 179–187, May 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290318S.

Temam:1996:MMS

[Tem96] Roger Temam. Multilevel methods for the simulation of turbulence: a simple model. Journal of Computational Physics, 127(2):309–315, September 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999196901771.

Tessmer:1992:MDCb

[Tes92] Ekkert Tessmer. Multi domain Chebyshev–Fourier method for the solution of the equations of motion of dynamic elasticity. Journal of Computational Physics, 99(2):351, April 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290217M.

Taylor:1991:URM

[TfNW91] Arthur C. Taylor III, Wing fai Ng, and Robert W. Walters. Upwind relaxation methods for the Navier–Stokes equations using inner iterations. Journal of Computational Physics, 97(2):580, December 1991. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919190017F.

Tang:1996:RBC

[TG96] Y. Tang and R. Grimshaw. Radiation boundary conditions in barotropic coastal ocean numerical models. Journal of Computational Physics, 123(1):96–110, January 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919690008X.

Tang:1996:SOA

[TH96a] H. S. Tang and D. Huang. A second-order accurate capturing scheme for 1D inviscid flows of gas and water with vacuum zones. Journal of Computational Physics, 128(2):301–318, October 15, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196902120.

Thiart:1996:NSF

[TH96b] Jacob J. Thiart and Vladimir Hlavacek. Numerical solution of free-boundary problems: Calculation of interface evolution during CVD growth. Journal of Computational Physics, 125(1):262– 276, April 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196900923.

Thompson:1990:TDB

[Tho90] Kevin W. Thompson. Time-dependent boundary conditions for hyperbolic systems, II. Journal of Computational Physics, 89(2):439–461, August 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919090152Q.

Thompson:1997:INS

[Tho97] Joe Thompson. Introduction to “Numerical Solution of the Quasilinear Poisson Equation in a Nonuniform Triangle Mesh”.

Journal of Computational Physics, 135(2):126–127, August 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197957260.

Thuburn:1996:MFL

[Thu96] John Thuburn. Multidimensional flux-limited advection schemes. Journal of Computational Physics, 123(1):74–83, January 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196900066.

Tidriri:1995:DDC

[Tid95] Moulay D. Tidriri. Domain decomposition for compressible Navier–Stokes equations with different discretizations and formulations. Journal of Computational Physics, 119(2):271–282, July 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185711357.

Tidriri:1997:PTN

[Tid97] M. D. Tidriri. Preconditioning techniques for the Newton– Krylov solution of compressible flows. Journal of Computational Physics, 132(1):51–61, March 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196956053.

Tiwari:1998:CBE

[Tiw98] S. Tiwari. Coupling of the Boltzmann and Euler equations with automatic domain decomposition. Journal of Computational Physics, 144(2):710–726, August 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960119.

Toumi:1996:ALR

[TK96] I. Toumi and A. Kumbaro. An approximate linearized Riemann solver for a two-fluid model. Journal of Computational Physics, 124(2):286–300, March 15, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900601.

Trujillo:1999:PMV

[TK99] James Trujillo and George Em Karniadakis. A penalty method for the vorticity-velocity formulation. Journal of Computational Physics, 149(1):32–58, February 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961356.

Tessmer:1992:MDCa

[TKKB92] Ekkehart Tessmer, David Kessler, Dan Kosloff, and Alfred Behle. Multi-domain Chebyshev–Fourier method for the solution of the equations of motion of dynamic elasticity. Journal of Computational Physics, 100(2):355–363, June 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290241P.

Tamura:1997:RCM

[TKT97] Atsuhiro Tamura, Kazuo Kikuchi, and Tadayasu Takahashi. Residual cutting method for elliptic boundary value problems. Journal of Computational Physics, 137(2):247–264, November 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197958071.

Tolstykh:1998:PMT

[TL98] Andrei I. Tolstykh and Michael V. Lipavskii. On performance of methods with third- and fifth-order compact upwind differencing. Journal of Computational Physics, 140(2):205–232, March 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198958879.

Turkington:1993:MVP

[TLES93] Bruce Turkington, Alexander Lifschitz, Alexander Eydeland, and Joel Spruck. Multiconstrained variational problems in magnetohydrodynamics: Equilibrium and slow evolution. Journal of Computational Physics, 106(2):269–285, June 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711071.

Thomas:1995:FSS

[TLW95] T. G. Thomas, D. C. Leslie, and J. J. R. Williams. Free surface simulations using a conservative 3D code. Journal of Computational Physics, 116(1):52–68, January 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710054.

Trapp:1993:DPM

[TM93] John A. Trapp and Glen A. Mortensen. A discrete particle model for bubble-slug two-phase flows. Journal of Computational Physics, 107(2):367–377, August 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711514.

Tome:1994:GCM

[TM94] Murilo F. Tome and Sean McKee. GENSMAC: a computational marker and cell method for free surface flows in general domains. Journal of Computational Physics, 110(1):171–186, January 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999184710138.

Tang:1992:SMNa

[TMR92a] Tao Tang, S. McKee, and M. W. Reeks. A spectral method for the numerical solutions of a kinetic equation describing the dispersion of small particles in a turbulent flow. Journal of Computational Physics, 101(1):228, July 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192900646.

Tang:1992:SMNb

[TMR92b] Tao Tang, S. McKee, and M. W. Reeks. A spectral method for the numerical solutions of a kinetic equation describing the dispersion of small particles in a turbulent flow. Journal of Computational Physics, 103(2):222–230, December 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290397H.

Thangaraj:1998:RMD

[TN98] D. Thangaraj and A. Nathan. A rotated monotone difference scheme for the two-dimensional anisotropic drift-diffusion equation. Journal of Computational Physics, 145(1):445–461, September 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198960338.

Taylor:1992:IUFa

[TNW92a] Arthur C. Taylor III, Wing Fai Ng, and Robert W. Walters. An improved upwind finite volume relaxation method for high speed viscous flows. Journal of Computational Physics, 98(2): 348, February 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290156S.

Taylor:1992:IUFb

[TNW92b] Arthur C. Taylor III, Wing-Fai Ng, and Robert W. Walters. An improved upwind finite volume relaxation method for high speed viscous flows. Journal of Computational Physics, 99(1):159– 168, March 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999192902835.

Taylor:1992:URMa

[TNW92c] Arthur C. Taylor III, Wing-Fai Ng, and Robert W. Walters. Upwind relaxation methods for the Navier–Stokes equations using inner iterations. Journal of Computational Physics, 98(2):348, February 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290155R.

Taylor:1992:URMb

[TNW92d] Arthur C. Taylor III, Wing-Fai Ng, and Robert W. Walters. Upwind relaxation methods for the Navier–Stokes equations using inner iterations. Journal of Computational Physics, 99(1):68– 78, March 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999192902754.

Toth:1996:CSF

[TO96] Gábor Tóth and Dusan Odstrcil. Comparison of some flux corrected transport and total variation diminishing numerical schemes for hydrodynamic and magnetohydrodynamic problems. Journal of Computational Physics, 128(1):82–100, October 1, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196901977.

Tomboulides:1998:QTD

[TO98] Ananias G. Tomboulides and Steven A. Orzag. A quasi-two-dimensional benchmark problem for low Mach number compressible codes. Journal of Computational Physics, 146(2):691–706, November 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199919896079X.

Tolstykh:1994:AFO

[Tol94] Mikhail A. Tolstykh. Application of fifth-order compact upwind differencing to moisture transport equation in atmosphere. Journal of Computational Physics, 112(2):394–403, June 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711090.

Ton:1996:ISC

[Ton96] Vinh T. Ton. Improved shock-capturing methods for multicomponent and reacting flows. Journal of Computational Physics, 128(1):237–253, October 1, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999196902065.

Toth:1997:LPA

[Tót97] Gábor Tóth. The LASY preprocessor and its application to general multidimensional codes. Journal of Computational Physics, 138(2):981–990, December 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958137.

Toumi:1992:WFR

[Tou92] I. Toumi. A weak formulation of Roe’s approximate Riemann solver. Journal of Computational Physics, 102(2):360–373, October 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290378C.

Tourrette:1997:ABC

[Tou97] Loïc Tourrette. Artificial boundary conditions for the linearized compressible Navier–Stokes equations. Journal of Computational Physics, 137(1):1–37, October 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957855.

Tourrette:1998:ABC

[Tou98] Loïc Tourrette. Artificial boundary conditions for the linearized compressible Navier–Stokes equations: II. The discrete approach. Journal of Computational Physics, 144(1):151–179, July 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198959964.

Trangenstein:1992:NASa

[TP92a] John A. Trangenstein and Richard B. Pember. Numerical algorithms for strong discontinuities in elastic-plastic solids. Journal of Computational Physics, 100(2):435, June 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290264Y.

Trangenstein:1992:NASb

[TP92b] John A. Trangenstein and Richard B. Pember. Numerical algorithms for strong discontinuities in elastic-plastic solids. Journal of Computational Physics, 103(1):63–89, November 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290326T.

Tiselj:1997:MTP

[TP97] Iztok Tiselj and Stojan Petelin. Modelling of two-phase flow with second-order accurate scheme. Journal of Computational Physics, 136(2):503–521, September 15, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957788.

Tang:1990:VEP

[TPL90] C. T. Tang, V. C. Patel, and L. Landweber. Viscous effects on propagation and reflection of solitary waves in shallow channels.

Journal of Computational Physics, 88(1):86–113, May 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090243T.

Trepanier:1996:CSF

[TPRC96] J.-Y. Trépanier, M. Paraschivoiu, M. Reggio, and R. Camarero. A conservative shock fitting method on unstructured grids. Journal of Computational Physics, 126(2):421–433, July 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901473.

Tzivion:1999:NSK

[TRL99] Shalva Tzivion, Tamir G. Reisin, and Zev Levin. A numerical solution of the kinetic collection equation using high spectral grid resolution: a proposed reference. Journal of Computational Physics, 148(2):527–544, January 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961289.

Tahara:1996:LDA

[TS96] Y. Tahara and F. Stern. A large-domain approach for calculating ship boundary layers and wakes and wave fields for nonzero Froude number. Journal of Computational Physics, 127(2):398– 411, September 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196901837.

Todo:1998:PAL

[TS98] Y. Todo and T. Sato. A particle algorithm for linear kinetic analysis in Tokamak plasmas. Journal of Computational Physics, 141(1):37–45, March 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959058.

Tang:1999:SOG

[TS99] H. S. Tang and F. Sotiropoulos. A second-order Godunov method for wave problems in coupled solid-water-gas systems. Journal of Computational Physics, 151(2):790–815, May 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962204.

Tahara:1992:IAC

[TSR92] Y. Tahara, F. Stern, and B. Rosen. An interactive approach for calculating ship boundary layers and wakes for nonzero Froude number. Journal of Computational Physics, 98(1):33–53, January 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290171T.

Tourigny:1992:NSB

[TSS92] Y. Tourigny and J. M. Sanz-Serna. The numerical study of blowup with application to a nonlinear Schrödinger equation. Journal of Computational Physics, 102(2):407–416, October 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999192903829.

Tan:1995:NFE

[TSW95] Shaofen Tan, Nestor Sepulveda, and John P. Wikswo, Jr. A new finite-element approach to reconstruct a bounded and discontinuous two-dimensional current image from a magnetic field map. Journal of Computational Physics, 122(1):150–164, November 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185712041.

Tsynkov:1995:ANE

[Tsy95] S. V. Tsynkov. An application of nonlocal external conditions to viscous flow computations. Journal of Computational Physics, 116(2):212–225, February 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710224.

Takeda:1990:CME

[TT90] Tatsuoki Takeda and Shinji Tokuda. Computation of MHD equilibrium of tokamak plasma (review article). Journal of Computational Physics, 89(2):488, August 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090156U.

Takeda:1991:CME

[TT91] Tatsuoki Takeda and Shinji Tokuda. Computation of MHD equilibrium of tokamak plasma. Journal of Computational Physics, 93(1):1–107, March 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919190074U.

Taylor:1997:SEM

[TTI97] Mark Taylor, Joseph Tribbia, and Mohamed Iskandarani. The spectral element method for the shallow water equations on the sphere. Journal of Computational Physics, 130(1):92–108, January 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196955540.

Tan:1994:MBE

[TV94] Zhiqiang Tan and Philip L. Varghese. The Δ–ε method for the Boltzmann equation. Journal of Computational Physics, 110(2):327–340, February 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999184710308.

teVelde:1992:NIPa

[tVB92a] G. te Velde and E. J. Baerends. Numerical integration for polyatomic systems. Journal of Computational Physics, 98(1):179, January 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290190A.

teVelde:1992:NIPb

[tVB92b] G. te Velde and E. J. Baerends. Numerical integration for polyatomic systems. Journal of Computational Physics, 99(1):84– 98, March 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999192902776.

Tam:1993:DRP

[TW93] Christopher K. W. Tam and Jay C. Webb. Dispersion-relation-preserving finite difference schemes for computational acoustics. Journal of Computational Physics, 107(2):262–281, August 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999183711423.

Tam:1994:RBC

[TW94] Christopher K. W. Tam and Jay C. Webb. Radiation boundary condition and anisotropy correction for finite difference solutions of the Helmholtz equation. Journal of Computational Physics, 113(1):122–133, July 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711223.

Tan:1997:MDR

[TWV97] Zhiqiang Tan, Dennis Wilson, and Philip L. Varghese. The mass-damped Riemann problem and the aerodynamic surface force calculation for an accelerating body. Journal of Computational Physics, 131(1):48–53, February 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196954169.

Tai:1995:NHG

[TYS95] C. H. Tai, S. L. Yin, and C. Y. Soong. A novel hyperbolic grid generation procedure with inherent adaptive dissipation. Journal of Computational Physics, 116(1):173–179, January 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710157.

Tamamidis:1996:CPB

[TZA96] Panos Tamamidis, Guoqing Zhang, and Dennis N. Assanis. Comparison of pressure-based and artificial compressibility methods for solving 3D steady incompressible viscous flows. Journal of Computational Physics, 124(1):1–13, March 1, 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900418.

Tseng:1992:NMMb

[TZWH92a] A. A. Tseng, J. Zou, H. P. Wang, and S. R. H. Hoole. Numerical modeling of macro and micro behaviors of materials in processing: a review. Journal of Computational Physics, 102(1):1–17, September 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105800014.

Tseng:1992:NMMa

[TZWH92b] A. A. Tseng, J. Zou, H. P. Wang, and S. R. H. Hoole. Numerical modelling of macro and micro behaviors of materials in processing: a review. Journal of Computational Physics, 100(2): 433, June 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290253U.

Ubbink:1999:MCS

[UI99] O. Ubbink and R. I. Issa. A method for capturing sharp fluid interfaces on arbitrary meshes. Journal of Computational Physics, 153(1):26–50, July 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962769.

Udaykumar:1997:MDA

[UKSTST97] H. S. Udaykumar, Heng-Chuan Kan, Wei Shyy, and Roger Tran-Son-Tay. Multiphase dynamics in arbitrary geometries on fixed Cartesian grids. Journal of Computational Physics, 137(2):366–405, November 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999197958058.

Ueki:1998:KTN

[UL98] Taro Ueki and Edward W. Larsen. A kinetic theory for nonanalog Monte Carlo particle transport algorithms: Exponential transform with angular biasing in planar-geometry anisotropically scattering media. Journal of Computational Physics, 145 (1):406–431, September 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999198960399.

Udaykumar:1999:CSL

[UMS99] H. S. Udaykumar, R. Mittal, and Wei Shyy. Computation of solid-liquid phase fronts in the sharp interface limit on fixed grids. Journal of Computational Physics, 153(2):535–574, August 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999199962940.

Ushijima:1996:ALE

[Ush96] Satoru Ushijima. Arbitrary Lagrangian–Eulerian numerical prediction for local scour caused by turbulent flows. Journal of Computational Physics, 125(1):71–82, April 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900807.

Unverdi:1991:FTM

[UT91] Salih Ozen Unverdi and Gretar Tryggvason. A front tracking method for viscous, incompressible, multi-fluid flows. Journal of Computational Physics, 97(2):583–584, December 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190029K.

Unverdi:1992:FTMb

[UT92a] Salih Ozen Unverdi and Grétar Tryggvason. A front-tracking method for viscous, incompressible, multi-fluid flows. Journal of Computational Physics, 99(1):180, March 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999192902949.

Unverdi:1992:FTMa

[UT92b] Salih Ozen Unverdi and Grétar Tryggvason. A front-tracking method for viscous, incompressible, multi-fluid flows. Journal of Computational Physics, 100(1):25–37, May 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290307K.

Umlauf:1999:CTT

[UWH99] L. Umlauf, Y. Wang, and K. Hutter. Comparing two topographyfollowing primitive equation models for lake circulation. Journal of Computational Physics, 153(2):638–659, August 10, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962988.

Umar:1990:BSC

[UWSB90] A. S. Umar, J. Wu, M. R. Strayer, and C. Bottcher. Basis-spline collocation method for the lattice solution of boundary value problems. Journal of Computational Physics, 89(2):490, August 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919090162T.

Umar:1991:BSC

[UWSB91] A. S. Umar, J. Wu, M. R. Strayer, and C. Bottcher. Basisspline collocation method for the lattice solution of boundary value problems. Journal of Computational Physics, 93(2):426– 448, April 1991. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919190193O.

Vadillo:1995:SBN

[Vad95] F. Vadillo. Spurious behavior for a numerical scheme of nonlinear elliptic equations. Journal of Computational Physics, 121(1):94– 101, October 1, 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999185711813.

Vadillo:1997:SFP

[Vad97] F. Vadillo. On spurious fixed points of Runge–Kutta methods. Journal of Computational Physics, 132(1):78–90, March 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196956156.

Valanju:1990:NFD

[Val90] P. M. Valanju. NUT: a fast 3-dimensional neutral transport code. Journal of Computational Physics, 88(1):114–130, May 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919090244U.

Valleau:1991:DSNa

[Val91a] J. P. Valleau. Density-scaling: a new Monte Carlo technique in statistical mechanics. Journal of Computational Physics, 94 (1):251, May 1991. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919190149F.

Valleau:1991:DSNb

[Val91b] J. P. Valleau. Density-scaling: a new Monte Carlo technique in statistical mechanics. Journal of Computational Physics, 96(1):193–216, September 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919190271L.

vanPutten:1993:NIM

[van93]  Maurice H. P. M. vanPutten.  A numerical implemen-

tation of MHD in divergence form. Journal of Computational Physics, 105(2):339–353, April 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710806.

VanLeer:1999:IAR

[Van99] Bram Van Leer. An introduction to the article “Reminiscences about Difference Schemes” by S. K. Godunov. Journal of Computational Physics, 153(1):1–5, July 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962708. See [God99].

Verboncoeur:1992:SPC

[VAVB92] J. P. Verboncoeur, M. V. Alves, V. Vahedi, and C. K. Birdsall. Simultaneous potential and circuit solution for 1D bounded plasma particle simulation codes. Journal of Computational Physics, 102(2):424, October 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290386D.

Verboncoeur:1993:SPC

[VAVB93] J. P. Verboncoeur, M. V. Alves, V. Vahedi, and C. K. Birdsall. Simultaneous potential and circuit solution for 1D bounded plasma particle simulation codes. Journal of Computational Physics, 104(2):321–328, February 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918371034X.

Vu:1995:ANS

[VB95] H. X. Vu and J. U. Brackbill. Accurate numerical solution of charged particle motion in a magnetic field. Journal of Computational Physics, 116(2):384–387, February 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710376.

Vrahatis:1990:CII

[VBB90] Michael N. Vrahatis, Tassos Bountis, and Nurit Budinsky. A convergence-improving iterative method for computing periodic orbits near bifurcation points. Journal of Computational Physics, 88(1):1–14, May 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090239W.

Vu:1999:AFK

[VBD99] H. X. Vu, B. Bezzerides, and D. F. DuBois. ASPEN: a fully kinetic, reduced-description particle-in-cell model for simulating parametric instabilities. Journal of Computational Physics, 156(1):12–42, November 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963507.

vanBuuren:1997:ISI

[vBKG97] R. van Buuren, J. G. M. Kuerten, and B. J. Geurts. Instabilities of stationary inviscid compressible flow around an airfoil. Journal of Computational Physics, 138(2):520–539, December 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958307.

vanBeek:1995:ADG

[vBvNW95] P. van Beek, R. R. P. van Nooyen, and P. Wesseling. Accurate discretization of gradients on non-uniform curvilinear staggered grids. Journal of Computational Physics, 117(2):364–367, March 15, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185710728.

Viecelli:1990:FRP

[VC90] J. A. Viecelli and E. H. Canfield, Jr. Functional representation of power-law random fields and times series. Journal of Computational Physics, 90(1):269–270, September 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090216N.

Viecelli:1991:FRP

[VC91] J. A. Viecelli and E. H. Canfield, Jr. Functional representation of power-law random fields and time series. Journal of Computational Physics, 95(1):29–39, July 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190251F.

Vazquez-Cendon:1999:ITS

[VC99] María Elena Vázquez-Cendón. Improved treatment of source terms in upwind schemes for the shallow water equations in channels with irregular geometry. Journal of Computational Physics, 148(2):497–526, January 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961277.

Valorani:1990:FMS

[VD90] M. Valorani and M. Di Giacinto. Fast moving sub-subsonic shocks in closed-end tubes. Journal of Computational Physics, 88(2): 409–432, June 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999190901876.

Vahedi:1997:SPC

[VD97a] Vahid Vahedi and G. DiPeso. Simultaneous potential and circuit solution for two-dimensional bounded plasma simulation codes. Journal of Computational Physics, 131(1):149–163, February 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196955916.

vanDorsselaer:1997:CEO

[vD97b] Jos L. M. van Dorsselaer. Computing eigenvalues occurring in continuation methods with the Jacobi–Davidson QZ method.

Journal of Computational Physics, 138(2):714–733, December 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197958447.

VanDooren:1993:NEF

[VDJ93] R. Van Dooren, M. De Groote, and H. Janssen. Numerical evidence of Feigenbaum’s number δ in non-linear oscillations. Journal of Computational Physics, 105(1):173–177, March 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710636.

vandeVorst:1992:BESb

[vdVMK92a] G. A. L. van de Vorst, R. M. M. Mattheij, and H. K. Kuiken. A boundary element solution for two-dimensional viscous sintering. Journal of Computational Physics, 99(1):180–181, March 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290297C.

vandeVorst:1992:BESa

[vdVMK92b] G. A. L. van de Vorst, R. M. M. Mattheij, and H. K. Kuiken. A boundary element solution for two-dimensional viscous sintering. Journal of Computational Physics, 100(1):50–63, May 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290309M.

vanderVegt:1998:DGF

[vdVvdV98] J. J. W. van der Vegt and H. van der Ven. Discontinuous Galerkin finite element method with anisotropic local grid refinement for inviscid compressible flows. Journal of Computational Physics, 141(1):46–77, March 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959046.

Velarde:1993:FMS

[Vel93] P. M. Velarde. A FCT method for staggered mesh. Journal of Computational Physics, 108(1):27–37, September 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711605.

Venkatakrishnan:1995:CSS

[Ven95] V. Venkatakrishnan. Convergence to steady state solutions of the Euler equations on unstructured grids with limiters. Journal of Computational Physics, 118(1):120–130, April 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710844.

Verstappen:1991:SNA

[Ver91] R. Verstappen. A simple numerical algorithm for elastohydrodynamic lubrication, based on a dynamic variation principle. Journal of Computational Physics, 97(2):460–488, December 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190009A.

Verkley:1997:SMTa

[Ver97a] W. T. M. Verkley. A spectral model for two-dimensional incompressible fluid flow in a circular basin. Journal of Computational Physics, 136(1):100–114, September 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957478.

Verkley:1997:SMTb

[Ver97b] W. T. M. Verkley. A spectral model for two-dimensional incompressible fluid flow in a circular basin. Journal of Computational Physics, 136(1):115–131, September 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919795748X.

Vicanek:1992:TGAb

[VG92a] M. Vicanek and N. M. Ghoniem. Two-group approach to the kinetics of particle cluster aggregation. Journal of Computational Physics, 99(2):351, April 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290215K.

Vicanek:1992:TGAa

[VG92b] M. Vicanek and N. M. Ghoniem. Two-group approach to the kinetics of particle cluster aggregation. Journal of Computational Physics, 101(1):1–10, July 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290038Z.

vanGijzen:1998:FED

[vGVO98] M. B. van Gijzen, C. B. Vreugdenhil, and H. Oksuzoglu. The finite element discretization for stream-function problems on multiply connected domains. Journal of Computational Physics, 140(1):30–46, February 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198958867.

Viera:1994:BSD

[Vie94] Fernando Viera. Boundary surface dynamics: An algorithm for stratified geostrophic flows. Journal of Computational Physics, 111(2):336–346, April 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710679.

VanderHeyden:1998:CFV

[VK98] W. B. VanderHeyden and B. A. Kashiwa. Compatible fluxes for van leer advection. Journal of Computational Physics, 146(1):1– 28, October 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198960703.

vanLeer:1997:GMGa

[vL97a] Bram van Leer. Godunov’s method for gas-dynamics: Current applications and future developments. Journal of Computational Physics, 132(1):1–2, March 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956655.

vanLeer:1997:GMGb

[vL97b] Bram van Leer. Godunov’s method for gas-dynamics: Current applications and future developments: Volume 132, number 1 (1997), pages 1–2. Journal of Computational Physics, 134(1): 199, June 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999197956965.

vanLeer:1997:TUC

[vL97c] Bram van Leer. Towards the ultimate conservative difference scheme. Journal of Computational Physics, 135(2):229–248, August 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197957041.

Vasilyev:1998:GCC

[VLM98] Oleg V. Vasilyev, Thomas S. Lund, and Parviz Moin. A general class of commutative filters for LES in complex geometries. Journal of Computational Physics, 146(1):82–104, October 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198960600.

Vinokur:1990:GFV

[VM90] Marcel Vinokur and Jean-Louis Montagné. Generalized fluxvector splitting and Roe average for an equilibrium real gas. Journal of Computational Physics, 89(2):276–300, August 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090145Q.

Venkatakrishnan:1993:ISU

[VM93] V. Venkatakrishnan and Dimitri J. Mavriplis. Implicit solvers for unstructured meshes. Journal of Computational Physics, 105(1): 83–91, March 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999183710557.

VanBuskirk:1994:SAC

[VM94] R. D. VanBuskirk and P. S. Marcus. Spectrally accurate contour dynamics. Journal of Computational Physics, 115(2):302–318, December 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184711983.

vandeVorst:1995:BBS

[vM95] G. A. L. vandeVorst and R. M. M. Mattheu. A BEM– BDF scheme for curvature driven moving Stokes flows. Journal of Computational Physics, 120(1):1–14, August 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918571145X.

Varadarajan:1996:LIA

[VM96a] V. Varadarajan and G. H. Miley. Linear instability analysis for toroidal plasma flow equilibria. Journal of Computational Physics, 123(2):415–434, February 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900340.

Venkatakrishnan:1996:IMC

[VM96b] V. Venkatakrishnan and D. J. Mavriplis. Implicit method for the computation of unsteady flows on unstructured grids. Journal of Computational Physics, 127(2):380–397, September 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901825.

Vosbeek:1997:CDS

[VM97] P. W. C. Vosbeek and R. M. M. Mattheij. Contour dynamics with symplectic time integration. Journal of Computational Physics, 133(2):222–234, May 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956680.

VandeVorst:1991:BES

[VMK91] G. A. L. Van de Vorst, R. M. M. Mattheij, and H. K. Kuiken. A boundary element solution for two-dimensional viscous sintering. Journal of Computational Physics, 97(2):581, December 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190021C.

Vold:1992:CICa

[VNC92a] E. L. Vold, F. Najmabadi, and R. W. Conn. Computational implementation of a coupled plasma-neutral fluid model. Journal of Computational Physics, 101(1):229, July 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290070F.

Vold:1992:CICb

[VNC92b] E. L. Vold, F. Najmabadi, and R. W. Conn. Computational implementation of a coupled plasma-neutral fluid model. Journal of Computational Physics, 103(2):300–319, December 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290403L.

vanVeldhuizen:1994:NLS

[vNZ94] M. vanVeldhuizen, R. Nieuwenhuizen, and W. Zijl. A note on log scale Hankel transforms. Journal of Computational Physics, 110(1):196–199, January 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710163.

Verzicco:1996:FDS

[VO96] R. Verzicco and P. Orlandi. A finite-difference scheme for three-dimensional incompressible flows in cylindrical coordinates. Journal of Computational Physics, 123(2):402–414, February 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196900339.

Vasilyev:1996:DAM

[VP96] Oleg V. Vasilyev and Samuel Paolucci. A dynamically adaptive multilevel wavelet collocation method for solving partial differential equations in a finite domain. Journal of Computational Physics, 125(2):498–512, May 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901114.

Vasilyev:1997:FAW

[VP97] Oleg V. Vasilyev and Samuel Paolucci. A fast adaptive wavelet collocation algorithm for multidimensional PDEs. Journal of Computational Physics, 138(1):16–56, November 15, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958149.

Vasilyev:1995:MWC

[VPS95] Oleg V. Vasilyev, Samuel Paolucci, and Mihir Sen. A multilevel wavelet collocation method for solving partial differential equations in a finite domain. Journal of Computational Physics, 120(1):33–47, August 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711473.

Vrahatis:1995:EML

[Vra95] Michael N. Vrahatis. An efficient method for locating and computing periodic orbits of nonlinear mappings. Journal of Computational Physics, 119(1):105–119, June 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711199.

Vierendeels:1999:MSI

[VRD99] Jan Vierendeels, Kris Riemslagh, and Erik Dick. A multigrid semi-implicit line-method for viscous incompressible and low-Mach-number flows on high aspect ratio grids. Journal of Computational Physics, 154(2):310–341, September 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963155.

Vreugdenhil:1991:APFa

[Vre91a] C. B. Vreugdenhil. Accuracy of product-formula algorithms. Journal of Computational Physics, 93(1):252, March 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999191900872.

Vreugdenhil:1991:APFb

[Vre91b] C. B. Vreugdenhil. Accuracy of product-formula algorithms. Journal of Computational Physics, 97(2):337–351, December 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999191900034.

Vesey:1995:TDF

[VS95] Roger A. Vesey and Don Steiner. A two-dimensional finite element model of the edge plasma. Journal of Computational Physics, 116(2):300–313, February 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710297.

Vorozhtsov:1995:SNM

[VSG95] E. V. Vorozhtsov, B. Yu. Scobelev, and V. G. Ganzha. Symbolic-numerical method for the stability analysis of difference schemes on the basis of the catastrophe theory. Journal of Computational Physics, 116(1):26–38, January 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710030.

Vorozhtsov:1996:C

[VSG96] E. V. Vorozhtsov, B. Yu. Scobelev, and V. G. Ganzha. Corrigendum. Journal of Computational Physics, 123(1):231–232, January 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196900194.

Vuik:1996:PKS

[VSH96] Kees Vuik, Agur G. J. Sevink, and Gérard C. Herman. A preconditioned Krylov subspace method for the solution of least squares problems in inverse scattering. Journal of Computational Physics, 123(2):330–340, February 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900273.

Vuik:1999:EPC

[VSM99] C. Vuik, A. Segal, and J. A. Meijerink. An efficient preconditioned CG method for the solution of a class of layered problems with extreme contrasts in the coefficients. Journal of Computational Physics, 152(1):385–403, June 10, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962551.

Vu:1996:AFE

[Vu96] H. X. Vu. An adiabatic fluid electron particle-in-cell code for simulating ion-driven parametric instabilities. Journal of Computational Physics, 124(2):417–430, March 15, 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900698.

Vu:1998:MPT

[Vu98] H. X. Vu. A massively parallel three-dimensional hybrid code for simulating ion-driven parametric instabilities. Journal of Computational Physics, 144(2):257–279, August 10, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919895803X.

vanWachem:1997:SFB

[vWBS^{+}97] B. G. M. van Wachem, A. F. Bakker, J. C. Schouten, M. W. Heemels, and S. W. de Leeuw. Simulation of fluidized beds with lattice gas cellular automata. Journal of Computational Physics, 135(1):1–7, July 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957193.

Vesier:1991:CMS

[VY91] C. Cockerham Vesier and A. P. Yoganathan. A computer method for simulation of cardiovascular flow fields: Validation of approach. Journal of Computational Physics, 94(1):252, May 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190153C.

Vesier:1992:CMS

[VY92] C. Cockerham Vesier and A. P. Yoganathan. A computer method for simulation of cardiovascular flow fields: Validation of approach. Journal of Computational Physics, 99(2):271–287, April 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290207F.

Williams:1995:MWS

[WA95] John R. Williams and Kevin Amaratunga. A multiscale wavelet solver with O(n) complexity. Journal of Computational Physics, 122(1):30–38, November 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711941.

Wajsowicz:1993:CFA

[Waj93] Roxana C. Wajsowicz. A consistent formulation of the anisotropic stress tensor for use in models of the large-scale ocean circulation. Journal of Computational Physics, 105(2):333–338, April 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002199918371079X.

Walsteijn:1994:RNM

[Wal94] Fred H. Walsteijn. Robust numerical methods for 2D turbulence. Journal of Computational Physics, 114(1):129–145, September 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999184711545.

Wang:1995:FCI

[Wan95] Z. J. Wang. A fully conservative interface algorithm for overlapped grids. Journal of Computational Physics, 122(1):96–106, November 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999185711990.

Wang:1997:NEF

[Wan97] Song Wang. A novel exponentially fitted triangular finite element method for an advection-diffusion problem with boundary layers. Journal of Computational Physics, 134(2):253–260, July 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197956916.

Wang:1999:EIE

[Wan99] Z. Jane Wang. Efficient implementation of the exact numerical far field boundary condition for Poisson equation on an infinite domain. Journal of Computational Physics, 153(2):666–670, August 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999199962897.

Wathen:1992:OMGb

[Wat92a] A. J. Wathen. Optimal moving grids for time-dependent partial differential equations. Journal of Computational Physics, 99(2): 352, April 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290221J.

Wathen:1992:OMGa

[Wat92b] A. J. Wathen. Optimal moving grids for time-dependent partial differential equations. Journal of Computational Physics, 101(1): 51–54, July 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290041V.

Wu:1994:NSA

[WCO94] Han-Ming Wu, G. F. Carey, and M. E. Oakes. Numerical simulation of AC plasma arc thermodynamics. Journal of Computational Physics, 112(1):24–30, May 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184710795.

Wu:1999:TRL

[WCSW99] K. Wu, A. Canning, H. D. Simon, and L.-W Wang. Thick-restart Lanczos method for electronic structure calculations. Journal of Computational Physics, 154(1):156–173, September 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963064.

Werner:1997:RAS

[WD97] M. J. Werner and P. D. Drummond. Robust algorithms for solving stochastic partial differential equations. Journal of Computational Physics, 132(2):312–326, April 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196956387.

Warschkow:1998:DCI

[WDE98] Oliver Warschkow, John M. Dyke, and Donald E. Ellis. A divide-and-conquer implementation of the discrete variational DFT method for large molecular and solid systems. Journal of Computational Physics, 143(1):70–89, June 10, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959745.

Williamson:1992:STSa

[WDH^{+}92a] David L. Williamson, John B. Drake, James J. Hack, Rüdiger Jakob, and Paul N. Swarztrauber. A standard test set for numerical approximations to the shallow water equations in spherical geometry. Journal of Computational Physics, 101(1):227–228, July 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290060C.

Williamson:1992:STSb

[WDH^{+}92b] David L. Williamson, John B. Drake, James J. Hack, Rüdiger Jakob, and Paul N. Swarztrauber. A standard test set for numerical approximations to the shallow water equations in spherical geometry. Journal of Computational Physics, 102(1):211–224, September 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999105800166.

Weber:1998:CEC

[Web98] C. F. Weber. Convergence of the equilibrium code SOLGASMIX. Journal of Computational Physics, 145(2):655–670, September 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198960466.

Wegmann:1997:UDS

[Weg97] Rudolf Wegmann. An upwind difference scheme for the double-adiabatic equations. Journal of Computational Physics, 131(1):199–215, February 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999196956028.

Weiland:1992:SMR

[Wei92] C. Weiland. A split-matrix Runge–Kutta type space marching procedure. Journal of Computational Physics, 102(2):319– 335, October 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999192903759.

Welch:1995:LST

[Wel95] Samuel W. J. Welch. Local simulation of two-phase flows including interface tracking with mass transfer. Journal of Computational Physics, 121(1):142–154, October 1, 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711850.

Welton:1998:TDP

[Wel98]  Walter C. Welton.  Two-dimensional PDF/SPH simula-

tions of compressible turbulent flows. Journal of Computational Physics, 139(2):410–443, January 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958782.

Winterhalter:1999:AAD

[WEMH99] J. Winterhalter, D. G. Ebling, D. Maier, and J. Honerkamp. Analysis of admittance data: Comparison of a parametric and a nonparametric method. Journal of Computational Physics, 153(1):139–159, July 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962691.

Wang:1999:FEL

[WEQ^{+}99] Hong Wang, Richard E. Ewing, Guan Qin, Stephen L. Lyons, Mohamed Al-Lawatia, and Shushuang Man. A family of Eulerian–Lagrangian localized adjoint methods for multidimensional advection-reaction equations. Journal of Computational Physics, 152(1):120–163, June 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962393.

Werne:1995:INS

[Wer95] J. Werne. Incompressibility and no-slip boundaries in the Chebyshev–Tau approximation: Correction to Kleiser and Schumann’s influence-matrix solution. Journal of Computational Physics, 120(2):260–265, September 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711631.

Werne:1998:LEC

[Wer98] J. Werne. Letter to the Editor: Comment on “There Is No Error in the Kleiser–Schumann Influence Matrix Method”. Journal of Computational Physics, 141(1):88–89, March 20, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959277. See [KHW98].

Westermann:1992:LSBa

[Wes92a] Thomas Westermann. Localization schemes in 2D boundaryfitted grids. Journal of Computational Physics, 100(2):433, June 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290251S.

Westermann:1992:LSBb

[Wes92b] Thomas Westermann. Localization schemes in 2D boundaryfitted grids. Journal of Computational Physics, 101(2):307– 313, August 1992. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919290008M.

Westermann:1994:PCS

[Wes94] T. Westermann. Particle-in-cell simulations with moving boundaries-adaptive mesh generation. Journal of Computational Physics, 114(2):161–175, October 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711570.

Wang:1999:CMC

[WF99] Nien-Tzu Wang and Aaron L. Fogelson. Computational methods for continuum models of platelet aggregation. Journal of Computational Physics, 151(2):649–675, May 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962125.

Wu:1997:AOC

[WH97] Xudong T. Wu and Edward F. Hayes. Algorithms for obtaining cumulative reaction probabilities for chemical reactions. Journal of Computational Physics, 130(1):136–147, January 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196955758.

Wang:1998:SIS

[WH98] Yongqi Wang and Kolumban Hutter. A semi-implicit semispectral primitive equation model for lake circulation dynamics and its stability performance. Journal of Computational Physics, 139(1):209–241, January 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958502.

Whalen:1996:ALT

[Wha96] Paul P. Whalen. Algebraic limitations on two-dimensional hydrodynamics simulations. Journal of Computational Physics, 124(1):46–54, March 1, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999196900431.

Whitaker:1990:NSHa

[Whi90a] Nathaniel Whitaker. Numerical solution of the Hele–Shaw equations. Journal of Computational Physics, 87(2):494, April 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919090263Z.

Whitaker:1990:NSHb

[Whi90b] Nathaniel Whitaker. Numerical solution of the Hele–Shaw equations. Journal of Computational Physics, 90(1):176–199, September 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919090202C.

Whitaker:1994:SNM

[Whi94] Nathaniel Whitaker. Some numerical methods for the Hele– Shaw equations. Journal of Computational Physics, 111(1):81– 88, March 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184710461.

Winslow:1995:MGM

[Win95] Alan M. Winslow. Multifrequency-gray method for radiation diffusion with Compton scattering. Journal of Computational Physics, 117(2):262–273, March 15, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185710649.

Winslow:1997:NSQ

[Win97] Alan M. Winslow. Numerical solution of the quasilinear Poisson equation in a nonuniform triangle mesh. Journal of Computational Physics, 135(2):128–138, August 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956989.

Wirth:1997:ESM

[Wir97] A. Wirth. An extension of spectral methods to quasi-periodic and multiscale problems. Journal of Computational Physics, 132(2): 285–290, April 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196956284.

Witwit:1996:AIP

[Wit96a] M. R. M. Witwit. Application of the inner product technique to some nonpolynomial potentials for multidimensional quantum systems. Journal of Computational Physics, 129(1):220–232, November 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196902454.

Witwit:1996:ELN

[Wit96b] M. R. M. Witwit. Energy levels for nonsymmetric double-well potentials in several dimensions: Hill determinant approach. Journal of Computational Physics, 123(2):369–378, February 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196900303.

Ward:1993:CAS

[WJC93] D. J. Ward, S. C. Jardin, and C. Z. Cheng. Calculations of axisymmetric stability of Tokamak plasmas with active and passive feedback. Journal of Computational Physics, 104(1):221– 240, January 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999183710223.

Watson:1996:FEP

[WJTP96] Willie R. Watson, Michael G. Jones, Sharon E. Tanner, and Tony L. Parrott. A finite element propagation model for extracting normal incidence impedance in nonprogressive acoustic wave fields. Journal of Computational Physics, 125(1):177–186, April 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002199919690087X.

Worlikar:1996:NST

[WK96] Aniruddha S. Worlikar and Omar M. Knio. Numerical simulation of a thermoacoustic refrigerator: I. Unsteady adiabatic flow around the stack. Journal of Computational Physics, 127(2):424– 451, September 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196901850.

Warburton:1999:DGM

[WK99a] T. C. Warburton and G. E. Karniadakis. A discontinuous Galerkin method for the viscous MHD equations. Journal of Computational Physics, 152(2):608–641, July 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962484.

Wood:1999:DCF

[WK99b] William A. Wood and William L. Kleb. Diffusion characteristics of finite volume and fluctuation splitting schemes. Journal of Computational Physics, 153(2):353–377, August 10, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962812.

Wu:1997:CRE

[WKHS97] Xudong T. Wu, Prakashan P. Korambath, Edward F. Hayes, and Danny C. Sorensen. Computation of rovibrational eigenvalues of van der Waals molecules on a CRAY T3D. Journal of Computational Physics, 138(2):286–301, December 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958198.

Worlikar:1998:NST

[WKK98] Aniruddha S. Worlikar, Omar M. Knio, and Rupert Klein. Numerical simulation of a thermoacoustic refrigerator: II. Stratified flow around the stack. Journal of Computational Physics, 144(2):299–324, August 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958162.

Weese:1997:USS

[WKMH97] J. Weese, E. Korat, D. Maier, and J. Honerkamp. Unfolding sphere size distributions with a density estimator based on Tikhonov regularization. Journal of Computational Physics, 138(2):331–353, December 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958228.

Winckelmans:1993:CVP

[WL93] G. S. Winckelmans and A. Leonard. Contributions to vortex particle methods for the computation of three-dimensional incompressible unsteady flows. Journal of Computational Physics, 109(2):247–273, December 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999183712167.

Webster:1996:PBL

[WL96] Frank Webster and Gen-Ching Lo. Projective block Lanczos algorithm for dense, Hermitian eigensystems. Journal of Computational Physics, 124(1):146–161, March 1, 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196900509.

Wissink:1996:EIM

[WLC96] Andrew M. Wissink, Anastasios S. Lyrintzis, and Anthony T. Chronopoulos. Efficient iterative methods applied to the solution of transonic flows. Journal of Computational Physics, 123(2):379– 393, February 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196900315.

Wilcoxson:1994:IES

[WM94] Mark Wilcoxson and Vasilios Manousiouthakis. On an implicit ENO scheme. Journal of Computational Physics, 115(2):376–389, December 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184712034.

Wineberg:1991:ISMa

[WMG^{+}91a] Stephan B. Wineberg, Joseph F. McGrath, Edward F. Gabl, L. Ridgway Scott, and Charles E. Southwell. Implicit spectral methods for wave propagation problems. Journal of Computational Physics, 93(2):486, April 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999191902005.

Wineberg:1991:ISMb

[WMG^{+}91b] Stephen B. Wineberg, Joseph F. McGrath, Edward F. Gabl, L. Ridgway Scott, and Charles E. Southwell. Implicit spectral methods for wave propagation problems. Journal of Computational Physics, 97(2):311–336, December 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999191900023.

Wang:1993:GCA

[WNY93] Jian-Ping Wang, Yoshiaki Nakamura, and Michiru Yasuhara. Global coefficient adjustment method for Neumann condition in explicit Chebyshev collocation method and its application to compressible Navier–Stokes equations. Journal of Computational Physics, 107(1):160–175, July 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711332.

Wang:1996:VIN

[WONM96] W. X. Wang, M. Okamoto, N. Nakajima, and S. Murakami. Vector implementation of nonlinear Monte Carlo Coulomb collisions. Journal of Computational Physics, 128(1):209–222, October 1, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196902041.

Welton:1997:PMC

[WP97] Walter C. Welton and Stephen B. Pope. PDF model calculations of compressible turbulent flows using smoothed particle hydrodynamics. Journal of Computational Physics, 134(1):150–168, June 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197956801.

Wang:1991:HRS

[WR91] Z. Wang and B. E. Richards. High resolution schemes for steady flow computation. Journal of Computational Physics, 97(1):53– 72, November 1991. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199919190038M.

Wright:1998:NID

[Wri98] Joseph P. Wright. Numerical instability due to varying time steps in explicit wave propagation and mechanics calculations. Journal of Computational Physics, 140(2):421–431, March 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0021999198959022. See comments [Ske98].

Williams:1991:AAM

[WS91] H. T. Williams and Richard R. Silbar. Automated angular momentum recoupling algebra. Journal of Computational Physics, 97(2):584, December 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919190032G.

Williams:1992:AAMa

[WS92a] H. T. Williams and Richard R. Silbar. Automated angular momentum recoupling algebra. Journal of Computational Physics, 99(1):180, March 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999192902916.

Williams:1992:AAMb

[WS92b] H. T. Williams and Richard R. Silbar. Automated angular momentum recoupling algebra. Journal of Computational Physics, 99(2):299–309, April 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919290209H.

Wright:1993:PBC

[WS93] J. A. Wright and W. Shyy. A pressure-based composite grid method for the Navier–Stokes equations. Journal of Computational Physics, 107(2):225–238, August 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711393.

Wang:1996:APF

[WS96] Shun-Lien Wang and Robert F. Sekerka. Algorithms for phase field computation of the dendritic operating state at large supercoolings. Journal of Computational Physics, 127(1):110–117, August 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196901618.

Wesseling:1999:CFG

[WSK99] P. Wesseling, A. Segal, and C. G. M. Kassels. Computing flows on general three-dimensional nonsmooth staggered grids. Journal of Computational Physics, 149(2):333–362, March 1, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961563.

Watanabe:1996:IEI

[WTH^{+}96] T. H. Watanabe, Y. Todo, R. Horiuchi, K. Watanabe, and T. Sato. Implementation of an electrostatic implicit particle simulation scheme. Journal of Computational Physics, 127(2):473– 481, September 1996. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999196901886.

Willemsen:1998:CDP

[WVHS98] S. M. Willemsen, T. J. H. Vlugt, H. C. J. Hoefsloot, and B. Smit. Combining dissipative particle dynamics and Monte Carlo techniques. Journal of Computational Physics, 147(2):507–517, December 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198960995.

Wu:1995:EVV

[WWW95] X. H. Wu, J. Z. Wu, and J. M. Wu. Effective vorticityvelocity formulations for three-dimensional incompressible viscous flows. Journal of Computational Physics, 122(1):68–82, November 1995. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999185711977.

Xu:1998:PDG

[XH98] Kun Xu and Jishan Hu. Projection dynamics in Godunov-type schemes. Journal of Computational Physics, 142(2):412–427, May 20, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002199919895923X.

Xiao:1999:CMS

[Xia99] Feng Xiao. A computational model for suspended large rigid bodies in 3D unsteady viscous flows. Journal of Computational Physics, 155(2):348–379, November 1, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963404.

Xu:1995:GKF

[XMJ95] Kun Xu, Luigi Martinelli, and Antony Jameson. Gas-kinetic finite volume methods, flux-vector splitting, and artificial diffusion. Journal of Computational Physics, 120(1):48–65, August 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999185711485.

Xu:1997:FDS

[XMP97] H. Y. Xu, M. D. Matovic, and A. Pollard. Finite difference schemes for three-dimensional time-dependent convectiondiffusion equation using full global discretization. Journal of Computational Physics, 130(1):109–122, January 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196955643.

Xu:1994:NNS

[XP94] Kun Xu and Kevin H. Prendergast. Numerical Navier– Stokes solutions from gas kinetic theory. Journal of Computational Physics, 114(1):9–17, September 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711454.

Xu:1999:ANA

[XP99] J. Xu and S. B. Pope. Assessment of numerical accuracy of PDF/Monte Carlo methods for turbulent reacting flows. Journal of Computational Physics, 152(1):192–230, June 10, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199962411.

Xia:1990:NSC

[XPK90] Yongwei Xia, Georges Pananakakis, and Georges Kamarinos. Numerical simulation of the current-voltage characteristics of MIS tunnel devices. Journal of Computational Physics, 91(2):478– 485, December 1990. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999190900497.

Xu:1993:APV

[XS93] Yuenong Xu and Mitchell D. Smooke. Application of a primitive variable Newton’s method for the calculation of an axisymmetric laminar diffusion flame. Journal of Computational Physics, 104(1):99–109, January 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710120.

Xu:1997:BBS

[Xu97] Kun Xu. BGK-based scheme for multicomponent flow calculations. Journal of Computational Physics, 134(1):122–133, June 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999197956771.

Xu:1999:GKT

[Xu99] Kun Xu. Gas-kinetic theory-based flux splitting method for ideal magnetohydrodynamics. Journal of Computational Physics, 153(2):334–352, August 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999199962800.

Yakimiw:1996:ACW

[Yak96] E. Yakimiw. Accurate computation of weights in classical Gauss–Christoffel quadrature rules. Journal of Computational Physics, 129(2):406–430, December 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196902582.

Yang:1990:ACM

[Yan90] Huanan Yang. An artificial compression method for ENO schemes: The slope modification method. Journal of Computational Physics, 89(1):125–160, July 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://

www.sciencedirect.com/science/article/pii/002199919090120P.

Yavneh:1997:AFO

[Yav97] Irad Yavneh. Analysis of a fourth-order compact scheme for convection-diffusion. Journal of Computational Physics, 133(2):361–364, May 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199919795659X.

Yamaleev:1998:SMM

[YB98] Nail K. Yamaleev and Josef Ballmann. Space-marching method for calculating steady supersonic flows on a grid adapted to the solution. Journal of Computational Physics, 146(1):436–463, October 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198960739.

Yang:1997:KBS

[YCTC97] Jaw-Yen Yang, Min-Hung Chen, I-Nan Tsai, and Jer-Wei Chang. A kinetic beam scheme for relativistic gas dynamics. Journal of Computational Physics, 136(1):19–40, September 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957673.

Yepes:1993:CSD

[YDT93] G. Yepes and R. Domínguez-Tenreiro. A computational study of the dynamical evolution of self-gravitating systems. Journal of Computational Physics, 104(1):75–85, January 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710090.

Yee:1997:EIM

[Yee97] H. C. Yee. Explicit and implicit multidimensional compact highresolution shock-capturing methods: Formulation. Journal of Computational Physics, 131(1):216–232, February 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196956089.

Yao:1994:HAA

[YG94] Jian Hua Yao and R. J. Gooding. A high-accuracy algorithm for solving nonlinear PDEs with high-order spatial derivatives in 1 + 1 dimensions. Journal of Computational Physics, 112(2): 382–393, June 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184711089.

Yiu:1995:SVI

[YG95] K. F. C. Yiu and M. B. Giles. Simultaneous viscousinviscid coupling via transpiration. Journal of Computational Physics, 120(2):157–170, September 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711564.

Yang:1997:SSE

[YGH97] B. Yang, D. Gottlieb, and J. S. Hesthaven. Spectral simulations of electromagnetic wave scattering. Journal of Computational Physics, 134(2):216–230, July 1, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956862.

Guo:1992:CSFa

[yGmC92] Ben yu Guo and Wei ming Cao. A combined spectral-finite element method for solving two-dimensional unsteady Navier– Stokes equations. Journal of Computational Physics, 100(2):434, June 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290258Z.

Yang:1995:RFC

[YH95] J. Y. Yang and J. C. Huang. Rarefied flow computations using nonlinear model Boltzmann equations. Journal of Computational Physics, 120(2):323–339, September 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711680.

Yoshino:1990:MSP

[YHW90] Fumio Yoshino, Tatsuo Hayashi, and Ryoji Waka. A method of self-pursued boundary value on a body and the Magnus effect calculated with this method. Journal of Computational Physics, 89(2):488, August 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919090155T.

Yoshino:1991:MSP

[YHW91] Fumio Yoshino, Tatsuo Hayashi, and Ryoti Waka. A method of self-pursued boundary value on a body and the Magnus effect calculated with this method. Journal of Computational Physics, 93(1):224–249, March 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999191900805.

Yin:1996:NMS

[Yin96] Z.-M. Yin. New methods for simulation of fractional Brownian motion. Journal of Computational Physics, 127(1):66–72, August 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196901588.

Yee:1990:HRS

[YKM90] H. C. Yee, G. H. Klopfer, and J.-L. Montagné. High-resolution shock-capturing schemes for inviscid and viscous hypersonic flows. Journal of Computational Physics, 88(1):31–61, May 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919090241R.

Yang:1998:EAT

[YL98] Ping Yang and K. N. Liou. An efficient algorithm for truncating spatial domain in modeling light scattering by finite-difference technique. Journal of Computational Physics, 140(2):346–369, March 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999198958971.

Yasar:1991:EAG

[YM91] Osman Yasar and Gregory A. Moses. Explicit adaptivegrid radiation magnetohydrodynamics. Journal of Computational Physics, 97(2):581, December 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190022D.

Yasar:1992:EAGb

[YM92a] Osman Yasar and Gregory A. Moses. Explicit adaptive-grid radiation magnetohydrodynamics. Journal of Computational Physics, 99(1):180, March 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919290295A.

Yasar:1992:EAGa

[YM92b] Osman Yasar and Gregory A. Moses. Explicit adaptive-grid radiation-magnetohydrodynamics. Journal of Computational Physics, 100(1):38–49, May 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199919290308L.

Yavneh:1995:RMS

[YM95] Irad Yavneh and James C. McWilliams. Robust multigrid solution of the shallow-water balance equations. Journal of Computational Physics, 119(1):1–25, June 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711126.

Younes:1999:NFM

[YMAC99] Anis Younes, Robert Mose, Philippe Ackerer, and Guy Chavent. A new formulation of the mixed finite element method for solving elliptic and parabolic PDE with triangular elements. Journal of Computational Physics, 149(1):148–167, February 10, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961502.

Young:1991:LRR

[YMB^{+}91] David P. Young, Robin G. Melvin, Michael B. Bieterman, Forrester T. Johnson, Satish S. Samant, and John E. Bussoletti. A locally refined rectangular grid finite element method: Application to computational fluid dynamics and computational physics. Journal of Computational Physics, 92(1):1–66, January 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190291R.

Ye:1999:ACG

[YMUS99] T. Ye, R. Mittal, H. S. Udaykumar, and W. Shyy. An accurate Cartesian grid method for viscous incompressible flows with complex immersed boundaries. Journal of Computational Physics, 156(2):209–240, December 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999199963568.

Yorkey:1990:EIT

[Yor90] Thomas J. Yorkey. Electrical impedance tomography with piecewise polynomial conductivities. Journal of Computational Physics, 91(2):344–360, December 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919090041X.

Yang:1992:CSAb

[YP92a] H. Q. Yang and A. J. Przekwas. A comparative study of advanced shock-capturing shcemes applied to Burgers’ equation. Journal of Computational Physics, 102(1):139–159, September 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999105800129.

Yang:1992:CSAa

[YP92b] H. Q. Yang and A. J. Przekwas. A comparative study of advanced shocl-capturing schemes applied to Burger’s equation. Journal of Computational Physics, 100(2):434, June 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290256X.

Yang:1998:PWA

[YP98] Baolin Yang and Peter G. Petropoulos. Plane-wave analysis and comparison of split-field, biaxial, and uniaxial PML methods as ABCs for pseudospectral electromagnetic wave simulations in curvilinear coordinates. Journal of Computational Physics, 146(2):747–774, November 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919896082X.

Yousfi:1994:FEM

[YPH94] M. Yousfi, A. Poinsignon, and A. Hamani. Finite element method for conservation equations in electrical gas discharge areas. Journal of Computational Physics, 113(2):268–278, August 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199918471134X.

Yeung:1993:CFE

[YR93] Y. Y. Yeung and C. Rudowicz. Crystal field energy levels and state vectors for the 3d^{N} ions at orthorhombic or higher symmetry sites. Journal of Computational Physics, 109(1):150–152, November 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999183712064.

Yarvin:1998:GOD

[YR98] Norman Yarvin and Vladimir Rokhlin. A generalized onedimensional fast multipole method with application to filtering of spherical harmonics. Journal of Computational Physics, 147(2):594–609, December 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999198961046.

Yaremchuk:1998:SAS

[YS98a] A. I. Yaremchuk and J. Schröter. Spectral analysis of symmetric operators: Application to the Laplace tidal model. Journal of Computational Physics, 147(1):1–21, November 20, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960211.

Yu:1998:ELG

[YS98b] Hua-Gen Yu and Sean C. Smith. The elimination of Lanczos ghosting effects by MINRES filter diagonalization. Journal of Computational Physics, 143(2):484–494, July 1, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919895971X.

Yee:1999:LDH

[YSD99] H. C. Yee, N. D. Sandham, and M. J. Djomehri. Low-dissipative high-order shock-capturing methods using characteristic-based filters. Journal of Computational Physics, 150(1):199–238, March 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198961770.

Yee:1991:DAS

[YSG91] H. C. Yee, P. K. Sweby, and D. F. Griffiths. Dynamical approach study of spurious steady-state numerical solutions of nonlinear differential equations. I. The dynamics of time discretization and its implications for algorithm development in computational fluid dynamics. Journal of Computational Physics, 97(2):249–310, December 1991. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999191900012.

Yavneh:1997:MSR

[YSMG97] Irad Yavneh, Alexander F. Shchepetkin, James C. McWilliams, and Lee Paul Graves. Multigrid solution of rotating, stably stratified flows. Journal of Computational Physics, 136(2):245–262, September 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999197957752.

Yu:1992:TDC

[YTS92] Sheng-Tao Yu, Y.-L Peter Tsai, and Jian-Shun Shuen. Threedimensional calculations of supersonic reacting flows using an LU scheme. Journal of Computational Physics, 101(2):276–286, August 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290005J.

Ying:1996:SRT

[YWS96] Jinpin Ying, Feng Wu, and Wei Sun. Simultaneous reconstruction of two parameters for transport equation in a stratified halfspace. Journal of Computational Physics, 125(2):434–439, May 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196901059.

Yang:1998:IWE

[YYCH98] Jaw-Yen Yang, Shih-Chang Yang, Yih-Nan Chen, and Chiang-An Hsu. Implicit weighted ENO schemes for the three-dimensional incompressible Navier–Stokes equations. Journal of Computational Physics, 146(1):464–487, October 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960624.

Yang:1998:AMI

[YZ98] George Z. Yang and Nicholas Zabaras. The adjoint method for an inverse design problem in the directional solidification of binary alloys. Journal of Computational Physics, 140(2):432–452, March 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198958934.

Zabusky:1997:INS

[Zab97] Norman J. Zabusky. Introduction to “Numerical Simulation of Hydrodynamics by the Method of Point Vortices”. Journal of Computational Physics, 135(2):187–188, August 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957521.

Zalesak:1997:IFC

[Zal97] Steven T. Zalesak. Introduction to “Flux-Corrected Transport. I. SHASTA, A Fluid Transport Algorithm That Works”. Journal of Computational Physics, 135(2):170–171, August 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957569.

Zaninetti:1991:ATE

[Zan91] L. Zaninetti. About the time evolving Voronoi tessellation. Journal of Computational Physics, 97(2):559–565, December 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919190014C.

Zanino:1992:HMPa

[Zan92a] R. Zanino. Hydrodynamic modeling of particle and angular momentum transport in rotating tokamak plasmas with impurities. Journal of Computational Physics, 98(1):179–180, January 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999192901933.

Zanino:1992:HMPb

[Zan92b] R. Zanino. Hydrodynamic modeling of particle and angular momentum transport in rotating tokamak plasmas with impurities. Journal of Computational Physics, 98(2):301–316, February 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290144N.

Zanino:1997:AFE

[Zan97] R. Zanino. Advanced finite element modeling of the Tokamak plasma edge. Journal of Computational Physics, 138(2):881–906, December 1997. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999197958484.

Zaninetti:1999:ARW

[Zan99] L. Zaninetti. About the random walk from many injection points. Journal of Computational Physics, 156(2):382–392, December 10, 1999. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002199919996357X.

Zegeling:1992:EGWa

[ZB92a] P. A. Zegeling and J. G. Blom. An evaluation of the gradientweighted moving-finite-element method in one space dimension. Journal of Computational Physics, 101(2):452, August 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290024S.

Zegeling:1992:EGWb

[ZB92b] P. A. Zegeling and J. G. Blom. An evaluation of the gradientweighted moving-finite-element method in one space dimension. Journal of Computational Physics, 103(2):422–441, December 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290413S.

Zha:1996:NST

[ZB96] G.-C. Zha and E. Bilgen. Numerical study of three-dimensional flows using unfactored upwind-relaxation sweeping algorithm. Journal of Computational Physics, 125(2):425–433, May 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0021999196901047.

Zachary:1991:HOG

[ZC91] Andrew L. Zachary and Phillip Colella. A higher-order Godunov method for the equations of ideal magnetohydrodynamics. Journal of Computational Physics, 97(2):585, December 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919190033H.

Zachary:1992:HOG

[ZC92] Andrew L. Zachary and Phillip Colellaz. A higher-order Godunov method for the equations of ideal magnetohydrodynamics. Journal of Computational Physics, 99(2):341–347, April 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199919290213I.

Zerroukat:1994:FDA

[ZC94] Mohamed Zerroukat and Chris R. Chatwin. A finite-difference algorithm for multiple moving boundary problems using real and virtual grid networks. Journal of Computational Physics, 112(2): 298–307, June 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184711028.

Zhu:1996:EBI

[ZCH96] Jingyi Zhu, Xinfu Chen, and Thomas Y. Hou. An efficient boundary integral method for the Mullins–Sekerka problem. Journal of Computational Physics, 127(2):246–267, September 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901734.

Zhao:1996:VLS

[ZCMO96] Hong-Kai Zhao, T. Chan, B. Merriman, and S. Osher. A variational level set approach to multiphase motion. Journal of Computational Physics, 127(1):179–195, August 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901679.

Zhang:1999:SSD

[ZDD99] Juhua Zhang, Zhuping Duan, and Jing Ding. Simulating shock to detonation transition: Algorithm and results. Journal of Computational Physics, 150(1):128–142, March 20, 1999. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198961691.

Zhuang:1995:SQS

[ZDG95] X. Zhuang, A. K. Didwania, and J. D. Goddard. Simulation of the quasi-static mechanics and scalar transport properties of ideal granular assemblages. Journal of Computational Physics, 121(2):331–346, October 15, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999195901876.

Zhang:1994:MSS

[Zha94] Lin-Bo Zhang. A multigrid solver for the steady incompressible Navier–Stokes equations on curvilinear coordinate systems. Journal of Computational Physics, 113(1):26–34, July 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711156.

Zhang:1998:FHA

[Zha98] Jun Zhang. Fast and high accuracy multigrid solution of the three dimensional Poisson equation. Journal of Computational Physics, 143(2):449–461, July 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198959824.

Zhong:1996:CMP

[ZHL96] Xiaoguang Zhong, Thomas Y. Hou, and Philippe G. LeFloch. Computational methods for propagating phase boundaries. Journal of Computational Physics, 124(1):192–216, March 1, 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999196900534.

Zhong:1996:ASI

[Zho96] Xiaolin Zhong. Additive semi-implicit Runge–Kutta methods for computing high-speed nonequilibrium reactive flows. Journal of Computational Physics, 128(1):19–31, October 1, 1996. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919690193X.

Zhong:1998:HOF

[Zho98] Xiaolin Zhong. High-order finite-difference schemes for numerical simulation of hypersonic boundary-layer transition. Journal of Computational Physics, 144(2):662–709, August 10, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960107.

Zabusky:1997:CDE

[ZHR97] Norman J. Zabusky, M. H. Hughes, and K. V. Roberts. Contour dynamics for the Euler equations in two dimensions. Journal of Computational Physics, 135(2):220–226, August 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002199919795703X.

Zhu:1995:SOP

[Zhu95] Jingyi Zhu. The second-order projection method for the backward-facing step flow. Journal of Computational Physics, 117(2):318–331, March 15, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999185710698.

Zinchenko:1994:ARC

[Zin94a] Alexander Z. Zinchenko. Algorithm for random close packing of spheres with periodic boundary conditions. Journal of Computational Physics, 114(2):298–307, October 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999184711685.

Zinchenko:1994:EAC

[Zin94b] Alexander Z. Zinchenko. An efficient algorithm for calculating multiparticle thermal interaction in a concentrated dispersion of spheres. Journal of Computational Physics, 111(1):120– 135, March 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999184710497.

Zingg:1993:FDS

[ZL93]  David W. Zingg and Harvard Lomax.  Finite-difference

schemes on regular triangular grids. Journal of Computational Physics, 108(2):306–313, October 1993. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183711848.

Zaitsev:1998:DST

[ZLOT98] F. S. Zaitsev, V. V. Longinov, M. R. O’Brien, and R. Tanner. Difference schemes for the time evolution of three-dimensional kinetic equations. Journal of Computational Physics, 147 (2):239–264, December 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999198960752.

Zhu:1997:AFA

[ZLP97] Xiaojun Zhu, Guangtsai Lei, and Guangwen Pan. On application of fast and adaptive periodic Battle–Lemarie wavelets to modeling of multiple lossy transmission lines. Journal of Computational Physics, 132(2):299–311, April 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196956375.

Zhao:1998:CBB

[ZMOW98] Hong-Kai Zhao, Barry Merriman, Stanley Osher, and Lihe Wang. Capturing the behavior of bubbles and drops using the variational level set approach. Journal of Computational Physics, 143(2):495–518, July 1, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958101.

Zrahia:1997:HSE

[ZOI97] U. Zrahia, S. A. Orszag, and M. Israeli. Hybrid spectral element/asymptotic method for boundary layers problems. Journal of Computational Physics, 138(2):858–880, December 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197956631.

Zhou:1995:ASM

[ZP95] Hua Zhou and C. Pozrikidis. Adaptive singularity method for Stokes flow past particles. Journal of Computational Physics, 117(1):79–89, March 1, 1995. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999185710467.

Zwart:1999:IST

[ZRR99] P. J. Zwart, G. D. Raithby, and M. J. Raw. The integrated space– time finite volume method and its application to moving boundary problems. Journal of Computational Physics, 154(2):497–519, September 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999199963246.

Zhu:1992:PMCa

[ZS92a] Jingyi Zhu and James Sethian. Projection methods coupled to level set interface techniques. Journal of Computational Physics, 100(2):435, June 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999192902673.

Zhu:1992:PMCb

[ZS92b] Jingyi Zhu and James Sethian. Projection methods coupled to level set interface techniques. Journal of Computational Physics, 102(1):128–138, September 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999105800117.

Zhou:1997:NMW

[ZS97] Jianwei Zhou and Nader Saffari. Numerical modelling of wave propagation in elastic rectangular block media. Journal of Computational Physics, 131(2):299–309, March 1, 1997. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196955965.

Zang:1994:NSG

[ZSK94] Yan Zang, Robert L. Street, and Jeffrey R. Koseff. A nonstaggered grid, fractional step method for time-dependent incompressible Navier–Stokes equations in curvilinear coordinates. Journal of Computational Physics, 114(1):18–33, September 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999184711466.

Zutic:1997:NIP

[ZV97] Igor Zutić and Oriol T. Valls. Numerically implemented perturbation method for the nonlinear magnetic moment of an anisotropic superconductor. Journal of Computational Physics, 136(2):337–353, September 15, 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197957399.

Zerroukat:1997:BEM

[ZW97] M. Zerroukat and L. C. Wrobel. A boundary element method for multiple moving boundary problems. Journal of Computational Physics, 138(2):501–519, December 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999197958290.

Zhou:1998:SNS

[ZWS98] Liqun Zhou, Vincent B. Wickwar, and Robert W. Schunk. Solving the Navier–Stokes systems with weak viscosity and strong heat conduction using the flux-corrected transport technique and the alternating-directional explicit method. Journal of Computational Physics, 144(2):379–401, August 10, 1998. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999198960168.

Zhao:1994:NIC

[ZY94] Shengkai Zhao and Matthew J. Yedlin. A new iterative Chebyshev spectral method for solving the elliptic equation ∇·(σ∇u) ≈ f. Journal of Computational Physics, 113(2):215–223, August 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999184711314.

Zhang:1998:TDC

[ZYKC98] Y. L. Zhang, K. S. Yeo, B. C. Khoo, and W. K. Chong. Three-dimensional computation of bubbles near a free surface. Journal of Computational Physics, 146(1):105–123, October 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999198960429.

Zhang:1998:IMI

[ZZ98] Guanquan Zhang and Yu Zhang. An iterative method for the inversion of the two-dimensional wave equation with a potential. Journal of Computational Physics, 147(2):485–506, December 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002199919899996X.